A Complete Bibliography of Publications in Computer Physics Communications: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

11 April 2024 Version 1.21

Title word cross-reference

(a, a′) [SR84a, SR84a]. (a, b) [SS84b, SMD84, SR84a]. (a, bγ) [SS84b, SMD84, SR84a]. (a, bγ − γ) [SS84b, SMD84, SR84a]. (a, γ) [SR84a]. (a, γ − γ) [SR84a]. (η, π) [BDG^{+}84]. (n, γ) [Har81, Har84a]. (np)^{q} [CLS^{+}84]. + [LH88, SAC87, vdM84a]. 0 [Col80a, Col84a, Coo84b]. 1 [ELS88, ERB88, GS89, KT81, Ods84]. 1.6 [NAS87b]. 1/2 [KF84]. 10 [Col80a, Col84a]. $11.50 [Eas80]. 15 [And84b, And84a]. 19 [And84b, And84a]. 1p [MNJL84]. 2 [Atz86, BPW84a, BPW84b, Cro88, ERB88, KT81, MCJ81, MCJ84, MN84b, RBMC86]. 2.5 [HL88]. 27 [And84b, And84a]. 3 [And84b, And84a, Car89, Cat89, CN87, JN85, Kir87, Moh89, SS81, Syk81, Via87, Wan87, YpMzCeCg85, vdV89]. 3/2 [Coo84b]. $36.80 [Eas84a]. 3n [Sha84]. 4 [BLR82, BLR84, YpMzCeCg85]. $45 [Taj86]. 5/2 [Cat89]. 6.3 [SB87]. 7 [And84b, And84a]. 8 [BS86b]. 9 [AS83, AS84, KA87, RRC89, SA83, SA84]. > [BKJ84]. ^{+} [BS84d, GS89, LH88]. _{2}^{+} [WMCH84]. ^{+}e^{−} [SB87]. ^{10} [VBD88]. ^{2}Σ [Nai84, Nai86]. ^{3} [MZZ89]. ^{k} [Lew84b]. _{1} [MZZ89]. _{2} [DST80, DST84a, DST84b, JB80, Joh88, LNB^{+}88, LH88, Mag84b, Rob80b, Rob84c, RS84a, RRG88, WMCH84]. _{2}^{+} [RS83, RS84d]. _{3} [LH88]. A [BF84, Kac84]. A(γ_{v}, D)R [TW83, TW84]. α [KKP82, KPH84, KKP84, Klo84a]. B [Cha84c]. b_{n} [BS84d]. β [LCH^{+}81, Sch81, Sch83]. · [LCS84]. D [Hel80, All84c, Chi84b, DM87]. δW [GTR^{+}81, TBG80]. e [SB87]. e^{+} [TS89]. e^{+−} → q [BKJ84]. e^{+}e^{−} [BKJ84, FGK89, Rit84b, Sjö83, Sjö84a, Sjö86]. e^{+}e^{−}μ^{+}μ^{−} [BDK86a]. e^{+}e^{−} → e^{+}e^{−}μ^{+}μ^{−} [BDK86c]. e^{+}e^{−} → μ^{+}μ^{−} [BKJ83]. e^{+}e^{−} → qq [BKJ83]. e^{+}e^{−} → τ^{+}τ^{−} [JW85]. e^{−} [TS89]. η [BFSG84]. f [AM84a]. F_{0}(x) [SC84]. F_{m}(z) [EMR84, Jak84]. g [CDN^{+}87]. H_{n}^{(1)}(z) [AM84b]. H_{n}^{(2)}(z) [AM84b]. Hx = eSx [Nas84]. I [GPS84, LJLC84, Hub84a]. I = 2 [Cat89]. I = 3/2 [DG84]. I nu(z) [Tho04a, Cam81, Cam84a, Cam84b, TB87]. IL [KPP84a]. J [CH88, Moo81a, Moo84a, Hub84a, Hub84b, RRC89, SG84a, Sha84]. J_{0} [AM84c]. J_{1} [AM84c]. j_{n} [AM84e]. J_{n}(x) [Col80a, Col84a]. J_{n}(z) [AM84b]. J_{ν}(x) [Cam84c]. jj [Dya86, Gra84a, PGB84]. k [Sim80, Sim84]. K_{ν}(z) [Cam81, Cam84a, Cam84b, TB87, Tho04a]. L [Gas84]. l(r) [Del84b]. λ [Bar81b, Bar84c]. λϕ^{4} [Bai87]. LDU [Oya86]. LU [KPP84a]. m [MPIM86]. m = 0 [Köl81b, Köl84b]. m = 1 [Köl81b, Köl84b]. μ [MS87b]. μ^{+}μ^{−} [BKJ84]. N [DW89, De 88, Duc80, KK84a, Mur88, VČ88, CN84, Man87a, Mor84e]. ν [Cam84d]. O(3) [MPS84]. O(α^{3}) [JW85]. P [Ben82, All84c, FBS88, FBS89]. P_{−1/2+iτ}^{m} (x) [Köl81b, Köl84b]. P_{L}^{M}(x) [Del84a]. p_{T} [BI85]. π [Sch86c]. p ⊥ [IW87, Ing87]. pT [Ben84b]. QR [BM85b]. R [BBM82, BBM84, BBC^{+}84, Cla84, DK87b, Lew84b, Mor84c, SB82, ST82a, ST84c, Sho81, Sho84, TBB87]. r^{−n} [Rud84]. ρ [BFSG84]. s = 1 [MGTD84]. S_{n} [SM84c]. SU(2) [AM81, AM84d, CM85, KLS^{+}89, Sha84, CGGK85]. SU(3) [AD84, Bra84d, Bra84e, BS86b, CN84, CDL88, Hoe86a, KN86, MHP84, VBD88, BMR84, KM85]. SU(N) [AMC83, AMC84, CM83, CM84e, DMM83, DMM84b, MT86]. SU3 [MNJL84]. τ^{±} → X^{±} [JW85]. → [SAC87, vdM84a, Le 89]. U [BRB89]. U(1) [CMS83, CMS84, DMM84a, EMM81]. U(3) [DLPL89, MPS84]. U(5) ⊃ O(5) ⊃ O(3) [YP88, YP89]. U(N) [AMC83, AMC84, DLPL89, DMM83, DMM84b]. U(n) × SU(m) [KS84a]. U_{A}1 [Via87]. V [Ano88t]. W^{±} [LO86, Odo84d]. y′′ = g(x)y + r(x) [Col80b]. Y_{2}^{0} [Hir84]. Y_{4}^{0} [Hir84]. y_{n} [AM84e]. Y_{n}(z) [AM84b]. Z [Lew84c]. Z^{0} [LO86, Odo84d]. Z_{0} [BKJ83, BKJ84].

-annihilation [Rit84b]. -body [Man87a, YpMzCeCg85, DW89, Mur88]. -centroids [DK87b]. -coupling [Gra84a, PGB84]. -d [BLR82, BLR84, ELS88, Atz86, Car89, CN87, Cro88, ERB88, Moh89, Ods84]. -dimensional [DM87, Kir87, De 88, VČ88]. -expansion [Lew84c]. -factors [CDN^{+}87]. -fermion [Duc80]. -harmonics [Gas84]. -He [Rob84c]. -Matrix [ST84c, Sho84, BBM82, BBM84, BBC^{+}84, Cla84, Mor84c, SB82, ST82a, Sho81, TBB87]. -particle [Mor84e]. -plane [BDG^{+}84]. -point [AS83, AS84, KA87, SA83, SA84]. -shell [MNJL84]. -space [Sim80, Sim84]. -spin [Hub84a]. -spline [Cha84c]. -square [Moo81a, Moo84a]. -systems [Sch86c]. -transform [BRB89]. -version [FBS88, FBS89]. -zeros [Cam84d].

0-19-853614-3 [Gri87]. 0-387-19466-5 [Ano89h]. 0-7458-0019-X [Jes87].

1 [Art85a, Dic82, DR82, FR81, Fin82, Fin84a, KJ85, PCM85, SG82, SB82]. 1/2 [Hei84, Lan84a, SDH84]. 10 [KC81a, KC81b, MKRJ89]. 10-P [MKRJ89]. 100 [Ano89x]. 100/200 [MMM85]. 1000 [KKLS82]. 108 [Sta87b]. 11 [DGK^{+}81]. 11-banded [CKA89]. 110.00 [Man86]. 12.95/$20.00 [Duc86]. 14.95 [Jes87]. 16/32 [BBE^{+}81]. 164 [MB83]. 169 [Sch86a]. 1983 [Ano82a]. 1986 [Gri87]. 1987-1997 [New87]. 1A [CRPL81, CRPL84]. 1s [THH82]. 1st [Wai87].

2 [BO80, Eas89, RWZ87, Rob85]. 2.0 [PCL89]. 200 [MMM85, VH87]. 205 [BM82b, BM82c, BCM83a, BM84a, BM86, BSDM88, BA86, Car89, CMO86, Dic82, HVS87, KW84a, Pur82, Tol85, VSH83, VBD88]. 205-ETA [VBD88]. 21-25 [Ano82-30]. 22.50 [Eas82]. 225.00 [Bra86]. 230-75 [MTN^{+}82]. 27.20 [And86]. 2nd [Bra86, Wai87]. 2O [Laa86].

3 [Ano89h, Eas84a, Eas87a, Eas89]. 3-540-11147-6 [Eas84a]. 3-540-13401-8 [Sta87b]. 3-540-16966-0 [Fin87]. 3-540-18909-2 [Eas89]. 3-540-19466-5 [Ano89h]. 3-D [Eas87a]. 3081/E [Sph87]. 3090/VF [Suc87]. 31 [KS08]. 32-bit [VGD^{+}89]. 34A [DCC84]. 36 [Tho04b]. 370/165 [HRR84]. 3d [And83b, And83a]. 3rd [Ano89a].

41 [Ran87a]. 47 [Tho04a]. 488 [KKLS82].

5 [AGH^{+}88, Ano89h, HR86a, HR86b]. 50 [RR84].

6 [Car80, Eas84a]. 64 [Ost85]. 68000 [BvEF85].

7 [OCC80]. 77 [Bru86, Nad86a]. 780 [VGD^{+}89]. 780/E [VGD^{+}89]. 7th [Ano88-27].

8088/8086 [CG87]. 88 [CC89a, CC89b]. 8x [Met87, RW85].

9.95 [Nad86a]. 90 [Art85a, Rob85]. 90-277-1591-2 [Rob85]. 90-277-1597-1 [Art85a].

= [CH88, MPIM86].

A-three [Aue84]. AAKF [FS84c]. AAKL [Sax84c]. AAKP [Sax84b, Sax84a]. ab-initio [OCC84, OCC80]. Abacus [LU83, LU84].

ABEL [BD82, FC84a, BD84b, FS84a]. Abelian [KR88]. ablation [NF83, NF84]. Abscissae [Tak88a, Tak88b]. Absorption [GS83, All84b, AHVV86, CHMM84, DPH82, DPH84, GH83, GH84a, JBWW86, Le 89, MGK86, Mor84a, VSP86]. abuse [BBD^{+}89]. accelerated [Bas87c, BKK86]. accelerating [BBR89a]. acceleration [IP88]. Accelerator [Gou88, And88, Ano81f, ABD82, ABD84, BT84a, Kon84, Ods80]. access [Ghi87]. accessible [BvSW84]. accompanying [TW89a]. according [AI86, Wil84b]. account [HK83, HK84b]. accretion [Rob87b]. Accuracy [Hay89a, Bar81a, Bar81b, Bar84c, BD82, BD84b, Eas87b, GB87, Jam88, Lis84, PS88, Sea84b, TB87, TV86]. Accurate [AM84b, Poo83, Di 86, Di 87]. acetylene [CH88, FH88a]. Acoustic [Nor82, Nor84b]. ACP [BAA^{+}87, GAA^{+}87, Sli89]. ACQI [Klo82, Klo84a]. ACQN [HBJ84, NA84]. acquisition [CMH87, CDF^{+}89, CSMHT88, CH84c, CHJZ89, Gai89, LBK88, PF86, SEF^{+}89, SBB^{+}89, SdAD^{+}88, Tha89, Whi89b]. ACRL [BTS84a]. acrz [God84]. ACTIV [Zlo82, Zlo84a]. Ada [Zal88, Pyl85, WHD86]. Adam [Duc86, Eas82, Lis88]. Adam-Hilger [Eas82]. ADAMO [KPST89, QTZ^{+}87]. Adams [Moo84e]. Adams-Gaunt [Moo84e]. adapt [Jac85b, Jac85a]. Adaptation [CT84b, DM86a, GH84b, HBJ84, Hib84b, Hib84a, LT87, MRW84, Sax84b, Sax84a, SCF86, Deu84c, God84, Hub84f, Klo82, Klo84a, MW84a, MW84b, MW84c, MT84b, Wil84a]. Adaptations [Ten84b, HA83, HA84a]. Adapted [FS84c, JT84b, BTS84a, IB87, JT83, NA84, PZCC89, Sax84c, Ská81, Ská84]. Adaptive [Fri88, MN80, WMN85, Lew82, Lew84a]. Adaptive-grid [WMN85]. addendum [Blo82b]. addressable [KC81b, KC84a, Köb89]. adiabatic [BMV83, BMV84, Hen84, Joh88, NW84, ONP89, WW84b]. adjoint [AMC83, AMC84]. adjuster [Han84f]. adsorbates [SP86]. adsorbed [SP87]. advanced [Vie88]. advances [Van81b]. advantages [HW84a]. adverse [LM84]. afterglows [LNS84]. Aftran [CRV^{+}89]. aggregation [SM86]. aid [DRD88, FC84b]. aided [JB80, Vie88]. air [GHO^{+}89, Goo84]. Airy [Moo81b, Moo84b]. Akima [DU89]. Alacant [Abr88]. ALAM [Mor80, Mor84d]. albedo [Moh89]. ALEPH [CMH87, CGMS89, DFOP89, Kel87, QTZ^{+}87]. Alfven [WBG84, Ter88]. Algebra [Har89, DVD87, FOW87, IKRT89, Eas89]. Algebraic [Stu88, YP89, AT84, BO80, FE88, KRSV82, KP81b, Pro82b, Pro82a, Pro84, SC89, III88, YP88]. algebraically [CW80, CW84a]. algebras [CT89, Ran86, Ran87a, RWZ87, Ran87b]. Algol [Jab84, KPPR84, SMD84]. Algorithm [CM84a, Moo80, Moo84c, AB86, BTWN87, BM82b, BM82c, BM84a, Bar81a, BB84c, BVL84, BLR82, BLR84, BMK89, BH89, Can81, CSC89, CGGK85, CAP86b, Cli82, CM83, CMM84, CM84e, Das83, DM87, FS85a, FS85b, HDK84, HP83, HP84, Hel82, JSB82, KN86, Klu84b, KRSV82, KM85, Mag81, MF81, MMS87, MHP84, MKRJ89, NFH^{+}88, Nex89, Par89b, SSM85, SCC^{+}88, SM85f, SM86, TB85, Tho04b, TBG80, Ver82, Wan87, WMH82, Wie86b].

Algorithms [Cha84a, GLC89, LBY86, Nes84, Sal87, WH81, BES82, BMBW85, CO84, Cla82a, Dru83, Eas87b, FR82, Hau87, Hoc82, Kra86a, MMBW88, MMM85, MN84a, McC89, Reg81, Sam85, Sco89, Tou89, Ves89, WH85, Wil88, WSV88, Eas82]. aligned [RSL^{+}84]. alkali [Moo84h]. alkali-like [Moo84h]. all-purpose [vM84a]. alliant [Par89b]. Allison [Moh80, Moh84]. allocation [Dob85, Mat87]. Allowed [Hub84a]. allowing [KS86b, KK84b]. alloys [BP84c, Isl81, Isl84a]. along [CCR88]. Alpha [PIM82, PIM84a, SDH84]. Alpha-decay [PIM82, PIM84a]. alternating [TDS81, TDS84]. ambiguous [Bas87b]. AMDS [HS84]. Amending [Cre84c, Cre80b]. amorphous [WW84b]. amount [BLMU81, TTT87]. amplification [RS84a]. amplitude [BWBL88, BB84b, Bas87a, FBW82]. Amplitudes [BS83, BS84a, Moo84h, AMK81, AMK84, Bar83, Bar84e, GPS84, KKS87, Pia84, PS84b, PP85, Tak85a]. AMPTE [BBH^{+}88]. Amsterdam [Bra86, Tru88, Her82b]. analogue [BGMM81]. analyser [CHMM84]. analysing [CSC^{+}86, KEMP81, KE84b, KEMP84, KE84a, Puf83, Puf84, RA82, RA84, VDM84b, War84]. Analysis [ASM80, ASM84, BB84a, BCM84a, Cot87, Eas87a, Gri87, HEL84, IA84, Nur85, RW86, Sal84b, AI88, AK84, AWG84, BLMU81, BSK88, BES82, BPA84, Ber87c, BR85, BBC^{+}87, BE89, BCVZ89, Bur87, CKT89, CT84a, CH84c, DG89, De 84a, Dun87, EHH88, FP89, Gal84, GV87, Gir81, Gre88, Hla86, Hoe89, IA80, Irv82, JR84, KN86, Ken89b, KLGT85, KK85, KB86a, KN84a, KAR81, KAR84b, Lis84, MGD81, Mur88, NM84, PS84c, RM89, Sch82a, Sea84a, Sko89, Szé85, TAK^{+}85b, Ver82, Via87, Vie88, WW84a, WW84c, WV80, WV84, Zlo84b, vM84d, vM84e, vM84b]. Analytic [BOT86, Bou87, CN84, BCM83b, BCM84b, CC84, Goo84, MP81b, MP84b, Sea82, Sea84c, Tro84, Lab84c]. Analytical [Ger80, Nes88, YB86, MN84b]. analyze [Eks84, Ver83, Ver84a, WC85]. analyzing [FMS87]. and/or [TS89]. angle [BDK86c, BKD83, BKD84, Deu84a, Deu84b, Deu84c, Deu84d, WML^{+}82]. Angular [SAC87, SB82, Tam84, AB83, BRTB84, BRT84, BGMP84, BC83a, BC84a, Bra87a, CW80, CDW82, CW84a, CDW84, Eks84, Gra84b, Gra84c, Hib84b, Hib84a, Hib84d, Hib84e, Hub84a, KEMP81, KEMP84, Lab84b, Lan84b, Rao81, RV84, RRC89, RSL^{+}84, SS84b, SDH84, TTW84a, TTW84b, Tak85a, TC87, VVM84, Wie85]. anharmonic [ST84b]. animals [RL80]. anisotropic [Cap86a, LB84, Pet84]. anisotropy [Nor82, Nor84b]. Anna [AK84]. annealing [SNH^{+}88]. annihilation [Câm82, Câm84g, DLB83, DLB84, FGK89, Odo81, Odo84c, Rit84b]. annihilations [ANP^{+}85]. Announcement [Ano80a, Ano82b, Ano82c, Ano83a, Ano84a, Ano84b, Ano85a, Ano87a, Ano88a, Ano88b, Ano80s]. Answer [Pag87]. antenna [FM84]. anti [LBY86, YB86]. anti-stokes [LBY86, YB86]. anticommuting [HD89]. antigenic [Fra83a, Fra84b]. any [Jud87, Ten84b]. AP120B [ASS82]. APE [Ano87b, Ano89x]. APE-100 [Ano89x]. Application [AWB82, AB88, BO80, CL84, CH87, FC84a, Gad80, GR81, LBY86, She84, WML^{+}82, ASS82, Ano87c, BM85a, BS86b, Cha82, Cha84c, Dob89, Kat84, Nav89, Oya86, PPHS86, SFB87, SBFI87, Sko89, Tou89, Zal88]. Applications [BM86, BCMR87, Ben88, Buz85, Jac81, KB86a, Kos80, NO89, SG82, Ber87a, BBC^{+}89, CDN^{+}87, Dag82, Gro89b, KK85, LCS84, PD89, PFGP82, Poh87a, Rea89, Sch88b, Sch88c, SM85a, Sne85, Whi89a, WMCH84, Zlá80, Art85a]. applied [AHVV86, RRG88, Ruf85, TSD84, WC85]. approach [AO88, BCF^{+}89, But89, CH84c, EP86, Fie80, GB87, HLM^{+}87, Jam88, KS84a, KS87a, KK83a, ML81, Phi86, SMT88, Vaj82]. approaches [PT82]. approx [Eas84a]. approxiamtion [CM81a]. approximants [GMR84a, GMR84b, Sta84]. Approximate [WL84, BKvLSE80, BKvLSE84]. Approximation [Jab84, Lab84d, Bin87, Bok84, BW87, CR84a, CC89b, CM84c, DO84, FAK82, GPS84, Hen84, Jak84, JA80, JA84, LZ84, MC84a, MCA87, NHK82, NK84, PPHS86, Pia84, PS84b, TCR84, WSV88, ZL84, Zlo89]. approximations [DNR83, DNR84, GWL89, Lab84c, ONP89, Sea84b]. APU [MTN^{+}82]. Aqueous [Laa86]. arbitrarily [Car84a]. Arbitrary [Sha84, Cap86a, DSK83, DSK84b, Deu84c, Deu84d, GT83, GT84, Hub84d, IKRT89, KS87a, Rys87, SS84b, VPB82, VPB84, WwYpCg88]. Architectural [Sch82a, JMS^{+}89, Nic89]. Architecture [Eas82, BMBW85, HBF^{+}85, Hey88, Jes85, Sli89]. architectures [Blo82a, Blo82b, BC88b, Gai89, Hal88, LK89, McC89, Sco89, TA89]. archival [TVD87]. archive [RHM^{+}89]. arcs [FCW84, NF83, NF84, SF84a, ZNF87]. Area [Ger88, BCC^{+}89, CM88, Flu89, ST84d]. argument [AM84b, AM84c, AM84e, BD80a, BD84a, Cam81, Cam84c, Cam84a, Cam84b, Di 86, Jak84, Köl84c, KB86b, TB87, Tho04a]. arguments [Bar82, Bar84b, Mas83, Moo81b, Moo84b, TB85, Tho04b]. ARGUS [TTT84]. ARGUS-V4 [TTT84]. Ariadne [CC89a]. Ariadne-88 [CC89a]. arising [AS83, And83b, And83a, AS84, And84b, And84a, AKS88a, AKS88b, Jes84, KA87, SA83, SA84, ST82b, ST84f]. arithmetic [Rys87]. AROVMI [AF84a, AF84b]. Array [Jes87, Par83, Ano87b, Bai87, CNRS82, Gro87, Gro89a, HAG^{+}89, JSB82, Jes85, KSG82, Män89, Nas84, Rap85, RW85, Sch82a, SM85a, Vis84a, CRV^{+}89]. arrays [ACC^{+}86, GH87, HRW85, Hey88, Hey89, Vis84b]. Art [Gri87]. articles [Van89]. Artificial [BE89, BD84d]. Asilomar [Nas87a, Ano87d, Wic87]. aspects [Sch86b]. assemblers [Tar86]. assignments [SDH84]. assistance [BR85]. assisted [AWG84]. Associated [Bra84a, ABG^{+}81a, AGF87, Del84a, YP88, YP89]. Association [SCF86, LH88]. associations [Fra83b, Fra84a]. associative [Par89a]. assurance [DP89, Sch88b, Sch88c]. astronomical [BS84c, MD82, RHM^{+}89]. astronomy [Jas84]. astrophysical [Ben88, Dor86, Har81, Har84a]. astrophysics [Jam82, Rob87b]. ASYM [TCS85]. asymmetric [And83b, And84b, AKS88a, AKS88b, JT83, JT84a, JT84b, KA87, Sea84a, SA83, SA84, WMN85]. asymmetrical [DM89a]. asymmetry [DG84, PI84, TCS85]. Asymptotic [KS84d, Nor84a, Chi84a, Cre80a, Cre84a, CSBB82, CSBB84, NN84]. asynchronous [BES82]. Asypck [Cre84a, Cre80a, Cre81, Cre84b]. ASYPCK2 [Cre81, Cre84b]. asyrot [Jud87]. ASYTOP [Sea84a]. Athene [CRPL81, CRL84, CRPL84]. atmospheres [Hub88]. Atom [NM84, AT84, AV84, Bar83, Bar84e, BGS81, BGS84c, Bil84, CSW84, Nov87, ONP89, Rin84, Sal82, Sal84e, ST84e, VČ88, Yat84]. atom-diatom [Bil84, Nov87, ONP89]. ATOMDIAT [Ten83, Ten84a, Ten84b]. Atomdiat2 [Ten84b]. Atomic [BT82, CT82, MR84b, AGF87, Arn84b, BS83, BS84a, BBC^{+}84, BBL^{+}84, Bla84, CCR84a, CLS^{+}84, DSK84a, DHN84, DH84a, DWVH89, DGJ^{+}89, FKM^{+}87, GPS84, GMN^{+}80, GMN^{+}84, Gra84b, Gra84c, HRS81, HRS84, Hib84b, Hib84a, Hib84d, Hib84e, HS84, HSB^{+}84, HKS83, HKS84, Kat84, KKLP84, Kla84, Lab84a, Lab84b, Lew84c, LHP87, MRC85, MEB88, Moo84h, Nob80, Nob84, Pia84, PS84b, PP85, Poo83, RFSG84a, SS89, ST82a, ST84c, STN89, Tal89, TW89a, WB84]. atoms [Aur84a, Aur84b, BBH^{+}84, BZ84, CCR84b, EA83, EA84, Kar80, Kar84a, Lew84d, LCW84, LZ84, Mal84, MBW89, Nus84, Sil84a, Sil84b, WS80, Wil84c, ZL84]. Attached [Bur85b, Dit89, Hoe89, PP82]. Auger [FV82, Rid82, Rid84b]. Auger-transitions [Rid82, Rid84b]. August [Bra86]. auroral [CCR88]. Austria [Kil80]. Author [Ano80b, Ano80c, Ano81a, Ano81b, Ano81c, Ano81d, Ano82d, Ano82e, Ano82f, Ano83b, Ano83c, Ano83d, Ano84c, Ano84d, Ano84e, Ano84f, Ano85b, Ano85c, Ano85d, Ano85e, Ano86a, Ano86b, Ano86c, Ano86d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano88c, Ano88d, Ano88e, Ano88f, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g]. authors [Ano80g, Ano82i, Ano86f]. automata [Den88, Suc87]. automated [Car84a]. Automatic [BLMU81, BF89, KKS87, LT87, Pie82, Pie84a, Pis84a, CM84d, Moh80, Moh84, NP84, Rap83, Zlo82, Zlo84a, vM84d, Gro89b]. automation [KKLS82]. Autonomous [Sch86d]. auxiliary [EMR84]. available [And86, SCF86]. AVANTI [HL88]. averaged [Ver89]. Averaging [Cha80]. axial [PI84, dBM82, dBM84]. axially [Cam84f, CDN^{+}87, DEW84, ZC84]. Axisymmetric [KJ84, KW84b, Rob87b, Shu87, BHM81, Cap86a, CO86, DM86b, Goe84, GPS^{+}81, HP83, HP84, KS84g, KS84h, MGD81, PT86, TAK^{+}85b, TW89b, TBG80]. axisymmetrical [Ugn80].

B [Eas84b, Taj86, Cha82]. B-spline [Cha82]. Back [Vio84]. Bäcklund [FK86a]. backward [CL84, MT84b]. Bag [CM84f, CM81b, CH84d, HGW86, Sal83b, Sal84f]. Balanced [PD89]. BALDUR [Red88, SPM^{+}88]. ballooning [TT89]. balloonings [GR81]. bamjet [Rit84a]. Band [Gru84a, MV81, MV83, MV84a, MV84b, De 84b, Gal81, Hof84b, HM84c, HM84d, HM84e, Sal83a, Sal84d]. banded [CKA89]. bands [BR85, HRB84, MW84a, MW84b, MW84c, WC84, Wil84a]. bank [Tau84a]. barbara [SS84b, SMD84]. Bardeen [MVV83, MVV84]. barrier [MR80, MR84a]. barriers [Jam86b, Sch82a]. baryon [HR84a]. base [Ber84a, Eld84, Gau84b, Gau84c, Gli84, Put89, Rob84f, Ros84a, Ros84b, SN84, SG89, Bel84b]. based [AO88, BD80b, BD81, BD84c, BKK86, BCF^{+}89, Cam86, FV81, Hla80, Hla84, Hof84a, HLM^{+}87, Joh89, KKLS82, KH89, KBMV86, Kon84, KRS83, Mat84, Mik86, RHM^{+}89, Ska82, Stu85, VM89, WBG84, WSV88]. bases [Fia84, Gau81, Har84b, Kat84, McI84, MPS84, Mee84]. Basic [Dew82, CW84b]. basis [AT84, DU89, Hay89a, Kar80, Kar84a, KS84a, Kol81a, Kol84a, MG84, MV81, MV84a, Poo83, Ran87b, ST84e, Tau84a, Ten85]. bcc [DM89b]. BCH [CGGK85]. be [BM82a, De 84c]. Beam [Mae88, ACR88, ABD82, ABD84, DLB83, DLB84, IL84, MT81, MMF81, OM88, Rit86, Ryn82, Ryn84]. beam-beam [Rit86]. beam-foil [IL84]. Beam-plasma [Mae88]. Belfast [HSB^{+}84]. below [AGH^{+}88, HR86a, HR86b]. Benchmark [CRH^{+}89]. Beniaminy [FV81]. Berlin [Eas84a, Eas89, Fin87, Man86, Sch86a, Sta87b]. Berlin-Heidelberg- [Eas84a]. Bernoulli [Di 87]. BESM [Car80]. BESM-6 [Car80]. Bessel [Tho04a, AM84b, AM84c, AM84e, Bar81a, Bar82, Bar84b, Cam81, Cam84c, Cam84a, Cam84b, Can81, Col80a, Col84a, Del84b, KB86b, Mas83, Nes84, Pie82, Pie84a, RC87, SZ84, Tal83, Tal84c, TB87]. beta [GTG81, GTG84, GTM85, GTMA88, GKM^{+}81]. beta-gamma [GTMA88]. betatron [AB88]. Bethe [BW87]. Betrt [BW87]. Better [DM86a]. between [BGMP84, DK87b, DEW84, DF84, Dya86, FT84a, Gia84, Jam84, Mal84, NAS87b, Ras80b, Ras84b, Sch87b, Ten85, Ter88]. Beyond [Gam89]. Bhabha [KKT^{+}88, TS89]. biconjugate [CKA89, KA87]. bicubic [CM84b]. biharmonic [Hou87]. Bilinear [Ito88]. Bin [ASM84, ASM80]. binary [BP84c, Gol84a, Isl81, Isl84a, Joh85, MEB88, OM83, OM84, PD89]. binding [Cug84]. biological [CNRS82]. Biophysical [Ber87a]. Birdsall [Taj86]. Birkhoff [ST84d]. bit [BBE^{+}81, VGD^{+}89]. bivariate [DU89]. black [PNO89]. Bloch [MN80]. Block [FM85, KP89, KWPP89, MP81a, MP84a, Nex89, Sun88]. blocked [Gru80, Gru84b]. BNDPKG [WC84]. BNL [Sau85]. Board [Ano81m, Ano81n, Ano82p, Ano82q, Ano82r, Ano82s, Ano83i, Ano83j, Ano84m, Ano84n, Ano84o, Ano84p, Ano85l, Ano85m, Ano85n, Ano86p, Ano86q, Ano86r, Ano86s, Ano86t, Ano87q, Ano87r, Ano87s, Ano87t, Ano88q, Ano88r, Ano88s, Ano88t, Ano89o, Ano89p, Ano89q, Ano89r, Ano89s, Ano80k, Ano80l, Ano81l]. Bochum [PF86]. bodies [DM89a]. body [AMK81, AMK84, Ass84, Bas87a, Bra84c, Bra84b, CM85, CgYpXh85, DW89, EEV89, Gir81, HIM84, Her88a, Man87a, MBW89, Mue84, Mur88, Rap85, Sil84a, Sil84b, WwYpCg88, WS80, Wil84c, YpMzCeCg85, YFL87]. body-centered [CM85]. Boeing [LP89]. Boltz [TSD84]. Boltzmann [Das89, RG80, RG84, WW80]. Boltzmann-equation [WW80]. bombardment [DM86a, MRW84, MM84a]. bond [FH88a, NW84, WW84b]. bond-mode [FH88a]. Book [And86, Ano89h, Art85a, Art85b, Bra86, Duc86, Eas80, Eas82, Eas84b, Eas84a, Eas89, Fin87, Gri87, Jes87, Lis88, Man86, Nad86a, Rob85, Sch86a, Sta87b, Taj86]. Boreas [Ods84]. born [DO84, NK84, TCR84]. bosons [LO86]. Boston [Duc86]. both [Nai86].

Bound [DHD84, Le 89, LR84a, Bat84, DW88, DeV84, FGSH85, Fie80, HDK84, HKT85, Kar80, Kar84a, KP85, LW84b, LT87, SS84c, SS84d, Smi84c]. bound-bound [Kar80, Kar84a]. bound-state [Fie80]. boundaries [BKD83, BKD84]. boundary [AP80, Bun86, CR84a, Cam84f, CM81b, CM84f, DM89a, HvRM86, Hoe86b, Hof88, MP81a, MP84a, O’C84a, VM89]. boundary-layer [AP80]. Bounding [BT81]. box [PNO89]. brackets [CgYpXh85, Dob84, FT84b, LJ84, SG84c, WwYpCg88]. brain [Vir89]. branch [HDK84]. branch-and-bound [HDK84]. Branches [EBS88]. Braunschweig [And86]. Breit [BGMP84, MGN80, MGN84, ST82a, ST84c, TN84]. Bremsstrahlung [BRTB84, BRT84, Ber80, DLB83, DLB84, GH83, GH84a, MC85, Odo84a, Odo84d]. Bringing [Boc89]. Bristol [Duc86, Eas82]. broad [Gal81]. broadened [KM86]. Brody [LJ84]. Brownian [SM86]. Brussels [Van81a]. bubble [BSP84, BD84d, BdCP81]. bubbles [Bor89]. bulk [Mor84a]. bunch [ABD82, ABD84]. bus [HGJ^{+}89].

C [Ano89h, Eas80, Eas82, Taj86, Bai86]. C.-Y [Eas80]. Cabannes [Eas84b]. caesium [ÖW84]. Calcomp [CA84]. calculate [All84a, BCFK85, Bar83, Bar84e, BGMP84, BBC^{+}84, BBL^{+}84, BTS84a, Bru85, Bru86, Bur84, Car89, CgYpXh85, De 84c, DLB83, DLB84, DHN84, DS83, DS84a, DH84a, Duf84, DO84, Dya86, EM84a, EM84b, FS84b, GTG81, GTG84, God84, Gra84b, Gra84c, HBJ84, HRS81, HRS84, Hib84b, Hib84a, Hib84c, Hib84e, JT84a, JA80, JA84, Jam84, KF84, Klo82, Klo84a, Klo84b, Lad86, LO86, LHP87, MLL86, Mar89c, MGN80, MGN84, MT84b, NA84, Odo82, Odo84b, PV84, Ran86, Ran87a, RBHW84, Sal82, Sal84e, San82, San84, SM84a, ST82a, ST84c, Ste84c, TSD84, YpMzCeCg85, YFL87]. calculated [Hen84]. calculating [BKvLSE80, BKvLSE84, BRTB84, BRT84, Bat84, Bor82, Bor84, BTS84b, BT84b, CDW82, CDW84, Cre80a, Cre84a, DK87a, DK87b, DSK83, DSK84b, FT84a, FE88, Gia84, Gru84d, HP83, HP84, Hib84d, Hof84b, Hof84a, HRB84, Hub88, Isl81, Isl84a, Isl84b, Kam86, KM84a, KM82, KM84b, Le 89, LBY86, MRCL84, MG84, MW84a, MW84b, MW84c, May81, May84, MS84, MKS83, MKS84, MT84a, OTB80, OTB84, PM80, PM84a, PGB84, RR84, Ras80b, Ras84b, RSL^{+}84, SAC87, Sei84, Tak85a, Tal84b, Tal83, Tal84c, Ten83, Ten84a, Vaz81, Vaz84, VPB82, VPB84, Wil84a, Yat84]. Calculation [AP82, AP84, BB84b, Bla84, BM84b, DR85, DPH82, DPH84, EBS88, EMR84, GMR84a, Har81, Har84a, HPT80, JP84a, KM84c, LLC84, MS85, MRW84, Mik84, Nog83, OM83, OM84, PS84a, PT84, PS84b, Pie84b, Pre84a, RFSG84a, RG86, SP86, SP87, Sal84c, SS84a, SB82, TK84, TK80, WwYpCg88, WV80, WV84, ZNO88, ZC84, All84b, ACR88, And84c, BSK88, BMR85a, BAG^{+}87, Bie83, Bie84, Bin87, CWS83, CWS84, CC89a, CM88, CD86, CLS^{+}84, Coo82, Coo84a, Cop81, Cop84a, Cop84d, Cop84c, Cop84b, CVvWF86, DL87, Eld84, ENSC86, FV86, FHMR84, FFH84, FK84, Fer84, GPS84, Gel85, GMR84b, Han84c, Han84e, HMS84, HR87, HK81, HK83, HK84a, HK84b, IFI84, JBWW86, Jam86b, Kac84, KS86b, MS87a, MR80, MR84a, Mar84a, MZM^{+}82, MZM^{+}84]. calculation [Mor80, Mor84d, Nai84, Nai86, Nex89, NN84, NKS88, Pia84, Pic89, PP85, RFSG84b, RM89, RS83, RS84d, RS84e, RS84f, She84, Sta84, SW80b, SW84, SMT88, Ten85, Ten86, TC87, TM89, VSP86, WC84, WGM83, WGM84, ZDN84, vdB84]. Calculations [Dob84, Ede84, FT84b, LB82b, SG82, ST89b, AT84, AFH84, AF84a, AF84b, AXG87, AAH84, Aue84, BM82a, BM82b, BM82c, BCM83a, BM84a, BZ84, BDK86b, BDK86c, BT82, BES^{+}87, Bet84, BT84a, BP84d, CC89b, CRH^{+}89, Cio89, CB88, CM83, CM84e, DVD87, Del85, Des84a, Duc80, EP86, FKU^{+}84, FB84, Fle81, Fri88, Ger80, GLST89, Han84a, HD89, HA82, HA83, HA84b, HA84a, HvRM86, HSV84, IKRT89, ITR^{+}87, Jad84, Jam87, Kas82, Ken89a, KP81b, KPP^{+}89, LLA^{+}89, LM81, LCS84, Lux80, MMBW88, MV81, MV83, MV84a, MV84b, MGM87, MBW89, MPIM86, MB83, MB84, MR87, Net84, OCS86, Pon81, Poo83, Ric82, Ric84, Rys87, SW80a, Ská81, Ská84, SK86, TTAT81, TUB83, TUB84, TL84, TBB87, WB88, Wro82, Wro84, Zwa85, vdM84a]. Calculus [Har89]. calibration [BO80, BG81, Dob89]. call [Ano84b]. calorimeter [BGGW^{+}89, KPST89]. CAMAC [Ska82]. Cambridge [Bel84b]. Can [BM82a, BGMM81]. canonical [Kra86a, Ran87b, III88, SM85c]. Canterbury [GMR84a]. capture [AGF87, BGS81, BGS83, BGS84c, BGS84a, BGS84b]. carbon [RS83, RS84d, ZR89]. card [Cre80b, Cre84c]. cards [Fem84]. Carla [SM84a]. Carlo [AS89, ANP^{+}85, AM81, AMC83, AM84d, AMC84, BM82a, BM82b, BM82c, BMR84, BM84a, Ben84b, BI85, BS87, BKJ83, BKJ84, BDK86a, BDK86b, BDK86c, BLR82, BLR84, CMR86, CFG^{+}88, CM85, CGGK85, Cha85, CGG87, Chr84b, CMS83, CMS84, Cop81, Cop84a, Cop84d, Cop84c, Cop84b, CVvWF86, CGB^{+}89, CM83, CM84e, DMM84a, DLB83, DLB84, Dec89, Del85, DM87, DMM83, DMM84b, DKA85, DS84b, DD86, EMM81, EA83, EA84, Fer89, FS85b, GTM87, GHO^{+}89, Gup88a, HR84a, HR86a, HR86b, HP83, HP84, HMTO84, IW87, Ing87, Jad84, JW85, Kas84, Kaw86, KVW84, KSE86, LM81, Lux80, MMS87, Mar89a, Mar89b, Miu87, MGM87, MB83, MB84, MT86, Mor84e, NFH^{+}88, NAS87b, Odo81, Odo82, Odo84a, Odo84b, Odo84c, Odo84d, RR84, RSS85, RL80, Rit84a, Rit84b]. Carlo [RS83, RS84d, Ryn82, Ryn84, SMMP86, Sal85, San82, San84, Sjö82, Sjö83, Sjö84b, Sjö84a, Sjö86, SB87, Tat80, TKS^{+}85, Tou89, VH87, Voh88, Vrb89, Wan87, YMWW89, ZHR81]. Cars [SLH84, YB86]. Cartan [ADO83, ADO84b]. cascade [AV84, Miu87]. cascades [RR84, SV84]. case [GPS84, PS84b, ST89a]. cases [Red88]. Casimir [Bos83, Bos84]. Castor [CW84c]. casual [Pag87]. cavities [PJ89]. cavity [Kon84, MMF81, MPM83, MPM84]. CCD [AN88]. CCDEF [FNDL89]. CCFUS [DL87]. CDC [BM82b, BM82c, BCM83a, BM84a, BM86, BSDM88, Com80a, Com80b, Com84, CMO86, HRL84, KW84a, LS84]. CDF [CHS87, QAD^{+}89, Sli89]. CdS [JB80]. CEAN [IA86]. CELESTE [SG83, SG84b]. celestial [SG83, SG84b]. cell [BKR88, BLB88, FR84, GDK89, JP84a, MRT89]. CELLO [Beh81].

CELLO-experiment [Beh81]. cells [JB80]. Cellular [Suc87, Den88]. center [Dit89, Ebe84, GTL^{+}86, HB85, Yur84, HB85]. centered [CM85]. centers [Bra87b, Pou88b]. Central [Dru83, BGG^{+}89, MN84a, Nic89, Zac81]. centralized [BCC^{+}89]. centre [CDN^{+}87, GRRP83, GRRP84, Jun81, Jun84, MLL86, Rea84, Ste84d]. centrifugal [SAC87]. centroids [DK87b]. CERN [BBE^{+}81, BvEF85, CFG^{+}87, DGK^{+}81, Jac81]. certain [DVD87]. CFASYM [NN84]. CFP [Bra84e]. CFPJJ [Gra84a]. CFRX [HWL89]. CG [PZCC89]. CGM [Osl88]. CH [LH88]. chain [Kra86a, YP88, YP89]. challenge [Sch87a]. challenges [Her89]. chamber [AN88, BLMU81, BO80, BSP84, BD84d, BdCP81, Dob89, GM89, VKM^{+}82]. chambers [GDH^{+}89, Per89, TCS85]. change [Fem84, Rus87]. changes [MEB88]. channel [DL87, DP88, DRRvB82, FNDL89, HT88, HLS84, NK84, SDH84]. channel-spin-1 [SDH84]. channel-spin-1/2 [SDH84]. channels [RSA84]. Characteristics [DR85, DLB83, DLB84, Mir84]. Characterization [Hoc82]. characters [EC84]. charge [DE84, Jun81, Jun84, KLS^{+}89, Mal84, MT86, Mor80, Mor84d, NW84, Nob80, Nob84, PIM80, PIM84b, SS89, SNC80, SNC84, TDS81, TDS84, WW84b, vdB84]. charge-exchange [vdB84]. charged [CJ84]. Chebyshev [CWS83, CWS84, CM81a, CM84c, DNR83, DNR84, Del84a, Del84b, DM88a, Nex80b]. Checking [Woo85, Hib84a, Sax84a]. checks [NM84]. chemical [BvSW84, Eck84, Fia84, GG89, ITR^{+}87, OBLR84, Rob84d, Rob84e, Rob84f, SAC87]. chemically [HH82]. chemist [BvSW84]. Chemistry [SG82, BMV83, BMV84, DK82, SWHB88]. CHEP [Ano87d, Ano89-27]. Cherenkov [Män89]. Chichester [Jes87]. Chirikov [CM88]. chloride [ÖW84]. Choice [Fin86]. choices [JMS^{+}89]. Cholesky [KPP81, KPP84b]. Chow [Eas80]. CHPACK [DNR83, DNR84]. chromodynamics [NFH^{+}88]. CINDY [SR84a]. circle [CO84]. circuit [KK85, LJN85]. circular [HP83, HP84]. CIV3 [GH84b, Hib84c]. Cl [SAC87]. Clarendon [Gri87, Nad86a]. Class [EBS88, CL84, De 88, DGPM85]. classes [ADO83, ADO84b]. Classical [BBH^{+}84, BPW84a, BPW84b, BPW84c, FH88b, GH83, GH84a, Gup88a, KK83b, Lad86, Man87a, OM83, OM84, Sch87b]. Classification [Rid82, Rid84b, GSZ85]. classify [AI86]. CLEBSCH [RB84, SM84c]. CLEBSCH-GORDAN [RB84]. ClH [SAC87]. Clifford [Ber87b]. climatology [WB82]. clipping [BGMM81]. close [BES^{+}87, LLA^{+}89, Sar87]. close-coupling [BES^{+}87, LLA^{+}89]. closed [BZ84, CC89a, CW80, CW84a, CW84b, Lee82, Lee84, Ras80a, Ras84a, Sin84, Tro84]. Cloth [Art85a, Rob85]. clouds [BS84d, SL83, SL84]. clouds-computation [BS84d]. cluster [BH89, FAK82, HA82, HA84b, KM82, KM84b, Mue84, Mur88]. cluster-core [HA82, HA84b]. Clustering [You87, JSB82]. Clusters [BBE^{+}81, Bun86, KM84c, SCC^{+}88]. CMS [CHS87, McC87]. CMZ [BBR89b]. CNDO [CC89b]. CNDO-like [CC89b]. CO [DST80, DST84a, DST84b, Mag84b, Rob80b, Rob84c, RS84a, RRG88]. CO-N [Rob84c]. coarse [BSA86, Dec89]. coarse-mesh [BSA86]. coarse-to-fine [Dec89]. coaxial [Ods80]. Cochase [HLS84]. CODATA [Lid84a]. Code [KK89, Aue84, BMR84, BSdlT87, BHM81, BSV^{+}85, BLT81, Bor82, Bor84, Bor89, BBR89b, Buc83, Buc84, CMR86, CMPR88, CHS87, Cha85, CRPL81, CRL84, CRPL84, CW84c, CAR84b, CGMS89, CM86, CM81b, CH84d, CM84f, DL87, DP88, DTE84, DST80, DST84a, DST84b, DM89b, DH84b, DKA85, ETH84, ELS88, ET84, ERB88, Fem84, FB84, FNDL89, FSR84, GS89, GDK89, GA84, GTB^{+}81, GTR^{+}81, GTB^{+}84, HR84a, HR86a, HR86b, HL88, HLS84, HL86, Hof88, HWL89, IGB^{+}89, Kir87, Kon84, KLS^{+}89, LWHH82, LWHH84, LK88, LM85, LNS84, Mac89a, MVV83, MVV84, Mar89b, MPM83, MPM84, MN84b, MGK86, MBOK86, MMO89, Miu87, MMR89, MPM85, MP89, NK84, PD89, PM80, PM83, PM84a, PM84b, RSMTV88, Red88, Rit84a, Rit84b, Rit86, Rob80b, Rob84c, RSA84, SM84a, SS81]. code [SG81, SSF^{+}89, Shu87, SPM^{+}88, SL83, SL84, Syk81, Tal82, Tal84b, Tal84a, TSD84, TW89b, TTM^{+}80, Var89, Vaz81, Vaz84, WBG84, WB88, Wis87, vM84d]. Codebase [MP89]. Codes [DR82, HR87, Kil80, BES^{+}87, BT88, DSK84a, GTL^{+}86, HW84b, LLA^{+}89, MTN^{+}82, RBMC86, Rin84]. coding [CGG87, Wan87, ZHR81]. Codnum [Fem84]. coefficient [Bur84, RRC89]. Coefficients [Hub84b, AD84, All84c, AM84a, BSK88, BGMP84, Bie83, Bie84, CN84, Chi84b, DHN84, DW84, DH84a, FS84b, Gra84a, Gra84b, Gra84c, KS86a, Le 89, LB84, MRCL84, Mar89c, Nus84, Ogi83, Ogi84, PR84, PZCC89, Rao81, RV84, Rob84f, RSL^{+}84, RB84, SM84c, SG84a, SMR87, SK86, SW80b, SW84, SF84b, Tam84, TSD84, VVM84, YpMzCeCg85, YFL87, Zoh84]. coherent [Cop81, Cop84b, LBY86, SDH84, YB86]. coincidences [Hla80, Hla84]. coincident [IA86]. COJETS [Odo84a]. COLASE [Rob80b, Rob84c]. cold [BS84d, FBW82, FM84, SAVV87, SR84b]. Coleman [IB87]. collaboration [BLMU81]. collect [vM84f]. Collection [FV82]. collective [Szy81]. colliders [Rit86]. colliding [MS84]. Collision [KM86, Sar84a, BKvLSE80, BKvLSE84, BT82, Cla84, CB88, Cre80a, Cre84a, CSW84, Das89, Gal84, HBJ84, Hof84a, Hub84c, Jam87, KS86b, MEB88, NM84, NA84, NN84, OS84, Sal84c, SK82, SK84, WW80, Yat84]. collisional [Nov87]. collisions [Arn84b, BBH^{+}84, Bar83, Bar84e, BGS81, BGS83, BGS84c, BGS84a, BGS84b, BDK86b, Bil84, Bil87, BW87, FR86a, GDH^{+}89, GPS84, Irv82, KVW84, MEB88, Nob80, Nob84, OM83, OM84, Par84, Pia84, PS84b]. colloquial [PT85]. Collrad [Tal84b]. color [CN87, Dun87, KR88]. Colour [Ano87j, BP84d]. COLRAD [LHP87]. COLRAD-A [LHP87]. columns [MCA87]. combination [Wil84b]. combinatorial [HR87]. combined [HMS84, KBMV86]. Cometary [IP88, MV88, SWHB88, SW80a]. Comm [Ran87a]. Comment [FV81, BD81, Jam88]. Comments [Gup88a, Mag81, MF81, MN84a, Sch86c, TT85]. Commercial [Kri89]. commercially [MP89]. committee [Ano84-86]. Committees [Ano80d, Ano81e, Ano82g, Ano82h, Ano86e, Ano88g, Ano89i, Ano88w]. common [Klu84b, CT89, GSS81]. Commun [KS08, Tho04a, Tho04b]. communications [ABLB^{+}89, Cis87, Ano80e, Ano80g, Ano80f, Ano82i, Ano86f, Ano88h, Ano88i]. community [McC87]. commutators [DFD81, DFD84, SM85d]. commute [DFD81, DFD84]. Compact [TS84b, CMS83, CMS84, DM87, KKP82, KPH84, KKP84, Nas84, Rap81a, Vis84a]. Compactified [DR85]. comparative [Sta87a]. compared [CT84b]. comparing [ABMT84]. Comparison [FBS88, HVS87, IR85, KPP^{+}89, PT82, AMB84, BES^{+}87, KJ85, LLA^{+}89, Whi89a]. compatible [BSK88, Bas88, Jud87]. compilers [CKV85]. Complete [BDK86b, BDK86c, Bas87a, Bas87b, Han84a, JP80, JP84b, Vaz81, Vaz84]. completeness [AI86]. Complex [BCM83b, BCM84b, TTW84a, AM84b, AM84c, BD80a, BD84a, BH85a, Cam81, Cam84a, Cam84b, De 84c, GS83, Jak84, KF84, KLS85, KB86a, Köl84c, KB86b, Man87b, Mas83, Moo81b, Moo84b, Nus84, SF84b, TTW84b, TR84, TB85, TB87, Tho04a, Tho04b]. complexity [Nas87a]. complicated [AB88, CWS83, CWS84]. component [O’C84a]. composition [MEB88]. compositor [HR83b, HR84c]. compound [Fia84, HMS84, SM84a, SS84b, SMD84, SR84a, SDH84]. compound-nuclear [SDH84]. compounds [KN84c, KN84b, MV83, MV84b, NJ82, NJ84]. compressible [Bra87a, Cro88]. Compton [SS84d]. Comput [KS08, Ran87a, Tho04a, Tho04b]. Computation [BGS81, BGS83, BGS84c, BGS84a, BGS84b, Bos83, Bos84, BS84c, Cug84, Deu84b, Deu84d, Eğe82, Ege84, HIM84, Hei84, Jun81, Jun84, Kno84, LJ84, LB84, Mal84, MSFR89, Rao81, Ros85, SS84b, SMD84, SR84a, SR84b, SM81a, SM84d, SM85b, AB88, BSdlT87, BS84d, Buc83, Buc84, CMR86, CGM89, Cli82, Cop81, Cop84b, DG84, DE84, DK82, FCW84, Gue89, Han84c, HC86, HSV84, Hub89, Key89, Klu84a, KMS88, KLS^{+}89, KB86b, Lab84a, Len80, LU83, LU84, ML81, New86, NF83, NF84, Ost85, Pet84, Pie82, Pie84a, Pol88, Reg81, SMR87, Ste84b, VČ88, ZNF87, Zoh84]. Computational [Ano89h, Bra86, DGO83, Eri85, HW84a, Jes86, MPR80, Mül87, Nex80a, Sch86a, Sta87b, SAC^{+}86, Wie86a, vD87b, And86, Ano81t, Ano84-87, DGL89, Eas80, Eas84b, GA84, Hey89, MN80, Nad86b, Nex80b, SMT88, Ano89h, Eas84a, Lis88, Sta87b]. computations [BPR^{+}86, CCD85, Kla84, Lab84c, OCC80, OCC84, SG83, SG84b, TL89, YSHK85]. compute [Câm82, Câm84g, Chi84a, Chi84b, GTM85, GTM87, HC83, HC84, MM88b, STN89, Tal89, TBG80, WB84]. Computed [OS86, Klu84b]. Computer [AP89, Ano80e, AGG^{+}82, Art85b, AWG84, BK84a, BR85, CT84a, Chr84a, CCS89, DBB84, Das89, FR81, Fre87, FOW87, Gro84, HB88, Kil80, KR81, Laa86, LB82b, Mak80, MMS86, MGTD84, PZCC89, Rap84, Taj86, WD84, WP84, Ano87b, AGF87, BB84a, Bas87a, Blo82a, Blo82b, Boc89, BPSB88, BA86, Bun86, BMBW85, Cha82, Cha84c, Cis87, CLS^{+}84, DCC84, DL86, DSK83, DSK84b, DM87, Dit89, DRD88, Eas89, EM84a, EM84b, Eks84, Fer89, Fle81, Gel85, Ger80, Gio84, Gli84, GWTM83, HH82, Han84a, Han84b, Han84c, Han84e, Her89, Hoa89, HVS87, HMTO84, HT84, Hub88, IKRT89, ITR^{+}87, IHS^{+}88, JB80, JM82, Jud87, Kac84, Kas82, KKLS82, KP81b, LWHH82, LWHH84, Le 89, LJN85, Mir84, MC81, MC84b, Mye88, PR84, PM83, PM84b, RWK^{+}85, RSMTV88, Rob87a]. computer [RD80, RD84, ST84d, Szé85, Tar86, Tol85, VM89, Vie88, VSH83, VH87, WMH82, WDM84, YSHK85, Zac81, Ano80g, Ano80f, Ano82i, Ano86f, Ano88h, Ano88i, Eas89, Fin87]. computer-aided [JB80, Vie88]. Computer-algebra [FOW87]. Computer-assisted [AWG84]. computer-controlled [BPSB88]. computer-generated [HB88]. computer-graphics-based [VM89]. computer-simulated [WDM84]. Computers [EBS88, And86, And88, Bas88, BTS84a, Com80a, Com80b, Com84, CKV85, DTE84, Dic82, ETH84, Eas82, EA89, Fla82, Gen82, Hoc82, JMS^{+}89, KC81a, KC81b, KC84a, KC84b, Kas82, LT85, MSC^{+}88, Mat87, NA84, Nev82, OY80, Qua87, Sch89c, SSM85, Sch85c, TL82, Tru88, vdV89, vM84d]. computes [MMO89]. Computing [AN87, Ano81f, Ano88-27, AHS^{+}86, Boc89, KLS85, New87, O’C84a, Szy81, Wil89, AHKM86, BD80a, BDM81, BDM84, BD84a, BDJ^{+}89, Bra87b, Cam86, CGGK85, Dow82c, Goe84, Han84d, Joh87b, Köl81b, Köl84b, Köl84c, KR88, Lao84, LNB^{+}88, Mac89a, MPM83, MPM84, Nag89, Nas89, Nic89, Pat82, RS81, RS84b, Sch87a, SNC80, SNC84, Sch84d, ST89a, Tel84, TV89, Tru88, Vio84, Wat89, Wic87, Wie86b, WH89, Zac81]. concept [Com80a, Com84, KC81a, KC84b]. Concepts [Tau85, Sch86b, Sch88b]. Concluding [Gen88]. CONCON [Buc83, Buc84]. Concurrency [PBKW85, ACSB85]. Concurrent [BBC^{+}89, MMBW88, Cis87, FS85a]. Conditional [EBS88]. conditioning [DeG88]. conditions [BDK86c, Bun86, Cam84f, Kas82, LM84, MP81a, MP84a]. Condon [All84a, DK87b, TT82, TT84]. conducting [DM89a, VM89]. conduction [GH83, GH84a]. conductivity [EK86]. Conference [Boc89, Bra86, Gri87, Wic87, Ano82w, Ano87d, Ano89-27, Ano89a, Nas87a]. configuration [BZ84, CDW82, CDW84, CLS^{+}84, DK87c, FKM^{+}87, Fro84a, Fro84b, Fro84c, Hib84a, Hib84c, HFG88, HWL89, KH89, Poo83, TBG80, TTL88, WH89]. configurations [BDD84, Cha84e, CO86, GPS^{+}81, Hoe86a, Rid84a, TS89]. Confined [Sta87b, MKRJ89]. confinement [KP84, MPM83, MPM84]. Conformal [Goe81]. confront [BMG^{+}85]. conical [Köl81b, Köl84b]. conics [CCR88]. Conjugate [GLC89, KPP81, OJK89, AKS88a, AKS88b, BMR85b, EA89, KJ85, KPP84a, KPP84b, Meu89, Oya86, Sun88]. connected [KS89, Rap84]. connection [AS89, CKT89]. Connectionist [TA89]. Conserved [IK85]. considerations [Lab84b]. consistency [NM84]. consistent [BLT81, CCR84b, Des84a, LCW84, SSS84, WB88]. constant [Han84f, ONP89, SNH^{+}88]. constants [Bil84, Bil87, ENSC86, Gel85, JP84a]. Constrained [Sha87, Buc83, Buc84, LM81, Pro82b, Pro82a, Pro84]. constraints [HVS87, Zlo89]. Construction [SM81b, SM84e, SM85c, Bri86, Hla80, Hla84, HM84e]. Consultant [Kil80].

Contact [EFK85]. contained [AI88, BM84b, IA80, IA84]. containing [Jac85b, Jac85a, MR84b, Rob84f, WSV88]. content [Köb89, Rob84d]. Contents [Ano80i, Ano80j, Ano80h, Ano81h, Ano81i, Ano81j, Ano81k, Ano81g, Ano82m, Ano82n, Ano82o, Ano82j, Ano82k, Ano82l, Ano83e, Ano83f, Ano83g, Ano84i, Ano84j, Ano84k, Ano84l, Ano84g, Ano84h, Ano85h, Ano85i, Ano85j, Ano85k, Ano85f, Ano85g, Ano86j, Ano86k, Ano86l, Ano86m, Ano86g, Ano86h, Ano86i, Ano87k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano88m, Ano88n, Ano88o, Ano88p, Ano88j, Ano88k, Ano88l, Ano89l, Ano89m, Ano89n, Ano89j, Ano89k]. CONTIN [Pro82b, Pro84]. continued [Lee89, Nex84, TB85, Tho04b]. continued-fraction [TB85, Tho04b]. continuous [ABG^{+}81a, CCR84a, CM81a, CM84c, CSC^{+}86, Her82a, Her84b]. Continuum [AT84, AXG87, Bat84, BBC^{+}84, BBL^{+}84, BS84c, DF84, HRS81, HRS84, IR80, Le 89, ST82a, ST84c, TUB83, TUB84, TBB87]. contracted [Tau84a]. contribute [JMS^{+}89]. contribution [AP82, AP84, Nog83, TBG80]. contributions [Mue84]. contributors [Ano82x, Ano85u, Ano88x, Ano88y]. Control [Gou88, And88, AGG^{+}82, BCC^{+}89, CR84d, CM84d, DGL89, DSSS86, FS82, FLM84, HRL84, LEF^{+}89, Moh80, Moh84, Rap83, Rom81]. controlled [BPSB88, VKM^{+}82]. controllers [Ska82]. conventions [HR83a]. Convergence [TTM^{+}80, BBR89a, BT84a, Hub84f, Lan80]. conversational [HHK86]. Conversion [Gol84a, AHVV86, DHN84, DH84a, PR84, PS88, RD80, RD84]. Converter [Sal84a]. converting [Gro89b]. convex [AN87]. convolution [DM88a]. Cooling [Hoe86a]. Cooper [MVV83, MVV84]. cooperation [Lid84a]. coordinate [BP84d, De 84a, Moo83, Var89]. Coordinates [Eas87a, Har89, AL81, CH88, Cla82b, DW89, Dav81, Dav84, FH88a, Han84a, Hug84, KP84, MT81, MR84b]. core [Bat84, BZ84, CCR84a, Fre87, HA82, HA84b, HH84, Hir84, LS84, Moh89, Joh89]. CORE-based [Joh89]. Coriolis [HK81, HK83, HK84a, HK84b, Kac84]. coronal [Duf84]. corrected [EFES83]. correction [Deu84a, Deu84b, Deu84c, Deu84d, Han84b, KP89, PP85, Rud84, Ver82]. corrections [BKJ83, BKJ84, BDK86a, DHN84, FKM^{+}87, JW85, MGN80, MGN84]. correctness [DK87c]. corrector [BP82a, BP82b, BP84a]. correlated [DKA85, WD84]. correlating [KS81, KS84b]. Correlation [BD87, DFD81, DFD84, HSV84, KS84a, KEMP81, KEMP84, MRCL84, RSL^{+}84, WDM84]. correlations [AB83, CMS83, CMS84, Eks84, HD88, JR84]. correlator [KMS88, Män89]. corresponding [Sch84a]. cosmic [MMR89]. cost [JHRR87, Mye89]. Costs [CNRS82]. COULCC [Tho04b, TB85]. Could [JMS^{+}89]. COULFG [Bar82, Bar84b]. COULN [NT84]. Coulomb [Tho04b, BDG^{+}84, Bar81a, Bar81b, Bar82, Bar84b, BFSG84, Bar84c, Bar84d, BS80, BS84b, BH88, DF84, FB84, HKT85, LJLC84, LT87, MC85, MKS83, MKS84, Nes84, Nes88, NT84, Pia84, Pol88, RSA84, SS84a, Sea82, Sea84b, Sea84c, SF84b, TTW84b, TR84, TB85, Tol86]. Coulomb-nuclear [FB84]. counter [MMF81]. counter-beam [MMF81]. counters [Män89]. Counting [Aur84a, Aur84b, EA83, EA84, AGF87, GTG81, GTG84, GTM85, GTM87, GTMA88]. Coupled [GHU81, AKS88a, AKS88b, Bai87, BBM82, BBM84, BP82a, Cha84d, CW80, CDW82, CW84a, CDW84, Chi84a, Cla84, Cla82a, Cre80a, Cre84a, CSW84, DL87, DP88, DRRvB82, FNDL89, GPS84, Gro87, Gro89a, HD86, HBF^{+}85, HT88, HH84, Hir84, HLS84, Hub84a, Mor84c, NK84, NN84, Nor84a, Pia84, RSA84, SK82, SK84, Tol85, WMCH84]. coupled-channel [DL87, FNDL89]. coupling [BES^{+}87, Bla88, CN84, Cha84e, Cla84, DW84, Dya86, EPS81, ENSC86, Gra84a, Hub84a, Hub84b, KBMV86, LLA^{+}89, LJN85, Nus84, PGB84, Rid82, Rid84b, Rud84, Sar87, SCF86, SW80b, SW84, Tam84, VVM84, WS82, WS84a]. couplings [AMC83, AMC84]. covariant [KMS88]. cover [Eas89, Fin87, Gri87, Tru88]. CPC [Ano86n, Jac80, MLL86]. CPDES2 [AKS88a]. CPDES3 [AKS88b]. CRAY [KJ85, RS85, Dic82, DR82, EBS88, FR81, Fin82, Fin84a, KM85, LB82b, NR86a, PCM85, SG82, Sch89c, SB82, THH82]. CRAY-1 [KJ85, Dic82, DR82, FR81, Fin82, Fin84a, PCM85, SG82, SB82]. Cray-1s [THH82]. create [Rob84f]. created [TDS81, TDS84]. creation [CM84a]. Cross [FGK89, All84b, AMK81, AMK84, BKvLSE80, BKvLSE84, BGS81, BGS83, BGS84c, BGS84a, BGS84b, Bet84, Bie83, Bie84, Bil84, BTS84a, BTS84b, BW87, Cap86a, CLS^{+}84, DL87, FNDL89, Har81, Har84a, Hei84, Hen84, HMS84, JT84a, KS86b, LW84a, LJLC84, MMM84, MPR80, MT84b, Nov87, OM83, OM84, ONP89, OTB80, OTB84, PDNE88, PV84, PS84b, RS81, RS84b, Sal82, Sal84e, Sar84b, Sar84c, Sar87, SAC87, SS84b, SMD84, SR84a, SG89, Tar86, Ver89, WL84, WMCH84, vdB84]. cross-assemblers [Tar86]. cross-section [Bet84, LW84a]. cross-sections [BGS81, BGS83, BGS84c, BGS84a, BGS84b, Har81, Har84a, OM83, OM84, PS84b, Ver89, vdB84]. crossings [Vio84]. crystal [Ber84a, DRD88, FR84, Han84d, MS85, MS87a, MS88, PS84a, Pre84a, SP87, Ste84a]. crystalline [BT84a, Han84b, Han84c, Han84e]. Crystallographic [Bel84b, CA84]. Crystals [Ric84, BMS87, BL86, CLL89, De 84a, Fre87, Han84a, Han84f, LB84, Ric82, Sal83a, Sal84d, SF84b, WD84, WDM84, Wor84a]. CsCI [Ano84-83]. CSDUST3 [ELS88]. Cu [JB80]. Cube [FAK82]. Cubic [Ric84, AAMB84, BMS87, BDD84, Bok84, DS83, DS84a, FAK82, LB84, MV81, MV83, MV84a, MV84b, PDF82, PDF84, Rei87, Ric82, WD84, WDM84, vM84i]. Culham [Syk81]. Cumulative [SS89]. Current [CHL^{+}81, AHS^{+}86, OSF86, SSF^{+}89, Uga88]. current-drive [SSF^{+}89]. Current-driven [CHL^{+}81]. currents [Lee82, Lee84]. curve [CAP86b, HE84b, Puf83, Puf84, vM84a, vM84c]. curve-fitting [CAP86b, vM84a]. curves [ASM80, ASM84, AMB84, De 84c, KEMP81, KEMP84, RS84e, Tel84]. Curvilinear [Eas87a, Har89]. CW [Smi84a]. CYBER [BM86, BM82b, BM82c, BCM83a, BM84a, BSDM88, CMO86, KW84a, HD86, HVS87, VH87, BA86, Car89, Dic82, Pur82, Tol85, VSH83, VBD88]. cycle [Tau85]. Cyclic [Joh85]. cyclotron [ABD82, ABD84, CCR88, MKH86, McN81, McN84, MGK86, SSF^{+}89]. cylinder [Cop84c, Kir87, MKRJ89]. Cylindrical [Mas83, AL81, CD86, ELS88, Hug84, Jad84, Kar80, Kar84a, KLGT85, KB86b, MBOK86, Pie84b, TTAT81, Vrb89, Wie86b]. cylindrically [SF84a].

D [Art85a, Eas84a, Eas89, Fin87, Gri87, Lis88, MN84a, MN84b, MZZ89, Rob85, Sch86a, BLR82, BLR84, ELS88, And84b, And84a, Atz86, Car89, CN87, Cro88, Eas87a, ERB88, GS89, HL88, JN85, KT81, MCJ81, MCJ84, Moh89, Ods84, RBMC86, SS81, Syk81, Via87, Wan87, vdV89]. D0 [Ano87-29, CHJZ89, CHS^{+}89]. damage [Cha84b]. DAP [Par83, AWB82, BM85b, Del85, DD85b, Hel82, JN85, LB82a, MP82, Par82, RSS85, SRS85, WB82]. Daresbury [HSB^{+}84, THH82]. Data [Bel84b, Ber84a, CHJZ89, Eld84, Jas84, KS87a, Mee84, OLR^{+}87, Poh87a, Put89, Rea89, Ros84a, SG89, Tha89, AI88, AAMB84, AMK81, AMK84, BGMM81, BD80b, BD81, BD84c, BvSW84, Bro83, BCVZ89, Bur87, CR84b, CM84a, CMH87, CM84b, CDF^{+}89, CSMHT88, Com80a, Com80b, Com84, CH84c, CP81, CP84, Del87, DS82, DK82, Dun87, Ezh84, Fia84, FP89, FM85, FV81, FMS87, Gai89, Gau81, Gau84b, Gau84a, Gau84c, Gli84, Han84b, Han84c, Han84e, Har84b, Her84a, Hib84a, Hin81, HSA82, HSA84, IA80, IA84, IA86, IL84, Jac85b, Jac85a, Kal83, KC81a, KC81b, KC84a, KC84b, Kat84, KM84a, KR81, Klo82, Klo84a, KN84a, KRS83, LCS84, Lid84b, LBK88, Mas81, McI84, Mir84, Moo84d, Mor84a, NM84, OY80, PF86, PM80, PM84a, PS88, Pro82a, QTZ^{+}87, Rea84]. data [Rim81, Rob84f, Ros84b, RHM^{+}89, Sal84a, SEF^{+}89, SN84, Sax84a, Sch89a, SSM85, SBB^{+}89, SM89b, Sko89, SdAD^{+}88, Ste84d, Tep84, TTT87, Tor82, Tor84, Ver84b, Ver89, Whi89b, Wil89, WC85, Woo86b, WSV88, Yat84, vMF81, vMF84, vM84g, vMT84, vM84e, vM84i]. Data-acquisition [Tha89, CDF^{+}89, SEF^{+}89, Whi89b]. data-base [Put89]. data-taking [Wil89]. Database [Mou87, TTT87, HS84, HSB^{+}84, MS87b, Rus86, Sko89, Ste87, Wha89]. databases [LSW87]. Dataflow [Gur85, YSHK85]. Datstor [Rob84f]. Davydov [AFH84]. DBLCON [War84]. dc [ABD84, OSF86, ABD82, EK86]. DCSCH3 [FC84b]. DCSCH4 [Ano84-84, FC84b]. debugging [Sch85b]. Debye [Rei87]. DEC [CHS87, Com80b, KC81a, KC81b, KC84a, KC84b]. DEC-10 [KC81a, KC81b, KC84a, KC84b]. decay [AB83, BMG^{+}85, FC84b, HR84a, Hla80, Hla84, IL84, Kal83, KL84b, PIM82, PIM84a]. decays [LO86]. December [Kil80]. decision [Fra81]. decisions [Chi89]. deck [Fem84]. decks [BJ81a, BJ81b, BJ84a, BJ84b, Jac85b, Jac85a]. DECNET [Ghi87]. DECNET/IBM [Ghi87]. Decomp [Wie86a]. decompose [RWZ87]. decomposition [Key89, KPP84a, MSC^{+}88, MP82, Oya86, Ran87b, Ryn82, Ryn84, RF83, RF84, Zlo84d]. decomposition-conjugate [KPP84a]. Deconvolution [DS82, WSV88, BD80b, BD81, BD84c, Ber80, FV81, KRS83]. dedicated [Gli82, JMS^{+}89]. deep [IR80]. defect [KM84c, Ric82, Ric84]. defects [BAG^{+}87, DBB84, LB84]. deferred [GB87, Jam88]. defined [Car84a]. definite [Gad85]. deformable [NJ82, NJ84]. deformation [KN84c, PI84]. deformation-dipole [KN84c]. deformations [FNDL89, Hir84]. Deformed [HH84, CDN^{+}87, HR87, KP85]. degenerate [MW84a, MW84b, MW84c, Wil84a]. degradation [LNV84]. degree [Dob89, Jes82]. degrees [SK82, SK84]. Dekker [Art85b]. delayed [PDNE88]. delegates [Ano82y, Ano85v]. DELPHI [BGG^{+}89, Ber87c, ABLB^{+}89]. Delphine [ABD82, ABD84]. Delsqphi [CH84a]. Delta [Eks84]. Densities [IK85, BD87, DP84, HR87, LHP87, Mor80, Mor84d, PIM80, PIM84b, Rob84e, Tal84b]. density [Chr84a, Chr84b, De 88, KS84d, LZ84, Pol88, SNC80, SNC84, ZL84]. DENTS [CSC^{+}86]. dependence [Fer89]. dependent [CMR86, IK81, GS89, Has84, HLM^{+}87, LZ84, LW84b, Mag81, Mag84a, MV87, MF81, MC84a, Ran86, Ran87a, RWZ87, ZL84, Irv82]. DEPOS [LW85]. deposition [DM86a, LW85, MRW84, MM84a, PSBG84]. Depth [MM84a, DM86a, MRW84]. Derivation [He84a, DGPM85, FR84]. derivative [Buc83, Buc84, DK87a]. derivatives [Bar82, Bar84b, JP84a, KK83a, Lee82, Lee84]. describing [SS89, SK82, SK84]. description [Ano87c, GHU81, WP84]. Design [DSSS86, KC81b, KC84a, RHM^{+}89, Ano81f, BF89, Buc85, Car88, CFG^{+}88, CMH87, Dow82b, Eck84, FP89, Gat89, Hoa89, Kas82, Poe85, Reg81, Ric86, Suf89, Vie88]. designed [DTE84, DRD88]. Designing [SSM85, ZS87]. DESY [Beh81, Sch89b]. detailed [BF89, SWHB88, Sea84a]. Detecting [Kal82]. detection [Bir85, Kal83, TN84]. detector [Ano87c, BGGW^{+}89, CFG^{+}88, CGMS89, HAG^{+}89, MS87b, YMWW89]. detectors [Boc87, Kan87, OVQT^{+}84]. determinant [KH89]. determinantal [Lab84a]. determinants [Fra83a, Fra84b]. Determination [Ata85, Bas87b, Cam84e, Cam84d, CF85, FK86a, Fra83a, Fra84b, Rid84a, RRG88, WDM84, CH87, CH88, GPS^{+}81, RP89b, RP89a, RD80, RD84, VH87, DG84]. determinations [DRD88]. Determining [EFK85, Sch82b, Sch86d, Bas87a, Sch88a, Sch84c]. Deutsch [FV81]. developed [DU89, MP89, WH81]. Developing [BvEF85, Cam85, Man87b, Uga88]. Development [BGM87, Kat84, KP89, Kel87, Ano87-29, BM86, BCMR87, Bri86, BCC^{+}85, Car80, CNRS82, HGJ^{+}89, IKM85, Krá86b, Nie88, Pla86, Sar84c, Vie88, Was85, Wis87]. Developments [Syk81, Ken89b]. deviations [Kal82]. Device [BCD^{+}89, Dew82, VKM^{+}82]. devices [BO80, CO86, KP84]. Dfl [Art85a, Bra86, Rob85]. Dfl.180.00 [Tru88]. DFPOT [Coo82, Coo84a]. DIAGEN [SUZ89]. diagnosis [BC87, BCF^{+}89, KK85]. diagnostics [KB86a]. diagonal [GMR84b]. Diagrammatic [MBW89, Sil84a, Sil84b, WS80, Wil84c]. diagrams [CW80, CW84a, CW84b, KR88, PT85, SUZ89, Stu88, WS80, WP84]. diamond [Wie85]. diatom [Bil84, Bil87, Nov87, ONP89]. diatom-diatom [Bil84, Bil87]. diatomic [BB84a, BOT86, Bou87, DK87a, DK87b, FT84a, Gia84, Lad86, MLL86, Nai84, Nai86, Ras80a, Ras80b, Ras84a, Ras84b, Sal84c, ST81, ST84a, SLH84, Sin84, TT82, TT84, Tel84]. dicon [TKS^{+}85]. dielectrics [LVPG84]. difference [AZK83, CSC89, Dru83, GV87, GS83, Kal82, Lis84, Man87a, MN84a, Sch84d]. differenced [Wie85]. different [PIM80, PIM84b]. Differential [Art85a, EFK85, IK85, LJLC84, RW86, Sch82b, Sch86d, AKS88a, AKS88b, AMK81, AMK84, BBM82, BBM84, BTS84a, BTS84b, Car84a, Cha84d, CSW84, DGL89, FK86a, HD86, Jes84, Mar89a, MPR80, Mor84c, OM83, OM84, OTB80, OTB84, Par84, PS84b, RWK^{+}85, Rud84, Sal82, Sal84e, Sch84c, SS84b, SMD84, SR84a, SK82, SK84]. differentiation [BKK86, GJ84, MP81b, MP84b, Wai87, WSV88]. differing [Fun84a, Fun84b]. diffracted [TK84, TK80]. diffraction [Bla84, FR84, Gru84d, Han84b, Han84c, Han84e, MGMZ84, RP89b, RP89a, vM84i]. diffractive [Kas84]. diffuse [SP86]. Diffusion [EFES83, RBMC86, BH81, BH84a, BSA86, BLT81, Câm82, Câm84g, CRL84, Del85, Sch87b, WD84, WDM84, vMF81, vMF84, vM84g]. diffusive [ACR88, WBG84]. digamma [Köl84c]. Digital [FS82, BGMM81, Dew82, Ruf85]. digitized [SCC^{+}88]. digitizing [BdCP81]. dilute [CGG87, OS84, Tho84]. diluted [NHK82]. dimension [HLM^{+}87, IKRT89]. Dimensional [CW84c, ABG^{+}84, AWB82, AK84, BC88a, BCM83a, BWD84, BL88, CR84c, Cho84, CRPL81, CRL84, CRPL84, CH84a, CAR84b, CMM84, CM86, CMO86, De 88, DM87, DM89b, DH84b, EHL80, EHL84, IK81, GJ84, GTL^{+}86, Goe81, Has84, Hay89b, HL86, HvRM86, Hou87, HWL89, KK88, KJ85, KR81, Kir87, Lan80, LK88, Lew82, Lew84a, LNS84, LB82b, Luc84, MMS87, MMK81, MMK84, MKH86, MBOK86, MMTK88, NR86b, NRS88, Ods80, OT80, OT84, Per89, Phi80, Pis84a, Pis84b, RSMTV88, RR84, RC85, RG86, Sch84d, SM89b, She84, SPM^{+}88, TW89b, Tre88, VČ88, WBG84, Wie85, YP86, Zlo84b]. dimensionalities [DV85]. dimensions [AKS88a, AKS88b, BSDM88, DP84, HH82, Sch81, SM82, SM84f, Wai87, YFL87]. dinuclear [Ogi83, Ogi84]. DIP [SM89a]. DIP-moveout [SM89a]. dipolar [LVPG84]. Dipole [DF84, Hib84c, KN84c, Sho81, Sho84]. diquark [Rit84a]. DIRAC [DSK84a, BDM81, BDM84, Des84b, FR86b, GMN^{+}80, GMN^{+}84, IKRT89, KMS88, MGN80, MGN84, Pic89]. Direct [Klu84a, BH85a, CC89a, FS84a, HL88, MG84, RWZ87, SC89]. directional [MRCL84, RSL^{+}84]. directions [Pat82]. Director [MG84]. Dirichlet [MP81a, Cam84f, MP84a]. DIRIGE [DK87a]. discharge [Abr88, DE84]. discharges [DEW84]. discontinuities [Moo84d]. discrete [CRS87, CCR84a, Dav81, Dav84, Hel80, Joh87b, NR86b, RA82, RA84, Wie85]. discretization [AS83, And83b, And83a, AS84, And84b, And84a, SA83, SA84]. discretizations [KA87]. discretized [KP89]. discussion [Ano88-35, WBV^{+}89]. disk [RHM^{+}89]. disk-based [RHM^{+}89]. dislocations [Pet84]. Disordered [Vir89, MZM^{+}82, MZM^{+}84, SP86]. dispersion [EFES83, Han84d, New86]. displacement [Cha84b, DS83, DS84a]. display [BGMM81, CGMS89, PDF82, PDF84, SdAD^{+}88]. dissipation [BKR88, Eas88]. dissociation [RS83, RS84d]. dissys [Sch86c]. Distorted [TCR84, BKvLSE80, BKvLSE84, DE84, SAC87, WV80, WV84, ZNO88, vM84h]. distortion [HE84b, May81, May84]. Distributed [Jes87, Nas89, Whi89b, Dow82c, JSB82, KSG82, Ott89, SM85a, Ste87, Par83]. distributed-memory [Ott89]. distribution [BB84a, BRTB84, BRT84, DM86a, DL86, Han84c, Han84e, Jun81, Jun84, KS84d, KS84e, KS08, Lan84b, MRW84, MM84a, MM88b, Pie84b, Rib80, RSL^{+}84, SS89]. distributions [AB83, Les89, Sal87, SAC87, Sch84a, SDH84, TC87, TSD84, ZC84]. divergence [RM89]. DLXANES [VSP86]. DM [Ano89h, Eas84a, Eas89, Fin87, Man86, Sch86a, Sta87b]. Dnestrovskii [Sch86a]. Do [Jam86a]. documents [BF89]. Does [Her89]. Domain [Key89, MSC^{+}88, AS89]. domains [KS89]. dominated [Cli82, WW80]. Domus [Zlo84b]. Dordrecht [Art85a, Rob85]. dose [Hub84c]. dosimetry [Hub84d]. double [Coo82, Coo84a, HEL84, He84a, MSFR89, SS84b, SMD84]. double-differential [SS84b, SMD84]. doubled [Ata85]. doublet [KK84b]. drag [WHW87]. drawings [PDF82, PDF84]. Dresden [AGG^{+}82]. drift [Dob89]. drifting [FM84]. drive [AHS^{+}86, OSF86, SSF^{+}89]. driven [BVL84, CHL^{+}81, Poh87a]. Drooling [JSB82]. drop [PI84]. Drummond [MN84a]. DSAM [MC81, MC84b]. DST [Blo89]. Dubna [BLMU81]. Ducksbury [Jes87]. due [Lee82, Lee84, May81, May84, McN81, McN84]. Duff [Bra86]. Dumbell [Lad86]. Durham [Wha89]. during [OBLR84]. dust [SL83, SL84]. DUSTCD [SL83, SL84]. dusty [ELS88]. DVSR [Gli84]. DWBA [But84, PV84, TUB83, TUB84, TL84, TUWA84]. DWPI [EM84a, Fun84a]. dye [BLB88]. dye-laser [BLB88]. dynamic [AF82, AF84c, Bin87, DSK84a, Fin82, Fin84a, MEB88, Rin84, Sho81, Sho84]. dynamical [CMPR88, HD88]. Dynamics [FR81, LH88, Ber87a, DGL89, Fin80, Fin84b, Fin86, Gad80, Gal85, GV89, GG89, GDK89, Gup88b, HSV84, HVS87, HT84, JL84, KN84c, KN84b, MKRJ89, NJ82, NW84, NJ84, Ods84, RFS89, RMV89, SBM^{+}86, Sch89c, SM86, Ter88, VSH83, WW84a, ZR89, Ano89h, And86]. dynamo [Moo84e]. Dyson [BCCM87]. DZDISP [Sch89a].

E1 [HK83, HK84b]. E2 [HK83, HK84b]. E687 [MS87b]. Early [RS85]. earth [Moo83, Swi88]. EBEGA [GTMA88]. echo [vMF81, vMF84, vM84g]. economic [McI84]. Economizing [AXG87]. Ecsimpact [LS84]. ed [Rob85]. edge [DPH82, DPH84, Pet84, VSP86]. Edinburgh [BBC^{+}89]. edition [BPW84a, BPW84c, Duc86]. Editor [Bas87d, Cre84c, Day84b, Tau84a, Cre80b]. Editorial [Ano81m, Ano81n, Ano82p, Ano82q, Ano82r, Ano82s, Ano83h, Ano83i, Ano83j, Ano84m, Ano84n, Ano84o, Ano84p, Ano85l, Ano85m, Ano85n, Ano86o, Ano86p, Ano86q, Ano86r, Ano86s, Ano86t, Ano87q, Ano87r, Ano87s, Ano87t, Ano88q, Ano88r, Ano88s, Ano88t, Ano89o, Ano89p, Ano89q, Ano89r, Ano89s, Bur85a, Rob80a, Rob81, Ano80k, Ano80l, Ano81l]. Editors [Eas84b, Ano80e, Ano80f, Ano88h]. Eds [And86, Bra86, Gri87]. Edwin [BKvLSE80, BKvLSE84]. effect [CKV85, DG84, HK81, HK84a, Kac84, Nai84, Vio84, WS84b]. Effective [CO84, HIM84, LR84b, MN80, MGTD84, SMT88, Tal89, Nad86a]. effective-range [LR84b]. effectively [BM82a]. Effects [EPS81, TTAT81, BS83, BS84a, Cop81, Cop84a, Cop84d, Cop84c, Cop84b, CVvWF86, Hoe86b, JW85, Odo82, Odo84b, San82, San84, Sch81, Sch83, Sei84]. Efficiency [MRT89, AGF87, GTG81, GTG84, GTM85, GTM87, GTMA88, Lux80, MTN^{+}82, Sch82a]. Efficient [KM85, KB86b, MB83, MHP84, MB84, New86, NR86a, And86, BSK88, BVL84, CHMM84, KMS88, NP84, SH82, SH84, SM86, Ver82]. efforts [Sch88b]. Effy [GTG84, GTM85, GTG81]. EFYGA [GTM87]. EGS [CN87]. EGS4 [Miu87]. EGS4V [Miu87]. eigenfrequency [Cap86a]. Eigenfunction [PZCC89]. eigenfunctions [DK87a, JT83, JT84b, JA80, JA84, ST84b]. eigenproblem [Nas84]. eigenproblems [DW88]. Eigenstates [MPS84, FV86, FH88b]. Eigenvalue [Gru80, Gru84b, BN81, BH82, BMK89, Cli82, FJ89, Gru84a, Jam87, Rap83, Saa89, YP86]. eigenvalues [Bos83, Bos84, CH87, CLS88, DK87a, FV86, Joh88, Kar80, Kar84a, KLS85, MPS84, MP82, NKS88, Sch88a]. eigenvectors [KS86a]. eikonal [MPR80]. EIV [Shu87]. elastic [All84b, BB84b, Cop81, Cop84b, DS83, DS84a, EM84b, Fun84b, Gje81, JT84a, KF84, Keg81, Keg84, Lan82, Lan84a, LK88, LB84, MS84, OM83, OM84, PSL88, Pet84, RM89, SP86, Smi84d, TTW84a, TW89a, Yat84]. elastic-plastic-hydro [LK88]. elastic-scattering [PSL88]. electric [Arn84b, God84, Hib84c, LJLC84, Nai86, OSF86, ÖW84, SS84a, Tol86]. electric-dipole [Hib84c]. electrical [EK86, Isl84b]. electrodes [DEW84]. electrodynamics [SMM89]. Electromagnetic [BF84, Cap86a, CCR88, OM88, DM89a, VM89, YMWW89]. Electron [MMM84, PSBG84, Ras80a, Ras84a, Sin84, AT84, AP82, AP84, ANP^{+}85, AGF87, Bar83, Bar84e, BGS81, BGS83, BGS84c, BGS84a, BGS84b, Ben82, Ben84a, BDK86b, BDD84, BW87, Chi84a, CLS^{+}84, Cre80a, Cre84a, CSW84, EBS88, FT84a, FV82, GTL^{+}86, Gia84, HC83, HC84, Hib84b, JT84a, JA80, JA84, Lan84b, LW85, Len80, Lew84d, MLL86, Män89, MKH86, MS84, MGK86, NM84, NN84, Nor84a, Odo81, Odo84c, OM88, PP85, Ras80b, Ras84b, Rit86, RP89b, RP89a, SP87, Sal84c, ST81, Sal82, Sal84e, ST84a, SNC80, SNC84, SSF^{+}89, ST84e, SV84, ST82b, ST84f, TW89a, TSD84, TW89b, WL84, WW80, WB84, Yat84]. electron-Atom [NM84, AT84, Bar83, Bar84e, CSW84, Sal82, Sal84e, ST84e, Yat84]. electron-capture [AGF87]. electron-diatomic-molecule [MLL86]. electron-hydrogen [BW87]. electron-molecule [SNC80]. electron-photon [SV84]. electron-positron [ANP^{+}85, BDK86b, Odo81, Odo84c, Rit86]. Electronic [BAG^{+}87, BB84a, CC89a, DK87b, DWVH89, HRB84, SK82, SK84, WC84]. Electronic-structure [BAG^{+}87]. electrons [All84c, AM84a, BS83, BS84a, BGMP84, Chi84b, Gra84a, LW84a, MR80, MR84a, Moo84h, SMMP86, SS84c, SS84d]. electrostatic [HM81, HM83, HM84a, HM84b, OM88]. electrotransport [OBLR84]. element [AHVV86, Bur85b, DM89a, DD85b, EP86, FM85, KS84c, KS84g, KS84h, Kon84, Len80, Man86, Man83a, Par89b, PB84b, Pis84a, Pis84b, SFB87, SMC81, SMC84, Sun88, TTAT81, TL89, VM89, Zlá80]. element-by-element [TL89]. elementary [CC89a, MV83, MV84b]. Elements [BH88, FBS88, FBS89, ABG^{+}84, Arn84b, Ben82, Ben84a, BOT86, Bou87, Bra84d, Bru85, Bru86, Cha84e, Cla84, Das89, EEV89, FS84c, Hod84, Jam84, Klo84c, Kol81a, Kol84a, KM82, KM84b, Lab84a, Lew84d, MRC85, Par84, PGB84, Rob84a, Sal84c, SS84a, Sax84b, Sax84a, Sax84c, ST84b, Ten85, Tol86, YP89]. elgenvalue [BH84b]. elimination [Vaj82]. Ellips [Ber84b]. ellipsometer [Ber84b]. elliptic [CKT89, Jes84, KS89, Vaj82]. elliptical [Moo83]. ellipticity [Moo80, Moo84c]. Ellis [Jes87]. elongated [Hof88]. Elsasser [Moo84e]. EMAP [Wie86b]. Embedded [Nie88]. emerging [Met87]. emission [Duf84, Le 89]. emissions [OM88]. emitter [GTMA88]. emitters [Pie84b]. emphasis [And86]. Empirical [Krá86b]. emulator [VGD^{+}89]. emulsion [GDH^{+}89, Per89]. energies [BS80, BS84b, CDN^{+}87, DS83, DS84a, FE88, HR86a, HR86b, IR80, KSE86, Lew84c, Mag84b, Odo81, Odo84c, PI84, Ric82, Ric84, Sil84b, TR84, TTL88, Wil84b, Wil84c, WB84]. Energy [AFH84, AF84a, AF84b, CJ84, Kar80, Kar84a, Wic87, AP82, AP84, BMV83, BMV84, BBE^{+}81, BMR85c, Bar81b, Bar84c, BGMP84, BGS81, BGS83, BGS84c, BGS84a, BGS84b, Bla84, Boc89, BE89, Bur87, CC89a, CH87, CH88, Cha84b, Cug84, DM86a, Den88, Dun87, FV86, Gal84, Ger80, Gro89a, HIM84, Hin81, HRB84, Jac81, JT83, JT84b, JMS^{+}89, JP80, JP84b, KAH^{+}89, KP81a, Kaw86, Keg85, KKLP84, KVW84, KK84b, KKT^{+}88, Lan87, LW85, LSW87, LB84, LLC84, MRW84, MM84a, MGN80, MGN84, MBW89, Mon87, Nas87a, Nas89, New89, NT84, Ogi83, Ogi84, OLR^{+}87, PSBG84, PFGP82, Poh87a, Poh87b, Put89, Qua87, RFSG84a, RR84, Rim81, Rit86, Rob87a, RRG88, RP89b, RP89a, SMMP86, SN84, Sea82, Sea84c, Sko84, Van81b, Ves89, VČ88, WC84, Wat89, Wha89]. energy [Whi89a, Wil81, Wis87, vD87b]. Energy-level [AF84b]. Energy-loss [CJ84]. engineering [BCC^{+}85, Cam86, CMH87, Eck84, Jam86a, Nad86b, Sch86b, Sto89, Stu85, Wes87]. Enhanced [PCL89]. enhancement [GSS81]. enhancing [KWPP89]. ENSDF [Mir84, Tep84]. ensemble [HD88, Kra86a]. entire [BDG^{+}84]. entropy [DL86]. enveloping [DVD87]. environment [ACSB85, BGM87, Bur87, Dit89, May89, MGN80, MGN84, Nag85, SN84, Sta87a, WHD86, Wic87, Wis87]. environments [KSS86, Stu85, Was85]. ephemeris [SGS^{+}89]. epic [HD89, Eas87b]. EPR [DSM^{+}81, DSM^{+}84, Ste84b, dBM82, dBM84]. EPR-spectra [DSM^{+}81, DSM^{+}84]. equal [Jab84]. equation [AS83, And83b, And83a, AS84, And84b, And84a, AZK83, BH85a, Bel84a, BVL84, BD82, BD84b, BSA86, BL88, CR84c, CKA89, Cla82b, DGL89, DRRvB82, ERB88, FV86, FGSH85, FS84a, FBW82, FC84a, Fog84, GS83, Hay80, Hug84, IR80, Ixa80, IR85, IB87, IGB^{+}89, Jam87, KK83b, KP85, LWHH82, LWHH84, LR84a, Mac89a, Mar84a, MP81a, MP84a, Moo83, Mor84b, MZZ89, NBI^{+}85, PR87, PM80, PM83, PM84a, PM84b, Rap81b, Rap83, RC85, RC87, RG80, RG84, SA83, SA84, SFB87, SBFI87, Ste84c, TSD84, TV86, TT89, VPB82, VPB84, Wie85, WW80, vTBF84]. Equations [Art85a, EFK85, IK85, RW86, Sch82b, Sch86d, AKS88a, AKS88b, BBM82, BBM84, BP82a, Bur85b, BCCM87, CSC89, Car84a, Cha84d, Chi84a, Cla82a, Cre80a, Cre84a, CSW84, DWC^{+}89, Ede84, FK86a, IK81, FOW87, Gaj85, GPS84, GS88, GSZ85, HD86, Has84, HLS84, Hou87, Ito86, Jes84, KP89, KA87, KRSV82, Kos80, LM85, Mar89a, MP81b, MP84b, Mik80, Mor84c, NN84, Nor84a, Par84, Pia84, Pro82b, Pro82a, Pro84, RWK^{+}85, RW84, Rud84, SC89, Sch85c, Sch84b, Sch84c, SK82, SK84, Sun88, TF89, Tol85, WMH82, WBG84, tR85]. equilibration [RBMC86]. equilibria [BWD84, CR84a, Goe84, GKM^{+}81, HL86, HC86, KJ84, Lao84, RG86, Sch84d, She84, TS84b, Var89, ZC84]. Equilibrium [SS81, Bet84, BLT81, BM84b, CRL84, Hof88, MN84b, Mik84, Shu87, WH81]. equilibrium-diffusion [CRL84]. equilibrium/compound [HMS84]. equivalence [PT82, You87]. equivalent [All84c, AM84a, Ata85, Chi84b, Gra84a]. era [WBV^{+}89]. Erato [GST^{+}81, GTR^{+}81, TTM^{+}80, GTB^{+}81, GTB^{+}84, SG81]. Erika [RF83, RF84]. EROTVIB [JT84a]. Erratum [Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano81o, Ano81p, Ano81q, Ano81r, Ano81s, Ano82t, Ano82u, Ano82v, Ano83k, Ano83l, Ano83m, Ano83n, Ano83o, Ano84x, Ano84-30, Ano84-72, Ano84-31, Ano84-32, Ano84-73, Ano84-33, Ano84-34, Ano84-35, Ano84-74, Ano84-36, Ano84-37, Ano84-38, Ano84-39, Ano84-40, Ano84-41, Ano84-42, Ano84-43, Ano84-44, Ano84-45, Ano84-46, Ano84-47, Ano84-48, Ano84-49, Ano84-50, Ano84-51, Ano84-52, Ano84y, Ano84z, Ano84-27, Ano84-28, Ano84-29, Ano84-53, Ano84-54, Ano84-55, Ano84-56, Ano84-57, Ano84-58, Ano84-59, Ano84-60, Ano84-61, Ano84-62, Ano84-63, Ano84-64, Ano84-65, Ano84-66, Ano84-67, Ano84-68, Ano84-69, Ano84-70, Ano84-71, Ano84-75, Ano84q, Ano84r, Ano84s, Ano84t, Ano84u, Ano84v, Ano84w, Ano85o, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano86u, Ano86v, Ano86w]. Erratum [Ano86x, Ano86y, Ano86z, Ano86-27, Ano87u, Ano87v, Ano87w, Ano88u, Ano89t, Ano89u, Ano89v, Ano89w, KS08, Ran87a, Tho04a, Tho04b]. error [CM84d, Moh80, Moh84, Rap83, Ver82]. Errors [MLL86, EFES83, Hay83, JR84]. ES-computer [Tar86]. ESA [Wen88]. ESECT [Wie86b]. ESECT/EMAP [Wie86b]. ESOP [DGK^{+}81, Jac81, Jac82]. especially [Aue84]. essential [Col88, Mou89]. establishing [FC84b]. estimate [GH83, GH84a]. Estimation [DP84, DL86, HB85]. ETA [BCMR87, MMR89, MKRJ89, VBD88]. ETA-10 [BCMR87, MMR89]. Ethernet [Gam89]. Euclidean [GV89]. Euler [Di 87]. Eulerian [IGB^{+}89, WHW87]. Europe [Duf85]. European [Ano88-27, CFG^{+}87]. eV [vdM84a]. eV-region [vdM84a]. eva2 [BPW84a]. evaluate [AGF87, BW87, CW80, CW84a, Cha84e, GH84b, Nov87, Rob84a, TT82, TT84]. Evaluated [LCS84, Mir84, Tep84, Ver84b, Ver89]. evaluates [HM81, HM83, HM84a, HM84b]. Evaluating [Ito85, NT84, Tal82, Tal84a]. Evaluation [Com82, EEV89, Rob87a, SG83, SG84b, BDJ^{+}89, DFD81, DFD84, Di 86, GDH^{+}89, HKK84, KK84a, KP84, Moo84e, Mue84, Nes88, Nob80, Nob84, ONP89, Sax84b, SG84a, SC84, ST84e, Tau80, Tau84c]. evaporation [DKA85]. evar [BPW84c]. eve [Wil89]. Event [Nas89, CDF^{+}89, CHS^{+}89, DFOP89, FR86a, Kaw86, O’N89, Odo81, Odo84c]. events [AI86, BD84d, Gal84, MPS81, Rit84b]. Everett [Her82a, Her84b]. everyday [Ano88-27]. Evolution [Van81a, Cla88, EPS81, GSZ85, Ito86, Mik80, Odo82, Odo84b, San82, San84, SVP86]. Ewald [AF82, AF84c, Fin82, Fin84a]. ex [LS84]. ex-core [LS84]. Exact [BH88, DG84, FKU^{+}84, Gol84b, Jam86b, LW84a, NKS88, TL84, TUWA84, TW83, TW84, HDK84, Mar84a, NFH^{+}88, SW80b, SW84, Ten83, Ten84a, WL84]. EXAFS [AI88, IA80, IA84]. examine [TCS85]. example [AB88, CNRS82]. examples [RW85]. exchange [AT84, Nob80, Nob84, Par84, ST84e, WGM83, WGM84, vdB84]. excitation [DM89a, Kas84, KK84b, LJLC84, MMM84, Nog83, RSA84, ST81, ST84a, SS84a, TSD84, Tol86, Vaz81, Vaz84]. excitations [Szy81]. excited [KM86, Lab84b, LHP87, Mal84, RA82, RA84, Tal84b, WS80]. exec [Rea84]. Exlam [WGM84, WGM83]. expanding [Tal82, Tal84a]. expansion [BCCM87, But89, CWS83, CWS84, Del84a, Lew84c, Lew84d, MLL86, Mar89c, MBW89, Mor84b, Mor80, Mor84d, SNC80, SNC84, She84, Sil84a, Sil84b, VČ88, WGM83, WGM84, WS80, Wil84c]. expansions [Gro89b, KS84d]. expectation [EHH88]. Experience [Del87, DR82, Jac82, JMS^{+}89, RS85, Sli89, Sph87]. experiment [ABLB^{+}89, AP89, BLMU81, Beh81, BE89, DGK^{+}81, Gje81, Kos80, PFGP82, PS88, SN84, Sko89]. Experimental [LGG^{+}87, ABMT84, BCM84a, BD80b, BD81, BD84c, CH84c, Den88, Eld84, FV81, IL84, KRS83, LCS84, Poh87a, Reg81, Rei89]. experimentation [CF85, Col88]. Experiments [MBW89, BBE^{+}81, Bun86, Cop81, Cop84a, Cop84d, Cop84c, Cop84b, CVvWF86, GH83, GH84a, Jac81, Mae88, Mak80, Man83b, Mou87, Nas89, New87, Nur85, OLR^{+}87, Rim81, Ryn82, Ryn84, RD80, RD84, Whi89b, Wil81, ZS87]. expert [BC87]. Explicit [RB84, CSC89]. explosion [MPM83, MPM84]. exponent [Com82]. Exponential [RC87, BKvLSE80, BKvLSE84, Kal83, PIM80, PIM84b, Tor82, Tor84]. exponentiality [Kal82]. Exponentially [Rap83]. Exponentially-fitted [Rap83]. exponentials [Bro83]. Expression [ACSB85]. expressions [YB86]. extended [AF84b, Cre81, Cre84b, DBB84, DW88, KE84b]. Extension [BRTB84, Var89]. extensions [GST^{+}81]. extensive [GHO^{+}89, Goo84]. External [DK82, Wor84a, VČ88]. extraction [IL84, Mor84a]. extrapolation [CC84]. extrapolations [CC84].

F [Eas89, Sta87b]. F.S.U. [LGG^{+}87]. Fabry [MMF81]. facilities [HR84e, Zac81]. facility [BGMM81, Day81, Day84a, DFOP89, MP89, Nic89]. FACOM [MTN^{+}82]. factor [Han84b, Isl84b, Poe85, SS84c, SS84d, TUWA84, WDM84]. factorization [BM85b]. factors [All84a, BCFK85, BS84d, CDN^{+}87, DK87b, DW84, EE89, Fun84a, GT83, GT84, Han84d, Isl81, Isl84a, Rei87, SSS84]. FADO [WPVS89]. failure [KK85, KB86a]. FAMP [Her82b, HGK^{+}81]. Faradaic [BCM84a]. farming [BDJ^{+}89]. farms [GH87]. Fast [AZK83, BKHJ83, Che89, FLM84, Her82b, MMS87, NFH^{+}88, SCC^{+}88, Tem82, AGR88, Beh81, BN81, BLR82, BLR84, BSDM88, CWS83, CWS84, Car89, CMM84, DGK^{+}81, FSR84, Gje81, IGB^{+}89, Jac81, Jac82, KS89, Keg85, NRS88, PJ89, PPHS86, RS83, RS84d, SM89b, Tel84, Uga88, Vaj82, VBD88, Voh88, CH84b, GT83, GT84, Mar84b, NR86a]. FASTBUS [ABG81b, BC87]. FASTF [GT83, GT84]. fault [HAG^{+}89, MD81]. fault-tolerant [MD81]. FBT [Hof88]. fcc [DM89b, NHK82, Sal83a, Sal84d]. FCFRKR [TT82, TT85, TT84]. FD [FR86b]. FDG [FR86b]. FDH [FR86b]. features [RW85, WHD86]. FEDGROUP [Ver84b]. FEDGROUP-3 [Ver84b]. FEDMIX [Ver89]. FEL [TW89b]. Fermi [BDM81, BDM84, FR86b, PDF82, PDF84, Pic89]. Fermi-surface [PDF82, PDF84]. Fermilab [GAA^{+}87, Joh89, Nic89, Whi89b]. fermion [DV85, Duc80]. fermions [DeG88, Oya86]. ferromagnet [NHK82]. Feshbach [SS84b, AXG87, Smi84b]. few [Gir81, HR87]. Feynman [KKS87, KR88, Stu88, WP84]. FFC [DK87b]. FGM [Szé85]. fiber [BCC^{+}89]. Field [Ano84-85, Mat84, SBM^{+}86, ST89b, BB84b, BM84b, CCR88, CCR84b, CGG87, CT82, Des84a, EPS81, Frü81, GLST89, Hoe86a, HWL89, Kar80, Kar84a, Klu84b, KS84g, KS84h, Lee82, Lee84, LCW84, May81, May84, McN81, McN84, MMO89, ML81, MR86, MR87, O’C84a, OSF86, ÖW84, Ste84a, VČ88, WB88, vMF81, vMF84, vM84g]. field-gradient [vMF81, vMF84, vM84g]. field-reversed [HWL89]. fields [CD86, DE84, DS83, DS84a, JBWW86, Mar89c, Mik84, Pet84, PT86, TSD84]. FIFPC [FSR84]. Fifth [Ano82w, Ano86f, TL82, LT85]. figs [Eas84a]. figures [AO88]. file [Ben82, Ben84a, CR84b, Mir84, Tep84]. files [BJ81a, BJ81b, BJ84a, BJ84b, Cre80b, Cre84c, LCS84]. filing [KS86a]. films [Hub84e, Hub84f]. filter [Gje81]. filtering [CHS^{+}89, WMN85]. filtration [Zlo81, Zlo84c]. final [ANP^{+}85, Bra84c, Bra84b, TS89]. Finding [IK85, Hau87, Ito86, LWHH82, LWHH84]. Fine [Sar84b, Sar84c, Dec89]. Finite [DD85b, FBS88, FBS89, Man87a, Zlá80, ABG^{+}84, AHVV86, BLR82, BLR84, Bur85b, CSC89, EP86, FKU^{+}84, FM85, GS83, Hod84, KS84c, KS84g, KS84h, Kon84, Lan80, Len80, Lis84, MVV83, MVV84, MV87, Man83a, MSFR89, NK84, PV84, Pis84a, Pis84b, SFB87, SMC81, SMC84, Sun88, TTAT81, TL84, TUWA84, ZNO88, Man86]. finite-and [Lan80]. finite-element [KS84c]. finite-range [FKU^{+}84, PV84]. FIRE [MPM85, MPM83, MPM84]. First [Hub84c, Sch86c, Sch86d, CRH^{+}89, DM88b, IH86, JP84a, Köl81b, Köl84b, Kos80, Moo80, Moo84c, tR85]. fission [Jam86b, TCS85]. fit [Han84f, Les89, MC85, RVD^{+}81, vMF81, vMF84]. fits [KL84a]. fitted [Rap83]. fitting [AL81, AMK81, AMK84, BB86, Bil89, Bro83, CO84, CAP86b, EHH88, GJ84, Hau87, MGMZ84, O’C84b, RC87, Sci89, Sha87, Ste84a, TN84, Wil84b, WS84b, Woo86b, vM84a, vM84c]. five [FGK89]. five-parton [FGK89]. fixed [Bat84]. FIXSRC [BH85b]. flavours [Odo84a]. Fletcher [Ano89h]. flexible [Eck84, MD81, MP81b, MP84b, Szé85, WS84b, vM84a]. flight [HAG^{+}89, Keg81, Keg84, Sne85]. Flip [BKR88]. floating [MB84, Bur85b, MB83]. floppy [Ten83, Ten84b, Ten84a]. flow [Bra87a, BKR88, Del87, FCW84, For89, KJ84, KW84b, SWHB88, SSM85, TL89, TF89, Eas84a]. Flows [Art85b, Cro88, Fri88, HH82, Key89, KL84b, Rob87b]. Fluences [Keg85]. Fluid [Ano89h, MT81, And86, Ano82w, Bra87a, BKR88, DGL89, Eas80, Eas84a, Eas88, FBW82, Gal85, Lad86, Ods84, RBMC86, Ros85]. fluids [Gup88a, Gup88b]. flux [EFES83, MMR89, Rei89, SF84a, Var89]. fluxes [Wie85]. Fluxoids [Kam84]. FMEHAM [Fle81]. Fock [Des84a, AAH84, ABH^{+}84, Bat84, CCR84a, CCR84b, Des84b, Fle81, Fro84a, Fro84b, Fro84c, Fro87, GMN^{+}80, GMN^{+}84, HKS83, HKS84, Irv82, KKLP84, Lab84c, MGN80, MGN84]. focusing [Nor82, Nor84b]. foil [IL84]. foils [RS83, RS84d]. Fokker [ERB88, FSR84, MMK81, MMK84, MGK86, MMTK88, OCS86, SFB87, SSF^{+}89]. Folded [PIM80, PIM84b, Coo82, Coo84a]. force [Bin87, CHL^{+}81, Gel85, Han84a, Han84f, May81, May84]. force-free [CHL^{+}81]. forces [LR84a, LR84b]. Forecasting [Dic82]. Form [Ito85, SS84c, BCFK85, BP82a, BM84b, EE89, Fun84a, Isl84b, Sal84a, SSS84, SS84d, TUWA84]. form-factor [Isl84b]. formac [FK86a]. Formal [Bri86, Gór88, Hoa89, Suf89, HKK84, Dob85]. formalism [Hof84b, MV87, NBI^{+}85, Sal82, Sal84e, SS84b]. format [Jac85b, Jac85a, Sal84a]. formation [KL84b, Ric82, Ric84]. formatting [Tau81, Tau84b]. formed [DKA85]. Formint [GSZ85]. forms [Sko89, TT89]. formula [BK84b, Jab84, KB84, Tak88a, Tak88b]. formulae [LU83, LU84, Sei84]. formulas [RB84]. formulation [CN84, MT81, Nob80, Nob84]. FORSIM [Car84a]. Fortran [Nad86a, AI88, Bai86, BSdlT87, BS86a, Bru86, CSC^{+}86, CM86, DFD81, DFD84, DM89b, GH87, Gro89b, MMM85, MD88, Pag87, RW85, Rom81, Sch89c, VH87, dB82, AD84, Bar84a, Ber84b, BDJ^{+}89, BH85b, BJ81a, BJ81b, BJ84a, BJ84b, CRV^{+}89, CDW82, CDW84, CR84d, Col80a, Col84a, Com80b, Dob85, GDK89, HM81, HM84b, HR83b, HR83c, HR84c, HR84d, IA80, IA84, Jac85b, Jac85a, Jes82, Joh87a, KK89, MW84a, MW84b, May89, Met85, Met87, Met89, Nai84, Nai86, PV84, PS84c, RV84, RRC89, Rob83, RS81, RS84b, Sal84a, SMD84, SR84a, SDH84, VSH83, Zoh84, dB84, vMF81, vM84d, vM84f, vMF84, vM84g, vMT84, vM84e, vM84h, vM84i]. Fortran-IV [SDH84]. Forward [Ano88v, Dob89]. Four [MC84a, NRS88, BCM83a, BDK86b, IR85, Mue84, WwYpCg88]. four-body [Mue84, WwYpCg88]. four-dimensional [BCM83a]. four-lepton [BDK86b]. Four-stage [MC84a]. four-step [IR85]. Four67 [CH84b]. Fourgen [Mar84b]. Fourier [CH84b, GT83, GT84, Mar84b, NR86a, She84, AI88, CRS87, Can81, CKT89, CT84b, Col84b, DM88a, IA80, IA84, IGB^{+}89, Jun81, Jun84, KK83b, NR86b, NRS88, PS84c, Sch84d, SM89a, SM89b, WSV88]. Fourier-expansion [She84]. fourth [Ano82i, Col80b, Nog83, Vaj82, WS80]. fourth-order [Col80b, Nog83, Vaj82, WS80]. FPPAC [MMK81, MMK84]. FPPAC88 [MMTK88]. FPS [Bur85b, Gro87, Gro89a, MB83, MB84, Rob87a]. FPS-164 [Bur85b, MB83, MB84, Rob87a]. fractal [GV89]. fraction [Nex84, SW80b, SW84, TB85, Tho04b]. Fractional [All84c, AM84a, Gra84a, Chi84b, Hub84b, MNJL84, SK86]. fractions [Lee89]. fragment [TCS85]. Fragmentation [GV89, BTWN87, Rit84a, Sjö82, Sjö84b, Sjö86, SB87]. fragments [DKA85]. frames [SM84b]. framework [Car80, GSS81]. Franck [All84a, DK87b, TT82, TT84]. FRANPIE [Vaz84, Vaz81]. Fredholm [tR85]. Fredrikze [BD81]. free [CR84a, CHL^{+}81, HvRM86, Hof88, Jac85b, Jac85a, Ran86, Ran87a, RWZ87, TW89b, Vaz81, Vaz84, Cro88, Fri88, Tre88]. free-boundary [Hof88]. free-electron-laser [TW89b]. Free-Lagrange [Cro88, Tre88]. Free-Lagrangian [Fri88]. freedom [SK82, SK84]. FREESCO [FR86a]. Freint [Gru84d]. frequencies [BRTB84, CCR88, Han84f]. Frequency [Lee89]. Fresnel [Gru84d]. Freudenthal [KB84]. friction [Vaz81, Vaz84]. frictional [Has84]. friendly [Vis84b]. FRITIOF [NAS87b]. Frozen [CCR84a]. Fscript [Tau84b, Tau81]. FT [AP89]. FT-NMR [AP89]. FTIDY [BS86a]. Fujitsu [MHP84]. full [JBWW86, KH89, Tau81, Tau84b]. full-wave [JBWW86]. function [BD80a, BD84a, Bar84a, Bru85, Bru86, Car89, Col80a, Col84a, DG89, De 84c, DL86, Del84b, Deu84b, Deu84c, Deu84d, Fie80, Fie81, Fie84, HvRM86, Jam80, JR84, Köl84c, LM85, LB84, MS85, MS87a, MS88, Moo81b, Moo84b, New86, Pie82, Pie84a, Tau80, Tau84c, Ten84b, Woo86a]. functional [Gir81, GSS81, Ric86]. Functions [Cam84c, SM84b, AT84, And84c, AM84b, AM84c, AM84e, BDM81, BDM84, BDG^{+}84, Bar81a, Bar81b, Bar82, Bar84b, BFSG84, Bar84c, BS80, BS84b, BD87, BC83a, BC83b, BC84a, BC84b, BCM83b, BCM84b, But89, BC88b, CWS83, CWS84, Cam81, Cam84a, Cam84b, Cam84d, Cat89, CgYpXh85, CM81a, CM84c, Cug84, CDN^{+}87, De 88, DGPM85, DFD81, DFD84, DS83, DS84a, Di 86, DF84, Eğe82, Ege84, EMR84, FS84d, Gas81, Gas84, Gro89b, Han84c, Han84e, Hel80, Her82a, Her84b, Hib84c, HFG88, HSV84, Jak84, Jam84, Kar80, Kar84a, KM84a, Köl81b, KS84d, Köl84b, KB86b, Lab84d, Les89, Mas83, MM88b, MKS83, MKS84, NP84, Nes84, Nes88, NT84, Odo82, Odo84b, PB84a, Rao81, Rob84b, RBHW84, Sal84c, San82, San84, Sar87, Sea82, Sea84b, Sea84c, Ská81, Ská84, Smi84c, ST84e]. functions [SG84c, Ste84c, TTW84b, TR84, Tau84a, Ten85, TB85, TB87, Tho04a, Tho04b, Vaz81, Vaz84, Ver89, Yur84, Zlo89]. fundamental [AMC83, AMC84]. FURI [Bar84a]. Fusion [Kil80, PDNE88, Atz86, CRPL81, CRPL84, CAR84b, DL87, FNDL89, MPM83, MPM84, Vaz81, Vaz84]. Future [Gam89, Hut89, Wen88, Bra87b, Gai89, Gri89, JMS^{+}89, Kas82, Lev89, Met85, Nis88, Pat82, Str87]. Fuzzy [KK85, KB86a]. FX [Par89b]. FX/8 [Par89b].

G [Eas89, Jes87, Sta87b, Rea84]. G-exec [Rea84]. GAEA [TTT87]. galactic [JN85]. game [Tat80]. GAMM [Ano82w]. Gamma [Hub84d, AWG84, Eks84, FC84b, GTM87, GTMA88, Hub84c, Köl84c, KAR81, KAR84b, Lea87, MRCL84, RSL^{+}84, SS84c, SS84d, Szé85, Zlo82, Zlo84a]. gamma-gamma [MRCL84, RSL^{+}84]. Gamma-radiation [Hub84d]. gamma-ray [AWG84, FC84b, Hub84c, Lea87, Zlo82, Zlo84a]. gamma-spectrum [Szé85]. Gamow [VPB82, VPB84]. gas [HP83, HP84, MPM83, MPM84, OS84, Rom81, SWHB88, Sea84a, Tho84]. gaseous [DE84, PSBG84]. gases [Câm82, Câm84g, EK86, PM80, PM84a, TSD84]. gateway [Ghi87]. gather [Voh88]. GATO [BHM81]. gauge [Ano87b, Ano89x, AM81, AMC83, AM84d, AMC84, BM82a, BM82b, BM82c, BCM83a, BMR84, BM84a, BMR85a, BLR82, BLR84, CM85, CGGK85, CCS89, CMS83, CMS84, CM83, CM84e, DMM84a, DM87, DMM83, DMM84b, EMM81, Hoe86a, Ken89a, KLS^{+}89, KR88, KM85, Mac89b, MMS86, MB83, MHP84, MB84, MT86, VBD88]. gauges [Chr84a, Chr84b]. Gaunt [Moo84e]. Gauss [Rys84, Tak88a, Tak88b]. Gauss-integration [Rys84]. Gaussian [AAH84, ACR88, HKS83, HKS84, Kol81a, Kol84a, Mat84, Puf83, Puf84, Tau84a, War84]. Gaussian-type [HKS83, HKS84]. Gaussians [vM84c]. GEANT3 [Ano87c]. General [Fro87, RS80, RS84c, Sil84a, ACD88, And84c, Aue84, Bai86, BSK88, BBC^{+}84, BBL^{+}84, BCVZ89, Bur84, Cha84d, DS84b, DWVH89, DGJ^{+}89, Fro84a, Gra84b, Gup88a, HC83, HC84, HRS81, HRS84, Hib84b, Hib84a, Hib84c, Hib84d, Hib84e, ITR^{+}87, JT84a, Jun81, Jun84, Klo84b, Pro82b, Pro84, ST84b, ST82a, ST84c, Ten84b, WB88, WS84b, vM84b]. general-purpose [BCVZ89, DGJ^{+}89, vM84b]. generalised [GV87, Rao81]. Generalization [WW84b]. Generalized [CKW89, FK86b, FR86b, SK86, SG84c, BHSS84, CL84, Das83, DW89, Dob84, FE88, Gel85, KP89, Mor84c, Pic89, SS84b, YpMzCeCg85, Zlo89, Zoh84]. generate [CDW82, CW84b, CDW84, JT83, JT84b, Rob84b, Rob84d]. generated [BK84a, CCR88, HSV84, HB88, PZCC89]. Generating [Gas81, Gas84, ANP^{+}85, BB84c, Ram85, Tau84d]. Generation [KS84a, SM84c, Ská81, Ská84, Woo86a, AP80, BF89, BD84d, Bro89, De 88, Di 87, GJ84, Gro89b, Han84a, KKS87, KVW84, LT85, Pis84a, Rap84, SRS85, TBB87, TL82, WP84]. Generator [Lab84b, SUZ89, ABM89a, Bro84, CG87, FR86a, FS82, HR83c, HR84d, Jad84, KW84a, Kaw86, KPP81, KPP84a, KPP84b, KCA84, Mar84b, Pis84b, Lab84a]. genpot [Ten84b]. Gentzsch [And86]. geodynamic [Moo83]. Geomagnetic [Klu84b, KL84a, Moo84e]. geometric [AO88, KK83a]. Geometrical [DR85]. geometries [BH85a, HKK84]. geometry [Ano87c, Cop84d, Cop84c, Ede84, ELS88, Hub84d, KBMV86, MGD81, MBOK86, Wie86b]. geophysics [Har84b]. GEP [Bas87d]. Germany [Run89]. GeV [AGH^{+}88, HR86a, HR86b, RR84]. GFIT [Zlo89]. Gflops [Ano89x]. GFSR [ABM89a]. giant [CGB^{+}89]. Gilat [Sim80, Sim84]. Ginzburg [Kam84]. GISP [Rys84]. given [Rob84d]. GKS [Osl88]. glass [BSDM88, FAK82, HDK84, NHK82]. glasses [BC88a]. GLE200 [SGS^{+}89]. global [KBMV86, WSV88]. glow [Abr88]. Glowcode [LNS84]. glueball [KMS88]. GMRES [Nav89, Wal89]. Godunov [ACD88]. golden [Le 89]. GORDAN [RB84, SM84c]. Goudsmit [Vio84]. Grad [LWHH82, LWHH84]. Gradient [Ano84-85, GLC89, KPP81, OJK89, AKS88a, AKS88b, BMR85b, CKA89, CLL89, EA89, KJ85, Klu84b, KA87, KPP84a, KPP84b, Mat84, Meu89, Sun88, vMF81, vMF84, vM84g]. gradients [BLB88]. grain [BKD83, BKD84]. Gram [O’C84b]. grand [BMG^{+}85]. Graphic [Art85b, MPS81, TTT84]. Graphical [Bas87d, BSK88, Bas88, Car88, FMS87, KS87a]. Graphics [Hop85, Rue89, Sch89b, Wil88, CGMS89, CN87, Die88, Gou88, HB88, Joh89, KR81, Mye88, Mye89, Osl88, Sch89a, Sta87a, VM89, Via87]. graphs [KKS87, Rap84]. GRASP [DGJ^{+}89]. Grass [HD89]. gravitational [Cla82b, CH84d, Man87a, ML81]. gray [Dun87]. gray-scale [Dun87]. Green [DS83, DS84a, HvRM86, LM85, LB84, MS85, MS87a, MS88, MKS83, MKS84, RBHW84]. Grenade [BSA86]. grid [AP80, Bas87c, WMN85]. grid-search [Bas87c]. gridding [Fri88]. Gridless [LTK88, AGR88]. gridpoint [LT87]. grids [AZK83]. ground [And84c, BC88a, HDK84]. ground-state [And84c]. Group [WW84c, EC84, Gau84a, HM84d, KS84a, KS87a, KR88, MG84, NW84, SM84b, SM81a, SM81b, SM84d, SM84e, SM85b, SM85d, SM85e, WW84a, WW84b]. group-theoretic [KR88]. Group-theoretical [WW84c, WW84a]. Grouped [TL89]. groups [ADO84a, BK84a, BLR82, BLR84, GS88, MSFR89, Net84, PZCC89, SM81a, SM84d, SM85c, TS84a, Tho87, YP88, YP89]. Growing [RL80]. growth [AS89, DE84, DEW84, KM84c]. Gruber [Man86]. Grüneisen [BMS87]. guaranteed [Bri86]. guide [AD84, CC84]. guiding [GTL^{+}86]. GVSCF [WB88].

H [Art85b, Eas84b, BS84d, GS89, Joh88, LNB^{+}88, RS83, RS84d, WMCH84, vdM84a]. H. [BD81]. hadrin [HR86a]. hadron [BB84b, HR86a, HR86b, RR84]. hadron-nucleon [HR86a]. hadronic [ANP^{+}85, BS87, Odo84a, Odo84d]. hadronization [BT89]. hadrons [KVW84, NAS87b]. half [Di 86, PIM82, PIM84a]. half-integer [Di 86]. half-life [PIM82, PIM84a]. Halftime [Ott89]. HAMILTONIAN [Ste84b, CH88, DK87c, EEV89, FH88a, GWL89, Hib84b, HM84e, MGTD84, ST82a, ST84c]. Hamiltonians [Hay89b]. handle [Fun84b]. handling [SdAD^{+}88]. Hankel [Cam84d, Joh87b]. hard [Fin87, Fre87, Gri87, Lad86, Tru88]. hard-core [Fre87]. Hardcover [Ano89h, Lis88]. Hardware [ABM89a, BGMM81, Hoe89, LEF^{+}89, Poe85, WH89, Zac81]. Harmonic [Rei87, CgYpXh85, Dob84, FT84b, Mar89c, SG84c]. harmonics [Bra84a, CRL^{+}89, Gas81, Gas84, WwYpCg88, WS82, WS84a, dB82, dB84]. Hartree [Fle81, HKS83, HKS84, Irv82, AAH84, ABH^{+}84, Bat84, CCR84a, CCR84b, Des84a, Fro84a, Fro84b, Fro84c, Fro87, KKLP84, Lab84c]. HASE [MN84b]. Hauser [AXG87, SS84b, Smi84b]. having [BZ84, Wai87]. hazards [Bro83]. HBOOK [Les89]. HBr [vdM84a]. HCl [SAC87]. heat [GH83, GH84a]. Heater [MCJ81]. heating [AHS^{+}86, EP86, MKH86, OCS86, SSF^{+}89]. heavy [Aue84, But84, CJ84, DL87, DP88, EE89, FRS83, FRS84, FKU^{+}84, Irv82, Odo82, Odo84a, Odo84b, Par84, Sta89, TL84, TUWA84, Vaz81, Vaz84]. heavy-ion [Aue84, DL87, EE89, FKU^{+}84, TL84, TUWA84]. Heermann [Fin87]. Hehl [Eas89]. Heidelberg [Eas84a, Fin87, Man86, MD81]. Height [Deu84d, AK84, Deu84a, Deu84b, Deu84c]. HEIZ [GH83, GH84a]. helical [GKM^{+}81]. helically [HC86]. helium [Kno84, RMV89]. hemispherical [TCS85]. HEP [AW89, BDJ^{+}89, BBC^{+}89, CFG^{+}87, DP89, DL89, Flu89, Hau87, Her89, LGG^{+}87, McC87, Mou87, Mye89, New87, Run89, Sne85, Wie89]. HEPDATA [Wha89]. HEPVM [McC87]. Hera [GST^{+}81, Var89]. hereditary [FOW87]. Hermes [Coo84b]. Hermite [Tak88a]. Hermitian [KLS85, Nas84]. Hevol [Odo82, Odo84b]. Hevol2 [San82, San84]. Hewlett [KKLS82]. hexagonal [May81, May84]. Hfater [MCJ84]. Hierarchical [Her88a, Mur88, HBF^{+}85, OY80]. Higgs [Man83b]. High [BCCM87, IKRT89, LK89, Sch81, Sch83, SC84, UI85, Wat89, Wic87, BBE^{+}81, BRTB84, BMR85c, Bar81b, Bar84c, BGS81, BGS83, BGS84c, BGS84a, BGS84b, Ben84b, BI85, BD82, BD84b, BR85, BKD83, BKD84, Boc89, BPSB88, BE89, Bor82, Bor84, Bur87, CR84c, CKV85, Den88, Dob89, Dun87, Gal84, Ger80, Gro89a, HIM84, Hin81, IW87, Ing87, Jac81, JMS^{+}89, Jes82, KAH^{+}89, KP81a, Kaw86, KVW84, KSE86, KGAL88, KKT^{+}88, Lan87, LSW87, LCH^{+}81, MT81, Mon87, Nas87a, Nas89, New89, OLR^{+}87, PFGP82, Poh87a, Poh87b, PS88, Put89, Qua87, RR84, RSS85, Rim81, Rit86, Rob87a, RP89b, RP89a, Run89, SM84a, SN84, SDH84, SRS85, Sun88, Van81b, Ves89, WPVS89, Wha89, Whi89a, Wil81, Wis87, vD87b]. High- [Sch83, IW87, Ing87, Ben84b, BI85]. high-accuracy [PS88]. high-energy [BMR85c, Dun87, Gal84, Lan87, PFGP82, RR84, SN84]. high-intensity-laser [Bor82, Bor84]. high-latitude [KGAL88]. high-lying [SDH84]. high-order [Sun88]. high-speed [CKV85]. higher [BB84b, DHN84, FOW87, Ros85]. higher-order [DHN84, Ros85]. highly [BMR84, Hof88, Kri89, Ost85, RA82, RA84]. HIGZ [Joh89]. Hilbert [GWL89, Lan80, Tau80, Tau84c]. Hilbert-space [GWL89]. Hilger [Duc86, Eas82, Lis88]. Hill [Taj86]. Hirota [Ito88]. histogram [BEM84]. histograms [Cha80, vM84f]. Historical [You89]. HOBO [Cla88]. Hockney [Eas82]. hold [Her89]. Holland [Bra86, Tru88]. Holt [Eas84b]. homogeneous [KS84f, Ugn80]. homonuclear [Lad86, SLH84]. HONDO [DWVH89]. horizontal [Cop84c]. Horwood [Jes87]. hot [Ber80, FM84, KBMV86]. hot-plasma [KBMV86]. Householder [BM85b]. HP [Jud87, KKLS82]. HP-1000 [KKLS82]. HPLAS [WMCH84]. HTR [KK85]. huge [BLMU81]. Hull [Duc86, AN87]. human [BJM89, Nag89]. hump [Jam86b]. Hurwltz [GV87]. hybrid [BT88, CCR88, OSF86, SAVV87, SSF^{+}89, TTAT81, TS84b]. hydro [LK88]. hydrodynamic [KL84b]. hydrodynamical [FM84]. Hydrodynamics [Ben88, JN85, MPM85, RSMTV88, Tre88]. hydrogen [BBH^{+}84, BW87, CT82, Kar80, Kar84a, LHP87, MLS84, MLo84, Mal84, SG89, Tal84b, VČ88]. hydrogen-like [LHP87, MLS84, MLo84, Tal84b]. Hydrogenic [Lew84b, FS84b, Jam84, Moo84h]. hydrostatic [Moo80, Moo84c]. Hymnia [Gru84a]. Hymniablock [Gru84b]. HYMNISBLOCK [Gru80]. hypercube [Sco89]. hypercubic [CM85]. hyperfine [GH84b, Nai84, Nai86]. hypergeometric [Rao81]. hyperspherical [CgYpXh85, KPP^{+}89, WwYpCg88]. hypersurface [BP84b]. hypersurfaces [BMV83, BMV84]. hypervirial [VČ88].

IAEA [Kil80]. IBM [Bas88, BTS84a, CHS87, Com80b, Ghi87, Gri89, Gro89a, HRR84, Jud87, LBK88, NA84, Suc87]. IBM-FPS [Gro89a]. Ibmol [OCC84, OCC80]. Ibmol-7 [OCC84, OCC80]. ICCG [KPP81, KPP84b, vdV89]. ICCG2 [AS83, AS84]. ICCG3 [And83a, And84a]. ICL [AWB82, Del85, JN85, KSG82, Par82, WB82]. ICL-DAP [WB82]. Iconic [Ves89]. ICRF [EP86, IFI84, JBWW86, KM86, PT86]. ICRH [KBMV86]. IDA [Bur87]. ideal [DM86b, FBW82, GPS^{+}81, HW84b, KW84b, Man86, Ods84, TBG80, WH81]. identical [BB84c]. Identification [Bun86, OVQT^{+}84, Fia84, IA86, KAR81, KAR84b, Ran86, Ran87a, RWZ87, Ran87b]. identifying [SCC^{+}88]. IEEE [KKLS82]. IEEE-488 [KKLS82]. IHEP [Ezh84]. II [AF82, AF84c, BGS83, BGS84a, BDK86b, Bra84d, BGGW^{+}89, DSK84a, Deu84b, Fin82, Fin84a, Han84c, Han84d, HM84d, KKP82, KKP84, KN84b, KPPR84, Lab84b, MLo84, MW84b, MW84c, Nai86, PS84b, Rob84e, RP89b, SS84b, Sil84b, WP84]. III [BGS84b, BDK86c, Deu84c, Han84e, Han84f, HM84e, Lab84c, Pia84, Rob84f, SMD84, Wil84c]. illustrations [Ano87j]. ILTHII [Sal84b]. ILUBCG2 [CKA89, KA87]. ILUBCG2-11 [CKA89]. ILUCG2 [SA83, SA84]. ILUCG3 [And83b, And84b]. IMA [Gri87]. image [AWB82, BC88b, Män89]. images [Cre80b, Cre84c, MD82, RSP81, RSP84, SCC^{+}88]. imaginary [FRS83, FRS84]. immersed [FM84]. Impact [CSW84, Pou88b, AP82, AP84, GPS84, MMM84, Nob80, Nob84, Pia84, PS84b, ST81, ST84a, LS84]. impacts [Wil81]. impedance [BCM84a]. Implementation [Bas88, BM85b, Com80a, Com84, CGMS89, CM83, CM84e, FBS88, KC81a, KC84b, KN86, NR86a, ABM89a, ABG81b, Car88, DM87, DGO83, DSSS86, FS80, Hel82, Jes85, Joh89, KC81b, KC84a, MD88, MB83, MHP84, MB84, Par89b, Put89, RHM^{+}89, Suf89, WH89, Zie88, DSK84a]. Implementations [Wal89]. implemented [CG87, KW84a, MMM85]. Implementing [PCM85, Rus86, SG81, Sco89, KS84a]. implications [Krá86b]. implicit [HL88, She84]. importance [KVW84]. Improved [BI85, BT89, Cop81, Cop84b, CVvWF86, WW84a, Ber80, BL88, CCS89, Di 86, Fro84c, Joh87b, KRSV82]. improving [PIM82, PIM84a]. impurities [BSV^{+}85]. impurity [SVP86]. inclination [Fer89]. including [BS83, BS84a, Coo84b, DHN84, FNDL89, JW85]. Inclusion [Hib84b, Hib84a, Sax84b, Sax84a, Odo82, Odo84a, Odo84b, Odo84d, San82, San84, Tat80]. inclusive [MT84b, MT84a]. incoherent [SR84b]. incomplete [KPP84a, KPP84b, Oya86, KPP81]. incompressible [For89, Kir87, TL89]. incorporate [Fun84a]. incorporating [ST82a, ST84c]. increase [Hub84f]. incremental [Nag85]. indefinite [Nav89]. independent [BT89]. Index [EBS88, Ano80b, Ano80c, Ano80t, Ano80u, Ano81a, Ano81b, Ano81c, Ano81d, Ano81u, Ano81y, Ano81v, Ano81w, Ano81x, Ano82d, Ano82e, Ano82f, Ano82-28, Ano82-30, Ano82-29, Ano83b, Ano83c, Ano83d, Ano83p, Ano83q, Ano83r, Ano84c, Ano84d, Ano84e, Ano84f, Ano84i, Ano84-78, Ano84-79, Ano84-80, Ano84-81, Ano84-87, Ano85b, Ano85c, Ano85d, Ano85e, Ano85z, Ano85-27, Ano85-28, Ano86a, Ano86b, Ano86c, Ano86d, Ano86-31, Ano86-32, Ano86-33, Ano86-34, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano87x, Ano87y, Ano87z, Ano87-27, Ano87-28, Ano88c, Ano88d, Ano88e, Ano88f, Ano88-31, Ano88-32, Ano88-33, Ano88-34, Ano88-36, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89y, Ano89-29, Ano89-30, Ano89-31, Ano89-32, BLB88]. induced [BLB88, FKU^{+}84, TL84, TUWA84]. induction [Klu84b]. industrial [Fli89, KB86a]. industry [Sto89]. inelastic [AGH^{+}88, Arn84b, BKvLSE80, BKvLSE84, Bar83, Bar84e, EM84a, EE89, Fun84a, Han84d, HR86a, HR86b, Hof84a, JT84a, Keg81, Keg84, Nov87, SUZ89, SR84b]. inequalities [KS84f]. inertia [ADV81, ADV84]. inertial [MPM83, MPM84]. infinite [Hay89b, Lan80]. infinite-dimensional [Lan80]. Influence [Reg81, Vie88]. information [BHSS84, Chi89, DSK84a, Ebe84, Eck84, HM84d, Rob84f]. informative [Mir84]. infrared [MW84a, MW84b, MW84c, Wil84a]. inherent [Bir85]. inhomogeneous [Cap86a, Cha84d, Che89]. Initial [ADV81, ADV84, CR84d, DK87a, Odo84a, Odo84d, RD80, RD84]. initial-value [CR84d]. initio [CC89a, Cio89, OCC84, OCC80]. injection [ABD82, ABD84, DST80, DST84a]. INJLOK [DST80, DST84a]. Inner [Hay83, TS84a, Tho87, KB84]. inorganic [Ber84a]. INPLO [Die88]. input [Sax84a]. instabilities [KGAL88]. instability [CHL^{+}81, HR81]. instructions [Ano80g, Ano82i, Ano86f]. instrument [Han84b]. instrumentation [HR84b, Ska82]. integer [AM84b, AM84e, Di 86, KCA84]. integers [Poe85]. integrable [GSZ85]. Integral [Ano84-85, BD82, BD84b, BTS84a, BTS84b, FS84a, FC84a, Jab84, Kos80, Mor84e, MR84b, OTB80, OTB84, Pro82b, Pro82a, Pro84, Ste84c, Yur84, dB82, dB84, tR85]. Integrals [EBS88, MLS84, MLo84, Sch86d, SB82, AT84, AB88, Ben82, Ben84a, BDD84, Com82, DF84, FR86b, Gio84, Gol84b, GRRP83, GRRP84, HC83, HC84, HBJ84, Hib84b, Hib84a, Hib84d, Hib84e, Lew84b, Mat84, Moo84e, Nes88, NA84, Nob80, Nob84, OS84, Pic89, Pie82, Pie84a, Sch86c, ST84e, Stu88, TT82, TT84, WB84, Yur84, ZDN84]. Integration [GPS84, Gli82, HD86, IGB^{+}89, NBI^{+}85, Pia84, AEH87, CR84c, CR87, CM81b, CM84f, DS89, DM88b, Fog82, Fog84, GJ84, Gen82, Gru84d, Kal82, KK88, Luc84, Mor84e, NP84, Phi80, PB84b, RC85, Rys84, Sim80, Sim84, TV86, Vrb89]. Integrative [TTT84]. integrator [NP84]. integro [CSW84]. integro-differential [CSW84]. intelligence [BE89]. intensities [Bla84, Bor82, Bor84, Cop81, Cop84b, FHMR84, FK84, Fer84, Hof84b, Hof84a, Le 89, SP86, TK84, TK80]. Intensity [Deu84a, Deu84c, BB84a, BRT84, Bor82, Bor84, EHH88, Fer89, He84a, IL84, Lan84b, MT81, ÖW84, RS84e, RS84f]. Intensity-correction [Deu84c]. Inter [BDD84, KS84a]. inter-correlation [KS84a]. Inter-electron [BDD84]. interact [Sch89a]. interacting [MM88b, Rid84a]. Interaction [BH88, BGMP84, BZ84, Cis87, DS83, DS84a, DK87c, FRS83, FRS84, FT84a, FKM^{+}87, GTM87, GH83, GH84a, Gia84, GWL89, HKT85, Hib84c, HFG88, Jam84, Klo84b, KH89, KS86b, LT87, MV83, MV84b, Nai86, Poo83, Ras80b, Ras84b, Rit86, Sax84c, SNC80, SNC84, SUZ89, TTL88]. interaction-prepared [GWL89]. Interactions [Fin84b, NAS87b, Ter88, AGH^{+}88, Fin80, FS85b, HR86a, HR86b, KGAL88, Laa86, Ste84c, SAC^{+}86, YMWW89, vM84f]. Interactive [Mye88, VDM84b, AK84, BGMM81, Ber87c, BE89, Bur87, Die88, MPS81, VM89]. interchange [Van89]. interdisciplinary [Lid84a]. interelectronic [MR84b]. interface [ACR88, BH81, BH84a, BJM89, Les89, MS85, MS87a, Nag89, Zal85]. interference [FB84]. interferometer [He84a]. interior [CC84]. Intermediate [Cla84]. internal [BHM81, CH88, DHN84, DH84a, MPIM86, PR84, Sch83]. International [Bra86, Lid84a, Pou88b]. interpolation [AAMB84, BKK86, CR87, Cha82, Cha84c, CM84b, CP81, CP84, DU89, Her82a, Her84b, Moo84d, Tor82, Tor84, Wai87]. interpret [vM84g, vM84i]. Interpretation [Jam80]. interpreter [Rom81]. intersection [Wie86b]. interstellar [BS84d, SG83, SG84b]. interval [Sch85c]. intrinsic [ABH^{+}84]. introduce [OCC80, OCC84]. Introduction [Ros84b, Tou89, Eas80, FMS87, Mon88]. Invariance [DK87c]. invariant [Jab84, Mor84e]. Invariants [SM85d]. inverse [Das83, Di 87, Gel85, GH83, GH84a, KMS88, MC85]. Inversion [FS84a, FC84a, LM81, BD82, BD84b, DeG88, Gae81, Hof84a, LL84, MN88, MNR89, RRG88]. inverting [Pro82b, Pro82a, Pro84]. Investigation [KPST89, Dec89, FOW87, MLS81]. investigations [CMM84]. involved [Hei84, MLS84, MLo84]. involving [AHVV86, Bas87a, WS80]. Ion [CCR88, Aue84, BGS81, BGS84c, But84, DL87, DP88, DM86a, EE89, FRS83, FRS84, FKU^{+}84, FSR84, GS89, IP88, Irv82, KGAL88, KN84c, LH88, MRW84, MM84a, NJ82, NJ84, Sal82, Sal84e, SAVV87, SG89, TL84, TUWA84, TS84b]. ion-atom [BGS81, BGS84c]. ion-ion [SAVV87]. ion-molecule [LH88]. Ionic [Ric84, BMS87, Ric82, SF84b]. ionisation [Bor82, Bor84, Mag84a]. ionization [Mag81, MF81, MRC85, MC84a]. ionized [EK86, WW80]. IONMIX [Mac89a]. ionographic [OVQT^{+}84]. ionosphere [Bla88]. ions [AF82, AF84c, Fin82, Fin84a, FS84b, LCW84, LHP87, RS83, RS84d, Sko84, Ste84b, Tal84b, Ter88, Vaz81, Vaz84]. IPPP [ENSC86]. iron [McN81, McN84]. irradiated [Bor84, Keg85]. irradiation [Bor82]. Irreducible [Wor84b, EC84, Net84, PZCC89]. irregular [Bar81a, Cis87, Nes84]. ISAS [Nis88]. ISBN [Ano89h, Art85a, Eas84a, Eas89, Fin87, Gri87, Jes87, Rob85, Sta87b]. Iserles [Gri87]. ISING [CMO86, ABM89b, AS89, BC88a, BSDM88, Car89, CGG87, CMM84, CM86, DM89b, HDK84, HD88, NHK82, Wan87]. islands [RG86]. isoelectronic [Kno84]. Isometric [Cho84]. isospin [Hub84b]. issues [Boc87, Kun89]. iterated [Sea84b]. iteration [Cam84e, TL89]. Iterations [BBR89a, KPP^{+}89]. Iterative [BCD^{+}89, DWC^{+}89, You89, BT84a, For89, KRSV82, LP89, SC89, She84, ST82b, ST84f, WMH82]. iterative-variational [SC89]. iteratively [Tol85]. IV [Deu84d, Lab84d, LH88, NW84, SR84a, SDH84, WW84b]. Ixaru [Art85a].

J [And86, Ano89h, Bra86, Duc86, Eas84b, Eas89, Gri87, Jes87, Lis88, Sta87b]. Jacobi [CC89a, DW89, MP82]. Jacques [Man86]. JADJAD [AGH^{+}88]. JAJOM [Sar84c]. Japan [KAH^{+}89, Wat89, UI85]. Jesshope [Eas82]. jet [Gal84, Sjö82, Sjö83, Sjö84b, Sjö84a, Sjö86, SB87, Str81, TW81]. jets [Odo84a, Rit84a, Rit84b, Gal84]. jetset [Sjö86, SB87]. JJ [Cha84e]. JJ-coupling [Cha84e]. journals [Van89]. jump [Moo84d]. JWBK [Sea84b].

K-L-atomic [AGF87]. K.d.V. [vTBF84]. Kac [Cha84a]. Kadomtsev [MLS81]. Kalman [Bil89, Sta89]. Kalman-like [Bil89]. kaon [Lan84a, Lan82]. Kármán [Sch84b]. KBr [vdM84a]. Keller [Eas84b]. Kerbel [Sta87b]. kernels [Eri85, LM85]. Kerr [ACR88]. Kerr-like [ACR88]. keV [SMMP86]. key [WHD86]. Killeen [Eas84b, Sta87b]. Killingbeck [Duc86]. kind [Köl81b, Köl84b, Kos80, PJ89, tR85]. Kinematical [WMC84]. Kinematics [Ass84, Bra84c, Bra84b, DM87]. Kinetic [BLM^{+}88, Sta87b, AS89]. kinetics [Das83, DST84b, HLM^{+}87, Rob80b, Rob84c]. kink [Sch83]. Kirchhoff [Art85b]. Klein [Bar81b, Bar84c]. Knowledge [Cam86, Stu85, BCF^{+}89, DG89, Mik86]. Knowledge-based [Cam86, Stu85, BCF^{+}89, Mik86]. Korteweg [SBFI87]. Kostomarov [Sch86a]. Kotani [MG84]. Kramers [CT84b, Col84b, KN84a]. Kronig [CT84b, Col84b, KN84a]. Kubik [Pis84b, Pis84a]. KWIC [Ano89y].

L [AGF87, Gas81]. L-harmonics [Gas81]. labelling [MPS84]. laboratory [CSMHT88, HR86a, HR86b, KKLS82, THH82]. ladder [Sil84b]. Lagrange [Cro88, TBB87, Tre88]. Lagrangian [Atz86, Fri88, MT81, WHW87]. Laguerre [Tak88b]. LAN [Gam89]. Lanczos [CKW89, GWL89, HW84a, KPP^{+}89, KWPP89, Nex89, NO89, PNO89, Sco89].

Lanczos-like [Sco89]. Landau [KS08, Bie83, Bie84, Kam84, KS84d, KS84e, Sch84a]. Landau-Fluxoids [Kam84]. Langdon [Taj86]. Langevin [GS88]. Language [PCM85, WHD86, Zal85, BF89, CRV^{+}89, CPL89, LK89, LU83, LU84, WPVS89]. Languages [IKM85, Per82, Dow82a, LT85, Rus87, Sch85a, Sch88c, Sch85b, Whi89a, Wil86]. Lanthanide [Ste84a]. Laplace [MN88]. Laplacian [O’C84a]. Large [Blo82a, Blo82b, BBC^{+}89, Duc80, FBW82, FS80, GWTM83, Kas82, MR87, Pat82, BWBL88, BMR85b, BLM^{+}88, BDK86c, Boc87, CH87, CGG87, CH84c, GDK89, Hoe86b, KLS85, Man87b, Mas83, MKRJ89, Poe85, Saa89, SN84, Sch87a, SC89, Sch85c, SM89b, TTT87, Tol85]. Large-scale [Kas82, MR87, BLM^{+}88, CH84c, SN84]. largest [HD86]. laser [Atz86, Bor82, Bor84, BLB88, CW84c, CAR84b, DST80, DST84a, DST84b, DH84b, GH83, GH84a, MT81, MCJ81, MCJ84, Rob80b, Rob84c, RS84a, Smi84a, TGA88, TW89b]. LATEN [JP84b, JP80]. Lateral [Ugn80, MRW84]. latitude [KGAL88]. Lattice [Ano84-83, BMR85a, KN84c, KN84b, NJ82, NW84, NJ84, Sch87a, Ano87b, Ano89x, AM81, AMC83, AM84d, AMC84, Bai87, BM82a, BM82b, BM82c, BCM83a, BMR84, BM84a, BLR82, BLR84, BKHJ83, BPR^{+}86, BS86b, BCCM87, CM85, CGGK85, CCS89, CMS83, CMS84, CM83, CM84e, DMM84a, DM87, DMM83, DMM84b, EMM81, FAK82, HM81, HM83, HM84a, HM84b, Hoe86a, HM84c, IHS^{+}88, JP80, JP84b, Ken89a, KLS^{+}89, KM85, Mac89b, MMS86, May81, May84, MB83, MHP84, MB84, MT86, NHK82, NFH^{+}88, ÖW84, Oya86, RL80, Sal85, VBD88, WW84a, WW84c]. lattices [DM89b]. Laue [CA84, Pre84a, Pre84b]. laws [Krá86b]. Lax [EFES81, Ito85]. layer [AP80, RS84e]. layer-by-layer [RS84e]. layered [Chr84b]. layering [Ric86]. layers [GS83, Sko84]. layout [Ano84i]. LCGO [WC84]. LCGTO [BT84a]. LEAD [Bla84]. leading [San82, San84, Vrb89]. learned [FP89]. learning [Köb89]. least [Bok84, CAP86b, Han84f, LM84, O’C84b, RFSG84b, Sha87, WS84b, Woo86b, vM84c]. least-squares [CAP86b, RFSG84b, vM84c]. Leed [RS84e, Hof84b, Hof84a, RP89b, RP89a, RS84f, SP86, TK84, TK80]. Legendre [Bra84a, Del84a, SNC80, SNC84]. Leiss [ASM84, ASM80]. lengths [JP84a]. lenses [Len80]. LEP [Odo81, Odo84c, Wil89]. lepton [BDK86b]. leptonic [LO86]. letter [Ram85]. level [AFH84, AF84a, AF84b, BCKO89, CDF^{+}89, LK89, MN80, Par89b, RFSG84a, RFSG84b, Vio84, WPVS89, Wil84b]. levels [CH87, CH88, DK87b, JT83, JT84b, LHP87, MGN80, MGN84, SDH84, Ten83, Ten84b, Ten84a, Ten86]. Levenberg [BP84b]. LEVI [Ran87b]. Levin [BRB89, Gro89b]. Li [LH88]. liana [Smi84b]. libraries [Her84a, Nev82, Sta87a]. library [Ano86n, BJ81a, BJ81b, BJ84a, BJ84b, DD85a, Jac80, Jac85b, Jac85a, LB82a, SFB87]. Lie [Ran87a, Sch84c, CT89, DVD87, EFK85, FK86a, Ran86, RWZ87, Ran87b, Sch82b, Sch84b]. lies [Woo85]. life [PIM82, PIM84a, Tau85]. lifetime [KEMP81, KE84b, KEMP84, KE84a, Puf83, Puf84, VDM84b, War84]. lifetimes [IL84, MC81, MC84b]. light [HE84b, RS84a]. like [ACR88, Bil89, CC89b, IR80, LHP87, MLS84, MLo84, Moo84h, Sch85c, Sco89, Tal84b]. likelihood [Jam80]. limit [BGMP84, GB87, GKM^{+}81, Jam88]. limitations [Kal83]. limited [Han84c, Han84e]. limits [GPS^{+}81]. LINDA [DKA85]. line [BS84c, CRH^{+}89, CGM89, CGGK85, CGMS89, DG89, Duf84, GPS84, Kri89, KK84b, Mas81, PF86, PS84b, RF83, RF84, SN84, SdAD^{+}88, Sta87a, Tal82, Tal84a, Van81a]. Linear [Bir85, Gou88, HR81, HW84b, AS83, And83b, And83a, AS84, And84b, And84a, AKS88a, AKS88b, AHVV86, CL84, CKA89, EPS81, Ede84, FR82, HVS87, Jes84, Kir87, KS84f, KA87, Kon84, KRSV82, Lan80, MGD81, Man86, MCA87, MGK86, MBOK86, Nav89, Pro82b, Pro82a, Pro84, Rit86, RRG88, RS80, RS84c, SC89, SA83, SA84, SK82, SK84, TC87, WMH82]. linearization [SFP86]. lines [Ata85, BS84d, CCR88, MMO89]. link [GDK89, MRT89]. link-cell [MRT89]. Linkage [Com80b]. linked [WS80]. links [BBD^{+}89]. Liquid [PI84, AGF87, CLL89, Fre87, GTG81, GTG84, GTM85, GTM87, GTMA88, Isl81, Isl84a, Isl84b, ZR89]. liquids [AF82, AF84c, Fin80, Fin82, Fin84a, Fin84b, Laa86, VH87]. LISP [BGM87, CT89]. List [Ano80e, Ano82x, Ano82y, Ano85u, Ano85v, Ano86-28, Ano88x, Ano88w, Ano88y, Ano88z, Ano89z, Ano80f, Ano88h]. lists [KS81, KS84b]. lithium [Keg85]. living [Phi86]. Liviu [Art85a]. loading [KL84b]. Local [Ger88, HC88, Hel80, Cis87, FH88a, KKLP84, LZ84, OCC80, OCC84, Ste84c, WGM83, WGM84, You87, ZL84]. localization [KM86, Sch87b]. Localized [BWBL88]. Locating [BP84b]. location [HAG^{+}89]. locking [DST80, DST84a]. log [DK87a]. logarithm [Köl84c, MC85]. logic [DGK^{+}81, Gje81, KK85]. London [And86, Ano89h, Fin87]. long [FS82, FS85b, MKRJ89]. loop [BCM83a, RFS89, Stu88]. loops [CMS83, CMS84, ST89b]. loosely [Bai87, Gro87, Gro89a]. Lorentz [Jab84]. Lorentz-invariant [Jab84]. Lorentzians [vM84c]. loss [AP82, AP84, CJ84, SMMP86, Sko84]. LOUHI78 [RS80, RS84c]. Low [BMS87, Eas88, JHRR87, BGMP84, Bla84, BKR88, Fri88, MRC85, RP89b, RP89a]. low-dissipation [BKR88]. Löwdin [MV81, MV84a]. lower [CCR88, OSF86, SSF^{+}89]. lower-hybrid [SSF^{+}89]. lowest [BDK86b, BDK86c]. LPOTp [PSL88]. LPOTT [Lan82, Lan84a]. LS [Cla84, Dya86, Rid82, Rid84b]. LS-coupling [Rid82, Rid84b]. LSFBTR [Tal83, Tal84c]. LSI [Ska82]. LSI-based [Ska82]. Ltd [Duc86]. LTE [Hub88, Mac89a]. LUCIFER [IW87]. LUCY [MD88]. lumped [Ver89]. Lund [BS87, NAS87b, Sjö82, Sjö83, Ben84b, BI85, Sjö84b, Sjö84a, Sjö86, SB87]. lying [SDH84].

M [Eas84b, FV81, Gri87, Nad86a, Sta87b]. M.Q.D.T. [RA82, RA84]. M1 [BF84, HK83, HK84b]. MAC [Del87]. machine [AS89, Bar81a, BGM87, CKT89, Ost85, Gur85]. Machines [Mac89b, Buz85, DD85a, Köb89, LS84, Tem82]. Mackeown [Lis88]. macpar [Rit86]. macroassembler [Tar86]. macromolecules [LNV84]. macroparticle [Rit86]. macroscopic [Tat80]. MACSYMA [Ros85, RW86]. made [He84a]. Madelung [JP84a, ÖW84]. magnet [ABD82, ABD84]. Magnetic [CO86, Lee82, Lee84, LB82b, BCFK85, BM84b, BJ81a, BJ84a, CT82, EPS81, Frü81, Gol84a, Hub84e, Hub84f, JBWW86, Kar80, Kar84a, Klu84a, KP84, McN81, McN84, Mik84, MMO89, ML81, Sea84a, vM84e, Nai86]. Magnetically [Sta87b]. magnetisation [BC88a]. magneto [TDS81, TDS84]. magneto-plasma [TDS81, TDS84]. magnetohydrodynamic [ABG^{+}84, IK81, HH82, HL86, Lao84]. magnetohydrodynamics [Goe81, Man86]. magnetoinductive [GTL^{+}86]. magnetosonic [PPHS86]. Magnetosphere [Bla88, KGAL88]. Magnetosphere-ionosphere [Bla88]. magnetotail [Swi88]. magnetotelluric [KS84c]. mainframe [Ast87]. mainframes [Mee84]. maintain [BJ81a, BJ81b, BJ84a, BJ84b, Jac85b, Jac85a]. maintenance [CHS87, PS88, Tau84a]. Making [BvSW84, HR84b, JHRR87, SDH84]. man [Ano89a]. management [BBR89b, Eld84, Gau84c, Kun89, LK89, Maz89, MP89, OLR^{+}87, Pla86, QTZ^{+}87, Ros84b, Rus86]. managing [Man87b, Nas87a]. Manchester [Gur85]. mandy [SS84b, SMD84]. Manipulation [FK86b, BO80, BEM84, DNR83, DNR84, SMM89, III88, YP89]. manipulations [CT89]. Manual [DFD81, DFD84, EK84, Pis84b]. Many [Rap85, Gir81, Lew84d, MGM87, MBW89, NP84, SMM89, Sil84a, Sil84b, WS80, Wil84c]. Many-body [Rap85, MBW89, Sil84a, Sil84b, WS80, Wil84c]. many-electron [Lew84d]. map [CM88]. Maple [Gro89b]. Mapping [Wie86b, Goe81, KK83a]. maps [CM88, ST84d]. Marcel [Art85b]. March [And86]. mark [BGGW^{+}89]. Markov [Kra86a]. Marquardt [BP84b]. Maser [Le 88]. mass [DSSS86, JW85, KL84b, TCS85, Wil81]. masses [Jab84, KMS88, WwYpCg88]. massively [LTK88, Mat87, Ves89, Gal85]. master [Ano84i]. matching [BSP84, Hof84b]. materials [DS83, DS84a, Han84b, Han84c, Han84e, Rei87, Sta89]. Mathematical [Nev82, But89, HHK86, LU83, LU84]. mathematics [Ano89a]. matrices [BH82, BH84b, BTS84b, BS86b, Cho84, Gad85, Gru80, Gru84b, Han84a, Hen84, KLS85, MG84, Nex80a, OTB80, OTB84, Ruf85, Sar84a, Sar84b, Sar84c, SM89b, SM81b, SM84e, SM85b]. Matrix [BH88, DR82, Hod84, ST84c, SFP86, Sho84, AS83, And83b, And83a, AS84, And84b, And84a, AKS88a, AKS88b, Arn84b, BBM82, BBM84, Bas87b, BBC^{+}84, BBL^{+}84, BOT86, Bou87, BC84c, Bra84d, Bru85, Bru86, BW87, CH87, Cha84e, Cla84, Das89, Dav89, DeG88, DK87c, EEV89, FH88a, FS84c, Gel85, Gru84a, HM84e, Jam84, Klo84b, Klo84c, Kol81a, Kol84a, KA87, KM82, KM84b, Lab84a, Lew84d, MP81a, MP84a, Mor84c, Nas84, Nex89, Par84, PT82, PB84b, Pol88, PGB84, RW84, Rob84a, SP87, Sal84c, SS84a, Sax84b, Sax84a, Sax84c, ST84b, SG81, SB82, ST82a, SA83, SA84, Sho81, SMC81, SMC84, Sta84, Sun88, TAK^{+}85b, Ten85, TBB87, Tol86, YP89, Hof84a]. MATSUP [Ben82, Ben84a]. matter [RR84, SMMP86, Sko84]. Maxentwdf [DL86]. maximally [IB87]. maximum [DL86, SM82, SM84f]. Maxwell [LW84a, MN80]. Maxwellian [LW84a, TDS81, TDS84, WL84]. may [De 84c]. MB [Nog83]. MC [PBKW85]. MCBP [BGMP84]. McCoy [Sta87b]. McGraw [Taj86]. McGraw-Hill [Taj86]. MCHF [FS84d, HFG88]. MCL [CPL89]. McMOLDYNH [Laa86]. McMOLDYNH/2O [Laa86]. MD [AB86, BA86, MRT89, PBKW85]. MDAI [NP84]. MDIONS [AF82, Fin82, Fin84a]. me [Lev89]. means [Hel80, IKRT89, tR85]. Measured [Cha85]. Measurement [Per89, FS82]. measurements [AGG^{+}82, BCM84a, FR84, Frü81, HR84b, MC81, MC84b, PDNE88]. Measuring [MT86, TV86, VKM^{+}82]. mechanical [ACC^{+}86, BTS84b, DWC^{+}89, NBI^{+}85, OTB80, OTB84, RSA84]. mechanics [And86, Ano82w, CCD85, Lee89, LL84, Duc86, Eas80]. mechanisms [ENSC86]. media [Che89, Chr84a, Chr84b, KK84b, Sch87b, Ugn80]. medium [Car80, ELS88, vM84d]. medium-scale [Car80, vM84d]. Medusa [CAR84b]. Meeting [Kil80]. Meiko [BDJ^{+}89]. Mellin [MN88]. memories [Köb89, Par89a]. memory [Dob85, GWTM83, LK89, LS84, Nas89, Ott89]. meridional [Cap86a]. MERLIN [PCL89, ERLD87]. MERLIN-2.0 [PCL89]. mesh [BSA86, EHL80, EHL84, Luc84, Pis84a, Pis84b]. meshes [Wie86b]. meson [HR84a]. metal [BCFK85, CR84b]. metals [DBB84, Isl81, Isl84a, Isl84b, MV83, MV84b, RMV89]. Metcalf [Nad86a]. meteorology [Nav89]. Method [FBS88, FBS89, Rib80, ACD88, ASM80, ASM84, AGR88, ABG^{+}84, AI86, Ata85, AZK83, BM82a, Bar82, Bar84b, Bas87b, Bas87c, BP82a, BP82b, BP84b, BP84a, BD80b, BD81, BD84c, Ber80, BBC^{+}84, BBL^{+}84, BWD84, BL88, BdCP81, BKR88, CL84, CH87, CH88, CR84c, CR87, Col80b, CM84d, CT84b, Col84b, Cro88, DW88, DHD84, DGO83, DU89, DRRvB82, EFES81, EFES83, EA89, FS84a, FV81, Gad80, GR81, Got83, Hay83, Hay89a, HRS81, HRS84, HM81, HM83, HM84a, HM84b, HvRM86, Hof84a, HVS87, HLM^{+}87, IB87, Joh87b, Joh88, KKP82, KPH84, KKP84, KVW84, Kla84, Kon84, KRS83, KK83b, KWPP89, KMS88, KPP81, KPP84a, KPP84b, Lao84, LK88, LNB^{+}88, Len80, LT87, Meu89, Mik80, Moh80, Moh84, Moo84d, MB83, MB84, Nex84, NO89]. method [PR87, PPHS86, PT82, PZCC89, Pro82a, RW84, RC85, RRG88, Rud84, RS84e, SAC87, SC89, ST82a, ST84c, She84, Sim80, Sim84, ST82b, ST84f, Sun88, TAK^{+}85b, Tau80, Tau84c, Tel84, Voh88, Wal89, WC84, Zlá80, vTBF84]. methodically [Cam85]. methodologies [Dow82b, Was85]. Methods [Ano81t, Art85a, BCD^{+}89, Buc85, DM86b, Fin87, FBS88, Sta87b, You89, Ano82w, Ano87-29, AHVV86, BKvLSE80, BKvLSE84, BT81, CLL89, Dav89, DGPM85, Dor86, DWC^{+}89, Eas84a, Eas86, Fla82, Fog82, For89, FOW87, GWL89, Goe81, GS83, HW84a, Her88a, Hoa89, Ixa80, IR85, JL84, Key89, Lan80, LP89, Lis84, Man86, Man83a, Man87a, Mar89a, MPR80, MP82, MRT89, Mur88, Nex80a, Oya86, RFS89, RC87, RL80, Sch85c, Sch84d, Suf89, TTW84a, Tem82, WMN85, ZNO88, vD87b, vdV89]. metrological [PS88]. Metropolis [KN86]. metropolitan [BCC^{+}89]. MeV [RR84]. MF [MPM85]. MF-FIRE [MPM85]. MFFT [NR86b]. MFFT4 [NRS88]. MFP [PM80, PM84a]. MHD [BLM^{+}88, BHM81, CR84a, DM86b, Eas88, Ede84, GPS^{+}81, HW84b, KJ84, KW84b, Kir87, LCH^{+}81, MGD81, MTN^{+}82, MN84b, MBOK86, Pou88a, RG86, SS81, SW80a, Sch84d, Str81, Syk81, TTAT81, TAK^{+}85b, TBG80, TTM^{+}80, TT89, TW81, Uga88, WH81, WBG84]. MHD-calculations [SW80a]. Micro [DSK83, DSK84b]. Micro/mini [DSK83, DSK84b]. microcanonical [CM83, CM84e, HD88]. Microcomputer [AN88, FV82, Gli82, Hal88, KS81, KS84b, WML^{+}82, Duc86]. Microcomputers [Ano88-27, CG87, Dow82a, Dow82b, Ger88, MGM87]. Microprocessor [Dag82, Her88b, Ska82, BvEF85, Rom81, VKM^{+}82]. Microprocessor- [Ska82]. Microprocessors [PS88, Wil86]. microprogrammable [Jac81, Jac82]. microprogrammed [BGMM81]. Microscopic [EE89, FKU^{+}84, Fun84a, vM84f]. Mida [MS87b]. migrating [WD84]. migration [KSG82]. Milano [BLMU81]. mild [NF83, NF84]. millimeter [Smi84a]. MIMD [ACSB85, Buz85, Hey89, JM82, RS85]. Mincer [GLST89]. mini [DSK83, DSK84b, KKLS82]. Minicomputer [KAR81, KAR84b]. minicomputers [BBE^{+}81, FS80, HBJ84]. minima [BP84b]. minimal [Cio89]. minimalisation [LM81]. minimization [Buc83, Buc84, ERLD87, JR84, Tro84]. minimum [Jam80]. MINUIT [Les89, DCC84]. Minut [JR84]. Mirin [Sta87b]. mirror [CO86, JBWW86]. mirrors [PT86]. MISHA [Rys87]. missing [TS89]. mission [Vie88]. missions [Wen88]. mixed [BF84, MP81a, MP84a, RSL^{+}84]. MIXERG [PM83, PM84b]. mixing [HK83, HK84b, KF84, MRCL84]. mixture [PSBG84, Zlo84d]. mixtures [Tho84]. mode [FH88a, HC88, Han84a, KT81, KLGT85, Mir84, MPIM86, Sch83]. Model [DV85, FJ89, AFH84, AF84a, AF84b, AXG87, AS89, Aue84, BTWN87, Bet84, BJM89, BCCM87, Car89, Cha84b, CHMM84, Coo84b, CMM84, CM86, CMO86, DTE84, DM89b, ETH84, FP89, FH88a, Gib85, GE88, GA84, Har81, Har84a, HD88, HMS84, Hof84a, HGW86, HMTO84, HT84, Hub88, Kas84, KBMV86, KN84c, KN84b, LJN85, MMBW88, Man83b, Nic89, NJ82, NW84, NJ84, Odo84d, Ods80, PIM80, PIM84b, PI84, Ric82, Ric84, RMV89, Sal83b, Sal84f, Sal85, ST82a, ST84c, SK86, Smi84a, Vaj82, Vaz81, Vaz84, Wan87, WGM83, WGM84, WW84b, Zwa85, SBB^{+}89]. Modeling [Abr88, Suc87, Wie86a, vD87a]. Modelling [Put89, ASS82, BT89, Bra87a, JB80, KS84c, Rob84d, Rob84e, Rob84f, Ros84a, SAVV87, Tha89, VM89]. Models [Dic82, Sta87b, AEH87, ADV81, ADV84, BPR^{+}86, CGG87, Fre87, HDK84, HA82, HA84b, Klu84b, LW84b, MGD81, McI84, Par89a, SS89, SAC^{+}86, Vir89, vMF81, vMF84]. Modern [Was85, Zac81, Sch85b]. modes [BLM^{+}88, EPS81, Han84d, Wor84a]. Modification [Cop84d, Cop84c, Bar84d, BP84b, HM83, HM84a]. modifications [GH83, GH84a]. Modified [TB87, BMR85b, Bra84a, O’C84b, SSS84, SBFI87, Tho87, Tho04a]. Modula [Nie88, Pyl85]. Modula-2 [Nie88]. modular [Ska82, Zal85]. modulation [Ber84b]. module [BCF^{+}89, LW85]. Moions [AF84c]. Moldauer [SS84b].

Molecular [CC89b, FR81, Fra83b, Fra84a, GG89, ZR89, Ano81t, BWBL88, BKvLSE80, BKvLSE84, Ber87a, Bin87, BR85, BL86, CB88, CLL89, DWVH89, Fin80, Fin84b, Fin86, GPS84, Gel85, GDK89, Gup88b, HSV84, HVS87, HT84, HS84, HSB^{+}84, Jam87, JL84, Kas82, Mor80, Mor84d, MKRJ89, Pia84, PS84b, RFS89, RMV89, Ryn82, Ryn84, SS89, SNC80, SNC84, Sch89c, Ská81, Ská84, TBB87, VSH83, Wor84a]. molecule [DK87a, DK87b, FT84a, Gia84, LH88, MLL86, Nai84, Nai86, Ras80b, Ras84b, SP87, SNC80, SNC84]. Molecules [Fin84b, BB84a, BOT86, Bou87, Fin80, Gup88b, HC88, Han84a, Han84f, Han84d, HVS87, JT83, JT84a, JT84b, Lad86, LBY86, MBW89, Nog83, Ogi83, Ogi84, OCC80, OCC84, Ras80a, Ras84a, Sal84c, ST81, ST84a, Sch88a, Sea84a, SLH84, Sil84a, Sil84b, Sin84, SMT88, TT82, TT84, Ten86, TC87, TM89, WS80, Wil84c]. Molforce [Gel85]. Molsimil [CC89b]. Molsimil-88 [CC89b]. MOMCON [HL86]. moment [ADV81, ADV84, LWHH82, LWHH84, Lao84, Lee89]. momenta [SS84b, TTW84b]. moments [CDN^{+}87, DK87c, Gae81, MN88, MNR89, SS89]. momentum [Bra87a, CW80, CDW82, CW84a, CDW84, ET84, Gra84b, Gra84c, HKT85, Hib84b, Hib84a, Hib84d, Hib84e, Hub84a, Jab84, Lan82, Lan84a, LW84b, LT87, PSL88, Rao81, RV84, RRC89, Sal82, Sal84e, TTW84a, Tak85a, Tam84, VVM84]. momentum-space [Jab84]. Monitoring [Ste87]. monochromator [BPSB88]. monoconfigurational [CC89a]. monoenergetic [Lan84b]. monoxide [ZR89]. Monte [Ben84b, BI85, Sjö84b, Sjö84a, Sjö86, SB87, AS89, ANP^{+}85, AM81, AMC83, AM84d, AMC84, BM82a, BM82b, BM82c, BMR84, BM84a, BS87, BKJ83, BKJ84, BDK86a, BDK86b, BDK86c, BLR82, BLR84, CMR86, CFG^{+}88, CM85, CGGK85, Cha85, CGG87, Chr84b, CMS83, CMS84, Cop81, Cop84a, Cop84d, Cop84c, Cop84b, CVvWF86, CGB^{+}89, CM83, CM84e, DMM84a, DLB83, DLB84, Dec89, Del85, DM87, DMM83, DMM84b, DKA85, DS84b, DD86, EMM81, EA83, EA84, Fer89, FS85b, GTM87, GHO^{+}89, Gup88a, HR84a, HR86a, HR86b, HP83, HP84, HMTO84, IW87, Ing87, Jad84, JW85, Kas84, Kaw86, KVW84, KSE86, LM81, Lux80, MMS87, Mar89a, Mar89b, Miu87, MGM87, MB83, MB84, MT86, Mor84e, NFH^{+}88, NAS87b, Odo81, Odo82, Odo84a, Odo84b, Odo84c, Odo84d]. Monte [RR84, RSS85, RL80, Rit84a, Rit84b, RS83, RS84d, Ryn82, Ryn84, SMMP86, Sal85, San82, San84, Sjö82, Sjö83, Tat80, TKS^{+}85, Tou89, VH87, Voh88, Vrb89, Wan87, YMWW89, ZHR81]. Monte-Carlo [AMC84, HP84]. Morang [DRD88]. morphology [DRD88]. Morse [PB84a]. Moshinsky [LJ84, YpMzCeCg85]. Mössbauer [BPA84, CT84a, Gro84, MGTD84, Ste84d, Ver83, Ver84a, WS84b]. motion [AB86, BPW84a, BPW84b, BPW84c, BMBW85, GTL^{+}86, Gio84, WDM84]. Motorola [BvEF85]. moveout [SM89a]. movie [JHRR87]. movies [HB88]. moving [BH81, BH84a]. MP [KM85, NR86a, RS85]. MPS [SEF^{+}89]. MRVAC [Rud84]. MS [Klo84a]. MS-X [Klo84a]. MSXALPHA [Klo82]. Multi [Cla82a, CGG87, Fro84a, Fro84b, Fro84c, Gib85, GHO^{+}89, KK88, PB84b, ZHR81, RSMTV88, TGA88]. multi-configuration [Fro84a, Fro84b, Fro84c]. multi-dimensional [KK88]. multi-particle [PB84b]. Multi-point [Cla82a]. multi-spin [CGG87]. multi-spin-coding [ZHR81]. multi-tasking [Gib85]. multi-transputer [GHO^{+}89]. multichannel [BB86, FJ89]. multiconfiguration [Des84b, MGN80, MGN84]. multiconfigurational [GMN^{+}80, GMN^{+}84, Lab84c]. Multidimensional [BKK86, BT88, DR85, NP84, NR86a, ERLD87, Fog82, JSB82, NKS88, Phi80, Vis84a, Vis84b, ZDN84]. multifrequency [MPM85]. multigrid [Dec89, Par89b]. Multigroup [WMCH84, RSMTV88]. multilayer [TGA88]. Multilevel [FBS88, FBS89]. Multiloop [ST89b, GLST89]. Multimicroprocessor [PFGP82]. Multiparabolic [MCA87]. Multiparameter [BT84a, Ska82]. multiparticle [DV85, KSE86, Sau85]. multiperipheral [BDK86a]. multiphase [TF89]. Multiphonon [Rei85]. multiphoton [BSdlT87]. multiple [AP89, CD86, Cop81, Cop84a, Cop84d, Cop84c, Cop84b, CVvWF86, CH84d, Gen82, LO86, RFS89]. multiplication [Ber87b, Cli82]. multiplicity [ADO83, ADO84a, ADO84b, BK84b, KB84, TS84a, Tho87]. multiplier [Wor84b]. multipolarities [RSL^{+}84]. multipole [AGR88, Arn84b, Com82, SS89, Tol86]. Multiprocessing [ABG81b, MSC^{+}88]. Multiprocessor [Her82b, BAA^{+}87, CSMHT88, GAA^{+}87, HBF^{+}85, Sli89, SdAD^{+}88]. Multiprocessors [Poh87b]. MULTIQUARK [Wro82, Wro84]. multispecies [MMK81, MMK84, MMTK88]. multispin [Wan87]. Multistate [GPS84, Pia84, PS84b]. multistep [JL84]. multitasking [AHKM86, Cha85, KM85]. Multivariate [O’C84b, FMS87]. muon [Rin84]. Muonic [AV84, Rin84]. muonic-atom [Rin84]. Murnaghan [EC84]. Muxworthy [Rob85].

N [Rob80b, Sch86a, Her88a, Rob84c, SM82, SM84f]. N-body [Her88a]. N.F.E [Isl84b]. NA-11 [DGK^{+}81]. NAG [Bas88]. namelist [Day81, Day84a]. nanosecond [RS84a]. national [Bra87b, Ebe84, Lid84b, Pou88b]. natural [LU83, LU84]. Navier [Gaj85, KP89]. NBCS [MV87]. near [DPH82, DPH84, VSP86]. near-edge [DPH82, DPH84, VSP86]. Nearly [Mar84a, Kal83]. necessary [Han84c, Kas82]. need [Jam86a, Jes86]. Néel [Hub84e, Hub84f]. Negative [BS84b, BS80, NT84]. neighbors [Voh88]. net [DL89]. Network [Per87, BCC^{+}89, BCF^{+}89, CH84c, Cot87, Ghi87, Gre88, Run89]. Networking [KAH^{+}89, Str87, CFG^{+}87, Flu89, Hut89, Mon87]. Networks [Ger88, New89, CHS^{+}89, Den88, Köb89, Par89a, Pou88b, Rue89]. Neumann [MP81a, MP84a]. Neural [CHS^{+}89, Den88, Köb89, Par89a, TV89, WH89, DL89]. neutral [HP83, HP84]. neutrino [AI86]. Neutron [Ver89, Cha85, CSC^{+}86, Cop81, Cop84a, Cop84d, Cop84c, Cop84b, CVvWF86, Del85, FHMR84, Fun84a, Fun84b, Han84b, Han84d, Keg81, Keg84, Moh89, SR84b]. neutrons [BSA86, Keg85]. Newman [Lis88]. Newton [Her82a, Her84b, Lan80, Sch85c]. Newton-like [Sch85c]. next [San82, San84]. Nicer [BN81]. NIEM [HRS81, HRS84]. nil [Ran86, Ran87a]. Nilsson [BF84, Hir84, MVV83, MVV84]. nitrogen [Wie86a]. NJGRAF [BSK88]. NJSYM [BSK88, SH82, SH84]. NMR [AP89, Gal81, WDM84, vMT84, vM84h]. no [Bor89, PV84, Ran87a]. no-recoil [PV84]. noble [PM80, PM84a]. nodes [Ghi87]. Noise [Kal83, WSV88]. noise-containing [WSV88]. noisy [Bro83, Pro82b, Pro84]. Non [BMV83, BMV84, Fla82, HFG88, SS84d, BZ84, BvSW84, DWC^{+}89, EPS81, Ede84, FR82, FH88a, Han84b, Han84c, Han84e, Hay89a, HP83, HP84, Hub88, JSB82, KW84b, KLS85, Kir87, KK83b, KR88, Lab84a, Mac89a, OY80, Puf83, Puf84, Rap81a, ST84e, ST82b, ST84f, Ugn80, Wai87, War84, WW80, WB84, Zlo81, Zlo84c]. non-Abelian [KR88]. Non-adiabatic [BMV83, BMV84]. non-chemist [BvSW84]. non-circular [HP83, HP84]. non-compact [Rap81a]. non-crystalline [Han84b, Han84c, Han84e]. non-exchange [ST84e]. non-Gaussian [Puf83, Puf84, War84]. non-Hermitian [KLS85]. non-hierarchical [OY80]. non-homogeneous [Ugn80]. non-ideal [KW84b]. non-iterative [ST82b, ST84f]. non-linear [EPS81, Ede84, FR82, Kir87]. non-local [FH88a]. non-LTE [Hub88, Mac89a]. Non-numerical [Fla82]. Non-orthogonal [HFG88]. non-orthogonality [Hay89a]. non-parametric [JSB82]. non-periodic [KK83b]. Non-relativistic [SS84d, BZ84, Lab84a, WB84]. non-sparse [DWC^{+}89]. non-stationary [WW80, Zlo81, Zlo84c]. non-uniformly [Wai87]. nonequilibrium [Lee89]. nonhyperbolicity [TF89]. nonideal [CHL^{+}81]. Nonlinear [Pet84, RW84, Ter88, ACR88, BVL84, BBR89a, FBW82, GSZ85, HR81, HW84b, Ito86, LM84, Mar89a, MMF81, MMK81, MMK84, MMTK88, RWK^{+}85, RS80, RS84c, Sch88a, Sch85c, Sha87, Syk81, TF89]. nonlinearity [ACR88]. nonlocal [Fie80, Fie81, Fie84, LM85, MZZ89]. nonparametric [Chi89]. nonproblem [TF89]. nonrelativistic [BRTB84, BRT84, ZL84]. nonspherical [Gup88b]. Nonsymmetric [CKW89, OJK89, BMK89, CKA89, Saa89]. Nonuniform [AP80]. Normal [De 84a, KLGT85, Han84a, Han84d]. normalised [PB84a]. normalization [Han84b]. North [Bra86, Tru88]. North-Holland [Bra86, Tru88]. notation [FS84c, Lab84a, Sax84c]. note [Ano86-35, Ano88i]. Notes [And86]. Notice [Ano80r, Ano83o, Ano84-66, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano81o, Ano81p, Ano81q, Ano81r, Ano81s, Ano82t, Ano82u, Ano82v, Ano83k, Ano83l, Ano83m, Ano83n, Ano84x, Ano84-30, Ano84-31, Ano84-32, Ano84-33, Ano84-34, Ano84-35, Ano84-36, Ano84-37, Ano84-38, Ano84-39, Ano84-40, Ano84-41, Ano84-42, Ano84-43, Ano84-44, Ano84-45, Ano84-46, Ano84-47, Ano84-48, Ano84-49, Ano84-50, Ano84-51, Ano84-52, Ano84y, Ano84z, Ano84-27, Ano84-28, Ano84-29, Ano84-53, Ano84-54, Ano84-55, Ano84-56, Ano84-57, Ano84-58, Ano84-59, Ano84-60, Ano84-61, Ano84-62, Ano84-63, Ano84-64, Ano84-65, Ano84-67, Ano84-68, Ano84-69, Ano84-70, Ano84-71, Ano85o, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano86u, Ano86v, Ano86w, Ano86x, Ano86y, Ano86z, Ano86-27, Ano87u, Ano87v, Ano87w, Ano88u, Ano89t, Ano89u, Ano89v]. notice [Ano89w, Ran87a]. notices [Ano84-72, Ano84-73, Ano84-74, Ano84-75]. nozzle [FCW84, NF83, NF84]. NSF [Bra87b]. NSPCG [OJK89]. NUCADA [HA83, HA84a]. Nuclear [Hla80, Hla84, LCS84, Sko84, Smi84c, Smi84d, Smi84e, AAH84, ABH^{+}84, AGG^{+}82, Arn84b, BTWN87, BT81, Bas87a, Bet84, CDN^{+}87, DG84, Das83, DTE84, DKA85, ETH84, FR86a, FB84, FNDL89, Gad80, GDH^{+}89, Gir81, Har81, Har84a, Hei84, HA82, HA83, HA84b, HA84a, HMS84, Her84a, HLM^{+}87, KK85, Lea87, LW84b, MMBW88, MVV83, MVV84, MV87, Mir84, MC81, MC84b, OVQT^{+}84, Rim81, RD80, RD84, Sei84, SS84b, SMD84, SR84a, SDH84, Smi84b, TUB83, TUB84, TCR84, TL84, TUWA84, Tep84, Ver84b, WC85, WMC84]. nucleation [KM84c]. nuclei [BF84, Cat89, EM84a, EM84b, Hen84, HR87, Hod84, HK81, HK83, HK84a, HK84b, Kac84, KM82, KM84b, Lan82, Lan84a, Nai86, NAS87b, PSL88, PIM80, PIM84b, PI84, RSL^{+}84, Szy81]. Nucleon [PSL88, DHD84, FRS83, FRS84, FKU^{+}84, HR86a, LR84a, LR84b, SDH84, Tak85a]. nucleon-nucleon [FRS83, FRS84]. nucleons [Hub84a, MZZ89]. nucleus [AGH^{+}88, Fun84a, Fun84b, GDH^{+}89, GE88, HR86b, HMS84, SM84a, SUZ89, TW89a, ZNO88]. nucleus-nucleus [AGH^{+}88, GDH^{+}89, SUZ89]. nuclide [KAR81, KAR84b]. nuclides [AGF87]. nucore [HA84a, HA82, HA84b, HA83]. NUCPAR [MV87]. nucrin [HR86b]. NUDENS [MVV83, MVV84]. Nullijn [De 84c]. Number [SM82, SM84f, SM85e, ABM89a, Aur84a, Aur84b, Bro84, Bro89, CG87, EA83, EA84, Fem84, Fog82, GDK89, KW84a, MRC85, MKRJ89, OY80, RFSG84b, Rob84e, SRS85]. numbers [Di 87, KS81, KS84b, Rib80, Sch84a]. Numerical [ABG^{+}81a, Arn84a, Bel84a, BMG^{+}85, CT84b, Col84b, Dec89, DS89, Dic82, Dor86, GJ84, Gae81, Gen82, Gri87, HH82, HC86, HKK84, IFI84, Jam87, KL84a, KP84, Len80, MLS81, MN88, MNR89, Moo84e, Net84, OSF86, PPHS86, Pic89, Pou88a, Rob84d, Rob84e, Rob84f, RG80, RG84, Saa89, SAVV87, SBM^{+}86, And86, Ano82w, CR84a, CR84c, CR87, Col88, CM81b, CH84d, CM84f, DG89, DGPM85, DM86b, DRRvB82, Fie80, Fla82, Fog82, Fog84, FS84d, GV87, GS88, Gib85, Gir81, Gro89b, GPS^{+}81, GRRP83, GRRP84, HKT85, IR80, Ixa80, IR85, Jak84, Kal82, Lab84d, Luc84, LU83, LU84, Man83b, McN81, McN84, Mic87, MMR89, New86, Ost85, Phi80, PT86, PB84b, RW84, Rap81b, RC85, RC87, Rob84b, SFB87, SSF^{+}89, SZ84, Tau80, Tau84c]. numerical [TTT87, TV86, Wai87, WH81, WH85, YP86, ZDN84, vTBF84, Art85a, Sch86a]. Numerov [BP82a, BP82b, BP84a, DW88, IR80]. Numerov-like [IR80].

O [Joh88, LNB^{+}88, LH88]. Obituary [Ano85w]. Object [Mar85, Ric86]. object-oriented [Ric86]. objects [BB84c]. observable [Bar83, Bar84e]. observables [CGGK85, MT84a]. observation [Ugn80]. observations [BBH^{+}88, He84a, KJRS88, KGAL88, MV88, O’C84a]. observed [Han84f]. obtain [Ben82, Ben84a, Han84b, Han84f, WH81]. obtained [TV86]. obtaining [HL86, Rib80]. Occam [HBM89]. Octahedron [NHK82]. odd [BF84, HK81, HK83, HK84b, Kac84, MR84b, HK84a]. odd- [Kac84]. odd-odd [HK81, HK83, HK84b, HK84a]. ODE [DM88b, PP82]. Odepainleve [RW86]. ODRIC [MBOK86]. off [CGM89, GMR84b]. off-diagonal [GMR84b]. off-line [CGM89]. offline [CHS87, Hoe89]. OLYDES [Car80]. Olympus [HRL84, HRR84, HR84e, Car80, HR83a, HR83b, HR83c, CR84d, HR84c, HR84d]. on-line [CGMS89, Kri89, Mas81, PF86, SN84, SdAD^{+}88, Sta87a, Van81a]. on-resonance [SDH84]. One [Has84, LB82b, YP86, ABG^{+}84, AK84, BL88, CR84c, Cha84e, CRPL81, CRL84, CRPL84, CAR84b, De 84c, DHD84, DP84, GJ84, Hib84b, HWL89, HLM^{+}87, Klo84c, Lew84d, LNS84, MMS87, MS84, MBOK86, Phi80, PGB84, RSMTV88, RC85, Rob84a, Sal84c, SPM^{+}88, Stu88, Wie85, WB84, Yur84]. one-and-a-quarter-dimensional [HWL89]. one-center [Yur84]. One-dimensional [Has84, LB82b, YP86, ABG^{+}84, AK84, BL88, CR84c, CRPL81, CRL84, CRPL84, CAR84b, GJ84, LNS84, MMS87, MBOK86, Phi80, RSMTV88, RC85, SPM^{+}88, Wie85]. one-electron [Hib84b, Lew84d, MS84, Sal84c, WB84]. one-loop [Stu88]. one-particle [Cha84e, Klo84c, PGB84, Rob84a]. Online [ABLB^{+}89, QAD^{+}89, WV89]. only [Hei84, O’C84a]. opacity [LLA^{+}89, PM83, PM84b]. opal [Ano87c, O’N89]. open [BDD84, CC89a, Ols83, Ols84, Ras80a, Ras84a]. operating [CHS87, MM88a, Per87]. operation [Ano81f, Ste84d]. Operations [Ito88]. operative [PS88]. Operator [Ito88, Bos83, Bos84, KMS88, Rap81a]. operators [Bru85, Bru86, CW80, CDW82, CW84a, Cha84e, CW84b, CDW84, FOW87, Lew84d, MPS84, PGB84, Rob84a]. optic [BCC^{+}89]. Optical [Lan87, All84b, Aue84, Coo84b, DTE84, ETH84, ET84, FRS83, FRS84, GE88, Has84, LZ84, PM80, PM84a, RHM^{+}89, TTW84a, TVD87, VPB82, VPB84, ZL84]. optically [KK84b, Tal82, Tal84a]. optimal [CC84, Ste84b]. optime [EK84]. Optimisation [LB82b, DD85b]. Optimising [Jes85]. Optimization [CPL89, Bas87c, HHK86]. optimized [Ano87b, BMR84, BA86, CRS87, Sch89c, ST84b, STN89, Tal89]. Optimizing [Dic82]. optimum [Gad85]. orbit [He84a, Klo84b, MV83, MV84b, PP85, Sax84c, Tro84]. orbital [FSP87, HE84b, HKS83, HKS84, Rob84b]. orbitals [AT84, Bat84, BZ84, HFG88, Lab84c, Lab84d, Mat84, MSFR89, ST84e, TBB87]. orbits [CM88, Hir84]. order [AM84b, AM84e, BB84b, BBM82, BBM84, Bar81a, BDK86b, BDK86c, BP84c, Bru85, Bru86, Cam81, Cam84c, Cam84a, Cam84b, CR84c, Cha84d, Cla82a, Col80a, Col80b, Col84a, DM88b, DHN84, DO84, FOW87, Lew84c, MBW89, Moo80, Moo84c, Mor84c, Nog83, Ros85, San82, San84, Sil84b, Sun88, TB85, TB87, Tho04a, Tho04b, Vaj82, Wai87, WS80, Wil84c]. ordered [CM88, KS81, KS84b, MZM^{+}82, MZM^{+}84]. Ordering [EA89]. Ordinary [RW86, Sch82b, Sch86d, Bra84a, Sch84c]. ordinates [Wie85]. organisation [Rea89]. organization [CHS87, Sil84a]. organized [WH89]. organizing [Ano84-86]. orientation [CLL89]. orientations [Pre84a]. oriented [CPL89, CSMHT88, DSM^{+}81, DSM^{+}84, Gal81, Mar85, Ric86, Sko89, Zlo84d]. origins [FH88b]. Orszag [Eas84b]. ORT1 [GJ84]. ORT1- [GJ84]. Orthogonal [DW89, Eas87a, Har89, Hel80, HFG88, Öpi87]. orthogonalisation [TBB87]. orthogonality [Hay89a]. Orthogonalization [Dav81, Dav84]. orthonormal [BG81, GJ84, RVD^{+}81]. orthonormalization [YP89]. ORTHOVEC [Eas87a]. ORTOCARTAN [KP81b]. oscillations [Mik84, OS86]. Oscillator [FS84d, Sar87, Dob84, FT84b, God84, Hib84c, SG84c]. OSF [WBV^{+}89]. other [FBS88, GST^{+}81, LCS84]. outer [BK84b, Eğe82, Ege84]. output [CDF^{+}89]. overlaid [Wie86b]. overlap [BGMM81, DF84, GRRP83, GRRP84]. overlapping [vM84c]. Overview [Flu89, Mou89, OJK89, You89, Gue89, TV89, vD87a]. Oxford [Bra86, Gri87, Nad86a, Ano89-27, Boc89].

P [BD81, Duc86, Jes87, Lis88, Sch86a, Nad86a, CM84a, MKRJ89]. P-Tau [CM84a]. P-X [CM84a]. P. [MN84a]. P3M3DP [EHL80, EHL84]. P3M3DP-The [EHL80, EHL84]. PAC [Cat89]. pack [Rys84]. Package [BH88, OJK89, RW86, Sch82b, Sch86d, AI88, AMB84, ABMT84, Ano87c, BKD83, BKD84, CRS87, Cam84f, Car84a, CM88, CH84b, CR84d, CSBB82, CSBB84, DNR83, DNR84, DM89a, GJ84, GMN^{+}80, GMN^{+}84, Gru84a, HB85, HRL84, HRR84, IA80, IA84, KK88, KS84c, KS84e, KS08, KS84g, KS84h, KM82, KM84b, Laa86, Mag84a, MGMZ84, MMK81, MMK84, MMTK88, NR86b, PDF82, PDF84, PT85, Sch84c, Sci89, SW80b, SW84, TTT84, WC84, Wie86a, Yur84, vdB84]. packages [FM85, SH82, SH84]. Packard [KKLS82]. packed [VVM84]. packet [Has84]. padé [Sta84]. Pages [Eas89, Ano89h, Art85a, Art85b, Bra86, Eas80, Eas82, Eas84a, Fin87, Lis88, Man86, Rob85, Sch86a, Taj86]. Painlevé [Hla86, RW86]. Pair [Ito85, Han84c, Han84e, Pol88]. pairs [Män89]. PAN [Whi89b]. PAN-DA [Whi89b]. Panel [WBV^{+}89]. paperback [Lis88]. papers [Ano84b, Ano84-87]. parabolic [PPHS86]. Parallel [AHKM86, BPR^{+}86, Bra86, CSC89, CKT89, DFOP89, Dit89, FR82, FBS88, Gal85, Hub89, McC89, MGM87, PCM85, Ruf85, ST89a, WH85, ABM89b, AW89, BES82, BMK89, Bro89, Cli82, DEW84, EA89, Fla82, FS85b, Gaj85, Gen82, GHO^{+}89, Gue89, Hoc82, HMTO84, HT84, IHS^{+}88, JM82, Key89, Kri89, LK89, LTK88, Mat87, May89, Meu89, MBW89, Nas89, Ost85, Per82, Sam85, SMR87, Sli89, Tem82, Vaj82, Ves89, WMH82, WH89, Eas82, Jes87]. Parallelism [CCD85, ACSB85, Bir85, Dob89, Jes82, Jes86, Nas89, PBKW85, Poh87a]. Parallelization [BH89]. Parallelizing [GLC89]. Parameter [Smi84f, Bas87c, BMS87, BS86b, DG84, GPS84, JR84, Klu84a, KL84a, Nob80, Nob84, Pia84, PS84b, Ran86, Ran87a, RWZ87, Smi84d]. parameter-dependent [Ran86, Ran87a, RWZ87]. parameter-free [Ran86, Ran87a, RWZ87]. parameters [ADV81, ADV84, Cam84e, FR84, Han84c, KF84, LR84b, PR84, PIM82, PIM84a, RD80, RD84, Ste84b, WDM84, WMC84]. Parametric [LW85, JSB82, Kla84, Les89]. Parametrization [Sal83a, Sal84d]. paraxial [MN80]. parentage [All84c, AM84a, Chi84b, Gra84a, Hub84b, MNJL84, SK86]. Paris [Fin87]. parity [MV87, SDH84]. parity-dependent [MV87]. parjet [Rit84b]. Part [Fer84, Ran86, Ran87b, Sch88b, Sch88c, VKM^{+}82, Zal85, FRS83, FRS84, FHMR84, Gje81, Hib84b, CC84, RWZ87]. Partial [IK85, Sch82b, AKS88a, AKS88b, CB88, CHMM84, DGL89, Dru83, Jes84, Mar89a, MN84a, MKS83, MKS84, Sch84c]. partial-wave [CB88, CHMM84]. partially [EK86, RSL^{+}84]. participants [Ano86-28, Ano87d, Ano88z, Ano89z, Ano89-27]. Particle [Ben88, Eas86, Ezh84, JN85, Sta89, Swi88, AGR88, AEH87, BMR85a, BKR88, Cha84e, CDN^{+}87, EHL80, EHL84, Gau84a, GTL^{+}86, GH87, GWTM83, Hir84, IA86, Jam82, Klo84c, KGAL88, Lew84c, LTK88, MSC^{+}88, Man87a, MT84a, Mor84e, Par84, PR84, Per89, PB84b, PGB84, Rob84a, SMM89, SAC^{+}86, Tri84, WHW87, Zac81, ZS87]. particle-in-cell [BKR88]. particle-mesh [EHL80, EHL84]. particle-particle [EHL80, EHL84, Man87a]. particle-particle/particle-mesh [EHL80, EHL84]. particle/particle [EHL80, EHL84]. Particles [LNB^{+}88, AB83, BPW84a, BPW84b, BPW84c, Bas87a, BT88, CN84, Cha84d, CJ84, GTG81, GTG84, GTM85, GDK89, Hei84, Jab84, KK84a, KF84, MKRJ89, Vrb89]. Particles-on-a-sphere [LNB^{+}88]. particular [WH81]. partition [BGG^{+}89, Car89]. partitioned [Com80a, Com80b, Com84, KC81a, KC84b]. partitioned-data-set [Com80a, Com84]. parton [BT89, FGK89]. parts [Jam84]. PASCAL [Ran87a, Ran86, RWZ87, Ran87b, MKS83, MKS84, Pyl85]. Past [Bra87b]. path [Bas87b, KPP^{+}89]. pations [MRCL84]. Patiwen [FB84]. pattern [AO88, Rob87a]. patterns [MGMZ84, Pre84a, Pre84b, RS84f, vM84h]. Pauli [ST84c, ST82a]. PAW [BBC^{+}87, BCVZ89, Joh89]. PAX [HMTO84, HT84]. PC [LBK88]. PDE [SSM85, ST82b, ST84f]. PDG [Gau84c]. PDP [DTE84]. PDP-15 [DTE84]. PDP11 [DCC84]. PDP11/34A [DCC84]. PEAD [TC87]. peak [vM84b]. peaks [CAP86b]. penetrability [Jam86b, Smi84e]. penetrating [Sko84]. penetration [SMMP86]. Penfold [ASM84, ASM80]. Penfold-Leiss [ASM84]. perform [DVD87, WB88]. Performance [BES82, DD85a, CKT89, CKV85, KPST89, KWPP89, MMM85, Mye89, Ros84a, Smi84a, SRS85, Sta87a]. performances [BC88b, Cha85]. performing [Fog84, MP81b, MP84b]. Periodic [JL84, Bun86, CM88, EHL80, EHL84, Kir87, KK83b, MP81a, MP84a]. periods [Bor82, Bor84]. peripheral [PB84b]. perm [SM85f]. permutations [BB84c]. Perot [MMF81]. Personal [And88, Qua87, Ast87, Bas88]. perspective [CFG^{+}87, Flu89, OT80, Ols83, OT84, Ols84, PDF82, PDF84].

PERTURB [FE88]. Perturbation [Cat89, VČ88, Gad80, MLS84, MLo84, MV81, MV84a, MS84, Mik80, MBW89, RS84e, III88, Sil84a, Sil84b, WS80, Wil84c, WP84]. Perturbative [Ixa80]. pervaded [Mik84]. PES [PIM80, PIM84b]. petroleum [BCC^{+}85]. Petviashvili [MLS81]. Peyret [Eas84a]. PFPOSFIT [Puf83, Puf84]. Phase [NM84, CM81b, CM84f, HSV84, Jad84, JA80, JA84, KK84a, KF84, KSE86, MS84, Mor84e, PB84b, Sea84a, Smi84e, Ste84c, Vrb89]. phase-space [CM81b, CM84f, Vrb89]. PHIGS [Mye88]. Phocha [DLB83, DLB84]. phoenical [EFES81]. phone [Str87]. phonon [Han84d, Nor82, Nor84b, Rei86, SR84b]. photoelectric [KM84a]. photoelectron [TC87]. photoemission [HPT80, PT84]. photographic [RSP81, RSP84]. photographs [BSP84, CA84]. photoionization [CLS^{+}84, KS86b, Sar87]. photon [BDK86a, BDK86b, BDK86c, BSdlT87, DO84, LW84a, SV84, TW83, TW84, WL84, WV80, WV84, ZNO88]. photon-Maxwellian [WL84]. photons [Chr84a, Chr84b, DLB83, DLB84]. photonuclear [ASM80, ASM84]. photoproduction [IW87]. PHOTUC [Sar87]. Phys [KS08, Ran87a, Tho04a, Tho04b]. physical [Bas87b, Zlá80]. physicist [Boc89]. physicists [Ano88-27, Jam86a]. Physics [Ano80g, Ano80f, Ano82i, Ano86f, Ano88h, Ano88i, Ano89h, CSMHT88, Gau81, Sch86a, Sta87b, Taj86, Ano81t, Ano84-87, AGG^{+}82, Arn84b, BBE^{+}81, BMR85c, BT81, Ben84b, BI85, BSA86, BBC^{+}87, BPR^{+}86, BE89, Bur87, Den88, Dun87, Eas86, Ezh84, Gau84b, Ger80, Gir81, GH87, Gre88, Gro89a, Hey89, Hin81, HS84, Jac81, JMS^{+}89, Jes86, KAH^{+}89, KP81a, Kaw86, Kos80, Lan87, LSW87, Mar89b, Mon87, Nad86b, Nas87a, Nas89, New89, Nex80b, OLR^{+}87, PBKW85, PFGP82, Poh87a, Poh87b, Put89, Qua87, Rim81, Rob87a, SN84, Sjö83, Sjö84a, Sjö86, SB87, Van89, Van81b, Ves89, Wat89, Wha89, Whi89a, Wil81, Wis87, Zac81, ZS87, vD87b, Ano80e, Ano88-27, Boc89, Wic87, Fin87, Lis88, Eas84b]. PIC [DH84b, Har88, HL88, LK88]. pickups [AN88]. pictorial [vD87a]. Picture [Col88, BO80]. pictures [BLMU81, BdCP81, VKM^{+}82]. Pinch [SBM^{+}86]. Pion [Lan84a, Fun84a, Fun84b, GE88, Lan82]. pion-nucleus [Fun84a, Fun84b, GE88]. pions [EM84a, EM84b, ET84]. PIPIT [ET84]. Pirk [EM84b, Fun84b]. place [Rob84d]. planar [AO88, BCCM87, ELS88, KBMV86]. Planck [ERB88, FSR84, MMK81, MMK84, MGK86, MMTK88, OCS86, SFB87, SSF^{+}89]. plane [BDG^{+}84, CM84a, DEW84, Fer89]. planes [CM84a]. planet [Moo80, Moo84c]. planewise [HM81, HM83, HM84a, HM84b]. planning [ASS82]. plant [KK85]. plants [KB86a]. Plaskem [Rob84e]. plasma [BV80, BV84, BSV^{+}85, CHL^{+}81, CD86, Eas86, FLM84, FBW82, FM84, FC84a, GH83, GH84a, GHU81, KBMV86, LHP87, LNS84, LJN85, Mae88, MTN^{+}82, MCA87, MV88, MGK86, New86, Ods80, OM88, Rob84d, Rob84e, Rob84f, SWHB88, Shu87, SPM^{+}88, SVP86, Swi88, TTAT81, TAK^{+}85b, TDS81, TDS84, Wen88, WW80, ZC84, Taj86]. plasma-circuit [LJN85]. plasma-wall [GHU81]. plasmas [BHM81, Ber80, Bor82, Bor84, Cap86a, CM81b, CM84f, DM86b, Dor86, EPS81, EP86, HP83, HP84, Hof88, IFI84, KLGT85, LCH^{+}81, Mac89a, MCJ81, MCJ84, OS86, SW80a, Tal82, Tal84a, WMCH84, Sta87b, Sch86a]. plastic [LK88]. PLATTSUM [HM83, HM81, HM84b, HM84a]. Plot [Pre84b, CA84]. plots [Dun87, Kal82]. plotter [Moo84f, Moo84g, RSP81, RSP84]. plotting [AMB84, ABMT84, AMK81, AMK84, PT85]. plus [HKT85, LT87, PIM80, PIM84b]. Point [Bur85b, EFK85, MB83, AS83, And83b, And83a, AS84, And84b, And84a, BAG^{+}87, Cla82a, FM84, Got83, KA87, LB84, May81, May84, MSFR89, MB84, Ric82, Ric84, SS89, SA83, SA84, TDS81, TDS84, Woo85]. points [AN87, AAMB84, AMB84, CC84, Tor82, Tor84, Wai87]. Poisson [AZK83, BH85a, CH84a, Cla82b, Hou87, Hug84]. Poisson-solver [CH84a]. polar [Laa86]. polarisation [Lan84b]. polarizabilities [Sho81, Sho84]. polarization [AMK81, AMK84, Ber84b, DGO83, Nur85, Sei84]. polarization-modulation [Ber84b]. polarizations [Hei84]. poletips [McN81, McN84]. pollution [Rap81a]. poloidal [BM84b]. poly [MD81]. poly-processor [MD81]. Polyatomic [EHH88]. polycrystal [SR84b]. Polyfit [Sci89]. polygamma [Di 86]. polygon [Woo85]. polygons [Lee82, Lee84]. Polyman [OL88]. polymer [Mic87]. Polymol [And84c]. Polynomial [MC85, Del84a, Sci89, Ten85]. polynomials [BG81, Bra84a, CWS83, CWS84, GJ84, Hel80, OL88, Öpi87, RVD^{+}81]. POLYP [MD81]. POLYRATE [ITR^{+}87]. polytropic [BM84b, MM82, Mik84, MM84b]. population [LHP87, SM84a, Tal84b]. portability [Mye89]. Portable [Car88, May89, BCVZ89, Day81, Day84b, Day84a, ERLD87, Tau81, Tau84b]. POS [GS89]. POSDIF [Câm82, Câm84g]. positions [FFH84]. positive [Bar81b, Bar84c, Gad85]. positive-definite [Gad85]. positron [ANP^{+}85, BDK86b, Câm82, Câm84g, JT84a, KEMP81, KE84b, KEMP84, KE84a, Odo81, Odo84c, Puf83, Puf84, Rit86, VDM84b, War84]. positronfit [War84, KE84b, KE84a, VDM84b]. positrons [DLB83, DLB84]. possible [BBD^{+}89, Lan87, Wil81]. post [KS86b]. POT4A [BH85a]. potential [All84b, BT84b, Cug84, CDN^{+}87, ET84, FRS83, FRS84, FT84a, FM84, Gia84, Has84, HH84, Hir84, Kla84, Klu84b, KP85, Lee82, Lee84, MLL86, MR80, MR84a, ML81, MZZ89, ÖW84, PP85, Ras80b, Ras84b, RRG88, SNC80, SNC84, TTW84a, Tel84, Ten84b, TDS81, TDS84, VPB82, VPB84, Art85b]. potentials [AAH84, CMR86, Coo82, Coo84a, Coo84b, Dom86, Fie80, Fie81, Fie84, FJ89, LW84b, PS84a, ST84b, ST82a, ST84c, STN89, Tal89, WGM83, WGM84, YP86]. powder [FHMR84, FFH84, FR84, MZM^{+}82, MZM^{+}84, MGMZ84, dBM82, dBM84, vM84h, vM84i]. powders [DG84]. Powell [Gri87]. Power [FK86b, Jam84, KK85]. powers [DM86a, Di 87, MR84b]. pp [Duc86]. Practical [Meu89, LM84]. practice [Ano88-27]. PRAVDA [MPS81]. pre [Bet84, HMS84]. pre-equilibrium [Bet84, HMS84]. pre-equilibrium/compound [HMS84]. precision [DSK83, DSK84b, Rys87]. precompiler [WC85]. Preconditioned [GLC89, OJK89, AKS88a, AKS88b, CKA89, EA89, KA87, Sun88]. preconditioners [CKT89]. preconditionings [KJ85]. predict [Lev89, Rob84e, RS84a, Smi84a]. prediction [Gib85]. predictor [BP82a, BP82b, BP84a]. predictor-corrector [BP82a, BP82b]. predissociation [Le 89]. Preface [Ano82z, Ano82-27, Ano84-76, Ano84-77, Ano85x, Ano85y, Ano86-30, Ano86-29, Ano88-28, Ano88-29, Ano88-30, Ano89-28, BM88, DD89, FPR89, Gru81, Gru84c, Jam81, KS87b, Nad80, Nüh86, TSH87]. Preliminary [Ano80s, OLR^{+}87]. preorthonormalization [HT88]. prepared [GWL89]. Preparing [Frü81]. Preprocessor [KK89, BH85b, HRR84, KPP81, KPP84b]. preprocessors [Joh87a]. presence [ACR88, ABG^{+}81a, OSF86, SSF^{+}89]. present [Bra87b]. presentation [BCVZ89]. preserving [CM88, ST84d, SBFI87]. Press [Gri87, Nad86a]. pressure [ZC84]. Preuss [Yur84]. price [Ano89h]. primary [KK85, MM82, MM84b]. prime [DW84]. primitive [Sch88a]. primitives [AB88, SMC81, SMC84]. principle [Pon81, Wil84b]. Principles [Cio89, CRH^{+}89, Man87b, RW84]. Printer [Moo84g, Moo84f]. Printer-plotter [Moo84g, Moo84f]. priory [Chi89]. prism [Hou87]. probabilities [BF84, HK83, HK84b, vdB84]. probability [BD87, De 88, DP84, GTM87, Mal84]. probe [Cat89]. problem [AZK83, BH81, BH84a, Bas87b, BMK89, CR84a, Cam84f, Com82, Das83, Gae81, GR81, GV87, Gir81, KBMV86, Man87a, SSF^{+}89, Vaj82, WH81, YP86]. Problems [GK80, Har84b, KW84a, AP80, ABG^{+}81a, AHVV86, BO80, BV80, BV84, Ben88, BG81, BTS84b, Cha85, Cis87, CDN^{+}87, DGL89, Fie80, For89, FJ89, GJ84, Gal85, Gru84a, HT88, HKK84, Jam82, KJ85, KS84g, KS84h, LW84b, Mül87, Nav89, OTB80, OTB84, Rea89, Rob87a, Saa89, Str81, vdV89]. Procedure [CKW89, Bil89, BKK86, BW87, DG89, Di 86, DU89, GMR84b, HT88, Jak84, LM84, TT82, TT84]. Procedures [FK86b, ST89b, KP89]. Proceedings [Bra86]. process [Eck84, FS82, JW85, SK82, SK84]. processes [BS87, BKJ83, BKJ84, BDK86a, BDK86b, BDK86c, BSdlT87, BBC^{+}84, BBL^{+}84, GHU81, HRS81, HRS84, IP88, Le 88, LVPG84, MC84a, ST82a, ST84c, Swi88, vdB84]. Processing [FR81, HT84, MD82, Nex84, AWB82, AW89, BO80, BC88b, CSMHT88, DD85a, DFOP89, Dit89, HRW85, HSA82, HSA84, IA86, KP81a, Mas81, MTN^{+}82, MBW89, OY80, Rim81, Ruf85, Sch89c, SM85a, Tol85, TGA88, Ver84b, VSH83, VH87, WH85, Zlo82, Zlo84a, Jes87]. Processor [Par83, ABM89b, Ano87b, CNRS82, DGK^{+}81, Gaj85, Gje81, HD86, Hou87, JSB82, Jac81, Jac82, Jes85, KK88, KSG82, LW84a, LTK88, MD81, Män89, MHP84, PP82, Rap85, Sch82a, SM85a, VM89, Ves89, VGD^{+}89, WMCH84, Bur85b, Gal85]. processor-emulator [VGD^{+}89]. Processors [Bra86, Bai87, BES82, Bro89, Dit89, Fra81, Gro87, Gro89a, Gue89, Hoe86b, Kri89, NFH^{+}88, Per82, Sam85, UI85]. Produced [KK89, DLB83, DLB84]. Producer [Flu89]. Producing [JN87]. product [CW80, CDW82, CW84a, CW84b, CDW84, MG84, Mar89c]. production [BDK86a, BDK86b, FGK89, Gib85, MT84a, Odo84a, Odo84d]. products [DW84, Eğe82, Ege84, Hay83, Klo84c, Sax84b]. profile [Tal82, Tal84a, Tal82, Tal84a]. profiles [DG84, OBLR84, SLH84, YB86]. Profiling [Sch85b]. progeny [Har88]. Program [Ano80t, Ano80u, Ano81u, Ano81y, Ano81v, Ano81w, Ano81x, Ano82-28, Ano82-30, Ano82-29, Ano83p, Ano83q, Ano83r, Ano84-78, Ano84-79, Ano84-80, Ano84-81, Ano84-82, Ano84-83, Ano84-84, Ano84-85, Ano85z, Ano85-27, Ano85-28, Ano86-31, Ano86-32, Ano86-33, Ano86-34, Ano87x, Ano87y, Ano87z, Ano87-27, Ano88-31, Ano88-32, Ano88-33, Ano88-34, Ano89-29, Ano89-30, Ano89-31, ABD82, ABD84, Bar83, Bar84e, Bas87d, Bet84, BTS84a, BTS84b, Cha82, Cha84c, CgYpXh85, DR85, DS83, DS84a, Eas87a, EFK85, GLST89, Har89, Hof84b, Hof84a, Ito85, IK85, Ito88, KM84a, KEMP81, KEMP84, KM82, KM84b, KR88, Lis84, Mar89c, MT84b, NA84, PP85, ST84e, Ste84b, Wil84b, WP84, YpMzCeCg85, YFL87, AV84, AI88, ABH^{+}84, All84a, And84c, ANP^{+}85, Ano86n, ABG^{+}84, AGF87, AMK81, AMK84, AK84]. program [Aur84b, BV80, BH81, BV84, BH84a, BKvLSE80, BKvLSE84, BBM82, BBM84, BD80a, BDM81, BRTB84, BRT84, BDM84, BD84a, BSK88, BCFK85, Bar84d, BB86, Bas87a, BSP84, Bat84, BP82b, BP84a, BH85a, BGMP84, Ben82, Ben84a, BD82, BD84b, BPA84, BBC^{+}84, BBL^{+}84, BD84d, BSA86, BSDM88, Bie83, Bie84, BP84c, BEM84, BM85a, BS86a, BA86, BJ81a, BJ81b, BJ84a, BJ84b, Bur84, BT84b, But84, Cam84f, Câm82, Câm84g, Car89, Car84a, CFG^{+}88, Cha84d, CW80, CW84a, Cha84e, CW84b, CCR84a, CCR84b, Chi84a, Chi84b, CH84a, CSC^{+}86, CC84, CF85, CLS^{+}84, Coo82, Coo84a, Coo84b, CGB^{+}89, Cre80a, Cre80b, Cre84a, Cre84c, CSW84, CMO86, CKV85, DK87a, DK87b, DM89a, DM86a, De 84c, DFD81, DFD84, DLB83, DLB84, DCC84, DL86, DSK83, DSK84b, DHN84, Des84b]. program [Deu84a, Deu84c, DRD88, DH84a, DS84b, Duf84, DO84, Dya86, DGJ^{+}89, EHL80, EHL84, EM84a, EM84b, EA83, EA84, Eks84, ENSC86, EHH88, FV86, FT84a, FS82, Fem84, FHMR84, FFH84, FK84, Fer84, FR84, Fie81, Fie84, FM85, Fle81, FS84b, Fog84, FS80, FKM^{+}87, FE88, Fro84a, Fro84b, Fro84c, Fro87, Gal84, GDH^{+}89, GTG81, GTG84, GTM85, GTM87, Gat89, Gel85, GH83, GH84a, GSZ85, Gia84, GE88, Gio84, Goo84, Gra84b, Gra84c, HC83, HC84, Han84a, Han84b, Han84c, Han84e, Han84f, Han84d, HHK86, HBJ84, HRS81, HRS84, HMS84, HM81, HM83, HM84a, HM84b, Hib84b, Hib84a, Hib84c, Hib84d, Hib84e, HRB84, HK81, HK83, HK84a, HK84b, Hub88, HKS83, HKS84, IA80, IA84, IL84, ITR^{+}87, Isl81, Isl84a, Ito86, Jab84]. program [Jac80, Jac85b, Jac85a, JT83, JT84a, JT84b, JA80, JA84, Jam84, JP80, JP84b, Kac84, Kam86, KF84, KKP82, KPH84, KKP84, KS81, KS84b, KE84b, KE84a, KS84c, Kla84, Klo82, Klo84a, Klo84b, KH89, Köl81b, Köl84b, KS84e, KS84f, KS08, KS84g, KS84h, KAR81, KAR84b, KPST89, KP81b, KPP81, KPP84a, KPP84b, KPPR84, Lad86, Le 89, LCW84, LZ84, LHP87, LS84, MS87a, MRCL84, MLL86, MR80, MR84a, MG84, MRW84, MW84a, MZM^{+}82, MZM^{+}84, Mar84b, MW84b, MS84, MGN80, MGN84, MP81b, MP84b, MKS83, MKS84, Mor84a, MC81, MC84b, Mor84c, MT84a, Mor80, Mor84d, Mor84e, Nai84, Nai86, NN84, NT84, Nov87, O’C84b, O’C84c, Odo81, Odo82, Odo84a, Odo84b, Odo84c, Odo84d, Ods84, OT80, Ols83, OT84, Ols84, OTB80, OTB84]. program [Öpi87, OCC80, OCC84, PB84a, PNO89, PR84, PV84, Pis84a, PS84c, Pre84b, Pro82b, Pro84, Puf83, Puf84, PGB84, Ram85, Ran86, Ran87a, RWZ87, RRC89, Ras80b, Ras84b, RA82, RA84, Rob84a, Rob84b, Rob84d, Rob84e, Rob84f, RS84a, RBHW84, RSL^{+}84, RS80, RS81, RS84b, RS84c, Ryn82, Ryn84, RD80, RF83, RD84, RF84, SMMP86, Sal82, Sal84e, San82, San84, Sar84c, Sax84b, Sax84a, SGS^{+}89, Sch88a, SNC80, SNC84, Sch89c, Sch86c, ST82a, ST84c, Sea84a, Sei84, ST84d, STN89, SR84a, SDH84, III88, SS84c, SS84d, Smi84d, SK82, SK84, SVP86, Ste84c, Szé85, Tak85a, Tal89, TCS85, Tau80, Tau81, Tau84b, Tau84c, Tau84d, TT85, Ten83, Ten84a, TN84, TM89, TKS^{+}85, TGA88, Tro84, TTL88, Ver83, Ver84a, VVM84, Ver84b, VPB82, VPB84, Vio84, VDM84b]. program [VBD88, WGM83, WGM84, WB84, YP88, Yat84, YMWW89, ZL84, Zlo81, Zlo82, Zlo84a, Zlo84b, Zlo84c, Zoh84, Zwa85, dB82, dB84, tR85, vMF81, vM84a, vM84f, vMF84, vM84g, vMT84, vM84e, vM84h, vM84i, MW84c, Wil84a]. program-package [AI88, IA80, IA84]. programmable [Män89, PCL89]. programme [VH87, Wen88]. Programming [Dow82a, Jes82, KSS86, LT85, Rus87, BKD83, BKD84, BM85a, CRV^{+}89, CPL89, Eas82, Mar85, May89, Nag85, Sch86b, Sch85b, Tar86, WHD86, Whi89a, vdB84, Rob85]. Programs [AF82, AF84c, Fin80, Fin82, Fin84a, Fin84b, FC84b, Köl84c, LO86, Sch84a, SG83, SG84b, WS82, WS84a, AD84, BF89, Bir85, BvEF85, BS86a, Bri86, Car80, CDW82, CDW84, CM88, CR84d, Com80b, Das89, DVD87, DK87c, Gro89a, Gro89b, Nis88, Ran86, Ran87a, RWZ87, Ran87b, RR84, RV84, SCF86, SS84b, SMD84, VSH83, WC84, WC85]. Progress [CGB^{+}89, CFG^{+}87, HL88, Met89, Ste87]. Progressive [Bil89]. PROION [Bor82, Bor84]. Project [Pla86, Maz89, Ste87, BBC^{+}89]. project-progress [Ste87]. Projected [Fle81]. projectile [Coo84b]. projecting [Tak85a]. projection [ABH^{+}84, GM89]. projections [CA84, Pre84b]. projects [Man87b]. Prolate [BC83a, BC83b, BC84a, BC84b]. prompt [Puf83, Puf84, War84]. Propagation [Hoe86b, AHVV86, BBM82, BBM84, Dru83, DH84b, Fer89, JBWW86, MT81, MMF81, MN84a, MCJ81, MCJ84, Mor84c, PPHS86]. propagator [DGO83]. Properties [Pol88, BSdlT87, Bri86, DK87c, Mac89a, Mic87, Nog83, Tho84, Tri84, VH87]. property [Eck84]. proposals [OLR^{+}87]. proposed [But89]. proteinic [CF85]. proteins [SCF86]. Proton [BMG^{+}85, Fun84a, Fun84b, Keg85, MT84b]. proton-irradiated [Keg85]. protons [BBH^{+}84, Mal84]. Prova [MZM^{+}84, MZM^{+}82]. provided [THH82]. pseudo [ABM89a, Rib80, SRS85]. pseudo-random [ABM89a, SRS85]. pseudofermion [CMPR88]. pseudofermions [FS85a]. Pseudopotential [Kol81a, Kol84a]. public [IH86]. publication [Blo89]. publisher [Ano88i, Ano86-35]. Pulsamp [RS84a]. pulse [AP89, AK84, DH84b]. pulse-height [AK84]. pulsed [vM84g]. pulses [RS84a]. punching [Fem84]. pure [AMC83, AMC84, BF84, CM85, DMM83, DMM84b]. purpose [BCVZ89, DS84b, DGJ^{+}89, Gup88a, Poe85, Pro82b, Pro84, RWK^{+}85, RS80, RS84c, vM84a, vM84b]. pushing [MSC^{+}88]. PWBA [TW83, TW84]. PYTHIA [BS87].

QCD [BS86b, CMR86, CMPR88, Ing87, IHS^{+}88, Odo81, Odo84a, Odo84c, Odo84d, Rit84b, Sch87a, Tou87, Tou89]. QCDPAX [IHS^{+}88]. QED [JW85, KKS87, LO86, MGN80, MGN84]. QR [BMK89]. quadratic [Zlo89]. quadrature [Lew82, Lew84a, ST84b, Tak88a, Tak88b, Tel84]. quadrupole [CDN^{+}87, DG84, DSSS86, God84, LJLC84, Nai86, SS84a, vM84h]. quadrupole-distorted [vM84h]. qualitative [Hla80, Hla84]. quality [DP89, Sch88b, Sch88c]. quantification [Bro83]. quantities [Bar83, Bar84e]. quantization [FE88, MD88]. Quantum [CDL88, Duc86, FH88b, MR86, RSA84, SG82, ST89b, BMV83, BMV84, BTS84b, CCD85, DWC^{+}89, GLST89, LL84, MMS87, MR87, NFH^{+}88, NBI^{+}85, OTB80, OTB84, Sch87b]. quantum-chemistry [BMV83, BMV84]. quark [Odo82, Odo84b, Rit84a]. quarks [CMPR88]. quarter [HWL89]. Quasi [DeV84, Gad85, RM89, Lan80, MN84b, MGK86, Tal84b]. quasi-analytical [MN84b]. Quasi-bound [DeV84]. Quasi-elastic [RM89]. quasi-linear [MGK86]. quasi-Newton [Lan80]. Quasi-optimum [Gad85]. quasi-steady [Tal84b]. quasicontinuum [Lea87]. quasilinear [SSF^{+}89]. quasiparticle [HR87, HH84]. QUB [BES^{+}87]. Query [Sko89]. questions [Gir81].

R [Eas82, Eas84a, BBL^{+}84]. R-matrix [BBL^{+}84]. Racah [AD84, BK84b, SMR87]. Radial [Arn84b, SF84a, Bel84a, BC83b, BC84b, FS84d, Jam84, Lab84c, Mar84a, Pie84b, SS84a, Sea84b, Ste84c, Tol86, VPB82, VPB84]. Radiation [FM84, BS84c, CFG^{+}88, Cha84b, ELS88, Hub84d, KK84b, Lan84b, LO86, RSMTV88, Rei89, TW89a]. Radiative [BDK86a, TS89, BKJ83, BKJ84, FS84b, HKK84, IL84, JW85, Mac89a, MPM85, SF84a, SL83, SL84, WBG84]. RADICAL [Ran87a, Ran86, Ran86]. radii [Fun84a, Fun84b]. Radio [BS84d, BS84c]. radioactive [Aur84a, Aur84b, EA83, EA84]. radioastronomical [Sal84b]. RAL [Wha89]. Ralf [Man86]. RAM [Ska82]. Raman [LBY86, MW84a, MW84b, MW84c, Wil84a, YB86]. Ramses [DH84b]. Random [Bro89, ABM89a, Bro84, Cha84b, CG87, CGG87, De 88, JA80, JA84, KW84a, KCA84, Rib80, Sal87, Sch87b, Sch84a, SRS85, WDM84]. random-field [CGG87]. Random-number [Bro89, Bro84]. randomly [DSM^{+}81, DSM^{+}84, Gal81, WD84]. RANF [KW84a]. range [BP84c, CCR88, Com82, FKU^{+}84, FS85b, Hub84f, LR84b, Mas83, NK84, PV84, Ste84c, TL84, TUWA84]. Ranged [Fin84b, Fin80]. ranges [MRW84]. Rank [IK85, Ito86]. Rapid [Mue84, GMR84b]. Rapidity [Jad84]. Rappaz [Man86]. Raptis [Moh80, Moh84]. rare [Câm82, Câm84g]. Rate [Bil84, Bil87, Bie83, Bie84, ONP89, Rob84f, WMCH84]. rates [Har81, Har84a, ITR^{+}87, Le 89, MM88b, SG89, TSD84]. rational [GMR84b, Gro89b, SW80b, SW84]. rationals [DSK83, DSK84b]. Raubenheimer [Sim84, Sim80]. ray [AWG84, Deu84a, Deu84b, Deu84c, Deu84d, DPH82, DPH84, FHMR84, FFH84, FK84, Fer84, FR84, FC84b, Hub84c, KBMV86, Lea87, MZM^{+}82, MZM^{+}84, MGMZ84, MCA87, PJ89, Rei85, Rei86, RM89, VSP86, WML^{+}82, Zlo82, Zlo84a, vM84i]. ray-tracing [KBMV86, PJ89]. Rayleigh [SSS84, SS84c]. Raynal [YFL87]. rays [GTM87, PDNE88, RSL^{+}84, SS84c, SS84d]. RCWFF [Bar84d]. RCWFN [Bar84d]. Rdlist [Day84a, Day81]. reacs [Rob84d]. reactance [BTS84b, BW87, Hen84, OTB80, OTB84, Sar84a, Sar84b, Sar84c]. reacting [Key89]. reaction [Bas87a, BvSW84, DKA85, GG89, Har81, Har84a, Hod84, ITR^{+}87, KPP^{+}89, MM88b, RS81, RS84b, SG89, WMCH84, vdM84a]. reactions [AMK81, AMK84, Ass84, Bas87a, Bet84, But84, DL87, DP88, DKA85, DD86, EE89, FKU^{+}84, HIM84, Hei84, HMS84, MT84a, ONP89, OVQT^{+}84, Rob84d, Rob84f, SM84a, SAC87, Sei84, SS84b, SMD84, SR84a, SDH84, Tak85a, TUB83, TUB84, TCR84, TL84, TUWA84, WMC84]. reactive [DWC^{+}89, HH82]. reactivity [Kam86]. reactor [BSA86, Das83, Gad80, HLM^{+}87, LCS84, Mar89b, Moh89]. ready [Wil89]. Real [FRS83, FRS84, SM85a, Sne85, AM84e, Bar81a, Bar81b, Bar82, Bar84b, BFSG84, Bar84c, Bar84d, Cam81, Cam84c, Cam84a, Cam84b, CFG^{+}87, Cli82, CRL^{+}89, EP86, FMS87, GSS81, Gli82, Gór88, Ken89b, Sch86b, Sch88b, Sch88c, Tau80, Tau84c, TB87, Tho04a, Zal88]. real-time [FMS87, GSS81, Gli82, Sch86b, Sch88b, Sch88c]. realistic [FRS83, FRS84]. realization [CDL88]. realize [PT85]. reciprocal [HM84c]. recognition [AO88, Bil89, Rob87a]. recoil [PV84]. recombination [BS84d, FS84b, LH88, Mag84a]. Reconnection [KJRS88, Son88, Uga88]. reconstruction [DFOP89, GDH^{+}89, KPST89, Per89, Sta89, TA89]. recoupling [BSK88, Bur84]. rectangular [Hou87, Lew82, Lew84a]. Recurrence [MP81a, MP84a, Tol86, Arn84a, Lee89]. recursion [Hay83, HW84a, Hay89a, Nex84, PT82]. Recursive [Hay89b, PB84b, SG84a, HDK84, Hay80]. reduce [Sch84c, BO80, BH88, DR85, Eas87a, EFK85, GK80, Har89, Ito85, IK85, Ito88, KK89, Sch82b, Sch86d, ST89b, FK86b, Ito86]. REDUCE-2 [BO80]. Reduced [Bra84e, FS84c, KKLP84, Klo84c, BC84c, BF84, Cha84e, MKS83, MKS84, Rob84a, Sax84b, Sax84a, Bra84d]. Reduction [FSP87, Joh85, Stu88, vMT84]. reference [Lid84b]. refinement [LM84]. reflection [ACR88, FK84, Fer84, KN84a, MR80, MR84a, PI84]. reflections [FHMR84, FFH84]. reflective [PJ89]. reflector [Moh89]. refraction [ACR88]. refractive [BLB88]. refresh [BGMM81]. regarding [Gup88a]. regeneration [Rob83]. Regge [BT84b, HIM84]. reggeon [BB84b]. region [BKJ83, BKJ84, vdM84a]. regions [BS84d, Lew82, Lew84a, Moh89, RG86]. register [CG87]. registered [OVQT^{+}84]. regression [DP84]. regular [Bar81a, Nes84]. regularity [BRB89]. regularization [Pro82b, Pro82a, Pro84, tR85]. regularizations [RS80, RS84c]. Reid [Bra86]. Reidel [Art85a, Rob85]. related [BLMU81, BKvLSE80, BKvLSE84, CR84a, vdV89]. relational [LSW87, Sko89]. Relations [Mur88, Arn84a, Han84d, Lee89, Tol86, You87]. Relationship [Sch87b]. relationships [PIM82, PIM84a]. relative [BPW84b, BPW84c]. Relativistic [BZ84, Bra84c, Bra84b, LCW84, BS83, BS84a, DLB83, DLB84, Des84b, DGJ^{+}89, FKM^{+}87, GE88, Gra84b, Gra84c, Lab84a, LW84a, LZ84, Man83a, MSFR89, Mor84e, PP85, SSS84, STN89, SFB87, SS84d, WB84, Yat84]. relativity [SG83, SG84b]. Relaxation [CLL89, Got83, AZK83, vMT84]. release [IH86]. relevant [Str81]. reliable [Han84c, Han84e]. Remark [SM85f]. Remarks [GS88, LB82b, Gen88, Goe84]. remote [Ghi87]. renaming [Nas87a]. rendering [vD87a]. renormalisation [vTBF84]. renormalised [SP87]. renormalized [DW88]. Renyi [Cha84a]. reorientation [BL86]. Reply [BD81, Mag81]. Report [Kil80]. Representation [FH88a, BS86b, Cho84, Hel80, Hub84b, MG84, SM81b, SM84e, SM85b]. Representations [Dom86, DLPL89, BWBL88, KS87a, MPS84, Net84, PZCC89, SM82, SM84f, SM85d, SM85e, Wor84b]. represented [Pro82a]. Repulsion [EBS88, Ben82, Ben84a, BDD84, HC83, HC84]. required [ST84e]. requirements [Cio89]. rescue [MPS81]. Research [FKM^{+}87, Kil80, DL89, LP89]. Resfit [BB86]. residual [Oya86]. residues [DKA85]. Resistive [LCH^{+}81, Ede84, IK81, HR81, KLGT85, Kir87, MBOK86, MPIM86, TAK^{+}85b, ZC84]. resistivity [Isl84b]. Resolution [SMM89, Ver82, BR85, BPSB88]. resolvents [Nex89]. RESON [TN84]. resonance [BB86, DG84, KM86, SAVV87, Sea84a, SDH84, TTL88]. resonances [HR84a, Hla86, TN84]. Resonant [KP85, KK84b, VPB82, VPB84]. resources [ASS82, McI84]. respect [JP84a]. respectively [Ano89h]. response [GM89, LZ84, LLC84, MPM83, MPM84, ZL84]. restart [HR84e]. restricted [Jam87]. restructing [CKV85]. results [ABMT84, BBH^{+}88, Eld84]. Retrans [Kam86]. retrieval [DSK84a, HM84d, Jac85a, Vis84b, BJ81b, BJ84b]. retrieve [Jac85a]. retrieving [Sko89]. Revai [YFL87]. Reversed [SBM^{+}86, HWL89]. Reversed-Field [SBM^{+}86]. Review [And86, Ano89h, Art85a, Art85b, Bra86, Duc86, Eas80, Eas82, Eas84b, Eas84a, Eas89, Fin87, Gri87, Jes87, Lis88, Man86, Nad86a, Rob85, Sch86a, Sta87b, Taj86, Van81b, MPIM86, Tri84]. review-of-particle-properties [Tri84]. Revised [Gra84c]. revision [Ano80g, Ano82i, Ano86f]. revolution [DM89a]. rezoning [MN80]. rf [OCS86, AHS^{+}86]. Riccati [RW84]. rich [Män89]. Richardson [GB87, Jam88]. Riemann [BD80a, BD84a]. Riemannian [DR85]. Rigid [AF82, AF84c, Fin82, Fin84a, AB86, KN84c]. ring [Män89, Sam85, TS84b, Wil84c]. RISC [Hey88]. RITSSCHIL [Zwa85]. RKR [Tel84]. ro [CRH^{+}89, Ten83, Ten84b, Ten84a, Ten86, TM89]. ro-vibrational [CRH^{+}89, Ten83, Ten84b, Ten84a, Ten86, TM89]. Robert [Art85b]. role [Ebe84, Hey89, LSW87, Lid84a, Zac81]. ROMPIN [GE88]. Root [SW80b, SW84, DSK83, DSK84b]. Root-rational-fraction [SW80b, SW84]. Roothaan [HKS83, HKS84]. rotating [BV80, BV84, Sch88a]. Rotation [CLS88, CRL^{+}89, BC84c, Bru85, Bru86, RRG88, SMT88]. Rotation-vibration [CLS88]. rotational [BB84a, Bou87, CRH^{+}89, CCS89, DK87b, EHH88, HH84, Hir84, JT83, JT84b, LNB^{+}88, Nai84, Nai86, Ogi83, Ogi84, BR85]. rotational-vibrational [LNB^{+}88]. rotationally [JT84a]. rotations [CC89a, Tau84d]. rotator [BCCM87]. rotlev [Ten86]. rotor [CDL88, Jam87]. ROTTRA [Tau84d]. Round [Ano88-35]. Round-table [Ano88-35]. Rounding [Hay83]. Routh [GV87]. Routh-Hurwltz [GV87]. routine [Bas87c, CKA89, Gru84d, Hou87, HR84b, Isl84b, KA87, MM88b, Moo84f, Moo84g, NP84, Ste84a, WS84b, Zlo84d, vM84e, vM84b, vM84c]. Routines [Wai87, BN81, Vis84b]. Rovibrational [Hen84]. RPA [ENSC86, Jes85]. RRKM [Ryn82, Ryn84]. RSPT [Nog83]. rule [AO88, EC84, Le 89]. rules [Lee89]. rumpled [MS88]. run [BTS84a, DTE84, NA84, VH87]. running [JMS^{+}89]. rupp [Rin84]. RWSYST [KS86a]. Rydberg [RA82, RA84].

S [Bra86, Eas84b, JB80, MZZ89, Cug84, SM82, SM84f]. S-state [Cug84]. SA [Ano87-29]. SA/SD [Ano87-29]. Saclay [Gou88]. safety [Sch88b, Sch88c]. SAMP080 [KAR84b]. samples [DSM^{+}81, DSM^{+}84, Gal81]. sampling [KVW84, Sal87]. SAMPO80 [KAR81]. Sanibel [Ano82a]. SASD [Bor89, Gat89]. SASD-tools [Gat89]. SATDSK [McN81, McN84]. Satellite [Hin81, WB82]. satellites [Gli82]. saturated [McN81, McN84]. sawtooth [OS86]. Saxon [Cug84, CDN^{+}87, DHD84, HH84, Hir84]. Saxon-Wood [Cug84]. Scalar [EK86, Klu84b, KS84g, Mar89c, Stu88, vM84f]. scale [BLM^{+}88, BBC^{+}89, Car80, CH84c, Duc80, Dun87, GWTM83, Kas82, Man87b, MR87, Pat82, SN84, Sch87a, vM84d]. scales [MRC85]. scaling [Gad85]. scan [MPS81]. scatter [Dun87]. scatter-plots [Dun87]. scattered [KKP82, KPH84, KKP84]. scattering [AT84, AP82, AP84, BB84b, BS83, Bar83, BS84a, Bar84e, Bas87b, BTS84b, BT84b, Cha84d, Chi84a, CLS^{+}84, Cop81, Cop84a, Cop84d, Cop84c, Cop84b, CVvWF86, Deu84a, Deu84b, Deu84c, Deu84d, DWC^{+}89, EM84a, EM84b, Fie80, Fun84a, Fun84b, GE88, Gje81, Gro84, Han84d, Ing87, JT84a, KF84, Keg81, Keg84, KKT^{+}88, Lan82, Lan84a, MS84, Moo84h, Nor84a, OM83, OM84, OTB80, OTB84, PSL88, PP85, Ras80a, Ras84a, Rei85, Rei86, RM89, SP87, Sal82, Sal84e, Sin84, Sko84, SS84c, SS84d, Smi84b, Smi84d, ST84e, SR84b, Ste84c, ST82b, ST84f, TTW84a, TW89a, TS89, VM89, WML^{+}82]. sced [BLT81]. SCF [Cio89, KKP82, KPH84, KKP84, WB84]. SCF-X [KPH84, KKP84]. scheme [Hla80, Hla84, IR80, ST84b, Wil84b]. schemes [Dya86, FC84b, GV87, TL89, WL84]. Schmidt [O’C84b]. Schonauer [And86, Tru88]. School [Ano88-27, Beh84]. SCHOONSCHIP [Wro82, Wro84, GLST89]. Schrieffer [MVV83, MVV84]. Schrödinger [Bel84a, BVL84, BL88, CR84c, DRRvB82, FV86, FGSH85, Fog84, GS83, Has84, Hay80, HLS84, IR80, Ixa80, IR85, IB87, KP85, LM85, LR84a, Mar84a, MZZ89, PR87, Rap81b, Rap83, RC85, RC87, TV86, VPB82, VPB84]. Schur [Eğe82, Ege84]. Schwinger [BCCM87]. science [Her84a, Her89, Kas82, Wen88, Bra86]. science-design [Kas82]. Scientific [Tru88, BHSS84, BGM87, Gue89, HGJ^{+}89, Hut89, JHRR87, Nag89, PD89, Pat82, Pla86, PT85, Rea89, YSHK85]. scientist [Boc89]. scintillation [AGF87]. scintillators [GTG81, GTG84, GTM85, GTM87, GTMA88]. score [Ott89]. screening [DH84a]. SD [Ano87-29]. Search [O’C84c, AB88, Bas87c, Coo84b, Man83a, Mir84, Smi84d, Smi84f]. Second [Duc86, Sil84b, AI88, Aur84b, BBM82, BBM84, Cha84d, Cla82a, DO84, Kal83, Lew84c, MBW89, Mor84c]. Second-order [Sil84b, BBM82, BBM84, Cha84d, Cla82a, Lew84c, MBW89, Mor84c]. section [AMK81, AMK84, Bet84, BW87, Cap86a, FGK89, KS86b, LW84a, ONP89, SR84a, WMCH84]. sections [All84b, BKvLSE80, BKvLSE84, BGS81, BGS83, BGS84c, BGS84a, BGS84b, Bie83, Bie84, Bil84, BTS84a, BTS84b, CLS^{+}84, DL87, FNDL89, Har81, Har84a, Hei84, Hen84, HMS84, JT84a, LJLC84, MMM84, MPR80, MT84b, Nov87, OM83, OM84, OTB80, OTB84, PDNE88, PV84, PS84b, RS81, RS84b, Sal82, Sal84e, Sar84b, Sar84c, Sar87, SAC87, SS84b, SMD84, SG89, Ver89, WL84, vdB84]. segregation [BH81, BH84a]. Seismic [KSG82, DS82]. select [Ten86]. selected [TB87]. selection [Whi89a, Zal85, HM84c]. Self [CCR84b, AI88, BLT81, Des84a, IA80, IA84, LCW84, MC81, MC84b, PIM82, PIM84a, SSS84, WB88, WD84, ZNF87]. Self-consistent [CCR84b, BLT81, LCW84, SSS84, WB88]. self-contained [AI88, IA80, IA84]. self-diffusion [WD84]. self-improving [PIM82, PIM84a]. self-similar [ZNF87]. self-supporting [MC81, MC84b]. Semianalytic [Wie85]. Semiclassical [Joh88, DP88, ITR^{+}87, MD88, NKS88, Sch88a, Vaz81, Vaz84]. Semiconductor [BCD^{+}89, MV83, MV84b]. semiconductors [MV81, MV83, MV84a, MV84b, NW84, WW84b]. semiempirical [BTWN87, PIM82, PIM84a]. sense [Bok84]. separable [LM85]. separation [MR84b]. separatrices [BHM81]. Sequence [EBS88, AZK83, CG87, Kno84]. Sequential [KCA84, Bir85]. SERC [THH82]. Series [Ano89h, FK86b, Gri87, Sch86a, Sta87b, Ano82a, CM81a, CM84c, CB88, Del84b, DM88a, Eas84b, Gro89b, Jam84, Nex80b, RA82, RA84, ST84d, Tru88, Ver83, Ver84a, Zlo81, Zlo84c]. server [O’N89]. service [Ano89a, THH82]. SESM [McI84]. set [AN87, Ano89h, Bas87b, Com80a, Com84, HD86, Han84a, KC81a, KC81b, KC84a, KC84b, Öpi87, SK82, SK84, SV84, Tau84a]. sets [Com80b, FSP87, KS84a, Moo84d, Poo83, Rao81, Rib80, SC89, Tol85]. SEURAT [HP83, HP84]. several [LBY86, SK82, SK84]. Shafranov [LWHH82, LWHH84]. shallow [CSC89]. Shape [MM82, MM84b, DG89, Jam80, SBFI87]. shaped [Hof88]. shapes [Tal82, Tal84a, vM84b]. sharable [JN87]. Shared [Ott89]. Shared-memory [Ott89]. Sharing [Rob85]. shasta [WBG84]. sheet [Uga88]. Shell [Ric82, Ric84, All84c, AM84a, AGF87, BDD84, CC89a, Chi84b, DS89, GA84, Hub84a, Klu84a, KL84a, LW84b, MMBW88, MNJL84, Ras80a, Ras84a, Sin84, SK86, Zwa85, KN84b]. shell-model [SK86, Zwa85]. shells [DHN84, DH84a]. shift [CG87, NM84, Smi84e]. shift-register [CG87]. shifted [Cha80]. shifts [KF84, MS84, Ste84c]. Shock [KL84b, Ter88]. shocks [Ods84]. shooting [BL88, CD86]. short [BP84c, Fin80, Fin84b, Ste84c]. short-range [Ste84c]. short-Ranged [Fin84b, Fin80]. shortest [Bas87b]. shortest-path [Bas87b]. shower [Miu87]. showers [Che89, GHO^{+}89, Goo84]. SI [Cam84e]. SIGMA [YSHK85]. SIGMA-1 [YSHK85]. Signal [HRW85, Kri89, SM85a]. significant [O’C84c]. SIGV5D [MM88b]. silicon [BAG^{+}87]. SIMD [BMBW85, WMH82]. similar [ZNF87]. similarity [CC89b]. Simple [Vis84a, FAK82, GSS81, Gli82, GMR84b, Jak84, Man83a, Man87b, Sch89c, Ver82, vM84i]. simplicial [DMM83, DMM84b]. simplified [DL87, FNDL89]. simply [KS89]. Simula [Phi86]. simulate [BSDM88, HR84a, HR86a, HR86b, Rit84a, Rit84b, Rit86, SMMP86, vM84h]. Simulated [SNH^{+}88, WDM84]. simulating [BS86b, MGMZ84, Odo84a, Odo84d, ZS87]. Simulation [AGH^{+}88, AB83, ACC^{+}86, Bai87, BCD^{+}89, BL86, DSM^{+}81, DSM^{+}84, DEW84, FR81, Fin87, Gal81, Jam82, Lea87, LVPG84, LNV84, LB82b, OBLR84, RSP81, RSP84, dBM82, dBM84, ABM89b, AP89, AS89, AGR88, AF82, AF84c, AM81, AMC83, AM84d, AMC84, BB84a, BKJ83, BKJ84, BDK86a, BDK86b, BDK86c, Ber84b, BBH^{+}88, BKHJ83, CMPR88, CNRS82, CM85, Che89, CGG87, Chr84a, Chr84b, CCS89, CMS83, CMS84, CGMS89, CM83, CM84e, DBB84, DMM84a, DM86b, DMM83, DMM84b, DKA85, DD86, Eas86, Eas88, Eas89, EMM81, Fer89, Fin80, Fin82, Fin84a, Fin84b, Fin86, Fre87, GM89, Goo84, GG89, Gro84, Gup88a, Gup88b, HMTO84, IHS^{+}88, JW85, JMS^{+}89, JN85, Kas84, Kir87, Kra86a, Laa86, LTK88, LNS84, MN80, McI84, McN81, McN84, Miu87, MEB88, MT86, MMR89]. simulation [MKRJ89, NFH^{+}88, Odo81, Odo84c, RFS89, RSS85, Ryn82, Ryn84, Sal85, Sch86a, SBM^{+}86, Sne85, SV84, SM86, Swi88, Tat80, Tha89, TGA88, TW89b, TW81, Wan87, WD84, YMWW89, Taj86]. simulations [AB86, Ano87b, Ano89x, BMR84, BLR82, BLR84, Bin87, BT88, BMG^{+}85, DM87, GHO^{+}89, GWTM83, GDK89, HH82, MMS87, MKH86, Mic87, Moh89, Mül87, MGTD84, Mur88, PBKW85, Pou88a, Rap85, Tou89, Uga88, Voh88, ZHR81]. simulator [Keg81, Keg84]. sinc [BKK86]. since [Van81a]. Singer [HVS87]. Single [CDN^{+}87, Lew84c, Rei86, DW88, Hub84a, Kal82, Kal83, MLL86, MT84a, Sal87, SDH84]. Single-particle [CDN^{+}87, Lew84c]. single-particle-inclusive [MT84a]. Single-phonon [Rei86]. single-variate [Sal87]. singlet [BR85]. singlet-singlet [BR85]. singular [Gad80, Mik80, MP82]. Sipsol [Jes84]. site [NHK82]. Skew [PP85]. Skyrme [Sal85]. slab [Cop84d, Ede84]. slabs [LW85]. SLAM [ACR88]. Slater [Yur84, AT84, Des84a, DF84, Gol84b, Lab84d, ST84e, WB84].

Slater-Transform-Preuss [Yur84]. Slater-type [AT84, DF84, ST84e]. SLC [BGGW^{+}89]. slightly [Moo83]. Slit [Deu84a, Deu84b, Deu84c, Deu84d]. slot [DM89a]. Small [BC88b, KKLS82, Aur84a, Aur84b, DTE84, Deu84a, Deu84b, Deu84c, Deu84d, ETH84, EA83, EA84, Gli84, LBY86, Rib80, Ste84d, Szé85]. small-computer [Gli84]. smearing [Deu84a, Deu84b, Deu84c, Deu84d]. SMOOS [Zlo81, Zlo84c]. Smooth [Ben88]. Smoothing [Woo86b, CP81, CP84, DGPM85, Woo86a]. SNEX [Wie85]. Socio [McI84]. Socio-economic [McI84]. sodium [ÖW84]. soft [Eas89]. Software [BAA^{+}87, Boc87, BCC^{+}85, DP89, Dow82c, Dow82b, Kan87, Kun89, Maz89, Nad86b, Ric86, Rob85, Sch86b, Sch88b, Sch88c, Sto89, Wes87, Wil86, WH89, Ano87-29, AW89, BF89, BGGW^{+}89, BGM87, BCVZ89, Cam86, Car88, CNRS82, CMH87, Dob89, FP89, Gau84c, Gou88, HGJ^{+}89, HB85, HR84b, IKM85, Jam86a, JN87, Kel87, Krá86b, LSW87, Mik86, Nev82, Phi86, Pla86, Rob83, Sau85, SEF^{+}89, SSM85, SBB^{+}89, Stu85, Tau85, Vie88, Was85, Zie88, Sch88b]. Software-safety [Sch88c]. solar [JB80]. Solids [vD87a, AF82, AF84c, Fin82, Fin84a, HRB84, HPT80, Laa86, MEB88, PT84]. Soliton [HGW86, Sal83b, Sal84f, FOW87, Mak80, MLS81]. solute [BH81, BH84a]. Solution [CKA89, Cla82b, DW88, IK81, FBS88, FBS89, Hug84, LW84b, Moo83, MZZ89, Par84, WW80, AS83, And83b, And83a, AS84, And84b, And84a, BH85a, Bel84a, BVL84, Bur85b, CSC89, Car84a, Chi84a, CT84b, Col84b, CSW84, Das83, ERB88, GJ84, GV87, GS88, GS83, Hay80, Hou87, HKK84, IR80, IR85, Jam87, Jes84, KP89, KBMV86, KA87, KRSV82, KK83b, KP85, LNV84, LR84a, Mar84a, MP81a, MP84a, Mar89a, Nor84a, PR87, PT86, Rap81b, RC87, Saa89, SA83, SA84, SFB87, SSF^{+}89, TV86, Vaj82, VPB82, VPB84, WMH82, Wie85]. solutions [Arn84a, Bas87b, Cre80a, Cre84a, DGL89, Ede84, FBW82, Gaj85, Has84, HKT85, LWHH82, LWHH84, LT87, NN84, Rap83, RG80, RG84, TTM^{+}80, WH81, Art85b]. solve [BV80, BH81, BV84, BH84a, BSA86, CL84, Ixa80, SK82, SK84, TSD84]. solver [AKS88a, AKS88b, BMR85b, BH82, BH84b, CH84a, Gru80, Gru84b, KS89, PP82]. Solving [FGSH85, Tol85, AZK83, BBM82, BBM84, Cha84d, DRRvB82, DWC^{+}89, Gru84a, KS84f, Mor84c, RW84, Rud84, SC89, ST82b, ST84f, Sun88, WBG84, tR85]. Some [For89, Goe84, KWPP89, LB82b, Man83b, Man87b, Gaj85, IR85, Sam85, WHD86]. SOS [NKS88]. sound [Mor84a]. Source [KK89, BBR89b, GS89, Hub84d, Pie84b, Sal84b]. sources [BS84c, SG89]. Space [DR85, CCR88, CM81b, CM84f, DE84, DV85, ET84, GWL89, Gre88, HKT85, HSV84, HLM^{+}87, IKRT89, Jab84, Jad84, KK84a, KSE86, Lan82, Lan84a, LT87, MD82, Mor84e, Net84, OM88, PSL88, PZCC89, PB84b, Sim80, Sim84, Vie88, Vrb89, Wen88]. space-charge [DE84]. spacecraft [SG83, SG84b]. spaced [AAMB84, CM84b, Tor82, Tor84, Wai87]. spaces [GS88, Lan80]. spall [RS84b, RS81]. spallation [RS81, RS84b]. spark [BLMU81]. Sparse [DR82, Ben82, Ben84a, BH82, BH84b, DWC^{+}89, Nex80a, SG81]. spatial [Nog83]. Special [RWK^{+}85, Rea89, KS87a, LS84, Poe85, SG83, TS89, Tru88, SG84b]. specialised [VGD^{+}89]. species [Rob84e]. specific [CNRS82]. Specification [Sch85a, Bri86, Gór88]. spectra [AP82, AP84, ABG^{+}81a, AK84, AWG84, BB84a, BPA84, Ber80, Bet84, BR85, CT84a, CSC^{+}86, Cli82, DSM^{+}81, DSM^{+}84, FV82, Gal81, GWL89, Gro84, HMS84, HPT80, Keg85, KEMP81, KE84b, KEMP84, KE84a, Lea87, LBY86, MZM^{+}82, MZM^{+}84, MGTD84, PT84, Puf83, Puf84, RA82, RA84, RRG88, RF83, RF84, Sea84a, SMT88, TUB83, TUB84, TM89, Ver83, Ver84a, VDM84b, War84, WS84b, Zlo82, Zlo84a, Zlo84b, dBM82, dBM84]. Spectral [DG89, Fia84, Lan84b, Rap81a, HL86, HvRM86, Mar89c, MD88, SLH84, YB86, vM84c]. Spectro [Zlo84d]. Spectro-oriented [Zlo84d]. spectrometer [BPSB88, DSSS86, LLC84, Sau85]. Spectroscopic [Ogi83, Ogi84, FS84c, HEL84, Lab84a, Sax84c, Tak85a, vM84b]. spectroscopy [FC84a, LBK88, Ste84d]. spectrum [DO84, HSA82, HSA84, IR80, Keg81, Keg84, KAR81, KAR84b, LNB^{+}88, Sal84b, SR84b, Szé85, TW83, TW84, WV80, WV84, ZNO88, vM84d]. speed [CKV85, Fri88, IKRT89, RSS85, RP89b, RP89a, Run89, SNH^{+}88, SC84, UI85]. SPH [Ben88, Mon88]. sphere [LNB^{+}88]. spheric [Chr84a]. Spherical [AM84e, Bra84a, Cla82b, CRL^{+}89, DS89, Del84b, ELS88, HR87, HKK84, JT84a, KP85, MW84b, MW84c, Tal83, Tal84c, WS82, WS84a, dB82, dB84, Fin80, Fin84b]. spherically [SL83, SL84, VČ88]. spheroidal [BC83a, BC83b, BC84a, BC84b]. Spin [Lan84a, NHK82, ABM89b, AB83, BC88a, Bas87a, BZ84, BSDM88, CMR86, CGG87, DG84, ENSC86, FAK82, HDK84, Hei84, HMTO84, Hub84a, JW85, KF84, KS87a, Klo84b, Lan82, MV83, MV84b, MGTD84, PSL88, PP85, SM84a, Sax84c, SDH84, ZHR81, vMF81, vMF84, vM84g, Ste84b]. spin-dependent [CMR86]. spin-echo [vM84g]. Spin-glass [NHK82, HDK84]. SPIN-HAMILTONIAN [Ste84b]. spin-orbit [Klo84b, MV83, MV84b, PP85, Sax84c]. spin-orbitals [BZ84]. spin-spin [ENSC86]. spins [Coo84b]. Spitbol [WC85]. Spline [CP81, CP84, AAMB84, BD80b, BD81, BD84c, Cha82, Cha84c, CM84b, Fie80, Fie81, Fie84, FV81, KRS83, Tor82, Tor84]. spline-based [BD80b, BD84c, FV81, KRS83]. splines [Bok84, PJ89, PS84c, SBFI87, WSV88]. SPLIT [RWZ87]. SPMD [JMS^{+}89]. Sponsors [Ano84-86, Ano85-29, Ano86-36, Ano89i]. spontaneously [Uga88]. Springer [Ano89h, Eas84a, Eas89, Fin87, Man86, Sch86a, Sta87b, Eas84b]. Springer-Verlag [Ano89h, Eas84a, Eas89, Fin87, Man86, Sch86a, Sta87b]. sputtering [Abr88]. SQSIMUL [BSdlT87]. square [DSK83, DSK84b, Luc84, Moo81a, Moo84a]. squares [Bok84, CAP86b, Han84f, LM84, O’C84b, RFSG84b, Sha87, WS84b, Woo86b, vM84c]. squeezing [BSdlT87]. SSOR [Cam84e]. SSOR-SI [Cam84e]. Stability [GKM^{+}81, ABG^{+}84, Atz86, BLM^{+}88, BHM81, DM86b, Eas87b, Fro84c, GR81, GV87, GTB^{+}81, GPS^{+}81, GTB^{+}84, KT81, Lis84, MLS81, MGD81, SS81, TTAT81, TAK^{+}85b, TTM^{+}80]. Stability-beta [GKM^{+}81]. stabilization [Arn84a]. Stable [BD82, BD84b, Man83a]. stage [MC84a, VKM^{+}82]. Standard [Hal88, Red88, Sta87a, CR84d, DK82, KKLS82, Lid84b, Met87, Odo84d, Sal84a, Tar86]. standardization [Met89, Sch88b]. standardized [Ver89]. Standards [Mye89, Rim81, Hop85, Osl88, Van89]. star [BM84b, HEL84, HE84b, He84a]. stark [Nai84]. Starnet [BCC^{+}89]. stars [Mik84]. State [Gri87, And84c, Cug84, DW88, DeV84, FCW84, Fie80, HDK84, HR87, ITR^{+}87, KP85, LW84b, LR84a, Mac89a, Mor84b, MZZ89, NF83, NF84, PM80, PM83, PM84a, PM84b, Smi84c, Son88, Tal84b, VPB82, VPB84, ZC84]. statements [Sal84a]. states [ANP^{+}85, BC88a, Bra84c, Bra84b, BF84, CCR84a, DK87b, DHD84, Eks84, FGSH85, GWL89, HKT85, HH84, Hir84, KW84b, Lab84a, Lid84b, LT87, Ogi83, Ogi84, SM84a, SM81b, SM84e, SM85c, SM85e, Tal84b, WS80, YP88, YP89]. Static [Mat87, Rin84, BSA86, CMR86, FT84a, FNDL89, Gia84, Han84b, MLL86, Ras80b, Ras84b, SNC80, SNC84, VH87]. stationary [Ede84, WW80, Zlo81, Zlo84c]. stations [Rei89]. Statistical [Chi89, FR86a, ACC^{+}86, Cha84b, CCD85, Har81, Har84a, KN86, Lee89, Zlo81, Zlo84c, ZHR81]. statistics [BSdlT87, Cot87]. Status [DL89, Ken89a, Mon87]. Stauffer [Eas89]. steady [FCW84, KW84b, NF83, NF84, Son88, Tal84b, ZC84]. steady-state [ZC84]. Steed [Bar82, Bar84b]. steering [BE89]. STELLA [SG83, SG84b]. stellar [Hub88]. stellarators [Dom86]. step [IR85, RC85, Zal88]. stereographic [Pre84b]. stereophotographs [AN88]. stereoscopy [KR81]. stiff [PP82]. still [Her89]. stimulated [LVPG84]. Stlplt [CA84]. STO [Jun81, Jun84]. Stochastic [ASS82, Pon81, HLM^{+}87, RG86]. stochastisation [EPS81]. stokes [LBY86, Gaj85, KP89, YB86]. stopping [DM86a]. storage [Cio89, Lan87, Nas84, VVM84, Vis84a, Vis84b, Wil81, BJ81a, BJ84a, HM84d]. store [Jac85b, TVD87]. straggling [CJ84]. straight [GPS84, Pet84, PS84b]. strategies [Fri88, KWPP89]. strengths [Ata85, CRH^{+}89, Cla84, Duf84, FS84d, God84, Hib84c, Sar84a, Sar87]. stretching [MN80]. Strimp [SVP86]. String [CMS84, CMS83]. strings [MMR89]. strip [MS87b]. strong [HKT85, LT87]. strongly [vM84c]. Structure [BMBW85, Sch89c, BB84a, BAG^{+}87, BR85, De 84b, DWVH89, DPH82, DPH84, DGJ^{+}89, FKM^{+}87, GH84b, Gra84b, Gra84c, Han84b, Han84d, HA82, HA83, HA84b, HA84a, Hib84b, Hib84a, Hib84d, Hib84e, Hof84b, HM84c, HM84d, HM84e, HRB84, Isl81, Isl84a, Kan87, KS87a, KN84c, KN84b, Lad86, MV81, MV83, MV84a, MV84b, Mik84, Mir84, Nai84, Nai86, NJ82, NJ84, Odo82, Odo84b, RP89b, RP89a, Sal83a, Sal84d, San82, San84, Sar84b, Sar84c, Tep84, VSP86]. Structure-from-motion [BMBW85]. structure-matching [Hof84b]. structured [Kel87, Ken89b]. structures [Ber84a, CF85, DK82, KM82, KM84b, MZM^{+}82, MZM^{+}84, Put89, Rea89, Sch89a, Sch88c, TGA88, Wil88]. studies [Atz86, HR81, HWL89, LCH^{+}81, SS81]. Study [MP82, BKD83, BKD84, BP84c, Cha84d, Mic87, OSF86, RMV89, ST89a, Sta87a, ZR89]. studying [SVP86]. Stuttgart [Rue89]. sub [Smi84a]. sub-millimeter [Smi84a]. subgroup [PZCC89]. subgroup-symmetry [PZCC89]. Subject [Ano84-87, Ano87-28, Ano88-36, Ano89-32, HE84b]. submitting [Van89]. SUBMMW [Smi84a]. subprogram [Bru85, Bru86]. Subprograms [AS83, And83b, And83a, AS84, And84b, And84a, SA83, SA84, FM85, Jes84]. Subroutine [Sta84, Bok84, Col80a, Col84a, DD85a, GMR84b, LB82a, MCJ81, MCJ84, Mor84b, Rys84, Smi84b, Smi84c, Smi84e, Smi84f, TTT84, Tal83, Tal84c]. Subroutines [DM88a, DM88b, Nob80, Nob84, May81, May84, SV84]. subset [Rom81]. subspace [KPP^{+}89, TV86]. substitution [CL84, Lew84c]. substitutions [MP81b, MP84b]. Successful [Mar89b]. Successive [Got83]. sudden [ONP89, SAC87]. suitable [CWS83, CWS84, De 88]. suite [Jes84, SBB^{+}89, TM89]. suited [Aue84]. sum [AF82, AF84c, Fin82, Fin84a, Lee89]. Summary [Ano84-82, Ano84-83, Ano84-84, Ano84-85, Beh84, Ano88-27, Boc89, Nas87a, Wic87]. summation [CT84b, Di 87, HM81, HM83, HM84a, HM84b]. summations [Klo84c, Rob84a]. Summer [Ano88-27, Beh84]. Summing [CB88]. sums [Bro83, DM88a, HM81, HM83, HM84a, HM84b, Lew84c, May81, May84, RWZ87]. Super [Dav89]. Super-matrix [Dav89]. Supercomputations [Tou87]. Supercomputer [BBC^{+}89, DGL89, BM85a, Pou88b]. Supercomputers [BMR85c, Duf85, Gup88a, IKM85, Meu89, MR86, MR87]. supercomputing [LGG^{+}87, Tru88]. superconducting [ABD82, ABD84]. superconductors [MMR89]. Supergravity [GK80]. supermatrix [Ben82, Ben84a]. supernode [Fli89]. supernova [Mül87]. superposition [RW84]. Supersymmetry [GK80]. Supervector [MMM85]. supplement [Bas88]. support [Nag85]. supporting [JN87, MC81, MC84b]. Surface [MS88, BDJ^{+}89, DD86, EHH88, FLM84, GS83, MS85, MS87a, MR80, MR84a, PDF84, RP89b, RP89a, SP87, PDF82]. surfaces [Cha82, Cha84c, DP84, HPT80, OT80, Ols83, OT84, Ols84, PDF82, PDF84, PT84, RRG88]. survey [Phi80]. susceptibility [vM84e]. Suydam [GR81]. switching [Joh88]. Symbolic [CT89, AB88, DVD87, DFD81, DFD84, HD89, KPPR84, SMM89, VČ88]. symbols [Sha84]. Symmetric [Hub84e, Hub84f, AS83, And83a, AS84, And84a, Cam84f, DEW84, EC84, Gad85, HC86, JT84a, MG84, MW84a, SM84b, SF84a, SM81a, SM81b, SM84d, SM84e, SL83, SL84, VČ88, Wil84a, ZC84]. Symmetries [EFK85, Sch82b, FK86a, Ito86, Ros85, Sch84b, Sch84c]. Symmetrization [RS84f]. Symmetrized [HRB84, HM84e]. symmetrizing [KPP81, KPP84b]. Symmetry [De 84b, HM84c, HM84d, HM84e, Atz86, BDD84, CCS89, JT83, JT84b, MSFR89, Nog83, OCC80, OCC84, PZCC89, PI84, Ská81, Ská84, Ste84b, dBM82, dBM84]. symmetry-Adapted [JT84b, JT83, Ská81, Ská84]. Symposia [Ano82a]. Synchrotron [Rei89, Lan84b]. synthesis [Col88]. System [ST89b, AN88, AW89, BGMM81, BO80, BHSS84, BCKO89, BAA^{+}87, BC87, BBR89b, Buc85, BJ81a, BJ81b, BJ84a, BJ84b, Car80, CDW82, CDW84, CMH87, CgYpXh85, Cis87, CDF^{+}89, CSMHT88, DCC84, Del87, DSSS86, Dob85, DWVH89, EK84, Eck84, ERLD87, GAA^{+}87, GHO^{+}89, GSS81, Gli82, Gli84, GLST89, HAG^{+}89, HHK86, HA82, HA83, HA84b, HA84a, HGK^{+}81, Her82b, Hoa89, HS84, HSA82, HSA84, IKRT89, IA86, Jac85b, Jac85a, JR84, JM82, Kas82, KKLS82, KEMP81, KEMP84, KRSV82, KS86a, Lid84b, LEF^{+}89, LBK88, MLS84, MLo84, MPS81, MP81b, MP84b, MM88a, Moo83, MP89, OL88, OY80, PF86, Per87, PFGP82, QTZ^{+}87, QAD^{+}89, Rom81, Rus86, Rys87, SEF^{+}89, Ska82, Sko89, Sli89, SdAD^{+}88, Str87, Suf89, TTT87, Tri84, Van81a, Ver84b, Via87, WV89, WH89, YFL87]. system [ZS87, Gou88]. systematic [Kal82]. systematization [Ezh84]. Systems [Bur85b, LB82b, Mik80, MB83, Sch86d, Abr88, ABM89b, ACC^{+}86, BMR85b, BCFK85, BS83, BS84a, Blo82a, Blo82b, BKHJ83, BCC^{+}85, BC88b, CL84, Cam85, Cam86, CHS87, CKA89, CLS^{+}84, CH84d, DW89, DM88b, DV85, Eld84, Ghi87, GV89, Gór88, Hoe86b, JA80, JA84, KS84f, Laa86, LSW87, Man83a, Man87b, May81, May84, Moo84h, MB84, Mou87, Nie88, NKS88, Nov87, Öpi87, PDF82, PDF84, Put89, Ros84a, Ros84b, Sch85c, Sch86c, Sph87, WwYpCg88, Whi89b, YpMzCeCg85, Zal85, ZHR81]. Systolic [RFS89, Män89].

T [Eas84a, Rob85, CM84a]. T-X [CM84a]. T.Node [Fli89]. table [Ano88-35]. tableaux [SM84b]. tables [Ber87b, SM81a, SM84d]. tagging [BDK86c, WPVS89]. tail [TW89a]. take [Rob84d]. taking [HK83, HK84b, Wil89]. Talmi [YpMzCeCg85, Zoh84]. TANAGRA [Ber87c]. tandem [ABD82, ABD84, CO86]. tape [BJ81a, BJ84a]. tapes [Gol84a]. target [CW84c, Hub84d, Keg85, MPM83, MPM84, MS84, MC81, MC84b]. targets [Atz86, KF84]. tasking [Gib85]. Tau [CM84a]. Taylor [Eas84a, CM88]. TDA [TW89b]. TDPOIS [Hou87]. Tearing [KT81, EPS81]. Technical [AGG^{+}82, Zlá80]. technique [CD86, DeG88, Got83, MV81, MV84a, RP89b, RP89a, Sta89, Tro84, ZHR81]. Techniques [Ano88-27, Ano89h, Lis88, Bri86, CMH87, EA89, FMS87, HB88, Kel87, KS87a, MMS86, MN80, SG81, Vie88]. technologies [Dew82]. technology [Her84a, Her88b, McI84, Wil81]. technology/resources [McI84]. telecommunications [Str87]. TEM [BKD83, BKD84]. temperature [BMS87, BCCM87, GH83, GH84a, MVV83, MVV84, MV87]. tension [CMS83, CMS84]. Tensor [RP89b, RP89a, Cha84e, Coo84b, FS84c, Klo84c, Klu84b, LB84, LR84a, LR84b, MZZ89, Rob84a, Sax84b]. terabyte [TVD87]. terminal [Ghi87]. terms [Lab84b, ZC84]. terrain [Ruf85]. test [DK87c, MRT89, Red88]. Tests [Hla86, ZHR81, Cha84a]. tetragonal [Ste84b]. Tetrahedral [SMC81, SMC84, HC88]. text [Day84b, Tau81, Tau84b, Van89]. Textbook [AHVV86]. Thalia [ABG^{+}84]. their [AI86, Bar82, Bar84b, Gau81, Krá86b, Lee82, Lee84, MKS83, MKS84, PJ89, RW85, Sch82a, Tou89, Was85]. theorem [VČ88]. theoretic [KR88]. Theoretical [Fin87, ABMT84, BMR85c, Fog82, HM84d, Kas82, Smi84a, WW84a, WW84c]. theories [AMC83, AMC84, BMR84, BLR82, BLR84, CCS89, DM87, DMM83, DMM84b, KR88, MMS86, MT86]. Theory [ST89b, Ano87b, Ano89x, AM81, AM84d, BB84b, Bai87, BM82a, BM82b, BM82c, BCM83a, BM84a, BMR85a, CM85, CGGK85, CMS83, CMS84, Cre80a, Cre84a, CSW84, CM83, CM84e, DMM84a, EMM81, GLST89, ITR^{+}87, Ken89a, KLS^{+}89, KM85, Lew84c, MLS84, MLo84, Mac89b, Mak80, MS84, MB83, MHP84, MB84, MR86, MR87, NN84, III88, ST84e, Son88, Tat80, VBD88, WP84]. thermal [BS84c, Cop81, Cop84a, Cop84d, Cop84c, Cop84b, CVvWF86, Sal84b]. Thermally [BLB88, LVPG84]. thermodynamic [SNH^{+}88]. thermodynamics [Lad86]. thermophysical [Eck84]. thick [Keg85, KK84b, Tal82, Tal84a]. thin [Hub84e, Hub84f, OM88, RS83, RS84d, Sko84, VM89]. Third [Ras80b, Ras84b, Ano80g, BCKO89, CDF^{+}89, Sil84b, Wil84c]. third-level [BCKO89]. third-order [Sil84b, Wil84c]. Three [HvRM86, MKH86, Tre88, AKS88b, Ass84, BSDM88, BWD84, BDD84, Bra84b, CgYpXh85, CMM84, CM86, CMO86, DM89b, EHL80, EHL84, FBW82, IK81, GTL^{+}86, HH82, HL86, Hou87, KJ85, KR81, MLL86, MR84b, NR86b, OT80, OT84, Per89, PT82, Pis84a, Pis84b, RR84, RG86, Sch81, Sch84d, She84, TW89b, YFL87, dB82, dB84, Aue84]. three-body [Ass84, Bra84b, YFL87]. Three-dimensional [HvRM86, MKH86, Tre88, BWD84, CMM84, CM86, CMO86, DM89b, EHL80, EHL84, IK81, GTL^{+}86, HL86, Hou87, KJ85, KR81, NR86b, OT80, OT84, Pis84a, Pis84b, RR84, RG86, Sch84d, She84, TW89b]. three-fluid [FBW82]. three-open-shell [BDD84]. threshold [Bor82, Bor84]. tidal [HE84b]. tightly [HBF^{+}85]. Time [AEH87, Rus87, vTBF84, Bir85, FS82, IK81, FMS87, GM89, GS89, GSS81, Gli82, Gór88, HAG^{+}89, Has84, Keg81, Keg84, Ken89b, LZ84, Mag81, Mag84a, MF81, MRC85, MC84a, Rob84e, Sch86b, Sch88b, Sch88c, SM85a, Sne85, WMN85, Zal88, ZL84, Irv82]. time-dependent [IK81, GS89, Has84, LZ84, Mag81, Mag84a, MF81, MC84a, ZL84, Irv82]. time-filtering [WMN85]. time-of-flight [Keg81, Keg84]. Timer [HR84b]. timesharing [MM88a]. timestep [Fin86]. timing [HR84b]. tissues [Wie86a]. Titel [Ram85]. Title [DW88]. TL [CAP86b]. Tlaser [DST84b]. TLEED1 [RP89a]. TLEED2 [RP89b]. together [Boc89]. toil [MMM85]. tokamak [BSV^{+}85, EP86, Hof88, IFI84, JBWW86, KBMV86, LCH^{+}81, MGK86, OS86, SVP86, OY80]. Tokamaks [AHS^{+}86, MPIM86, KM86, OCS86, PPHS86, Sch83]. Tokyo [Fin87, Man86]. tolerant [MD81]. tomography [Cha82, Cha84c]. tool [BCVZ89, Col88, Die88, Sch89a]. Toolpack [IH86]. tools [Buc85, Gat89, IKM85, Mik86, Mou89, Tau85]. top [JT83, JT84a, JT84b, Sea84a]. Topic [BSV^{+}85]. topics [Tru88]. topological [KLS^{+}89, MT86]. topology [ACD88, EPS81]. tops [MW84a, MW84b, MW84c, Wil84a]. TORINO [DP88]. tornadoes [CCR88]. toroidal [BWD84, Cap86a, CHL^{+}81, EPS81, GPS^{+}81, HP83, HP84, KP84, MGD81, Mik84, Sch81, Sch83, TAK^{+}85b]. torus [TS84b]. torus-ion [TS84b]. total [BGS81, BGS83, BGS84c, BGS84a, BGS84b, Sal82, Sal84e, SS84b, SMD84, SR84a, WB84]. traces [CDW82, CDW84]. tracing [KBMV86, MCA87, PJ89]. Track [AL81, BO80, BSP84, Bil89, GDH^{+}89, Hau87, Sta89, TA89, VKM^{+}82]. trackfinder [Beh81]. tracking [Ves89]. tracks [Per89]. traditional [MRT89]. trajectories [BT84b, HIM84, HSV84, Pia84]. Trajectory [vdM84a]. Tranal [KPPR84]. Transfer [KK84b, But84, FKU^{+}84, HKK84, Mal84, MPM85, MP89, Nob80, Nob84, PV84, Sal82, Sal84e, Tak85a]. Transform [Yur84, BRB89, Can81, Joh87b, Jun81, Jun84, SM89a, SM89b, Tau80, Tau84c, Zie88, CH84b, Dya86, GT83, GT84, Mar84b]. Transformation [GTR^{+}81, BMV83, BMV84, BM85a, CgYpXh85, Dec89, Gro89b, SZ84, SG84c, WwYpCg88, WSV88]. transformations [BM85b, Dya86, O’C84b]. transforms [BBR89a, CRS87, CT84b, Col84b, MN88, NR86b, NRS88, Tal83, Tal84c, NR86a]. Transient [MRC85, MMF81]. transients [Kam86]. transition [BCFK85, BTS84b, BF84, CFG^{+}88, GPS84, HK83, HK84b, ITR^{+}87, Nai84, Nai86, OTB80, OTB84, Pia84, PS84b, RMV89, Wil84b]. transitions [Bil84, Bil87, DK87b, Kar80, Kar84a, Rid82, Rid84b]. translation [KPPR84]. translations [SMD84, Tau84d]. translator [Bai86]. transmission [Deu84c, Deu84d, ENSC86, Hin81, MR80, MR84a, Ver89]. Transparent [CGM89, Ugn80]. Transport [OS84, Tho84, Cha85, ELS88, GHU81, HBJ84, HP83, HP84, HWL89, KK84b, Lux80, NA84, Red88, RG80, RG84, SPM^{+}88, SL83, SL84, TSD84, WMCH84, WH85, Wie85, Wie86a, WHW87]. Transputer [GH87, ACC^{+}86, BBD^{+}89, GHO^{+}89, HRW85, Hey88, Hoe89]. transputers [BBC^{+}89, CGM89, HD88, Hey88, Hey89, RFS89, SMR87, WV89, WH89]. transverse [MGN80, MGN84]. trapezoidal [CT84b]. TRAPS [HRW85]. traversing [RS83, RS84d]. Trax [Ves89]. Trax-I [Ves89]. treating [Fie81, Fie84]. treatment [AN88, Fie80, GPS84, KSE86, NFH^{+}88, Pia84, PS84b]. tree [Joh85, PD89, Vis84a, Vis84b]. trees [Mur88]. Trends [Gai89, Mas81, Wil81]. triangular [BCCM87]. Triatom [Ten86]. triatomic [Sch88a, SMT88, Ten84b, Ten86, TM89]. triatomics [RRG88, Ten83, Ten84a]. triaxial [Fle81]. tridiagonal [MP81a, MP84a]. tridiagonalization [Hay89b, PT82]. Tridyn [MEB88]. Tridyn-binary [MEB88]. Trigger [Fra81, BCKO89, BGGW^{+}89, CDF^{+}89, DGK^{+}81, Gje81, Jac82, Män89]. triggering [TA89]. TRIP [Mag84a]. triple [Jam86b, Nog83]. triple-excitation [Nog83]. triply [MW84c, WS80]. triply-excited [WS80]. TRISTAN [Kan87]. trivariate [DU89]. TRP [GDH^{+}89]. TSDC [CAP86b]. tube [SS81]. TUBE88 [MMO89]. tunneling [ITR^{+}87]. turbulence [CCR88, Pou88a]. twin [SDH84]. twist [CM88]. TWISTER [Ing87]. Two [Ano89h, CFG^{+}87, DVD87, EBS88, GTL^{+}86, GRRP83, GRRP84, HA83, HA84a, Lew82, Lew84a, Luc84, May81, May84, Ods80, SM89b, TVD87, TT89, AKS88a, AMK81, AMK84, AWB82, BC88a, Bas87a, BZ84, Ben82, Ben84a, BDK86a, BDK86b, BDK86c, Bra84c, Cho84, Chr84b, CW84c, CH84a, CDN^{+}87, DK87b, De 84c, DH84b, DP84, EEV89, FKU^{+}84, Goe81, HC83, HC84, HIM84, JA80, JA84, Jam84, Jun81, Jun84, KS81, KS84b, KM82, KM84b, LK88, LR84a, LR84b, Mar89c, MN80, MMK81, MMK84, MMTK88, MZZ89, NR86b, SCF86, Tak85a, WBG84, YFL87, Zlo84b]. Two- [GTL^{+}86, NR86b]. two-body [AMK81, AMK84, Bas87a, Bra84c, EEV89, HIM84]. two-centre [Jun81, Jun84]. two-centre-type [CDN^{+}87]. two-cluster [KM82, KM84b]. Two-dimensional [Lew82, Lew84a, Luc84, Ods80, CW84c, AWB82, Cho84, CH84a, DH84b, Goe81, LK88, MMK81, MMK84, MMTK88, WBG84, Zlo84b]. Two-electron [EBS88, Ben82, Ben84a]. two-layered [Chr84b]. two-level [MN80]. two-nucleon [FKU^{+}84, LR84a, LR84b, Tak85a]. two-nucleons [MZZ89]. two-photon [BDK86a, BDK86b, BDK86c]. two-volume [Ano89h]. TWODEPEP [SFB87]. type [AT84, CDN^{+}87, DF84, DH84b, HKS83, HKS84, Mat84, SS84b, SMD84, SR84a, ST84e, TT82, TT84].

UA1 [CDF^{+}89, Sph87]. ucg [KPP84a]. UCL [BES^{+}87, LLA^{+}89]. UFMulti [AW89]. UK [Gau84a, Gau84c]. Ultrafast [Wan87]. ultrasonic [LNV84]. unaligned [Eks84]. uncertainty [AGF87]. underdense [MCJ81, MCJ84]. unequally [AAMB84, CM84b, Tor82, Tor84]. unfolding [RS80, RS84c]. UNIDFT [CRS87]. unification [BMG^{+}85]. Uniform [IK85, CT82, Ito86, Rib80]. uniformly [Wai87]. UNIMOL [Ryn82, Ryn84]. unimolecular [Ryn82, Ryn84]. unit [CDF^{+}89, FR84]. unitary [ADO84a, GS88, KS84a, KS87a, SM85c, SM85b, SM85d, SM85e, TS84a, Tho87]. united [Lid84b]. universal [BPSB88]. University [AGG^{+}82, Duc86]. UNIX [CGMS89, MM88a, Wie89, WBV^{+}89]. update [Jac85b, SB87, VSP86]. updating [Dec89]. UPEAK [Zlo84d]. upstream [Ter88]. upwind [EFES83]. US$39.50 [Rob85]. US$45 [Art85b]. US$69 [Art85a]. USA [DL89]. Use [BG81, CMH87, CN87, DGK^{+}81, GB87, Gro89a, HBM89, Hoe89, Moh89, QTZ^{+}87, SN84, And86, Ano87-29, BBD^{+}89, Car88, CGM89, DG89, Duf85, Dun87, GJ84, Gau81, Gau84c, Her84a, Jam88, KBMV86, Lan87, MTN^{+}82, Meu89, Nex80b, PJ89, Pou88b, RW85, Tar86, Tau84a, Van89, WV89]. used [Put89]. useful [Got83]. User [Sko89, Vis84b, AD84]. User-friendly [Vis84b]. User-oriented [Sko89]. Users [EK84, Pis84b]. Using [FR81, FP89, GK80, Par82, Pur82, WB82, AT84, All84b, AF82, AF84c, ABG^{+}84, AN88, BKvLSE80, BKvLSE84, Bas87b, Bas87c, BBC^{+}84, BBL^{+}84, BES^{+}87, BKD83, BKD84, BM85b, BT88, CWS83, CWS84, CSC89, CC89a, CC89b, CH88, CT89, CGG87, CD86, CGMS89, DFOP89, FK86a, Fin82, Fin84a, Gaj85, Gel85, Gro89b, HD88, HRS81, HRS84, HvRM86, Hof84b, Jes85, Kal82, Kar80, Kar84a, Kel87, KMS88, KSG82, KN84c, LLA^{+}89, LZ84, LTK88, LBK88, Luc84, MT81, MS84, Mic87, MC81, MC84b, Nas84, NJ82, NW84, NJ84, Nie88, Nog83, NKS88, O’C84b, PP85, Pis84a, Pre84a, RFS89, Rao81, Rap85, Ros85, Rus86, Sal82, Sal84e, SMR87, ST84e, Tat80, Ten84b, TBB87, Ves89, VČ88, VSH83, WMH82, ZL84]. utility [BS86a, CR84d, HRL84]. UV [BPSB88]. UV-VIS [BPSB88].

V [Eas89, WS80]. V103 [HW84b]. V2 [Ost85]. V2/64 [Ost85]. V4 [TTT84]. vacancies [WD84, WDM84]. vacuum [Dom86, TBG80]. valence [BZ84]. value [CR84d, EHH88, MP82]. values [ADV81, ADV84, DK87a, Hub84a, RFSG84a, RFSG84b]. vanishing [Arn84a]. VANVLK [III88]. Variable [ASM84, ASM80, ADV81, ADV84, DW88, HEL84, HE84b, RC85]. variables [De 84c, HD89, Hel80, Mor84e, NP84, O’C84c]. Variance [Lux80]. variate [Sal87]. variation [Rob84e]. Variational [BWD84, Lao84, AT84, CH88, DHD84, ITR^{+}87, Pon81, Rud84, SC89, TT89]. variationally [STN89, Tal89, Ten83, Ten84a]. various [Dya86]. Vavilov [Sch84a]. VAX [CHS87, Ghi87, Gro87, Hoe89]. VAX-FPS [Gro87]. VAX/VMS [CHS87]. Vector [DW84, Eas87a, FR81, Har89, KP81a, MTN^{+}82, Nus84, And86, BA86, CNRS82, CN84, Cli82, DD85a, Gue89, HH82, HD86, Hoe86b, HVS87, Hou87, KK88, Klu84b, KS84h, Lee82, Lee84, MMM85, MBW89, MHP84, MMR89, NFH^{+}88, Nev82, Per82, Sch89c, SSM85, Sch85c, SW80b, SW84, Tem82, Tol85, Tru88, UI85, VSH83, VH87, vdV89, Bra86]. vector-coupling [SW80b, SW84]. vector-multiplication [Cli82]. vector-processing [DD85a]. vector-processor [Hou87]. Vector33 [Har89]. Vectorised [GM89, Fin82, Fin84a]. vectorizable [BH82, BH84b]. Vectorization [CMO86, Gup88b, Hau87, Miu87, TKS^{+}85, VSH83, Cha85, Mar89b]. Vectorized [GDK89, HSV84, ACR88, BM82a, BMR84, BSDM88, CMR86, CMPR88, Car89, CGGK85, CGB^{+}89, KLS^{+}89, MKRJ89, NR86b, NRS88, VH87, Voh88, Wan87, YMWW89]. Vectorizing [BM82b, BM82c, BM84a, EBS88]. vectors [CLS88, HM84c]. Vectra [Jud87]. Verkerk [BD81]. Verlag [Ano89h, Eas84a, Eas89, Fin87, Man86, Sch86a, Sta87b]. Versatile [BPA84, KE84a, Sar84c]. Version [DWVH89, KKP82, KKP84, AI88, BS87, BBL^{+}84, Bru86, Chi84a, Chi84b, CGB^{+}89, Cre81, Cre84b, DTE84, DCC84, Fin82, Fin84a, Fli89, FBS88, FBS89, FS84c, Fun84a, Fun84b, Gia84, Gra84c, Hib84e, Jud87, KE84b, LS84, MS87a, Moo84f, NAS87b, OTB80, OTB84, PCL89, Puf83, Puf84, Ras80b, Ras84b, Sax84b, Sax84a, Sax84c, SH82, SH84, Sjö86, SB87, Tho87, VDM84b, War84, WW84a]. versions [MLL86]. versus [Mye89, Ott89, Sch84d]. vertex [MS87b]. Very [SRS85, BMR85b, CH87, CGG87, Di 87, RSS85]. VF [Suc87]. VI [Car84a, MBW89]. via [BCCM87, Rit84b, Ste87, Taj86, WD84]. VIASKL [AGF87]. vibrating [Sch88a]. vibrating/rotating [Sch88a]. vibration [Bou87, CLS88, Ogi83, Ogi84, RRG88, SMT88]. vibration-rotation [RRG88, SMT88]. vibration-rotational [Bou87, Ogi83, Ogi84].

Vibrational [Mag84b, Bil84, Bil87, BOT86, CH87, CH88, CRH^{+}89, DK87b, EHH88, FH88b, FE88, Joh88, LNB^{+}88, ST81, ST84a, SK82, SK84, Ten83, Ten84b, Ten84a, Ten86, TM89, TSD84, TTL88, WB88, YP86]. vibrational-rotational [DK87b, EHH88]. vibrationally [JT84a, Nov87]. vibrations [BWBL88, HC88, WW84c]. video [JHRR87]. Vienna [Kil80]. view [Gri89, SG83, SG84b]. Vieweg [And86]. views [OT80, Ols83, OT84, Ols84]. Viking [KGAL88]. Ville [Zie88]. virtual [DO84, GWTM83, TW83, TW84, WV80, WV84, ZNO88]. VIS [BPSB88]. viscosities [Mor84a]. viscous [Ede84]. vision [BMBW85, HGJ^{+}89]. visual [AMB84]. visually [ABMT84]. VLAM [SNC80, SNC84]. Vlasov [ERB88, IGB^{+}89, NBI^{+}85]. Vleck [III88]. VLSI [Jes85]. VM [CHS87, McC87]. VM/CMS [CHS87, McC87]. VMOMS [LWHH82, LWHH84]. VMS [CHS87]. Vogelaere [CM84d]. volume [And86, Ano80b, Ano80c, Ano80i, Ano80j, Ano80t, Ano80u, Ano81a, Ano81b, Ano81c, Ano81d, Ano81h, Ano81i, Ano81j, Ano81k, Ano81u, Ano81v, Ano81w, Ano81x, Ano82d, Ano82e, Ano82f, Ano82m, Ano82n, Ano82o, Ano82-28, Ano82-29, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83p, Ano83q, Ano83r, Ano84d, Ano84e, Ano84f, Ano84j, Ano84k, Ano84l, Ano84-79, Ano84-80, Ano84-81, Ano85b, Ano85c, Ano85d, Ano85e, Ano85h, Ano85i, Ano85j, Ano85k, Ano85z, Ano85-27, Ano85-28, Ano86a, Ano86b, Ano86c, Ano86d, Ano86j, Ano86k, Ano86l, Ano86m, Ano86-31, Ano86-32, Ano86-33, Ano86-34, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87x, Ano87y, Ano87z, Ano87-27, Ano87-28, Ano88c, Ano88d]. volume [Ano88e, Ano88f, Ano88m, Ano88n, Ano88o, Ano88p, Ano88-31, Ano88-32, Ano88-33, Ano88-34, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89l, Ano89m, Ano89n, Ano89y, Ano89-29, Ano89-30, Ano89-31, Ano89-32, KK84a, Tru88]. volumes [Ano81y, Ano82-30, Ano84i, Ano89h, Wie86b]. vortices [MMR89]. vorticity [Bra87a]. VP [MMM85]. VP-100 [MMM85]. VP-100/200 [MMM85]. VP200 [MHP84]. VPM [CSBB82, CSBB84]. Vries [SBFI87]. vs [Pag87].

W [And86, Eas82, Eas89, Fin87]. W. [Tru88]. wall [GHU81, NF83, NF84]. Waller [Rei87]. walls [Hub84e, Hub84f]. want [Lev89]. water [ASS82, CSC89, CM81b, CH84d, CM84f]. WATERLOPP [Ost85]. Wave [RFSG84b, And84c, AHVV86, BKvLSE80, BKvLSE84, BFSG84, BC83a, BC83b, BC84a, BC84b, CD86, CB88, CHMM84, Cug84, CDN^{+}87, DL86, FBW82, Has84, Hib84c, HFG88, JBWW86, Jam84, KKP82, KPH84, KKP84, KGAL88, KK83b, MV88, Nes88, PPHS86, Sal84c, Sar87, SAC87, Sea84b, Smi84c, Smi84a, Ste84c, SAC^{+}86, TCR84, WV80, WV84, ZNO88]. wave-functions [Sal84c]. wave-particle [KGAL88, SAC^{+}86]. wavefunction [ABH^{+}84, Bar84d, Kla84]. wavefunctions [DeV84, GRRP83, GRRP84, KKLP84, Kno84]. waves [CCR88, GS83, IFI84, KM86, MGK86, MKS83, MKS84, Ter88]. weak [LO86]. weakly [WW80]. Weather [Dic82, Gib85]. Weber [WW84b].

Weight [ADO83, ADO84a, ADO84b, KR88]. Weighted [Les89, RVD^{+}81]. weights [SH82, Tak88a, Tak88b, SH84]. Weizmann [GA84]. well [DHD84]. Wendroff [EFES81]. Weyl [BK84a]. whether [Woo85]. which [BGMM81, De 84c, Hei84, MMO89, Rob84d]. whole [Moo83]. wide [Flu89, WML^{+}82]. widths [Ata85, TTL88]. Wigner [AD84, TN84, Zie88]. Wiley [And86, Eas80, Jes87]. Wilson [BCM83a, CGGK85, CMS83, CMS84, DeG88]. Winkelmann [Eas89]. wire [Lee82, Lee84]. wires [VM89]. within [Bet84, Car80, HMS84, QTZ^{+}87, Rob84e, ST82a, ST84c]. without [Buc83, Buc84, DG89, DH84a, MMM85]. WIZJET [Odo84d]. Wood [Cug84]. Woods [CDN^{+}87, DHD84, HH84, Hir84]. word [KC81b, KC84a]. work [KM84a]. Worker [BEM84]. workstation [BBC^{+}87]. workstations [DFOP89, vD87a]. world [Ano89a, Rea84]. writer [Bar84a]. written [WC85]. Wronskian [Ito88]. WZ [Hel82].

X [Jes87, KKP82, Rob85, CM84a, Deu84a, Deu84b, Deu84c, Deu84d, DPH82, DPH84, FHMR84, FFH84, FK84, Fer84, FR84, KPH84, KKP84, Klo84a, KM85, MZM^{+}82, MZM^{+}84, MGMZ84, NR86a, PDNE88, Rei85, Rei86, RM89, RS85, VSP86, WML^{+}82, vM84i]. X-MP [KM85, NR86a, RS85]. X-ray [Deu84a, Deu84b, Deu84c, Deu84d, DPH82, DPH84, FHMR84, FFH84, FK84, Fer84, FR84, MZM^{+}82, MZM^{+}84, MGMZ84, Rei85, Rei86, RM89, VSP86, WML^{+}82, vM84i]. X-rays [PDNE88]. xanes [DPH84, DPH82]. XFIT [MGMZ84]. xii [Duc86, Eas82, Nad86a]. xiv [Art85b]. XRF [Fer89]. xxi [Art85a]. xxiii [Taj86].

Yamanouchi [MG84]. years [CFG^{+}87]. yield [ASM80, ASM84]. York [Art85b, Eas84a, Fin87, Man86, Taj86]. Yukawa [PIM80, PIM84b]. Yukawa-plus-exponential [PIM80, PIM84b].

Z [Lew84d]. Z-expansion [Lew84d]. Zabolitzky [Eas89]. ZEBRA [Sch89a]. Zeeman [DG84]. Zener [Bie83, Bie84]. zero [Bas87a, De 84c]. zero-spin [Bas87a]. ZEROD [LJN85]. zeros [BCM83b, BCM84b, Cam84d]. Zeta [BD84a, BD80a]. ZEUS [BCKO89, HBM89, KPST89, LEF^{+}89, WV89, Dob89]. zincblende [KN84c, KN84b, NJ82, NJ84].

References

Allard:1984:HFN

[AAH84] J. F. Allard, A. Abzouzi, and B. Houssais. Hartree–Fock nuclear calculations with Gaussian potentials. Computer Physics Communications, 35(1–3):C–118, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823760.

Anderson:1984:CSI

[AAMB84] J. Anderson, R. W. B. Ardill, K. J. M. Moriarty, and R. C. Beckwith. A cubic spline interpolation of unequally spaced data points. Computer Physics Communications, 35(1–3):C–531–C– 532, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827113.

Amsler:1983:SAD

[AB83] C. Amsler and J. C. Bizot. Simulation of angular distributions and correlations in the decay of particles with spin. Computer Physics Communications, 30(1):21–30, July/August 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583901182.

Ahlrichs:1986:NRM

[AB86] Reinhart Ahlrichs and Stefan Brode. A new rigid motion algorithm for MD simulations. Computer Physics Communications, 42(1):59–64, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586902316.

Autin:1988:ASC

[AB88] B. Autin and J. Bengtsson. Application of symbolic computation to the search of complicated primitives; the example of the “betatron” integrals. Computer Physics Communications, 48(2):181–195, February 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588900392.

Antony:1982:DPB

[ABD82] M. S. Antony, J. M. Britz, and J. Denimal. Delphine: Program to bunch a DC beam through a tandem accelerator for injection into a cyclotron with superconducting magnet. Computer Physics Communications, 25(3):311–324, March 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900261.

Antony:1984:DPB

[ABD84] M. S. Antony, J. M. Britz, and J. Denimal. Delphine: Program to bunch a dc beam through a tandem accelerator for injection into a cyclotron with superconducting magnet. Computer Physics Communications, 35(1–3):C–793, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829173.

Appert:1981:NPA

[ABG^{+}81a] K. Appert, B. Balet, R. Gruber, F. Troyon, and J. Vaclavik. Numerical problems associated with the presence of continuous spectra. Computer Physics Communications, 24(3–4):329–335, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901557.

Azzolini:1981:MIF

[ABG81b] M. C. Azzolini, R. Biancastelli, and A. Giordana. Multiprocessing implementation on the FASTBUS. Computer Physics Communications, 22(2–3):211–213, April 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900515.

Appert:1984:TOD

[ABG^{+}84] K. Appert, D. Berger, R. Gruber, F. Troyon, and K. V. Roberts. Thalia — a one-dimensional magnetohydrodynamic stability program using the method of finite elements. Computer Physics Communications, 35(1–3):C–323–C–324, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825473.

Allard:1984:NHF

[ABH^{+}84] J. Allard, N. Boumahrat, B. Houssais, M. Hadj Hassan, and M. Lambert. A nuclear Hartree–Fock intrinsic wavefunction projection program. Computer Physics Communications, 35(1–3):C–156, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824078.

Adye:1989:OCD

[ABLB^{+}89] T. Adye, T. Berners-Lee, S. Brobecker, A. Camacho, D. Davids, F. Harris, and P. Lorenz. Online communications in the Delphi experiment. Computer Physics Communications, 57(1–3):466–471, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902646.

Aiello:1989:HIG

[ABM89a] G. R. Aiello, M. Budinich, and E. Milotti. Hardware implementation of a GFSR pseudo-random number generator. Computer Physics Communications, 56(2):135–139, December 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900143.

Aiello:1989:PPS

[ABM89b] G. R. Aiello, M. Budinich, and E. Milotti. A parallel processor for the simulation of Ising spin systems. Computer Physics Communications, 56(2):141–146, December 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900155.

Anderson:1984:PPVb

[ABMT84] J. Anderson, R. C. Beckwith, K. J. M. Moriarty, and J. H. Tabor. A plotting package for visually comparing theoretical and experimental results. Computer Physics Communications, 35(1–3):C–517, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826983.

Abril:1988:AMG

[Abr88] Isabel Abril. Alacant: Modeling of glow discharge sputtering systems. Computer Physics Communications, 51(3):413–422, November 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901543.

Askew:1986:SSM

[ACC^{+}86] C. R. Askew, D. B. Carpenter, J. T. Chalker, A. J. G. Hey, D. A. Nicole, and D. J. Pritchard. Simulation of statistical mechanical systems on transputer arrays. Computer Physics Communications, 42(1):21–26, September 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586902250.

Addessio:1988:GTG

[ACD88] Frank L. Addessio, Michael Cline, and John K. Dukowicz. A general topology, Godunov method. Computer Physics Communications, 48(1):65–73, January 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900240.

Andersen:1988:SVC

[ACR88] David R. Andersen, Robert Cuykendall, and Jeffrey J. Regan. SLAM — vectorized calculation of refraction and reflection for a Gaussian beam at a nonlinear interface in the presence of a diffusive kerr-like nonlinearity. Computer Physics Communications, 48(2):255–264, February 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900458.

Adelantado:1985:ECP

[ACSB85] M. Adelantado, D. Comte, P. Siron, and Ph. Berger. Expression of concurrency and parallelism in an MIMD environment. Computer Physics Communications, 37(1–3):63–67, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901365.

Akiyama:1984:UGF

[AD84] Yoshimi Akiyama and J. P. Draayer. A user’s guide to Fortran programs for Wigner and Racah coefficients of SU(3). Computer Physics Communications, 35(1–3):C–194, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482439X.

Amar:1983:WMC

[ADO83] V. Amar, U. Dozzio, and C. Oleari. Weight multiplicity for Cartan classes. Computer Physics Communications, 29(2):201– 209, April 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900747.

Amar:1984:WMU

[ADO84a] V. Amar, U. Dozzio, and C. Oleari. Weight multiplicity for unitary groups. Computer Physics Communications, 35(1–3): C–478, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584826740.

Amary:1984:WMC

[ADO84b] V. Amary, U. Dozzio, and C. Oleari. Weight multiplicity for Cartan classes. Computer Physics Communications, 35(1–3): C–875, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829793.

Anagnostatos:1981:IVP

[ADV81] G. S. Anagnostatos, K. Demakos, and A. Vassiliou. Initial values of parameters for variable moment of inertia models. Computer Physics Communications, 24(2):197–203, November 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900941.

Anagnostatos:1984:IVP

[ADV84] G. S. Anagnostatos, K. Demakos, and A. Vassiliou. Initial values of parameters for variable moment of inertia models. Computer Physics Communications, 35(1–3):C–760, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828908.

Amini:1987:TIP

[AEH87] M. Amini, J. W. Eastwood, and R. W. Hockney. Time integration in particle models. Computer Physics Communications, 44(1–2):83–93, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900191.

Anastasiou:1982:PDS

[AF82] N. Anastasiou and D. Fincham. Programs for the dynamic simulation of liquids and solids II. MDIONS: Rigid ions using the Ewald sum. Computer Physics Communications, 25(2):159–176, February 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900327.

Abecasis:1984:ELCb

[AF84a] S. M. Abecasis and F. R. Femenia. Energy level calculations with AROVMI model. Computer Physics Communications, 35(1–3):C–159, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824108.

Abecasis:1984:ELCc

[AF84b] S. M. Abecasis and F. R. Femenia. Energy-level calculations with the extended AROVMI model. Computer Physics Communications, 35(1–3):C–230, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824698.

Anastasiou:1984:PDS

[AF84c] N. Anastasiou and D. Fincham. Programs for the dynamic simulation of liquids and solids II. Moions: Rigid ions using the Ewald sum. Computer Physics Communications, 35(1–3): C–778, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829057.

Abecasis:1984:ELCa

[AFH84] S. M. Abecasis, F. R. Femenia, and E. S. Hernandez. Energy level calculations in Davydov model. Computer Physics Communications, 35(1–3):C–61–C–62, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823279.

Arcos:1987:VCP

[AGF87] JoséM. Los Arcos, Agustín Grau, and Antonio Fernandez. VIASKL: a computer program to evaluate the liquid scintillation counting efficiency and its associated uncertainty for K-L-atomic shell electron-capture nuclides. Computer Physics Communications, 44(1–2):209–220, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900282.

Arlt:1982:CCN

[AGG^{+}82] R. Arlt, F. Gleisberg, W. Grimm, R. Krause, H. Märten, W. Meiling, G. Ortlepp, G. Pausch, J. Pöthig, K. Seidel, R. Teichner, W. Wagner, and F. Weidhase. Computer control in nuclear physics measurements at the technical university of dresden. Computer Physics Communications, 26(1–2):195– 199, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901680.

Agakishiev:1988:JSI

[AGH^{+}88] H. N. Agakishiev, V. G. Grishin, K. Hänssgen, T. Kanarek, and R. M. Mechtiev. JADJAD: Simulation of inelastic nucleus-nucleus interactions below 5 GeV. Computer Physics Communications, 48(3):391–405, March 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902032.

Ambrosiano:1988:FMM

[AGR88] John Ambrosiano, Leslie Greengard, and Vladimir Rokhlin. The fast multipole method for gridless particle simulation. Computer Physics Communications, 48(1):117–125, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558890029X.

Anderson:1986:PCM

[AHKM86] David V. Anderson, Eric J. Horowitz, Alice E. Koniges, and Michael G. McCoy. Parallel computing and multitasking. Computer Physics Communications, 43(1):69–87, December 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900548.

Appert:1986:CRH

[AHS^{+}86] K. Appert, T. Hellsten, O. Sauter, S. Succi, J. Vaclavik, and L. Villard. Computing of RF heating and current drive in tokamaks. Computer Physics Communications, 43(1):125–141, December 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900573.

Appert:1986:TFE

[AHVV86] K. Appert, T. Hellsten, J. Vaclavik, and L. Villard. Textbook finite element methods applied to linear wave propagation problems involving conversion and absorption. Computer Physics Communications, 40(1):73–93, May 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901499.

Armenise:1986:MCN

[AI86] N. Armenise and G. Iaselli. A method to classify neutrino events according to their completeness. Computer Physics Communications, 39(3):333–341, April 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900949.

Aldea:1988:FAE

[AI88] N. Aldea and E. Indrea. Fourier analysis of EXAFS data — a self-contained FORTRAN program-package — a second version. Computer Physics Communications, 51(3):451–462, November 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901580.

Assimakopoulos:1984:AIP

[AK84] P. A. Assimakopoulos and S. Kossionides. Anna: An interactive program for analysis of one-dimensional pulse-height spectra. Computer Physics Communications, 35(1–3):C–359, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825783.

Anderson:1988:CPCa

[AKS88a] D. V. Anderson, A. E. Koniges, and D. E. Shumaker. CPDES2: a preconditioned conjugate gradient solver for linear asymmetric matrix equations arising from coupled partial differential equations in two dimensions. Computer Physics Communications, 51(3):391–403, November 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558890152X.

Anderson:1988:CPCb

[AKS88b] D. V. Anderson, A. E. Koniges, and D. E. Shumaker. CPDES3: a preconditioned conjugate gradient solver for linear asymmetric matrix equations arising from coupled partial differential equations in three dimensions. Computer Physics Communications, 51(3):405–412, November 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901531.

Almehed:1981:TFC

[AL81] S. Almehed and B. Lorstad. Track fitting in cylindrical coordinates. Computer Physics Communications, 22(2–3):209– 210, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900503.

Allison:1984:PCF

[All84a] A. C. Allison. A program to calculate Franck–Condon factors. Computer Physics Communications, 35(1–3):C–4, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822754.

Allison:1984:CAE

[All84b] Arthur C. Allison. The calculation of absorption and elastic cross sections using the optical potential. Computer Physics Communications, 35(1–3):C–135, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823899.

Allison:1984:FPCa

[All84c] D. C. S. Allison. Fractional parentage coefficients for equivalent p shell and equivalent d shell electrons. Computer Physics Communications, 35(1–3):C–2–C–3, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822742.

Ardill:1981:MCS

[AM81] R. W. B. Ardill and K. J. M. Moriarty. Monte Carlo simulation of SU(2) lattice gauge theory. Computer Physics Communications, 24(2):127–134, November 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900862.

Allison:1984:FPCb

[AM84a] D. C. S. Allison and J. E. McNulty. Fractional parentage coefficients for equivalent f shell electrons. Computer Physics Communications, 35(1–3):C–278, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825096.

Ardill:1984:ABF

[AM84b] R. W. B. Ardill and K. J. M. Moriarty. Accurate Bessel functions J_{n}(z), Y_{n}(z), H_{n}^{(1)}(z) and H_{n}^{(2)}(z) of integer order and complex argument. Computer Physics Communications, 35(1–3):C–559, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827344.

Ardill:1984:BFC

[AM84c] R. W. B. Ardill and K. J. M. Moriarty. The Bessel functions J_{0} and J_{1} of complex argument. Computer Physics Communications, 35(1–3):C–409, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826193.

Ardill:1984:MCSa

[AM84d] R. W. B. Ardill and K. J. M. Moriarty. Monte Carlo simulation of SU(2) lattice gauge theory. Computer Physics Communications, 35(1–3):C–751, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828829.

Ardill:1984:SBF

[AM84e] R. W. B. Ardill and K. J. M. Moriarty. Spherical Bessel functions j_{n} and y_{n} of integer order and real argument. Computer Physics Communications, 35(1–3):C–466, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826661.

Anderson:1984:PPVa

[AMB84] J. Anderson, K. J. M. Moriarty, and R. C. Beckwith. A plotting package for visual comparison of points and curves. Computer Physics Communications, 35(1–3):C–293, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825230.

Ardill:1983:MCS

[AMC83] R. W. B. Ardill, K. J. M. Moriarty, and Michael Creutz. Monte–Carlo simulation of pure U(N) and SU(N) lattice gauge theories with fundamental and adjoint couplings. Computer Physics Communications, 29(1):97–108, March 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900309.

Ardill:1984:MCSb

[AMC84] R. W. B. Ardill, K. J. M. Moriarty, and Michael Creutz. Monte-Carlo simulation of pure U(N) and SU(N) lattice gauge theories with fundamental and adjoint couplings. Computer Physics Communications, 35(1–3):C–867–C–868, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829732.

Ardill:1981:PFP

[AMK81] R. W. B. Ardill, K. J. M. Moriarty, and P. Koehler. A program for fitting and plotting amplitudes, polarization and differential cross section data for two-body reactions. Computer Physics Communications, 22(4):419–432, May 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901399.

Ardill:1984:PFP

[AMK84] R. W. B. Ardill, K. J. M. Moriarty, and P. Koehler. A program for fitting and plotting amplitudes, polarization and differential cross section data for two-body reactions. Computer Physics Communications, 35(1–3):C–693–C–694, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482843X.

Allison:1987:CCH

[AN87] D. C. S. Allison and M. T. Noga. Computing the convex hull of a set of points. Computer Physics Communications, 43(3):381–386, February/March 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900555.

Arzumanyan:1988:MST

[AN88] S. G. Arzumanyan and H. V. Navasardyan. Microcomputer system of treatment of chamber stereophotographs using CCD pickups. Computer Physics Communications, 50(1–2):213– 215, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901269.

Anderson:1983:ISSc

[And83a] D. V. Anderson. ICCG3: Subprograms for the solution of a linear symmetric matrix equation arising from A 7, 15, 19 or 27 point 3d discretization. Computer Physics Communications, 30(1):51–57, July/August 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583901224.

Anderson:1983:ISSb

[And83b] D. V. Anderson. ILUCG3: Subprograms for the solution of a linear asymmetric matrix equation arising from A 7, 15, 19 or 27 point 3d discretization. Computer Physics Communications, 30(1):43–49, July/August 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583901212.

Anderson:1984:ISSc

[And84a] D. V. Anderson. ICCG3: Subprograms for the solution of a linear symmetric matrix equation arising from a 7, 15, 19 or 27 point 3D discretization. Computer Physics Communications, 35(1–3):C–897–C–898, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829987.

Anderson:1984:ISSb

[And84b] D. V. Anderson. ILUCG3: Subprograms for the solution of a linear asymmetric matrix equation arising from a 7, 15, 19 or 27 point 3D discretization. Computer Physics Communications, 35(1–3):C–895–C–896, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829975.

Andre:1984:PGP

[And84c] Jean-Marie Andre. Polymol: a general program for the calculation of ground-state wave functions. Computer Physics Communications, 35(1–3):C–45, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823115.

Anderson:1986:BRB

[And86] D. Anderson. Book review: Notes on numerical fluid mechanics, volume 12 The efficient use of vector computers with emphasis on computational fluid dynamics: Eds. W. Schonauer and W. Gentzsch, Vieweg, Braunschweig (available from J. Wiley, London), March 1986. £27.20. Computer Physics Communications, 42(2):301–302, October 1986.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900445.

Anderssen:1988:PCA

[And88] P. S. Anderssen. Personal computers in accelerator control. Computer Physics Communications, 50(1–2):89–99, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558890118X.

Anonymous:1980:A

[Ano80a] Anonymous. Announcement. Computer Physics Communications, 19(1):159, January/March 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465580900727.

Anonymous:1980:AIVa

[Ano80b] Anonymous. Author index to volume 19. Computer Physics Communications, 19(3):399–401, July/August 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900934.

Anonymous:1980:AIVb

[Ano80c] Anonymous. Author index to volume 20. Computer Physics Communications, 20(3):465–467, November 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900272.

Anonymous:1980:Ca

[Ano80d] Anonymous. Committees. Computer Physics Communications, 20(1):x, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900971.

Anonymous:1980:CPCa

[Ano80e] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 19(1):v–vi, January/ March 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900570.

Anonymous:1980:CPCc

[Ano80f] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 21(1):v–vi, December 1, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900739.

Anonymous:1980:CPCb

[Ano80g] Anonymous. Computer Physics Communications instructions to authors (third revision). Computer Physics Communications, 19(1):1–10, January/March 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900594.

Anonymous:1980:Cb

[Ano80h] Anonymous. Contents. Computer Physics Communications, 20(1):xi–xii, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900983.

Anonymous:1980:CVa

[Ano80i] Anonymous. Contents to volume 19. Computer Physics Communications, 19(3):397, July/August 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465580900922.

Anonymous:1980:CVb

[Ano80j] Anonymous. Contents to volume 20. Computer Physics Communications, 20(3):463–464, November 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900260.

Anonymous:1980:EBa

[Ano80k] Anonymous. Editorial Board. Computer Physics Communications, 19(1):??, January/March 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900569.

Anonymous:1980:EBb

[Ano80l] Anonymous. Editorial Board. Computer Physics Communications, 20(1):v–vi, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900958.

Anonymous:1980:ENa

[Ano80m] Anonymous. Erratum notice. Computer Physics Communications, 19(2):271, April/June 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900545.

Anonymous:1980:ENb

[Ano80n] Anonymous. Erratum notice. Computer Physics Communications, 19(2):272, April/June 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900557.

Anonymous:1980:ENc

[Ano80o] Anonymous. Erratum notice. Computer Physics Communications, 19(3):395–396, July/August 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465580900910.

Anonymous:1980:ENd

[Ano80p] Anonymous. Erratum notice. Computer Physics Communications, 20(3):459–460, November 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900235.

Anonymous:1980:ENe

[Ano80q] Anonymous. Erratum notice. Computer Physics Communications, 20(3):461, November 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900247.

Anonymous:1980:ENf

[Ano80r] Anonymous. Erratum notice. Computer Physics Communications, 20(3):462, November 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900259.

Anonymous:1980:PA

[Ano80s] Anonymous. Preliminary announcement. Computer Physics Communications, 20(2):323, October 1980. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465580900119.

Anonymous:1980:PIVa

[Ano80t] Anonymous. Program index to volume 19. Computer Physics Communications, 19(3):402–404, July/August 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900946.

Anonymous:1980:PIVb

[Ano80u] Anonymous. Program index to volume 20. Computer Physics Communications, 20(3):469–470, November 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900284.

Anonymous:1981:AIVa

[Ano81a] Anonymous. Author index to volume 21. Computer Physics Communications, 21(3):441–443, January 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900229.

Anonymous:1981:AIVb

[Ano81b] Anonymous. Author index to volume 22. Computer Physics Communications, 22(4):477–480, May 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901454.

Anonymous:1981:AIVc

[Ano81c] Anonymous. Author index to volume 23. Computer Physics Communications, 23(4):437–440, August 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901843.

Anonymous:1981:AIVd

[Ano81d] Anonymous. Author index to volume 24. Computer Physics Communications, 24(3–4):481–484, December 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901727.

Anonymous:1981:Ca

[Ano81e] Anonymous. Committees. Computer Physics Communications, 22(2–3):vi, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900382.

Anonymous:1981:CAD

[Ano81f] Anonymous. Computing in accelerator design and operation. Computer Physics Communications, 24(2):234, November 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900990.

Anonymous:1981:Cb

[Ano81g] Anonymous. Contents. Computer Physics Communications, 24(3–4):vii–viii, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558190148X.

Anonymous:1981:CVa

[Ano81h] Anonymous. Contents to volume 21. Computer Physics Communications, 21(3):439–440, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465581900217.

Anonymous:1981:CVb

[Ano81i] Anonymous. Contents to volume 22. Computer Physics Communications, 22(4):473–475, May 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465581901442.

Anonymous:1981:CVc

[Ano81j] Anonymous. Contents to volume 23. Computer Physics Communications, 23(4):435–436, August 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465581901831.

Anonymous:1981:CVd

[Ano81k] Anonymous. Contents to volume 24. Computer Physics Communications, 24(3–4):477–479, December 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901715.

Anonymous:1981:EBa

[Ano81l] Anonymous. Editorial Board. Computer Physics Communications, 22(1):v–vi, February/March 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465581900746.

Anonymous:1981:EBb

[Ano81m] Anonymous. Editorial board. Computer Physics Communications, 23(1):v–vi, June 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581901223.

Anonymous:1981:EBc

[Ano81n] Anonymous. Editorial board. Computer Physics Communications, 24(1):v–vi, September/October 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901004.

Anonymous:1981:ENa

[Ano81o] Anonymous. Erratum notice. Computer Physics Communications, 21(3):437, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900205.

Anonymous:1981:ENb

[Ano81p] Anonymous. Erratum notice. Computer Physics Communications, 22(1):103–104, February/March 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900837.

Anonymous:1981:ENc

[Ano81q] Anonymous. Erratum notice. Computer Physics Communications, 23(1):109–113, June 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581901338.

Anonymous:1981:ENd

[Ano81r] Anonymous. Erratum notice. Computer Physics Communications, 23(2):221, July 1, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900357.

Anonymous:1981:ENe

[Ano81s] Anonymous. Erratum notice. Computer Physics Communications, 23(2):222, July 1, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900369.

Anonymous:1981:MCM

[Ano81t] Anonymous. Methods in computational molecular physics. Computer Physics Communications, 24(2):233, November 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900989.

Anonymous:1981:PIVa

[Ano81u] Anonymous. Program index to volume 21. Computer Physics Communications, 21(3):444–446, January 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900230.

Anonymous:1981:PIVc

[Ano81v] Anonymous. Program index to volume 22. Computer Physics Communications, 22(4):481–482, May 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901466.

Anonymous:1981:PIVd

[Ano81w] Anonymous. Program index to volume 23. Computer Physics Communications, 23(4):441–444, August 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901855.

Anonymous:1981:PIVe

[Ano81x] Anonymous. Program index to volume 24. Computer Physics Communications, 24(3–4):485–486, December 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901739.

Anonymous:1981:PIVb

[Ano81y] Anonymous. Program index to volumes 16–20. Computer Physics Communications, 21(3):447–458, January 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900242.

Anonymous:1982:SSS

[Ano82a] Anonymous. 1983 Sanibel Symposia series. Computer Physics Communications, 27(3):323, September 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901795.

Anonymous:1982:Aa

[Ano82b] Anonymous. Announcement. Computer Physics Communications, 25(1):112, January 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900509.

Anonymous:1982:Ab

[Ano82c] Anonymous. Announcement. Computer Physics Communications, 27(1):103, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900157.

Anonymous:1982:AIVa

[Ano82d] Anonymous. Author index to volume 25. Computer Physics Communications, 25(4):439–441, April 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900819.

Anonymous:1982:AIVb

[Ano82e] Anonymous. Author index to volume 26. Computer Physics Communications, 26(3–4):495–500, June 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901503.

Anonymous:1982:AIVc

[Ano82f] Anonymous. Author index to volume 27. Computer Physics Communications, 27(4):426–429, October 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901047.

Anonymous:1982:Ca

[Ano82g] Anonymous. Committees. Computer Physics Communications, 26(1–2):xii, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901539.

Anonymous:1982:Cc

[Ano82h] Anonymous. Committees. Computer Physics Communications, 26(1–2):xii, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901539.

Anonymous:1982:CPC

[Ano82i] Anonymous. Computer Physics Communications instructions to authors (fourth revision). Computer Physics Communications, 27(1):1–9, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900029.

Anonymous:1982:Cb

[Ano82j] Anonymous. Contents. Computer Physics Communications, 26(1–2):xiii, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901540.

Anonymous:1982:Cd

[Ano82k] Anonymous. Contents. Computer Physics Communications, 26(1–2):xiii, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901540.

Anonymous:1982:Ce

[Ano82l] Anonymous. Contents. Computer Physics Communications, 26(3–4):vii–viii, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901072.

Anonymous:1982:CVa

[Ano82m] Anonymous. Contents to volume 25. Computer Physics Communications, 25(4):437–438, April 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465582900807.

Anonymous:1982:CVb

[Ano82n] Anonymous. Contents to volume 26. Computer Physics Communications, 26(3–4):491–493, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465582901497.

Anonymous:1982:CVc

[Ano82o] Anonymous. Contents to volume 27. Computer Physics Communications, 27(4):423–425, October 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465582901035.

Anonymous:1982:EBaa

[Ano82p] Anonymous. Editorial board. Computer Physics Communications, 25(1):??, January 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900376.

Anonymous:1982:EBb

[Ano82q] Anonymous. Editorial board. Computer Physics Communications, 26(1–2):v–vii, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901515.

Anonymous:1982:EBc

[Ano82r] Anonymous. Editorial board. Computer Physics Communications, 27(1):v–vii, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900017.

Anonymous:1982:EBd

[Ano82s] Anonymous. Editorial board. Computer Physics Communications, 28(1):??, November 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900583.

Anonymous:1982:ENa

[Ano82t] Anonymous. Erratum notice. Computer Physics Communications, 25(1):111, January 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900492.

Anonymous:1982:ENb

[Ano82u] Anonymous. Erratum notice. Computer Physics Communications, 25(2):207, February 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900364.

Anonymous:1982:ENc

[Ano82v] Anonymous. Erratum notice. Computer Physics Communications, 28(2):217, December 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900571.

Anonymous:1982:FGC

[Ano82w] Anonymous. Fifth GAMM conference on numerical methods in fluid mechanics. Computer Physics Communications, 27(3):324, September 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901801.

Anonymous:1982:LC

[Ano82x] Anonymous. List of contributors. Computer Physics Communications, 26(3–4):489, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901485.

Anonymous:1982:LD

[Ano82y] Anonymous. List of delegates. Computer Physics Communications, 26(3–4):481–488, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901473.

Anonymous:1982:Pa

[Ano82z] Anonymous. Preface. Computer Physics Communications, 26(1–2):xi, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901527.

Anonymous:1982:Pb

[Ano82-27] Anonymous. Preface. Computer Physics Communications, 26(3–4):v–vi, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901060.

Anonymous:1982:PIVa

[Ano82-28] Anonymous. Program index to volume 25. Computer Physics Communications, 25(4):443–446, April 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900820.

Anonymous:1982:PIVc

[Ano82-29] Anonymous. Program index to volume 27. Computer Physics Communications, 27(4):430–432, October 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901059.

Anonymous:1982:PIVb

[Ano82-30] Anonymous. Program index to volumes 21-25. Computer Physics Communications, 25(4):447–460, April 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900832.

Anonymous:1983:A

[Ano83a] Anonymous. Announcement. Computer Physics Communications, 30(4):429, December 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558390084X.

Anonymous:1983:AIVa

[Ano83b] Anonymous. Author index to volume 28. Computer Physics Communications, 28(4):435–437, February 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900383.

Anonymous:1983:AIVb

[Ano83c] Anonymous. Author index to volume 29. Computer Physics Communications, 29(4):431–436, June 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900206.

Anonymous:1983:AIVc

[Ano83d] Anonymous. Author index to volume 30. Computer Physics Communications, 30(4):434–437, December 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900863.

Anonymous:1983:CVa

[Ano83e] Anonymous. Contents to volume 28. Computer Physics Communications, 28(4):433–434, February 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.com/science/article/pii/0010465583900371.

Anonymous:1983:CVb

[Ano83f] Anonymous. Contents to volume 29. Computer Physics Communications, 29(4):427–429, June 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/001046558390019X.

Anonymous:1983:CVc

[Ano83g] Anonymous. Contents to volume 30. Computer Physics Communications, 30(4):431–433, December 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900851.

Anonymous:1983:E

[Ano83h] Anonymous. Editorial. Computer Physics Communications, 30(2):??, September/October 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900528.

Anonymous:1983:EBa

[Ano83i] Anonymous. Editorial board. Computer Physics Communications, 29(1):??, March 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558390022X.

Anonymous:1983:EBb

[Ano83j] Anonymous. Editorial board. Computer Physics Communications, 30(1):i–iii, July/August 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583901157.

Anonymous:1983:ENa

[Ano83k] Anonymous. Erratum notice. Computer Physics Communications, 28(3):323, January 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900504.

Anonymous:1983:ENb

[Ano83l] Anonymous. Erratum notice. Computer Physics Communications, 28(3):324, January 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900516.

Anonymous:1983:ENc

[Ano83m] Anonymous. Erratum notice. Computer Physics Communications, 29(4):417–426, June 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900188.

Anonymous:1983:ENd

[Ano83n] Anonymous. Erratum notice. Computer Physics Communications, 30(1):107–108, July/August 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465583901285.

Anonymous:1983:ENe

[Ano83o] Anonymous. Erratum notice. Computer Physics Communications, 30(2):219, September/October 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900644.

Anonymous:1983:PIVa

[Ano83p] Anonymous. Program index to volume 28. Computer Physics Communications, 28(4):438–440, February 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900395.

Anonymous:1983:PIVb

[Ano83q] Anonymous. Program index to volume 29. Computer Physics Communications, 29(4):437–440, June 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900218.

Anonymous:1983:PIVc

[Ano83r] Anonymous. Program index to volume 30. Computer Physics Communications, 30(4):438–440, December 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900875.

Anonymous:1984:A

[Ano84a] Anonymous. Announcement. Computer Physics Communications, 32(2):229, May 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900754.

Anonymous:1984:ACP

[Ano84b] Anonymous. Announcement and call for papers. Computer Physics Communications, 31(4):441–442, March 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900328.

Anonymous:1984:AI

[Ano84c] Anonymous. Author index. Computer Physics Communications, 35(1–3):D–1–D–93, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584830271.

Anonymous:1984:AIVa

[Ano84d] Anonymous. Author index to volume 31. Computer Physics Communications, 31(4):445–447, March 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900341.

Anonymous:1984:AIVb

[Ano84e] Anonymous. Author index to volume 32. Computer Physics Communications, 32(4):443–445, July 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900626.

Anonymous:1984:AIVc

[Ano84f] Anonymous. Author index to volume 33. Computer Physics Communications, 33(4):447–449, October 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901504.

Anonymous:1984:Ca

[Ano84g] Anonymous. Contents. Computer Physics Communications, 31(2–3):vii, February 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900377.

Anonymous:1984:Cb

[Ano84h] Anonymous. Contents. Computer Physics Communications, 33(1–3):xiii–xiv, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900997.

Anonymous:1984:CLM

[Ano84i] Anonymous. Contents and layout of master index (volumes 1– 30). Computer Physics Communications, 35(1–3):v–viii, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822705.

Anonymous:1984:CVa

[Ano84j] Anonymous. Contents to volume 31. Computer Physics Communications, 31(4):443–444, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/001046558490033X.

Anonymous:1984:CVb

[Ano84k] Anonymous. Contents to volume 32. Computer Physics Communications, 32(4):441–442, July 1984. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465584900614.

Anonymous:1984:CVc

[Ano84l] Anonymous. Contents to volume 33. Computer Physics Communications, 33(4):445–446, October 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465584901498.

Anonymous:1984:EBa

[Ano84m] Anonymous. Editorial board. Computer Physics Communications, 31(1):v–vii, January 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900766.

Anonymous:1984:EBb

[Ano84n] Anonymous. Editorial board. Computer Physics Communications, 32(1):v–vii, April 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900018.

Anonymous:1984:EBc

[Ano84o] Anonymous. Editorial board. Computer Physics Communications, 33(1–3):v–vii, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900961.

Anonymous:1984:EBd

[Ano84p] Anonymous. Editorial board. Computer Physics Communications, 34(1–2):v–vii, November/December 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901528.

Anonymous:1984:Ea

[Ano84q] Anonymous. Erratum. Computer Physics Communications, 35(1–3):C–27, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482295X.

Anonymous:1984:Eb

[Ano84r] Anonymous. Erratum. Computer Physics Communications, 35(1–3):C–57, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823231.

Anonymous:1984:Ec

[Ano84s] Anonymous. Erratum. Computer Physics Communications, 35(1–3):C–165, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824169.

Anonymous:1984:Ed

[Ano84t] Anonymous. Erratum. Computer Physics Communications, 35(1–3):C–184, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824315.

Anonymous:1984:Ee

[Ano84u] Anonymous. Erratum. Computer Physics Communications, 35(1–3):C–211, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824546.

Anonymous:1984:Ef

[Ano84v] Anonymous. Erratum. Computer Physics Communications, 35(1–3):C–234, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482473X.

Anonymous:1984:Eg

[Ano84w] Anonymous. Erratum. Computer Physics Communications, 35(1–3):C–355, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825758.

Anonymous:1984:ENa

[Ano84x] Anonymous. Erratum notice. Computer Physics Communications, 31(4):433–434, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900286.

Anonymous:1984:ENb

[Ano84y] Anonymous. Erratum notice. Computer Physics Communications, 31(4):435–436, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900298.

Anonymous:1984:ENc

[Ano84z] Anonymous. Erratum notice. Computer Physics Communications, 31(4):437–438, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900304.

Anonymous:1984:ENd

[Ano84-27] Anonymous. Erratum notice. Computer Physics Communications, 31(4):439–440, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900316.

Anonymous:1984:ENe

[Ano84-28] Anonymous. Erratum notice. Computer Physics Communications, 34(1–2):223, November/December 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901747.

Anonymous:1984:ENf

[Ano84-29] Anonymous. Erratum notice. Computer Physics Communications, 34(1–2):224, November/December 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901759.

Anonymous:1984:ENaa

[Ano84-30] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–114, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823723.

Anonymous:1984:ENac 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824303.

Anonymous:1984:ENad

[Ano84-32] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–27, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584822948.

Anonymous:1984:ENaf

[Ano84-33] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–208, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824509.

Anonymous:1984:ENag

[Ano84-34] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–56, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482322X.

Anonymous:1984:ENah

[Ano84-35] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–242, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824807.

Anonymous:1984:ENaj

[Ano84-36] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–269, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825047.

Anonymous:1984:ENak

[Ano84-37] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–282, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825138.

Anonymous:1984:ENal 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825278.

Anonymous:1984:ENam

[Ano84-39] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–308, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825357.

Anonymous:1984:ENan

[Ano84-40] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–328, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825527.

Anonymous:1984:ENao

[Ano84-41] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–343, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825655.

Anonymous:1984:ENap

[Ano84-42] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–355, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825746.

Anonymous:1984:ENaq

[Ano84-43] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–392, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482603X.

Anonymous:1984:ENar

[Ano84-44] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–408, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826181.

Anonymous:1984:ENas 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826259.

Anonymous:1984:ENat

[Ano84-46] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–416, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826260.

Anonymous:1984:ENau

[Ano84-47] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–421, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826302.

Anonymous:1984:ENav

[Ano84-48] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–426, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826363.

Anonymous:1984:ENaw

[Ano84-49] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–431, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826417.

Anonymous:1984:ENax

[Ano84-50] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–445, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826521.

Anonymous:1984:ENay

[Ano84-51] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–472, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826715.

Anonymous:1984:ENaz 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826995.

Anonymous:1984:ENg

[Ano84-53] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–27, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584822948.

Anonymous:1984:ENh

[Ano84-54] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–571, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827459.

Anonymous:1984:ENi

[Ano84-55] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–571, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827460.

Anonymous:1984:ENj

[Ano84-56] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–607, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827745.

Anonymous:1984:ENk

[Ano84-57] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–621, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482785X.

Anonymous:1984:ENl

[Ano84-58] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–621, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827861.

Anonymous:1984:ENm 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827915.

Anonymous:1984:ENn

[Ano84-60] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–649, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482809X.

Anonymous:1984:ENo

[Ano84-61] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–679, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828325.

Anonymous:1984:ENp

[Ano84-62] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–689, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828404.

Anonymous:1984:ENq

[Ano84-63] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–710, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828544.

Anonymous:1984:ENr

[Ano84-64] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–722, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828623.

Anonymous:1984:ENs

[Ano84-65] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–774, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482901X.

Anonymous:1984:ENt 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829094.

Anonymous:1984:ENu

[Ano84-67] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–847, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829562.

Anonymous:1984:ENv

[Ano84-68] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–855, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829641.

Anonymous:1984:ENw

[Ano84-69] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–890, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482994X.

Anonymous:1984:ENx

[Ano84-70] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–912, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584830118.

Anonymous:1984:ENy

[Ano84-71] Anonymous. Erratum notice. Computer Physics Communications, 35(1–3):C–56, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482322X.

Anonymous:1984:ENab

[Ano84-72] Anonymous. Erratum notices. Computer Physics Communications, 35(1–3):C–141, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823954.

Anonymous:1984:ENae

[Ano84-73] Anonymous. Erratum notices. Computer Physics Communications, 35(1–3):C–193, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824388.

Anonymous:1984:ENai

[Ano84-74] Anonymous. Erratum notices. Computer Physics Communications, 35(1–3):C–78, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823413.

Anonymous:1984:ENz

[Ano84-75] Anonymous. Erratum notices. Computer Physics Communications, 35(1–3):C–78, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823413.

Anonymous:1984:Pa

[Ano84-76] Anonymous. Preface. Computer Physics Communications, 33 (1–3):xi, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900973.

Anonymous:1984:Pb

[Ano84-77] Anonymous. Preface. Computer Physics Communications, 35(1–3):??, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584822699.

Anonymous:1984:PI

[Ano84-78] Anonymous. Program index. Computer Physics Communications, 35(1–3):??, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584822729.

Anonymous:1984:PIVa

[Ano84-79] Anonymous. Program index to volume 31. Computer Physics Communications, 31(4):449–450, March 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900353.

Anonymous:1984:PIVb

[Ano84-80] Anonymous. Program index to volume 32. Computer Physics Communications, 32(4):446–448, July 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900638.

Anonymous:1984:PIVc

[Ano84-81] Anonymous. Program index to volume 33. Computer Physics Communications, 33(4):450, October 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901516.

Anonymous:1984:PS

[Ano84-82] Anonymous. Program summary. Computer Physics Communications, 35(1–3):C–9, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584822791.

Anonymous:1984:PSC

[Ano84-83] Anonymous. Program summary (CsCI lattice). Computer Physics Communications, 35(1–3):C–34, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823024.

Anonymous:1984:PSD

[Ano84-84] Anonymous. Program summary (DCSCH4). Computer Physics Communications, 35(1–3):C–95, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823541.

Anonymous:1984:PSF

[Ano84-85] Anonymous. Program summary (field gradient integral). Computer Physics Communications, 35(1–3):C–132, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823863.

Anonymous:1984:SOC

[Ano84-86] Anonymous. Sponsors and organizing committee. Computer Physics Communications, 33(1–3):xii, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900985.

Anonymous:1984:SIC

[Ano84-87] Anonymous. Subject index computational physics papers. Computer Physics Communications, 35(1–3):??, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822717.

Anonymous:1985:A

[Ano85a] Anonymous. Announcement. Computer Physics Communications, 38(3):449, December 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901134.

Anonymous:1985:AIVa

[Ano85b] Anonymous. Author index to volume 34. Computer Physics Communications, 34(4):442–445, February 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900748.

Anonymous:1985:AIVb

[Ano85c] Anonymous. Author index to volume 36. Computer Physics Communications, 36(4):446–450, June 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900360.

Anonymous:1985:AIVc

[Ano85d] Anonymous. Author index to volume 37. Computer Physics Communications, 37(1–3):383–386, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558590178X.

Anonymous:1985:AIVd

[Ano85e] Anonymous. Author index to volume 38. Computer Physics Communications, 38(3):454–458, December 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901158.

Anonymous:1985:Ca

[Ano85f] Anonymous. Contents. Computer Physics Communications, 37(1–3):xiii–xv, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901298.

Anonymous:1985:Cb

[Ano85g] Anonymous. Contents. Computer Physics Communications, 38 (2):vii–viii, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900785.

Anonymous:1985:CVa

[Ano85h] Anonymous. Contents to volume 34. Computer Physics Communications, 34(4):439–441, February 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.com/science/article/pii/0010465585900736.

Anonymous:1985:CVb

[Ano85i] Anonymous. Contents to volume 36. Computer Physics Communications, 36(4):443–445, June 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465585900359.

Anonymous:1985:CVc

[Ano85j] Anonymous. Contents to volume 37. Computer Physics Communications, 37(1–3):380–382, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465585901778.

Anonymous:1985:CVd

[Ano85k] Anonymous. Contents to volume 38. Computer Physics Communications, 38(3):451–453, December 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901146.

Anonymous:1985:EBa

[Ano85l] Anonymous. Editorial board. Computer Physics Communications, 36(1):v–vii, March 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900128.

Anonymous:1985:EBb

[Ano85m] Anonymous. Editorial board. Computer Physics Communications, 37(1–3):v–vii, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901274.

Anonymous:1985:EBc

[Ano85n] Anonymous. Editorial board. Computer Physics Communications, 38(1):v–vii, August/September 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900384.

Anonymous:1985:ENa

[Ano85o] Anonymous. Erratum notice. Computer Physics Communications, 34(4):427–429, February 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900700.

Anonymous:1985:ENb

[Ano85p] Anonymous. Erratum notice. Computer Physics Communications, 34(4):431–436, February 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900712.

Anonymous:1985:ENc

[Ano85q] Anonymous. Erratum notice. Computer Physics Communications, 34(4):437–438, February 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900724.

Anonymous:1985:ENd

[Ano85r] Anonymous. Erratum notice. Computer Physics Communications, 36(3):337, May 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900608.

Anonymous:1985:ENe

[Ano85s] Anonymous. Erratum notice. Computer Physics Communications, 36(4):441, June 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900347.

Anonymous:1985:ENf

[Ano85t] Anonymous. Erratum notice. Computer Physics Communications, 38(1):118, August/September 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900529.

Anonymous:1985:LC

[Ano85u] Anonymous. List of contributors. Computer Physics Communications, 37(1–3):379, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901766.

Anonymous:1985:LD

[Ano85v] Anonymous. List of delegates. Computer Physics Communications, 37(1–3):371–378, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901754.

Anonymous:1985:O

[Ano85w] Anonymous. Obituary. Computer Physics Communications, 38(3):i–ii, December 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900992.

Anonymous:1985:Pa

[Ano85x] Anonymous. Preface. Computer Physics Communications, 37 (1–3):xi, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585901286.

Anonymous:1985:Pb

[Ano85y] Anonymous. Preface. Computer Physics Communications, 38(2):v, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900761.

Anonymous:1985:PIVa

[Ano85z] Anonymous. Program index to volume 34. Computer Physics Communications, 34(4):446–448, February 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558590075X.

Anonymous:1985:PIVb

[Ano85-27] Anonymous. Program index to volume 36. Computer Physics Communications, 36(4):451–452, June 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900372.

Anonymous:1985:PIVc

[Ano85-28] Anonymous. Program index to volume 38. Computer Physics Communications, 38(3):459–461, December 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558590116X.

Anonymous:1985:S

[Ano85-29] Anonymous. Sponsors. Computer Physics Communications, 38(2):vi, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900773.

Anonymous:1986:AIVa

[Ano86a] Anonymous. Author index to volume 39. Computer Physics Communications, 39(3):452–456, April 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901037.

Anonymous:1986:AIVb

[Ano86b] Anonymous. Author index to volume 40. Computer Physics Communications, 40(2–3):443–447, June 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901244.

Anonymous:1986:AIVc

[Ano86c] Anonymous. Author index to volume 41. Computer Physics Communications, 41(2–3):430, August 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558690086X.

Anonymous:1986:AIVd

[Ano86d] Anonymous. Author index to volume 42. Computer Physics Communications, 42(3):446–448, November 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900147.

Anonymous:1986:Cc

[Ano86e] Anonymous. Committees. Computer Physics Communications, 43(1):xii, December 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900482.

Anonymous:1986:CPC

[Ano86f] Anonymous. Computer Physics Communications instructions to authors (fifth revision). Computer Physics Communications, 42(1):xv–xxiii, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586902213.

Anonymous:1986:Ca

[Ano86g] Anonymous. Contents. Computer Physics Communications, 40 (1):xiii, May 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558690144X.

Anonymous:1986:Cb

[Ano86h] Anonymous. Contents. Computer Physics Communications, 41(2–3):vii–viii, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900639.

Anonymous:1986:Cd

[Ano86i] Anonymous. Contents. Computer Physics Communications, 43(1):xiii, December 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900494.

Anonymous:1986:CVa

[Ano86j] Anonymous. Contents to volume 39. Computer Physics Communications, 39(3):449–451, April 1986. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465586901025.

Anonymous:1986:CVb

[Ano86k] Anonymous. Contents to volume 40. Computer Physics Communications, 40(2–3):441–442, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465586901232.

Anonymous:1986:CVc

[Ano86l] Anonymous. Contents to volume 41. Computer Physics Communications, 41(2–3):427–429, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.com/science/article/pii/0010465586900858.

Anonymous:1986:CVd

[Ano86m] Anonymous. Contents to volume 42. Computer Physics Communications, 42(3):443–445, November 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900135.

Anonymous:1986:CPL

[Ano86n] Anonymous. CPC program library. Computer Physics Communications, 42(1):xxv–xxvi, September 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586902225.

Anonymous:1986:E

[Ano86o] Anonymous. Editorial. Computer Physics Communications, 42(1):xi–xiii, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586902201.

Anonymous:1986:EBa

[Ano86p] Anonymous. Editorial board. Computer Physics Communications, 39(1):v–vii, January 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558690158X.

Anonymous:1986:EBb

[Ano86q] Anonymous. Editorial board. Computer Physics Communications, 40(1):v–vii, May 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901426.

Anonymous:1986:EBc

[Ano86r] Anonymous. Editorial board. Computer Physics Communications, 41(1):v–vii, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900160.

Anonymous:1986:EBd

[Ano86s] Anonymous. Editorial board. Computer Physics Communications, 42(1):v–vii, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586902183.

Anonymous:1986:EBe

[Ano86t] Anonymous. Editorial board. Computer Physics Communications, 43(1):v–vii, December 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900469.

Anonymous:1986:ENa

[Ano86u] Anonymous. Erratum notice. Computer Physics Communications, 39(1):153, January 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901700.

Anonymous:1986:ENb

[Ano86v] Anonymous. Erratum notice. Computer Physics Communications, 39(1):154, January 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901712.

Anonymous:1986:ENc

[Ano86w] Anonymous. Erratum notice. Computer Physics Communications, 39(2):303, February/March 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465586901414.

Anonymous:1986:ENd

[Ano86x] Anonymous. Erratum notice. Computer Physics Communications, 39(3):447, April 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901013.

Anonymous:1986:ENe

[Ano86y] Anonymous. Erratum notice. Computer Physics Communications, 41(1):195, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900317.

Anonymous:1986:ENf

[Ano86z] Anonymous. Erratum notice. Computer Physics Communications, 41(1):196, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900329.

Anonymous:1986:ENg

[Ano86-27] Anonymous. Erratum notice. Computer Physics Communications, 42(2):303, October 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900457.

Anonymous:1986:LP

[Ano86-28] Anonymous. List of participants. Computer Physics Communications, 43(1):169–170, December 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900603.

Anonymous:1986:Pb

[Ano86-29] Anonymous. Preface. Computer Physics Communications, 41(2–3):v, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900615.

Anonymous:1986:Pa

[Ano86-30] Anonymous. Preface 0. Computer Physics Communications, 40 (1):xi, May 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586901438.

Anonymous:1986:PIVa

[Ano86-31] Anonymous. Program index to volume 39. Computer Physics Communications, 39(3):457–460, April 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901049.

Anonymous:1986:PIVb

[Ano86-32] Anonymous. Program index to volume 40. Computer Physics Communications, 40(2–3):448, June 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901256.

Anonymous:1986:PIVc

[Ano86-33] Anonymous. Program index to volume 41. Computer Physics Communications, 41(2–3):433–434, August 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900871.

Anonymous:1986:PIVd

[Ano86-34] Anonymous. Program index to volume 42. Computer Physics Communications, 42(3):449–451, November 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900159.

Anonymous:1986:PN

[Ano86-35] Anonymous. Publisher’s note. Computer Physics Communications, 42(1):ix, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586902195.

Anonymous:1986:S

[Ano86-36] Anonymous. Sponsors. Computer Physics Communications, 41(2–3):vi, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900627.

Anonymous:1987:A

[Ano87a] Anonymous. Announcement. Computer Physics Communications, 46(2):336, August 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900099.

Anonymous:1987:ACA

[Ano87b] Anonymous. The APE computer: An array processor optimized for lattice gauge theory simulations. Computer Physics Communications, 45(1–3):345–353, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558790172X.

Anonymous:1987:AGG

[Ano87c] Anonymous. An application of the GEANT3 geometry package to the description of the opal detector. Computer Physics Communications, 47(1):55–74, October 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558790066X.

Anonymous:1987:ACC

[Ano87d] Anonymous. ASILOMAR CHEP conference participants. Computer Physics Communications, 45(1–3):487–490, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901998.

Anonymous:1987:AIVa

[Ano87e] Anonymous. Author index to volume 43. Computer Physics Communications, 43(3):442–444, February/March 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900609.

Anonymous:1987:AIVb

[Ano87f] Anonymous. Author index to volume 44. Computer Physics Communications, 44(3):311–313, June 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900889.

Anonymous:1987:AIVc

[Ano87g] Anonymous. Author index to volume 45. Computer Physics Communications, 45(1–3):495–499, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587902013.

Anonymous:1987:AIVd

[Ano87h] Anonymous. Author index to volume 46. Computer Physics Communications, 46(3):458–460, September 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901019.

Anonymous:1987:AIVe

[Ano87i] Anonymous. Author index to volume 47. Computer Physics Communications, 47(2–3):377–380, November/December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901275.

Anonymous:1987:CI

[Ano87j] Anonymous. Colour illustrations. Computer Physics Communications, 44(1–2):ix, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900117.

Anonymous:1987:C

[Ano87k] Anonymous. Contents. Computer Physics Communications, 44 (3):vii, June 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900774.

Anonymous:1987:CVa

[Ano87l] Anonymous. Contents to volume 43. Computer Physics Communications, 43(3):439–441, February/March 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900592.

Anonymous:1987:CVb

[Ano87m] Anonymous. Contents to volume 44. Computer Physics Communications, 44(3):309–310, June 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465587900877.

Anonymous:1987:CVc

[Ano87n] Anonymous. Contents to volume 45. Computer Physics Communications, 45(1–3):491–494, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587902001.

Anonymous:1987:CVd

[Ano87o] Anonymous. Contents to volume 46. Computer Physics Communications, 46(3):455–457, September 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901007.

Anonymous:1987:CVe

[Ano87p] Anonymous. Contents to volume 47. Computer Physics Communications, 47(2–3):373–375, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901263.

Anonymous:1987:EBa

[Ano87q] Anonymous. Editorial board. Computer Physics Communications, 44(1–2):v–vii, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900105.

Anonymous:1987:EBb

[Ano87r] Anonymous. Editorial board. Computer Physics Communications, 45(1–3):v–vii, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587901330.

Anonymous:1987:EBc

[Ano87s] Anonymous. Editorial board. Computer Physics Communications, 46(1):v–vii, July 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900312.

Anonymous:1987:EBd

[Ano87t] Anonymous. Editorial board. Computer Physics Communications, 47(1):v–vii, October 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900622.

Anonymous:1987:ENa

[Ano87u] Anonymous. Erratum notice. Computer Physics Communications, 44(1–2):227–231, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465587900300.

Anonymous:1987:ENb

[Ano87v] Anonymous. Erratum notice. Computer Physics Communications, 47(2–3):367, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901238.

Anonymous:1987:ENc

[Ano87w] Anonymous. Erratum notice. Computer Physics Communications, 47(2–3):368, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558790124X.

Anonymous:1987:PIVa

[Ano87x] Anonymous. Program index to volume 43. Computer Physics Communications, 43(3):445–446, February/March 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900610.

Anonymous:1987:PIVb

[Ano87y] Anonymous. Program index to volume 44. Computer Physics Communications, 44(3):315–316, June 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900890.

Anonymous:1987:PIVc

[Ano87z] Anonymous. Program index to volume 46. Computer Physics Communications, 46(3):461–463, September 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901020.

Anonymous:1987:PIVd

[Ano87-27] Anonymous. Program index to volume 47. Computer Physics Communications, 47(2–3):381–382, November/December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901287.

Anonymous:1987:SIV

[Ano87-28] Anonymous. Subject index to volume 45. Computer Physics Communications, 45(1–3):500–503, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587902025.

Anonymous:1987:USS

[Ano87-29] Anonymous. The use of SA/SD methods in D0 software development. Computer Physics Communications, 45(1–3):245–257, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901615.

Anonymous:1988:Aa

[Ano88a] Anonymous. Announcement. Computer Physics Communications, 48(1):??, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588900379.

Anonymous:1988:Ab

[Ano88b] Anonymous. Announcement. Computer Physics Communications, 50(3):415–416, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901968.

Anonymous:1988:AIVa

[Ano88c] Anonymous. Author index to volume 48. Computer Physics Communications, 48(3):462–464, March 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902068.

Anonymous:1988:AIVb

[Ano88d] Anonymous. Author index to volume 49. Computer Physics Communications, 49(3):511–513, June 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900100.

Anonymous:1988:AIVc

[Ano88e] Anonymous. Author index to volume 50. Computer Physics Communications, 50(3):420–422, August 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901981.

Anonymous:1988:AIVd

[Ano88f] Anonymous. Author index to volume 51. Computer Physics Communications, 51(3):480–483, November 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901610.

Anonymous:1988:Cc

[Ano88g] Anonymous. Committees. Computer Physics Communications, 50(1–2):xii, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901117.

Anonymous:1988:CPCa

[Ano88h] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 49(1):v–vii, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902081.

Anonymous:1988:CPCb

[Ano88i] Anonymous. Computer Physics Communications — publisher’s note. Computer Physics Communications, 50(3):??, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901828.

Anonymous:1988:Ca

[Ano88j] Anonymous. Contents. Computer Physics Communications, 48(1):xiii–xiv, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588900161.

Anonymous:1988:Cb

[Ano88k] Anonymous. Contents. Computer Physics Communications, 49(1):xiii–xiv, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588902111.

Anonymous:1988:Cd

[Ano88l] Anonymous. Contents. Computer Physics Communications, 51(1–2):xiii–xiv, September/October 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900562.

Anonymous:1988:CVa

[Ano88m] Anonymous. Contents to volume 48. Computer Physics Communications, 48(3):459–461, March 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465588902056.

Anonymous:1988:CVb

[Ano88n] Anonymous. Contents to volume 49. Computer Physics Communications, 49(3):507–509, June 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465588900094.

Anonymous:1988:CVc

[Ano88o] Anonymous. Contents to volume 50. Computer Physics Communications, 50(3):417–419, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/001046558890197X.

Anonymous:1988:CVd

[Ano88p] Anonymous. Contents to volume 51. Computer Physics Communications, 51(3):477–479, November 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901609.

Anonymous:1988:EBa

[Ano88q] Anonymous. Editorial board. Computer Physics Communications, 48(1):v–vii, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588900148.

Anonymous:1988:EBb

[Ano88r] Anonymous. Editorial board. Computer Physics Communications, 50(1–2):v–vii, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901099.

Anonymous:1988:EBc

[Ano88s] Anonymous. Editorial board. Computer Physics Communications, 51(1–2):v–vii, September/October 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900537.

Anonymous:1988:EBd

[Ano88t] Anonymous. Editorial board 1425 1101 V 2. Computer Physics Communications, 52(1):v–vii, December 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901634.

Anonymous:1988:EN

[Ano88u] Anonymous. Erratum notice. Computer Physics Communications, 52(1):165, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901816.

Anonymous:1988:F

[Ano88v] Anonymous. Forward. Computer Physics Communications, 51 (1–2):xi, September/October 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588900549.

Anonymous:1988:LCb

[Ano88w] Anonymous. List of committees. Computer Physics Communications, 49(1):xii, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558890210X.

Anonymous:1988:LCa

[Ano88x] Anonymous. List of contributors. Computer Physics Communications, 48(1):174, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588900367.

Anonymous:1988:LCc

[Ano88y] Anonymous. List of contributors. Computer Physics Communications, 50(1–2):285, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901397.

Anonymous:1988:LP

[Ano88z] Anonymous. List of participants. Computer Physics Communications, 49(1):271–274, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588902366.

Anonymous:1988:MEP

[Ano88-27] Anonymous. Microcomputers in everyday practice of physicists: (a summary of the 7th European Summer School on Computing Techniques in Physics). Computer Physics Communications, 50(1–2):281–283, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901385.

Anonymous:1988:Pa

[Ano88-28] Anonymous. Preface. Computer Physics Communications, 49 (1):xi, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588902093.

Anonymous:1988:Pb

[Ano88-29] Anonymous. Preface. Computer Physics Communications, 50 (1–2):xi, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901105.

Anonymous:1988:Pc

[Ano88-30] Anonymous. Preface. Computer Physics Communications, 51 (1–2):xii, September/October 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588900550.

Anonymous:1988:PIVa

[Ano88-31] Anonymous. Program index to volume 48. Computer Physics Communications, 48(3):465–466, March 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558890207X.

Anonymous:1988:PIVb

[Ano88-32] Anonymous. Program index to volume 49. Computer Physics Communications, 49(3):514, June 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465588900112.

Anonymous:1988:PIVc

[Ano88-33] Anonymous. Program index to volume 50. Computer Physics Communications, 50(3):423, August 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901993.

Anonymous:1988:PIVd

[Ano88-34] Anonymous. Program index to volume 51. Computer Physics Communications, 51(3):484–486, November 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901622.

Anonymous:1988:RTD

[Ano88-35] Anonymous. Round-table discussion. Computer Physics Communications, 48(1):167–173, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465588900355.

Anonymous:1988:SI

[Ano88-36] Anonymous. Subject index. Computer Physics Communications, 50(1–2):287–288, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901403.

Anonymous:1989:WCM

[Ano89a] Anonymous. 3rd world conference on mathematics at the service of man. Computer Physics Communications, 52(3):446, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901203.

Anonymous:1989:AI

[Ano89b] Anonymous. Author index. Computer Physics Communications, 56(1):104, November 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558990057X.

Anonymous:1989:AIVa

[Ano89c] Anonymous. Author index to volume 52. Computer Physics Communications, 52(3):450–453, March 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901227.

Anonymous:1989:AIVb

[Ano89d] Anonymous. Author index to volume 53. Computer Physics Communications, 53(1–3):479–480, May 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990180X.

Anonymous:1989:AIVc

[Ano89e] Anonymous. Author index to volume 54. Computer Physics Communications, 54(2–3):416–418, June/July 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901021.

Anonymous:1989:AIVd

[Ano89f] Anonymous. Author index to volume 55. Computer Physics Communications, 55(3):474–477, October 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901410.

Anonymous:1989:AIVe

[Ano89g] Anonymous. Author index to volume 57. Computer Physics Communications, 57(1–3):568–576, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902956.

Anonymous:1989:BRB

[Ano89h] Anonymous. Book review: Computational Techniques for Fluid Dynamics: C. A. J. Fletcher, Springer Series in Computational Physics, Springer-Verlag, London, 1988. Two volumes; 409 and 484 pages, respectively. Hardcover price for two-volume set DM 198.00. ISBN 0-387-19466-5 and 3-540-19466-5. Computer Physics Communications, 56(2):291–292, December 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900258.

Anonymous:1989:CS

[Ano89i] Anonymous. Committees and sponsors. Computer Physics Communications, 57(1–3):xiv, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901835.

Anonymous:1989:Ca

[Ano89j] Anonymous. Contents. Computer Physics Communications, 56(1):xiii, November 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900489.

Anonymous:1989:Cb

[Ano89k] Anonymous. Contents. Computer Physics Communications, 57(1–3):xv–xx, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901847.

Anonymous:1989:CVa

[Ano89l] Anonymous. Contents to volume 52. Computer Physics Communications, 52(3):447–449, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465589901215.

Anonymous:1989:CVb

[Ano89m] Anonymous. Contents to volume 54. Computer Physics Communications, 54(2–3):413–415, June/July 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990101X.

Anonymous:1989:CVc

[Ano89n] Anonymous. Contents to volume 55. Computer Physics Communications, 55(3):471–473, October 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465589901409.

Anonymous:1989:EBa

[Ano89o] Anonymous. Editorial board. Computer Physics Communications, 53(1–3):v–vii, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901434.

Anonymous:1989:EBb

[Ano89p] Anonymous. Editorial board. Computer Physics Communications, 54(1):v–vii, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558990026X.

Anonymous:1989:EBc

[Ano89q] Anonymous. Editorial board. Computer Physics Communications, 55(1):??, August 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900581.

Anonymous:1989:EBd

[Ano89r] Anonymous. Editorial board. Computer Physics Communications, 56(1):v–vii, November 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900465.

Anonymous:1989:EBe

[Ano89s] Anonymous. Editorial board. Computer Physics Communications, 57(1–3):??, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901811.

Anonymous:1989:ENa

[Ano89t] Anonymous. Erratum notice. Computer Physics Communications, 52(3):443–444, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901185.

Anonymous:1989:ENb

[Ano89u] Anonymous. Erratum notice. Computer Physics Communications, 52(3):445, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901197.

Anonymous:1989:ENc

[Ano89v] Anonymous. Erratum notice. Computer Physics Communications, 55(2):251, September 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900817.

Anonymous:1989:ENd

[Ano89w] Anonymous. Erratum notice. Computer Physics Communications, 55(3):469, October 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901392.

Anonymous:1989:AAG

[Ano89x] Anonymous. From APE to APE-100: From 1 to 100 Gflops in lattice gauge theory simulations. Computer Physics Communications, 57(1–3):285–289, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902294.

Anonymous:1989:KIV

[Ano89y] Anonymous. KWIC index to volume. Computer Physics Communications, 57(1–3):585–587, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990297X.

Anonymous:1989:LP

[Ano89z] Anonymous. List of participants. Computer Physics Communications, 56(1):101–103, November 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900568.

Anonymous:1989:OCC

[Ano89-27] Anonymous. Oxford CHEP conference participants. Computer Physics Communications, 57(1–3):563–567, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902944.

Anonymous:1989:P

[Ano89-28] Anonymous. Preface. Computer Physics Communications, 53(1–3):xi, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901446.

Anonymous:1989:PIVa

[Ano89-29] Anonymous. Program index to volume 52. Computer Physics Communications, 52(3):454–456, March 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901239.

Anonymous:1989:PIVb

[Ano89-30] Anonymous. Program index to volume 54. Computer Physics Communications, 54(2–3):419–422, June/July 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901033.

Anonymous:1989:PIVc

[Ano89-31] Anonymous. Program index to volume 55. Computer Physics Communications, 55(3):478–480, October 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901422.

Anonymous:1989:SIV

[Ano89-32] Anonymous. Subject index to volume 57. Computer Physics Communications, 57(1–3):577–584, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902968.

Angelini:1985:MCP

[ANP^{+}85] L. Angelini, L. Nitti, M. Pellicoro, G. Preparata, and G. Valenti. A Monte Carlo program for generating hadronic final states in electron-positron annihilations. Computer Physics Communications, 34(4):371–385, February 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900669.

Alvisi:1988:RBA

[AO88] L. Alvisi and R. Odorico. A rule based approach for pattern recognition in planar geometric figures. Computer Physics Communications, 51(3):443–450, November 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901579.

Agrawal:1980:NGG

[AP80] A. K. Agrawal and R. S. Peckover. Nonuniform grid generation for boundary-layer problems. Computer Physics Communications, 19(2):171–178, April/June 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900478.

Aers:1982:CIS

[AP82] G. C. Aers and J. B. Pendry. Calculation of the impact scattering contribution to electron energy loss spectra. Computer Physics Communications, 25(4):389–416, April 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900777.

Aers:1984:CIS

[AP84] G. C. Aers and J. B. Pendry. Calculation of the impact scattering contribution to electron energy loss spectra. Computer Physics Communications, 35(1–3):C–798, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829197.

Allouche:1989:CSF

[AP89] A. Allouche and G. Pouzard. Computer simulation of FT-NMR multiple pulse experiment. Computer Physics Communications, 54(1):171–176, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900428.

Arnoldus:1984:NSR

[Arn84a] H. F. Arnoldus. Numerical stabilization of recurrence relations with vanishing solutions. Computer Physics Communications, 33(4):347–352, October 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584901401.

Arnoldus:1984:REM

[Arn84b] H. F. Arnoldus. Radial electric multipole matrix elements for inelastic collisions in atomic and nuclear physics. Computer Physics Communications, 32(4):421–437, July 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900584.

Arter:1985:BRBa

[Art85a] W. Arter. Book review: Numerical Methods for Differential Equations and Applications: Liviu Ixaru, D. Reidel, Dordrecht, 1984. xxi + 337 pages Cloth Dfl 180, US$69 ISBN 90-277-1597-1. Computer Physics Communications, 34(3):335, January 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900116.

Arter:1985:BRBb

[Art85b] W. Arter. Book review: Potential Flows: Computer Graphic Solutions: Robert H. Kirchhoff, Marcel Dekker Inc., New York, 1985. xiv + 182 pages. US$45. Computer Physics Communications, 38(1):117, August/September 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900517.

Anderson:1983:ISSa

[AS83] D. V. Anderson and A. I. Shestakov. ICCG2: Subprograms for the solution of a linear symmetric matrix equation arising from a 9-point discretization. Computer Physics Communications, 30(1):37–42, July/August 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583901200.

Anderson:1984:ISSa

[AS84] D. V. Anderson and A. I. Shestakov. ICCG2: Subprograms for the solution of a linear symmetric matrix equation arising from a 9-point discretization. Computer Physics Communications, 35(1–3):C–893–C–894, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829963.

Amar:1989:MCS

[AS89] Jacques G. Amar and Francis Sullivan. Monte Carlo simulation of domain growth in the kinetic Ising model on the connection machine. Computer Physics Communications, 55(3):287–295, October 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901264.

Allen:1980:APY

[ASM80] P. D. Allen, Su Su, and E. G. Muirhead. Analysis of photonuclear yield curves by the variable bin Penfold–Leiss method. Computer Physics Communications, 21(2):163–184, December 2, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900387.

Allen:1984:APY

[ASM84] P. D. Allen, Su Su, and E. G. Muirhead. Analysis of photonuclear yield curves by the variable bin penfold-leiss method. Computer Physics Communications, 35(1–3):C–656– C–657, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828167.

Ahmad:1982:SMA

[ASS82] K. Ahmad, R. A. Scott, and M. Sur. Stochastic modelling on the AP120B: An application to water resources planning. Computer Physics Communications, 27(1):91–95, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900121.

Assimakopoulos:1984:KTB

[Ass84] P. A. Assimakopoulos. Kinematics of three-body reactions. Computer Physics Communications, 35(1–3):C–350, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825710.

Aston:1987:TPM

[Ast87] David Aston. Towards a personal mainframe. Computer Physics Communications, 45(1–3):191–194, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558790155X.

Abdallah:1984:CEI

[AT84] Joseph Abdallah, Jr. and Donald G. Truhlar. Continuum exchange integrals for algebraic variational calculations of electron-atom scattering using Slater-type orbitals as basis functions. Computer Physics Communications, 35(1–3):C–314– C–315, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825400.

Atakan:1985:DSD

[Ata85] Ahmet K. Atakan. Determination of the strengths of doubled lines by the method of equivalent widths. Computer Physics Communications, 36(1):9–17, March 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900153.

Atzeni:1986:DLS

[Atz86]  Stefano Atzeni.  2-D Lagrangian studies of symmetry

and stability of laser fusion targets. Computer Physics Communications, 43(1):107–124, December 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900561.

Auerbach:1984:TGO

[Aue84] E. H. Auerbach. A-three: a general optical model code especially suited to heavy-ion calculations. Computer Physics Communications, 35(1–3):C–494–C–495, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826831.

Aurela:1984:CSNa

[Aur84a] A. M. Aurela. Counting a small number of radioactive atoms. Computer Physics Communications, 35(1–3):C–430, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584826399.

Aurela:1984:CSNb

[Aur84b] A. M. Aurela. Counting a small number of radioactive atoms, second program. Computer Physics Communications, 35(1–3): C–555, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827307.

Akylas:1984:MAC

[AV84] V. R. Akylas and P. Vogel. Muonic atom cascade program. Computer Physics Communications, 35(1–3):C–506, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826910.

Avery:1989:UNP

[AW89] Paul Avery and Andrew White. UFMulti: a new parallel processing software system for HEP. Computer Physics Communications, 57(1–3):422–426, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902555.

Arnot:1982:AID

[AWB82] N. R. Arnot, G. G. Wilkinson, and R. E. Burge. Application of the ICL DAP for two-dimensional image processing. Computer Physics Communications, 26(3–4):455–457, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901424.

Atkinson:1984:CAA

[AWG84] G. D. Atkinson, Jr., J. B. Whitworth, and S. J. Gage. Computer-assisted analysis of gamma-ray spectra. Computer Physics Communications, 35(1–3):C–63, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823280.

Akkermans:1987:EHF

[AXG87] J. M. Akkermans, Shi Xiangjun, and H. Gruppelaar. Economizing Hauser–Feshbach model calculations in the continuum. Computer Physics Communications, 43(3):347–354, February/ March 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558790052X.

Ayrjan:1983:FRM

[AZK83] E. A. Ayrjan, E. P. Zhidkov, and B. N. Khoromsky. Fast relaxation method for solving the difference problem for the Poisson equation on a sequence of grids. Computer Physics Communications, 29(2):125–130, April 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900681.

Brode:1986:OMP

[BA86] Stefan Brode and Reinhart Ahlrichs. An optimized MD program for the vector computer Cyber 205. Computer Physics Communications, 42(1):51–57, September 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586902304.

Biel:1987:SAM

[BAA^{+}87] J. Biel, H. Areti, R. Atac, A. Cook, M. Fischler, I. Gaines, C. Kaliher, R. Hance, D. Husby, T. Nash, and T. Zmuda. Software for the ACP multiprocessor system. Computer Physics Communications, 45(1–3):331–337, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901706.

Beeler:1987:ESC

[BAG^{+}87] F. Beeler, O. K. Andersen, O. Gunnarsson, O. Jepsen, and M. Scheffler. Electronic-structure calculation of point defects in silicon. Computer Physics Communications, 44(3):297– 305, June 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900853.

Baillie:1986:GFC

[Bai86] C. F. Baillie. A general FORTRAN to C translator. Computer Physics Communications, 41(2–3):409–414, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900810.

Baig:1987:SLT

[Bai87] M. Baig. Simulation of the λϕ^{4} lattice theory on a loosely coupled array of processors. Computer Physics Communications, 47(2–3):181–186, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901044.

Barnett:1981:ARI

[Bar81a] A. R. Barnett. An algorithm for regular and irregular Coulomb and Bessel functions of real order to machine accuracy. Computer Physics Communications, 21(3):297–314, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Barnett:1981:KCF

[Bar81b] A. R. Barnett. Klein: Coulomb functions for real λ and positive energy to high accuracy. Computer Physics Communications, 24(2):141–159, November 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900886.

Barnett:1982:CCB

[Bar82] A. R. Barnett. COULFG: Coulomb and Bessel functions and their derivatives, for real arguments, by Steed’s method. Computer Physics Communications, 27(2):147–166, August 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900704.

Bartschat:1983:PCO

[Bar83] K. Bartschat. Program to calculate observable quantities from scattering amplitudes for inelastic electron-atom collisions. Computer Physics Communications, 30(4):383–396, December 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900796.

Barnard:1984:FFF

[Bar84a] A. J. Barnard. FURI — a Fortran function writer. Computer Physics Communications, 35(1–3):C–429, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826387.

Barnett:1984:CCB

[Bar84b] A. R. Barnett. Coulfg: Coulomb and Bessel functions and their derivatives, for real arguments, by Steed’s method. Computer Physics Communications, 35(1–3):C–812–C–813, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829306.

Barnett:1984:KCF

[Bar84c] A. R. Barnett. Klein: Coulomb functions for real λ and positive energy to high accuracy. Computer Physics Communications, 35(1–3):C–753, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828842.

Barnett:1984:RMR

[Bar84d] A. R. Barnett. RCWFF — a modification of the real Coulomb wavefunction program RCWFN. Computer Physics Communications, 35(1–3):C–370, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825850.

Bartschat:1984:PCO

[Bar84e] K. Bartschat. Program to calculate observable quantities from scattering amplitudes for inelastic electron-atom collisions. Computer Physics Communications, 35(1–3):C–924, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584830222.

Basrak:1987:CPD

[Bas87a] Z. Basrak. A computer program for determining the complete reaction amplitude for two-body nuclear reactions involving zero-spin particles. Computer Physics Communications, 46(1):155–178, July 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900439.

Basrak:1987:DPS

[Bas87b] Z. Basrak. Determination of the physical scattering matrix from a complete set of ambiguous solutions of the scattering problem by using the shortest-path method. Computer Physics Communications, 46(1):179–186, July 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900440.

Basrak:1987:RPO

[Bas87c] Z. Basrak. A routine for parameter optimization using an accelerated grid-search method. Computer Physics Communications, 46(1):149–154, July 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900427.

Bassler:1987:GEP

[Bas87d] E. Bassler. The graphical editor program: GEP. Computer Physics Communications, 45(1–3):201–205, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901573.

Baszenski:1988:ING

[Bas88] G. Baszenski. Implementation of the NAG graphical supplement for IBM compatible personal computers. Computer Physics Communications, 50(1–2):217–223, July 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901270.

Bates:1984:FCH

[Bat84] G. N. Bates. A fixed core Hartree–Fock program for calculating bound and continuum orbitals. Computer Physics Communications, 35(1–3):C–276, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825072.

Baas:1984:AID

[BB84a] R. Ch. Baas and C. I. M. Beenakker. Analysis of the intensity distribution in the rotational structure of the electronic spectra of diatomic molecules by computer simulation. Computer Physics Communications, 35(1–3):C–277, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825084.

Baig:1984:CHE

[BB84b] M. Baig and C. Bourrely. Calculation of hadron elastic scattering amplitude from higher order reggeon field theory. Computer Physics Communications, 32(3):281–289, June 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900912.

Basso:1984:AGP

[BB84c] P. Basso and C. Bourrely. An algorithm generating permutations with identical objects. Computer Physics Communications, 35(1–3):C–46, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823127.

Bartschat:1986:RMR

[BB86] K. Bartschat and P. G. Burke. Resfit — a multichannel resonance fitting program. Computer Physics Communications, 41(1):75–84, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900226.

Berrington:1984:GPC

[BBC^{+}84] K. A. Berrington, P. G. Burke, J. J. Chang, A. T. Chivers, W. D. Robb, and K. T. Taylor. A general program to calculate atomic continuum processes using the R-matrix method. Computer Physics Communications, 35(1–3):C–270–C–273, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825059.

Bock:1987:PTP

[BBC^{+}87] R. Bock, R. Brun, O. Couet, J. C. Marin, R. Nierhaus, L. Pape, N. Saumon, C. Vandoni, and P. Zanarini. PAW — towards a physics analysis workstation. Computer Physics Communications, 45(1–3):181–190, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901548.

Booth:1989:LSA

[BBC^{+}89] S. P. Booth, K. C. Bowler, D. J. Candlin, R. D. Kenway, B. J. Pendleton, A. M. Thornton, D. J. Wallace, J. Blair-Fish, and D. Roweth. Large scale applications of transputers in HEP: the Edinburgh Concurrent Supercomputer Project. Computer Physics Communications, 57(1–3):101–107, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901975.

Bizeau:1989:UPA

[BBD^{+}89] C. Bizeau, A. Bogaerts, R. W. Dobinson, D. R. N. Jeffery, W. Lu, C. Parkman, and Y. Perrin. The use and possible abuse of transputer links. Computer Physics Communications, 57(1–3):301–308, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902324.

Bahler:1981:CBM

[BBE^{+}81] P. Bahler, A. Bogaerts, C. Eck, M. Ferran, D. Jacobs, A. Lacourt, J. Ogilvie, H. Overas, and J. O. Petersen. Clusters of 16/32 bit minicomputers in high energy physics experiments at CERN. Computer Physics Communications, 22(2–3):285– 292, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900618.

Banks:1984:CCP

[BBH^{+}84] D. Banks, K. S. Barnes, P. E. Hughes, I. C. Percival, D. Richards, N. A. Valentine, and J. Mc. B. Wilson. Classical collisions of protons with hydrogen atoms. Computer Physics Communications, 35(1–3):C–427–C–428, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826375.

Bingham:1988:AOS

[BBH^{+}88] R. Bingham, D. A. Bryant, D. S. Hall, J. M. Dawson, F. Kazeminejad, J. J. Su, and C. M. C. Nairn. AMPTE observations and simulation results. Computer Physics Communications, 49(1):257–266, April 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902342.

Berrington:1984:NVG

[BBL^{+}84] K. A. Berrington, P. G. Burke, M. Le Dourneuf, W. D. Robb, K. T. Taylor, and Vo Ky Lan. A new version of the general program to calculate atomic continuum processes using the R-matrix method. Computer Physics Communications, 35(1–3):C–475–C–477, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826739.

Baluja:1982:MPP

[BBM82] K. L. Baluja, P. G. Burke, and L. A. Morgan. R-matrix propagation program for solving coupled second-order differential equations. Computer Physics Communications, 27(3):299–307, September 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901771.

Baluja:1984:MPP

[BBM84] K. L. Baluja, P. G. Burke, and L. A. Morgan. R-matrix propagation program for solving coupled second-order differential equations. Computer Physics Communications, 35(1–3): C–823, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829380.

Bhowmick:1989:ICA

[BBR89a] Siddhartha Bhowmick, Ranjan Bhattacharya, and Dhiranjan Roy. Iterations of convergence accelerating nonlinear transforms. Computer Physics Communications, 54(1):31–46, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900301.

Brun:1989:CSC

[BBR89b] M. Brun, R. Brun, and A. A. Rademakers. CMZ — a source code management system. Computer Physics Communications, 57(1–3):235–238, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902191.

Beu:1983:PAS

[BC83a] T. A. Beu and R. I. Câmpeanu. Prolate angular spheroidal wave functions. Computer Physics Communications, 30(2): 187–192, September/October 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900619.

Beu:1983:PRS

[BC83b] T. A. Beu and R. I. Câmpeanu. Prolate radial spheroidal wave functions. Computer Physics Communications, 30(2): 177–185, September/October 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900607.

Beu:1984:PAS

[BC84a] T. A. Beu and R. I. Câmpeanu. Prolate angular spheroidal wave functions. Computer Physics Communications, 35(1–3): C–909, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584830088.

Beu:1984:PRS

[BC84b] T. A. Beu and R. I. Câmpeanu. Prolate radial spheroidal wave functions. Computer Physics Communications, 35(1–3): C–908, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584830076.

Braithwaite:1984:RRM

[BC84c] W. J. Braithwaite and J. G. Cramer. The reduced rotation matrix. Computer Physics Communications, 35(1–3):C–142, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823966.

Booth:1987:ESF

[BC87] A. W. Booth and J. T. Carroll. An expert system for FASTBUS diagnosis. Computer Physics Communications, 45(1–3):67–76, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558790141X.

Barahona:1988:MGS

[BC88a] Francisco Barahona and Adolfo Casari. On the magnetisation of the ground states in two dimensional Ising spin glasses. Computer Physics Communications, 49(3):417–421, June 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588900021.

Buzuloiu:1988:SIP

[BC88b] V. Buzuloiu and D. Coltuc. Small image processing systems (performances, architectures, functions). Computer Physics Communications, 50(1–2):229–235, July 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901294.

Browning:1985:SSD

[BCC^{+}85] D. J. Browning, G. M. Cain, N. P. Carmichael, F. G. Gouldstone, A. W. Wadsley, S. J. Webb, and P. Winder. Software systems development in petroleum engineering. Computer Physics Communications, 38(2):301–308, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900943.

Bacilieri:1989:SFO

[BCC^{+}89] P. Bacilieri, B. Caccia, R. Cardarelli, G. P. Carlucci, O. Ciaffoni, M. Coli, G. Di Pirro, M. L. Ferrer, A. Ghiselli, A. Martini, G. Medici, G. Mirabelli, E. Pace, R. Santonico, L. Trasatti, E. Valente, and S. Valentini. Starnet, a fiber optic metropolitan area network with centralized control. Computer Physics Communications, 57(1–3):459–465, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902634.

Butera:1987:HTE

[BCCM87] P. Butera, R. Cabassi, M. Comi, and G. Marchesini. High temperature expansion via Schwinger–Dyson equations: The planar rotator model on a triangular lattice. Computer Physics Communications, 44(1–2):143–156, April/May 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900245.

Bank:1989:IMS

[BCD^{+}89] R. E. Bank, W. M. Coughran, Jr., M. A. Driscoll, R. K. Smith, and W. Fichtner. Iterative methods in semiconductor device simulation. Computer Physics Communications, 53(1–3):201– 212, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901604.

Booth:1989:KBA

[BCF^{+}89] Alexander W. Booth, John T. Carroll, Robert Forster, Geriann Goeransson, Leif Gustafsson, and Norman Ho. A knowledgebased approach to network and module diagnosis. Computer Physics Communications, 57(1–3):332–338, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902385.

Barnes:1985:PCM

[BCFK85] L. A. Barnes, G. S. Chandler, B. N. Figgis, and D. C. Khan. A program to calculate magnetic form factors for transition metal systems. Computer Physics Communications, 36(4):373– 382, June 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585900268.

Bhadra:1989:ZTL

[BCKO89] S. Bhadra, M. Crombie, D. Kirkby, and R. S. Orr. The ZEUS third-level trigger system. Computer Physics Communications, 57(1–3):321–324, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902361.

Barkai:1983:WLC

[BCM83a] D. Barkai, M. Creutz, and K. J. M. Moriarty. Wilson loop calculations in four-dimensional lattice gauge theory on the CDC Cyber 205. Computer Physics Communications, 30(1):13–19, July/August 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583901170.

Botten:1983:CZA

[BCM83b] L. C. Botten, M. S. Craig, and R. C. McPhedran. Complex zeros of analytic functions. Computer Physics Communications, 29(3):245–259, May 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558390005X.

Batana:1984:AFI

[BCM84a] Alicia Batana, Ernesto R. Conzalez, and Maria C. Monard. Analysis of Faradaic impedance experimental measurements. Computer Physics Communications, 35(1–3):C–574, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827484.

Botten:1984:CZA

[BCM84b] L. C. Botten, M. S. Craig, and R. C. McPhedran. Complex zeros of analytic functions. Computer Physics Com-

munications, 35(1–3):C–878–C–879, ????  1984.  CO-

DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829823.

Barkai:1987:ADE

[BCMR87] D. Barkai, M. Campostrini, K. J. M. Moriarty, and C. Rebbi. Applications development of the ETA-10. Computer Physics Communications, 46(1):13–33, July 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465587900348.

Brun:1989:PGP

[BCVZ89] René Brun, Olivier Couet, Carlo E. Vandoni, and Pietro Zanarini. PAW, a general-purpose portable software tool for data analysis and presentation. Computer Physics Communications, 57(1–3):432–437, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902579.

Banuelos:1980:PCR

[BD80a] Alicia Bañuelos and Ricardo Angel Depine. A program for computing the Riemann zeta function for complex argument. Computer Physics Communications, 20(3):441–445, November 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900211.

Beniaminy:1980:SBM

[BD80b] Israel Beniaminy and Moshe Deutsch. A spline-based method for experimental data deconvolution. Computer Physics Communications, 21(2):271–277, December 2, 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900454.

Beniaminy:1981:SBM

[BD81] I. Beniaminy and M. Deutsch. A spline based method for experimental data deconvolution: Reply to a comment by H. Fredrikze and P. Verkerk. Computer Physics Communications, 24(1):9–10, September/October 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558190103X.

Beniaminy:1982:ASH

[BD82] Israel Beniaminy and Moshe Deutsch. ABEL: Stable, high accuracy program for the inversion of Abel’s integral equation. Computer Physics Communications, 27(4):415–422, October 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901023.

Banuelos:1984:PCR

[BD84a] Alicia Bañuelos and Ricardo Angel Depine. A program for computing the Riemann zeta function for complex argument. Computer Physics Communications, 35(1–3):C–647, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828076.

Beniaminy:1984:ASH

[BD84b] Israel Beniaminy and Moshe Deutsch. Abel: Stable, high accuracy program for the inversion of Abel’s integral equation. Computer Physics Communications, 35(1–3):C–828–C–829, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829422.

Beniaminy:1984:SBM

[BD84c] Israel Beniaminy and Moshe Deutsch. A spline-based method for experimental data deconvolution. Computer Physics Communications, 35(1–3):C–668–C–669, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828234.

Bettels:1984:PGA

[BD84d] J. Bettels and P. Dodd. A program for the generation of artificial bubble chamber events. Computer Physics Communications, 35(1–3):C–133, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823875.

Berg:1987:CFP

[BD87] Bernd A. Berg and Alberto Devoto. Correlation functions from probability densities. Computer Physics Communications, 46(3):345–349, September 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900919.

Bocciolini:1981:MDB

[BdCP81] M. Bocciolini, G. di Caporiacco, and G. Parrini. A method for digitizing bubble chamber pictures. Computer Physics Communications, 22(2–3):375–382, April 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900734.

Bird:1984:IER

[BDD84] B. Bird, C. Daul, and P. Day. Inter-electron repulsion integrals for three-open-shell configurations in cubic symmetry. Computer Physics Communications, 35(1–3):C–468, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826685.

Bardin:1984:CFE

[BDG^{+}84] C. Bardin, Y. Dandeu, L. Gauthier, J. Guillermin, T. Lena, J.-M. Pernet, H. H. Wolter, and T. Tamura. Coulomb functions in entire (η, π)-plane. Computer Physics Communications, 35(1–3):C–125–C–126, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823826.

Booth:1989:EMC

[BDJ^{+}89] S. P. Booth, R. W. Dobinson, D. R. N. Jeffery, W. Lu, K. M. Storr, and A. Thornton. An evaluation of the Meiko computing surface for HEP Fortran farming. Computer Physics Communications, 57(1–3):486–491, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902695.

Berends:1986:MCSa

[BDK86a] F. A. Berends, P. H. Daverveldt, and R. Kleiss. Monte Carlo simulation of two-photon processes: I: Radiative corrections to multiperipheral e^{+}e^{−}μ^{+}μ^{−} production. Computer Physics Communications, 40(2–3):271–284, June 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901141.

Berends:1986:MCSb

[BDK86b] F. A. Berends, P. H. Daverveldt, and R. Kleiss. Monte Carlo simulation of two-photon processes: II: Complete lowest order calculations for four-lepton production processes in electron-positron collisions. Computer Physics Communications, 40(2–3):285–307, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901153.

Berends:1986:MCSc

[BDK86c] F. A. Berends, P. H. Daverveldt, and R. Kleiss. Monte Carlo simulation of two-photon processes: III: Complete lowest order calculations for e^{+}e^{−} → e^{+}e^{−}μ^{+}μ^{−} with large angle tagging conditions. Computer Physics Communications, 40(2–3):309– 326, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586901165.

Banuelos:1981:PCF

[BDM81] Alicia Bañuelos, Ricardo Angel Depine, and Roberto Claudio Mancini. A program for computing the Fermi–Dirac functions. Computer Physics Communications, 21(3):315–322, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900126.

Banuelos:1984:PCF

[BDM84] Alicia Bañuelos, Ricardo Angel Depine, and Roberto Claudio Mancini. A program for computing the Fermi–Dirac functions. Computer Physics Communications, 35(1–3):C–670, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828246.

Bonissent:1989:AIS

[BE89] A. Bonissent and F. Etienne. Artificial intelligence steering for the interactive analysis of a high energy physics experiment. Computer Physics Communications, 57(1–3):447–451, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589902609.

Behrend:1981:FTC

[Beh81] H.-J. Behrend. The fast trackfinder for the CELLO-experiment at DESY. Computer Physics Communications, 22(2–3):365– 374, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900722.

Behrens:1984:SSS

[Beh84] H. Behrens. Summary of this summer school. Computer Physics Communications, 33(1–3):291–297, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901346.

Beleznay:1984:NSR

[Bel84a] F. Beleznay. Numerical solution of the radial Schrödinger equation. Computer Physics Communications, 35(1–3):C–144, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482398X.

Bellard:1984:CCD

[Bel84b] Sharon Bellard. The Cambridge Crystallographic Data Base. Computer Physics Communications, 33(1–3):71–78, August/ September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901085.

Bolger:1984:WPH

[BEM84] Joseph E. Bolger, Hunter Ellinger, and C. Fred Moore. Worker, a program for histogram manipulation. Computer Physics Communications, 35(1–3):C–545–C–546, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827228.

Benard:1982:MPO

[Ben82] Marc Benard. MATSUP: a program to obtain two-electron repulsion integrals from a sparse file of P supermatrix elements. Computer Physics Communications, 27(1):79–86, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900108.

Benard:1984:MPO

[Ben84a] Marc Benard. Matsup: a program to obtain two-electron repulsion integrals from a sparse file of & supermatrix elements. Computer Physics Communications, 35(1–3):C–809–C– 810, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829288.

Bengtsson:1984:LMC

[Ben84b] H.-U. Bengtsson. The Lund Monte Carlo for high-pT physics. Computer Physics Communications, 31(4):323–355, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900183.

Benz:1988:ASP

[Ben88] Willy Benz. Applications of smooth particle hydrodynamics (SPH) to astrophysical problems. Computer Physics Communications, 48(1):97–105, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588900276.

Bernhardi:1980:IDM

[Ber80] Karl Bernhardi. An improved deconvolution method for Bremsstrahlung spectra from hot plasmas. Computer Physics Communications, 19(1):17–21, January/March 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900612.

Bergerhoff:1984:DBI

[Ber84a] G. Bergerhoff. Data base for inorganic crystal structures. Computer Physics Communications, 33(1–3):79–84, August/ September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901097.

Bermudez:1984:EFS

[Ber84b] Victor M. Bermudez. Ellips — a Fortran simulation of a polarization-modulation ellipsometer. Computer Physics Communications, 35(1–3):C–426, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0010465584826351.

Berendsen:1987:BAM

[Ber87a] H. J. C. Berendsen. Biophysical applications of molecular dynamics. Computer Physics Communications, 44(3):233–242, June 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900786.

Bergdolt:1987:CMT

[Ber87b] G. Bergdolt. Clifford multiplication tables. Computer Physics Communications, 44(1–2):137–142, April/May 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900233.

Bertrand:1987:DIA

[Ber87c] D. Bertrand. The DELPHI interactive analysis and TANA-GRA. Computer Physics Communications, 45(1–3):207–213, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901585.

Barlow:1982:PAA

[BES82] R. H. Barlow, D. J. Evans, and J. Shanehchi. Performance analysis of algorithms on asynchronous parallel processors. Computer Physics Communications, 26(3–4):233–236, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901102.

Berrington:1987:CCC

[BES^{+}87] K. A. Berrington, W. B. Eissner, H. E. Saraph, M. J. Seaton, and P. J. Storey. A comparison of close-coupling calculations using UCL and QUB codes. Computer Physics Communications, 44(1–2):105–119, April/May 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558790021X.

Betak:1984:PSC

[Bet84] Emil Beták. Program for spectra and cross-section calculations within the pre-equilibrium model of nuclear reactions. Computer Physics Communications, 35(1–3):C–294, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825242.

Browne:1984:EMR

[BF84] E. Browne and F. R. Femenia. Electromagnetic M1 reduced transition probabilities for pure and mixed Nilsson states in odd A nuclei. Computer Physics Communications, 35(1–3):C– 101, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584823590.

Berny:1989:AGS

[BF89] L. Berny and H. Frese. Automatic generation of software detailed design documents for language programs. Computer Physics Communications, 57(1–3):476–477, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990266X.

Barnett:1984:CWF

[BFSG84] A. R. Barnett, D. H. Feng, J. W. Steed, and L. J. B. Goldfarb. Coulomb wave functions for all real η and ρ. Computer Physics Communications, 35(1–3):C–285, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825151.

Bogdanova:1981:UOP

[BG81] N. Bogdanova and V. Gadjokov. Use of orthonormal polynomials in calibration problems. Computer Physics Communications, 24(2):225–229, November 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900965.

Barao:1989:DCP

[BGG^{+}89] F. Barao, C. Gaspar, Ph. Gavillet, J. Ph. Laugier, B. Martin, P. Moreau, M. Pimenta, M. Reis, and J. Varela. DEL-PHI’s central partition. Computer Physics Communications, 57(1–3):358–363, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589902439.

Briggs:1989:CST

[BGGW^{+}89] D. Briggs, T. Glanzman, P. Grosse-Wiesmann, J. Tinsman, S. Holmgren, and M. W. Schaad. A calorimeter software trigger for the mark II detector at SLC. Computer Physics Communications, 57(1–3):273–277, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902270.

Brower:1987:DSS

[BGM87] R. Brower, R. Giles, and G. Maturana. Development of scientific software in the LISP machine environment. Computer Physics Communications, 45(1–3):427–431, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901858.

Bacilieri:1981:MIR

[BGMM81] P. Bacilieri, A. Ghiselli, P. Matteuzzi, and M. Masetti. A microprogrammed interactive refresh display system which can overlap digital and analogue data with hardware clipping facility. Computer Physics Communications, 22(2–3):231–237, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900540.

Beatham:1984:MPC

[BGMP84] N. Beatham, I. P. Grant, B. J. McKenzie, and N. C. Pyper. MCBP — a program to calculate angular coefficients of the Breit interaction between electrons in the low energy limit. Computer Physics Communications, 35(1–3):C–591–C– 592, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827629.

Belkic:1981:CTC

[BGS81] Dz. Belkić, R. Gayet, and A. Salin. Computation of total cross-sections for electron capture in high energy ion-atom collisions. Computer Physics Communications, 23(2):153–167, July 1, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900308.

Belkic:1983:CTC

[BGS83] Dz. Belkić, R. Gayet, and A. Salin. Computation of total cross-sections for electron capture in high energy collisions. II. Computer Physics Communications, 30(2):193–205, September/October 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900620.

Belkic:1984:CTCb

[BGS84a] Dz. Belkić, R. Gayet, and A. Salin. Computation of total cross-sections for electron capture in high energy collisions. II. Computer Physics Communications, 35(1–3):C–910, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558483009X.

Belkic:1984:CTCc

[BGS84b] Dz. Belkić, R. Gayet, and A. Salin. Computation of total cross-sections for electron capture in high energy collisions. III. Computer Physics Communications, 32(4):385–397, July 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900559.

Belkic:1984:CTCa

[BGS84c] Dz. Belkić, R. Gayet, and A. Salin. Computation of total cross-sections for electron capture in high energy ion-atom collisions. Computer Physics Communications, 35(1–3):C–714–C– 715, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046558482857X.

Bakker:1981:PSS

[BH81] M. Bakker and D. Hoonhout. A program to solve a solute diffusion problem with segregation at a moving interface. Computer Physics Communications, 22(4):439–450, May 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901417.

Bernard:1982:VES

[BH82] L. C. Bernard and F. J. Helton. A vectorizable eigenvalue solver for sparse matrices. Computer Physics Communications, 25(1):73–80, January 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900455.

Bakker:1984:PSS

[BH84a] M. Bakker and D. Hoonhout. A program to solve a solute diffusion problem with segregation at a moving interface. Computer Physics Communications, 35(1–3):C–696, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828453.

Bernard:1984:VES

[BH84b] L. C. Bernard and F. J. Helton. A vectorizable elgenvalue solver for sparse matrices. Computer Physics Communications, 35(1–3):C–769, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828970.

Beard:1985:PPD

[BH85a] S. J. Beard and R. W. Hockney. POT4A — a program for the direct solution of Poisson’s equation in complex geometries. Computer Physics Communications, 36(1):25–57, March 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900177.

Brown:1985:FFP

[BH85b] James W. Brown and Stanford B. Hooker. FIXSRC: a Fortran preprocessor. Computer Physics Communications, 38 (3):435–440, December 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901110.

Bogdanova:1988:RPE

[BH88] N. Bogdanova and H. Hogreve. A REDUCE package for exact Coulomb interaction matrix elements. Computer Physics Communications, 48(2):319–326, February 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900513.

Burkitt:1989:PCA

[BH89] Anthony N. Burkitt and Dieter W. Heermann. Parallelization of a cluster algorithm. Computer Physics Communications, 54(2–3):201–209, June/July 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900829.

Bernard:1981:GMS

[BHM81] L. C. Bernard, F. J. Helton, and R. W. Moore. GATO: An MHD stability code for axisymmetric plasmas with internal separatrices. Computer Physics Communications, 24(3–4):377–380, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901600.

Bandyopadhyay:1984:GSI

[BHSS84] S. Bandyopadhyay, J. G. Hughes, F. J. Smith, and K. Sen. A generalized scientific information system. Computer Physics Communications, 33(1–3):49–53, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901061.

Bengtsson:1985:ILM

[BI85] H.-U. Bengtsson and G. Ingelman. Improved Lund Monte Carlo for high-p_{T} physics. Computer Physics Communications, 34(3):251–270, January 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900037.

Bienstock:1983:PCL

[Bie83] S. Bienstock. A program for the calculation of Landau–Zener cross sections and rate coefficients. Computer Physics Communications, 29(4):333–339, June 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900127.

Bienstock:1984:PCL

[Bie84] S. Bienstock. A program for the calculation of Landau–Zener cross sections and rate coefficients. Computer Physics Communications, 35(1–3):C–885, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829884.

Billing:1984:RCC

[Bil84] G. D. Billing. Rate constants and cross sections for vibrational transitions in atom-diatom and diatom-diatom collisions. Computer Physics Communications, 32(1):45–62, April 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900079.

Billing:1987:RCV

[Bil87] G. D. Billing. Rate constants for vibrational transitions in diatom-diatom collisions. Computer Physics Communications, 44(1–2):121–136, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900221.

Billoir:1989:PTR

[Bil89] Pierre Billoir. Progressive track recognition with a kalmanlike fitting procedure. Computer Physics Communications, 57 (1–3):390–394, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558990249X.

Bingham:1987:ACF

[Bin87] D. Bingham. An approximation to the calculation of the force in molecular dynamic simulations. Computer Physics Communications, 43(2):203–207, January 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587902050.

Bird:1985:LTD

[Bir85] Peter L. Bird. Linear time detection of inherent parallelism in sequential programs. Computer Physics Communications, 37(1–3):69–76, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901377.

Burke:1981:FSMa

[BJ81a] V. M. Burke and C. Jackson. A Fortran system to maintain a program library: 1. Storage of the program decks in magnetic tape files. Computer Physics Communications, 22(1):59–75, February/March 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900795.

Burke:1981:FSMb

[BJ81b] V. M. Burke and C. Jackson. A Fortran system to maintain a program library: 2. Retrieval of program decks from the library files. Computer Physics Communications, 22(1):77–84, February/March 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900801.

Burke:1984:FSMa

[BJ84a] V. M. Burke and C. Jackson. A Fortran system to maintain a program library: 1. Storage of the program decks in magnetic tape files. Computer Physics Communications, 35(1–3): C–685, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828362.

Burke:1984:FSMb

[BJ84b] V. M. Burke and C. Jackson. A Fortran system to maintain a program library: 2. Retrieval of program decks from the library files. Computer Physics Communications, 35(1–3):C–686, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828374.

Boissat:1989:MHI

[BJM89] C. Boissat, R. Jones, and G. Mornacchi. The model human interface. Computer Physics Communications, 57(1–3):512–515, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589902762.

Beck:1984:CGW

[BK84a] Robert E. Beck and Bernard Kolman. Computer generated Weyl groups. Computer Physics Communications, 35(1–3):C– 134, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584823887.

Beck:1984:ROM

[BK84b] Robert E. Beck and Bernard Kolman. Racah’s outer multiplicity formula. Computer Physics Communications, 35(1–3): C–265, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824996.

Bleris:1983:PPS

[BKD83] G. L. Bleris, Th. Karakostas, and P. Delavignette. A programming package for the study of high angle grain boundaries by using TEM. Computer Physics Communications, 28(3):287–297, January 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900462.

Bleris:1984:PPS

[BKD84] G. L. Bleris, Th. Karakostas, and P. Delavignette. A programming package for the study of high angle grain boundaries by using TEM. Computer Physics Communications, 35(1–3):C– 853, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829628.

Bohr:1983:FSL

[BKHJ83] H. Bohr, E. Katznelson, F. Hansen, and K. Jensen. Fast simulation of lattice systems. Computer Physics Communications, 30(4):337–347, December 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900759.

Berends:1983:MCS

[BKJ83] F. A. Berends, R. Kleiss, and S. Jadach. Monte Carlo simulation of radiative corrections to the processes e^{+}e^{−} → μ^{+}μ^{−} and e^{+}e^{−} → qq in the Z_{0} region. Computer Physics Communications, 29(2):185–200, April 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900735.

Berends:1984:MCS

[BKJ84] F. A. Berends, R. Kleiss, and S. Jadach. Monte Carlo simulation of radiative corrections to the processes e^{+}e^{−}- > μ^{+}μ^{−} and e^{+−} → q in the Z_{0} region. Computer Physics Communications, 35(1–3):C–874, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829781.

Bisseling:1986:MID

[BKK86] R. H. Bisseling, R. Kosloff, and D. Kosloff. Multidimensional interpolation and differentiation based on an accelerated sinc interpolation procedure. Computer Physics Communications, 39(3):313–332, April 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900937.

Brackbill:1988:FLD

[BKR88] J. U. Brackbill, D. B. Kothe, and H. M. Ruppel. Flip: a low-dissipation, particle-in-cell method for fluid flow. Computer Physics Communications, 48(1):25–38, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900203.

Balint-Kurti:1980:EPC

[BKvLSE80] G. G. Balint-Kurti, J. H. van Lenthe, R. Saktreger, and L. Eno. Edwin — a program for calculating inelastic molecular collision cross sections using the exponential distorted wave and related approximate methods. Computer Physics Communications, 19(3):359–375, July/August 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900892.

Balint-Kurti:1984:EPC

[BKvLSE84] G. G. Balint-Kurti, J. H. van Lenthe, R. Saktreger, and L. Eno. Edwin — a program for calculating inelastic molecular collision cross sections using the exponential distorted wave and related approximate methods. Computer Physics Communications, 35(1–3):C–625–C–626, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827897.

Boeyens:1986:SMR

[BL86] J. C. A. Boeyens and D. C. Levendis. Simulation of molecular reorientation in crystals. Computer Physics Communications, 39(2):221–231, February/March 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901335.

Blenski:1988:ISM

[BL88] Tomasz Błeński and Jacques Ligou. An improved shooting method for one-dimensional Schrödinger equation. Computer Physics Communications, 50(3):303–311, August 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901865.

Blake:1984:CLE

[Bla84] R. J. Blake. Calculation of low energy atomic diffraction (LEAD) intensities. Computer Physics Communications, 33 (4):425–444, October 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584901486.

Blanc:1988:MIC

[Bla88] Michel Blanc. Magnetosphere-ionosphere coupling. Computer Physics Communications, 49(1):103–118, April 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902196.

Burshtein:1988:TIR

[BLB88] Z. Burshtein, D. Levron, and G. Bialolenker. Thermally induced refractive index gradients in a dye-laser cell. Computer Physics Communications, 51(3):349–353, November 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901476.

Barnes:1988:KSL

[BLM^{+}88] D. C. Barnes, H. R. Lewis, Z. Mikic, J. L. Schwartzmeier, and J. Staudemeier. Kinetic stability of large-scale MHD modes. Computer Physics Communications, 48(1):145–147, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558890032X.

Bacilieri:1981:AAH

[BLMU81] P. Bacilieri, M. L. Luvisetto, M. Masetti, and E. Ugolini. Automatic analysis of a huge amount of pictures related to a spark chamber experiment of the Milano–Dubna collaboration. Computer Physics Communications, 22(2–3):273–277, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558190059X.

Bloch:1982:LCSa

[Blo82a] T. Bloch. Large computer systems and new architectures. Computer Physics Communications, 26(1–2):125–145, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901655.

Bloch:1982:LCSb

[Blo82b] T. Bloch. Large computer systems and new architectures addendum. Computer Physics Communications, 26(1–2):147– 152, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901667.

Blobel:1989:DP

[Blo89] Volker Blobel. From DST to publication. Computer Physics Communications, 57(1–3):148–155, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990204X.

Bhanot:1982:FAM

[BLR82] Gyan Bhanot, Christian B. Lang, and Claudio Rebbi. A fast algorithm for Monte Carlo simulations of 4-d lattice gauge theories with finite groups. Computer Physics Communications, 25(3):275–287, March 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900236.

Bhanot:1984:FAM

[BLR84] Gyan Bhanot, Christian B. Lang, and Claudio Rebbi. A fast algorithm for Monte Carlo simulations of 4-d lattice gauge theories with finite groups. Computer Physics Communications, 35(1–3):C–789–C–790, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829148.

Blum:1981:SCE

[BLT81] J. Blum, J. Le Foll, and B. Thooris. The self-consistent equilibrium and diffusion code sced. Computer Physics Communications, 24(3–4):235–254, December 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901491.

Barkai:1982:CMC

[BM82a] D. Barkai and K. J. M. Moriarty. Can the Monte Carlo method for lattice gauge theory calculations be effectively vectorized? Computer Physics Communications, 27(2):105–111, August 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900662.

Barkai:1982:VMCa

[BM82b] D. Barkai and K. J. M. Moriarty. Vectorizing the Monte Carlo algorithm for lattice gauge theory calculations on the CDC Cyber 205. Computer Physics Communications, 25(1):57–62, January 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900431.

Barkai:1982:VMCb

[BM82c] D. Barkai and K. J. M. Moriarty. Vectorizing the Monte Carlo algorithm for lattice gauge theory calculations on the CDC Cyber 205. Computer Physics Communications, 26(3–4):477– 479, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901461.

Barkai:1984:VMC

[BM84a] D. Barkai and K. J. M. Moriarty. Vectorizing the Monte Carlo algorithm for lattice gauge theory calculations on the CDC Cyber 205. Computer Physics Communications, 35(1–3):C–766– C–767, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828957.

Brundrit:1984:CFE

[BM84b] G. B. Brundrit and M. J. Miketinac. Calculation of the form of an equilibrium poloidal magnetic field contained in a polytropic star. Computer Physics Communications, 35(1–3):C–389, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826016.

Bossavit:1985:APT

[BM85a] Alain Bossavit and Bertrand Meyer. An application of program transformation to supercomputer programming. Computer Physics Communications, 37(1–3):27–38, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558590133X.

Bowgen:1985:IFD

[BM85b] G. S. J. Bowgen and J. J. Modi. Implementation of QR factorization on the DAP using Householder transformations. Computer Physics Communications, 37(1–3):167–170, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901493.

Barkai:1986:ADC

[BM86] D. Barkai and K. J. M. Moriarty. Applications development on the CDC CYBER 205. Computer Physics Communications, 40(2–3):159–171, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901050.

Brackbill:1988:P

[BM88] Jeremiah Brackbill and Joseph Monaghan. Preface. Computer Physics Communications, 48(1):xi–xii, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558890015X.

Buxton:1985:SMA

[BMBW85] B. F. Buxton, D. W. Murray, H. Buxton, and N. S. Williams. Structure-from-motion algorithms for computer vision on an SIMD architecture. Computer Physics Communications, 37 (1–3):273–280, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901626.

Brower:1985:PDN

[BMG^{+}85] R. C. Brower, G. Maturana, R. C. Giles, K. J. M. Moriarty, and S. Samuel. Proton decay: Numerical simulations confront grand unification. Computer Physics Communications, 38(1):9–14, August/September 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900402.

Boley:1989:PQA

[BMK89] Daniel Boley, Robert Maier, and Joung Kim. A parallel QR algorithm for the nonsymmetric eigenvalue problem. Computer Physics Communications, 53(1–3):61–70, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901483.

Barkai:1984:HOV

[BMR84] D. Barkai, K. J. M. Moriarty, and C. Rebbi. A highly optimized vectorized code for Monte Carlo simulations of Su(3) lattice gauge theories. Computer Physics Communications, 32(1):1– 9, April 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558490002X.

Barkai:1985:LGC

[BMR85a] D. Barkai, K. J. M. Moriarty, and C. Rebbi. Lattice gauge calculation in particle theory. Computer Physics Communications, 36(3):241–247, May 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900542.

Barkai:1985:MCG

[BMR85b] D. Barkai, K. J. M. Moriarty, and C. Rebbi. A modified conjugate gradient solver for very large systems. Computer Physics Communications, 36(1):1–8, March 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900141.

Barkai:1985:STH

[BMR85c] D. Barkai, K. J. M. Moriarty, and C. Rebbi. Supercomputers in theoretical high-energy physics. Computer Physics Communications, 38(1):1–7, August/September 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900396.

Batana:1987:LTG

[BMS87] Alicia Batana, María C. Monard, and María Rosario Soriano. Low temperature Grüneisen parameter of cubic ionic crystals. Computer Physics Communications, 43(3):399–411, February/ March 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900579.

Bacchus-Montabonel:1983:NAT

[BMV83] M. C. Bacchus-Montabonel and P. Vermeulin. Non-adiabatic transformation of quantum-chemistry energy hypersurfaces. Computer Physics Communications, 30(2):163–167, September/October 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900589.

Bacchus-Montabonel:1984:NAT

[BMV84] M. C. Bacchus-Montabonel and P. Vermeulin. Non-adiabatic transformation of quantum-chemistry energy hypersurfaces.

Computer Physics Communications, 35(1–3):C–906, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584830052.

Beppu:1981:NFE

[BN81] Yoshitaka Beppu and Ichizo Ninomiya. Nicer — fast eigenvalue routines. Computer Physics Communications, 23(2):123–126, July 1, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900266.

Bajla:1980:ARA

[BO80] I. Bajla and G. A. Ososkov. Application of REDUCE-2 algebraic manipulation system in calibration problems of track chamber picture processing devices. Computer Physics Communications, 20(1):81–83, September 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580901125.

Bock:1987:SIL

[Boc87] R. K. Bock. Software issues for large detectors. Computer Physics Communications, 45(1–3):15–25, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901378.

Bock:1989:BTH

[Boc89] R. K. Bock. Bringing together high energy physicist and computer scientist: a summary of the Oxford Conference on Computing in High Energy Physics. Computer Physics Communications, 57(1–3):1–7, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901859.

Bok:1984:SAC

[Bok84] J. Bok. A subroutine for approximation by cubic splines in the least squares sense. Computer Physics Communications, 35(1–3):C–524, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827058.

Boreham:1982:PCC

[Bor82] B. W. Boreham. PROION: a code for calculating ionisation threshold intensities and ionisation periods in high-intensity-laser irradiation plasmas. Computer Physics Communications, 27(1):65–71, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558290008X.

Boreham:1984:PCC

[Bor84] B. W. Boreham. Proion: a code for calculating ionisation threshold intensities and ionisation periods in high-intensity-laser irradiated plasmas. Computer Physics Communications, 35(1–3):C–807, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829264.

Bornat:1989:SAB

[Bor89] R. Bornat. SASD — all bubbles and no code. Computer Physics Communications, 57(1–3):562, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902932.

Bose:1983:CCO

[Bos83] A. K. Bose. Computation of Casimir operator eigenvalues. Computer Physics Communications, 28(3):271–274, January 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900449.

Bose:1984:CCO

[Bos84] A. K. Bose. Computation of Casimir operator eigenvalues. Computer Physics Communications, 35(1–3):C–851, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829604.

Bouanich:1986:AVM

[BOT86] J. P. Bouanich, J. F. Ogilvie, and R. H. Tipping. Analytic vibrational matrix elements for diatomic molecules. Computer Physics Communications, 39(3):439–446, April 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901001.

Bouanich:1987:AVR

[Bou87] J. P. Bouanich. Analytic vibration-rotational matrix elements for diatomic molecules. Computer Physics Communications, 47(2–3):259–266, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901123.

Baylis:1982:PCF

[BP82a] W. E. Baylis and S. J. Peel. A predictor-corrector form of the Numerov method for coupled equations. Computer Physics Communications, 25(1):7–19, January 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558290039X.

Baylis:1982:PPC

[BP82b] W. E. Baylis and S. J. Peel. A program for the predictorcorrector Numerov method. Computer Physics Communications, 25(1):21–28, January 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900406.

Baylis:1984:PPC

[BP84a] W. E. Baylis and S. J. Peel. A program for the predictor– corrector Numerov method. Computer Physics Communications, 35(1–3):C–762, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828921.

Baylis:1984:LMH

[BP84b] W. E. Baylis and Atul D. Pradhan. Locating minima of a hypersurface: a modification of the Levenberg–Marquardt method. Computer Physics Communications, 31(4):297–301, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900146.

Bleris:1984:PSS

[BP84c] G. L. Bleris and Ch. Polatoglou. A program for the study of short range order of binary alloys. Computer Physics Communications, 35(1–3):C–413, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826223.

Bradly:1984:CCC

[BP84d] D. L. Bradly and R. Perrin. Colour coordinate calculations. Computer Physics Communications, 35(1–3):C–312, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825382.

Bent:1984:VPA

[BPA84] Michael F. Bent, Börje I. Persson, and David G. Agresti. Versatile program for analysis of Mössbauer spectra. Computer Physics Communications, 35(1–3):C–11, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482281X.

Bohr:1986:PCL

[BPR^{+}86] H. Bohr, T. Petersen, B. Rathjen, E. Katznelson, and A. Nobile. Parallel computations of lattice models in physics. Computer Physics Communications, 42(1):11–19, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586902249.

Bok:1988:UCC

[BPSB88] J. Bok, P. Praus, J. Stepanek, and V. Baumruk. A universal computer-controlled UV-VIS spectrometer with high resolution monochromator. Computer Physics Communications, 50(1–2):225–228, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901282.

Banks:1984:CMP

[BPW84a] D. Banks, I. C. Percival, and J. McB. Wilson. Classical motion of 2 particles (eva2 edition 01). Computer Physics Communications, 35(1–3):C–137, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823917.

Banks:1984:CRMa

[BPW84b] D. Banks, I. C. Percival, and J. McB. Wilson. Classical relative motion of 2 particles. Computer Physics Communications, 35(1–3):C–71, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482336X.

Banks:1984:CRMb

[BPW84c] D. Banks, I. C. Percival, and J. McB. Wilson. Classical relative motion of 2 particles (evar edition 02). Computer Physics Communications, 35(1–3):C–136, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823905.

Birss:1985:CAA

[BR85] F. W. Birss and D. A. Ramsay. Computer assistance in the analysis of molecular spectra: I. Rotational structure of high resolution singlet-singlet bands. Computer Physics Communications, 38(1):83–112, August/September 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900487.

Braithwaite:1984:ALP

[Bra84a] W. J. Braithwaite. Associated Legendre polynomials, ordinary and modified spherical harmonics. Computer Physics Communications, 35(1–3):C–192, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824376.

Braithwaite:1984:RKTb

[Bra84b] W. J. Braithwaite. Relativistic kinematics for three-body final states. Computer Physics Communications, 35(1–3):C– 155, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824066.

Braithwaite:1984:RKTa

[Bra84c] W. J. Braithwaite. Relativistic kinematics for two-body final states. Computer Physics Communications, 35(1–3):C–153–C– 154, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824054.

Braunschweig:1984:IRM

[Bra84d] D. Braunschweig. II. Reduced SU(3) matrix elements. Computer Physics Communications, 35(1–3):C–502–C–503, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826880.

Braunschweig:1984:RC

[Bra84e] D. Braunschweig. Reduced SU(3) CFP’s. Computer Physics Communications, 35(1–3):C–455, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826582.

Bransden:1986:BRB

[Bra86] B. H. Bransden. Book review: Vector and Parallel Processors in Computational Science: Proceedings of the 2nd International Conference, Oxford, 28–31 August 1984. Eds. I. S. Duff and J. K. Reid, North-Holland, Amsterdam, 1985. 386 pages. Dfl. 225.00. Computer Physics Communications, 41(1):193, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900305.

Brackbill:1987:MAM

[Bra87a] J. U. Brackbill. On modelling angular momentum and vorticity in compressible fluid flow. Computer Physics Communications, 47(1):1–16, October 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900634.

Brandt:1987:NNC

[Bra87b] Lawrence E. Brandt. The NSF national computing centers: Past, present, and future. Computer Physics Communications, 45(1–3):147–148, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587901494.

Bhattacharya:1989:RLT

[BRB89] Ranjan Bhattacharya, Dhiranjan Roy, and Siddhartha Bhowmick. On the regularity of the Levin U-transform. Computer Physics Communications, 55(3):297–301, October 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901276.

Brix:1986:FSC

[Bri86] H. Brix. Formal specification and construction techniques for the development of programs with guaranteed properties. Computer Physics Communications, 41(2–3):245–257, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900688.

Bromage:1983:QHF

[Bro83] G. E. Bromage. A quantification of the hazards of fitting sums of exponentials to noisy data. Computer Physics Communications, 30(3):229–233, November 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900899.

Brody:1984:RNG

[Bro84] T. A. Brody. A random-number generator. Computer Physics Communications, 34(1–2):39–46, November/December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901577.

Brody:1989:RNG

[Bro89] T. A. Brody. Random-number generation for parallel processors. Computer Physics Communications, 56(2):147–153, December 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900167.

Banuelos:1984:PCA

[BRT84] Alicia Bañuelos and Felix Rodriguez-Trelles. A program for calculating the angular distribution of nonrelativistic Bremsstrahlung intensity. Computer Physics Communications, 35(1–3):C–489, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826806.

Banuelos:1984:EHF

[BRTB84] Alicia Bañuelos, Felix Rodriguez-Trelles, and Luis Bilbao. Extension to high frequencies of a program for calculating the angular distribution of nonrelativistic Bremsstrahlung. Computer Physics Communications, 35(1–3):C–556, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827319.

Brut:1985:FSO

[Bru85] F. Brut. A function subprogram in order to calculate the matrix elements of rotation operators. Computer Physics Communications, 36(2):213–217, April 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901249.

Brut:1986:FVF

[Bru86] F. Brut. A FORTRAN 77 version of “A function subprogram in order to calculate the matrix elements of rotation operators”. Computer Physics Communications, 39(2):297–299, February/ March 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586901396.

Bell:1980:CFN

[BS80] K. L. Bell and N. S. Scott. Coulomb functions (negative energies). Computer Physics Communications, 20(3):447–458, November 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900223.

Bartschat:1983:ASE

[BS83] K. Bartschat and N. S. Scott. Amplitudes for scattering of electrons by atomic systems including relativistic effects.

Computer Physics Communications, 30(4):369–381, December 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900784.

Bartschat:1984:ASE

[BS84a] K. Bartschat and N. S. Scott. Amplitudes for scattering of electrons by atomic systems including relativistic effects. Computer Physics Communications, 35(1–3):C–923, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584830210.

Bell:1984:CFN

[BS84b] K. L. Bell and N. S. Scott. Coulomb functions (negative energies). Computer Physics Communications, 35(1–3):C–648, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828088.

Brocklehurst:1984:CLC

[BS84c] M. Brocklehurst and M. Salem. Computation of line and continuum radiation from thermal radio astronomical sources. Computer Physics Communications, 35(1–3):C–306–C–307, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825345.

Brocklehurst:1984:RRL

[BS84d] M. Brocklehurst and M. Salem. Radio recombination lines from H^{+} regions and cold interstellar clouds-computation of the b_{n} factors. Computer Physics Communications, 35(1–3):C–411– C–412, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584826211.

Boyd:1986:FUP

[BS86a] K. M. H. Boyd and N. S. Scott. FTIDY — a utility program for FORTRAN 77 programs. Computer Physics Communications, 39(3):421–430, April 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900986.

Bunk:1986:PRM

[BS86b] B. Bunk and R. Sommer. An 8 parameter representation of SU(3) matrices and its application for simulating lattice QCD. Computer Physics Communications, 40(2–3):229–232, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901116.

Bengtsson:1987:LMC

[BS87] Hans-Uno Bengtsson and Torbjörn Sjöstrand. The lund Monte Carlo for hadronic processes — PYTHIA version 4.8. Computer Physics Communications, 46(1):43–82, July 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900361.

Beu:1986:GCM

[BSA86] T. A. Beu, D. I. Simionovici, and V. N. Anghel. Grenade — a coarse-mesh reactor physics program to solve the static diffusion equation for neutrons. Computer Physics Communications, 42(2):197–216, October 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900378.

Bermejo:1987:SFC

[BSdlT87] F. J. Bermejo, J. Santoro, and L. Sainz de los Terreros. SQSIMUL: a FORTRAN code for the computation of squeezing properties and photon statistics in multiphoton processes. Computer Physics Communications, 43(2):245–256, January 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587902098.

Bhanot:1988:FVP

[BSDM88] Gyan Bhanot, Román Salvador, Dennis Duke, and K. J. M. Moriarty. A fast vectorized program for the CDC Cyber 205 to simulate the Ising spin glass in three dimensions. Computer Physics Communications, 49(3):465–474, June 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900070.

Bar-Shalom:1988:NEP

[BSK88] A. Bar-Shalom and M. Klapisch. NJGRAF — an efficient program for calculation of general recoupling coefficients by graphical analysis, compatible with NJSYM. Computer Physics Communications, 50(3):375–393, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901920.

Bastien:1984:TMP

[BSP84] P. L. Bastien, J. N. Snyder, and V. Pless. A track matching program for bubble chamber photographs. Computer Physics Communications, 35(1–3):C–106, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823644.

Beu:1985:TTP

[BSV^{+}85] T. A. Beu, F. Spineanu, M. Vlad, R. I. Câmpeanu, and I. I. Popescu. Topic — a tokamak plasma impurities code. Computer Physics Communications, 36(2):161–176, April 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901213.

Barnsley:1981:BMN

[BT81] M. Barnsley and G. Turchetti. Bounding methods in nuclear physics. Computer Physics Communications, 22(2–3):325–343, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900680.

Berrington:1982:ACC

[BT82] K. A. Berrington and K. T. Taylor. Atomic collision calculations. Computer Physics Communications, 26(3–4):397– 410, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901345.

Boettger:1984:MIC

[BT84a] J. C. Boettger and S. B. Trickey. Multiparameter iterative convergence accelerator for crystalline LCGTO calculations. Computer Physics Communications, 32(4):361–365, July 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900535.

Burke:1984:PCR

[BT84b] P. G. Burke and C. Tate. A program for calculating Regge trajectories in potential scattering. Computer Physics Communications, 35(1–3):C–13, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822833.

Brecht:1988:MSU

[BT88] Stephen H. Brecht and Vincent A. Thomas. Multidimensional simulations using hybrid particles codes. Computer Physics Communications, 48(1):135–143, January 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900318.

Biddulph:1989:IMI

[BT89] Phillip Biddulph and Graham Thompson. Improved modelling of independent parton hadronization. Computer Physics Communications, 54(1):13–21, April 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900283.

Brandt:1984:PAC

[BTS84a] Maynard A. Brandt, Donald G. Truhlar, and Richard L. Smith. Program ACRL to calculate differential and integral cross sections adapted to run on IBM computers. Computer Physics Communications, 35(1–3):C–234, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824728.

Brandt:1984:PCD

[BTS84b] Maynard A. Brandt, Donald G. Truhlar, and Richard L. Smith. Program for calculating differential and integral cross sections for quantum mechanical scattering problems from reactance or transition matrices. Computer Physics Communications, 35(1–3):C–201–C–202, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824443.

Badavi:1987:ASN

[BTWN87] Forooz F. Badavi, Lawrence W. Townsend, John W. Wilson, and John W. Norbury. An algorithm for a semiempirical nuclear fragmentation model. Computer Physics Communications, 47(2–3):281–294, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901147.

Buckley:1983:CCC

[Buc83] A. Buckley. CONCON: a code for constrained minimization without derivative computation. Computer Physics Communications, 29(3):231–236, May 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900036.

Buckley:1984:CCC

[Buc84] A. Buckley. Concon: a code for constrained minimization without derivative computation. Computer Physics Communications, 35(1–3):C–876, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482980X.

Buckle:1985:MTS

[Buc85] J. K. Buckle. Methods and tools for system design. Computer Physics Communications, 38(2):165–171, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900827.

Bunz:1986:ICC

[Bun86] H. Bunz. Identification of clusters in computer experiments with periodic boundary conditions. Computer Physics Communications, 42(3):435–439, November 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900111.

Burke:1984:PCG

[Bur84] P. G. Burke. A program to calculate a general recoupling coefficient. Computer Physics Communications, 35(1–3):C–30, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822985.

Burke:1985:E

[Bur85a] P. G. Burke. Editorial. Computer Physics Communications, 36 (1):ix, March 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558590013X.

Burton:1985:SFE

[Bur85b] C. G. Burton. The solution of finite element equations on the Floating Point Systems FPS-164 Attached Processor. Computer Physics Communications, 37(1–3):171–180, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558590150X.

Burnett:1987:IID

[Bur87] T. H. Burnett. IDA: An interactive data analysis environment for high energy physics. Computer Physics Communications, 45(1–3):195–199, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901561.

Buttle:1984:DPH

[But84] P. J. A. Buttle. DWBA program for heavy ion transfer reactions. Computer Physics Communications, 35(1–3):C–458, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826600.

Butler:1989:PNA

[But89] J. W. Butler. A proposed new approach to the expansion of mathematical functions. Computer Physics Communications, 54(2–3):221–234, June/July 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900842.

Buzbee:1985:AMM

[Buz85] B. L. Buzbee. Applications of MIMD machines. Computer Physics Communications, 37(1–3):1–5, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901304.

Bakker:1980:PSR

[BV80] M. Bakker and M. S. Van Den Berg. A program to solve rotating plasma problems. Computer Physics Communications, 20(3):429–439, November 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558090020X.

Bakker:1984:PSR

[BV84] M. Bakker and M. S. Van Den Berg. A program to solve rotating plasma problems. Computer Physics Communications, 35(1–3): C–646, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828064.

Blake:1985:DPM

[BvEF85] Julian Blake, Horst von Eicken, and David Foster. Developing programs for the Motorola 68000 microprocessor at CERN. Computer Physics Communications, 38(2):315–318, October/ November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585900967.

Belic:1984:NEA

[BVL84] Milivoj R. Belić, Slobodan Vuković, and Melvin Lax. New efficient algorithm for solution of the driven nonlinear Schrödinger equation. Computer Physics Communications, 32(3):239–243, June 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900882.

Brandt:1984:MCR

[BvSW84] J. Brandt, A. von Scholley, and M. Wochner. Making chemical reaction data accessible to the non-chemist. Computer Physics Communications, 33(1–3):197–203, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901218.

Burgess:1987:BPE

[BW87] Alan Burgess and Colm T. Whelan. Betrt — a procedure to evaluate the cross section for electron-hydrogen collisions in the Bethe approximation to the reactance matrix. Computer Physics Communications, 47(2–3):295–304, November/ December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901159.

Bacic:1988:LRL

[BWBL88] Z. Bacić, R. M. Whitnell, D. Brown, and J. C. Light. Localized representations for large amplitude molecular vibrations. Computer Physics Communications, 51(1–2):35–47, September/October 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588900604.

Bhattacharjee:1984:VMT

[BWD84] A. Bhattacharjee, J. C. Wiley, and R. L. Dewar. Variational method for three-dimensional toroidal equilibria. Computer Physics Communications, 31(2–3):213–225, February 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900468.

Beck:1984:RNR

[BZ84] Donald R. Beck and Richard N. Zare. Relativistic and nonrelativistic configuration interaction calculations for atoms having a closed core and two valence spin-orbitals. Computer Physics Communications, 35(1–3):C–15–C–16, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822857.

Canut-Amoros:1984:SCP

[CA84] Marisa Canut-Amoros. Stlplt — Calcomp plot of crystallographic projections of Laue photographs. Computer Physics Communications, 35(1–3):C–35, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823036.

Campbell:1981:BFR

[Cam81] J. B. Campbell. Bessel functions I_{ν}(z) and K_{ν}(z) of real order and complex argument. Computer Physics Communications, 24(1):97–105, September/October 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901090.

Campeanu:1982:PPC

[Câm82] R. I. Câmpeanu. POSDIF — a program to compute positron diffusion and annihilation in rare gases. Computer Physics Communications, 25(4):433–436, April 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900790.

Campbell:1984:BFRb

[Cam84a] J. B. Campbell. Bessel functions I_{ν}(z) and K_{ν}(z) of real order and complex argument. Computer Physics Communications, 35 (1–3):C–747–C–748, 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Campbell:1984:BFZ

[Cam84b] J. B. Campbell. Bessel functions I_{ν}(z) and K_{ν}(z) of real order and complex argument. Computer Physics Communications, 35(1–3):C–747–C–748, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828805.

Campbell:1984:BFRa

[Cam84c] J. B. Campbell. Bessel functions J_{ν}(x) of real order and real argument. Computer Physics Communications, 35(1–3): C–583, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827563.

Campbell:1984:DZH

[Cam84d] J. B. Campbell. Determination of ν-zeros of Hankel functions. Computer Physics Communications, 32(3):333–339, June 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558490095X.

Campbell:1984:DSS

[Cam84e] J. B. Campbell. Determination of SSOR-SI iteration parameters. Computer Physics Communications, 35(1–3):C–339, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825618.

Campbell:1984:PPD

[Cam84f] J. B. Campbell. A program package for the Dirichlet problem with axially symmetric boundary conditions. Computer Physics Communications, 35(1–3):C–309, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825369.

Campeanu:1984:PPC

[Câm84g] R. I. Câmpeanu. Posdif — a program to compute positron diffusion and annihilation in rare gases. Computer Physics Communications, 35(1–3):C–800, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829215.

Cameron:1985:DSM

[Cam85] J. R. Cameron. Developing systems methodically. Computer Physics Communications, 38(2):149–163, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900815.

Campbell:1986:KBC

[Cam86] J. A. Campbell. Knowledge-based computing systems and software engineering. Computer Physics Communications, 41(2–3):285–290, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900718.

Candel:1981:AFB

[Can81] S. M. Candel. An algorithm for the Fourier–Bessel transform. Computer Physics Communications, 23(4):343–353, August 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901752.

Cap:1986:EEA

[Cap86a]  Ferdinand F. Cap.  Electromagnetic eigenfrequency of

anisotropic inhomogeneous axisymmetric toroidal plasmas of arbitrary meridional cross section. Computer Physics Communications, 40(1):99–103, May 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901517.

Christodoulides:1986:ALS

[CAP86b] C. Christodoulides, L. Apekis, and P. Pissis. An algorithm for least-squares curve-fitting of TL and TSDC peaks. Computer Physics Communications, 41(1):35–39, July 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900196.

Carl:1980:OBD

[Car80] P. Carl. OLYMPUS for a BESM-6 — the development of medium-scale programs within the OLYMPUS framework: system OLYDES. Computer Physics Communications, 20 (1):165–179, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580901216.

Carver:1984:FVP

[Car84a] M. B. Carver. FORSIM VI: a program package for the automated solution of arbitrarily defined differential equations. Computer Physics Communications, 35(1–3):C–552–C– 553, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827289.

Christiansen:1984:MOD

[CAR84b] J. P. Christiansen, D. E. T. F. Ashby, and K. V. Roberts. Medusa: a one-dimensional laser fusion code. Computer Physics Communications, 35(1–3):C–245–C–246, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824820.

Carpenter:1988:PGS

[Car88] Lesley Ann Carpenter. Portable graphical software — its design, implementation and use. Computer Physics Communications, 50(1–2):159–168, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901233.

Carter:1989:FVP

[Car89] Paul A. Carter. A fast vectorized program for the Cyber 205 to calculate the partition function of the 3-D Ising model. Computer Physics Communications, 54(1):103–108, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900362.

Catchen:1989:PFP

[Cat89] Gary L. Catchen. Perturbation functions: PAC probe nuclei, I = 2, 5/2, and 3. Computer Physics Communications, 55(1):85–90, August 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900660.

Clary:1988:SPW

[CB88] D. C. Clary and A. J. Banks. Summing the partialwave series in molecular collision calculations. Computer Physics Communications, 48(2):223–228, February 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900422.

Ciulli:1984:GAE

[CC84] M. Ciulli and S. Ciulli. A guide to analytic extrapolations: Part I: A program for optimal extrapolation to interior points. Computer Physics Communications, 35(1–3):C–589–C–590, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827617.

Carbo:1989:AIM

[CC89a] Ramon Carbó and Blanca Calabuig. Ariadne-88: An ab initio monoconfigurational closed and open shell direct electronic energy calculation using elementary Jacobi rotations.

Computer Physics Communications, 52(3):345–354, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901100.

Carbo:1989:MMS

[CC89b] Ramon Carbó and Blanca Calabuig. Molsimil-88: Molecular similarity calculations using a CNDO-like approximation. Computer Physics Communications, 55(1):117–126, August 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900702.

Clementi:1985:PCQ

[CCD85] E. Clementi, G. Corongiu, and J. H. Detrich. Parallelism in computations in quantum and statistical mechanics. Computer Physics Communications, 37(1–3):287–294, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558590164X.

Chernysheva:1984:FCH

[CCR84a] L. V. Chernysheva, N. A. Cherepkov, and V. Radojević. Frozen core Hartree–Fock program for atomic discrete and continuous states. Computer Physics Communications, 35(1–3):C–578–C– 579, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827526.

Chernysheva:1984:SCF

[CCR84b] L. V. Chernysheva, N. A. Cherepkov, and V. Radojević. Self-consistent field Hartree–Fock program for atoms. Computer Physics Communications, 35(1–3):C–361–C–362, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825801.

Chang:1988:ETS

[CCR88] Tom Chang, G. B. Crew, and J. M. Retterer. Electromagnetic tornadoes in space: Ion conics along auroral field lines generated by lower hybrid waves and electromagnetic turbulence in the ion cyclotron range of frequencies. Computer Physics Communications, 49(1):61–74, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902159.

Colangelo:1989:CSG

[CCS89] P. Colangelo, L. Cosmai, and E. Scrimieri. Computer simulation of gauge theories on a lattice with improved rotational symmetry. Computer Physics Communications, 54(2–3):235–237, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900854.

Clancy:1986:CWF

[CD86] B. E. Clancy and I. J. Donnelly. The calculation of wave fields in a cylindrical plasma using a multiple shooting technique. Computer Physics Communications, 42(2):153–167, October 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900330.

Cittolin:1989:TLT

[CDF^{+}89] S. Cittolin, M. Demoulin, A. Fucci, W. Haynes, B. Martin, J. P. Porte, and P. Sphicas. The third level trigger and output event unit of the UA1 data-acquisition system. Computer Physics Communications, 57(1–3):370–374, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902452.

Castanos:1988:QRR

[CDL88] O. Castaños, J. P. Draayer, and Y. Leschber. Quantum rotor and its SU(3) realization. Computer Physics Communications, 52(1):71–84, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901749.

Cwiok:1987:SPE

[CDN^{+}87] S. Cwiok, J. Dudek, W. Nazarewicz, J. Skalski, and T. Werner. Single-particle energies, wave functions, quadrupole moments and g-factors in an axially deformed Woods–Saxon potential with applications to the two-centre-type nuclear problems. Computer Physics Communications, 46(3):379–399, September 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900932.

Chang:1982:SGF

[CDW82] B. D. Chang, J. P. Draayer, and S. S. M. Wong. A system to generate Fortran programs for calculating configuration traces of angular momentum coupled product operators. Computer Physics Communications, 28(1):41–60, November 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900613.

Chang:1984:SGF

[CDW84] B. D. Chang, J. P. Draayer, and S. S. M. Wong. A system to generate Fortran programs for calculating configuration traces of angular momentum coupled product operators. Computer Physics Communications, 35(1–3):C–833–C–834, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829458.

Coghlan:1985:DPS

[CF85] Brant Coghlan and Serafin Fraga. Determination of proteinic structures: An experimentation program. Computer Physics Communications, 36(4):391–399, June 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900281.

Carpenter:1987:TYR

[CFG^{+}87] B. E. Carpenter, F. Fluckiger, J. M. Gerard, D. Lord, and B. Segal. Two years of real progress in European HEP networking: a CERN perspective. Computer Physics Communications, 45(1–3):83–92, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901433.

Castellano:1988:MCP

[CFG^{+}88] M. Castellano, C. Favuzzi, N. Giglietto, E. Nappi, and P. Spinelli. A Monte Carlo program to design a transition radiation detector. Computer Physics Communications, 51(3):431–441, November 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901567.

Chiu:1987:SRS

[CG87] Ting-Wai Chiu and Tian-Shin GUU. A shift-register sequence random number generator implemented on the microcomputers with 8088/8086 and 8087. Computer Physics Communications, 47(1):129–137, October 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900725.

Corden:1989:PTV

[CGB^{+}89] M. J. Corden, C. H. Georgiopoulos, R. Brun, F. Bruyant, and J.-L. Dekeyser. Progress towards a vectorized version of the giant Monte Carlo program. Computer Physics Communications, 57(1–3):268–272, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902269.

Chowdhury:1987:MCS

[CGG87] Debashish Chowdhury, E. T. Gawlinski, and J. D. Gunton. Monte Carlo simulation of very large dilute randomfield Ising models using multi-spin coding. Computer Physics Communications, 43(3):329–337, February/March 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900506.

Celmaster:1985:VMC

[CGGK85] W. Celmaster, F. Green, R. Gupta, and E. Kovacs. A vectorized Monte Carlo algorithm for computing Wilson line observables in Su(2) gauge theory on a BCH lattice. Computer Physics Communications, 36(4):409–415, June 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558590030X.

Carter:1989:TUT

[CGM89] J. M. Carter, M. G. Green, and T. Medcalf. Transparent use of transputers for off-line computation. Computer Physics Communications, 57(1–3):495–498, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902713.

Corden:1989:IAD

[CGMS89] M. J. Corden, C. H. Georgiopoulos, M. E. Mermikides, and J. Streets. Implementation of the ALEPH detector simulation code using UNIX with on-line graphics display. Computer Physics Communications, 57(1–3):260–262, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902245.

Cheng-guang:1985:PCT

[CgYpXh85] Bao Cheng-guang, Gan You-ping, and Lui Xian-hui. Program to calculate transformation brackets of hyperspherical harmonic functions of a three body system. Computer Physics Communications, 36(4):401–407, June 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900293.

Christiansen:1984:DTD

[CH84a] J. P. Christiansen and R. W. Hockney. Delsqphi, a twodimensional Poisson-solver program. Computer Physics Communications, 35(1–3):C–74–C–75, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823383.

Christiansen:1984:FFF

[CH84b] J. P. Christiansen and R. W. Hockney. Four67, a Fast Fourier Transform package. Computer Physics Communications, 35 (1–3):C–72–C–73, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823371.

Corbould:1984:NAL

[CH84c] M. A. Corbould and J. A. How. A network approach to large-scale experimental data acquisition and analysis. Computer Physics Communications, 32(3):231–237, June 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900870.

Cuperman:1984:NCM

[CH84d] S. Cuperman and A. Harten. A numerical code for multiple “water bag” gravitational systems. Computer Physics Communications, 35(1–3):C–279, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825102.

Carter:1987:MDE

[CH87] S. Carter and N. C. Handy. A method for the determination of the eigenvalues of a very large matrix: Application to vibrational energy levels. Computer Physics Communications, 44(1–2):1–9, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900129.

Carter:1988:VMD

[CH88] S. Carter and N. C. Handy. A variational method for the determination of the vibrational ( J = 0) energy levels of acetylene, using a Hamiltonian in internal coordinates. Computer Physics Communications, 51(1–2):49–58, September/October 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900616.

Chamayou:1980:ASH

[Cha80] J. M. F. Chamayou. Averaging shifted histograms. Computer Physics Communications, 21(2):145–161, December 2, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900375.

Chamayou:1982:PBS

[Cha82] J. M. F. Chamayou. Program for B-spline interpolation of surfaces with application to computer tomography. Computer Physics Communications, 27(2):187–199, August 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558290073X.

Chamayou:1984:AKR

[Cha84a] J. M. F. Chamayou. Algorithms for the Kac and Renyi tests. Computer Physics Communications, 35(1–3):C–395, ????

1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826065.

Chamayou:1984:SMR

[Cha84b] J. M. F. Chamayou. On the statistical model of radiation damage with a random displacement energy. Computer Physics Communications, 32(1):29–33, April 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900055.

Chamayou:1984:PSI

[Cha84c] J. M. F. Chamayou. Program for B-spline interpolation of surfaces with application to computer tomography. Computer Physics Communications, 35(1–3):C–816, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829331.

Chandra:1984:GPS

[Cha84d] N. Chandra. A general program to study the scattering of particles by solving coupled inhomogeneous second-order differential equations. Computer Physics Communications, 35(1–3): C–196, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824418.

Chang:1984:PER

[Cha84e] J. J. Chang. A program to evaluate the reduced matrix elements of one-particle tensor operators for the configurations in JJ-coupling. Computer Physics Communications, 35(1–3):C– 241, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824790.

Chauvet:1985:MPV

[Cha85] Yves Chauvet. Measured performances on vectorization and multitasking with a Monte Carlo code for neutron transport problems. Computer Physics Communications, 37(1–3):281– 285, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585901638.

Chen:1989:FSS

[Che89] He Sheng Chen. Fast simulation of showers in inhomogeneous media. Computer Physics Communications, 57(1–3):375–379, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589902464.

Chivers:1984:NVPa

[Chi84a] Alfred T. Chivers. A new version of the program to compute the asymptotic solution of coupled equations for electron scattering. Computer Physics Communications, 35(1–3):C–195, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824406.

Chivers:1984:NVPb

[Chi84b] Alfred T. Chivers. A new version of the program to compute the fractional parentage coefficients for equivalent d shell electrons. Computer Physics Communications, 35(1–3):C–210, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824522.

Chilingarian:1989:SDU

[Chi89] A. A. Chilingarian. Statistical decisions under nonparametric a priory information. Computer Physics Communications, 54(2–3):381–390, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900982.

Cutts:1989:DAD

[CHJZ89] David Cutts, Jan S. Hoftun, Christopher R. Johnson, and Raymond T. Zeller. Data acquisition at D0. Computer Physics Communications, 57(1–3):339–342, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902397.

Chu:1981:CDN

[CHL^{+}81] M. S. Chu, F. J. Helton, J. K. Lee, R. W. Moore, J. M. Greene, and T. H. Jensen. Current-driven nonideal instability in a force-free toroidal plasma. Computer Physics Communications, 24(3–4):399–406, December 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901636.

Collins:1984:EPW

[CHMM84] P. A. Collins, B. J. Hartley, R. W. Moore, and K. J. M. Moriarty. An efficient partial-wave analyser for the absorption model. Computer Physics Communications, 35(1–3):C–188–C– 189, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824340.

Choudry:1984:IRT

[Cho84] A. Choudry. Isometric representation of two-dimensional matrices. Computer Physics Communications, 35(1–3):C–32, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823000.

Christensen:1984:CSP

[Chr84a] E. R. Christensen. Computer simulation of photons in spheric media for density gauges. Computer Physics Communications, 35(1–3):C–236, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824753.

Christensen:1984:MCS

[Chr84b] E. R. Christensen. Monte Carlo simulation of photons in twolayered media for density gauges. Computer Physics Communications, 35(1–3):C–235, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824741.

Chadwick:1987:OMC

[CHS87] K. Chadwick, R. Hollebeek, and P. K. Sinervo. The organization and maintenance of the CDF offline code on IBM VM/CMS and DEC VAX/VMS operating systems. Computer Physics Communications, 45(1–3):409–415, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901822.

Cutts:1989:NNE

[CHS^{+}89] Dave Cutts, Jan S. Hoftun, Andrew Sornborger, Christopher R. Johnson, and Raymond T. Zeller. Neural networks for event filtering at D0. Computer Physics Communications, 57(1–3):478–482, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902671.

Cioslowski:1989:PIS

[Cio89] J. Cioslowski. Principles of AB initio SCF calculations with minimal storage requirements. Computer Physics Communications, 53(1–3):117–122, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901525.

Cisneros:1987:CSI

[Cis87] Arturo Cisneros. A communications system for irregular local interaction problems on a concurrent computer. Computer Physics Communications, 46(1):35–41, July 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558790035X.

Clarkson:1984:ELS

[CJ84] R. G. Clarkson and Nelson Jarmie. Energy-loss straggling of heavy charged particles. Computer Physics Communications, 35(1–3):C–109, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482367X.

Chen:1989:ISB

[CKA89] Y.-M. Chen, A. E. Koniges, and D. V. Anderson. ILUBCG2-11: Solution of 11-banded nonsymmetric linear equation systems by a preconditioned biconjugate gradient routine. Computer Physics Communications, 55(3):359–365, October 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990132X.

Chan:1989:PEP

[CKT89] Tony F. Chan, C.-C. Jay Kuo, and Charles Tong. Parallel elliptic preconditioners: Fourier analysis and performance on the connection machine. Computer Physics Communications, 53(1–3):237–252, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558990163X.

Cytron:1985:ERC

[CKV85] Ron Cytron, David J. Kuck, and Alex V. Veidenbaum. The effect of restructing compilers on program performance for highspeed computers. Computer Physics Communications, 37(1–3): 39–48, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585901341.

Cullum:1989:GNL

[CKW89] Jane Cullum, Wolfgang Kerner, and Ralph Willoughby. A generalized nonsymmetric Lanczos procedure. Computer Physics Communications, 53(1–3):19–48, May 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990146X.

Calinon:1984:AGB

[CL84] R. Calinon and J. Ligou. Application of the generalized backward substitution method to solve a class of linear systems. Computer Physics Communications, 35(1–3):C–558, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827332.

Clarke:1982:MPA

[Cla82a] N. M. Clarke. Multi-point algorithms for second-order coupled equations. Computer Physics Communications, 27(4):365–375, October 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900984.

Claudius:1982:SGP

[Cla82b] M. Claudius. Solution of the gravitational Poisson equation in spherical coordinates. Computer Physics Communications, 27(2):119–128, August 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900686.

Clark:1984:ICC

[Cla84] Robert E. H. Clark. Intermediate coupling collision strengths from LS coupled R-matrix elements. Computer Physics Communications, 35(1–3):C–525, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S001046558482706X.

Clark:1988:EH

[Cla88] R. A. Clark. The evolution of HOBO. Computer Physics Communications, 48(1):61–64, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465588900239.

Clint:1982:VMD

[Cli82] M. Clint. A vector-multiplication dominated parallel algorithm for the computation of real eigenvalue spectra. Computer Physics Communications, 26(3–4):373–376, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558290131X.

Cohen:1989:RGM

[CLL89] Robert Cohen, San-Yih Lin, and Mitchell Luskin. Relaxation and gradient methods for molecular orientation in liquid crystals. Computer Physics Communications, 53(1–3):455– 465, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901781.

Conneelly:1984:CPC

[CLS^{+}84] M. J. Conneelly, L. Lipsky, K. Smith, P. G. Burke, and R. J. W. Henry. A computer program for the calculation of electron scattering and photoionization cross sections of atomic systems with configuration (np)^{q}. Computer Physics Communications, 35(1–3):C–36–C–37, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823048.

Chang:1988:RVE

[CLS88] Bernard H. Chang, Jae Shin Lee, and Don Secrest. Rotationvibration eigenvalues and vectors. Computer Physics Communications, 51(1–2):195–205, September/October 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900719.

Christie:1981:CSA

[CM81a] M. A. Christie and K. J. M. Moriarty. A Chebyshev series approxiamtion to continuous functions. Computer Physics Communications, 23(2):145–152, July 1, 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900291.

Cuperman:1981:NCP

[CM81b] S. Cuperman and M. Mond. A numerical code for the phasespace boundary integration of water bag plasmas. Computer Physics Communications, 21(3):397–406, January 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900163.

Creutz:1983:IMM

[CM83] Michael Creutz and K. J. M. Moriarty. Implementation of the microcanonical Monte Carlo simulation algorithm for SU(N) lattice gauge theory calculations. Computer Physics Communications, 30(3):255–257, November 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900930.

Carswell:1984:ACP

[CM84a] A. Carswell and W. Moon. Algorithm for the creation of the P-tau and P-X planes from T-X plane data. Computer Physics Communications, 32(2):185–190, May 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900699.

Christie:1984:BSI

[CM84b] M. A. Christie and K. J. M. Moriarty. A bicubic spline interpolation of unequally spaced data. Computer Physics Communications, 35(1–3):C–563, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827381.

Christie:1984:CSA

[CM84c] M. A. Christie and K. J. M. Moriarty. A Chebyshev series approximation to continuous functions. Computer Physics Communications, 35(1–3):C–712–C–713, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828568.

Coleman:1984:VMA

[CM84d] John P. Coleman and Julie Mohamed. De Vogelaere’s method with automatic error control. Computer Physics Communications, 35(1–3):C–554, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827290.

Creutz:1984:IMM

[CM84e] Michael Creutz and K. J. M. Moriarty. Implementation of the microcanonical Monte Carlo simulation algorithm for SU(N) lattice gauge theory calculations. Computer Physics Communications, 35(1–3):C–914, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584830131.

Cuperman:1984:NCP

[CM84f] S. Cuperman and M. Mond. A numerical code for the phasespace boundary integration of water bag plasmas. Computer Physics Communications, 35(1–3):C–675, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828283.

Celmaster:1985:MCS

[CM85] W. Celmaster and K. J. M. Moriarty. Monte Carlo simulation of pure SU(2) gauge theory on a body-centered hypercubic lattice. Computer Physics Communications, 34(4):415–425, February 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900694.

Creutz:1986:FCT

[CM86] Michael Creutz and K. J. M. Moriarty. FORTRAN code for the three-dimensional Ising model. Computer Physics Communications, 39(2):173–180, February/March 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901281.

Chen:1988:TCM

[CM88] Q. Chen and B. D. Mestel. Taylor–Chirikov map package: a package of programs for the calculation of ordered periodic orbits of area preserving twist maps. Computer Physics Communications, 51(3):463–476, November 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901592.

Charity:1987:USE

[CMH87] T. Charity, R. McClatchey, and J. Harvey. Use of software engineering techniques in the design of the ALEPH data acquisition system. Computer Physics Communications, 45(1–3):433–441, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558790186X.

Creutz:1984:FAI

[CMM84] Michael Creutz, P. Mitra, and K. J. M. Moriarty. A fast algorithm for investigations on the three-dimensional Ising model. Computer Physics Communications, 33(4):361–366, October 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901425.

Creutz:1986:VTD

[CMO86] Michael Creutz, K. J. M. Moriarty, and M. O’Brien. Vectorization of the three-dimensional ISING model program on the CDC Cyber 205. Computer Physics Communications, 42(2):191–196, October 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900366.

Campostrini:1988:VCP

[CMPR88] Massimo Campostrini, Kevin J. M. Moriarty, Jean Potvin, and Claudio Rebbi. A vectorized code for the pseudofermion simulation of QCD with dynamical quarks. Computer Physics Communications, 50(3):395–411, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901932.

Campostrini:1986:VCM

[CMR86] M. Campostrini, K. J. M. Moriarty, and C. Rebbi. A vectorized code for the Monte Carlo computation of spin-dependent static potentials in QCD. Computer Physics Communications, 42(2):175–189, October 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900354.

Cole:1983:WLS

[CMS83] M. J. Cole, K. J. M. Moriarty, and P. E. Stolorz. Wilson loops, string tension and correlations in Monte Carlo simulation of compact U(1) lattice gauge theory. Computer Physics Communications, 30(4):421–428, December 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900838.

Cole:1984:WLS

[CMS84] M. J. Cole, K. J. M. Moriarty, and P. E. Stolorz. Wilson loops, string tension and correlations in Monte Carlo simulation of compact U(1) lattice gauge theory. Computer Physics Communications, 35(1–3):C–928, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558483026X.

Casilio:1984:AFV

[CN84] John M. Casilio and M. E. Noz. Analytic formulation of SU(3) vector coupling coefficients for n particles. Computer Physics Communications, 35(1–3):C–190, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824352.

Cowan:1987:UDC

[CN87] R. Cowan and W. R. Nelson. Use of 3-D color graphics with EGS. Computer Physics Communications, 45(1–3):485, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901962.

Cartledge:1982:CVA

[CNRS82] C. J. Cartledge, M. Narotam, N. D. H. Ross, and J. B. Slater. Costs of vector and array processor software development: a specific example from biological simulation. Computer Physics Communications, 26(3–4):247–248, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901126.

Chernov:1984:EAC

[CO84] N. I. Chernov and G. A. Ososkov. Effective algorithms for circle fitting. Computer Physics Communications, 33(4):329–333, October 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901371.

Cuperman:1986:MCA

[CO86] S. Cuperman and L. Ofman. Magnetic configurations for axisymmetric tandem mirror devices. Computer Physics Communications, 42(2):217–232, October 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/001046558690038X.

Coleman:1980:FSB

[Col80a] J. P. Coleman. A Fortran subroutine for the Bessel function J_{n}(x) of order 0 to 10. Computer Physics Communications, 21(1):109–118, December 1, 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900806.

Coleman:1980:NFO

[Col80b] John P. Coleman. A new fourth-order method for y′′ = g(x)y + r(x). Computer Physics Communications, 19(2):185–195, April/June 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900491.

Coleman:1984:FSB

[Col84a] J. P. Coleman. A Fortran subroutine for the Bessel function J_{n}(x) of order 0 to 10. Computer Physics Communications, 35(1–3):C–654, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828143.

Collocott:1984:NSK

[Col84b] S. J. Collocott. Numerical solution of Kramers–Kronig transforms by a Fourier method. Computer Physics Communications, 35(1–3):C–425, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482634X.

Colonna:1988:PSE

[Col88] Jean-François Colonna. Picture synthesis: An essential tool for numerical experimentation. Computer Physics Communications, 49(1):215–228, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588902299.

Comfort:1980:IPD

[Com80a] J. R. Comfort. Implementation of the partitioned-data-set concept for CDC computers. Computer Physics Communications, 19(1):51–61, January/March 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900661.

Comfort:1980:LFP

[Com80b] J. R. Comfort. Linkage of Fortran programs with partitioned data sets on IBM, CDC, and DEC computers. Computer Physics Communications, 19(1):43–49, January/March 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558090065X.

Comfort:1982:EMI

[Com82] J. R. Comfort. Evaluation of multipole integrals and the exponent range problem. Computer Physics Communications, 25(3):217–222, March 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900182.

Comfort:1984:IPD

[Com84] J. R. Comfort. Implementation of the partitioned-data-set concept for CDC computers. Computer Physics Communications, 35(1–3):C–608, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827757.

Cook:1982:DPC

[Coo82] J. Cook. DFPOT — a program for the calculation of double folded potentials. Computer Physics Communications, 25(2):125–139, February 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900297.

Cook:1984:DPC

[Coo84a] J. Cook. Dfpot — a program for the calculation of double folded potentials. Computer Physics Communications, 35(1–3): C–775, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829021.

Cook:1984:HOM

[Coo84b] J. Cook. Hermes — an optical model search program including tensor potentials for projectile spins 0 to 3/2. Computer Physics Communications, 31(4):363–383, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900201.

Copley:1981:MCC

[Cop81] J. R. D. Copley. Monte Carlo calculation of multiple scattering effects in thermal neutron scattering experiments: Improved computation of elastic coherent scattering intensities. Computer Physics Communications, 21(3):431–436, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900199.

Copley:1984:MCCa

[Cop84a] J. R. D. Copley. Monte Carlo calculation of multiple scattering effects in thermal neutron scattering experiments. Computer Physics Communications, 35(1–3):C–247, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824832.

Copley:1984:MCCd

[Cop84b] J. R. D. Copley. Monte Carlo calculation of multiple scattering effects in thermal neutron scattering experiments: Improved computation of elastic coherent scattering intensities. Computer Physics Communications, 35(1–3):C–678, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828313.

Copley:1984:MCCc

[Cop84c] J. R. D. Copley. Monte Carlo calculation of multiple scattering effects in thermal neutron scattering experiments: Modification to horizontal cylinder geometry. Computer Physics Communications, 35(1–3):C–292, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825229.

Copley:1984:MCCb

[Cop84d] J. R. D. Copley. Monte Carlo calculation of multiple scattering effects in thermal neutron scattering experiments: Modification to slab geometry. Computer Physics Communications, 35(1–3): C–291, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825217.

Cottrell:1987:ANS

[Cot87] R. L. A. Cottrell. Analysis of network statistics. Computer Physics Communications, 45(1–3):93–109, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901445.

Cutkosky:1981:SIS

[CP81] R. E. Cutkosky and C. Pomponiu. Spline interpolation and smoothing of data. Computer Physics Communications, 23(3):287–299, July 2, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900047.

Cutkosky:1984:SIS

[CP84] R. E. Cutkosky and C. Pomponiu. Spline interpolation and smoothing of data. Computer Physics Communications, 35 (1–3):C–730, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828659.

Chassapis:1989:MOO

[CPL89] C. S. Chassapis, D. G. Papageorgiou, and I. E. Lagaris. MCL — optimization oriented programming language. Computer Physics Communications, 52(2):223–239, January/February 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900064.

Caloz:1984:NAF

[CR84a] G. Caloz and J. Rappaz. On the numerical approximation of a free boundary problem related to MHD equilibria. Computer Physics Communications, 31(2–3):137–141, February 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900407.

Calvert:1984:MDF

[CR84b] L. D. Calvert and J. R. Rodgers. The metal data file. Computer Physics Communications, 33(1–3):93–98, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901115.

Cash:1984:HOM

[CR84c] J. R. Cash and A. D. Raptis. A high order method for the numerical integration of the one-dimensional Schrödinger equation. Computer Physics Communications, 33(4):299–304, October 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901358.

Christiansen:1984:OSC

[CR84d] J. P. Christiansen and K. V. Roberts. Olympus: a standard control and utility package for initial-value Fortran programs. Computer Physics Communications, 35(1–3):C–243–C– 244, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824819.

Catsaros:1987:NMI

[CR87] N. Catsaros and P. Ribon. A new method of interpolation and numerical integration. Computer Physics Communications, 43(3):339–346, February/March 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900518.

Crees:1980:APC

[Cre80a] M. A. Crees. ASYPCK, a program for calculating asymptotic solutions of the coupled equations of electron collision theory. Computer Physics Communications, 19(1):103–137, January/ March 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900703.

Crees:1980:EPA

[Cre80b] M. A. Crees. Editor, a program for amending files of card images. Computer Physics Communications, 19(1):139–157, January/March 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900715.

Crees:1981:AEV

[Cre81] M. A. Crees. ASYPCK2, an extended version of ASYPCK. Computer Physics Communications, 23(2):181–198, July 1, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900321.

Crees:1984:APC

[Cre84a] M. A. Crees. Asypck, a program for calculating asymptotic solutions of the coupled equations of electron collision theory. Computer Physics Communications, 35(1–3):C–612–C– 614, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827794.

Crees:1984:AEV

[Cre84b] M. A. Crees. ASYPCK2, an extended version of ASYPCK. Computer Physics Communications, 35(1–3):C–717–C–719, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828593.

Crees:1984:EPA

[Cre84c] M. A. Crees. Editor, A program for amending files of card images. Computer Physics Communications, 35(1–3):C–615, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827800.

Carter:1989:BCF

[CRH^{+}89] Stuart Carter, Pavel Rosmus, Nicholas C. Handy, Steven Miller, Jonathan Tennyson, and Brian T. Sutcliffe. Benchmark calculations of first principles rotational and ro-vibrational line strengths. Computer Physics Communications, 55(1):71–75, August 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900647.

Christiansen:1984:AODa

[CRL84] J. P. Christiansen, K. V. Roberts, and J. W. Long. Athene 1: a one-dimensional equilibrium-diffusion code. Computer Physics Communications, 35(1–3):C–479–C–480, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826752.

Collado:1989:RRS

[CRL^{+}89] JoséRamón Álvarez Collado, Jaime Fernández Rico, Rafael López, Miguel Paniagua, and Guillermo Ramírez. Rotation of real spherical harmonics. Computer Physics Communications, 52(3):323–331, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901070.

Crowley:1988:FLM

[Cro88] W. P. Crowley. A Free-Lagrange method for 2-D compressible flows. Computer Physics Communications, 48(1):51–60, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588900227.

Christiansen:1981:AOD

[CRPL81] J. P. Christiansen, K. V. Roberts, V. A. Piotrowicz, and J. W. Long. Athene 1A: a one-dimensional fusion code. Computer Physics Communications, 23(1):63–80, June 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901302.

Christiansen:1984:AODb

[CRPL84] J. P. Christiansen, K. V. Roberts, V. A. Piotrowicz, and J. W. Long. Athene 1A: a one-dimensional fusion code. Computer Physics Communications, 35(1–3):C–703–C–704, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828519.

Cabras:1987:UPO

[CRS87] G. Cabras, V. Roberto, and G. Salemi. UNIDFT: a package of optimized discrete Fourier transforms. Computer Physics Communications, 47(1):113–127, October 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900713.

Chandra:1989:AAF

[CRV^{+}89] U. Chandra, G. Riccardi, J. Vagi, J.-L. Dekeyser, and F. Hannedouche. Aftran: Array Fortran programming language. Computer Physics Communications, 57(1–3):263–267, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589902257.

Croskery:1982:VNA

[CSBB82] J. P. Croskery, N. S. Scott, K. L. Bell, and K. A. Berrington. VPM — a new asymptotic package. Computer Physics Communications, 27(4):385–401, October 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558290100X.

Croskery:1984:VNA

[CSBB84] J. P. Croskery, N. S. Scott, K. L. Bell, and K. A. Berrington. VPM — a new asymptotic package. Computer Physics Communications, 35(1–3):C–825, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829409.

Ciarcia:1986:DFP

[CSC^{+}86] C. A. Ciarcia, W. A. Schier, G. P. Couchell, D. J. Pullen, R. S. Tanczyn, M. H. Haghighi, and Q. Sharfuddin. DENTS: a FORTRAN program for analysing continuous neutron spectra. Computer Physics Communications, 39(2):233–243, February/March 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586901347.

Capotondi:1989:PSS

[CSC89] A. Capotondi, V. Sonnad, and S. Chin. Parallel solution of the shallow water equations using an explicit finite difference algorithm. Computer Physics Communications, 52(2):195–205, January/February 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900040.

Ciurel:1988:PLO

[CSMHT88] S. Ciurel, D. Stoica-Mann, R. Hritcu, and C. Tulpan. Physics laboratory oriented multiprocessor system for data acquisition and processing. Computer Physics Communications, 50 (1–2):237–240, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901300.

Crees:1984:IPS

[CSW84] M. A. Crees, M. J. Seaton, and P. M. H. Wilson. Impact, a program for the solution of the coupled integrodifferential equations of electron-atom collision theory. Computer Physics Communications, 35(1–3):C–481–C–482, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826764.

Clark:1982:AHU

[CT82] C. W. Clark and K. T. Taylor. Atomic hydrogen in a uniform magnetic field. Computer Physics Communications, 26 (3–4):415–417, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901369.

Chrisman:1984:CAM

[CT84a] B. L. Chrisman and T. A. Tumolillo. Computer analysis of Mössbauer spectra. Computer Physics Communications, 35(1–3):C–100, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823589.

Collocott:1984:ANS

[CT84b] S. J. Collocott and G. J. Troup. Adaptation: Numerical solution of the Kramers–Kronig transforms by trapezoidal summation as compared to a Fourier method. Computer Physics Communications, 35(1–3):C–567, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827423.

Cecchini:1989:SLA

[CT89] R. Cecchini and M. Tarlini. Symbolic Lie algebras manipulations using COMMON LISP. Computer Physics Communications, 52(2):283–289, January/February 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900106.

Cugnon:1984:CSB

[Cug84] J. Cugnon. Computation of S-state binding energy and wave functions in a saxon-wood potential. Computer Physics Communications, 35(1–3):C–204, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824467.

Copley:1986:IMC

[CVvWF86] J. R. D. Copley, P. Verkerk, A. A. van Well, and H. Fredrikze. Improved Monte Carlo calculation of multiple scattering effects in thermal neutron scattering experiments. Computer Physics Communications, 40(2–3):337–357, June 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901189.

Chang:1980:PEC

[CW80] B. D. Chang and S. S. M. Wong. A program to evaluate closed diagrams algebraically for angular momentum coupled product operators. Computer Physics Communications, 20(2):191–211, October 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900028.

Chang:1984:PEC

[CW84a] B. D. Chang and S. S. M. Wong. A program to evaluate closed diagrams algebraically for angular momentum coupled product operators. Computer Physics Communications, 35(1–3):C–629– C–630, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827927.

Chang:1984:PGC

[CW84b] B. D. Chang and S. S. M. Wong. A program to generate closed basic diagrams for product operators. Computer Physics Communications, 35(1–3):C–575, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827496.

Christiansen:1984:CTD

[CW84c] J. P. Christiansen and N. K. Winsor. Castor 2: a twodimensional laser target code. Computer Physics Communications, 35(1–3):C–568–C–569, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827435.

Caceres:1983:ECF

[CWS83] Manuel O. Cáceres, Horacio S. Wio, and R. J. J. Stamm’ler. An expansion of complicated functions using Chebyshev polynomials suitable for fast calculation. Computer Physics Communications, 29(3):261–267, May 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900061.

Caceres:1984:ECF

[CWS84] Manuel O. Cáceres, Horacio S. Wio, and R. J. J. Stamm’ler. An expansion of complicated functions using Chebyshev polynomials suitable for fast calculation. Computer Physics Communications, 35(1–3):C–880, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829835.

Dagless:1982:MA

[Dag82] E. L. Dagless. Microprocessor applications. Computer Physics Communications, 26(1–2):153–194, May 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901679.

Das:1983:GAS

[Das83] S. Das. A generalized algorithm for the solution of the inverse problem of nuclear reactor kinetics. Computer Physics Communications, 29(4):321–331, June 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900115.

Das:1989:CPB

[Das89] P. Das. Computer programs for the Boltzmann collision matrix elements. Computer Physics Communications, 55(2):177–187, September 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900751.

Davies:1981:ODC

[Dav81] Charles W. Davies. Orthogonalization of discrete coordinates. Computer Physics Communications, 23(4):427–433, August 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558190182X.

Davies:1984:ODC

[Dav84] Charles W. Davies. Orthogonalization of discrete coordinates. Computer Physics Communications, 35(1–3):C–740, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046558482874X.

Davidson:1989:SMM

[Dav89] Ernest R. Davidson. Super-matrix methods. Computer Physics Communications, 53(1–3):49–60, May 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901471.

Day:1981:RPN

[Day81] A. C. Day. RDLIST, a portable namelist facility. Computer Physics Communications, 22(4):403–410, May 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901375.

Day:1984:RPN

[Day84a] A. C. Day. Rdlist, a portable namelist facility. Computer Physics Communications, 35(1–3):C–690–C–691, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828416.

Day:1984:PTE

[Day84b] Colin Day. A portable text editor. Computer Physics Communications, 35(1–3):C–417, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826272.

deBrito:1982:FPI

[dB82] Adelsindo Liberato de Brito. FORTRAN program for the integral of three spherical harmonics. Computer Physics Communications, 25(1):81–85, January 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900467.

deBrito:1984:FPI

[dB84] Adelsindo Liberato de Brito. Fortran program for the integral of three spherical harmonics. Computer Physics Communications, 35(1–3):C–770, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828982.

Dahl:1984:CSE

[DBB84] R. E. Dahl, Jr., J. R. Beeler, Jr., and R. D. Bourquin. Computer simulation of extended defects in metals. Computer Physics Communications, 35(1–3):C–98–C–99, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823577.

deBiasi:1982:SPE

[dBM82] R. S. de Biasi and J. A. M. Mendonça. Simulation of powder EPR spectra with axial symmetry. Computer Physics Communications, 28(1):69–74, November 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900637.

deBiasi:1984:SPE

[dBM84] R. S. de Biasi and J. A. M. Mendonça. Simulation of powder EPR spectra with axial symmetry. Computer Physics Communications, 35(1–3):C–836, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829471.

Dedoussis:1984:NVM

[DCC84] Sp. Dedoussis, M. Chardalas, and Stef. Charalambous. A new version of minuit program for a PDP11/34A computer system. Computer Physics Communications, 31(1):29–39, January 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900791.

Daly:1985:PSL

[DD85a] C. Daly and J. J. Du Croz. Performance of a subroutine library on vector-processing machines. Computer Physics Communications, 37(1–3):181–186, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901511.

Dixon:1985:FEO

[DD85b] L. C. W. Dixon and P. G. Ducksbury. Finite element optimisation on the DAP. Computer Physics Communications, 37(1–3):187–193, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901523.

Dumont:1986:MCS

[DD86] Martine Dumont and P. Dufour. Monte Carlo simulation of surface reactions. Computer Physics Communications, 41(1): 1–19, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900172.

Devenish:1989:P

[DD89] R. C. E. Devenish and T. Daniels. Preface. Computer Physics Communications, 57(1–3):xiii, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901823.

DeHosson:1984:NCA

[De 84a] J. Th. M. De Hosson. Normal coordinate analysis of crystals. Computer Physics Communications, 35(1–3):C–331, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825552.

DeHosson:1984:SBS

[De 84b] J. Th. M. De Hosson. Symmetry and band structure. Computer Physics Communications, 35(1–3):C–328, ???? 1984.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825515.

DeJagher:1984:NPC

[De 84c] P. C. De Jagher. Nullijn, a program to calculate zero curves of a function of two variables of which one may be complex. Computer Physics Communications, 35(1–3):C–507, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826922.

DeAngelis:1988:CDP

[De 88] A. De Angelis. A class of N-dimensional probability density functions suitable for random generation. Computer Physics Communications, 52(1):61–64, December 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901725.

Davies:1984:CGG

[DE84] A. J. Davies and C. J. Evans. The computation of the growth of a gaseous discharge in space-charge distorted fields. Computer Physics Communications, 35(1–3):C–143, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823978.

Decker:1989:NIC

[Dec89] Karsten Decker. Numerical investigation of a coarse-to-fine transformation for multigrid Monte Carlo updating. Computer Physics Communications, 54(1):1–11, April 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900271.

DeGrand:1988:CTM

[DeG88] Thomas A. DeGrand. A conditioning technique for matrix inversion for Wilson fermions. Computer Physics Communications, 52(1):161–164, December 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901804.

Delic:1984:CEA

[Del84a] G. Delic. Chebyshev expansion of the associated Legendre polynomial P_{L}^{M}(x). Computer Physics Communications, 35(1–3): C–576, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827502.

Delic:1984:CSS

[Del84b] G. Delic. Chebyshev series for the spherical Bessel function l(r). Computer Physics Communications, 35(1–3):C–577, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827514.

Delves:1985:MCC

[Del85] L. M. Delves. Monte Carlo calculations of neutron diffusion on the ICL DAP. Computer Physics Communications, 37(1–3):295–301, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901651.

Delfino:1987:EMD

[Del87] Manuel C. Delfino. Experience with the MAC data flow system. Computer Physics Communications, 45(1–3):467–471, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901901.

Denby:1988:NNC

[Den88] B. Denby. Neural networks and cellular automata in experimental high energy physics. Computer Physics Communications, 49(3):429–448, June 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588900045.

Desclaux:1984:HFS

[Des84a] J. P. Desclaux. Hartree Fock Slater self consistent field calculations. Computer Physics Communications, 35(1–3):C–26, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822936.

Desclaux:1984:MRD

[Des84b] J. P. Desclaux. A multiconfiguration relativistic Dirac–Fock program. Computer Physics Communications, 35(1–3):C–288, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825187.

Deutsch:1984:SHSa

[Deu84a] Moshe Deutsch. Slit height smearing correction in small angle X-ray scattering I: Intensity correction program. Computer Physics Communications, 35(1–3):C–560, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827356.

Deutsch:1984:SHSb

[Deu84b] Moshe Deutsch. Slit height smearing correction in small angle X-ray scattering II: Computation of the correction function. Computer Physics Communications, 35(1–3):C–561, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827368.

Deutsch:1984:SHSc

[Deu84c] Moshe Deutsch. Slit height smearing correction in small angle X-ray scattering III: Intensity-correction program adaptation to arbitrary slit transmission function. Computer Physics Communications, 35(1–3):C–584–C–585, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827575.

Deutsch:1984:SHSd

[Deu84d] Moshe Deutsch. Slit height smearing correction in small angle X-ray scattering IV: Computation of the correction function for an arbitrary slit transmission function. Computer Physics Communications, 35(1–3):C–586, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827587.

DeVries:1984:QBS

[DeV84] R. M. DeVries. Quasi-bound state wavefunctions. Computer Physics Communications, 35(1–3):C–375, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825904.

Dewar:1982:BDD

[Dew82] P. A. Dewar. Basic digital device technologies. Computer Physics Communications, 26(1–2):1–33, May 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901552.

Davies:1984:SGA

[DEW84] A. J. Davies, C. J. Evans, and P. M. Woodison. Simulation of the growth of axially symmetric discharges between plane parallel electrodes. Computer Physics Communications, 35(1–3):C–469–C–470, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826697.

Docken:1984:DOI

[DF84] Kate K. Docken and A. Lewis Ford. Dipole and overlap integrals between Slater-type functions and continuum Coulomb functions. Computer Physics Communications, 35(1–3):C–360, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825795.

DeRaedt:1981:MCF

[DFD81] H. De Raedt, J. Fivez, and B. De Raedt. Manual for commute, a FORTRAN program for symbolic evaluation of commutators and correlation functions. Computer Physics Communications, 23(2):209–220, July 1, 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900345.

DeRaedt:1984:MCF

[DFD84] H. De Raedt, J. Fivez, and B. De Raedt. Manual for commute, a FORTRAN program for symbolic evaluation of commutators and correlation functions. Computer Physics Communications, 35(1–3):C–721, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828611.

Delfino:1989:AER

[DFOP89] M. Delfino, E. Fernández, S. Orteu, and A. Pacheco. The ALEPH event reconstruction facility: Parallel processing using workstations. Computer Physics Communications, 57(1– 3):401–406, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589902518.

Darville:1984:ECZ

[DG84] J. Darville and A. Gérard. Exact computation of the Zeeman effect on nuclear quadrupole resonance profiles for powders (spin I = 3/2). Determination of the asymmetry parameter. Computer Physics Communications, 35(1–3):C–300, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825291.

Damarowsky:1989:SAK

[DG89] M. Damarowsky and G. H. Guthöhrlein. Spectral analysis without knowledge of the line shape function by use of a new numerical procedure. Computer Physics Communications, 52(2):187–194, January/February 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900039.

Dyall:1989:GGP

[DGJ^{+}89] K. G. Dyall, I. P. Grant, C. T. Johnson, F. A. Parpia, and E. P. Plummer. GRASP: a general-purpose relativistic atomic structure program. Computer Physics Communications, 55(3):425– 456, October 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901367.

Damerell:1981:UEF

[DGK^{+}81] C. Damerell, D. Giddings, B. Kisielewski, G. Lutz, G. Lutjens, J.-P Melot, K. Rybicki, F. Wickens, and ACCMOR Collaboration. Use of the ESOP fast processor in the trigger logic of the CERN experiment NA-11. Computer Physics Communications, 22(2–3):349–352, April 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900709.

Dean:1989:SSP

[DGL89] E. Dean, R. Glowinski, and C. H. Li. Supercomputer solutions of partial differential equation problems in computational fluid dynamics and in control. Computer Physics Communications, 53(1–3):401–439, May 1, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901768.

Diercksen:1983:CIP

[DGO83] G. H. F. Diercksen, N. E. Grüner, and Jens Oddershede. Computational implementation of the polarization propagator method. Computer Physics Communications, 30(4):349–358, December 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900760.

DeBlasi:1985:CMS

[DGPM85] C. De Blasi, L. Guercia, C. Paracelli, and C. Manfredotti. A class of methods for smoothing and numerical derivation of functions. Computer Physics Communications, 36(4):339– 343, June 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585900220.

Dragoun:1984:PCI

[DH84a] O. Dragoun and G. Heuser. A program to calculate internal conversion coefficients for all atomic shells without screening. Computer Physics Communications, 35(1–3):C–108, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823668.

Dudder:1984:RTD

[DH84b] H. D. Düdder and D. B. Henderson. Ramses, a twodimensional PIC type, laser pulse propagation code. Computer Physics Communications, 35(1–3):C–334–C–335, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825588.

Delbrouck-Habaru:1984:BSO

[DHD84] J. M. Delbrouck-Habaru and Daniel M. Dubois. Bound states of one nucleon in a Woods–Saxon well from a variational method. Computer Physics Communications, 35(1–3):C–286, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825163.

Der:1984:PCI

[DHN84] R. Der, D. Hinneburg, and M. Nagel. A program to calculate internal conversion coefficients including higher-order corrections for all atomic shells. Computer Physics Communications, 35(1–3):C–603, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827721.

DiMarzio:1986:IPA

[Di 86] F. Di Marzio. An improved procedure for the accurate evaluation of polygamma functions with integer and half-integer argument. Computer Physics Communications, 39(3):343–345, April 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900950.

DiMarzio:1987:VAS

[Di 87] F. Di Marzio. The very accurate summation of inverse powers and the generation of Bernoulli and Euler numbers. Computer Physics Communications, 44(1–2):57–62, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900166.

Dickinson:1982:ONW

[Dic82] A. Dickinson. Optimizing numerical weather forecasting models for the Cray-1 and Cyber 205 computers. Computer Physics Communications, 26(3–4):459–468, June 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901436.

Dietz:1988:IIG

[Die88] F. Dietz. INPLO: An interactive graphics tool. Computer Physics Communications, 50(1–2):241–245, July 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901312.

Dittus:1989:PPA

[Dit89] Fridolin Dittus. Parallel processing with attached processors in a computer center environment. Computer Physics Communications, 57(1–3):395–400, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902506.

Diercksen:1982:ESD

[DK82] G. H. F. Diercksen and W. P. Kraemer. External standard data structures for computation in chemistry. Computer Physics Communications, 25(1):1–6, January 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900388.

Dagher:1987:DPC

[DK87a] Mounzer Dagher and Hafez Kobeissi. DIRIGE — a program for calculating eigenvalues and initial values of log derivative eigenfunctions for a diatomic molecule. Computer Physics Communications, 46(3):445–451, September 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900981.

Dagher:1987:FPC

[DK87b] Mounzer Dagher and Hafez Kobeissi. FFC — a program for calculating Franck–Condon factors and R-centroids for transitions between the vibrational-rotational levels of two electronic states of a diatomic molecule. Computer Physics Communications, 47(2–3):305–309, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901160.

Diercksen:1987:IPM

[DK87c] Geerd H. F. Diercksen and Jacek Karwowski. Invariance properties of the moments of the Hamiltonian matrix as a test of the correctness of configuration interaction programs. Computer Physics Communications, 47(1):83–89, October 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900683.

Duek:1985:CLM

[DKA85] E. Duek, L. Kowalski, and John M. Alexander. The code LINDA: a Monte Carlo reaction simulation for correlated fragments and evaporation residues formed in nuclear reactions. Computer Physics Communications, 34(4):395–413, February 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900682.

Delannoy:1986:MCP

[DL86] Claude Delannoy and François Lefeuvre. Maxentwdf — a computer program for the maximum entropy estimation of a wave distribution function. Computer Physics Communications, 40(2–3):389–419, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901219.

Dasso:1987:CSC

[DL87] C. H. Dasso and S. Landowne. CCFUS — a simplified coupledchannel code for calculation of fusion cross sections in heavyion reactions. Computer Physics Communications, 46(1):187– 191, July 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900452.

Denby:1989:SHN

[DL89] Bruce Denby and Stephen L. Linn. Status of HEP neural net research in the USA. Computer Physics Communications, 57(1–3):297–300, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902312.

DeSanctis:1983:PMC

[DLB83] E. De Sanctis, V. Lucherini, and V. Bellini. Phocha: a Monte Carlo program to calculate the characteristics of a beam of photons produced by annihilation and Bremsstrahlung of relativistic positrons. Computer Physics Communications, 30(1):71–85, July/August 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583901248.

DeSanctis:1984:PMC

[DLB84] E. De Sanctis, V. Lucherini, and V. Bellini. Phocha: a Monte Carlo program to calculate the characteristics of a beam of photons produced by annihilation and Bremsstrahlung of relativistic positrons. Computer Physics Communications, 35(1–3):C– 901, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584830003.

Draayer:1989:R

[DLPL89] J. P. Draayer, Y. Leschber, S. C. Park, and R. Lopez. Representations of U(3) in U(N). Computer Physics Communications, 56(2):279–290, December 1, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900246.

Davisson:1986:APD

[DM86a] C. M. Davisson and Irwin Manning. Adaptation of a program for depth distribution of energy deposition by ion bombardment: Better stopping powers. Computer Physics Communications, 42(1):137–147, September 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586902389.

Degtyarev:1986:MNS

[DM86b] L. M. Degtyarev and S. Yu. Medvedev. Methods for numerical simulation of ideal MHD stability of axisymmetric plasmas. Computer Physics Communications, 43(1):29–56, December 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900524.

DiGenova:1987:CAI

[DM87] Glauco Di Genova and Marco Matone. A compact algorithm for the implementation of Monte Carlo kinematics in computer simulations of d-dimensional lattice gauge theories. Computer Physics Communications, 47(2–3):235–244, November/ December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558790110X.

Delic:1988:SCS

[DM88a] G. Delic and S. M. Malherbe. Subroutines for convolution sums of Chebyshev and Fourier series. Computer Physics Communications, 48(2):305–312, February 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900495.

Delic:1988:SIS

[DM88b] G. Delic and S. M. Malherbe. Subroutines for integration of systems of first order ODE’s. Computer Physics Communications, 48(2):293–304, February 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588900483.

Davidson:1989:BEP

[DM89a] D. B. Davidson and D. A. McNamara. A boundary element program package for electromagnetic excitation of conducting bodies of revolution by an asymmetrical slot. Computer Physics Communications, 56(2):249–257, December 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900222.

Drouffe:1989:FCT

[DM89b] J.-M. Drouffe and K. J. M. Moriarty. FORTRAN code for the three-dimensional Ising model on bcc and fcc lattices. Computer Physics Communications, 52(2):249–259, January/ February 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900088.

Drouffe:1983:MCS

[DMM83] J.-M. Drouffe, K. J. M. Moriarty, and C. N. Mouhas. Monte Carlo simulation of pure U(N) and SU(N) gauge theories on a simplicial lattice. Computer Physics Communications, 30(3):249–254, November 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900929.

Ddgar:1984:MCS

[DMM84a] R. C. Ddgar, L. McCrossen, and K. J. M. Moriarty. Monte Carlo simulation of U(1) lattice gauge theory. Computer Physics Communications, 35(1–3):C–695, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828441.

Drouffe:1984:MCS

[DMM84b] J.-M. Drouffe, K. J. M. Moriarty, and C. N. Mouhas. Monte Carlo simulation of pure U(N) and SU(N) gauge theories on a simplicial lattice. Computer Physics Communications, 35(1–3): C–913, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046558483012X.

DAguanno:1983:CPM

[DNR83] B. D’Aguanno, A. Nobile, and E. Roman. CHPACK: a package for the manipulation of Chebyshev approximations. Computer Physics Communications, 29(4):361–374, June 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900152.

DAguanno:1984:CPM

[DNR84] B. D’Aguanno, A. Nobile, and E. Roman. Chpack: a package for the manipulation of Chebyshev approximations. Computer Physics Communications, 35(1–3):C–888, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829914.

Durgapal:1984:PCV

[DO84] P. Durgapal and D. S. Onley. A program to calculate virtual photon spectrum in second order born approximation. Computer Physics Communications, 32(3):291–307, June 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900924.

Dobes:1984:CGH

[Dob84] J. Dobes. Calculations of generalized harmonic oscillator brackets. Computer Physics Communications, 35(1–3):C–548, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827241.

Dobes:1985:FFM

[Dob85] K. Dobes. Formal — a Fortran memory allocation system. Computer Physics Communications, 36(2):147–159, April 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901201.

Dobberstein:1989:CZF

[Dob89] M. P. Dobberstein. The calibration of the Zeus forward drift chamber: An application for software with a high degree of parallelism. Computer Physics Communications, 57(1–3):483–485, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589902683.

Dommaschk:1986:RVP

[Dom86] W. Dommaschk. Representations for vacuum potentials in stellarators. Computer Physics Communications, 40(2–3):203– 218, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586901098.

Dorfi:1986:NMA

[Dor86] E. A. Dorfi. Numerical methods for astrophysical plasmas. Computer Physics Communications, 43(1):1–15, December 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900500.

Dowsing:1982:PLM

[Dow82a] R. D. Dowsing. Programming languages for microcomputers. Computer Physics Communications, 26(1–2):41–48, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901576.

Dowsing:1982:SDM

[Dow82b] R. D. Dowsing. Software design methodologies for microcomputers. Computer Physics Communications, 26(1–2):49– 56, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901588.

Dowsing:1982:SDC

[Dow82c] R. D. Dowsing. Software for distributed computing. Computer Physics Communications, 26(1–2):35–40, May 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901564.

Duhamel:1984:EDP

[DP84] Christian Duhamel and Gérard Parlant. Estimation of densities of probability and regression surfaces in one or two dimensions. Computer Physics Communications, 32(1):21–27, April 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900043.

Dasso:1988:TSC

[DP88] C. H. Dasso and G. Pollarolo. TORINO — a semiclassical coupled channel code for heavy ion reactions. Computer Physics Communications, 50(3):341–366, August 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901907.

DaRocha:1989:SQA

[DP89] Ana Regina C. Da Rocha and Simoni Palermo. Software quality assurance in HEP. Computer Physics Communications, 57(1–3):524–527, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902798.

Durham:1982:CXR

[DPH82] P. J. Durham, J. B. Pendry, and C. H. Hodges. Calculation of X-ray absorption near-edge structure, XANES. Computer Physics Communications, 25(2):193–205, February 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900352.

Durham:1984:CXR

[DPH84] P. J. Durham, J. B. Pendry, and C. H. Hodges. Calculation of X-ray absorption near-edge structure, xanes. Computer Physics Communications, 35(1–3):C–781–C–782, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829082.

Duff:1982:ESM

[DR82] I. S. Duff and J. K. Reid. Experience of sparse matrix codes on the Cray-1. Computer Physics Communications, 26(3–4):293– 302, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901199.

Demichev:1985:RPC

[DR85] A. P. Demichev and A. Ya. Rodionov. A REDUCE program for the calculation of geometrical characteristics of compactified multidimensional Riemannian space. Computer Physics Communications, 38(3):441–448, December 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901122.

Docherty:1988:MCP

[DRD88] R. Docherty, K. J. Roberts, and E. Dowty. Morang — a computer program designed to aid in the determinations of crystal morphology. Computer Physics Communications, 51(3):423–430, November 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901555.

Dullemond:1982:NMS

[DRRvB82] C. Dullemond, G. Rupp, T. A. Rijken, and E. van Beveren. A numerical method for solving a coupled channel Schrödinger equation. Computer Physics Communications, 27(4):377–384, October 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900996.

Drummond:1983:CPD

[Dru83] P. D. Drummond. Central partial difference propagation algorithms. Computer Physics Communications, 29(3):211–225, May 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900012.

Delves:1982:DSD

[DS82] L. M. Delves and A. Samba. Deconvolution of seismic data. Computer Physics Communications, 26(3–4):473–476, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558290145X.

Deutz:1983:PCE

[DS83] J. W. Deutz and H. R. Schober. Program to calculate elastic Green’s functions, displacement fields and interaction energies in cubic materials. Computer Physics Communications, 30(1):87–91, July/August 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558390125X.

Deutz:1984:PCE

[DS84a] J. W. Deutz and H. R. Schober. Program to calculate elastic Green’s functions, displacement fields and interaction energies in cubic materials. Computer Physics Communications, 35(1–3):C–902, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584830015.

Dufour:1984:GPM

[DS84b] P. Dufour and J. Schlesinger. A general purpose Monte–Carlo program. Computer Physics Communications, 35(1–3):C–317, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825424.

Defranceschi:1989:NIS

[DS89] M. Defranceschi and M. Sarrazin. Numerical integration over a spherical shell. Computer Physics Communications, 52(3):409– 414, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558990115X.

Demsky:1983:MMC

[DSK83] J. Demsky, M. Schlesinger, and R. D. Kent. Micro/mini computer program for calculating the square root of rationals at arbitrary precision. Computer Physics Communications, 29(3):237–244, May 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900048.

Davies:1984:DDI

[DSK84a] Alan R. Davies, Kenneth Smith, and K. L. Kwok. DIRAC dynamic information retrieval of atomic codes: II. Implementation. Computer Physics Communications, 35(1–3):C–216–C– 217, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824583.

Demsky:1984:MMC

[DSK84b] J. Demsky, M. Schlesinger, and R. D. Kent. Micro/mini computer program for calculating the square root of rationals at arbitrary precision. Computer Physics Communications, 35(1–3): C–877, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829811.

Daul:1981:SES

[DSM^{+}81] C. Daul, C. W. Schläpfer, B. Mohos, J. Ammeter, and E. Gamp. Simulation of EPR-spectra of randomly oriented samples. Computer Physics Communications, 21(3):385–395, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900151.

Daul:1984:SES

[DSM^{+}84] C. Daul, C. W. Schläpfer, B. Mohos, J. Ammeter, and E. Gamp. Simulation of EPR-spectra of randomly oriented samples. Computer Physics Communications, 35(1–3):C–674, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828271.

Dios:1986:DIC

[DSSS86] Z. Diós, K. Sepsy, I. Szabó, and G. Székely. Design and implementation of a control system to a quadrupole mass spectrometer. Computer Physics Communications, 41(2–3):423– 425, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900846.

Davies:1980:ICL

[DST80] A. R. Davies, Kenneth Smith, and R. M. Thomson. IN-JLOK — a CO_{2} laser injection locking code. Computer Physics Communications, 20(3):413–420, November 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900181.

Davies:1984:ICL

[DST84a] A. R. Davies, Kenneth Smith, and R. M. Thomson. Injlok — a CO_{2} laser injection locking code. Computer Physics Communications, 35(1–3):C–644, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828040.

Davies:1984:TCL

[DST84b] A. R. Davies, Kenneth Smith, and R. M. Thomson. Tlaser — a CO_{2} laser kinetics code. Computer Physics Communications, 35(1–3):C–332, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825564.

Datta:1984:VNO

[DTE84] S. K. Datta, W. J. Thompson, and D. O. Elliott. A version of a nuclear optical model code for small computers designed to run on a PDP-15. Computer Physics Communications, 35(1–3): C–250, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824868.

Dubey:1989:TIM

[DU89] I. P. Dubey and S. K. Upadhyay. A trivariate interpolation method developed on the basis of Akima’s bivariate interpolation procedure. Computer Physics Communications, 54(1): 23–29, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900295.

Duch:1980:LSF

[Duc80] Wlodzislaw Duch. Large scale N-fermion calculations. Computer Physics Communications, 20(1):49–52, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558090106X.

Duch:1986:BRB

[Duc86] Wlodzislaw Duch. Book review: Microcomputer Quantum Mechanics: Second edition, J. P. Killingbeck (University of Hull), Adam Hilger Ltd., Bristol and Boston, 1985. xii + 188 pp. £12.95/$20.00. Computer Physics Communications, 39(2):301–302, February/March 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901402.

Dufton:1984:PCC

[Duf84] P. L. Dufton. A program to calculate coronal emission line strengths. Computer Physics Communications, 35(1–3):C–410, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046558482620X.

Duff:1985:USE

[Duf85] Iain S. Duff. The use of supercomputers in Europe. Computer Physics Communications, 37(1–3):15–25, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901328.

Dunwoodie:1987:UCG

[Dun87] W. Dunwoodie. The use of color and gray-scale scatterplots in the analysis of high-energy physics data. Computer Physics Communications, 45(1–3):485, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901974.

Draayer:1985:MSD

[DV85] J. P. Draayer and H. T. Valdes. Model space dimensionalities for multiparticle fermion systems. Computer Physics Communications, 36(3):313–320, May 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558590058X.

DeMeyer:1987:TPP

[DVD87] H. De Meyer, G. Vanden Berghe, and P. De Wilde. Two programs to perform certain symbolic calculations in the enveloping algebra of a Lie algebra. Computer Physics Communications, 44(1–2):197–208, April/May 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900270.

Dodds:1984:VCC

[DW84] R. McD. Dodds and G. Wiechers. Vector coupling coefficients as products of prime factors. Computer Physics Communications, 35(1–3):C–160, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482411X.

Davie:1988:TSB

[DW88] K. Davie and R. Wallace. Title — solution of bound state single variable eigenproblems by the extended renormalized Numerov method. Computer Physics Communications, 51(1–2):217–223, September/October 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900732.

Davie:1989:OGJ

[DW89] K. Davie and R. Wallace. Orthogonal generalized Jacobi coordinates for N-body systems. Computer Physics Communications, 55(3):463–468, October 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901380.

Duneczky:1989:IMS

[DWC^{+}89] Csilla Duneczky, Robert E. Wyatt, David Chatfield, Kenneth Haug, David W. Schwenke, Donald G. Truhlar, Yan Sun, and Donald J. Kouri. Iterative methods for solving the non-sparse equations of quantum mechanical reactive scattering. Computer Physics Communications, 53(1–3):357–379, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901732.

Dupuis:1989:GAM

[DWVH89] M. Dupuis, J. D. Watts, H. O. Villar, and G. J. B. Hurst. The general atomic and molecular electronic structure system HONDO: Version 7.0. Computer Physics Communications, 52(3):415–425, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901161.

Dyall:1986:TPC

[Dya86] K. G. Dyall. Transform: a program to calculate transformations between various jj and LS coupling schemes. Computer Physics Communications, 39(1):141–152, January 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901694.

Ekholm:1983:CSN

[EA83] K. T. Ekholm and A. M. Aurela. Counting a small number of radioactive atoms Monte Carlo program. Computer Physics Communications, 29(2):163–170, April 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900711.

Ekholm:1984:CSN

[EA84] K. T. Ekholm and A. M. Aurela. Counting a small number of radioactive atoms Monte Carlo program. Computer Physics Communications, 35(1–3):C–872, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829768.

Elman:1989:OTP

[EA89] Howard C. Elman and Elvira Agrón. Ordering techniques for the preconditioned conjugate gradient method on parallel computers. Computer Physics Communications, 53(1–3):253– 269, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901641.

Eastwood:1980:BRB

[Eas80] J. W. Eastwood. Book review: An introduction to computational fluid mechanics: C.-Y. Chow, Wiley, 1979. 396 pages. $11.50. Computer Physics Communications, 20(2):321–322, October 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900107.

Eastwood:1982:BRB

[Eas82] J. W. Eastwood. Book review: Parallel computers: Architecture, programming and algorithms: R. W. Hockney and C. R. Jesshope, Adam-Hilger, Bristol, 1981. xii + 416 pages. £22.50. Computer Physics Communications, 27(1):104, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900169.

Eastwood:1984:BRBb

[Eas84a] J. W. Eastwood. Book review: Computational methods for fluid flow: R. Peyret and T. D. Taylor 1982, 125 figs., 358 pages, DM 92; approx. US $36.80 Springer-Verlag, Berlin-Heidelberg- New York ISBN 3-540-11147-6. Computer Physics Communications, 32(4):439–440, July 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900602.

Eastwood:1984:BRBa

[Eas84b] J. W. Eastwood. Book review: Springer series in computational physics: Editors: H. Cabannes, M. Holt, H. B. Keller, J. Killeen and S. A. Orszag. Computer Physics Communications, 32(4): 439, July 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900596.

Eastwood:1986:PSM

[Eas86] James W. Eastwood. Particle simulation methods in plasma physics. Computer Physics Communications, 43(1):89–106, December 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558690055X.

Eastwood:1987:ORP

[Eas87a] James W. Eastwood. ORTHOVEC: A REDUCE program for 3-D vector analysis in orthogonal curvilinear coordinates. Computer Physics Communications, 47(1):139–147, October 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900737.

Eastwood:1987:SAE

[Eas87b] James W. Eastwood. The stability and accuracy of EPIC algorithms. Computer Physics Communications, 44(1–2):73–82, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558790018X.

Eastwood:1988:LDF

[Eas88] James W. Eastwood. Low dissipation fluid and MHD simulation. Computer Physics Communications, 48(1):17–23, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900197.

Eastwood:1989:BRB

[Eas89] J. W. Eastwood. Book review: Computer simulation and computer algebra: D. Stauffer, F. W. Hehl, V. Winkelmann and J. G. Zabolitzky, Springer-Verlag, Berlin, 1988, 155 Pages. DM 34 (soft cover). ISBN 3-540-18909-2. Computer Physics Communications, 54(1):199, April 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900453.

Ebel:1984:RNI

[Ebe84] G. Ebel. The role of a national information center. Computer Physics Communications, 33(1–3):211–215, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901243.

Ernenwein:1988:VSC

[EBS88] René Ernenwein, Marc Benard, and Isaiah Shavitt. Vectorizing a sequence of conditional branches: The calculation of the class index of two-electron repulsion integrals on Cray computers. Computer Physics Communications, 48(2):175–180, February 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900380.

Egecioglu:1984:MRI

[EC84] Ömer Eğecioğlu and Geraldo M. Costa. Murnaghan’s rule and the irreducible characters of the symmetric group. Computer Physics Communications, 31(4):357–362, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900195.

Eckermann:1984:FTP

[Eck84] R. Eckermann. A flexible thermophysical property information system for chemical engineering and chemical process design. Computer Physics Communications, 33(1–3):245–256, August/ September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901309.

Edery:1984:CSS

[Ede84] D. Edery. Calculations of stationary solutions for the nonlinear viscous resistive MHD equations in slab geometry. Computer Physics Communications, 31(2–3):149–153, February 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900420.

Etchegoyen:1989:MIF

[EE89] A. Etchegoyen and M. C. Etchegoyen. Microscopic inelastic form factors for heavy-ion reactions. Computer Physics Communications, 55(2):217–226, September 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900787.

Etchegoyen:1989:EHT

[EEV89] A. Etchegoyen, M. C. Etchegoyen, and E. G. Vergini. Evaluation of Hamiltonian two-body matrix elements. Computer Physics Communications, 55(2):227–231, September 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900799.

El-Fayoumi:1981:PLW

[EFES81] M. K. K. El-Fayoumi and W. A. El-Sebaii. On the phoenical Lax–Wendroff method. Computer Physics Communications, 23(1):27–30, June 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581901260.

El-Fayoumi:1983:DDE

[EFES83] M. K. K. El-Fayoumi and W. A. El-Sebaii. Diffusion and dispersion errors of the flux corrected upwind method. Computer Physics Communications, 28(3):317–321, January 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900498.

Eliseev:1985:RPD

[EFK85] V. P. Eliseev, R. N. Fedorova, and V. V. Kornyak. A RE-DUCE program for determining point and contact Lie symmetries of differential equations. Computer Physics Communications, 36(4):383–389, June 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558590027X.

Egecioglu:1982:COP

[Eğe82] Ömer Eğecioğlu. Computation of outer products of Schur functions. Computer Physics Communications, 28(2):183–187, December 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900534.

Egecioglu:1984:COP

[Ege84] Ömer Egecioglu. Computation of outer products of Schur functions. Computer Physics Communications, 35(1–3):C–843, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829525.

Ermler:1988:PSF

[EHH88] Walter C. Ermler, Hsiuchin C. Hsieh, and Lawrence B. Harding. Polyatomic surface fitting, vibrational-rotational analysis, expectation value and intensity program. Computer Physics Communications, 51(1–2):257–284, September/ October 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588900768.

Eastwood:1980:PTD

[EHL80] J. W. Eastwood, R. W. Hockney, and D. N. Lawrence. P3M3DP-the three-dimensional periodic particle-particle/ particle-mesh program. Computer Physics Communications, 19(2):215–261, April/June 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900521.

Eastwood:1984:PTD

[EHL84] J. W. Eastwood, R. W. Hockney, and D. N. Lawrence. P3M3DP-the three-dimensional periodic particle-particle/particle-mesh program. Computer Physics Communications, 35(1– 3):C–618–C–619, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827836.

Eberhard:1984:UMO

[EK84] Philippe H. Eberhard and Werner O. Koellner. Users manual for the optime system. Computer Physics Com-

munications, 35(1–3):C–173–C–174, ????  1984.  CO-

DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824224.

Erwin:1986:SDE

[EK86] D. A. Erwin and J. A. Kunc. Scalar DC electrical conductivity of partially ionized gases. Computer Physics Communications, 42(1):119–125, September 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586902365.

Ekstrom:1984:DCP

[Eks84] L. P. Ekström. Delta — a computer program to analyze gamma–gamma angular correlations from unaligned states. Computer Physics Communications, 32(4):399–411, July 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900560.

Elder:1984:DBM

[Eld84] Michael Elder. Data base management systems for experimental and calculation results. Computer Physics Communications, 33(1–3):41–47, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558490105X.

Egan:1988:CRT

[ELS88] Michael P. Egan, Chun Ming Leung, and George F. Spagna, Jr. CSDUST3: a radiation transport code for a dusty medium with 1-d planar, spherical or cylindrical geometry. Computer Physics Communications, 48(2):271–292, February 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900471.

Eisenstein:1984:DCP

[EM84a] R. A. Eisenstein and G. A. Miller. DWPI: a computer program to calculate the inelastic scattering of pions from nuclei. Computer Physics Communications, 35(1–3):C–365–C–366, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825825.

Eisenstein:1984:PCP

[EM84b] R. A. Eisenstein and G. A. Miller. Pirk: a computer program to calculate the elastic scattering of pions from nuclei. Computer Physics Communications, 35(1–3):C–269, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825035.

Edgar:1981:MCS

[EMM81] R. C. Edgar, L. McCrossen, and K. J. M. Moriarty. Monte Carlo simulation of U(1) lattice gauge theory. Computer Physics Communications, 22(4):433–437, May 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901405.

Errea:1984:CAF

[EMR84] L. F. Errea, L. Méndez, and A. Riera. Calculation of the auxiliary functions F_{m}(z). Computer Physics Communications, 31(1):47–52, January 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558490081X.

Engelmann:1986:IPR

[ENSC86] Alejandro R. Engelmann, Mario A. Natiello, Gustavo E. Scuseria, and Rubén H. Contreras. IPPP — a program for the RPA calculation of transmission mechanisms of spin-spin coupling constants. Computer Physics Communications, 39(3):409–420, April 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900974.

Edery:1986:FEA

[EP86] D. Edery and H. Picq. A finite element approach for ICRF heating calculations in real tokamak plasmas. Computer Physics Communications, 40(1):95–98, May 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901505.

Edery:1981:ETC

[EPS81] D. Edery, R. Pellat, and J. L. Soulé. Effects of toroidal coupling on the non-linear evolution of tearing modes and on the stochastisation of the magnetic field topology in plasmas. Computer Physics Communications, 24(3–4):427–436, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901673.

Epperlein:1988:CSV

[ERB88] E. M. Epperlein, G. J. Rickard, and A. R. Bell. A code for the solution of the Vlasov–Fokker–Planck equation in 1-D or 2-D. Computer Physics Communications, 52(1):7–13, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901658.

Erisman:1985:CK

[Eri85] A. M. Erisman. Computational kernels. Computer Physics Communications, 37(1–3):149–157, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.com/science/article/pii/001046558590147X.

Evangelakis:1987:MPS

[ERLD87] G. A. Evangelakis, J. P. Rizos, I. E. Lagaris, and I. N. Demetropoulos. Merlin — a portable system for multidimensional minimization. Computer Physics Communications, 46(3):401–415, September 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900944.

Eisenstein:1984:PMS

[ET84] R. A. Eisenstein and Frank Tabakin. PIPIT: a momentum space optical potential code for pions. Computer Physics Communications, 35(1–3):C–402, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826120.

Eastgate:1984:NOM

[ETH84] R. J. Eastgate, W. J. Thompson, and R. A. Hardekopf. A nuclear optical model code for small computers. Computer Physics Communications, 35(1–3):C–168–C–169, ????

1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824194.

Ezhela:1984:PPD

[Ezh84] V. V. Ezhela. Particle physics data systematization at IHEP. Computer Physics Communications, 33(1–3):225–228, August/ September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901279.

Fujiki:1982:CCA

[FAK82] Sumiyoshi Fujiki, Yoshihiko Abe, and Shigetoshi Katsura. Cube cluster approximation for the spin glass in the simple cubic lattice. Computer Physics Communications, 25(2):119–123, February 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900285.

Feng:1984:PCC

[FB84] Da Hsuan Feng and A. R. Barnett. Patiwen — a code for Coulomb-nuclear interference calculations. Computer Physics Communications, 35(1–3):C–351–C–352, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825722.

Foresti:1988:MSM

[FBS88] S. Foresti, G. Brussino, and V. Sonnad. Multilevel solution method for the p-version of finite elements: Parallel implementation and comparison with other solution methods. Technical Report KGN-137, IBM Corporation, Center for Scientific & Engineering Computations, Dept. 48B/428, Neighborhood Road, Kingston NY 12401, 1988. This technical report is an extended version of a published paper [FBS89]. The IBM Kingston Center for Scientific & Engineering Computations was closed. Please contact the author at his present address.

Foresti:1989:MSM

[FBS89] S. Foresti, G. Brussino, and V. Sonnad. Multilevel solution method for the p-version of finite elements. Computer Physics Communications, 53(1–3):349–355, May 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901720.  The technical report

[FBS88] is an extended version of this paper.

Festeau-Barrioz:1982:LAS

[FBW82] M. C. Festeau-Barrioz and E. S. Weibel. Large amplitude solutions of the nonlinear wave equation for an ideal, cold three-fluid plasma. Computer Physics Communications, 27(1):11–23, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900030.

Fleurier:1984:IAI

[FC84a] C. Fleurier and J. Chapelle. Inversion of ABEL’s integral equation — application to plasma spectroscopy. Computer Physics Communications, 35(1–3):C–237, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824765.

Foster:1984:PAE

[FC84b] B. P. Foster and D. C. Camp. Programs to aid in establishing gamma-ray decay schemes. DCSCH3 and DC-SCH4. Computer Physics Communications, 35(1–3):C–94, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482353X.

Fang:1984:CSS

[FCW84] M. T. C. Fang, S. K. Chan, and R. D. Wright. The computation of steady state arcs in nozzle flow. Computer Physics Communications, 35(1–3):C–437, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826454.

Fried:1988:PPC

[FE88] Laurence E. Fried and Gregory S. Ezra. PERTURB: a program for calculating vibrational energies by generalized algebraic quantization. Computer Physics Communications, 51(1–2):103–114, September/October 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900653.

Femenia:1984:CPC

[Fem84] F. R. Femenia. Codnum, a program to change the punching code and to number a deck of cards. Computer Physics Communications, 35(1–3):C–77, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823401.

Ferguson:1984:PCXb

[Fer84] I. F. Ferguson. A program for the calculation of X-ray reflection intensities, part 2. Computer Physics Communications, 35(1–3):C–326, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825497.

Fernandez:1989:MCC

[Fer89] J. E. Fernandez. Monte Carlo computer simulation of the XRF intensity dependence on the propagation plane inclination. Computer Physics Communications, 54(2–3):211–220, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900830.

Ferguson:1984:PCP

[FFH84] I. F. Ferguson, R. S. Fox, and T. E. Hughes. A program for the calculation of the positions of X-ray powder reflections. Computer Physics Communications, 35(1–3):C–406, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826156.

Falck:1989:CSF

[FGK89] N. K. Falck, D. Graudenz, and G. Kramer. Cross section for five-parton production in e^{+}e^{−} annihilation. Computer Physics Communications, 56(2):181–198, December 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900192.

Falkensteiner:1985:SSE

[FGSH85] P. Falkensteiner, H. Grosse, F. Schöberl, and P. Hertel. Solving the Schrödinger equation for bound states. Computer Physics Communications, 34(3):287–293, January 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900050.

Fleming:1988:RHM

[FH88a] Patrick R. Fleming and John S. Hutchinson. Representation of the Hamiltonian matrix in non-local coordinates for an acetylene bond-mode model. Computer Physics Communications, 51(1–2):59–71, September/October 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900628.

Frederick:1988:QVE

[FH88b] John H. Frederick and Eric J. Heller. Quantum vibrational eigenstates from classical origins. Computer Physics Communications, 51(1–2):83–102, September/October 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900641.

Ferguson:1984:PCI

[FHMR84] I. F. Ferguson, A. D. Hardy, M. U. Modi, and A. H. Rogerson. A program for the calculation of the intensities of X-ray or neutron powder reflections, part 3. Computer Physics Communications, 32(1):83–94, April 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900109.

Fiala:1984:SDB

[Fia84] Jaroslav Fiala. Spectral data bases for chemical compound identification. Computer Physics Communications, 33(1–3): 85–92, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584901103.

Fiebig:1980:SFA

[Fie80] H. R. Fiebig. A spline function approach to the numerical treatment of scattering and bound-state problems for nonlocal potentials. Computer Physics Communications, 20(2):181–189, October 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900016.

Fiebig:1981:SFP

[Fie81] H. R. Fiebig. A spline function program for treating nonlocal potentials. Computer Physics Communications, 23(2):135– 143, July 1, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558190028X.

Fiebig:1984:SFP

[Fie84] H. R. Fiebig. A spline function program for treating nonlocal potentials. Computer Physics Communications, 35(1–3): C–711, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828556.

Fincham:1980:PMD

[Fin80] David Fincham. Programs for the molecular dynamics simulation of liquids: I. Spherical molecules with short-ranged interactions. Computer Physics Communications, 21(2):247–256, December 2, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900430.

Fincham:1982:PDS

[Fin82] D. Fincham. Programs for the dynamic simulation of liquids and solids II. MDIONS: Rigid ions using the Ewald sum (vectorised version on the Cray-1). Computer Physics Communications, 25(2):177–179, February 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900339.

Fincham:1984:PDS

[Fin84a] D. Fincham. Programs for the dynamic simulation of liquids and solids II. Mdions: Rigid ions using the Ewald sum (vectorised version on the Cray-1). Computer Physics Communications, 35(1–3):C–779, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829069.

Fincham:1984:PMD

[Fin84b] David Fincham. Programs for the molecular dynamics simulation of liquids: I. Spherical molecules with short-ranged interactions. Computer Physics Communications, 35(1–3):C–665, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828210.

Fincham:1986:CTM

[Fin86] David Fincham. Choice of timestep in molecular dynamics simulation. Computer Physics Communications, 40(2–3):263– 269, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558690113X.

Fincham:1987:BRB

[Fin87] David Fincham. Book review: Computer Simulation Methods in Theoretical Physics: D. W. Heermann, Springer-Verlag, Berlin, Heidelberg, New York, London, Paris, Tokyo, 1986. 148 pages. ISBN 3-540-16966-0. DM 54 (hard cover). Computer Physics Communications, 46(1):203–204, July 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900476.

Friedman:1989:MPM

[FJ89] R. S. Friedman and M. J. Jamieson. Model potentials for multichannel eigenvalue problems. Computer Physics Communications, 55(2):137–140, September 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900726.

Ferguson:1984:PCXa

[FK84] I. F. Ferguson and J. E. Kirwan. A program for the calculation of X-ray reflection intensities. Computer Physics Communications, 35(1–3):C–187, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824339.

Fedorova:1986:DLB

[FK86a] R. N. Fedorova and V. V. Kornyak. Determination of Lie– Bäcklund symmetries of differential equations using formac. Computer Physics Communications, 39(1):93–103, January 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901657.

Feldmar:1986:RPM

[FK86b] E. Feldmar and K. S. Kölbig. Reduce procedures for the manipulation of generalized power series. Computer Physics Communications, 39(2):267–284, February/March 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901372.

Fraga:1987:RAS

[FKM^{+}87] Serafin Fraga, Mariusz Klobukowski, Janina Muszynska, K. M. S. Saxena, Jose A. Sordo, John D. Climenhaga, and Paul Clark. Research in atomic structure: a configuration interaction program with relativistic corrections. Computer Physics Communications, 47(1):159–172, October 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900750.

Feng:1984:EFR

[FKU^{+}84] D. H. Feng, B. T. Kim, T. Udagawa, T. Tamura, and K. S. Low. Exact finite-range microscopic calculations for heavy-ion induced two-nucleon transfer reactions. Computer Physics Communications, 35(1–3):C–404–C–405, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826144.

Flanders:1982:NNM

[Fla82] P. M. Flanders. Non-numerical methods on parallel computers. Computer Physics Communications, 26(3–4):363–371, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901308.

Flerackers:1981:FCP

[Fle81] E. Flerackers. FMEHAM: a computer program for triaxial Projected Hartree–Fock calculations. Computer Physics Communications, 22(2–3):215–221, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900527.

Flieller:1989:NIV

[Fli89] Sylvain Flieller. T.node, industrial version of supernode. Computer Physics Communications, 57(1–3):492–494, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902701.

Feneberg:1984:FCP

[FLM84] W. Feneberg, K. Lackner, and P. Martin. Fast control of the plasma surface. Computer Physics Communications, 31 (2–3):143–148, February 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900419.

Fluckiger:1989:OHW

[Flu89] Francois Fluckiger. Overview of HEP wide area networking: Producer perspective. Computer Physics Communications, 57(1–3):183–187, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902099.

Fijalkow:1984:RPP

[FM84] E. Fijalkow and G. Mourgues. Radiation potential of a point antenna immersed in drifting cold or hot (hydrodynamical) plasma. Computer Physics Communications, 35(1–3):C–598, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827678.

Flach:1985:BDS

[FM85] L. Flach and D. A. McNamara. Block data subprograms for finite element program packages. Computer Physics Communications, 36(2):223–224, April 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901262.

Friedman:1987:IRT

[FMS87] Jerome H. Friedman, John Alan McDonald, and Werner Stuetzle. An introduction to real-time graphical techniques for analyzing multivariate data. Computer Physics Communications, 45(1–3):161–167, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901512.

Fernandez-Niello:1989:CSC

[FNDL89] J. Fernández-Niello, C. H. Dasso, and S. Landowne. CCDEF — a simplified coupled-channel code for fusion cross sections including static nuclear deformations. Computer Physics Communications, 54(2–3):409–412, June/July 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901008.

Foglia:1982:NTM

[Fog82] C. Foglia. On number theoretical methods for multidimensional numerical integration. Computer Physics Communications, 25(2):113–118, February 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900273.

Foglia:1984:PPN

[Fog84] C. Foglia. A program for performing a numerical integration of the Schrödinger equation. Computer Physics Communications, 32(2):209–213, May 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900729.

Fortin:1989:SIM

[For89] M. Fortin. Some iterative methods for incompressible flow problems. Computer Physics Communications, 53(1–3):393– 399, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901756.

Fuchssteiner:1987:CAM

[FOW87] Benno Fuchssteiner, Walter Oevel, and Waldemar Wiwianka. Computer-algebra methods for investigation of hereditary operators of higher order soliton equations. Computer Physics Communications, 44(1–2):47–55, April/May 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900154.

Fisher:1989:UDM

[FP89] S. M. Fisher and P. Palazzi. Using a data model from software design to data analysis: What have we learned? Computer Physics Communications, 57(1–3):169–175, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902075.

Ferrero:1989:P

[FPR89] Armando M. J. Ferrero, Roberto P. J. Perazzo, and Silvia L. Reich. Preface. Computer Physics Communications, 56(1):xi, November 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900477.

Fincham:1981:MDS

[FR81] David Fincham and B. J. Ralston. Molecular dynamics simulation using the Cray-1 vector processing computer. Computer Physics Communications, 23(2):127–134, July 1, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900278.

Feilmeier:1982:PNL

[FR82] M. Feilmeier and W. Rönsch. Parallel non-linear algorithms. Computer Physics Communications, 26(3–4):335–348, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901266.

Ferguson:1984:PDC

[FR84] I. F. Ferguson and A. H. Rogerson. A program for the derivation of crystal unit cell parameters from X-ray powder diffraction measurements. Computer Physics Communications, 32(1):95– 107, April 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900110.

Fai:1986:FSE

[FR86a] George Fai and Jørgen Randrup. FREESCO: Statistical event generator for nuclear collisions. Computer Physics Communications, 42(3):385–397, November 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558690007X.

Fullerton:1986:GFD

[FR86b] L. W. Fullerton and G. A. Rinker. Generalized Fermi– Dirac integrals — FD, FDG, FDH. Computer Physics Communications, 39(2):181–185, February/March 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901293.

Franke:1981:TDP

[Fra81] G. Franke. Trigger and decision processors. Computer Physics Communications, 22(2–3):105–113, April 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900394.

Fraga:1983:DAD

[Fra83a] Serafin Fraga. Determination of antigenic determinants. Computer Physics Communications, 30(3):325–331, November 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900991.

Fraga:1983:MA

[Fra83b] Serafin Fraga. Molecular associations. Computer Physics Communications, 29(4):351–359, June 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465583900140.

Fraga:1984:MA

[Fra84a] Serafin Fraga. Molecular associations. Computer Physics Communications, 35(1–3):C–887, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0010465584829902.

Frage:1984:DAD

[Fra84b] Serafin Frage. Determination of antigenic determinants. Computer Physics Communications, 35(1–3):C–921, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584830192.

Frenkel:1987:CSH

[Fre87] D. Frenkel. Computer simulation of hard-core models for liquid crystals. Computer Physics Communications, 44(3):243– 253, June 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900798.

Fritts:1988:AGS

[Fri88] Martin J. Fritts. Adaptive gridding strategies for Free-Lagrangian calculations of low speed flows. Computer Physics Communications, 48(1):75–88, January 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900252.

FroeseFischer:1984:GMC

[Fro84a] Charlotte Froese Fischer. A general multi-configuration Hartree–Fock program. Computer Physics Communications, 35(1–3):C–459–C–460, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826612.

FroeseFischer:1984:MCHa

[Fro84b] Charlotte Froese Fischer. A multi-configuration Hartree–Fock program. Computer Physics Communications, 35(1–3):C–18, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584822870.

FroeseFischer:1984:MCHb

[Fro84c] Charlotte Froese Fischer. A multi-configuration Hartree– Fock program with improved stability. Computer Physics Communications, 35(1–3):C–146–C–147, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824005.

FroeseFischer:1987:GHF

[Fro87] Charlotte Froese Fischer. General Hartree–Fock program. Computer Physics Communications, 43(3):355–365, February/ March 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900531.

Faessler:1983:RIP

[FRS83] Amand Faessler, L. Rikus, and R. Sartor. Real and imaginary part of the heavy ion optical potential from a realistic nucleon-nucleon interaction. Computer Physics Communications, 28(3):275–286, January 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900450.

Faessler:1984:RIP

[FRS84] Amand Faessler, L. Rikus, and R. Sartor. Real and imaginary part of the heavy ion optical potential from a realistic nucleon-nucleon interaction. Computer Physics Communications, 35(1–3):C–852, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829616.

Fruhwirth:1981:PMF

[Frü81] R. Frühwirth. Preparing magnetic field measurements. Computer Physics Communications, 22(2–3):223–229, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900539.

Fonti:1980:LPI

[FS80] Luigi Fonti and Luciana Simoni. Large program implementation on minicomputers. Computer Physics Communications, 20(3):325–328, November 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900120.

Farkas:1982:DPG

[FS82] László Farkas and Péter Szöke. Digital program generator for long time measurement and process control. Computer Physics Communications, 26(1–2):205–207, May 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901709.

Fan:1984:IAI

[FS84a] L. S. Fan and W. Squire. Inversion of Abel’s integral equation by a direct method. Computer Physics Communications, 35(1–3):C–330, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825540.

Flower:1984:PCR

[FS84b] D. R. Flower and M. J. Seaton. A program to calculate radiative recombination coefficients of hydrogenic ions. Computer Physics Communications, 35(1–3):C–6, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822778.

FroeseFischer:1984:NVA

[FS84c] C. Froese Fischer and K. M. S. Saxena. A new version of AAKF (reduced tensor matrix elements) adapted to spectroscopic notation. Computer Physics Communications, 35(1–3): C–318, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825436.

FroeseFischer:1984:OSN

[FS84d] C. Froese Fischer and K. M. S. Saxena. Oscillator strengths from numerical MCHF radial functions. Computer Physics Communications, 35(1–3):C–319, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0010465584825448.

Fucito:1985:CPA

[FS85a] F. Fucito and S. Solomon. A concurrent pseudofermions algorithm. Computer Physics Communications, 36(2):141–145, April 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585901195.

Fucito:1985:MCP

[FS85b] F. Fucito and S. Solomon. Monte Carlo parallel algorithm for long range interactions. Computer Physics Communications, 34(3):225–230, January 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900013.

Francisco:1987:ROS

[FSP87] E. Francisco, L. Seijo, and L. Pueyo. Reduction of orbital sets. Computer Physics Communications, 43(2):269–277, January 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587902116.

Fowler:1984:FFI

[FSR84] R. H. Fowler, J. Smith, and J. A. Rome. FIFPC — a fast ion Fokker–Planck code. Computer Physics Communications, 35(1–3):C–432–C–433, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826429.

Faisal:1984:PCS

[FT84a] F. H. M. Faisal and A. L. V. Tench. A program for calculating the static interaction potential between an electron and a diatomic molecule. Computer Physics Communications, 35(1–3): C–90, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584823516.

Feng:1984:CHO

[FT84b] Da Hsuan Feng and T. Tamura. Calculations of harmonic oscillator brackets. Computer Physics Communications, 35(1–3): C–329, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825539.

Funsten:1984:NVD

[Fun84a] H. O. Funsten. A new version of DWPI (inelastic pionnucleus scattering) to incorporate microscopic form factors and differing proton and neutron radii. Computer Physics Communications, 35(1–3):C–551, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827277.

Funsten:1984:NVP

[Fun84b] H. O. Funsten. A new version of pirk (elastic pionnucleus scattering) to handle differing proton and neutron radii. Computer Physics Communications, 35(1–3):C–550, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827265.

Fredrikze:1981:CSB

[FV81] Henk Fredrikze and Peter Verkerk. Comment on “A splinebased method for experimental data deconvolution” by I. Beniaminy and M. Deutsch. Computer Physics Communications, 24(1):5–7, September/October 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901028.

Frank:1982:CAE

[FV82] Ludek Frank and Petr Vasina. Collection of Auger electron spectra with a microcomputer. Computer Physics Communications, 26(1–2):113–119, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901631.

Fack:1986:PCE

[FV86] V. Fack and G. Vanden Berghe. A program for the calculation of energy eigenvalues and eigenstates of a Schrödinger equation. Computer Physics Communications, 39(2):187–196, February/ March 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558690130X.

Gross:1984:WSM

[GA84] R. Gross and Y. Accad. Weizmann shell model computational code. Computer Physics Communications, 35(1–3):C–266, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482500X.

Gaines:1987:AMS

[GAA^{+}87] I. Gaines, H. Areti, R. Atac, J. Biel, A. Cook, M. Fischler, R. Hance, D. Husby, T. Nash, and T. Zmuda. The ACP multiprocessor system at Fermilab. Computer Physics Communications, 45(1–3):323–329, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558790169X.

Gadomski:1980:ASP

[Gad80] Adam M. Gadomski. Application of the singular perturbation method in nuclear reactor dynamics. Computer Physics Communications, 20(1):53–60, September 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580901071.

Gadjokov:1985:QOS

[Gad85] V. Gadjokov. Quasi-optimum scaling of symmetric positivedefinite matrices. Computer Physics Communications, 36 (1):19–24, March 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900165.

Gaemers:1981:NIP

[Gae81] K. J. F. Gaemers. Numerical inversion of the problem of moments. Computer Physics Communications, 22(2–3):115– 123, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900400.

Gaines:1989:TFD

[Gai89] I. Gaines. (trends in future) data acquisition architectures. Computer Physics Communications, 57(1–3):560, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902877.

Gajjar:1985:SSN

[Gaj85] J. S. B. Gajjar. On some solutions of the Navier– Stokes equations using a parallel processor. Computer Physics Communications, 37(1–3):303–309, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901663.

Galindo:1981:SBB

[Gal81] S. Galindo. Simulation of broad band NMR spectra of randomly oriented samples. Computer Physics Communications, 24(2):231–232, November 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900977.

Galli:1984:JPJ

[Gal84] M. Galli. JETS: a program for jet analysis of highenergy collision events. Computer Physics Communications, 34(1–2):135–143, November/December 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901656.

Gallopoulos:1985:MPP

[Gal85] E. J. Gallopoulos. The Massively Parallel Processor for problems in fluid dynamics. Computer Physics Communications, 37(1–3):311–315, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901675.

Gamgle:1989:BEF

[Gam89] John. N. Gamgle. Beyond Ethernet — future LAN’s. Computer Physics Communications, 57(1–3):129–133, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902014.

Gaskell:1981:GFH

[Gas81] R. W. Gaskell. Generating functions for L-harmonics. Computer Physics Communications, 24(2):191–196, November 1981.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558190093X.

Gaskell:1984:GFH

[Gas84] R. W. Gaskell. Generating functions for L-harmonics. Computer Physics Communications, 35(1–3):C–759, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828891.

Gather:1989:STP

[Gat89] Karl S. Gather. SASD-tools for program design. Computer Physics Communications, 57(1–3):29–36, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901896.

Gault:1981:PDB

[Gau81] F. D. Gault. Physics data bases and their use. Computer Physics Communications, 22(2–3):125–132, April 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900412.

Gault:1984:PDG

[Gau84a] F. D. Gault. The particle data group in the UK. Computer Physics Communications, 33(1–3):217–219, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901255.

Gault:1984:PDB

[Gau84b] F. D. Gault. The physics data base. Computer Physics Communications, 33(1–3):1–4, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901000.

Gault:1984:UDB

[Gau84c] F. D. Gault. The use of data base management software by the PDG(UK). Computer Physics Communications, 33(1–3):221–223, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901267.

Goto:1987:UAR

[GB87] F. Goto and H. C. Bolton. Use and accuracy of the Richardson deferred approach to the limit. Computer Physics Communications, 46(1):7–11, July 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900336.

Ganssauge:1989:TRP

[GDH^{+}89] E. Ganssauge, B. Dressel, S. Hackel, H. Kallies, E. Koch, Ch. Müller, J. T. Rhee, W. Schulz, and K. Ständecke. A track reconstruction program (TRP) for evaluation of nucleus-nucleus collisions in nuclear track emulsion chambers. Computer Physics Communications, 55(2):233–249, September 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900805.

Grest:1989:VLC

[GDK89] Gary S. Grest, Burkhard Dünweg, and Kurt Kremer. Vectorized link cell Fortran code for molecular dynamics simulations for a large number of particles. Computer Physics Communications, 55(3):269–285, October 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901252.

Giebink:1988:RRO

[GE88] David R. Giebink and David J. Ernst. ROMPIN — a relativistic optical model program for pion-nucleus scattering. Computer Physics Communications, 48(3):407–458, March 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902044.

Gellai:1985:MCP

[Gel85] Barbara Gellai. Molforce — a computer program for calculation of molecular force constants using the generalized inverse matrix. Computer Physics Communications, 36(2):177–189, April 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901225.

Genz:1982:NMI

[Gen82] Alan C. Genz. Numerical multiple integration on parallel computers. Computer Physics Communications, 26(3–4):349– 352, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901278.

Gendrin:1988:CR

[Gen88] R. Gendrin. Concluding remarks. Computer Physics Communications, 49(1):267–269, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588902354.

Gerdt:1980:ACH

[Ger80] V. P. Gerdt. Analytical calculations in high energy physics by computer. Computer Physics Communications, 20(1):85–90, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580901137.

Gerard:1988:LAN

[Ger88] J. Michael Gerard. Local area networks for microcomputers. Computer Physics Communications, 50(1–2):101–120, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901191.

Gorecki:1989:MDS

[GG89] J. Gorecki and J. Gryko. Molecular dynamics simulation of a chemical reaction. Computer Physics Communications, 54 (2–3):245–249, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900878.

Gellert:1983:HPE

[GH83] B. Gellert and J. Handke. HEIZ — a program to estimate temperature modifications in laser plasma interaction experiments by inverse Bremsstrahlung absorption and classical heat conduction. Computer Physics Communications, 30(2):169–175, September/October 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900590.

Gellert:1984:HPE

[GH84a] B. Gellert and J. Handke. HEIZ — a program to estimate temperature modifications in laser plasma interaction experiments by inverse Bremsstrahlung absorption and classical heat conduction. Computer Physics Communications, 35(1–3):C– 907, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584830064.

Glass:1984:ACE

[GH84b] R. Glass and A. Hibbert. Adaptation of CIV3 to evaluate hyperfine structure. Computer Physics Communications, 35 (1–3):C–368–C–369, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825849.

Glendinning:1987:TAF

[GH87] Ian Glendinning and Anthony Hey. Transputer arrays as FORTRAN farms for particle physics. Computer Physics Communications, 45(1–3):367–371, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901755.

Ghiselli:1987:DIG

[Ghi87] A. Ghiselli. A DECNET/IBM gateway for 3270 remote terminal access to IBM systems from VAX nodes of a DECNET network. Computer Physics Communications, 45(1–3):447–453, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901883.

Gils:1989:MTS

[GHO^{+}89] H. J. Gils, D. Heck, J. Oehlschläger, G. Schatz, T. Thouw, and A. Merkel. A multi-transputer system for parallel Monte Carlo simulations of extensive air showers. Computer Physics Communications, 56(2):105–113, December 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900118.

Gillet:1981:CDP

[GHU81] C. Gillet, J. Hackmann, and J. Uhlenbusch. Coupled description of plasma transport and plasma-wall processes. Computer Physics Communications, 24(3–4):301–309, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901521.

Gianturco:1984:NVP

[Gia84] F. A. Gianturco. A new version of a program calculating the static interaction potential between an electron and a diatomic molecule. Computer Physics Communications, 35(1–3): C–374, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825898.

Gibson:1985:PMT

[Gib85] J. K. Gibson. A production multi-tasking numerical weather prediction model. Computer Physics Communications, 37 (1–3):317–327, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901687.

Giorgilli:1984:CPI

[Gio84] Antonio Giorgilli. A computer program for integrals of motion. Computer Physics Communications, 35(1–3):C–544, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827216.

Giraud:1981:FQN

[Gir81] B. G. Giraud. A few questions of numerical and functional analysis for the many body problem in nuclear physics. Computer Physics Communications, 22(2–3):133–138, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900424.

Gadjokov:1984:NGU

[GJ84] V. Gadjokov and J. Jordanova. Numerical generation and use of orthonormal polynomials: I. ORT1- A one-dimensional package for the solution of fitting, differentiation and integration problems. Computer Physics Communications, 31(1):53–73, January 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900821.

Gjerpe:1981:FFP

[Gje81] I. Kenyon Gjerpe. A fast filter processor as a part of the trigger logic in an elastic scattering experiment. Computer Physics Communications, 22(2–3):239–252, April 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900552.

Grimm:1980:URP

[GK80] R. Grimm and H. Kühnelt. Using REDUCE in problems of supersymmetry and supergravity. Computer Physics Communications, 20(1):77–79, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580901113.

Gruber:1981:SBL

[GKM^{+}81] R. Gruber, W. Kerner, P. Merkel, J. Nührenberg, W. Schneider, and F. Troyon. Stability-beta limit of helical equilibria. Computer Physics Communications, 24(3–4):389–398, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901624.

Greenbaum:1989:PPC

[GLC89] Anne Greenbaum, Congming Li, and Han Zheng Chao. Parallelizing preconditioned conjugate gradient algorithms. Computer Physics Communications, 53(1–3):295–309, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901677.

Gliss:1982:IDM

[Gli82] B. Gliss. Integration of dedicated microcomputer satellites into a simple real-time system. Computer Physics Communications, 26(1–2):57–67, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558290159X.

Gliss:1984:SCD

[Gli84] B. Gliss. The small-computer data base system DVSR. Computer Physics Communications, 33(1–3):31–39, August/ September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901048.

Gorishny:1989:MPM

[GLST89] S. G. Gorishny, S. A. Larin, L. R. Surguladze, and F. V. Tkachov. Mincer: Program for multiloop calculations in quantum field theory for the Schoonschip system. Computer Physics Communications, 55(3):381–408, October 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901343.

Georgiopoulos:1989:VSR

[GM89] C. H. Georgiopoulos and M. E. Mermikides. Vectorised simulation of the response of a time projection chamber. Computer Physics Communications, 57(1–3):255–259, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902233.

Grant:1980:AMD

[GMN^{+}80] I. P. Grant, B. J. McKenzie, P. H. Norrington, D. F. Mayers, and N. C. Pyper. An atomic multiconfigurational Dirac–Fock package. Computer Physics Communications, 21(2):207–231, December 2, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900417.

Grant:1984:AMD

[GMN^{+}84] I. P. Grant, B. J. McKenzie, P. H. Norrington, D. F. Mayers, and N. C. Pyper. An atomic multiconfigurational Dirac–Fock package. Computer Physics Communications, 35(1–3):C–661– C–662, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828192.

Graves-Morris:1984:CCA

[GMR84a] P. R. Graves-Morris and D. E. Roberts. Calculation of Canterbury approximants. Computer Physics Communications, 35(1–3):C–341–C–342, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825631.

Graves-Morris:1984:SPR

[GMR84b] P. R. Graves-Morris and D. E. Roberts. A subroutine and procedure for the rapid calculation of simple off-diagonal rational approximants. Computer Physics Communications, 35(1–3):C– 289, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825199.

Godefroid:1984:AAC

[God84] M. Godefroid. An adaptation of acrz to calculate electric quadrupole oscillator strengths. Computer Physics Communications, 35(1–3):C–504, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826892.

Goedbloed:1981:CMM

[Goe81] J. P. Goedbloed. Conformal mapping methods in twodimensional magnetohydrodynamics. Computer Physics Communications, 24(3–4):311–321, December 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901533.

Goedbloed:1984:SRC

[Goe84] J. P. Goedbloed. Some remarks on computing axisymmetric equilibria. Computer Physics Communications, 31(2–3): 123–135, February 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900390.

Goldberg:1984:CBM

[Gol84a] J. Goldberg. Conversion of binary magnetic tapes. Computer Physics Communications, 35(1–3):C–47, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823139.

Golden:1984:ESI

[Gol84b] L. B. Golden. Exact Slater integrals. Computer Physics Communications, 35(1–3):C–465, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S001046558482665X.

Goorevich:1984:PAS

[Goo84] L. Goorevich. A program for the analytic simulation of extensive air showers. Computer Physics Communications, 35(1–3): C–251, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046558482487X.

Gorski:1988:FSR

[Gór88] Janusz Górski. Formal specification of real time systems. Computer Physics Communications, 50(1–2):71–88, July 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901178.

Goto:1983:UTP

[Got83] Fumiaki Goto. A useful technique for the point successive over relaxation method. Computer Physics Communications, 30(2):121–125, September/October 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900541.

Gournay:1988:GSS

[Gou88] J. F. Gournay. The graphics software of the Saclay Linear Accelerator Control System. Computer Physics Communications, 50(1–2):247–254, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901324.

Gruber:1981:NDI

[GPS^{+}81] R. Gruber, Ch. Pfersisch, S. Semenzato, F. Troyon, and T. Tsunematsu. On the numerical determination of ideal MHD limits of stability of axisymmetric toroidal configurations. Computer Physics Communications, 24(3–4):381–387, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901612.

Gaussorgues:1984:MMT

[GPS84] C. Gaussorgues, R. D. Piacentini, and A. Salin. Multistate molecular treatment of atomic collisions in the impact parameter approximation. I- integration of coupled equations and calculation of transition amplitudes for the straight line case. Computer Physics Communications, 35(1–3):C–340, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482562X.

Galvao:1981:ASM

[GR81] R. M. O. Galvão and J. Rem. Application of the Suydam method to the balloonings stability problem. Computer Physics Communications, 22(4):399–402, May 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901363.

Grant:1984:CFP

[Gra84a] I. P. Grant. CFPJJ — fractional parentage coefficients for equivalent electrons in jj-coupling. Computer Physics Communications, 35(1–3):C–164, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824157.

Grant:1984:GPC

[Gra84b] I. P. Grant. A general program to calculate angular momentum coefficients in relativistic atomic structure. Computer Physics Communications, 35(1–3):C–179–C–180, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824261.

Grant:1984:PCA

[Gra84c] I. P. Grant. A program to calculate angular momentum coefficients in relativistic atomic structure — revised version. Computer Physics Communications, 35(1–3):C–390–C–391, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826028.

Green:1988:SPA

[Gre88] James L. Green. The space physics analysis network. Computer Physics Communications, 49(1):205–213, April 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902287.

Griffiths:1987:BRB

[Gri87] David F. Griffiths. Book review: The State of the Art in Numerical Analysis, 1986: A. Iserles and M. J. D. Powell, eds., IMA Conference Series (New Series), Clarendon Press, Oxford, 1987. ISBN 0-19-853614-3. £55 (hard cover). Computer Physics Communications, 47(2–3):365–366, November/ December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901226.

Grimison:1989:IVF

[Gri89] A. Grimison. IBM’s view of the future. Computer Physics Communications, 57(1–3):560, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902865.

Groves:1984:CSM

[Gro84] J. L. Groves. Computer simulation of Mössbauer scattering spectra. Computer Physics Communications, 35(1–3):C–145, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584823991.

Grosdidier:1987:VFL

[Gro87] G. Grosdidier. A VAX-FPS loosely coupled array of processors. Computer Physics Communications, 45(1–3):361–365, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901743.

Grosdidier:1989:UIF

[Gro89a] G. Grosdidier. Use of an IBM-FPS loosely coupled array of processors for high energy physics programs. Computer Physics Communications, 52(2):207–222, January/February 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900052.

Grotendorst:1989:MPC

[Gro89b] J. Grotendorst. Maple programs for converting series expansions to rational functions using the Levin transformation. Automatic generation of FORTRAN functions for numerical applications. Computer Physics Communications, 55(3):325–335, October 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901306.

Gupta:1983:TCO

[GRRP83] R. P. Gupta, Rashmi-Rekha, and Satya Pal. Two centre overlap integrals of numerical wavefunctions. Computer Physics Communications, 29(1):87–95, March 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900292.

Gupta:1984:TCO

[GRRP84] R. P. Gupta, Rashmi-Rekha, and Satya Pal. Two centre overlap integrals of numerical wavefunctions. Computer Physics Communications, 35(1–3):C–866, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829720.

Gruber:1980:HES

[Gru80] R. Gruber. HYMNISBLOCK — eigenvalue solver for blocked matrices. Computer Physics Communications, 20(3):421–428, November 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900193.

Gruber:1981:P

[Gru81] Ralf Gruber. Preface. Computer Physics Communications, 24(3–4):v, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581901478.

Gruber:1984:HBM

[Gru84a] R. Gruber. Hymnia — band matrix package for solving eigenvalue problems. Computer Physics Communications, 35(1–3): C–325, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825485.

Gruber:1984:HES

[Gru84b] R. Gruber. Hymniablock — eigenvalue solver for blocked matrices. Computer Physics Communications, 35(1–3):C–645, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828052.

Gruber:1984:P

[Gru84c] Ralf Gruber. Preface. Computer Physics Communications, 31(2–3):v, February 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900365.

Gruschel:1984:FIR

[Gru84d] W. J. Gruschel. Freint, an integration routine calculating Fresnel diffraction. Computer Physics Communications, 35(1–3): C–528, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827095.

Grun:1983:AWC

[GS83] Norbert Grün and Werner Scheid. Absorption of waves by complex surface layers in the solution of the Schrödinger equation by finite difference methods. Computer Physics Communications, 30(3):243–247, November 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900917.

Gausterer:1988:RNS

[GS88] H. Gausterer and S. Sanielevici. Remarks on the numerical solution of Langevin equations on unitary groups spaces. Computer Physics Communications, 52(1):43–48, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901701.

Glasser:1989:PTD

[GS89] Alan H. Glasser and Kenneth Smith. POS — a 1D timedependent H^{+} ion source code. Computer Physics Communications, 55(3):409–424, October 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901355.

Gliss:1981:CFF

[GSS81] B. Gliss, H. Swars, and W. Schmettow. Common: a framework for the functional enhancement of a simple real-time system. Computer Physics Communications, 22(4):389–397, May 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901351.

Gruber:1981:HOE

[GST^{+}81] R. Gruber, S. Semenzato, F. Troyon, T. Tsunematsu, W. Kerner, P. Merkel, and W. Schneider. Hera and other extensions of erato. Computer Physics Communications, 24 (3–4):363–376, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581901594.

Gerdt:1985:FPC

[GSZ85] V. P. Gerdt, A. B. Shvachka, and A. Yu. Zharkov. Formint — a program for the classification of integrable nonlinear evolution equations. Computer Physics Communications, 34(3):303–311, January 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585900074.

Garcia-Torano:1983:FFF

[GT83] Eduardo Garcia-Toraño. FASTF: Fast Fourier Transform with arbitrary factors. Computer Physics Communications, 30(4):397–402, December 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900802.

Garcia-Torano:1984:FFF

[GT84] Eduardo Garcia-Toraño. Fastf: Fast Fourier Transform with arbitrary factors. Computer Physics Communications, 35(1–3): C–925, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584830234.

Gruber:1981:ESC

[GTB^{+}81] R. Gruber, F. Troyon, D. Berger, L. C. Bernard, S. Rousset, R. Schreiber, W. Kerner, W. Schneider, and K. V. Roberts. Erato stability code. Computer Physics Communications, 21(3):323–371, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900138.

Gruber:1984:ESC

[GTB^{+}84] R. Gruber, F. Troyon, D. Berger, L. C. Bernard, S. Rousset, R. Schreiber, W. Kerner, W. Schneider, and K. V. Roberts. Erato stability code. Computer Physics Communications, 35(1–3):C–671–C–672, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828258.

Garcia-Torano:1981:EPC

[GTG81] Eduardo Garcia-Toraño and Agustin Grau. Effy, a program to calculate the counting efficiency of beta particles in liquid scintillators. Computer Physics Communications, 23(4):385– 391, August 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558190179X.

Garcia-Torano:1984:EPC

[GTG84] Eduardo Garcia-Toraño and Agustín Grau. Effy, a program to calculate the counting efficiency of beta particles in liquid scintillators. Computer Physics Communications, 35(1–3):C– 737, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828714.

Geary:1986:TTD

[GTL^{+}86] J. L. Geary, T. Tajima, J-N. Leboeuf, E. G. Zaidman, and J. H. Han. Two- and three-dimensional magnetoinductive particle codes with guiding center electron motion. Computer Physics Communications, 42(3):313–331, November 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900020.

Garcia-Torano:1985:ENP

[GTM85] Eduardo Garcia-Toraño and Agustín Grau Malonda. EFFY, a new program to compute the counting efficiency of beta particles in liquid scintillators. Computer Physics Communications, 36(3):307–312, May 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900578.

Garcia-Torano:1987:EMC

[GTM87] E. García-Toraño and A. Grau Malonda. EFYGA, a Monte Carlo program to compute the interaction probability and the counting efficiency of gamma rays in liquid scintillators. Computer Physics Communications, 47(2–3):341–347, November/ December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901184.

Garcia-Torano:1988:ECE

[GTMA88] E. García-Toraño, A. Grau Malonda, and J. M. Los Arcos. EBEGA — the counting efficiency of a beta-gamma emitter in liquid scintillators. Computer Physics Communications, 50(3):313–319, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901877.

Gruber:1981:TEC

[GTR^{+}81] R. Gruber, F. Troyon, S. Rousset, W. Kerner, and L. C. Bernard. Transformation of ERATO into a δW code. Computer Physics Communications, 22(4):383–387, May 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558190134X.

Guest:1989:OVP

[Gue89] M. Guest. An overview of vector and parallel processors in scientific computation. Computer Physics Communications, 57(1–3):560, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589902853.

Gupta:1988:CRM

[Gup88a] Sumnesh Gupta. Comments regarding Monte Carlo simulation of classical fluids on general purpose supercomputers. Computer Physics Communications, 50(3):293–295, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901841.

Gupta:1988:VMD

[Gup88b] Sumnesh Gupta. Vectorization of molecular dynamics simulation for fluids of nonspherical molecules. Computer Physics Communications, 48(2):197–206, February 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900409.

Gurd:1985:MDM

[Gur85] J. R. Gurd. The Manchester Dataflow Machine. Computer Physics Communications, 37(1–3):49–62, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901353.

Ganzha:1987:SAD

[GV87] V. G. Ganzha and E. V. Vorozhtsov. The stability analysis of difference schemes by numerical solution of the generalised routh-hurwltz problem. Computer Physics Communications, 43(2):209–216, January 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587902062.

Gomes:1989:FDF

[GV89] M. A. F. Gomes and G. L. Vasconcelos. Fragmentation dynamics in fractal and Euclidean systems. Computer Physics Communications, 54(2–3):257–261, June/July 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900891.

Gil:1989:LMH

[GWL89] T. J. Gil, C. L. Winstead, and P. W. Langhoff. Lanczos methods for Hamiltonian spectra: Hilbert-space approximations to interaction-prepared states. Computer Physics Communications, 53(1–3):123–131, May 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901537.

Gray:1983:LSP

[GWTM83] P. C. Gray, J. S. Wagner, T. Tajima, and R. Million. Large scale particle simulations in a virtual memory computer. Computer Physics Communications, 30(2):109–120, September/October 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558390053X.

Heras:1982:NSN

[HA82] C. A. Heras and S. M. Abecasis. Nucore — a system for nuclear structure calculations with cluster-core models. Computer Physics Communications, 25(3):237–252, March 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900200.

Heras:1983:NTA

[HA83] C. A. Heras and S. M. Abecasis. NUCADA — two adaptations of the system nucore for nuclear structure calculations. Computer Physics Communications, 29(1):73–85, March 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900280.

Heras:1984:NTA

[HA84a] C. A. Heras and S. M. Abecasis. Nucada — two adaptations of the system nucore for nuclear structure calculations. Computer Physics Communications, 35(1–3):C–864–C–865, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829719.

Heras:1984:NSN

[HA84b] C. A. Heras and S. M. Abecasis. Nucore — a system for nuclear structure calculations with cluster-core models. Computer Physics Communications, 35(1–3):C–785–C–786, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829112.

Hall:1989:FLS

[HAG^{+}89] D. E. Hall, A. Agogino, W. H. Greiman, W. F. Johnston, D. Olson, R. Paasch, A. Padgaonkar, and D. W. Robertson. A fault location system for a time of flight detector array. Computer Physics Communications, 57(1–3):499–502, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902725.

Halling:1988:SMA

[Hal88] H. Halling. Standard microcomputer architectures. Computer Physics Communications, 50(1–2):33–49, July 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901142.

Hansen:1984:CPG

[Han84a] Flemming Yssing Hansen. I. A computer program for generation of a complete set of coordinates and force matrices for normal mode calculations of crystals and molecules. Computer Physics Communications, 35(1–3):C–461, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826624.

Hansen:1984:CPN

[Han84b] Flemming Yssing Hansen. I: a computer program for normalization and instrument correction of neutron diffraction data on non-crystalline materials to obtain the static structure factor. Computer Physics Communications, 35(1–3):C–511–C– 512, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584826958.

Hansen:1984:ICPa

[Han84c] Flemming Yssing Hansen. II. A computer program for calculation of parameters necessary for the computation of reliable pair distribution functions of non-crystalline materials from limited diffraction data. Computer Physics Communications, 35(1–3):C–513–C–514, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482696X.

Hansen:1984:IPC

[Han84d] Flemming Yssing Hansen. II. A program for computing normal modes of molecules, crystal phonon dispersion relations and structure factors for neutron inelastic scattering. Computer Physics Communications, 35(1–3):C–462–C–463, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826636.

Hansen:1984:ICPb

[Han84e] Flemming Yssing Hansen. III. A computer program for calculation of reliable pair distribution functions of noncrystalline materials from limited diffraction data. Computer Physics Communications, 35(1–3):C–515–C–516, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826971.

Hansen:1984:IFC

[Han84f] Flemming Yssing Hansen. III. A force constant adjuster program to obtain least squares fit to observed frequencies of molecules and crystals. Computer Physics Communications, 35(1–3):C–464, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826648.

Harris:1981:CCS

[Har81] M. J. Harris. Calculation of (n, γ) cross-sections and astrophysical reaction rates by the nuclear statistical model. Computer Physics Communications, 21(3):407–419, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900175.

Harris:1984:CCS

[Har84a] M. J. Harris. Calculation of (n, γ) cross-sections and astrophysical reaction rates by the nuclear statistical model. Computer Physics Communications, 35(1–3):C–676, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828295.

Hartmann:1984:PDB

[Har84b] G. K. Hartmann. Problems of data bases in geophysics. Computer Physics Communications, 33(1–3):267–287, August/ September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901322.

Harlow:1988:PP

[Har88] Francis H. Harlow. PIC and its progeny. Computer Physics Communications, 48(1):1–10, January 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900173.

Harper:1989:VRP

[Har89] David Harper. Vector33: a REDUCE program for vector algebra and calculus in orthogonal curvilinear coordinates. Computer Physics Communications, 54(2–3):295–305, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900921.

Hasse:1984:ODW

[Has84] Rainer W. Hasse. One-dimensional wave packet solutions of time-dependent frictional or optical potential Schrödinger equations. Computer Physics Communications, 35(1–3):C–385, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825989.

Hauser:1987:VTF

[Hau87] Jay Hauser. Vectorization of track finding and fitting algorithms for HEP. Computer Physics Communications, 45(1–3):121–125, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901469.

Haydock:1980:RSS

[Hay80] Roger Haydock. The recursive solution of the Schrödinger equation. Computer Physics Communications, 20(1):11–16, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580901010.

Haydock:1983:RER

[Hay83] Roger Haydock. Rounding errors in the recursion method: Inner products. Computer Physics Communications, 30(3): 221–228, November 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900887.

Haydock:1989:ARM

[Hay89a] Roger Haydock. Accuracy of the recursion method and basis non-orthogonality. Computer Physics Communications, 53 (1–3):133–139, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901549.

Haydock:1989:RTI

[Hay89b] Roger Haydock. Recursive tridiagonalization of infinite dimensional Hamiltonians. Computer Physics Communications, 55(1):1–3, August 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900593.

Hooker:1985:CSP

[HB85] Stanford B. Hooker and James W. Brown. CENTER: a software package for center estimation. Computer Physics Communications, 38(3):421–433, December 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901109.

Holzman:1988:CGT

[HB88] Robert E. Holzman and James F. Blinn. Computer graphics techniques and computer-generated movies. Computer Physics Communications, 49(1):229–233, April 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902305.

Handler:1985:TCH

[HBF^{+}85] W. Händler, A. Bode, G. Fritsch, W. Henning, and J. Volkert. A tightly coupled and hierarchical multiprocessor architecture. Computer Physics Communications, 37(1–3):87–93, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901390.

Hegyi:1984:AAP

[HBJ84] György Hegyi, Liviu Barbu, and L. Jakab. Adaptation of the ACQN program to calculate transport collision integrals on minicomputers. Computer Physics Communications, 34(1–2):219–222, November/December 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901735.

Hallam-Baker:1989:UOZ

[HBM89] P. M. Hallam-Baker and I. C. McArthur. Use of Occam in ZEUS. Computer Physics Communications, 57(1–3):520–523, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589902786.

Habitz:1983:GPC

[HC83] Peter Habitz and Enrico Clementi. A general program to compute two electron repulsion integrals. Computer Physics Communications, 29(3):301–306, May 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900097.

Habitz:1984:GPC

[HC84] Peter Habitz and Enrico Clementi. A general program to compute two electron repulsion integrals. Computer Physics Communications, 35(1–3):C–883, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829860.

Hirshman:1986:NCH

[HC86] S. P. Hirshman and W. A. Cooper. Numerical computation of helically symmetric equilibria. Computer Physics Communications, 42(1):37–41, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586902286.

Halonen:1988:LMV

[HC88] L. Halonen and M. S. Child. Local mode vibrations in tetrahedral molecules. Computer Physics Communications, 51(1–2):173–193, September/October 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900707.

Halcomb:1986:ILS

[HD86] Lawrence L. Halcomb and Dennis J. Diestler. Integration of a largest set of coupled differential equations on the CY-BER 205 vector processor. Computer Physics Communications, 39(1):27–35, January 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558690161X.

Heermann:1988:IMT

[HD88] Dieter W. Heermann and Rashmi C. Desai. Ising model, transputers and dynamical correlations using a microcanonical ensemble. Computer Physics Communications, 50(3):297–302, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901853.

Hartmann:1989:EGS

[HD89] U. Hartmann and E. D. Davis. epic grass — symbolic calculations with anticommuting variables. Computer Physics Communications, 54(2–3):353–369, June/July 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900969.

Hartwig:1984:RBB

[HDK84] A. Hartwig, F. Daske, and S. Kobe. A recursive branchand-bound algorithm for the exact ground state of Ising spinglass models. Computer Physics Communications, 32(2):133– 138, May 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900663.

Herbison-evans:1984:DOD

[He84a] D. Herbison-evans. Derivation of the orbit of a double star from observations made with an intensity interferometer. Computer Physics Communications, 35(1–3):C–66, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823310.

Herbison-Evans:1984:LCV

[HE84b] D. Herbison-Evans. The light curve of a variable star subject to orbital tidal distortion. Computer Physics Communications, 35(1–3):C–185–C–186, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824327.

Heiss:1984:CCS

[Hei84] P. Heiss. Computation of cross sections and polarizations for nuclear reactions, in which only spin 1/2 particles are involved. Computer Physics Communications, 35(1–3):C–163, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824145.

Helfrich:1980:LRF

[Hel80] K. Helfrich. Local representation of functions of D variables by means of discrete orthogonal polynomials. Computer Physics Communications, 20(1):61–63, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580901083.

Hellier:1982:DIW

[Hel82] Richard L. Hellier. DAP implementation of the WZ algorithm. Computer Physics Communications, 26(3–4):321–323, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901230.

Herbison-Evans:1984:AVS

[HEL84] D. Herbison-Evans and N. R. Lomb. Analysis of a variable spectroscopic double star. Computer Physics Communications, 35(1–3):C–105, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823632.

Henry:1984:RCS

[Hen84] Ronald J. W. Henry. Rovibrational cross sections from reactance matrices calculated in adiabatic nuclei approximation. Computer Physics Communications, 35(1–3):C–349, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825709.

Hernando:1982:NEI

[Her82a] J. A. Hernando. Newton–Everett interpolation of continuous functions. Computer Physics Communications, 27(1):73– 78, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900091.

Hertzberger:1982:FAM

[Her82b] L. O. Hertzberger. The fast Amsterdam multiprocessor (FAMP) system. Computer Physics Communications, 26(1–2): 79–97, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901618.

Hermsdorf:1984:UND

[Her84a] D. Hermsdorf. On the use of nuclear data libraries in science and technology. Computer Physics Communications, 33(1–3):147–153, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901164.

Hernando:1984:NEI

[Her84b] J. A. Hernando. Newton–Everett interpolation of continuous functions. Computer Physics Communications, 35(1–3): C–808, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829276.

Hernquist:1988:HBM

[Her88a] Lars Hernquist. Hierarchical N-body methods. Computer Physics Communications, 48(1):107–115, January 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900288.

Hertzberger:1988:MT

[Her88b] L. O. Hertzberger. Microprocessor technology. Computer Physics Communications, 50(1–2):1–21, July 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901129.

Hertzberger:1989:DHS

[Her89] L. O. Hertzberger. Does HEP still hold challenges for computer science? Computer Physics Communications, 57(1–3):15–22, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901872.

Hey:1988:RAT

[Hey88] Anthony J. G. Hey. RISC architecture in transputers and transputer arrays. Computer Physics Communications, 50(1–2):23– 31, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901130.

Hey:1989:RMA

[Hey89] Anthony J. G. Hey. The role of MIMD arrays of transputers in computational physics. Computer Physics Communications, 56(1):1–24, November 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900490.

Hibbert:1988:NOO

[HFG88] A. Hibbert, C. Froese Fischer, and M. R. Godefroid. Nonorthogonal orbitals in MCHF or configuration interaction wave functions. Computer Physics Communications, 51(3):285–293, November 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901415.

Hall:1989:SBV

[HGJ^{+}89] D. E. Hall, W. H. Greiman, W. F. Johnston, A. X. Merola, S. C. Loken, and D. W. Robertson. The software bus: a vision for scientific software development. Computer Physics Communications, 57(1–3):211–216, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902142.

Hertzberger:1981:FS

[HGK^{+}81] L. O. Hertzberger, D. Gosman, G. Kieft, G. J. A. Por, M. Schoorel, and L. W. Wiggers. Famp system. Computer Physics Communications, 22(2–3):253–260, April 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900564.

Horn:1986:SBM

[HGW86] Richard Horn, Rudolf Goldflam, and Lawrence Wilets. Soliton bag model. Computer Physics Communications, 42(1):105–117, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586902353.

Hain:1982:NSM

[HH82] Klaus Hain and Gertrud Hain. Numerical simulations of magnetohydrodynamic chemically reactive flows in three dimensions on a vector computer. Computer Physics Communications, 27(4):325–333, October 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900959.

Hird:1984:DQS

[HH84] B. Hird and K. H. Huang. Deformed quasiparticle states in a Woods–Saxon potential and coupled to rotational states of the core. Computer Physics Communications, 35(1–3):C–344, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825667.

Hanzalek:1986:CPS

[HHK86] Petr Hanzálek, Jirí Hrebícek, and Jan Kucera. A conversational program system for mathematical optimization. Computer Physics Communications, 41(2–3):403–408, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900809.

Hibbert:1984:AGPb

[Hib84a] A. Hibbert. Adaptation of a general program to calculate angular momentum integrals in atomic structure: Inclusion of the checking of the configuration data. Computer Physics Communications, 35(1–3):C–281, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825126.

Hibbert:1984:AGPa

[Hib84b] A. Hibbert. Adaptation of a general program to calculate angular momentum integrals in atomic structure: Inclusion of the one-electron part of the Hamiltonian. Computer Physics Communications, 35(1–3):C–248, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824844.

Hibbert:1984:CGP

[Hib84c] A. Hibbert. CIV3 — a general program to calculate configuration interaction wave functions and electric-dipole oscillator strengths. Computer Physics Communications, 35(1–3):C–298– C–299, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046558482528X.

Hibbert:1984:GPC

[Hib84d] A. Hibbert. A general program for calculating angular momentum integrals in atomic structure. Computer Physics Communications, 35(1–3):C–43, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823097.

Hibbert:1984:NVG

[Hib84e] A. Hibbert. A new version of a general program to calculate angular momentum integrals in atomic structure. Computer Physics Communications, 35(1–3):C–80–C–81, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823437.

Harrison:1984:CER

[HIM84] D. J. Harrison, A. C. Irving, and A. D. Martin. Computation of effective Regge trajectories for high energy two-body reactions. Computer Physics Communications, 35(1–3):C–172, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824212.

Hine:1981:SDT

[Hin81] M. G. N. Hine. Satellite data transmission in high energy physics. Computer Physics Communications, 22(2–3):139–148, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900436.

Hird:1984:NOP

[Hir84] B. Hird. Nilsson orbits for a particle in a Woods–Saxon potential with Y_{2}^{0} and Y_{4}^{0} deformations, and coupled to core rotational states. Computer Physics Communications, 35(1–3): C–206, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824480.

Hons:1981:PCC

[HK81] Z. Hons and J. Kvasil. A program for calculation of the Coriolis effect in odd-odd nuclei. Computer Physics Communications, 24(2):161–172, November 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900898.

Hons:1983:PCE

[HK83] Z. Hons and J. Kvasil. A program for calculation of the E1, E2 and M1 transition probabilities in odd-odd nuclei taking the Coriolis mixing into account. Computer Physics Communications, 30(1):59–69, July/August 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583901236.

Hons:1984:PCC

[HK84a] Z. Hons and J. Kvasil. A program for calculation of the Coriolis effect in ODD-ODD nuclei. Computer Physics Communications, 35(1–3):C–754–C–755, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828854.

Hons:1984:PCE

[HK84b] Z. Hons and J. Kvasil. A program for calculation of the E1, E2 and M1 transition probabilities in odd-odd nuclei taking the Coriolis mixing into account. Computer Physics Communications, 35(1–3):C–899–C–900, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829999.

Hummer:1984:NEF

[HKK84] D. G. Hummer, C. V. Kunasz, and P. B. Kunasz. Numerical evaluation of the formal solution of radiative transfer problems in spherical geometries. Computer Physics Communications, 35(1–3):C–207, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824492.

Huzinaga:1983:AGT

[HKS83] S. Huzinaga, M. Klobukowski, and Y. Sakai. An atomic Gaussian-type orbital Roothaan–Hartree–Fock program. Computer Physics Communications, 30(3):311–323, November 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558390098X.

Huzinaga:1984:AGT

[HKS84] S. Huzinaga, M. Klobukowski, and Y. Sakai. An atomic Gaussian-type orbital Roothaan–Hartree–Fock program. Computer Physics Communications, 35(1–3):C–919–C–920, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584830180.

Heddle:1985:CPS

[HKT85] David P. Heddle, Yong Rae Kwon, and F. Tabakin. Coulomb plus strong interaction bound states — momentum space numerical solutions. Computer Physics Communications, 38(1): 71–82, August/September 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900475.

Hirshman:1986:MSC

[HL86] S. P. Hirshman and D. K. Lee. MOMCON: a spectral code for obtaining three-dimensional magnetohydrodynamic equilibria. Computer Physics Communications, 39(2):161–172, February/March 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558690127X.

Hewett:1988:RPA

[HL88] Dennis W. Hewett and A. Bruce Langdon. Recent progress with AVANTI: a 2.5D EM direct implicit PIC code. Computer Physics Communications, 48(1):127–133, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900306.

Hlavaty:1980:NDS

[Hla80] Ladislav Hlavatý. Nuclear decay scheme construction based on qualitative coincidences. Computer Physics Communications, 19(2):197–204, April/June 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900508.

Hlavaty:1984:NDS

[Hla84] Ladislav Hlavatý. Nuclear decay scheme construction based on qualitative coincidences. Computer Physics Communications, 35(1–3):C–616, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827812.

Hlavaty:1986:TRP

[Hla86] Ladislav Hlavatý. Tests of resonances in the Painlevé analysis. Computer Physics Communications, 42(3):427–433, November 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558690010X.

Huard:1987:MBS

[HLM^{+}87] A. Huard, P. Laigle, V. Mastrangelo, M. Talbi, and S. Xhemalce. A method based on a stochastic approach for space dependent nuclear reactor kinetics in one dimension. Computer Physics Communications, 46(3):351–377, September 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900920.

Hirschi:1984:CCC

[HLS84] S. Hirschi, E. Lomon, and N. Spencer. Cochase, a code for coupled channel Schrödinger equations. Computer Physics Communications, 35(1–3):C–287, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825175.

Hernando:1981:PFP

[HM81] J. A. Hernando and V. Massidda. Plattsum: a Fortran program that evaluates electrostatic lattice sums by the planewise summation method. Computer Physics Communications, 22(1):13–31, February/March 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558190076X.

Hernando:1983:MPP

[HM83] J. A. Hernando and V. Massidda. A modification to PLATTSUM, a program that evaluates electrostatic lattice sums by the planewise summation method. Computer Physics Communications, 30(4):403–409, December 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900814.

Hernando:1984:MPP

[HM84a] J. A. Hernando and V. Massidda. A modification to plattsum, a program that evaluates electrostatic lattice sums by the planewise summation method. Computer Physics Communications, 35(1–3):C–926, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584830246.

Hernando:1984:PFP

[HM84b] J. A. Hernando and V. Massidda. Plattsum: a Fortran program that evaluates electrostatic lattice sums by the planewise summation method. Computer Physics Communications, 35(1–3):C–680–C–681, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828337.

Hoffstein:1984:SBSa

[HM84c] V. Hoffstein and O. Moller. Symmetry and band structure I. Selection of reciprocal lattice vectors. Computer Physics Communications, 35(1–3):C–58, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823243.

Hoffstein:1984:SBSb

[HM84d] V. Hoffstein and O. Moller. Symmetry and band structure II. Storage and retrieval of group theoretical information. Computer Physics Communications, 35(1–3):C–59, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823255.

Hoffstein:1984:SBSc

[HM84e] V. Hoffstein and O. Moller. Symmetry and band structure III. Construction of symmetrized Hamiltonian matrix. Computer Physics Communications, 35(1–3):C–60, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823267.

Herman:1984:PCS

[HMS84] M. Herman, A. Marcinkowski, and K. Stankiewicz. A program for calculation of spectra and cross sections within the combined pre-equilibrium/compound nucleus model of nuclear reactions. Computer Physics Communications, 33(4):373–398, October 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901449.

Hoshino:1984:MCS

[HMTO84] Tsutomu Hoshino, Sumiko Majima, Kiyo Takenouchi, and Yoshio Oyanagi. Monte Carlo simulation of a spin model on the parallel computer “PAX”. Computer Physics Communications, 34(1–2):31–38, November/December 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901565.

Hoare:1989:FMC

[Hoa89] C. A. R. Hoare. Formal methods in computer system design. Computer Physics Communications, 57(1–3):206–210, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589902130.

Hockney:1982:CPC

[Hoc82] R. W. Hockney. Characterization of parallel computers and algorithms. Computer Physics Communications, 26(3–4):285– 291, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901187.

Hodgson:1984:MER

[Hod84] R. J. W. Hodgson. Matrix elements of the reaction matrix in finite nuclei. Computer Physics Communications, 35 (1–3):C–366–C–367, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825837.

Hoek:1986:CLG

[Hoe86a] Jaap Hoek. Cooling of SU(3) lattice gauge field configurations. Computer Physics Communications, 39(1):21–26, January 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586901608.

Hoek:1986:PBE

[Hoe86b] Jaap Hoek. Propagation of boundary effects in large systems on vector processors. Computer Physics Communications, 42(2):169–173, October 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900342.

Hoek:1989:UAT

[Hoe89] Jaap Hoek. Use of attached transputer hardware to VAX’s for offline analysis. Computer Physics Communications, 57 (1–3):503–504, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589902737.

Hoffstein:1984:PCLb

[Hof84a] V. Hoffstein. Program for calculating LEED intensities based on the inelastic collision model: I. Matrix inversion method. Computer Physics Communications, 35(1–3):C–226, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824650.

Hoffstein:1984:PCLa

[Hof84b]  V. Hoffstein.  Program for calculating LEED intensi-

ties using band structure-matching formalism. Computer Physics Communications, 35(1–3):C–102, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823607.

Hofmann:1988:FFB

[Hof88] F. Hofmann. FBT — a free-boundary tokamak equilibrium code for highly elongated and shaped plasmas. Computer Physics Communications, 48(2):207–221, February 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900410.

Hopgood:1985:GS

[Hop85] F. R. A. Hopgood. Graphics standards. Computer Physics Communications, 38(2):233–243, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558590089X.

Houseman:1987:TVP

[Hou87] Gregory A. Houseman. TDPOIS, a vector-processor routine for the solution of the three-dimensional Poisson and biharmonic equations in a rectangular prism. Computer Physics Communications, 43(2):257–267, January 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587902104.

Heifetz:1983:SMC

[HP83] D. B. Heifetz and D. E. Post. SEURAT: a Monte– Carlo algorithm for calculating neutral gas transport in non-circular axisymmetric toroidal plasmas. Computer Physics Communications, 29(3):287–299, May 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900085.

Heifetz:1984:SMC

[HP84] D. B. Heifetz and D. E. Post. Seurat: a Monte-Carlo algorithm for calculating neutral gas transport in non-circular axisymmetric toroidal plasmas. Computer Physics Communications, 35(1–3):C–882, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829859.

Hopkinson:1980:CPS

[HPT80] J. F. L. Hopkinson, J. B. Pendry, and D. J. Titterington. Calculation of photoemission spectra for surfaces of solids. Computer Physics Communications, 19(1):69–92, January/March 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900685.

Hender:1981:LNR

[HR81] T. C. Hender and D. C. Robinson. Linear and nonlinear resistive instability studies. Computer Physics Communications, 24(3–4):413–419, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558190165X.

Hughes:1983:OC

[HR83a] M. H. Hughes and K. V. Roberts. OLYMPUS conventions. Computer Physics Communications, 29(1):15–43, March 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900255.

Hughes:1983:OFC

[HR83b] M. H. Hughes and K. V. Roberts. The OLYMPUS Fortran compositor. Computer Physics Communications, 29(1):45–57, March 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900267.

Hughes:1983:OFG

[HR83c] M. H. Hughes and K. V. Roberts. The OLYMPUS Fortran generator. Computer Physics Communications, 29(1):59–71, March 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900279.

Hanssgen:1984:MCC

[HR84a] K. Hänssgen and S. Ritter. The Monte Carlo code decay to simulate the decay of baryon and meson resonances. Computer Physics Communications, 31(4):411–418, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900249.

Hughes:1984:TSI

[HR84b] M. H. Hughes and A. P. V. Roberts. Timer — a software instrumentation routine for making timing measurements. Computer Physics Communications, 35(1–3):C–267, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825011.

Hughes:1984:OFC

[HR84c] M. H. Hughes and K. V. Roberts. The Olympus Fortran compositor. Computer Physics Communications, 35(1–3):C– 862, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829690.

Hughes:1984:OFG

[HR84d] M. H. Hughes and K. V. Roberts. The Olympus Fortran generator. Computer Physics Communications, 35(1–3):C–863, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829707.

Hughes:1984:ORF

[HR84e] M. H. Hughes and K. V. Roberts. Olympus restart facilities. Computer Physics Communications, 35(1–3):C–268, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825023.

Hanssgen:1986:MCCa

[HR86a] K. Hänssgen and J. Ranft. The Monte Carlo code hadrin to simulate inelastic hadron-nucleon interactions at laboratory energies below 5 GeV. Computer Physics Communications, 39(1):37–51, January 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901621.

Hanssgen:1986:MCCb

[HR86b] K. Hänssgen and J. Ranft. The Monte Carlo code nucrin to simulate inelastic hadron–nucleus interactions at laboratory energies below 5 GeV. Computer Physics Communications, 39(1):53–70, January 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901633.

Herman:1987:CCC

[HR87] M. Herman and G. Reffo. Codes for the combinatorial calculation of few quasiparticle state densities in spherical and deformed nuclei. Computer Physics Communications, 47(1):103– 111, October 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900701.

Hoffstein:1984:SPC

[HRB84] V. Hoffstein, D. K. Ray, and M. Belakhovsky. Symmetrized program for calculating energy bands and electronic structure of solids. Computer Physics Communications, 35(1–3):C–162, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824133.

Hughes:1984:OCU

[HRL84] M. H. Hughes, K. V. Roberts, and G. G. Lister. Olympus control and utility package for the CDC 6500. Computer Physics Communications, 35(1–3):C–336–C–337, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482559X.

Hughes:1984:OPP

[HRR84] M. H. Hughes, K. V. Roberts, and P. D. Roberts. Olympus and preprocessor package for an IBM 370/165. Computer Physics Communications, 35(1–3):C–290, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825205.

Henry:1981:GPC

[HRS81] Ronald J. W. Henry, S. P. Rountree, and Ed R. Smith. A general program to calculate atomic continuum processes using the NIEM method. Computer Physics Communications, 23(3):233– 273, July 2, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900023.

Henry:1984:GPC

[HRS84] Ronald J. W. Henry, S. P. Rountree, and Ed R. Smith. A general program to calculate atomic continuum processes using the NIEM method. Computer Physics Communications, 35(1–3):C–723–C–727, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828635.

Harp:1985:SPT

[HRW85] J. G. Harp, J. B. G. Roberts, and J. S. Ward. Signal processing with transputer arrays (TRAPS). Computer Physics Communications, 37(1–3):77–86, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901389.

Hughes:1984:ADS

[HS84] J. G. Hughes and F. J. Smith. AMDS: a database system for atomic and molecular physics. Computer Physics Communications, 32(3):317–331, June 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900948.

Hult:1982:SDP

[HSA82] T. P. Hult, S. P. Svensson, and T. G. Andersson. A spectrum data processing system. Computer Physics Communications, 25(4):417–431, April 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900789.

Hult:1984:SDP

[HSA84] T. P. Hult, S. P. Svensson, and T. G. Andersson. A spectrum data processing system. Computer Physics Communications, 35(1–3):C–799, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829203.

Hughes:1984:AMD

[HSB^{+}84] J. G. Hughes, F. J. Smith, K. A. Berrington, K. M. Aggarwal, and M. Elder. The atomic and molecular database at Belfast and Daresbury. Computer Physics Communications, 33(1–3):99–103, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901127.

Hoheisel:1984:VCC

[HSV84] C. Hoheisel, M. Schoen, and R. Vogelsang. Vectorized computation of correlation functions from phase space trajectories generated by molecular dynamics calculations. Computer Physics Communications, 34(1–2):9–13, November/December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901541.

Hoshino:1984:PMD

[HT84] Tsutomu Hoshino and Kiyo Takenouchi. Processing of the molecular dynamics model by the parallel computer pax. Computer Physics Communications, 31(4):287–296, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900134.

Hansen:1988:PPC

[HT88] J. P. Hansen and K. Taulbjerg. A preorthonormalization procedure for coupled channel problems. Computer Physics Communications, 51(3):317–321, November 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901440.

Hubbard:1984:AVC

[Hub84a] Lincoln B. Hubbard. Allowed values of coupled angular momentum and i-spin for nucleons in a single shell in j-j coupling. Computer Physics Communications, 35(1–3):C–53, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823188.

Hubbard:1984:CFP

[Hub84b] Lincoln B. Hubbard. Coefficients of fractional parentage in j-j coupling in the isospin representation. Computer Physics Communications, 35(1–3):C–28, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822961.

Hubbard:1984:FCG

[Hub84c] Lincoln B. Hubbard. First collision gamma-ray dose. Computer Physics Communications, 35(1–3):C–222, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824625.

Hubbard:1984:GRD

[Hub84d] Lincoln B. Hubbard. Gamma-radiation dosimetry for arbitrary source and target geometry. Computer Physics Communications, 35(1–3):C–111–C–112, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823693.

Hubert:1984:SNWa

[Hub84e] A. Hubert. Symmetric Néel walls in thin magnetic films. Computer Physics Communications, 35(1–3):C–40, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823073.

Hubert:1984:SNWb

[Hub84f] A. Hubert. Symmetric Néel walls in thin magnetic films. An adaptation to increase the range of convergence. Computer Physics Communications, 35(1–3):C–56, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823218.

Hubeny:1988:CPC

[Hub88] I. Hubeny. A computer program for calculating non-LTE model stellar atmospheres. Computer Physics Communications, 52(1):103–132, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901774.

Huberman:1989:PC

[Hub89] Bernardo A. Huberman. Parallel computation. Computer Physics Communications, 56(1):25–42, November 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900507.

Hughes:1984:SPE

[Hug84] M. H. Hughes. Solution of Poisson’s equation in cylindrical coordinates. Computer Physics Communications, 35(1–3):C– 76, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584823395.

Hutton:1989:FSN

[Hut89] James Hutton. Future scientific networking. Computer Physics Communications, 57(1–3):188–190, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902105.

Hirshman:1986:TDF

[HvRM86] S. P. Hirshman, W. I. van RIJ, and P. Merkel. Threedimensional free boundary calculations using a spectral Green’s function method. Computer Physics Communications, 43 (1):143–155, December 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900585.

Hoheisel:1987:CSM

[HVS87] C. Hoheisel, R. Vogelsang, and M. Schoen. Comparison of the Singer method and the constraints method for molecular dynamics with linear molecules on the vector computer CY-BER 205. Computer Physics Communications, 43(2):217–229, January 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587902074.

Haydock:1984:CAR

[HW84a] Roger Haydock and D. Weaire. Computational advantages of the recursion and Lanczos methods. Computer Physics Communications, 31(4):431–432, March 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900274.

Hicks:1984:LNI

[HW84b] H. R. Hicks and J. W. Wooten. Linear and nonlinear ideal MHD codes — V103. Computer Physics Communications, 35(1–3):C–419–C–420, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826296.

Hsiao:1989:COQ

[HWL89] Ming-Yuan Hsiao, K. A. Werley, and K. M. Ling. CFRX, a one-and-a-quarter-dimensional transport code for field-reversed configuration studies. Computer Physics Communications, 54 (2–3):329–352, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900957.

Indrea:1980:FAE

[IA80] E. Indrea and N. Aldea. Fourier analysis of EXAFS data, a self-contained Fortran program-package. Computer Physics Communications, 21(1):91–96, December 1, 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900788.

Indrea:1984:FAE

[IA84] E. Indrea and N. Aldea. Fourier analysis of exafs data, a selfcontained Fortran program-package. Computer Physics Communications, 35(1–3):C–652, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S001046558482812X.

Ioannides:1986:CSP

[IA86] K. Ioannides and P. A. Assimakopoulos. CEAN — a system for processing coincident data and particle identification. Computer Physics Communications, 39(2):245–257, February/ March 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586901359.

Ixaru:1987:CMM

[IB87] L. Gr. Ixaru and S. Berceanu. Coleman’s method maximally adapted to the Schrödinger equation. Computer Physics Communications, 44(1–2):11–20, April/May 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900130.

Itoh:1984:NCI

[IFI84] Kimitaka Itoh, Atsushi Fukuyama, and Sanae-Inoue Itoh. Numerical calculation of ICRF waves in tokamak plasmas. Computer Physics Communications, 32(1):35–44, April 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900067.

Izrar:1989:IVE

[IGB^{+}89] B. Izrar, A. Ghizzo, P. Bertrand, E. Fijalkow, and M. R. Feix. Integration of Vlasov equation by a fast Fourier Eulerian code. Computer Physics Communications, 52(3):375– 382, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901124.

Iles:1986:TFP

[IH86] R. M. J. Iles and S. J. Hague. Toolpack: The first public release. Computer Physics Communications, 41(2–3):259–270, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558690069X.

Iwasaki:1988:QPC

[IHS^{+}88] Y. Iwasaki, T. Hoshino, T. Shirakawa, Y. Oyanagi, and T. Kawai. QCDPAX: a parallel computer for lattice QCD simulation. Computer Physics Communications, 49(3):449– 455, June 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588900057.

Sibert:1988:VAM

[III88] Edwin L. Sibert III. VANVLK: An algebraic manipulation program for canonical van vleck perturbation theory. Computer Physics Communications, 51(1–2):149–160, September/ October 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588900689.

Finan:1981:STD

[IK81] C. H. Finan III and J. Killeen. Solution of the time-dependent, three-dimensional resistive magnetohydrodynamic equations. Computer Physics Communications, 24(3–4):441–463, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465581901697.

Ito:1985:RPF

[IK85] Masaaki Ito and Fujio Kako. A REDUCE program for finding conserved densities of partial differential equations with uniform rank. Computer Physics Communications, 38(3):415–419, December 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585901092.

Ina:1985:LSD

[IKM85] Hiroshi Ina, Sachio Kamiya, and Jiro Mikami. Languages and software development tools for supercomputers. Computer Physics Communications, 38(2):211–219, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900876.

Ilyin:1989:HSD

[IKRT89] V. A. Ilyin, A. P. Kryukov, A. Ya. Rodioniov, and A. Yu. Taranov. High speed Dirac algebra calculations in a space of arbitrary dimension by means of a computer algebra system. Computer Physics Communications, 57(1–3):505–506, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902749.

Irwin:1984:PER

[IL84] D. J. G. Irwin and A. E. Livingston. A program for the extraction of radiative lifetimes from experimental beam-foil intensity decay data. Computer Physics Communications, 35(1–3): C–229, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824686.

Ingelman:1987:TMC

[Ing87] G. Ingelman. TWISTER — a Monte Carlo for QCD high-p ⊥ scattering. Computer Physics Communications, 46(2):217–240, August 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900038.

IP:1988:CIA

[IP88] W.-H. IP. Cometary ion acceleration processes. Computer Physics Communications, 49(1):1–7, April 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902123.

Ixaru:1980:NLS

[IR80] L. Gr. Ixaru and M. Rizea. A Numerov-like scheme for the numerical solution of the Schrödinger equation in the deep continuum spectrum of energies. Computer Physics Communications, 19(1):23–27, January/March 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900624.

Ixaru:1985:CSF

[IR85] L. Gr. Ixaru and M. Rizea. Comparison of some four-step methods for the numerical solution of the Schrödinger equation. Computer Physics Communications, 38(3):329–337, December 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901006.

Irvine:1982:TDH

[Irv82] J. M. Irvine. Time-Dependent Hartree–Fock analysis of heavy ion collisions. Computer Physics Communications, 26(3– 4):433–440, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901394.

Islam:1981:PCS

[Isl81] Md. M. Islam. A program for calculating the structure factors of liquid metals and binary liquid alloys. Computer Physics Communications, 23(1):43–50, June 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901284.

Islam:1984:PCS

[Isl84a] Md. M. Islam. A program for calculating the structure factors of liquid metals and binary liquid alloys. Computer Physics Communications, 35(1–3):C–701, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828490.

Islam:1984:RCF

[Isl84b] Md. M. Islam. A routine for calculating the form-factor and the electrical resistivity of liquid N.F.E. metals. Computer Physics Communications, 35(1–3):C–183, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824297.

Ito:1985:RPE

[Ito85] Masaaski Ito. A REDUCE program for evaluating a Lax pair form. Computer Physics Communications, 34(3):325–331, January 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585900098.

Ito:1986:RPF

[Ito86] Masaaki Ito. A Reduce program for finding symmetries of nonlinear evolution equations with uniform rank. Computer Physics Communications, 42(3):351–357, November 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900056.

Ito:1988:RPH

[Ito88] Masaaki Ito. A REDUCE program for Hirota’s bilinear operator and Wronskian operations. Computer Physics Communications, 50(3):321–330, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901889.

Isaacson:1987:PGC

[ITR^{+}87] Alan D. Isaacson, Donald G. Truhlar, Sachchida N. Rai, Rozeanne Steckler, Gene C. Hancock, Bruce C. Garrett, and Michael J. Redmon. POLYRATE: a general computer program for variational transition state theory and semiclassical tunneling calculations of chemical reaction rates. Computer Physics Communications, 47(1):91–102, October 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900695.

Ingelman:1987:LMC

[IW87] G. Ingelman and A. Weigend. LUCIFER — a Monte Carlo for high-p ⊥ photoproduction. Computer Physics Communications, 46(2):241–261, August 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558790004X.

Ixaru:1980:PNM

[Ixa80] L. Gr Ixaru. Perturbative numerical methods to solve the Schrödinger equation. Computer Physics Communications, 20(1):97–112, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580901150.

Jamieson:1980:PCE

[JA80] M. J. Jamieson and I. H. K. Aldeen. A program to calculate the eigenfunctions of the random phase approximation for two electron systems. Computer Physics Communications, 20(2):213– 219, October 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558090003X.

Jamieson:1984:PCE

[JA84] M. J. Jamieson and I. H. K. Aldeen. A program to calculate the eigenfunctions of the random phase approximation for two electron systems. Computer Physics Communications, 35(1–3): C–631, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827939.

Jabs:1984:AFA

[Jab84]  Arthur Jabs.  Approximation formula and Algol pro-

gram of the Lorentz-invariant momentum-space integral for particles of equal masses. Computer Physics Communications, 35(1–3):C–175–C–176, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824236.

Jackson:1980:CPL

[Jac80] C. Jackson. The CPC program library. Computer Physics Communications, 19(1):11–15, January/March 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900600.

Jacobs:1981:AEF

[Jac81] David A. Jacobs. Applications of ESOP, a fast microprogrammable processor, in high energy physics experiments at CERN. Computer Physics Communications, 22(2–3):261–267, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900576.

Jacobs:1982:EEF

[Jac82] D. A. Jacobs. Experience with ESOP, a fast microprogrammable trigger processor. Computer Physics Communications, 26(1–2):69–77, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901606.

Jackson:1985:FSMb

[Jac85a] C. Jackson. A Fortran system to maintain a program library, adapt retrieve for retrieval of decks containing free format data. Computer Physics Communications, 38(1):115, August/ September 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585900505.

Jackson:1985:FSMa

[Jac85b] C. Jackson. A Fortran system to maintain a program library, adapt update to store decks containing free format data. Computer Physics Communications, 38(1):113–114, August/ September 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585900499.

Jadach:1984:RGM

[Jad84] S. Jadach. Rapidity generator for Monte–Carlo calculations of cylindrical phase space. Computer Physics Communications, 35(1–3):C–310–C–311, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825370.

Jakab:1984:SPN

[Jak84] L. Jakab. A simple procedure for numerical approximation of the F_{m}(z) functions with complex argument. Computer Physics Communications, 31(1):89–95, January 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900845.

James:1980:ISL

[Jam80] F. James. Interpretation of the shape of the likelihood function around its minimum. Computer Physics Communications, 20(1):29–35, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580901034.

James:1981:P

[Jam81] Frederick James. Preface. Computer Physics Communications, 22(2–3):v, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900370.

James:1982:SPP

[Jam82] R. A. James. Simulation of particle problems in astrophysics. Computer Physics Communications, 26(3–4):423–431, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901382.

Jamieson:1984:PCR

[Jam84] M. J. Jamieson. A program to calculate the radial parts of interaction matrix elements between two hydrogenic wave functions as power series. Computer Physics Communications, 35(1–3): C–50, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584823152.

James:1986:DPN

[Jam86a] F. James. Do physicists need software engineering? Computer Physics Communications, 41(2–3):205–216, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900652.

James:1986:ECP

[Jam86b] G. D. James. Exact calculation of the penetrability through triple hump fission barriers. Computer Physics Communications, 40(2–3):375–387, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901207.

Jamieson:1987:NSR

[Jam87] M. J. Jamieson. Numerical solution of the restricted rotor eigenvalue equation in molecular collision calculations. Computer Physics Communications, 47(2–3):229–233, November/ December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901093.

Jamieson:1988:CUA

[Jam88] M. J. Jamieson. A comment on the use and accuracy of the Richardson deferred approach to the limit. Computer Physics Communications, 50(3):289–292, August 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558890183X.

Jaschek:1984:DA

[Jas84] C. Jaschek. Data in astronomy. Computer Physics Communications, 33(1–3):289–290, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901334.

Jacquemin:1980:CAM

[JB80] J. L. Jacquemin and G. Bordure. A computer-aided modelling of Cu_{2}S–CdS solar cells. Computer Physics Communications, 21(1):51–61, December 1, 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900752.

Jaeger:1986:IWP

[JBWW86] E. F. Jaeger, D. B. Batchelor, H. Weitzner, and J. H. Whealton. ICRF wave propagation and absorption in tokamak and mirror magnetic fields — a full-wave calculation. Computer Physics Communications, 40(1):33–64, May 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901475.

Jesshope:1982:PHD

[Jes82] C. R. Jesshope. Programming with a high degree of parallelism in Fortran. Computer Physics Communications, 26(3–4):237– 246, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901114.

Jesshope:1984:SSS

[Jes84] C. R. Jesshope. Sipsol — a suite of subprograms for the solution of the linear equations arising from elliptic partial differential equations. Computer Physics Communications, 35(1–3): C–566, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827411.

Jesshope:1985:ROP

[Jes85] C. R. Jesshope. The RPA — optimising a processor array architecture for implementation using VLSI. Computer Physics Communications, 37(1–3):95–100, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901407.

Jesshope:1986:CPN

[Jes86] C. R. Jesshope. Computational physics and the need for parallelism. Computer Physics Communications, 41(2–3):363–375, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900755.

Jesshope:1987:BRB

[Jes87] C. R. Jesshope. Book review: Parallel Array Processing: P. G. Ducksbury, Ellis Horwood, Chichester, 1986 (Distributed by J. Wiley). ISBN 0-7458-0019-X. £14.95. Computer Physics Communications, 43(2):313, January 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587902153.

Johnston:1987:LCS

[JHRR87] William E. Johnston, Dennis E. Hall, Fritz Renema, and David Robertson. Low cost scientific video movie making. Computer Physics Communications, 45(1–3):479–484, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901937.

Janzen:1984:PMM

[JL84] A. R. Janzen and J. W. Leech. Periodic multistep methods in molecular dynamics. Computer Physics Communications, 32(4):349–359, July 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900523.

Joubert:1982:MPC

[JM82] G. R. Joubert and A. J. Maeder. An MIMD parallel computer system. Computer Physics Communications, 26(3–4):253–257, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558290114X.

Jejcic:1989:CRE

[JMS^{+}89] Alain Jejcic, Jacques Maillard, Jorge Silva, Michel Auguin, and Fernand Boeri. Could running experience on SPMD computers contribute to the architectural choices for future dedicated computers for high energy physics simulation? Computer Physics Communications, 57(1–3):507–511, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902750.

Johns:1985:PSG

[JN85] T. C. Johns and A. H. Nelson. Particle simulation of 3D galactic hydrodynamics on the ICL DAP. Computer Physics Communications, 37(1–3):329–336, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901699.

Johnstad:1987:PSS

[JN87] H. Johnstad and J. Nicholls. Producing and supporting sharable software. Computer Physics Communications, 45(1–3):473–474, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901913.

Johnsson:1985:CRB

[Joh85] S. Lennart Johnsson. Cyclic reduction on a binary tree. Computer Physics Communications, 37(1–3):195–203, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901535.

Johnson:1987:FP

[Joh87a] A. S. Johnson. Fortran preprocessors. Computer Physics Communications, 45(1–3):275–281, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901640.

Johnson:1987:IMC

[Joh87b] H. Fisk Johnson. An improved method for computing a discrete Hankel transform. Computer Physics Communications, 43(2):181–202, January 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587902049.

Johnson:1988:SVE

[Joh88] B. R. Johnson. Semiclassical vibrational eigenvalues of H_{2}O and SO_{2} by the adiabatic switching method. Computer Physics Communications, 51(1–2):1–10, September/October 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900574.

Johnstad:1989:PFC

[Joh89] Harald Johnstad. PAW at Fermilab CORE-based graphics implementation of HIGZ. Computer Physics Communications, 57(1–3):438–442, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902580.

Jenkins:1980:LCL

[JP80] H. D. B. Jenkins and K. F. Pratt. Laten — a complete lattice energy program. Computer Physics Communications, 21 (2):257–269, December 2, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900442.

Jenkins:1984:CFD

[JP84a] H. D. B. Jenkins and K. F. Pratt. Calculation of the first derivatives of Madelung constants with respect to cell lengths. Computer Physics Communications, 35(1–3):C–434, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584826430.

Jenkins:1984:LCL

[JP84b] H. D. B. Jenkins and K. F. Pratt. LATEN — a complete lattice energy program. Computer Physics Communications, 35(1–3):C–666–C–667, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828222.

James:1984:MSF

[JR84] F. James and M. Roos. Minut — a system for function minimization and analysis of the parameter errors and correlations. Computer Physics Communications, 35(1–3):C–346–C– 347, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825680.

Jackson:1982:DNP

[JSB82] A. A. Jackson, H. M. Sykes, and R. S. Blake. Drooling — a non-parametric multidimensional clustering algorithm for distributed array processor. Computer Physics Communications, 27(4):351–364, October 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900972.

Jain:1983:PGS

[JT83] Ashok Jain and D. G. Thompson. A program to generate the symmetry-adapted rotational eigenfunctions and energy levels for asymmetric top molecules. Computer Physics Communications, 30(3):301–309, November 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900978.

Jain:1984:EGP

[JT84a] Ashok Jain and D. G. Thompson. EROTVIB, a general program to calculate rotationally and / or vibrationally elastic and inelastic cross sections for electron (positron) scattering by spherical, symmetric and asymmetric top molecules. Computer Physics Communications, 32(4):367–383, July 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900547.

Jain:1984:PGS

[JT84b] Ashok Jain and D. G. Thompson. A program to generate the symmetry-adapted rotational eigenfunctions and energy levels for asymmetric top molecules. Computer Physics Communications, 35(1–3):C–918, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584830179.

Judge:1987:NVA

[Jud87] R. H. Judge. A new version of ‘asyrot’ for the HP Vectra or any IBM AT compatible computer. Computer Physics Communications, 47(2–3):361–363, November/December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901214.

Junker:1981:CFT

[Jun81] B. R. Junker. Computation of Fourier transform of a general two-centre STO charge distribution. Computer Physics Communications, 23(4):377–384, August 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901788.

Junker:1984:CFT

[Jun84] B. R. Junker. Computation of Fourier transform of a general two-centre STO charge distribution. Computer Physics Communications, 35(1–3):C–735–C–736, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828702.

Jadach:1985:MCS

[JW85] S. Jadach and Z. Was. Monte Carlo simulation of the process e^{+}e^{−} → τ^{+}τ^{−}, τ^{±} → X^{±} including radiative O(α^{3}) QED corrections, mass and spin effects. Computer Physics Communications, 36(2):191–211, April 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901237.

Koniges:1987:IPB

[KA87] A. E. Koniges and D. V. Anderson. ILUBCG2: a preconditioned biconjugate gradient routine for the solution of linear asymmetric matrix equations arising from 9-point discretizations. Computer Physics Communications, 43(2):297– 302, January 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558790213X.

Kaczarowski:1984:CPC

[Kac84] R. Kaczarowski. A computer program for calculation of the Coriolis effect in odd-A nuclei. Computer Physics Communications, 35(1–3):C–415, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826247.

Karita:1989:NHE

[KAH^{+}89] Yukio Karita, Fumio Abe, Hitoshi Hirose, Hiroyuki Goto, Ryusuke Ogasawara, Fukuko Yuasa, Yoshiaki Banno, and Yoshiji Yasu. Networking for high energy physics in japan. Computer Physics Communications, 57(1–3):455–458, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589902622.

Kalantar:1982:DSD

[Kal82] A. H. Kalantar. Detecting systematic deviations from single exponentiality, using numerical integration and difference plots. Computer Physics Communications, 25(3):209–216, March 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900170.

Kalantar:1983:NLD

[Kal83] A. H. Kalantar. Noise limitations on the detection of a second exponential in nearly single exponential decay data. Computer Physics Communications, 28(3):315–316, January 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900486.

Kammerer:1984:GLF

[Kam84] U. Kammerer. Ginzburg–Landau-fluxoids. Computer Physics Communications, 35(1–3):C–1, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0010465584822730.

Kamelander:1986:RPC

[Kam86] G. Kamelander. Retrans — a program for calculating reactivity transients. Computer Physics Communications, 39(1):105–125, January 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586901669.

Kanzaki:1987:SSD

[Kan87] J. Kanzaki. Software structure for detectors at TRISTAN. Computer Physics Communications, 45(1–3):485, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901950.

Kara:1980:EEB

[Kar80] S. M. Kara. Energy eigenvalues and bound-bound transitions of hydrogen atoms in a magnetic field using cylindrical basis functions. Computer Physics Communications, 20(2):221–235, October 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900041.

Koskelo:1981:SMP

[KAR81] M. J. Koskelo, P. A. Aarnio, and J. T. Routti. SAMPO80: Minicomputer program for gamma spectrum analysis with nuclide identification. Computer Physics Communications, 24(1): 11–35, September/October 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581901041.

Kara:1984:EEB

[Kar84a] S. M. Kara. Energy eigenvalues and bound-bound transitions of hydrogen atoms in a magnetic field using cylindrical basis functions. Computer Physics Communications, 35(1–3):C–632, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827940.

Koskelo:1984:SMP

[KAR84b] M. J. Koskelo, P. A. Aarnio, and J. T. Routti. SAMP080: Minicomputer program for gamma spectrum analysis with nuclide identification. Computer Physics Communications, 35(1–3):C– 741, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828751.

Kashiwagi:1982:LST

[Kas82] Hiroshi Kashiwagi. Large-scale theoretical calculations in molecular science-design of a large computer system for molecular science and necessary conditions for future computers. Computer Physics Communications, 26(3–4):411–414, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901357.

Kasman:1984:MCS

[Kas84] J. Kasman. Monte Carlo simulation of the diffractive excitation model. Computer Physics Communications, 35(1–3):C– 301, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825308.

Katsonis:1984:DAA

[Kat84] Konstantinos Katsonis. Development and application of atomic data bases. Computer Physics Communications, 33(1–3):115– 128, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584901140.

Kawabata:1986:NMC

[Kaw86] S. Kawabata. A new Monte Carlo event generator for high energy physics. Computer Physics Communications, 41(1):127– 153, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900251.

Kolman:1984:FIM

[KB84] Bernard Kolman and Robert E. Beck. Freudenthal’s inner multiplicity formula. Computer Physics Communications, 35(1–3): C–205, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824479.

Kitowski:1986:AFD

[KB86a] Jacek Kitowski and Monika Bargiel. Applications of fuzzy diagnostics to failure analysis of industrial complex plants. Computer Physics Communications, 41(2–3):419–421, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900834.

Kushner:1986:ECC

[KB86b] Ed Kushner and Rick Broussard. Efficient computation of the cylindrical Bessel functions of complex argument. Computer Physics Communications, 42(3):345–349, November 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900044.

Koch:1986:GSI

[KBMV86] R. Koch, V. P. Bhatnagar, A. M. Messiaen, and D. Van Eester. A global solution of the ICRH problem based on the combined use of a planar coupling model and hot-plasma ray-tracing in tokamak geometry. Computer Physics Communications, 40(1): 1–22, May 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586901451.

Kapp:1981:IPD

[KC81a] B. C. Kapp and J. R. Comfort. Implementation of the partitioned data set concept for DEC-10 computers. Computer Physics Communications, 23(4):365–375, August 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901776.

Karp:1981:DIW

[KC81b] B. C. Karp and J. R. Comfort. Design and implementation of a word addressable data set for DEC-10 computers. Computer Physics Communications, 23(4):355–364, August 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901764.

Karp:1984:DIW

[KC84a] B. C. Karp and J. R. Comfort. Design and implementation of a word addressable data set for DEC-10 computers. Computer Physics Communications, 35(1–3):C–733, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828684.

Karp:1984:IPD

[KC84b] B. C. Karp and J. R. Comfort. Implementation of the partitioned data set concept for DEC-10 computers. Computer Physics Communications, 35(1–3):C–734, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828696.

Kuo:1984:SRI

[KCA84] C. T. K. Kuo, T. W. Cadman, and R. J. Arsenault. Sequential random integer generator. Computer Physics Communications, 35(1–3):C–394, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826053.

Kirkegaard:1984:PVP

[KE84a] Peter Kirkegaard and Morten Eldrup. Positronfit: a versatile program for analysing positron lifetime spectra. Computer Physics Communications, 35(1–3):C–138, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823929.

Kirkegaard:1984:PEN

[KE84b] Peter Kirkegaard and Morten Eldrup. Positronfit extended: a new version of a program for analysing positron lifetime spectra. Computer Physics Communications, 35(1–3):C–255, ????

1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824911.

Kegel:1981:TFS

[Keg81] Gunter H. R. Kegel. A time-of-flight spectrum simulator for neutron elastic and inelastic scattering. Computer Physics Communications, 24(2):205–223, November 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900953.

Kegel:1984:TFS

[Keg84] Gunter H. R. Kegel. A time-of-flight spectrum simulator for neutron elastic and inelastic scattering. Computer Physics Communications, 35(1–3):C–761, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482891X.

Kegel:1985:FES

[Keg85] Gunter H. R. Kegel. Fluences and energy spectra of fast neutrons from a proton-irradiated thick lithium target. Computer Physics Communications, 36(3):321–336, May 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900591.

Kellner:1987:DSA

[Kel87] G. Kellner. Development of software for ALEPH using structured techniques. Computer Physics Communications, 45 (1–3):229–243, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587901603.

Kirkegaard:1981:PSA

[KEMP81] Peter Kirkegaard, Morten Eldrup, Ole E. Mogensen, and Niels J. Pedersen. Program system for analysing positron lifetime spectra and angular correlation curves. Computer Physics Communications, 23(3):307–335, July 2, 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900060.

Kirkegaard:1984:PSA

[KEMP84] Peter Kirkegaard, Morten Eldrup, Ole E. Mogensen, and Niels J. Pedersen. Program system for analysing positron lifetime spectra and angular correlation curves. Computer Physics Communications, 35(1–3):C–732, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828672.

Kennedy:1989:SLG

[Ken89a] A. D. Kennedy. Status of lattice gauge theory calculations. Computer Physics Communications, 57(1–3):57–67, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901926.

Kennedy:1989:RDR

[Ken89b] A. S. Kennedy. Recent developments in real time structured analysis. Computer Physics Communications, 57(1–3):561, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589902920.

Keyes:1989:DDM

[Key89]  David E. Keyes.  Domain decomposition methods for

the parallel computation of reacting flows. Computer Physics Communications, 53(1–3):181–200, May 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901598.

Kankowsky:1984:PCC

[KF84] R. Kankowsky and D. Fick. A program to calculate complex phase shifts and mixing parameters of elastic scattering of spin 1/2 particles on spin 1/2 targets. Computer Physics Communications, 35(1–3):C–86–C–87, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823486.

Koskinen:1988:VOW

[KGAL88] Hannu E. J. Koskinen, Georg Gustafsson, Mats André, and Rickard Lundin. Viking observations of wave-particle interactions and ion wave instabilities in the high-latitude magnetosphere. Computer Physics Communications, 49(1):75–83, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902160.

Knowles:1989:DBF

[KH89] Peter J. Knowles and Nicholas C. Handy. A determinant based full configuration interaction program. Computer Physics Communications, 54(1):75–83, April 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900337.

Killeen:1980:RIC

[Kil80] J. Killeen. Report of the IAEA Consultants’ Meeting on Computer Codes for Fusion Research 5–7 December 1979, Vienna, Austria. Computer Physics Communications, 19(3):273–291, July/August 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558090082X.

Kirby:1987:SCD

[Kir87] P. Kirby. A simulation code for 3-dimensional non-linear incompressible resistive MHD in a periodic cylinder. Computer Physics Communications, 47(1):17–43, October 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900646.

Kerner:1984:AME

[KJ84] W. Kerner and O. Jandl. Axisymmetric MHD equilibria with flow. Computer Physics Communications, 31(2–3):269–285, February 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558490050X.

Kightley:1985:CCG

[KJ85] J. R. Kightley and I. P. Jones. A comparison of conjugate gradient preconditionings for three-dimensional problems on a CRAY-1. Computer Physics Communications, 37(1–3):205– 214, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585901547.

Kessel:1988:RO

[KJRS88] R. L. Kessel, A. D. Johnstone, D. J. Rodgers, and M. F. Smith. Reconnection observations. Computer Physics Communications, 49(1):161–172, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588902238.

Keren:1983:GAM

[KK83a] E. Keren and O. Kafri. A geometric approach for mapping of derivatives. Computer Physics Communications, 29(2):109– 112, April 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900656.

Kosloff:1983:NPF

[KK83b] D. Kosloff and R. Kosloff. A non-periodic Fourier method for solution of the classical wave equation. Computer Physics Communications, 30(3):333–336, November 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583901005.

Kajantie:1984:EVP

[KK84a] K. Kajantie and V. Karimäki. The evaluation of the volume of the phase space of N particles. Computer Physics Communications, 35(1–3):C–84, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823462.

Kunasz:1984:TLR

[KK84b] C. V. Kunasz and P. B. Kunasz. Transfer of line radiation in optically thick media allowing for transport of excitation energy: The resonant doublet. Computer Physics Communications, 35(1–3):C–345, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825679.

Kitowski:1985:FLA

[KK85] J. Kitowski and E. Ksiazek. Fuzzy logic applications for failure analysis and diagnosis of a primary circuit of the HTR nuclear power plant. Computer Physics Communications, 38(2):323–327, October/November 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900980.

Kawabata:1988:MDI

[KK88] S. Kawabata and T. Kaneko. A multi-dimensional integration package for a vector processor. Computer Physics Communications, 48(3):353–365, March 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588902019.

Kaneko:1989:PFS

[KK89] Toshiaki Kaneko and Setsuya Kawabata. A preprocessor for Fortran source code produced by REDUCE. Computer Physics Communications, 55(2):141–147, September 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900738.

King:1984:RLE

[KKLP84] Frederick W. King, Michael K. Kelly, Mary A. LeGore, and Martin E. Poitzsch. Reduced local energy for atomic Hartree–Fock wavefunctions. Computer Physics Communications, 32(2):215–224, May 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900730.

Kirillov:1982:SLC

[KKLS82] V. I. Kirillov, E. L. Kosarev, L. B. Luganskii, and A. M. Schastlivtzev. Small laboratory computer automation system based on Hewlett–Packard HP-1000 mini computer and IEEE-488 standard. Computer Physics Communications, 26 (1–2):201–204, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901692.

Katsuki:1982:CPS

[KKP82] Shinichi Katsuki, Mariusz Klobukowski, and Pancracio Palting. A compact program of the SCF–Xα scattered wave method: Version II. Computer Physics Communications, 25(1):39–55, January 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558290042X.

Katsuki:1984:CPSb

[KKP84] Shinichi Katsuki, Mariusz Klobukowski, and Pancracio Palting. A compact program of the SCF-Xα scattered wave method: Version II. Computer Physics Communications, 35 (1–3):C–764–C–765, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828945.

Kaneko:1987:AGF

[KKS87] T. Kaneko, S. Kawabata, and Y. Shimizu. Automatic generation of Feynman graphs and amplitudes in QED. Computer Physics Communications, 43(2):279–295, January 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587902128.

Kuroda:1988:BSH

[KKT^{+}88] S. Kuroda, T. Kamitani, K. Tobimatsu, S. Kawabata, and Y. Shimizu. Bhabha scattering at high energy. Computer Physics Communications, 48(3):335–351, March 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902007.

Kluge:1984:NFG

[KL84a] G. Kluge and K. G. Lenhart. Numerical fits for the geomagnetic shell parameter. Computer Physics Communications, 35(1–3): C–120, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584823784.

Komle:1984:SFD

[KL84b] Norbert I. Kömle and Herbert I. M. Lichtenegger. Shock formation and decay in hydrodynamic flows with mass loading.

Computer Physics Communications, 34(1–2):47–55, November/ December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901589.

Klapisch:1984:PAW

[Kla84] M. Klapisch. A program for atomic wavefunction computations by the parametric potential method. Computer Physics Communications, 35(1–3):C–89, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823504.

Kerner:1985:NMA

[KLGT85] W. Kerner, K. Lerbinger, R. Gruber, and T. Tsunematsu. Normal mode analysis for resistive cylindrical plasmas. Computer Physics Communications, 36(3):225–240, May 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900530.

Klobukowski:1982:AAC

[Klo82] M. Klobukowski. An adaptation of ACQI to calculate the data for MSXALPHA program. Computer Physics Communications, 25(1):29–38, January 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900418.

Klobukowski:1984:AAC

[Klo84a] M. Klobukowski. An adaptation of acqi to calculate the data for MS-Xα program. Computer Physics Communications, 35(1–3):C–763, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828933.

Klotz:1984:GPC

[Klo84b] W.-D. Klotz. A general program to calculate the matrix of the spin-orbit interaction. Computer Physics Communications, 35(1–3):C–295, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825254.

Klotz:1984:RME

[Klo84c] Wolf-Dieter Klotz. Reduced matrix elements of summations of one-particle tensor products. Computer Physics Communications, 35(1–3):C–327, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825503.

Kerner:1985:CCE

[KLS85] W. Kerner, K. Lerbinger, and J. Steuerwald. Computing complex eigenvalues of large non-Hermitian matrices. Computer Physics Communications, 38(1):27–37, August/ September 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585900426.

Kronfeld:1989:VCC

[KLS^{+}89] A. S. Kronfeld, M. L. Laursen, G. Schierholz, C. Schleiermacher, and U.-J. Wiese. A vectorized code for the computation of the topological charge in SU(2) lattice gauge theory. Computer Physics Communications, 54(1):109–124, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900374.

Kluge:1984:DCM

[Klu84a] G. Kluge. Direct computation of the magnetic shell parameter. Computer Physics Communications, 35(1–3):C–119, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823772.

Kluge:1984:GFM

[Klu84b] G. Kluge. Geomagnetic field models: Scalar and vector potential, induction vector and its gradient tensor computed by a common algorithm. Computer Physics Communications, 35(1–3):C–161, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824121.

Krivec:1982:PPC

[KM82] R. Krivec and M. V. Mihailović. Program package for calculating matrix elements of two-cluster structures in nuclei.

Computer Physics Communications, 28(2):153–182, December 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900522.

Kerr:1984:PCW

[KM84a] I. F. Kerr and C. H. B. Mee. Program for calculating work functions from photoelectric data. Computer Physics Communications, 35(1–3):C–257, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824935.

Krivec:1984:PPC

[KM84b] R. Krivec and M. V. Mihailović. Program package for calculating matrix elements of two-cluster structures in nuclei. Computer Physics Communications, 35(1–3):C–841–C–842, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829513.

Kruger:1984:CNG

[KM84c] P. B. Kruger and R. M. Mayer. Calculation of the nucleation and growth of defect clusters. Computer Physics Communications, 35(1–3):C–602, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482771X.

Kuba:1985:EML

[KM85] D. W. Kuba and K. J. M. Moriarty. Efficient multitasking of the Su(3) lattice gauge theory algorithm on the Cray X-MP. Computer Physics Communications, 36(4):351–362, June 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900244.

Kerbel:1986:CBR

[KM86] G. D. Kerbel and M. G. McCoy. Collision broadened resonance localization in tokamaks excited with ICRF waves. Computer Physics Communications, 40(1):105–113, May 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901529.

Kronfeld:1988:EMC

[KMS88] A. S. Kronfeld, K. J. M. Moriarty, and G. Schierholz. An efficient method for the computation of glueball masses using the inverse of the covariant Dirac operator as correlator. Computer Physics Communications, 52(1):1–6, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901646.

Klucker:1984:KKA

[KN84a] R. Klucker and U. Nielsen. Kramers–Kronig analysis of reflection data. Computer Physics Communications, 35(1–3):C–218– C–219, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824595.

Kunc:1984:LDZb

[KN84b] K. Kunc and O. H. Nielsen. Lattice dynamics of zincblende structure compounds II. Shell model. Computer Physics Communications, 35(1–3):C–570, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0010465584827447.

Kunc:1984:LDZa

[KN84c] K. Kunc and O. Holm Nielsen. Lattice dynamics of zincblende structure compounds using deformation-dipole model and rigid ion model. Computer Physics Communications, 35(1–3):C–529–C–530, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827101.

Katznelson:1986:ISA

[KN86] E. Katznelson and A. Nobile. Implementation and statistical analysis of Metropolis algorithm for SU(3). Computer Physics Communications, 39(1):1–19, January 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901591.

Knox:1984:CWH

[Kno84] H. O. Knox. Computation of wavefunctions for the helium isoelectronic sequence. Computer Physics Communications, 35(1–3):C–19, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584822882.

Koberle:1989:NNC

[Köb89] Roland Köberle. Neural networks as content addressable memories and learning machines. Computer Physics Communications, 56(1):43–50, November 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900519.

Kolar:1981:PME

[Kol81a] M. Kolár. Pseudopotential matrix elements in the Gaussian basis. Computer Physics Communications, 23(3):275–286, July 2, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900035.

Kolbig:1981:PCC

[Köl81b] K. S. Kölbig. A program for computing the conical functions of the first kind P_{−1/2+iτ}^{m} (x) for m = 0 and m = 1. Computer Physics Communications, 23(1):51–61, June 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901296.

Kolar:1984:PME

[Kol84a] M. Kolár. Pseudopotential matrix elements in the Gaussian basis. Computer Physics Communications, 35(1–3):C–728–C– 729, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828647.

Kolbig:1984:PCC

[Köl84b] K. S. Kölbig. A program for computing the conical functions of the first kind P_{−1/2+iτ}^{m} (x) for m = 0 and m = 1. Computer Physics Communications, 35(1–3):C–702, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828507.

Kolbig:1984:PCL

[Köl84c] K. S. Kölbig. Programs for computing the logarithm of the gamma function, and the digamma function, for complex argument. Computer Physics Communications, 35(1–3):C–152, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824042.

Konrad:1984:LAC

[Kon84] A. Konrad. A linear accelerator cavity code based on the finite element method. Computer Physics Communications, 35 (1–3):C–435–C–436, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826442.

Kosarev:1980:AIE

[Kos80] E. L. Kosarev. Applications of integral equations of the first kind in experiment physics. Computer Physics Communications, 20(1):69–75, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580901101.

Kascic:1981:VPH

[KP81a] Michael J. Kascic and Charles J. Purcell. Vector processing for high energy physics. Computer Physics Communications, 22(2–3):293–295, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558190062X.

Krasinski:1981:ONC

[KP81b] Andrzej Krasiński and Marek Perkowski. ORTOCARTAN — a new computer program for algebraic calculations. Computer Physics Communications, 22(2–3):269–271, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900588.

Kuo-Petravic:1984:NEM

[KP84] Gioietta Kuo-Petravic. Numerical evaluation of magnetic coordinates for toroidal magnetic confinement devices. Computer Physics Communications, 33(4):353–359, October 1984.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901413.

Kruppa:1985:RBS

[KP85] A. T. Kruppa and Z. Papp. Resonant or bound state solution of the Schrödinger equation in deformed or spherical potential. Computer Physics Communications, 36(1):59–78, March 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900189.

Kelkar:1989:DGB

[KP89] Kanchan M. Kelkar and Suhas V. Patankar. Development of generalized block correction procedures for the solution of discretized Navier–Stokes equations. Computer Physics Communications, 53(1–3):329–336, May 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901707.

Katsuki:1984:CPSa

[KPH84] S. Katsuki, P. Palting, and S. Huzinaga. A compact program of the SCF-Xα scattered wave method. Computer Physics Communications, 35(1–3):C–446–C–447, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826533.

Kuo-Petravic:1981:PGI

[KPP81] G. Kuo-Petravic and M. Petravic. A program generator for the Incomplete Cholesky Conjugate Gradient (ICCG) method with a symmetrizing preprocessor. Computer Physics Communications, 22(1):33–48, February/March 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900771.

Kuo-Petravic:1984:PGIa

[KPP84a] G. Kuo-Petravic and M. Petravic. A program generator for the incomplete LU decomposition-conjugate gradient (IL ucg) method. Computer Physics Communications, 35(1–3):C–572– C–573, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827472.

Kuo-Petravic:1984:PGIb

[KPP84b] G. Kuo-Petravic and M. Petravic. A program generator for the incomplete Cholesky conjugate gradient (ICCG) method with a symmetrizing preprocessor. Computer Physics Communications, 35(1–3):C–682–C–683, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828349.

Kress:1989:CLS

[KPP^{+}89] J. D. Kress, G. A. Parker, R. T. Pack, B. J. Archer, and W. A. Cook. Comparison of Lanczos and subspace iterations for hyperspherical reaction path calculations. Computer Physics Communications, 53(1–3):91–108, May 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901501.

Kuo-Petravic:1984:TPT

[KPPR84] L. G. Kuo-Petravic, M. Petravic, and K. V. Roberts. Tranal — a program for the translation of symbolic Algol I into symbolic Algol II. Computer Physics Communications, 35(1–3):C–356, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482576X.

Kowalski:1989:IAP

[KPST89] H. Kowalski, T. Poser, L. Stanco, and E. Tscheslog. Investigation of ADAMO performance in the ZEUS calorimeter reconstruction program. Computer Physics Communications, 57 (1–3):222–224, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589902166.

Kirby:1981:CGS

[KR81] G. H. Kirby and A. Rixon. Computer graphics and stereoscopy for three-dimensional data. Computer Physics Communications, 21(3):287–291, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900084.

Kryukov:1988:PCC

[KR88] A. P. Kryukov and A. Ya. Rodionov. Program “color” for computing the group-theoretic weight of Feynman diagrams in non-Abelian gauge theories. Computer Physics Communications, 48(2):327–334, February 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588900525.

Krajci:1986:MCA

[Kra86a] M. Krajcí. Markov chain algorithms for canonical ensemble simulation. Computer Physics Communications, 42(1):29–35, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586902274.

Kral:1986:ELS

[Krá86b] J. Král. Empirical laws of software development and their implications. Computer Physics Communications, 41(2–3):385– 391, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900779.

Krischer:1989:CHP

[Kri89] Werner Krischer. Commercial highly parallel signal processors on-line? Computer Physics Communications, 57(1–3):121–128, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589902002.

Kosachevskaya:1983:SBM

[KRS83] L. L. Kosachevskaya, V. V. Romanovtsev, and I. E. Shparlinskiy. On the spline-based method for experimental data deconvolution. Computer Physics Communications, 29(3):227– 230, May 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900024.

Kosachevskaya:1982:NIA

[KRSV82] L. L. Kosachevskaya, V. V. Romanovtsev, I. E. Shparlinskiy, and A. N. Vystavkin. New improved algorithm for the iterative solution of a system of linear algebraic equations. Computer Physics Communications, 27(1):87–89, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558290011X.

Kent:1981:MPC

[KS81] R. D. Kent and M. Schlesinger. A microcomputer program for the correlating of two ordered lists of numbers. Computer Physics Communications, 23(3):301–305, July 2, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900059.

Kent:1984:GIC

[KS84a] R. D. Kent and M. Schlesinger. Generation and intercorrelation of basis sets in implementing the unitary group approach to U(n) × SU(m). Computer Physics Communications, 33(4):367–372, October 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584901437.

Kent:1984:MPC

[KS84b] R. D. Kent and M. Schlesinger. A microcomputer program for the correlating of two ordered lists of numbers. Computer Physics Communications, 35(1–3):C–731, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828660.

Kisak:1984:FEP

[KS84c] E. Kisak and P. Silvester. A finite-element program package for magnetotelluric modelling. Computer Physics Communications, 35(1–3):C–353–C–354, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825734.

Kolbig:1984:AEL

[KS84d] K. S. Kölbig and B. Schorr. Asymptotic expansions for the Landau density and distribution functions. Computer Physics Communications, 32(2):121–131, May 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900651.

Kolbig:1984:PPL

[KS84e] K. S. Kölbig and B. Schorr. A program package for the Landau distribution. Computer Physics Communications, 31(1):97– 111, January 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900857. See erratum

[KS08].

Kolbig:1984:PSS

[KS84f] K. S. Kölbig and F. Schwarz. A program for solving systems of homogeneous linear inequalities. Computer Physics Communications, 35(1–3):C–565, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482740X.

Konrad:1984:FEPa

[KS84g] A. Konrad and P. Silvester. A finite element program package for axisymmetric scalar field problems. Computer Physics Communications, 35(1–3):C–198–C–200, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824431.

Konrad:1984:FEPb

[KS84h] A. Konrad and P. Silvester. A finite element program package for axisymmetric vector field problems. Computer Physics Communications, 35(1–3):C–302–C–303, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482531X.

Kossionides:1986:RFS

[KS86a] S. Kossionides and L. D. Skouras. RWSYST — a filing system for coefficients and eigenvectors. Computer Physics Communications, 39(2):213–219, February/March 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901323.

Kuchiev:1986:CPC

[KS86b]  M. Yu. Kuchiev and S. A. Sheinerman.  The calcu-

lation of the photoionization cross section allowing for the post collision interaction. Computer Physics Communications, 39(2):155–160, February/March 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901268.

Kent:1987:DST

[KS87a] R. D. Kent and M. Schlesinger. Data structure techniques for the graphical special unitary group approach to arbitrary spin representations. Computer Physics Communications, 43(3):413–437, February/March 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900580.

Kunz:1987:P

[KS87b] Paul Kunz and Terry Schalk. Preface. Computer Physics Communications, 45(1–3):xiii, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465587901342.

Kamgnia:1989:FES

[KS89] E. Kamgnia and A. Sameh. A fast elliptic solver for simply connected domains. Computer Physics Communications, 55(1):43– 69, August 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900635.

Kolbig:2008:EPP

[KS08] K. S. Kölbig and B. Schorr. Erratum to: “A program package for the Landau distribution” [Comput. Phys. Commun. 31 (1984) 97–111]. Computer Physics Communications, 178(12): 972, June 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508001136. See [KS84e].

Kleiss:1986:NMC

[KSE86] R. Kleiss, W. J. Stirling, and S. D. Ellis. A new Monte Carlo treatment of multiparticle phase space at high energies. Computer Physics Communications, 40(2–3):359–373, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901190.

Krzeczkowski:1982:SMU

[KSG82] A. J. Krzeczkowski, E. A. Smith, and T. Gethin. Seismic migration using the ICL distributed array processor. Computer Physics Communications, 26(3–4):447–453, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901412.

Kolb:1986:PE

[KSS86] Dieter Kolb, Manfred Sommer, and Manfred Stadel. Programming environments. Computer Physics Communications, 41(2–3):227–244, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900676.

Kerner:1981:TMS

[KT81] W. Kerner and H. Tasso. Tearing mode stability in 1D and 2D. Computer Physics Communications, 24(3–4):407–411, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901648.

Kunz:1989:SMI

[Kun89] Paul F. Kunz. Software management issues. Computer Physics Communications, 57(1–3):191–197, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902117.

Kittel:1984:MCG

[KVW84] W. Kittel, L. Van Hove, and W. Wojcik. A Monte Carlo generation method with importance sampling for high energy collisions of hadrons. Computer Physics Communications, 35(1–3):C–48–C–49, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823140.

Kalle:1984:PRN

[KW84a] Claus Kalle and Stephan Wansleben. Problems with the random number generator RANF implemented on the CDC Cyber 205. Computer Physics Communications, 33(4):343–346, October 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901395.

Kerner:1984:ANI

[KW84b] W. Kerner and H. Weitzner. Axisymmetric, non-ideal MHD states with steady flow. Computer Physics Communications, 31(2–3):249–267, February 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900493.

Kress:1989:SSE

[KWPP89] J. D. Kress, S. B. Woodruff, G. A. Parker, and R. T. Pack. Some strategies for enhancing the performance of the block Lanczos method. Computer Physics Communications, 53(1–3):109– 115, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901513.

Laaksonen:1986:CSP

[Laa86] Aatto Laaksonen. Computer simulation package for liquids and solids with polar interactions: I. McMOLDYNH/2O: Aqueous systems. Computer Physics Communications, 42(2):271–300, October 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900433.

Labarthe:1984:GDN

[Lab84a] J. J. Labarthe. I. Generator of determinantal non-relativistic atomic states from spectroscopic notation, computation of matrix elements. Computer Physics Communications, 35(1–3):C– 540, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827174.

Labarthe:1984:IGA

[Lab84b] J. J. Labarthe. II. generator of atomic excited terms from angular considerations. Computer Physics Communications, 35(1–3):C–541, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827186.

Labarthe:1984:IAA

[Lab84c] J. J. Labarthe. III. Analytic approximations of radial orbitals for multiconfigurational Hartree–Fock computations. Computer Physics Communications, 35(1–3):C–542, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827198.

Labarthe:1984:IAN

[Lab84d] J. J. Labarthe. IV. approximation of numerical orbitals by Slater functions. Computer Physics Communications, 35(1–3): C–543, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827204.

Lado:1986:DPC

[Lad86] F. Lado. Dumbell — a program to calculate the structure and thermodynamics of a classical fluid of hard, homonuclear diatomic molecules. Computer Physics Communications, 39(1):133–140, January 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901682.

Lannes:1980:LCQ

[Lan80] A. Lannes. On the linear convergence of quasi-Newton methods in finite-and infinite-dimensional Hilbert spaces. Computer Physics Communications, 20(1):91–95, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580901149.

Landau:1982:LPK

[Lan82] Rubin H. Landau. LPOTT — PION and KAON elastic scattering from spin 12 nuclei in momentum space. Computer Physics Communications, 28(2):109–151, December 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900510.

Landau:1984:LPK

[Lan84a] Rubin H. Landau. Lpott — pion and kaon elastic scattering from spin 1/2 nuclei in momentum space. Computer Physics Communications, 35(1–3):C–839–C–840, ????

1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829501.

Lang:1984:SIA

[Lan84b] James Lang. Spectral intensity, angular distribution and polarisation of synchrotron radiation from a monoenergetic electron. Computer Physics Communications, 35(1–3):C–51, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823164.

Langeveld:1987:OSP

[Lan87] W. G. J. Langeveld. Optical storage and its possible use in high-energy physics. Computer Physics Communications, 45 (1–3):395–402, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587901809.

Lao:1984:VMM

[Lao84]  L. L. Lao.  Variational moment method for comput-

ing magnetohydrodynamic equilibria. Computer Physics Communications, 31(2–3):201–212, February 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900456.

Liddell:1982:DSL

[LB82a] Heather M. Liddell and G. S. J. Bowgen. The DAP subroutine library. Computer Physics Communications, 26(3–4):311– 315, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901217.

Loveluck:1982:CSC

[LB82b] J. M. Loveluck and E. Balcar. Computer simulation calculations for one-dimensional magnetic systems, with some remarks on optimisation on the Cray computer. Computer Physics Communications, 27(4):335–350, October 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900960.

Leutz:1984:CAC

[LB84] R. K. Leutz and R. Bauer. Computation of the anisotropic cubic elastic Green’s tensor function and the elastic energy coefficients of point defects in crystals. Computer Physics Communications, 35(1–3):C–384, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825977.

Lowe:1988:DAS

[LBK88] Mary Lowe, Stephan Blumenroeder, and Peter H. Kutt. A data acquisition system for spectroscopy using an IBM PC. Computer Physics Communications, 50(3):367–373, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901919.

Luthe:1986:ACC

[LBY86] J. C. Luthe, E. J. Beiting, and F. Y. Yueh. Algorithms for calculating coherent anti-stokes Raman spectra: Application to several small molecules. Computer Physics Communications, 42(1):73–92, September 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558690233X.

Lynch:1981:RMS

[LCH^{+}81] V. E. Lynch, B. A. Carreras, H. R. Hicks, J. A. Holmes, and L. Garcia. Resistive MHD studies of high β tokamak plasmas. Computer Physics Communications, 24(3–4):465–476, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901703.

Lemmel:1984:NDF

[LCS84] H. D. Lemmel, D. E. Cullen, and J. J. Schmidt. Nuclear data files for reactor calculations and other applications· experimental data — evaluated data. Computer Physics Communications, 33(1–3):161–171, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558490119X.

Liberman:1984:RSC

[LCW84] D. A. Liberman, D. T. Cromer, and J. T. Waber. Relativistic self-consistent field program for atoms and ions. Computer Physics Communications, 35(1–3):C–70, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823358.

LeQueau:1988:MP

[Le 88] D. Le Queau. Maser processes. Computer Physics Communications, 49(1):85–95, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588902172.

LeRoy:1989:BCI

[Le 89] Robert J. Le Roy. Bound → continuum intensities — a computer program for calculating absorption coefficients, emission intensities or (golden rule) predissociation rates. Computer Physics Communications, 52(3):383–395, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901136.

Leander:1987:SNQ

[Lea87] G. A. Leander. Simulation of nuclear quasicontinuum gammaray spectra. Computer Physics Communications, 47(2–3):311– 340, November/December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587901172.

Lee:1982:MFV

[Lee82] D. K. Lee. Magnetic field, vector potential and their derivatives due to currents in closed polygons of wire. Computer Physics Communications, 25(2):181–191, February 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900340.

Lee:1984:MFV

[Lee84] D. K. Lee. Magnetic field, vector potential and their derivatives due to currents in closed polygons of wire. Computer Physics Communications, 35(1–3):C–780, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829070.

Lee:1989:FMS

[Lee89] M. Howard Lee. Frequency moment sum rules, recurrence relations and continued fractions in nonequilibrium statistical mechanics. Computer Physics Communications, 53(1–3):147– 155, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901562.

Loveless:1989:ZHC

[LEF^{+}89] R. Loveless, P. Erhard, J. Ficenec, K. Gather, G. Heath, M. Iacovacci, J. Kehres, M. Mobayyen, D. Notz, R. Orr, R. Orr, A. Sephton, R. Stroili, K. Tokushuku, W. Vogel, J. Whitmore, and L. Wiggers. ZEUS hardware control system. Computer Physics Communications, 57(1–3):313–315, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902348.

Lencova:1980:NCE

[Len80] Bohumila Lencová. Numerical computation of electron lenses by the finite element method. Computer Physics Communications, 20(1):127–132, September 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580901186.

Lessner:1989:WFP

[Les89] Eliane S. Lessner. Weighted fit of parametric functions to distributions: The new interface of HBOOK with MINUIT. Computer Physics Communications, 57(1–3):385–389, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902488.

Levine:1989:YWM

[Lev89] Randolph H. Levine. You want me to predict the future? Computer Physics Communications, 57(1–3):118–120, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901999.

Lewellen:1982:TDA

[Lew82] Philip C. Lewellen. Two-dimensional adaptive quadrature over rectangular regions. Computer Physics Communications, 27(2):167–178, August 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900716.

Lewellen:1984:TDA

[Lew84a] Philip C. Lewellen. Two-dimensional adaptive quadrature over rectangular regions. Computer Physics Communications, 35(1–3):C–814, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829318.

Lewis:1984:HI

[Lew84b] M. N. Lewis. Hydrogenic R ^{k} integrals. Computer Physics Communications, 35(1–3):C–38, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482305X.

Lewis:1984:SPS

[Lew84c] M. N. Lewis. Single-particle substitution sums in the secondorder Z-expansion theory of atomic energies. Computer Physics Communications, 35(1–3):C–31, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822997.

Lewis:1984:ZEM

[Lew84d] M. N. Lewis. Z-expansion of matrix elements of oneelectron operators for many-electron atoms. Computer Physics Communications, 35(1–3):C–581, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482754X.

Levinthal:1987:EHS

[LGG^{+}87] D. Levinthal, H. Goldman, C. Georgiopoulos, J. L. DeKeyser, S. Linn, S. Youssef, and M. F. Hodous. Experimental HEP supercomputing at F.S.U. Computer Physics Communications, 45(1–3):137–146, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901482.

Linde:1988:DIM

[LH88] Scott R. Vande Linde and William L. Hase. Dynamics of ion-molecule recombination IV. Li^{+} + (CH_{3} )_{2}O association. Computer Physics Communications, 51(1–2):17–34, September/October 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588900598.

Ljepojevic:1987:CPC

[LHP87] N. N. Ljepojevic, R. J. Hutcheon, and J. Payne. COLRAD-A program to calculate population densities of the excited atomic levels of hydrogen-like ions in a plasma. Computer Physics Communications, 44(1–2):157–176, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900257.

Lide:1984:CRI

[Lid84a] David R. Lide, Jr. The CODATA role in international and interdisciplinary cooperation. Computer Physics Communications, 33(1–3):205–206, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558490122X.

Lide:1984:NSR

[Lid84b] David R. Lide, Jr. The national standard reference data system of the united states. Computer Physics Communications, 33(1–3):207–210, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901231.

Liska:1984:PSA

[Lis84] R. Liska. Program for stability and accuracy analysis of finite difference methods. Computer Physics Communications, 34(1–2):175–186, November/December 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901693.

Lister:1988:BRB

[Lis88] G. G. Lister. Book review: Computational Techniques in Physics: P. K. Mackeown and D. J. Newman, Adam Hilger, 1987. 217 pages. £30.00 (hardcover), £12.95 (paperback). Computer Physics Communications, 50(3):414, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901956.

Lejeune:1984:CBM

[LJ84] A. Lejeune and J. P. Jeukenne. Computation of Brody– Moshinsky brackets. Computer Physics Communications, 35 (1–3):C–88, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823498.

Lea:1984:DCS

[LJLC84] Suzanne M. Lea, Vinaya Joshi, and A. B. Lopez-Cepero. Differential cross sections for electric quadrupole Coulomb excitation I. Computer Physics Communications, 35(1–3):C–129–C– 130, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046558482384X.

Long:1985:ZCM

[LJN85] J. W. Long, J. W. Johnston, and A. A. Newton. ZEROD — a computer model for plasma-circuit coupling. Computer Physics Communications, 34(3):231–250, January 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900025.

Lee:1988:PMT

[LK88] Wen Ho Lee and Dochan Kwak. PIC method for a twodimensional elastic-plastic-hydro code. Computer Physics Communications, 48(1):11–16, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465588900185.

Lebrun:1989:HLL

[LK89] P. Lebrun and A. Kreymer. High level language memory management on parallel architectures. Computer Physics Communications, 57(1–3):231–234, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990218X.

Lubkin:1984:IQM

[LL84] Elihu Lubkin and Thelma Lubkin. An inversion of quantum mechanics. Computer Physics Communications, 35(1–3):C– 533, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827125.

Lan:1989:CCC

[LLA^{+}89] Vo Ky Lan, M. Le Dourneuf, N. F. Allard, H. E. Saraph, and W. Eissner. On the comparison of close-coupling calculations using the UCL and the opacity codes. Computer Physics Communications, 55(3):303–310, October 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901288.

Lotrian:1984:CER

[LLC84] J. Lotrian, M. Leriche, and J. Cariou. Calculation of the energy response of a spectrometer. Computer Physics Communications, 35(1–3):C–401, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826119.

Langbein:1981:IMC

[LM81] W. Langbein and A. Mattheus. Inversion of Monte Carlo calculations and constrained minimalisation. Computer Physics Communications, 21(3):279–286, January 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900072.

Lang:1984:PNL

[LM84] J. Lang and R. Müller. A procedure for nonlinear least squares refinement in adverse practical conditions. Computer Physics Communications, 35(1–3):C–67, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823322.

Leeb:1985:GFC

[LM85] H. Leeb and H. Markum. A Green’s function code for Schrödinger equations with nonlocal separable kernels. Computer Physics Communications, 34(3):271–286, January 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900049.

Leitner:1988:PSM

[LNB^{+}88] David M. Leitner, Grigory A. Natanson, R. Stephen Berry, Pablo Villarreal, and Gerardo Delgado-Barrio. Particles-on-a-sphere method for computing the rotational-vibrational spectrum of H_{2}O. Computer Physics Communications, 51(1–2): 207–216, September/October 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588900720.

Long:1984:GOD

[LNS84] J. W. Long, A. A. Newton, and M. C. Sexton. Glowcode: a onedimensional code for the simulation of plasma afterglows. Computer Physics Communications, 35(1–3):C–399–C–400, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826107.

Linkens:1984:SUD

[LNV84] A. Linkens, J. Niezette, and J. Vanderschueren. Simulation of ultrasonic degradation of macromolecules in solution. Computer Physics Communications, 35(1–3):C–508, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826934.

Laporta:1986:PCM

[LO86] S. Laporta and R. Odorico. Programs to calculate multiple QED radiation in leptonic decays of the Z^{0} and W^{±} weak bosons. Computer Physics Communications, 39(1):127–131, January 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586901670.

Lewis:1989:RRI

[LP89] John G. Lewis and Daniel J. Pierce. Recent research in iterative methods at boeing. Computer Physics Communications, 53(1–3):213–221, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901616.

Lovitch:1984:BSS

[LR84a] L. Lovitch and S. Rosati. Bound state solution of the twonucleon Schrödinger equation with tensor forces. Computer Physics Communications, 35(1–3):C–103, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823619.

Lovitch:1984:TNE

[LR84b] L. Lovitch and S. Rosati. The two-nucleon effectiverange parameters with tensor forces. Computer Physics Communications, 35(1–3):C–150–C–151, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824030.

Luke:1984:ESV

[LS84] T. M. Luke and H. E. Saraph. Ecsimpact, a special version of program impact for CDC machines with ex-core memory. Computer Physics Communications, 35(1–3):C–596–C–597, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827666.

Lauer:1987:RRD

[LSW87] R. Lauer, A. J. Slaughter, and E. Wolin. A role for relational databases in high energy physics software systems. Computer Physics Communications, 45(1–3):373–377, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901767.

Lima:1985:PLF

[LT85] Isabel Gouveia Lima and Philip C. Treleaven. Programming languages for fifth generation computers. Computer Physics Communications, 38(2):221–231, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900888.

Luce:1987:ACP

[LT87] R. J. Luce and F. Tabakin. Adaptation of Coulomb plus strong interaction bound states — momentum space solutions: Automatic gridpoint method. Computer Physics Communications, 46(1):193–201, July 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900464.

Lin:1988:GPS

[LTK88] C. S. Lin, A. L. Thring, and J. Koga. Gridless particle simulation using the massively parallel processor. Computer Physics Communications, 48(1):149–154, January 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900331.

Luvisetto:1983:ANL

[LU83] M. L. Luvisetto and E. Ugolini. Abacus: a natural language for the numerical computation of mathematical formulae. Computer Physics Communications, 30(3):277–299, November 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900966.

Luvisetto:1984:ANL

[LU84] M. L. Luvisetto and E. Ugolini. Abacus: a natural language for the numerical computation of mathematical formulae. Computer Physics Communications, 35(1–3):C–917, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584830167.

Luchini:1984:TDN

[Luc84] P. Luchini. Two-dimensional numerical integration using a square mesh. Computer Physics Communications, 31(4):303– 310, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900158.

Lux:1980:VEM

[Lux80] Iván Lux. Variance and efficiency in Monte Carlo transport calculations. Computer Physics Communications, 20(1):119–125, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580901174.

Linkens:1984:STS

[LVPG84] A. Linkens, J. Vanderschueren, P. Parot, and J. Gasiot. Simulation of thermally stimulated dipolar processes in dielectrics. Computer Physics Communications, 35(1–3):C–443, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584826508.

Lathrop:1984:MEP

[LW84a] B. L. Lathrop and B. R. Wienke. Maxwell: Exact photon cross-section processor for relativistic Maxwellian electrons. Computer Physics Communications, 32(3):309–315, June 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900936.

Lodhi:1984:SBS

[LW84b] M. A. K. Lodhi and B. T. Waak. Solution of bound state problems in nuclear shell models with momentum dependent potentials. Computer Physics Communications, 35(1–3):C–338, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825606.

Lathrop:1985:DPE

[LW85] B. L. Lathrop and B. R. Wienke. DEPOS: Parametric electron energy deposition module in slabs. Computer Physics Communications, 38(3):389–396, December 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901067.

Lao:1982:VCC

[LWHH82] L. L. Lao, R. M. Wieland, W. A. Houlberg, and S. P. Hirshman. VMOMS — a computer code for finding moment solutions to the Grad–Shafranov equation. Computer Physics Communications, 27(2):129–146, August 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900698.

Lao:1984:VCC

[LWHH84] L. L. Lao, R. M. Wieland, W. A. Houlberg, and S. P. Hirshman. VMOMS — a computer code for finding moment solutions to the Grad–Shafranov equation. Computer Physics Communications, 35(1–3):C–811, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482929X.

Libermani:1984:RPO

[LZ84] D. A. Libermani and A. Zangwill. A relativistic program for optical response in atoms using a time-dependent local density approximation. Computer Physics Communications, 32(1):75– 82, April 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900092.

Macfarlane:1989:ICC

[Mac89a] J. J. Macfarlane. IONMIX — a code for computing the equation of state and radiative properties of LTE and non-LTE plasmas. Computer Physics Communications, 56(2):259–278, December 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900234.

Mackenzie:1989:MLG

[Mac89b] Paul B. Mackenzie. Machines for lattice gauge theory. Computer Physics Communications, 57(1–3):37–46, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901902.

Maehlum:1988:BPE

[Mae88] B. N. Maehlum. Beam-plasma experiments. Computer Physics Communications, 49(1):119–132, April 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902202.

Magill:1981:RCT

[Mag81] J. Magill. Reply to “Comments on a time-dependent ionization algorithm”. Computer Physics Communications, 21(3): 295, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900102. See [MF81].

Magill:1984:TTD

[Mag84a] J. Magill. TRIP 1, a time-dependent recombination ionisation package. Computer Physics Communications, 35(1–3):C– 526, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827071.

Maguire:1984:VEC

[Mag84b] William C. Maguire. Vibrational energies of CO_{2}. Computer Physics Communications, 35(1–3):C–348, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825692.

Makhankov:1980:CES

[Mak80] Vladimir Makhankov. Computer experiments in soliton theory. Computer Physics Communications, 21(1):1–49, December 1, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900740.

Malaviya:1984:CCT

[Mal84] V. Malaviya. Computation of charge transfer probability between protons and excited hydrogen atoms. Computer Physics Communications, 35(1–3):C–44, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823103.

Mann:1983:SSF

[Man83a] Patrick J. Mann. The search for stable finite element methods for simple relativistic systems. Computer Physics Communications, 30(2):127–137, September/October 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900553.

Manoranjan:1983:SNE

[Man83b] V. S. Manoranjan. Some numerical experiments on Higgs model. Computer Physics Communications, 29(1):1–5, March 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900231.

Manickam:1986:BRB

[Man86]  J. Manickam.  Book review: Finite element methods

in linear ideal magnetohydrodynamics: Ralf Gruber and Jacques Rappaz, Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1985. 180 pages. DM 110.00. Computer Physics Communications, 41(1):191–192, July 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900299.

Mann:1987:FDM

[Man87a] Patrick J. Mann. Finite difference methods for the classical particle-particle gravitational n-body problem. Computer Physics Communications, 47(2–3):213–228, November/ December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901081.

Manzo:1987:MLS

[Man87b] John Manzo. On managing large scale projects: Some simple principles for developing complex systems. Computer Physics Communications, 45(1–3):215–228, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901597.

Manner:1989:PSA

[Män89] R. Männer. A programmable systolic array correlator as a trigger processor for electron pairs in rich (ring image Cherenkov) counters. Computer Physics Communications, 57 (1–3):516–519, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589902774.

Marquez:1984:NEC

[Mar84a] L. Marquez. Nearly exact calculation of the solution of the radial Schrödinger equation. Computer Physics Communications, 35(1–3):C–191, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824364.

Maruhn:1984:FFF

[Mar84b] J. A. Maruhn. Fourgen: a Fast Fourier Transform program generator. Computer Physics Communications, 35(1–3):C–393, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584826041.

Marty:1985:OOP

[Mar85] Rudolf Marty. Object oriented programming. Computer Physics Communications, 38(2):181–190, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900840.

Marshall:1989:MCM

[Mar89a] Guillermo Marshall. Monte Carlo methods for the solution of nonlinear partial differential equations. Computer Physics Communications, 56(1):51–61, November 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900520.

Martin:1989:SVR

[Mar89b] William R. Martin. Successful vectorization — reactor physics Monte Carlo code. Computer Physics Communications, 57 (1–3):68–77, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901938.

Martinec:1989:PCS

[Mar89c] Zdenek Martinec. Program to calculate the spectral harmonic expansion coefficients of the two scalar fields product. Computer Physics Communications, 54(1):177–182, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990043X.

Masetti:1981:TLD

[Mas81] Massimo Masetti. Trends in on-line data processing. Computer Physics Communications, 22(2–3):149–157, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900448.

Mason:1983:CBF

[Mas83] Janet P. Mason. Cylindrical Bessel functions for a large range of complex arguments. Computer Physics Communications, 30(1):1–11, July/August 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583901169.

Matsuoka:1984:FFG

[Mat84] Osamu Matsuoka. Field and field gradient integrals based on Gaussian type orbitals. Computer Physics Communications, 35(1–3):C–131, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823851.

Maturana:1987:SAM

[Mat87] Guillermo Maturana. Static allocation in massively parallel computers. Computer Physics Communications, 45(1– 3):319–322, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587901688.

Maysenholder:1981:TSC

[May81] W. Maysenhölder. Two subroutines for calculating lattice sums and the distortion field due to a point force in hexagonal systems. Computer Physics Communications, 24(1): 89–95, September/October 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581901089.

Maysenholder:1984:TSC

[May84] W. Maysenhölder. Two subroutines for calculating lattice sums and the distortion field due to a point force in hexagonal systems. Computer Physics Communications, 35(1–3):C–746, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828799.

May:1989:PPP

[May89] Edward N. May. Portable parallel programming in a Fortran environment. Computer Physics Communications, 57 (1–3):278–284, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589902282.

Mazza:1989:SPM

[Maz89] C. Mazza. Software project management. Computer Physics Communications, 57(1–3):23–28, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901884.

Moriarty:1983:EIM

[MB83] K. J. M. Moriarty and J. E. Blackshaw. Efficient implementation of the Monte Carlo method for lattice gauge theory calculations on the Floating Point Systems FPS-164. Computer Physics Communications, 29(2):155–161, April 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558390070X.

Moriarty:1984:EIM

[MB84] K. J. M. Moriarty and J. E. Blackshaw. Efficient implementation of the Monte Carlo method for lattice gauge theory calculations on the floating point systems FPS-164. Computer Physics Communications, 35(1–3):C–870–C–871, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829756.

Mirin:1986:OOD

[MBOK86] A. A. Mirin, R. J. Bonugli, N. J. O’Neill, and J. Killeen. ODRIC — a one-dimensional linear resistive MHD code in cylindrical geometry. Computer Physics Communications, 41(1):85– 103, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900238.

Moncrieff:1989:DMB

[MBW89] David Moncrieff, David J. Baker, and Stephen Wilson. Diagrammatic many-body perturbation expansion for atoms and molecules: VI: Experiments in vector processing and parallel processing for second-order energy calculations. Computer Physics Communications, 55(1):31–42, August 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900623.

Morand:1981:CPN

[MC81] C. Morand and Tsan Ung Chan. A computer program for nuclear lifetimes measurements by DSAM using a self-supporting target. Computer Physics Communications, 23(4):393–410, August 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901806.

McMullin:1984:FSA

[MC84a] J. N. McMullin and C. E. Capjack. Four-stage approximation in time-dependent ionization processes. Computer Physics Communications, 32(1):11–19, April 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900031.

Morand:1984:CPN

[MC84b] C. Morand and Tsan Ung Chan. A computer program for nuclear lifetimes measurements by dsam using a self-supporting target. Computer Physics Communications, 35(1–3):C–738, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828726.

Marchand:1985:PFC

[MC85] R. Marchand and C. E. Capjack. Polynomial fit to the Coulomb logarithm for inverse Bremsstrahlung. Computer Physics Communications, 38(3):357–358, December 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901031.

McMullin:1987:MAR

[MCA87] J. N. McMullin, C. E. Capjack, and S. Au. Multiparabolic approximation for ray tracing in linear plasma columns. Computer Physics Communications, 47(2–3):187–193, November/ December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901056.

McCubbin:1987:VCH

[McC87] N. A. McCubbin. VM/CMS in the HEP community (HEPVM). Computer Physics Communications, 45(1–3):61–66, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901408.

McColl:1989:PAA

[McC89] W. F. McColl. Parallel algorithms and architectures. Computer Physics Communications, 57(1–3):84–94, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901951.

McInnis:1984:SES

[McI84] B. C. McInnis. Socio-economic simulation models (SESM) and data bases of technology/resources. Computer Physics Communications, 33(1–3):257–265, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901310.

McMullin:1981:HLP

[MCJ81] J. N. McMullin, C. E. Capjack, and C. R. James. Heater: a 2D laser propagation subroutine for underdense plasmas. Computer Physics Communications, 23(1):31–42, June 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901272.

McMullin:1984:HLP

[MCJ84] J. N. McMullin, C. E. Capjack, and C. R. James. Hfater: a 2D laser propagation subroutine for underdense plasmas. Computer Physics Communications, 35(1–3):C–700, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828489.

McNeilly:1981:SNS

[McN81] Gregory S. McNeilly. SATDSK: a numerical simulation of the magnetic field due to saturated iron in cyclotron poletips. Computer Physics Communications, 23(2):199–207, July 1, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900333.

McNeilly:1984:SNS

[McN84] Gregory S. McNeilly. SATDSK: a numerical simulation of the magnetic field due to saturated iron in cyclotron poletips. Computer Physics Communications, 35(1–3):C–720, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046558482860X.

Manner:1981:HPF

[MD81] R. Männer and B. Deluigi. The Heidelberg POLYP — a flexible and fault-tolerant poly-processor. Computer Physics Communications, 22(2–3):279–284, April 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900606.

Martin:1982:PAI

[MD82] B. R. Martin and E. Dunford. Processing astronomical images from space. Computer Physics Communications, 26(3–4):441– 446, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901400.

Mehta:1988:LFI

[MD88] Manish A. Mehta and N. De Leon. LUCY: a FORTRAN implementation of semiclassical spectral quantization. Computer Physics Communications, 51(1–2):115–134, September/ October 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588900665.

Moller:1988:TBC

[MEB88] W. Möller, W. Eckstein, and J. P. Biersack. Tridynbinary collision simulation of atomic collisions and dynamic composition changes in solids. Computer Physics Communications, 51(3):355–368, November 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901488.

Meersman:1984:DBM

[Mee84] R. Meersman. Data bases and mainframes. Computer Physics Communications, 33(1–3):23–29, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901036.

Metcalf:1985:FF

[Met85]  Michael Metcalf. Has Fortran a future?  Computer

Physics Communications, 38(2):199–210, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900864.

Metcalf:1987:FES

[Met87] M. Metcalf. Fortran 8x — the emerging standard. Computer Physics Communications, 45(1–3):259–268, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901627.

Metcalf:1989:RPF

[Met89] Michael Metcalf. Recent progress in Fortran standardization. Computer Physics Communications, 57(1–3):78–83, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990194X.

Meurant:1989:PUC

[Meu89] Gérard Meurant. Practical use of the conjugate gradient method on parallel supercomputers. Computer Physics Communications, 53(1–3):467–477, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465589901793.

Mancini:1981:CTD

[MF81] Roberto C. Mancini and Constantino Ferro Fontán. Comments on a time-dependent ionization algorithm. Computer Physics Communications, 21(3):293–294, January 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900096. See reply [Mag81].

Manley:1984:DPC

[MG84] J. C. Manley and J. Gerratt. Director: a program for calculating representation matrices of the symmetric group in the Yamanouchi Kotani basis or in a direct product basis. Computer Physics Communications, 31(1):75–88, January 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900833.

Manickam:1981:LSA

[MGD81] J. Manickam, R. C. Grimm, and R. L. Dewar. The linear stability analysis of MHD models in axisymmetric toroidal geometry. Computer Physics Communications, 24(3–4):355–361, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901582.

Meyer:1986:FPC

[MGK86] R. L. Meyer, G. Giruzzi, and V. Krivenski. Fokker–Planck code for the quasi-linear absorption of electron cyclotron waves in a tokamak plasma. Computer Physics Communications, 40(1):153–157, May 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901578.

Moatti:1987:PMC

[MGM87] A. Moatti, J. Goldberg, and G. Memmi. Parallel Monte Carlo calculations with many microcomputers. Computer Physics Communications, 45(1–3):355–359, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901731.

Martorana:1984:XPS

[MGMZ84] Antonio Martorana, Rosalba Gerbasi, Antonio Marigo, and Roberto Zannetti. XFIT: a package for simulating and fitting X-ray powder diffraction patterns. Computer Physics Communications, 34(1–2):145–151, November/December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901668.

McKenzie:1980:PCT

[MGN80] B. J. McKenzie, I. P. Grant, and P. H. Norrington. A program to calculate transverse Breit and QED corrections to energy levels in a multiconfiguration Dirac–Fock environment. Computer Physics Communications, 21(2):233–246, December 2, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900429.

McKenzie:1984:PCT

[MGN84] B. J. McKenzie, I. P. Grant, and P. H. Norrington. A program to calculate transverse Breit and QED corrections to energy levels in a multiconfiguration Dirac–Fock environment. Computer Physics Communications, 35(1–3):C–663–C–664, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828209.

Munck:1984:CSM

[MGTD84] E. Münck, J. L. Grove, T. A. Tumolillo, and P. G. Debrunner. Computer simulations of Mössbauer spectra for an effective spin s = 1 Hamiltonian. Computer Physics Communications, 35(1–3):C–177, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824248.

Moriarty:1984:EIL

[MHP84] K. J. M. Moriarty, M. Haraguchi, and C. Pangali. Efficient implementation of the SU(3) lattice gauge theory algorithm on the Fujitsu VP200 vector processor. Computer Physics Communications, 34(1–2):1–7, November/December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558490153X.

Michels:1987:SPP

[Mic87] J. P. J. Michels. A study of polymer properties using numerical simulations. Computer Physics Communications, 44(3):289– 295, June 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900841.

Mika:1980:SEE

[Mik80] Janusz Mika. Systems of evolution equations and the singular perturbation method. Computer Physics Communications, 20(1):43–47, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580901058.

Miketinac:1984:CES

[Mik84] M. J. Miketinac. Calculation of the equilibrium structure and oscillations of polytropic stars pervaded by toroidal magnetic fields. Computer Physics Communications, 35(1–3):C–256, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824923.

Mikulecky:1986:KBS

[Mik86] Peter Mikulecký. On knowledge-based software tools. Computer Physics Communications, 41(2–3):397–401, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900792.

Mircheva:1984:SNC

[Mir84] D. Mircheva. On the search for nuclear characteristics in the evaluated nuclear structure data file (ENSDF) in the informative mode on the ES computer. Computer Physics Communications, 33(1–3):157–160, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901188.

Miura:1987:EVM

[Miu87] Kenichi Miura. EGS4V: Vectorization of the Monte Carlo cascade shower simulation code EGS4. Computer Physics Communications, 45(1–3):127–136, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901470.

McCoy:1986:TDS

[MKH86] M. G. McCoy, G. D. Kerbel, and R. W. Harvey. Threedimensional simulations of electron cyclotron heating. Computer Physics Communications, 40(1):115–121, May 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901530.

Moscinski:1989:VAE

[MKRJ89] J. Mościński, J. Kitowski, Z. A. Rycerz, and P. W. M. Jacobs. A vectorized algorithm on the ETA 10-P for molecular dynamics simulation of large number of particles confined in a long cylinder. Computer Physics Communications, 54(1):47–54, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900313.

Mlodzki:1983:PPC

[MKS83] J. Mlodzki, J. Kuszkowski, and M. Suffczynski. A Pascal program for calculating the reduced Coulomb Green’s functions and their partial waves. Computer Physics Communications, 29(4):341–350, June 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900139.

Mlodzki:1984:PPC

[MKS84] J. Mlodzki, J. Kuszkowski, and M. Suffczynski. A Pascal program for calculating the reduced Coulomb Green’s functions and their partial waves. Computer Physics Communications, 35(1–3):C–886, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829896.

Moon:1981:NAM

[ML81] Wooil Moon and Patrick Lui. A new approach to magnetic and gravitational potential field computation. Computer Physics Communications, 24(2):107–112, November 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900849.

Malegat:1986:ETC

[MLL86] L. Malegat, M. Le Dourneuf, and Vo Ky Lan. Errors in the three CPC versions of the program to calculate the single centre expansion of the electron-diatomic-molecule static potential. Computer Physics Communications, 41(1):179–189, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900287.

MLodzki:1984:IIPb

[MLo84] J. MLodzki. Integrals involved in the perturbation theory of a hydrogen-like system. II. Computer Physics Communications, 34(1–2):211–217, November/December 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901723.

Makhankov:1981:NIK

[MLS81] V. G. Makhankov, E. I. Litvinenko, and A. B. Shvachka. Numerical investigation of the Kadomtsev–Petviashvili soliton stability. Computer Physics Communications, 23(3):223–232, July 2, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900011.

MLodzki:1984:IIPa

[MLS84] J. MLodzki, A. Lusakowski, and M. Suffczyński. Integrals involved in the perturbation theory of a hydrogen-like system. I. Computer Physics Communications, 34(1–2):199– 209, November/December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584901711.

Miketinac:1982:SPP

[MM82] M. J. Miketinac and J. Miketinac. Shape of a polytropic primary. Computer Physics Communications, 27(3):285–297, September 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558290176X.

Manning:1984:DDE

[MM84a] Irwin Manning and G. P. Mueller. Depth distribution of energy deposition by ion bombardment. Computer Physics Communications, 35(1–3):C–228, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824674.

Miketinac:1984:SPP

[MM84b] M. J. Miketinac and J. Miketinac. Shape of a polytropic primary. Computer Physics Communications, 35(1–3):C–821–C– 822, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829379.

Meyer:1988:UTO

[MM88a] Veronika Meyer and Walter Meyer. The UNIX timesharing operating system. Computer Physics Communications, 50(1–2): 51–57, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901154.

Mirin:1988:SRC

[MM88b] A. A. Mirin and M. G. McCoy. SIGV5D: a routine to compute the reaction rates of interacting distribution functions. Computer Physics Communications, 51(3):369–372, November 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558890149X.

Mackenzie:1988:CAN

[MMBW88] L. M. Mackenzie, A. M. Macleod, D. J. Berry, and R. R. Whitehead. Concurrent algorithms for nuclear shell model calculations. Computer Physics Communications, 48(2):229–240, February 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588900434.

Mattar:1981:TCB

[MMF81] F. P. Mattar, G. Moretti, and R. E. Franceour. Transient counter-beam propagation in a nonlinear Fabry–Perot cavity. Computer Physics Communications, 23(1):1–17, June 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901247.

McCoy:1981:FTD

[MMK81] M. G. McCoy, A. A. Mirin, and J. Killeen. FPPAC: a two-dimensional multispecies nonlinear Fokker–Planck package. Computer Physics Communications, 24(1):37–61, September/ October 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901053.

McCoy:1984:FTD

[MMK84] M. G. McCoy, A. A. Mirin, and J. Killeen. FPPAC: a two-dimensional multispecies nonlinear Fokker–Planck package. Computer Physics Communications, 35(1–3):C–742–C– 743, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828763.

McDowell:1984:EIE

[MMM84] M. R. C. McDowell, L. Morgan, and V. P. Myerscough. Electron impact excitation cross sections. Computer Physics Communications, 35(1–3):C–225, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824649.

Matsuura:1985:SPT

[MMM85] Toshihiko Matsuura, Kenichi Miura, and Mitsuhiro Makino. Supervector performance without toil: FORTRAN implemented vector algorithms on the VP-100/200. Computer Physics Communications, 37(1–3):101–107, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901419.

Mirin:1989:TCW

[MMO89] A. A. Mirin, D. R. Martin, and N. J. O’Neill. TUBE88 — a code which computes magnetic field lines. Computer Physics Communications, 54(1):183–198, April 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900441.

Moriarty:1989:VCN

[MMR89] K. J. M. Moriarty, Eric Myers, and Claudio Rebbi. A vector code for the numerical simulation of cosmic strings and flux vortices in superconductors on the ETA-10. Computer Physics Communications, 54(2–3):273–294, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990091X.

Martin:1986:CTL

[MMS86] Olivier Martin, K. J. M. Moriarty, and Stuart Samuel. Computer techniques for lattice gauge theories. Computer Physics Communications, 40(2–3):173–179, June 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901062.

Marcu:1987:FAO

[MMS87] Mihail Marcu, Jürgen Müller, and Franz-Karl Schmatzer. Fast algorithm for one-dimensional quantum Monte Carlo simulations. Computer Physics Communications, 44(1–2):63–71, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900178.

Mirin:1988:FTD

[MMTK88] A. A. Mirin, M. G. McCoy, G. P. Tomaschke, and J. Killeen. FPPAC88: a two-dimensional multispecies nonlinear Fokker– Planck package. Computer Physics Communications, 51(3): 373–380, November 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901506.

Mattar:1980:ASR

[MN80] F. P. Mattar and M. C. Newstein. Adaptive stretching and rezoning as effective computational techniques for twolevel paraxial Maxwell–Bloch simulation. Computer Physics Communications, 20(1):139–163, September 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580901204.

Mattar:1984:CPD

[MN84a] F. P. Mattar and M. C. Newstein. Comments on P. D. Drummond’s central partial difference propagation algorithms. Computer Physics Communications, 32(2):225–227, May 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900742.

Merkel:1984:HQA

[MN84b] P. Merkel and J. Nührenberg. HASE — a quasi-analytical 2D MHD equilibrium code. Computer Physics Communications, 31(2–3):115–122, February 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900389.

Migneron:1988:NIM

[MN88] R. Migneron and K. S. S. Narayanan. Numerical inversion of Mellin moments and Laplace transforms. Computer Physics Communications, 49(3):457–463, June 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900069.

Meyer:1984:FPS

[MNJL84] J. Meyer, R. S. Nahabetian, J. Joseph, and J. Lafoucriere. SU3 fractional parentage in the 1p-shell. Computer Physics Communications, 35(1–3):C–107, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823656.

Migneron:1989:NIM

[MNR89] R. Migneron, K. S. S. Narayanan, and H. Rasmussen. Numerical inversion of moments. Computer Physics Communications, 54(2–3):239–244, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900866.

Mohamed:1980:MRA

[Moh80] Julie Mohamed. The method of Raptis and Allison with automatic error control. Computer Physics Communications, 20(2):309–320, October 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900090.

Mohamed:1984:MRA

[Moh84] Julie Mohamed. The method of Raptis and Allison with automatic error control. Computer Physics Communications, 35(1–3):C–638, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482799X.

Mohanakrishnan:1989:UAN

[Moh89] P. Mohanakrishnan. Use of albedo for neutron reflector regions in reactor core 3-D simulations. Computer Physics Communications, 55(3):311–323, October 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990129X.

Montgomery:1987:SUH

[Mon87] H. E. Montgomery. Status of (US) high energy physics networking. Computer Physics Communications, 45(1–3):77–82, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901421.

Monaghan:1988:IS

[Mon88] J. J. Monaghan. An introduction to SPH. Computer Physics Communications, 48(1):89–96, January 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900264.

Moon:1980:AFO

[Moo80] Wooil Moon. Algorithm for the first order hydrostatic ellipticity of a planet. Computer Physics Communications, 19 (1):63–67, January/March 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900673.

Moon:1981:S

[Moo81a] Wooil Moon. J-square. Computer Physics Communications, 22(1):97–101, February/March 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900825.

Moon:1981:AFC

[Moo81b] Wooil Moon. Airy function with complex arguments. Computer Physics Communications, 22(4):411–417, May 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901387.

Moon:1983:SWE

[Moo83] Wooil Moon. Solution of the whole earth geodynamic equation in a slightly elliptical coordinate system. Computer Physics Communications, 28(3):219–227, January 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900401.

Moon:1984:S

[Moo84a] Wooil Moon. J-square. Computer Physics Communications, 35(1–3):C–688, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828398.

Moon:1984:AFC

[Moo84b] Wooil Moon. Airy function with complex arguments. Computer Physics Communications, 35(1–3):C–692, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828428.

Moon:1984:AFO

[Moo84c] Wooil Moon. Algorithm for the first order hydrostatic ellipticity of a planet. Computer Physics Communications, 35(1–3): C–609, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827769.

Moon:1984:IMD

[Moo84d] Wooil Moon. An interpolation method for data sets with jump discontinuities. Computer Physics Communications, 35(1–3): C–539, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827162.

Moon:1984:NEG

[Moo84e] Wooil Moon. Numerical evaluation of geomagnetic dynamo integrals (elsasser and adams-gaunt integrals). Computer Physics Communications, 35(1–3):C–538, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827150.

Moore:1984:NVP

[Moo84f] C. Fred Moore. A new version of a printer-plotter routine. Computer Physics Communications, 35(1–3):C–114, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823711.

Moore:1984:PPR

[Moo84g] C. Fred Moore. Printer-plotter routine. Computer Physics Communications, 35(1–3):C–65, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0010465584823309.

Moores:1984:ASE

[Moo84h] D. L. Moores. Amplitudes for scattering of electrons by hydrogenic and alkali-like atomic systems. Computer Physics Communications, 35(1–3):C–104, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823620.

Morrison:1980:APC

[Mor80] Michael A. Morrison. ALAM, a program for the calculation and expansion of molecular charge densities. Computer Physics Communications, 21(1):63–77, December 1, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900764.

Moraal:1984:PEB

[Mor84a] H. Moraal. A program for the extraction of bulk viscosities from sound absorption data. Computer Physics Communications, 35(1–3):C–115, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823735.

Morgan:1984:EES

[Mor84b] K. Morgan. An expansion equation of state subroutine. Computer Physics Communications, 35(1–3):C–167, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824182.

Morgan:1984:GMP

[Mor84c] Lesley A. Morgan. A generalized R-matrix propagation program for solving coupled second-order differential equations. Computer Physics Communications, 31(4):419–422, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900250.

Morrison:1984:APC

[Mor84d] Michael A. Morrison. ALAM, a program for the calculation and expansion of molecular charge densities. Computer Physics Communications, 35(1–3):C–650, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828106.

Morrow:1984:MCI

[Mor84e] Richard A. Morrow. Monte Carlo integration program for the n-particle relativistic phase space integral in invariant variables. Computer Physics Communications, 35(1–3):C–441–C– 442, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584826491.

Mount:1987:DSH

[Mou87] Richard P. Mount. Database systems for HEP experiments. Computer Physics Communications, 45(1–3):299–310, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901664.

Mount:1989:OET

[Mou89] R. P. Mount. Overview of the essential tools. Computer Physics Communications, 57(1–3):140–147, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902038.

Marsh:1981:RSB

[MP81a] F. Marsh and D. E. Potter. Recurrence solution of a block tridiagonal matrix equation with neumann, dirichlet, mixed or periodic boundary conditions. Computer Physics Communications, 24(2):185–190, November 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900928.

Merdes:1981:FPP

[MP81b] Daniel W. Merdes and Josef Pliva. A flexible program for performing analytic differentiation and substitutions on a system of equations. Computer Physics Communications, 24(2):113–125, November 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900850.

Modi:1982:SJM

[MP82] J. J. Modi and D. Parkinson. Study of Jacobi methods for eigenvalues and singular value decomposition on DAP. Computer Physics Communications, 26(3–4):317–320, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901229.

Marsh:1984:RSB

[MP84a] F. Marsh and D. E. Potter. Recurrence solution of a block tridiagonal matrix equation with Neumann, Dirichlet, mixed or periodic boundary conditions. Computer Physics Communications, 35(1–3):C–758, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482888X.

Merdes:1984:FPP

[MP84b] Daniel W. Merdes and Josef Pliva. A flexible program for performing analytic differentiation and substitutions on a system of equations. Computer Physics Communications, 35(1–3):C–749– C–750, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828817.

Mueller:1989:CCD

[MP89] K. Mueller and P. Pfeifer. Codebase: a commercially developed code management system and code transfer facility. Computer Physics Communications, 57(1–3):239–243, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902208.

Monticello:1986:RCR

[MPIM86] D. A. Monticello, W. Park, R. Izzo, and K. McGuire. A review of calculations of the resistive internal m = 1 mode in tokamaks. Computer Physics Communications, 43(1):57–67, December 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900536.

McCarville:1983:FCC

[MPM83] Thomas J. McCarville, Robert R. Peterson, and Gregory A. Moses. Fire — a code for computing the response of an inertial confinement fusion cavity gas to a target explosion. Computer Physics Communications, 28(4):367–403, February 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900346.

McCarville:1984:FCC

[MPM84] Thomas J. McCarville, Robert R. Peterson, and Gregory A. Moses. Fire — a code for computing the response on an inertial confinement fusion cavity gas to a target explosion. Computer Physics Communications, 35(1–3):C–858–C–859, ????

1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829677.

Moses:1985:MFM

[MPM85] Gregory A. Moses, Robert R. Peterson, and Thomas J. Mc-Carville. MF-FIRE — a multifrequency radiative transfer hydrodynamics code. Computer Physics Communications, 36(3):249–293, May 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900554.

Miraglia:1980:CME

[MPR80] J. E. Miraglia, R. D. Piacentini, and R. D. Rivarola. Computational methods for eikonal differential cross sections. Computer Physics Communications, 19(3):299–303, July/August 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900843.

Menasce:1981:PIG

[MPS81] D. Menasce, F. Palombo, and S. Sala. PRAVDA: An interactive graphic system to scan and rescue events. Computer Physics Communications, 22(2–3):317–323, April 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900679.

McKay:1984:EEL

[MPS84] W. McKay, J. Patera, and R. T. Sharp. Eigenstates and eigenvalues of labelling operators for O(3) bases of U(3) representations. Computer Physics Communications, 35(1–3):C–322, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825461.

Malmstrom:1980:PCR

[MR80] G. Malmström and J. Rundgren. A program for calculation of the reflection and transmission of electrons through a surface potential barrier. Computer Physics Communications, 19(2):263–270, April/June 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900533.

Malmstrom:1984:PCR

[MR84a] G. Malmström and J. Rundgren. A program for calculation of the reflection and transmission of electrons through a surface potential barrier. Computer Physics Communications, 35(1–3): C–620, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827848.

Moussa:1984:AIC

[MR84b] A. H. Moussa and H. M. A. Radi. Atomic integral containing three odd powers of interelectronic separation coordinates. Computer Physics Communications, 35(1–3):C–211, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824534.

Moriarty:1986:QFT

[MR86] K. J. M. Moriarty and C. Rebbi. Quantum field theory and supercomputers. Computer Physics Communications, 40(2–3):181–188, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901074.

Moriarty:1987:LSQ

[MR87] K. J. M. Moriarty and C. Rebbi. Large-scale quantum field theory calculations on supercomputers. Computer Physics Communications, 47(1):75–82, October 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900671.

Marchand:1985:TIT

[MRC85] R. Marchand, R. Rankin, and C. E. Capjack. Transient ionization time scales for low atomic number elements. Computer Physics Communications, 38(3):359–363, December 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901043.

Macias:1984:PCG

[MRCL84] E. S. Macias, W. D. Ruhter, D. C. Camp, and R. G. Lanier. A program for calculating gamma-gamma directional correlation coefficients and mixing pations. Computer Physics Communications, 35(1–3):C–363–C–364, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825813.

Morales:1989:ETT

[MRT89] Juan J. Morales, Luis F. Rull, and Søren Toxvaerd. Efficiency test of the traditional MD and the link-cell methods. Computer Physics Communications, 56(2):129–134, December 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900131.

Manning:1984:APD

[MRW84] Irwin Manning, Mervine Rosen, and J. E. Westmoreland. Adaptation of a program for depth distribution of energy deposition by ion bombardment: Calculation of ion lateral ranges. Computer Physics Communications, 35(1–3):C–408, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046558482617X.

McGreevy:1984:PCE

[MS84] E. McGreevy and A. L. Stewart. A program for calculating elastic scattering phase shifts for an electron colliding with a one-electron target using perturbation theory. Computer Physics Communications, 35(1–3):C–454, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826570.

Maca:1985:CGF

[MS85] Franti sek Máca and Matthias Scheffler. Calculation of the Green’s function for a crystal surface or interface. Computer Physics Communications, 38(3):403–413, December 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901080.

Maca:1987:NVP

[MS87a] Frantisek Máca and Matthias Scheffler. A new version of the program for the calculation of the Green’s function for a crystal surface or interface. Computer Physics Communications, 47(2–3):349–350, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901196.

Menasce:1987:MDE

[MS87b] D. Menasce and S. Sala. The Mida database of the E687 μ strip vertex detector. Computer Physics Communications, 45(1–3):385–389, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587901780.

Maca:1988:SGF

[MS88] Frantisek Máca and Matthias Scheffler. Surface Green’s function for a rumpled crystal surface. Computer Physics Communications, 51(3):381–390, November 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901518.

Mankofsky:1988:DDP

[MSC^{+}88] A. Mankofsky, J. L. Seftor, C. L. Chang, K. Ko, A. A. Mondelli, A. T. Drobot, J. Moura, W. Aimonetti, S. T. Brandon, D. E. Nielsen, Jr., and K. M. Dyer. Domain decomposition and particle pushing for multiprocessing computers. Computer Physics Communications, 48(1):155–165, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900343.

Meyer:1989:CRS

[MSFR89] J. Meyer, W.-D. Sepp, B. Fricke, and A. Rosén. Computation of relativistic symmetry orbitals for finite double point groups. Computer Physics Communications, 54(1):55–73, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900325.

Mattar:1981:FFH

[MT81] F. P. Mattar and J. Teichmann. Fluid formulation of high intensity laser beam propagation using Lagrangian coordinates. Computer Physics Communications, 22(1):1–11, February/March 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900758.

Moriarty:1984:PCO

[MT84a] K. J. M. Moriarty and J. H. Tabor. A program for calculating the observables for single-particle-inclusive production reactions. Computer Physics Communications, 35(1–3):C–403, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826132.

Moriarty:1984:PAC

[MT84b] K. J. M. Moriarty and H. N. Thompson. Program adaptation: To calculate inclusive backward proton cross sections. Computer Physics Communications, 35(1–3):C–587, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827599.

Moriarty:1986:MTC

[MT86] K. J. M. Moriarty and M. Teper. Measuring topological charge in Monte Carlo simulation of SU(N) lattice gauge theories. Computer Physics Communications, 42(1):1–9, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586902237.

Matsuura:1982:VPE

[MTN^{+}82] T. Matsuura, Y. Tanaka, K. Naraoka, T. Takizuka, T. Tsunematsu, S. Tokuda, M. Azumi, G. Kurita, and T. Takeda. Vector processing efficiency of plasma MHD codes by use of the FACOM 230-75 APU. Computer Physics Communications, 26(3–4):377–387, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901321.

Mueller:1984:REF

[Mue84] G. P. Mueller. Rapid evaluation of four-body cluster contributions. Computer Physics Communications, 35(1–3):C–85, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823474.

Muller:1987:CPS

[Mül87] Ewald Müller. Computational problems in supernova simulations. Computer Physics Communications, 44(3):271–277, June 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900828.

Murtagh:1988:HTB

[Mur88] F. D. Murtagh. Hierarchical trees in N-body simulations: Relations with cluster analysis methods. Computer Physics Communications, 52(1):15–18, December 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558890166X.

Masovic:1981:BSC

[MV81] D. R. Masović and F. Vukajlović. Band structure calculations of cubic semiconductors on the basis of Löwdin’s perturbation technique. Computer Physics Communications, 24(2):181–184, November 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900916.

Masovic:1983:BSC

[MV83] D. R. Mašović and F. R. Vukajlović. Band structure calculations of cubic metals, elementary semiconductors and semiconductor compounds with spin-orbit interaction. Computer Physics Communications, 30(2):207–217, September/October 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900632.

Masovic:1984:BSCa

[MV84a] D. R. Masović and F. Vukajlović. Band structure calculations of cubic semiconductors on the basis of Löwdin’s perturbation technique. Computer Physics Communications, 35(1–3):

C–757, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828878.

Masovic:1984:BSCb

[MV84b] D. R. Masović and F. R. Vukajlović. Band structure calculations of cubic metals, elementary semiconductors and semiconductor compounds with spin-orbit interaction. Computer Physics Communications, 35(1–3):C–911, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584830106.

Maino:1987:NPD

[MV87] G. Maino and A. Ventura. NUCPAR — a parity-dependent NBCS formalism at finite nuclear temperature. Computer Physics Communications, 43(2):303–312, January 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587902141.

Meyer-Vernet:1988:CPW

[MV88] N. Meyer-Vernet. Cometary plasma wave observations. Computer Physics Communications, 49(1):9–15, April 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902135.

Maino:1983:NNB

[MVV83] G. Maino, M. Vaccari, and A. Ventura. NUDENS: a Nilsson– Bardeen–Cooper–Schrieffer code at finite nuclear temperature. Computer Physics Communications, 29(4):375–390, June 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900164.

Maino:1984:NNB

[MVV84] G. Maino, M. Vaccari, and A. Ventura. NUDENS: a Nilsson– Bardeen–Cooper–Schrieffer code at finite nuclear temperature. Computer Physics Communications, 35(1–3):C–889, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829926.

Marsr:1984:FPC

[MW84a] F. N. Marsr and I. R. Williams. I. A Fortran program for calculating degenerate Raman bands of symmetric tops with an adaptation for infrared bands. Computer Physics Communications, 35(1–3):C–96, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823553.

Masri:1984:IFP

[MW84b] F. N. Masri and I. R. Williams. II. A Fortran program for calculating degenerate Raman bands of spherical tops with an adaptation for infrared bands. Computer Physics Communications, 35(1–3):C–97, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823565.

Masri:1984:IPC

[MW84c] F. N. Masri and I. R. Williams. II. Program for calculating triply degenerate Raman bands of spherical tops with an adaptation for infrared bands. Computer Physics Communications, 35(1–3):C–68, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823334.

Myers:1988:ICG

[Mye88] David R. Myers. Interactive computer graphics and PHIGS. Computer Physics Communications, 50(1–2):143–157, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901221.

Myers:1989:HGS

[Mye89] D. R. Myers. HEP graphics: Standards and portability versus performance and cost. Computer Physics Communications, 57(1–3):176–182, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902087.

Martorana:1982:PPC

[MZM^{+}82] Antonino Martorana, Roberto Zannetti, Antonio Marigo, David Ajò, and Viscardo Malta. PROVA, a program for the calculation of X-ray powder spectra (ordered and disordered structures). Computer Physics Communications, 27(1):49–55, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900066.

Martorana:1984:PPC

[MZM^{+}84] Antonino Martorana, Roberto Zannetti, Antonio Marigo, David Ajò, and Viscardo Malta. Prova, a program for the calculation of X-ray powder spectra (ordered and disordered structures). Computer Physics Communications, 35(1–3):C–804–C– 805, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829240.

Mustafa:1989:STN

[MZZ89] Mustafa M. Mustafa, Mark W. Zahran, and Elbadry S. Zahran. Solution of the two-nucleons Schrödinger equation with nonlocal tensor potential in the ^{3}S_{1} – ^{3}D_{1} state. Computer Physics Communications, 55(1):109–116, August 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900696.

Neufeld:1984:PAC

[NA84] P. D. Neufeld and R. A. Aziz. Program acqn to calculate transport collision integrals adapted to run on IBM computers. Computer Physics Communications, 35(1–3):C–140, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823942.

Nadrchal:1980:P

[Nad80] J. Nadrchal. Preface. Computer Physics Communications, 20(1):ix, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558090096X.

Nadrchal:1986:BRB

[Nad86a] J. Nadrchal. Book review: Effective Fortran 77: M. Metcalf, Clarendon Press, Oxford, 1985, xii + 231 p., £9.95. Computer Physics Communications, 42(1):149–150, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586902390.

Nadrchal:1986:SEC

[Nad86b] J. Nadrchal. Software engineering and computational physics. Computer Physics Communications, 41(2–3):197–203, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900640.

Nagl:1985:IPS

[Nag85] M. Nagl. An incremental programming support environment. Computer Physics Communications, 38(2):245–276, October/ November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585900906.

Nagel:1989:HIS

[Nag89] D. Nagel. The human interface in scientific computing. Computer Physics Communications, 57(1–3):561, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902907.

Nair:1984:FPC

[Nai84] K. P. Rajappan Nair. A Fortran program for the calculation of hyperfine structure and stark effect in the rotational transition of a ^{2}Σ diatomic molecule. Computer Physics Communications, 34(1–2):163–174, November/December 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901681.

Nair:1986:FPC

[Nai86] K. P. R. Nair. A Fortran program for the calculation of hyperfine structure in the rotational transition of a ^{2}Σ diatomic molecule: II. Magnetic and electric quadrupole interaction from both nuclei. Computer Physics Communications, 41(1):59– 73, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900214.

Nash:1984:HME

[Nas84] J. C. Nash. The Hermitian matrix eigenproblem Hx = eSx using compact array storage. Computer Physics Communications, 35(1–3):C–264, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824984.

Nash:1987:ACM

[Nas87a] Thomas Nash. Asilomar conference on managing complexity in high energy physics: a summary and renaming of the conference. Computer Physics Communications, 45(1–3):9–14, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901366.

Nilsson-Almqvist:1987:IBH

[NAS87b] Bo Nilsson-Almqvist and Evert Stenlund. Interactions between hadrons and nuclei: The lund Monte Carlo — FRITIOF version 1.6-. Computer Physics Communications, 43(3): 387–397, February/March 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900567.

Nash:1989:EPD

[Nas89] Thomas Nash. Event parallelism: Distributed memory parallel computing for high energy physics experiments. Computer Physics Communications, 57(1–3):47–56, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901914.

Navarra:1989:AGI

[Nav89] Antonio Navarra. An application of GMRES to indefinite linear problems in meteorology. Computer Physics Communications, 53(1–3):321–327, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901690.

Nguyen:1985:IVE

[NBI^{+}85] V. T. Nguyen, P. Bertrand, B. Izrar, E. Fijalkow, and M. R. Feix. Integration of Vlasov equation by quantum mechanical formalism. Computer Physics Communications, 34(3):295–301, January 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585900062.

Nesbet:1984:ARI

[Nes84] R. K. Nesbet. Algorithms for regular and irregular Coulomb and Bessel functions. Computer Physics Communications, 32(4):341–347, July 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900511.

Nesbet:1988:AEI

[Nes88] R. K. Nesbet. Analytical evaluation of integrals over Coulomb wave functions. Computer Physics Communications, 52(1): 29–33, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901683.

Neto:1984:NCI

[Net84] N. Neto. Numerical calculations of the irreducible representations of space groups. Computer Physics Communications, 35(1–3):C–304, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825321.

Neves:1982:MSL

[Nev82] Kenneth W. Neves. Mathematical software libraries for vector computers. Computer Physics Communications, 26(3–4):303– 310, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901205.

Newberger:1986:ENC

[New86]  Barry S. Newberger.  Efficient numerical computation

of the plasma dispersion function. Computer Physics Communications, 42(3):305–311, November 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900019.

Newman:1987:CHE

[New87] Harvey B. Newman. Computing for HEP experiments: 1987-1997. Computer Physics Communications, 45(1–3):27–45, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558790138X.

Newman:1989:NHE

[New89] H. B. Newman. Networks for high energy physics. Computer Physics Communications, 57(1–3):561, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902919.

Nex:1980:CMS

[Nex80a] C. M. M. Nex. Computational methods for sparse matrices. Computer Physics Communications, 20(1):7–10, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580901009.

Nex:1980:UCS

[Nex80b] C. M. M. Nex. The use of Chebyshev series in computational physics. Computer Physics Communications, 20(1):1–5, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900995.

Nex:1984:RMP

[Nex84] C. M. M. Nex. The recursion method: Processing the continued fraction. Computer Physics Communications, 34(1–2): 101–122, November/December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584901632.

Nex:1989:BLA

[Nex89] C. M. M. Nex. The block Lanczos algorithm and the calculation of matrix resolvents. Computer Physics Communications, 53(1–3):141–146, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901550.

Newland:1983:CSS

[NF83] D. B. Newland and M. T. C. Fang. The computation of steady state arcs with mild nozzle — wall ablation. Computer Physics Communications, 28(3):299–314, January 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900474.

Newland:1984:CSS

[NF84] D. B. Newland and M. T. C. Fang. The computation of steady state arcs with mild nozzle — Wall ablation. Computer Physics Communications, 35(1–3):C–854, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482963X.

Nakamura:1988:FAE

[NFH^{+}88] A. Nakamura, G. Feuer, H. C. Hege, V. Linke, and M. Haraguchi. Fast algorithm for an exact treatment of lattice quantum chromodynamics by Monte Carlo simulation on vector processors. Computer Physics Communications, 51(3):301–315, November 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901439.

Nagahara:1982:SGS

[NHK82] Izuru Nagahara, Shigeru Hosaka, and Shigetoshi Katsura. Spin-glass in the site diluted Ising ferromagnet: Octahedron approximation in the FCC lattice. Computer Physics Communications, 27(2):113–117, August 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900674.

Nicholls:1989:FCC

[Nic89] J. Nicholls. The Fermilab central computing facility architectural model. Computer Physics Communications, 57(1– 3):417–421, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589902543.

Nielsen:1988:ESD

[Nie88] Jan Højlund Nielsen. Embedded systems development using modula-2. Computer Physics Communications, 50(1–2):59– 70, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901166.

Nishida:1988:IFP

[Nis88] A. Nishida. ISAS future programs. Computer Physics Communications, 49(1):253–255, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588902330.

Nielsen:1982:LDZ

[NJ82] O. H. Nielsen and S. S. Jaswal. Lattice dynamics of zincblende structure compounds using a deformable ion model. Computer Physics Communications, 25(3):269–273, March 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900224.

Nielsen:1984:LDZ

[NJ84] O. H. Nielsen and S. S. Jaswal. Lattice dynamics of zincblende structure compounds using a deformable ion model. Computer Physics Communications, 35(1–3):C–788, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829136.

Nagel:1984:FRC

[NK84] Pierre Nagel and R. D. Koshel. A finite range coupled channel born approximation code. Computer Physics Communications, 35(1–3):C–496–C–497, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826843.

Noid:1988:ESC

[NKS88] D. W. Noid, S. K. Knudson, and B. G. Sumpter. Exact semiclassical calculation of eigenvalues for multidimensional systems using SOS. Computer Physics Communications, 51(1–2):11–15, September/October 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900586.

Naccache:1984:PSA

[NM84] P. F. Naccache and M. R. C. McDowell. Phase shift analysis and consistency checks on electron-atom collision data. Computer Physics Communications, 35(1–3):C–209, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824510.

Noble:1984:CPC

[NN84] C. J. Noble and R. K. Nesbet. CFASYM, a program for the calculation of the asymptotic solutions of the coupled equations of electron collision theory. Computer Physics Communications, 33(4):399–411, October 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901450.

Nour-Omid:1989:ALM

[NO89] B. Nour-Omid. Applications of the Lanczos method. Computer Physics Communications, 53(1–3):157–168, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901574.

Noble:1980:SEE

[Nob80] C. J. Noble. Subroutines for the evaluation of exchange integrals in the impact parameter formulation of atomic charge transfer collisions. Computer Physics Communications, 19(3):327–335, July/August 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900879.

Noble:1984:SEE

[Nob84] C. J. Noble. Subroutines for the evaluation of exchange integrals in the impact parameter formulation of atomic charge transfer collisions. Computer Physics Communications, 35(1–3):C–622– C–623, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827873.

Noga:1983:CFO

[Nog83] J. Noga. Calculation of the fourth-order triple-excitation contribution to MB RSPT by using spatial symmetry properties of molecules. Computer Physics Communications, 29(2):117– 124, April 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558390067X.

Northrop:1982:APA

[Nor82] G. A. Northrop. Acoustic phonon anisotropy: Phonon focusing. Computer Physics Communications, 28(1):103–107, November 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900650.

Norcross:1984:ASC

[Nor84a] David W. Norcross. Asymptotic solution of coupled equations for electron scattering. Computer Physics Communications, 35(1–3):C–12, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584822821.

Northrop:1984:APA

[Nor84b] G. A. Northrop. Acoustic phonon anisotropy: Phonon focusing. Computer Physics Communications, 35(1–3):C–838, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829495.

Novak:1987:PEV

[Nov87] Miroslav M. Novak. A program to evaluate vibrationally inelastic collisional cross sections of atom-diatom systems. Computer Physics Communications, 46(3):417–425, September 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900956.

Nazarewicz:1984:MAI

[NP84] W. Nazarewicz and M. Pindor. Multidimensional automatic integrator (MDAI) — an efficient routine for automatic integration of functions of many variables. Computer Physics Communications, 31(1):1–12, January 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900778.

Nobile:1986:EIM

[NR86a] A. Nobile and V. Roberto. Efficient implementation of multidimensional Fast Fourier Transforms on a Cray X-MP. Computer Physics Communications, 40(2–3):189–201, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901086.

Nobile:1986:MPT

[NR86b] A. Nobile and V. Roberto. MFFT: a package for twoand three-dimensional vectorized discrete Fourier transforms. Computer Physics Communications, 42(2):233–247, October 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900391.

Nobile:1988:MFD

[NRS88] A. Nobile, V. Roberto, and F. Saitta. MFFT4: Four dimensional vectorized fast Fourier transforms. Computer Physics Communications, 48(2):313–318, February 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900501.

Noble:1984:CPE

[NT84] C. J. Noble and I. J. Thompson. COULN, a program for evaluating negative energy Coulomb functions. Computer Physics Communications, 33(4):413–419, October 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901462.

Nuhrenberg:1986:P

[Nüh86] Jürgen Nührenberg. Preface. Computer Physics Communications, 43(1):xi, December 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900470.

Nurzynski:1985:APE

[Nur85] J. Nurzynski. Analysis of polarization experiments. Computer Physics Communications, 36(3):295–305, May 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900566.

Nussbaumer:1984:VCC

[Nus84] H. Nussbaumer. Vector coupling coefficients for complex atoms. Computer Physics Communications, 35(1–3):C–22, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822900.

Nielsen:1984:LDG

[NW84] O. H. Nielsen and W. Weber. Lattice dynamics of group IV semiconductors using an adiabatic bond charge model. Computer Physics Communications, 35(1–3):C–580, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827538.

Olson:1984:SCP

[OBLR84] D. L. Olson, J. L. Blough, T. S. Lakshmanan, and D. A. Rigney. Simulation of chemical profiles during electrotransport. Computer Physics Communications, 35(1–3):C–197, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482442X.

OConnell:1984:CLF

[O’C84a] M. J. O’Connell. Computing a Laplacian field component from boundary observations only. Computer Physics Communications, 35(1–3):C–373, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825886.

OConnell:1984:MLS

[O’C84b] M. J. O’Connell. Multivariate least squares fitting program using modified Gram–Schmidt transformations. Computer Physics Communications, 35(1–3):C–261, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824960.

OConnell:1984:SPS

[O’C84c] M. J. O’Connell. Search program for significant variables. Computer Physics Communications, 35(1–3):C–260, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824959.

Ortoleva:1980:PIL

[OCC80] Emanuele Ortoleva, Giorgio Castiglione, and Enrico Clementi. A program to introduce local symmetry in ab-initio computations of molecules: IBMOL-7. Computer Physics Communications, 19(3):337–357, July/August 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900880.

Ortoleva:1984:PIL

[OCC84] Emanuele Ortoleva, Giorgio Castiglione, and Enrico Clementi. A program to introduce local symmetry in ab-initio computations of molecules: Ibmol-7. Computer Physics Communications, 35(1–3):C–624, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827885.

OBrien:1986:FPC

[OCS86] M. R. O’Brien, M. Cox, and D. F. H. Start. Fokker– Planck calculations of rf heating in tokamaks. Computer Physics Communications, 40(1):123–129, May 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901542.

Odorico:1981:MCP

[Odo81] R. Odorico. A Monte Carlo program for QCD event simulation in electron-positron annihilation at LEP energies. Computer Physics Communications, 24(1):73–87, September/ October 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901077.

Odorico:1982:HMC

[Odo82] R. Odorico. Hevol: a Monte Carlo program to calculate the evolution of structure functions with the inclusion of heavy quark effects. Computer Physics Communications, 25(3):253– 268, March 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900212.

Odorico:1984:CMC

[Odo84a] R. Odorico. COJETS: a Monte Carlo program simulating QCD in hadronic production of jets and heavy flavours with inclusion of initial QCD Bremsstrahlung. Computer Physics Communications, 32(2):139–172, May 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900675.

Odorico:1984:HMC

[Odo84b] R. Odorico. Hevol: a Monte Carlo program to calculate the evolution of structure functions with the inclusion of heavy quark effects. Computer Physics Communications, 35(1–3):C– 787, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829124.

Odorico:1984:MCP

[Odo84c] R. Odorico. A Monte Carlo program for QCD event simulation in electron-positron annihilation at LEP energies. Computer Physics Communications, 35(1–3):C–745, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828787.

Odorico:1984:WMC

[Odo84d] R. Odorico. WIZJET: a Monte Carlo program for hadronic production of W^{±} and Z^{0} simulating the standard model with inclusion of initial QCD Bremsstrahlung. Computer Physics Communications, 32(2):173–184, May 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900687.

Odstrcil:1980:TDM

[Ods80] Dušan Odstrčil. Two-dimensional model of the plasma coaxial accelerator. Computer Physics Communications, 20(1):65–68, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580901095.

Odstrcil:1984:BPD

[Ods84] D. Odstrcil. Boreas: a program for 1-D ideal fluid dynamics with shocks. Computer Physics Communications, 31(1):13–28, January 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558490078X.

Ogilvie:1983:SEC

[Ogi83] J. F. Ogilvie. Spectroscopic energy coefficients for vibrationrotational states of dinuclear molecules. Computer Physics Communications, 30(1):101–105, July/August 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583901273.

Ogilvie:1984:SEC

[Ogi84] J. F. Ogilvie. Spectroscopic energy coefficients for vibrationrotational states of dinuclear molecules. Computer Physics Communications, 35(1–3):C–904, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0010465584830039.

Oppe:1989:ONN

[OJK89] Thomas C. Oppe, Wayne D. Joubert, and David R. Kincaid. An overview of NSPCG: a nonsymmetric preconditioned conjugate gradient package. Computer Physics Communications, 53(1–3):283–293, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901665.

Obdrzalek:1988:PSP

[OL88] J. Obdrzálek and J. Luzný. Polyman — a system for polynomials. Computer Physics Communications, 50(1–2):255, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901336.

Olken:1987:DMH

[OLR^{+}87] Frank Olken, Stewart C. Loken, Doron Rotem, Arie Shoshani, and Thomas G. Trippe. Data management for high energy physics experiments — preliminary proposals. Computer Physics Communications, 45(1–3):379–383, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901779.

Olszewski:1983:PPV

[Ols83] E. A. Olszewski. A program for perspective views of open surfaces. Computer Physics Communications, 30(3):259–269, November 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900942.

Olszewski:1984:PPVb

[Ols84] E. A. Olszewski. A program for perspective views of open surfaces. Computer Physics Communications, 35(1–3):C–915, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584830143.

ORaifeartaigh:1983:CDS

[OM83] C. O’Raifeartaigh and J. F. McGilp. Calculation of differential scattering cross-sections for classical binary elastic collisions. Computer Physics Communications, 28(3):255–264, January 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900425.

ORaifeartaigh:1984:CDS

[OM84] C. O’Raifeartaigh and J. F. McGilp. Calculation of differential scattering cross-sections for classical binary elastic collisions. Computer Physics Communications, 35(1–3):C–849, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829586.

Omura:1988:EEE

[OM88] Yoshiharu Omura and Hiroshi Matsumoto. Electromagnetic and electrostatic emissions from a thin electron beam in space plasma. Computer Physics Communications, 49(1):133–142, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588902214.

ONeal:1989:OES

[O’N89] S. W. O’Neal. The OPAL event server. Computer Physics Communications, 57(1–3):413–416, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902531.

Ohsaki:1989:ECS

[ONP89] Akihiko Ohsaki, Hiroki Nakamura, and Seung C. Park. On the evaluation of cross section and rate constant of atomdiatom reactions in the sudden and adiabatic approximations. Computer Physics Communications, 52(3):291–302, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901045.

Opik:1987:PSS

[Öpi87] U. Öpik. A program to set up systems of orthogonal polynomials. Computer Physics Communications, 46(2):263–296, August 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900051.

Ohara:1984:TCI

[OS84] H. O’hara and F. J. Smith. Transport collision integrals for a dilute gas. Computer Physics Communications, 35(1–3):C– 64, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584823292.

Olivares:1986:CSO

[OS86] Antonio G. Olivares and Mario Soler. Computed sawtooth oscillations for tokamak plasmas. Computer Physics Communications, 40(2–3):219–227, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901104.

Okazaki:1986:NSL

[OSF86] T. Okazaki, M. Sugihara, and N. Fujisawa. Numerical study of a lower hybrid current drive in the presence of a dc electric field. Computer Physics Communications, 40(1):131–136, May 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901554.

Osland:1988:GCG

[Osl88] C. D. Osland. GKS and CGM graphics standards. Computer Physics Communications, 50(1–2):129–141, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558890121X.

Ostlund:1985:WVH

[Ost85] Neil S. Ostlund. WATERLOPP V2/64: a highly parallel machine for numerical computation. Computer Physics Communications, 37(1–3):109–117, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901420.

Olszewski:1980:PPV

[OT80] E. A. Olszewski and W. J. Thompson. A program for perspective views of three-dimensional surfaces. Computer Physics Communications, 21(2):185–193, December 2, 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900399.

Olszewski:1984:PPVa

[OT84] E. A. Olszewski and W. J. Thompson. A program for perspective views of three-dimensional surfaces. Computer Physics Communications, 35(1–3):C–658–C–659, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828179.

Onda:1980:NVP

[OTB80] Kunizo Onda, Donald G. Truhlar, and Maynard A. Brandt. New version of program for calculating differential and integral cross sections for quantum mechanical scattering problems from reactance or transition matrices. Computer Physics Communications, 21(1):97–108, December 1, 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558090079X.

Onda:1984:NVP

[OTB84] Kunizo Onda, Donald G. Truhlar, and Maynard A. Brandt. New version of program for calculating differential and integral cross sections for quantum mechanical scattering problems from reactance or transition matrices. Computer Physics Communications, 35(1–3):C–653, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828131.

Otto:1989:SMV

[Ott89] Steve W. Otto. Shared-memory versus distributed-memory: Halftime score. Computer Physics Communications, 57(1– 3):95–100, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901963.

Ortega:1984:INR

[OVQT^{+}84] M. Ortega, A. Vidal-Quadras, M. Tomás, F. Fernández, V. Gandía, and C. Jacquot. Identification of nuclear reactions registered in ionographic detectors. Computer Physics Communications, 35(1–3):C–381, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825941.

Opik:1984:MPE

[ÖW84] U. Öpik and R. F. Wood. The Madelung potential and electric field intensity for a sodium chloride and for a caesium chloride lattice. Computer Physics Communications, 35(1–3):C–33, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823012.

Ogata:1980:NCN

[OY80] Atsushi Ogata and Kazuo Yuasa. The number of computers in a non-hierarchical system for Tokamak data processing. Computer Physics Communications, 19(1):35–41, January/March 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900648.

Oyanagi:1986:IDL

[Oya86] Yoshio Oyanagi. An incomplete LDU decomposition of lattice fermions and its application to conjugate residual methods. Computer Physics Communications, 42(3):333–343, November 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900032.

Pagiola:1987:FVA

[Pag87] E. Pagiola. FORTRAN vs? answer; casual. Computer Physics Communications, 45(1–3):485, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901949.

Parkinson:1982:UID

[Par82]  Dennis Parkinson.  Using the ICL DAP.  Computer

Physics Communications, 26(3–4):227–232, June 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901096.

Parkinson:1983:DAP

[Par83] D. Parkinson. The Distributed Array Processor (DAP). Computer Physics Communications, 28(4):325–336, February 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900310.

Parcell:1984:SDE

[Par84] L. A. Parcell. Solution of differential equations for exchange matrix elements in heavy particle collisions. Computer Physics Communications, 35(1–3):C–181, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824273.

Parga:1989:NNM

[Par89a] N. Parga. Neural networks as models of associative memories. Computer Physics Communications, 55(1):77–84, August 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900659.

Parsons:1989:IEL

[Par89b] I. D. Parsons. The implementation of an element level multigrid algorithm on the alliant FX/8. Computer Physics Communications, 53(1–3):337–348, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901719.

Patterson:1982:LSS

[Pat82] G. S. Patterson, Jr. Large scale scientific computing — future directions. Computer Physics Communications, 26(3–4):217– 225, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901084.

Parkinson:1984:PNM

[PB84a] J. R. Parkinson and D. T. Birtwistle. A program for normalised Morse functions. Computer Physics Communications, 35(1–3): C–158, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824091.

Pirala:1984:RNI

[PB84b] P. Piralä and E. Byckling. Recursive numerical integration of multi-particle phase space with peripheral matrix element. Computer Physics Communications, 35(1–3):C–148, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824017.

Pawley:1985:CPM

[PBKW85] G. S. Pawley, K. C. Bowler, R. D. Kenway, and D. J. Wallace. Concurrency and parallelism in MC and MD simulations in physics. Computer Physics Communications, 37(1–3):251– 260, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585901602.

Papageorgiou:1989:MEP

[PCL89] D. G. Papageorgiou, C. S. Chassapis, and I. E. Lagaris. MERLIN-2.0 — enhanced and programmable version. Computer Physics Communications, 52(2):241–247, January/ February 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900076.

Perrott:1985:IPL

[PCM85] R. H. Perrott, D. Crookes, and P. Milligan. Implementing a parallel language on the Cray-1. Computer Physics Communications, 37(1–3):119–124, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901432.

Park:1989:BBT

[PD89] S. C. Park and J. P. Draayer. Balanced binary tree code for scientific applications. Computer Physics Communications, 55(2):189–204, September 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900763.

Pattnaik:1982:FSP

[PDF82] Pratap C. Pattnaik, Philip H. Dickinson, and John L. Fry. Fermi-surface: a package to display perspective drawings of Fermi surfaces in cubic systems. Computer Physics Communications, 25(1):63–71, January 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900443.

Pattnaik:1984:FSP

[PDF84] Pratap C. Pattnaik, Philip H. Dickinson, and John L. Fry. Fermi-surface: a package to display perspective drawings of Fermi surfaces in cubic systems. Computer Physics Communications, 35(1–3):C–768, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828969.

Pacheco:1988:FCS

[PDNE88] A. J. Pacheco, D. E. Di Gregorio, J. O. Fernández Niello, and M. Elgue. Fusion cross sections from measurements of delayed X-rays. Computer Physics Communications, 52(1):93–102, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901762.

Perrott:1982:LVP

[Per82] R. H. Perrott. Languages for vector and parallel processors. Computer Physics Communications, 26(3–4):267–275, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901163.

Perotto:1987:NOS

[Per87] E. Perotto. Network operating system. Computer Physics Communications, 45(1–3):455–466, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901895.

Persson:1989:MTD

[Per89] Stefan Persson. Measurement and three dimensional reconstruction of particle tracks in emulsion chambers. Computer Physics Communications, 55(1):103–108, August 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900684.

Petrasch:1984:NCA

[Pet84] P. Petrasch. Nonlinear computation of anisotropic elastic fields about straight edge dislocations. Computer Physics Communications, 35(1–3):C–380, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482593X.

Paul:1986:BLD

[PF86] H.-J. Paul and H. Freiesleben. The Bochum on-line data acquisition system. Computer Physics Communications, 39(3):305– 311, April 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900925.

Piska:1982:MSH

[PFGP82] K. Píska, W. Falkenberg, C.-P. Glasneck, and W. Pflugbeil. Multimicroprocessor system for high-energy physics experiment applications. Computer Physics Communications, 26(1–2):213– 215, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901722.

Pyper:1984:NPC

[PGB84] N. C. Pyper, I. P. Grant, and N. Beatham. A new program for calculating matrix elements of one-particle operators in jjcoupling. Computer Physics Communications, 35(1–3):C–509– C–510, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584826946.

Phillips:1980:SOD

[Phi80] G. M. Phillips. A survey of one-dimensional and multidimensional numerical integration. Computer Physics Communications, 20(1):17–27, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580901022.

Philippot:1986:SAL

[Phi86] Georg P. Philippot. The Simula approach to living software. Computer Physics Communications, 41(2–3):291–294, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558690072X.

Poenaru:1984:LDM

[PI84] D. N. Poenaru and M. Ivascu. Liquid drop model deformation energies of nuclei with axial symmetry and reflection asymmetry. Computer Physics Communications, 35(1–3):C–522, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827034.

Piacentini:1984:MMTa

[Pia84] R. D. Piacentini. Multistate molecular treatment of atomic collisions in the impact parameter approximation. III — integration of coupled equations and calculation of transition amplitudes for Coulomb trajectories. Computer Physics Communications, 35(1–3):C–397, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826089.

Pichon:1989:NCG

[Pic89] Bernard Pichon. Numerical calculation of the generalized Fermi–Dirac integrals. Computer Physics Communications, 55(2):127–136, September 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900714.

Piessens:1982:ACB

[Pie82] R. Piessens. Automatic computation of Bessel function integrals. Computer Physics Communications, 25(3):289–295, March 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900248.

Piessens:1984:ACB

[Pie84a] R. Piessens. Automatic computation of Bessel function integrals. Computer Physics Communications, 35(1–3):C–791, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482915X.

Piessens:1984:CRD

[Pie84b] R. Piessens. Calculation of the radial distribution of emitters in a cylindrical source. Computer Physics Communications, 35(1–3):C–182, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824285.

Poenaru:1980:FYP

[PIM80] D. N. Poenaru, M. Ivascu, and D. Mazilu. Folded Yukawaplus-exponential model PES for nuclei with different charge densities. Computer Physics Communications, 19(2):205–214, April/June 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558090051X.

Poenaru:1982:ADH

[PIM82] D. N. Poenaru, M. Ivascu, and D. Mazilu. Alpha-decay half-life semiempirical relationships with self-improving parameters. Computer Physics Communications, 25(3):297–309, March 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558290025X.

Poenaru:1984:ADH

[PIM84a] D. N. Poenaru, M. Ivascu, and D. Mazilu. Alpha-decay halflife semiempirical relationships with self-improving parameters. Computer Physics Communications, 35(1–3):C–792, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829161.

Poenaru:1984:FYP

[PIM84b] D. N. Poenaru, M. Ivascu, and D. Mazilu. Folded Yukawa-plus-exponential model PES for nuclei with different charge densities. Computer Physics Communications, 35(1–3):C–617, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827824.

Pissanetzky:1984:ATD

[Pis84a] Sergio Pissanetzky. Automatic three-dimensional finite element mesh generation using the program kubik. Computer Physics Communications, 32(3):245–265, June 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900894.

Pissanetzky:1984:KTD

[Pis84b] Sergio Pissanetzky. Kubik: a three-dimensional finite element mesh generator users manual. Computer Physics Communications, 32(3):267–279, June 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900900.

Pantelic:1989:NKS

[PJ89] Dejan V. Pantelic and Zoran D. Janevski. A new kind of splines and their use for fast ray-tracing in reflective cavities. Computer Physics Communications, 55(1):5–11, August 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990060X.

Platz:1986:PMD

[Pla86] Jochen Platz. Project management in the development of scientific software. Computer Physics Communications, 41 (2–3):217–225, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900664.

Peterson:1980:MCC

[PM80] Robert R. Peterson and Gregory A. Moses. MFP — a code for calculating equation of state and optical data for noble gases. Computer Physics Communications, 20(3):353–371, November 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900168.

Peterson:1983:MES

[PM83] Robert R. Peterson and Gregory A. Moses. MIXERG — an equation of state and opacity computer code. Computer Physics Communications, 28(4):405–426, February 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900358.

Peterson:1984:MCC

[PM84a] Robert R. Peterson and Gregory A. Moses. MFP — a code for calculating equation of state and optical data for noble gases. Computer Physics Communications, 35(1–3):C–641–C– 642, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828027.

Peterson:1984:MES

[PM84b] Robert R. Peterson and Gregory A. Moses. Mixerg — an equation of state and opacity computer code. Computer Physics Communications, 35(1–3):C–860–C–861, ????

1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829689.

Parlett:1989:TBB

[PNO89] Beresford N. Parlett and Bahram Nour-Omid. Towards a black box Lanczos program. Computer Physics Communications, 53(1–3):169–179, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL https://www.

math.utah.edu/pub/bibnet/subjects/acc-stab-num-alg.
bib;   http://www.sciencedirect.com/science/article/
pii/0010465589901586.

Poet:1985:DSP

[Poe85] Ron Poet. The design of special purpose hardware to factor large integers. Computer Physics Communications, 37 (1–3):337–341, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901705.

Pohl:1987:DDP

[Poh87a] Martin Pohl. Data driven parallelism in experimental high energy physics applications. Computer Physics Communications, 45(1–3):311–317, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587901676.

Pohl:1987:MHE

[Poh87b] Martin Pohl. Multiprocessors for high energy physics. Computer Physics Communications, 45(1–3):47–60, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901391.

Pollock:1988:PCC

[Pol88] E. L. Pollock. Properties and computation of the Coulomb pair density matrix. Computer Physics Communications, 52(1):49–60, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901713.

Ponting:1981:SVP

[Pon81] D. K. Ponting. Stochastic variational principle calculations. Computer Physics Communications, 22(2–3):159–166, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558190045X.

Poon:1983:ABS

[Poo83] Ying Ming Poon. Accurate basis sets for atomic configuration interaction calculations. Computer Physics Communications, 29(2):113–116, April 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900668.

Pouquet:1988:NSM

[Pou88a] A. Pouquet. Numerical simulations of MHD turbulence. Computer Physics Communications, 49(1):97–102, April 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902184.

Pouzin:1988:INI

[Pou88b] Louis Pouzin. Impact of national and international networks on the use of supercomputer centers. Computer Physics Communications, 49(1):201–203, April 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902275.

Pryce:1982:SOS

[PP82] J. D. Pryce and J. W. Paine. A stiff ODE solver for an attached processor. Computer Physics Communications, 27(1): 97–100, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900133.

Pilipczuk:1985:SPC

[PP85] Edward Pilipczuk and Iwona Pilipczuk. Skew — program for calculation of the electron scattering amplitudes on an atomic potential using spin-orbit relativistic correction. Computer Physics Communications, 36(1):101–107, March 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900207.

Phillips:1986:NAP

[PPHS86] C. K. Phillips, F. W. Perkins, D. Q. Hwang, and D. G. Swanson. Numerical application of the parabolic approximation method for fast magnetosonic wave propagation in tokamaks. Computer Physics Communications, 40(1):23–31, May 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901463.

Pauli:1984:CPI

[PR84] H. C. Pauli and U. Raff. A computer program for internal conversion coefficients and particle parameters. Computer Physics Communications, 35(1–3):C–320–C–321, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482545X.

Papageorgiou:1987:MSS

[PR87] C. D. Papageorgiou and A. D. Raptis. A method for the solution of the Schrödinger equation. Computer Physics Communications, 43(3):325–328, February/March 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558790049X.

Preuss:1984:CCO

[Pre84a] E. Preuss. Calculation of crystal orientations using Laue patterns. Computer Physics Communications, 35(1–3):C–594, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827642.

Preuss:1984:PPL

[Pre84b] E. Preuss. Plot program for Laue patterns and stereographic projections. Computer Physics Communications, 35(1–3):C– 593, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827630.

Provencher:1982:CRM

[Pro82a] Stephen W. Provencher. A constrained regularization method for inverting data represented by linear algebraic or integral equations. Computer Physics Communications, 27(3):213–227, September 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901734.

Provencher:1982:CGP

[Pro82b] Stephen W. Provencher. CONTIN: a general purpose constrained regularization program for inverting noisy linear algebraic and integral equations. Computer Physics Communications, 27(3):229–242, September 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901746.

Provencher:1984:CGP

[Pro84] Stephen W. Provencher. Contin: a general purpose constrained regularization program for inverting noisy linear algebraic and integral equations. Computer Physics Communications, 35(1–3):C–818–C–819, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829355.

Papaconstantopoulos:1984:CCP

[PS84a] Dimitrios A. Papaconstantopoulos and Wayne R. Slaughter. Calculation of crystal potentials. Computer Physics Communications, 35(1–3):C–238, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824777.

Piacentini:1984:MMTb

[PS84b] R. D. Piacentini and A. Salin. Multistate molecular treatment of atomic collisions in the impact parameter approximation. II — calculation of differential cross-sections from the transition amplitudes for the straight line case. Computer Physics Communications, 35(1–3):C–414, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826235.

Pomponiu:1984:FAS

[PS84c] C. Pomponiu and M. Sararu. Fourier analysis with splines, a Fortran program. Computer Physics Communications, 35(1–3): C–523, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827046.

Popov:1988:MOM

[PS88] V. P. Popov and A. L. Semjenov. Microprocessors and operative metrological maintenance of high-accuracy data conversion in the experiment. Computer Physics Communications, 50(1–2):257–261, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901348.

Peterson:1984:EED

[PSBG84] L. R. Peterson, T. Sawada, J. N. Bass, and A. E. S. Green. Electron energy deposition in a gaseous mixture. Computer Physics Communications, 35(1–3):C–178, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482425X.

Paez:1988:LNE

[PSL88] Manuel J. Páez, Milton E. Sagen, and Rubin H. Landau. LPOTp: Nucleon elastic-scattering from spin 0 and 12 nuclei in momentum space. Computer Physics Communications, 52(1):141–159, December 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901798.

Pick:1982:ETA

[PT82] S. Pick and M. Tomásek. On the equivalence of three approaches to matrix tridiagonalization: Comparison with the recursion method. Computer Physics Communications, 27 (1):101–102, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900145.

Pendry:1984:CPS

[PT84] J. B. Pendry and D. J. Titterington. Calculation of photoemission spectra for surfaces of solids. Computer Physics Communications, 35(1–3):C–610, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827770.

Potenza:1985:CPP

[PT85] R. Potenza and C. Tuvè. A colloquial plotting package to realize scientific diagrams. Computer Physics Communications, 38(1):53–60, August/September 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900451.

Phillips:1986:NSI

[PT86] M. W. Phillips and A. M. M. Todd. The numerical solution of ICRF fields in axisymmetric mirrors. Computer Physics Communications, 40(1):65–72, May 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901487.

Puff:1983:PNV

[Puf83] Werner Puff. PFPOSFIT: a new version of a program for analysing positron lifetime spectra with non-Gaussian prompt curve. Computer Physics Communications, 30(4):359–368, December 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900772.

Puff:1984:PNV

[Puf84] Werner Puff. Pfposfit: a new version of a program for analysing positron lifetime spectra with non-Gaussian prompt curve. Computer Physics Communications, 35(1–3):C–922, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584830209.

Purcell:1982:UC

[Pur82] C. J. Purcell. Using the Cyber 205. Computer Physics Communications, 26(3–4):249–251, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465582901138.

Putzer:1989:DSD

[Put89] Alois Putzer. Data structures and data-base systems used in high energy physics: Modelling and implementation. Computer Physics Communications, 57(1–3):156–163, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902051.

Payne:1984:FPC

[PV84] G. L. Payne and P. L. Von Behren. Fortran program to calculate finite-range no-recoil DWBA transfer cross sections. Computer Physics Communications, 35(1–3):C–223–C–224, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824637.

Pyle:1985:PMA

[Pyl85] I. C. Pyle. Pascal, Modula and Ada. Computer Physics Communications, 38(2):191–197, October/November 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900852.

Ping:1989:CGS

[PZCC89] Jia-Lun Ping, Qing-Rong Zheng, Bing-Qing Chen, and Jin-Quan Chen. Computer generated subgroup-symmetry adapted irreducible representations and CG coefficients of space groups by the eigenfunction method. Computer Physics Communications, 52(3):355–373, March 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901112.

Quarrie:1989:COS

[QAD^{+}89] D. R. Quarrie, M. D. Anderson, C. T. Day, G. Goeransson, J. Patrick, M. Schmitz, E. Sexton, and B. Troemel. The CDF online system. Computer Physics Communications, 57 (1–3):325–331, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589902373.

Qian:1987:UAD

[QTZ^{+}87] Z. Qian, S. Tang, W. Zhao, S. M. Fisher, J. Harvey, M. Boano, J. Bunn, R. McClatchey, V. Emiliani, P. Palazzi, R. Brazioli, A. Putzer, M. G. Green, and R. Fantechi. Use of the ADAMO data management system within ALEPH. Computer Physics Communications, 45(1–3):283–298, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901652.

Quarrie:1987:PCH

[Qua87] David R. Quarrie. Personal computers in high energy physics. Computer Physics Communications, 45(1–3):175–179, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901536.

Robaux:1982:PAR

[RA82] O. Robaux and M. Aymar. A program for analysing the Rydberg series of highly excited discrete spectra by M.Q.D.T. Computer Physics Communications, 25(3):223–236, March 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900194.

Robaux:1984:PAR

[RA84] O. Robaux and M. Aymar. A program for analysing the Rydberg series of highly excited discrete spectra by M.Q.D.T. Computer Physics Communications, 35(1–3):C–784, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829100.

Ramek:1985:TLG

[Ram85] Michael Ramek. Titel — a letter generating program. Computer Physics Communications, 36(4):433–440, June 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900335.

Rand:1986:PPI

[Ran86] D. W. Rand. PASCAL programs for identification of Lie algebras: Part 1. Radical — a program to calculate the radical and nil radical of parameter-free and parameter-dependent Lie algebras. Computer Physics Communications, 41(1):105–125, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558690024X. See erratum [Ran87a].

Rand:1987:ENP

[Ran87a] D. W. Rand. Erratum notice: “PASCAL programs for identification of Lie algebras. I. RADICAL — a program to calculate the radical and nil radical of parameter-free and parameter-dependent Lie algebras” [Comput. Phys. Comm. 41 (1986), no. 1, 105–125]. Computer Physics Communications, 47(2–3):369–371, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901251. See [Ran86].

Rand:1987:PPIb

[Ran87b] D. W. Rand. PASCAL programs for identification of Lie algebras: Part 3: LEVI decomposition and canonical basis. Computer Physics Communications, 46(2):311–322, August 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900075.

Rao:1981:CAM

[Rao81] K. Srinivasa Rao. Computation of angular momentum coefficients using sets of generalised hypergeometric functions. Computer Physics Communications, 22(2–3):297–302, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900631.

Rappaz:1981:SPN

[Rap81a] J. Rappaz. Spectral pollution of a non-compact operator. Computer Physics Communications, 24(3–4):323–327, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901545.

Raptis:1981:NSS

[Rap81b] A. D. Raptis. On the numerical solution of the Schrödinger equation. Computer Physics Communications, 24(1):1–4, September/October 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581901016.

Raptis:1983:EFS

[Rap83] A. D. Raptis. Exponentially-fitted solutions of the eigenvalue Schrödinger equation with automatic error control. Computer Physics Communications, 28(4):427–431, February 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558390036X.

Rapaport:1984:CGC

[Rap84] D. C. Rapaport. Computer generation of connected graphs. Computer Physics Communications, 35(1–3):C–280, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825114.

Rapaport:1985:MBS

[Rap85] D. C. Rapaport. Many-body simulations using an array processor. Computer Physics Communications, 37(1–3):343–349, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901717.

Raseev:1980:ESC

[Ras80a] G. Raseev. Electron scattering by closed or open shell diatomic molecules. Computer Physics Communications, 20(2):275–289, October 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900077.

Raseev:1980:TVP

[Ras80b] G. Raseev. Third version of a program for calculating the static interaction potential between an electron and a diatomic molecule. Computer Physics Communications, 20(2):267–274, October 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900065.

Raseev:1984:ESC

[Ras84a] G. Raseev. Electron scattering by closed or open shell diatomic molecules. Computer Physics Communications, 35(1–3): C–636, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827976.

Raseev:1984:TVP

[Ras84b] G. Raseev. Third version of a program for calculating the static interaction potential between an electron and a diatomic molecule. Computer Physics Communications, 35(1–3):C–635, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827964.

Rudnicki-Bujnowski:1984:EFC

[RB84] Georges Rudnicki-Bujnowski. Explicit formulas for CLEBSCH-GORDAN coefficients. Computer Physics Communications, 35(1–3):C–343, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825643.

Rountree:1984:PCG

[RBHW84] S. P. Rountree, T. Burnett, R. J. W. Henry, and C. A. Weatherford. A program to calculate Green’s functions. Computer Physics Communications, 35(1–3):C–357–C–358, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825771.

Rankin:1986:DEF

[RBMC86] R. Rankin, A. Birnboim, R. Marchand, and C. E. Capjack. Diffusion and equilibration in 2D fluid codes. Computer Physics Communications, 41(1):21–34, July 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900184.

Raptis:1985:VSM

[RC85] A. D. Raptis and J. R. Cash. A variable step method for the numerical integration of the one-dimensional Schrödinger equation. Computer Physics Communications, 36(2):113–119, April 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901171.

Raptis:1987:EBF

[RC87] A. D. Raptis and J. R. Cash. Exponential and Bessel fitting methods for the numerical solution of the Schrödinger equation. Computer Physics Communications, 44(1–2):95–103, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900208.

Rysavy:1980:CPD

[RD80] M. Rysavý and O. Dragoun. A computer program for determination of nuclear parameters from initial conversion experiments. Computer Physics Communications, 19(1):93–101, January/March 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900697.

Rysavy:1984:CPD

[RD84] M. Rysavý and O. Dragoun. A computer program for determination of nuclear parameters from initial conversion experiments. Computer Physics Communications, 35(1–3):C–611, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827782.

Read:1984:GEW

[Rea84] Brian J. Read. G-exec and the world data centre. Computer Physics Communications, 33(1–3):235–244, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901292.

Read:1989:DSO

[Rea89] Brian J. Read. Data structures and organisation: Special problems in scientific applications. Computer Physics Communications, 57(1–3):164–168, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902063.

Redi:1988:STC

[Red88] Martha H. Redi. Standard test cases for the BALDUR transport code. Computer Physics Communications, 49(2):399– 407, May 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588900136.

Regler:1981:ICA

[Reg81] M. Regler. Influence of computation algorithms on experimental design. Computer Physics Communications, 22(2–3):167– 175, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900461.

Reid:1985:MXR

[Rei85] John S. Reid. Multiphonon X-ray scattering. Computer Physics Communications, 38(1):43–52, August/September 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558590044X.

Reid:1986:SPX

[Rei86] John S. Reid. Single-phonon X-ray scattering. Computer Physics Communications, 42(3):417–426, November 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900093.

Reid:1987:HDW

[Rei87] John S. Reid. Harmonic Debye–Waller factors for cubic materials. Computer Physics Communications, 46(1):141–148, July 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900415.

Reid:1989:SRF

[Rei89] John S. Reid. Synchrotron radiation flux at experimental stations. Computer Physics Communications, 54(2–3):307–314, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900933.

Rysavy:1983:EPD

[RF83] M. Rysavý and M. Fiser. Erika — a program for the decomposition of line spectra. Computer Physics Communications, 29(2):171–183, April 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900723.

Rysavy:1984:EPD

[RF84] M. Rysavý and M. Fiser. Erika — a program for the decomposition of line spectra. Computer Physics Communications, 35(1–3):C–873, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482977X.

Raine:1989:SLM

[RFS89] A. R. C. Raine, D. Fincham, and W. Smith. Systolic loop methods for molecular dynamics simulation using multiple transputers. Computer Physics Communications, 55(1):13–30, August 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900611.

Radziemski:1984:CAE

[RFSG84a] Leon J. Radziemski, Jr., Kay J. Fisher, David W. Steinhaus, and Aaron S. Goldman. Calculation of atomic energy level values. Computer Physics Communications, 35(1–3):C–116, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823747.

Radziemski:1984:WNC

[RFSG84b] Leon J. Radziemski, Jr., Kay J. Fisher, David W. Steinhaus, and Aaron S. Goldman. Wave number calculation from least-squares level values. Computer Physics Communications, 35(1–3):C–117, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823759.

Rockwood:1980:NSB

[RG80] S. D. Rockwood and A. E. Greene. Numerical solutions of the Boltzmann transport equation. Computer Physics Communications, 19(3):377–393, July/August 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900909.

Rockwood:1984:NSB

[RG84] S. D. Rockwood and A. E. Greene. Numerical solutions of the Boltzmann transport equation. Computer Physics Communications, 35(1–3):C–627, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827903.

Reiman:1986:CTD

[RG86] A. Reiman and H. Greenside. Calculation of three-dimensional MHD equilibria with islands and stochastic regions. Computer Physics Communications, 43(1):157–167, December 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900597.

Rushton:1989:DIO

[RHM^{+}89] A. M. Rushton, L. Hunt, T. McGlynn, F. Ochsenbein, B. Perrine, A. Richmond, F. Romelfanger, G. Russo, P. M. B. Shames, J. Travisano, L. Willard, and S. Zeller. Design and implementation of an optical disk-based astronomical data archive. Computer Physics Communications, 57(1–3):427–431, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902567.

Ribeiro:1980:MOS

[Rib80] C. A. Ribeiro. Method for obtaining small sets of pseudo random numbers with uniform distribution. Computer Physics Communications, 19(3):305–307, July/August 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900855.

Richardson:1982:SMC

[Ric82] D. D. Richardson. Shell model calculations of point defect formation energies in cubic ionic crystals. Computer Physics Communications, 28(1):75–101, November 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900649.

Richardson:1984:SMC

[Ric84] D. D. Richardson. Shell model calculations of point defect formation energies in cubic ionic crystals. Computer Physics Communications, 35(1–3):C–837, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829483.

Richmond:1986:SDO

[Ric86] Alan Richmond. Software design by object-oriented functional layering. Computer Physics Communications, 41(2–3):377–384, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900767.

Ridder:1982:CAT

[Rid82] Detlef Ridder. Classification of Auger-transitions in LScoupling. Computer Physics Communications, 28(2):201–205, December 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900558.

Ridder:1984:DIC

[Rid84a] Detlef Ridder. Determination of interacting configurations. Computer Physics Communications, 31(4):423–429, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900262.

Ridder:1984:CAT

[Rid84b] Detlet Ridder. Classification of Auger-transitions in LScoupling. Computer Physics Communications, 35(1–3):C–846, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829549.

Rimmer:1981:SDP

[Rim81] E. M. Rimmer. Standards for data processing in high energy and nuclear physics experiments. Computer Physics Communications, 22(2–3):303–305, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900643.

Rinker:1984:SDM

[Rin84] G. A. Rinker. Static and dynamic muonic-atom codes — muon and rupp. Computer Physics Communications, 35(1–3):C–534– C–535, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827137.

Ritter:1984:MCCa

[Rit84a] S. Ritter. Monte Carlo code bamjet to simulate the fragmentation of quark and diquark jets. Computer Physics Communications, 31(4):393–400, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900225.

Ritter:1984:MCCb

[Rit84b] S. Ritter. Monte Carlo code parjet to simulate e^{+}e^{−}- annihilation events via QCD jets. Computer Physics Communications, 31(4):401–409, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900237.

Ritter:1986:MCM

[Rit86] S. Ritter. The macroparticle code macpar to simulate the beambeam interaction of high energy linear electron-positron colliders. Computer Physics Communications, 39(1):71–91, January 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901645.

Reich:1980:GLA

[RL80] G. R. Reich and P. L. Leath. Growing lattice animals and Monte–Carlo methods. Computer Physics Communications, 19(1):29–33, January/March 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900636.

Reid:1989:QEX

[RM89] John S. Reid and Gordon J. Milne. Quasi-elastic Xray scattering divergence analysis calculation. Computer Physics Communications, 55(1):91–101, August 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900672.

Rosato:1989:MDM

[RMV89] V. Rosato, G. Maino, and A. Ventura. A molecular dynamics model for the study of helium on transition metals. Computer Physics Communications, 54(2–3):251–256, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990088X.

Roberts:1980:E

[Rob80a]  K. V. Roberts.  Editorial.  Computer Physics Com-

munications, 19(1):vii–viii, January/March 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900582.

Roberts:1980:CCL

[Rob80b] S. A. Roberts. COLASE — a CO–N_{2}–He laser kinetics code. Computer Physics Communications, 20(3):373–412, November 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558090017X.

Roberts:1981:E

[Rob81] K. V. Roberts. Editorial. Computer Physics Communications, 23(1):vii, June 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581901235.

Roberts:1983:RFS

[Rob83] K. V. Roberts. The regeneration of Fortran software. Computer Physics Communications, 29(1):7–13, March 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900243.

Robb:1984:PER

[Rob84a] W. D. Robb. A program to evaluate the reduced matrix elements of summations of one-particle tensor operators. Computer Physics Communications, 35(1–3):C–214, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482456X.

Robb:1984:PGN

[Rob84b] W. D. Robb. A program to generate numerical orbital functions. Computer Physics Communications, 35(1–3):C–54, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482319X.

Roberts:1984:CCL

[Rob84c] S. A. Roberts. Colase — a CO-N_{2}-He laser kinetics code. Computer Physics Communications, 35(1–3):C–643, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828039.

Roberts:1984:NMCa

[Rob84d] S. A. Roberts. Numerical modelling of a chemical plasma I. reacs: a program to generate all reactions which take place in a plasma of given chemical content. Computer Physics Communications, 35(1–3):C–599, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482768X.

Roberts:1984:NMCb

[Rob84e] S. A. Roberts. Numerical modelling of a chemical plasma II. plaskem: a program to predict the variation with time of the number densities of chemical species within a plasma. Computer Physics Communications, 35(1–3):C–600, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827691.

Roberts:1984:NMCc

[Rob84f] S. A. Roberts. Numerical modelling of a chemical plasma III. datstor: a program to create a data base containing information on rate coefficients of chemical reactions. Computer Physics Communications, 35(1–3):C–601, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827708.

Roberts:1985:BRB

[Rob85] K. V. Roberts. Book review: Programming for Software Sharing: D. T. Muxworthy, ed., D. Reidel, Dordrecht, 1983. x + 283 pages Cloth Dfl 90, US$39.50 ISBN 90-277-1591-2. Computer Physics Communications, 34(3):333–334, January 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900104.

Roberts:1987:EFC

[Rob87a] Lee Roberts. Evaluation of the FPS-164 computer for high energy physics pattern recognition problems. Computer Physics Communications, 47(2–3):195–206, November/ December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901068.

Robertson:1987:AAF

[Rob87b] J. A. Robertson. Axisymmetric accretion flows in astrophysics. Computer Physics Communications, 44(3):279–288, June 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558790083X.

Romaya:1981:FSI

[Rom81] A. M. Romaya. FORTRAN subset interpreter for a microprocessor gas control system. Computer Physics Communications, 22(2–3):307–309, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900655.

Rossiter:1984:DMP

[Ros84a] B. N. Rossiter. Data modelling and performance of data base systems. Computer Physics Communications, 33(1–3): 13–21, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584901024.

Rossiter:1984:IDB

[Ros84b] B. N. Rossiter. Introduction to data base management systems. Computer Physics Communications, 33(1–3):5–12, August/ September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901012.

Rosencrans:1985:CHO

[Ros85] S. I. Rosencrans. Computation of higher-order fluid symmetries using MACSYMA. Computer Physics Communications, 38(3):347–356, December 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558590102X.

Rous:1989:TLT

[RP89a] P. J. Rous and J. B. Pendry. Tensor LEED I: a technique for high speed surface structure determination by low energy electron diffraction. TLEED1. Computer Physics Communications, 54(1):137–156, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900398.

Rous:1989:TLI

[RP89b] P. J. Rous and J. B. Pendry. Tensor LEED II: a technique for high speed surface structure determination by low energy electron diffraction. TLEED2. Computer Physics Communications, 54(1):157–166, April 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900404.

Ranft:1984:MCP

[RR84] J. Ranft and J. T. Routti. Monte Carlo programs for calculating three-dimensional high-energy (50 MeV–500 GeV) hadron cascades in matter. Computer Physics Communications, 35(1–3): C–249, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824856.

Rao:1989:NFP

[RRC89] K. Srinivasa Rao, V. Rajeswari, and Charles B. Chiu. A new Fortran program for the 9-j angular momentum coefficient.

Computer Physics Communications, 56(2):231–248, December 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900210.

Romanowski:1988:DPE

[RRG88] Hubert Romanowski, Mark A. Ratner, and R. B. Gerber. Determination of potential energy surfaces of linear triatomics from vibration-rotation spectra: An inversion method applied to CO_{2}. Computer Physics Communications, 51(1–2):161–171, September/October 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900690.

Routti:1980:GPU

[RS80] J. T. Routti and J. V. Sandberg. General purpose unfolding program LOUHI78 with linear and nonlinear regularizations. Computer Physics Communications, 21(1):119–144, December 1, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900818.

Routti:1981:FPS

[RS81] J. T. Routti and J. V. Sandberg. Fortran program spall for computing spallation reaction cross sections. Computer Physics Communications, 23(4):411–426, August 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901818.

Ruden:1983:MCC

[RS83] W. D. Ruden and R. M. Schectman. A Monte Carlo calculation of the dissociation of fast H_{2}^{+} ions traversing thin carbon foils. Computer Physics Communications, 28(4):355–366, February 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900334.

Roberts:1984:PPP

[RS84a] S. A. Roberts and Kenneth Smith. Pulsamp: a program to predict the amplification of nanosecond CO_{2} laser light pulses. Computer Physics Communications, 35(1–3):C–407, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584826168.

Routti:1984:FPS

[RS84b] J. T. Routti and J. V. Sandberg. Fortran program spall for computing spallation reaction cross sections. Computer Physics Communications, 35(1–3):C–739, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828738.

Routti:1984:GPU

[RS84c] J. T. Routti and J. V. Sandberg. General purpose unfolding program LOUHI78 with linear and nonlinear regularizations. Computer Physics Communications, 35(1–3):C–655, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828155.

Ruden:1984:MCC

[RS84d] W. D. Ruden and R. M. Schectman. A Monte Carlo calculation of the dissociation of fast H_{2}^{+} ions traversing thin carbon foils. Computer Physics Communications, 35(1–3):C–857, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829665.

Rundgren:1984:LIC

[RS84e] J. Rundgren and A. Salwén. Leed intensity curves by the layer-by-layer method and perturbation calculation. Computer Physics Communications, 35(1–3):C–252, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824881.

Rundgren:1984:SCL

[RS84f] J. Rundgren and A. Salwén. Symmetrization and calculation of LEED intensity patterns. Computer Physics Communications, 35(1–3):C–313, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825394.

Rhoades:1985:EME

[RS85] Clifford E. Rhoades, Jr. and K. G. Stevens, Jr. Early MIMD experience on the CRAY X-MP. Computer Physics Communications, 37(1–3):215–221, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901559.

Rosel:1984:QMC

[RSA84] F. Rösel, J. X. Saladin, and K. Alder. Quantum mechanical coupled channels code for Coulomb excitation. Computer Physics Communications, 35(1–3):C–258–C–259, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824947.

Rouse:1984:PCG

[RSL^{+}84] Robert J. Rouse, Jr., Gordon L. Struble, Robert G. Lanier, Lloyd G. Mann, and Edward S. Macias. A program for calculating gamma-gamma directional correlation coefficients and angular distribution coefficients for gamma rays of mixed multipolarities from partially aligned nuclei. Computer Physics Communications, 35(1–3):C–487–C–488, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482679X.

Ramis:1988:MCC

[RSMTV88] R. Ramis, R. Schmalz, and J. Meyer-Ter-Vehn. MULTI — a computer code for one-dimensional multigroup radiation hydrodynamics. Computer Physics Communications, 49(3):475– 505, June 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588900082.

Robouch:1981:SPI

[RSP81] B. V. Robouch, A. Sestero, and S. Podda. Simulation of photographic images on a plotter. Computer Physics Communications, 24(1):63–71, September/October 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901065.

Robouch:1984:SPI

[RSP84] B. V. Robouch, A. Sestero, and S. Podda. Simulation of photographic images on a plotter. Computer Physics Communications, 35(1–3):C–744, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828775.

Reddaway:1985:VHS

[RSS85] S. F. Reddaway, D. M. Scott, and K. A. Smith. A very high speed Monte Carlo simulation on DAP. Computer Physics Communications, 37(1–3):351–356, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901729.

Rudge:1984:MVC

[Rud84] M. R. H. Rudge. MRVAC — a variational correction method for solving differential equations with r^{−n} coupling. Computer Physics Communications, 34(1–2):187–197, November/ December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558490170X.

Ruehle:1989:GNS

[Rue89] R. Ruehle. Graphics and networks at Stuttgart. Computer Physics Communications, 57(1–3):561, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902890.

Ruffhead:1985:PPA

[Ruf85] A. Ruffhead. Parallel processing applied to digital terrain matrices. Computer Physics Communications, 37(1–3):357–361, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901730.

Runge:1989:HSN

[Run89] K. Runge. A high speed network for HEP in germany. Computer Physics Communications, 57(1–3):452–454, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902610.

Russell:1986:IUD

[Rus86] J. E. Russell. Implementing and using the database management system. Computer Physics Communications, 41 (2–3):393–395, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900780.

Russell:1987:PLT

[Rus87] J. J. Russell. Programming languages: Time for a change? Computer Physics Communications, 45(1–3):269–273, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901639.

Rao:1984:NFP

[RV84] K. Srinivasa Rao and K. Venkatesh. New Fortran programs for angular momentum coefficients. Computer Physics Communications, 35(1–3):C–498–C–499, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826855.

Rosato:1981:WFO

[RVD^{+}81] E. Rosato, M. Vigilante, N. De Cesare, E. Perillo, and G. Spadaccini. Weighted fit by orthonormal polynomials. Computer Physics Communications, 22(2–3):311–315, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900667.

Rand:1984:NSP

[RW84] D. W. Rand and P. Winternitz. Nonlinear superposition principles: a new numerical method for solving matrix Riccati equations. Computer Physics Communications, 33(4):305–328, October 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558490136X.

Reid:1985:AFF

[RW85] J. K. Reid and A. Wilson. The array features in FORTRAN 8x with examples of their use. Computer Physics Communications, 37(1–3):125–132, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901444.

Rand:1986:OMP

[RW86] D. W. Rand and P. Winternitz. Odepainleve — a Macsyma package for Painlevé analysis of ordinary differential equations. Computer Physics Communications, 42(3):359–383, November 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900068.

Radehaus:1985:SPC

[RWK^{+}85] Ch. Radehaus, M. Waldowski, K. Kardell, J. Berkemeier, M. Wiesemann, and H. G. Purwins. Special purpose computer for nonlinear differential equations. Computer Physics Communications, 36(4):345–350, June 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900232.

Rand:1987:PPIa

[RWZ87] D. W. Rand, P. Winternitz, and H. Zassenhaus. PAS-CAL programs for identification of Lie algebras. Part 2: SPLIT — a program to decompose parameter-free and parameter-dependent Lie algebras into direct sums. Computer Physics Communications, 46(2):297–309, August 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900063.

Rynefors:1982:UPM

[Ryn82] K. Rynefors. UNIMOL: a program for Monte Carlo simulation of RRKM unimolecular decomposition in molecular beam experiments. Computer Physics Communications, 27(2):201– 212, August 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900741.

Rynefors:1984:UPM

[Ryn84] K. Rynefors. Unimol: a program for Monte Carlo simulation of RRKM unimolecular decomposition in molecular beam experiments. Computer Physics Communications, 35(1–3):C– 817, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829343.

Rysavy:1984:GGI

[Rys84] M. Rysavý. GISP — gauss-integration subroutine pack. Computer Physics Communications, 32(2):191–199, May 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900705.

Rysavy:1987:MSC

[Rys87] M. Rysavý. MISHA — a system for calculations with arbitrary arithmetic precision. Computer Physics Communications, 47(2–3):351–359, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901202.

Shestakov:1983:ISS

[SA83] A. I. Shestakov and D. V. Anderson. ILUCG2: Subprograms for the solution of a linear asymmetric matrix equation arising from a 9-point discretization. Computer Physics Communications, 30(1):31–36, July/August 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583901194.

Shestakov:1984:ISS

[SA84] A. I. Shestakov and D. V. Anderson. ILUCG2: Subprograms for the solution of a linear asymmetric matrix equation arising from a 9-point discretization. Computer Physics Communications, 35(1–3):C–891–C–892, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829951.

Saad:1989:NSL

[Saa89] Youcef Saad. Numerical solution of large nonsymmetric eigenvalue problems. Computer Physics Communications, 53(1–3): 71–90, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901495. Practical it-

erative methods for large scale computations (Minneapolis, MN, 1988).

Succi:1986:CMW

[SAC^{+}86] S. Succi, K. Appert, W. Core, H. Hamnén, T. Hellsten, and J. Vaclavik. Computational models for wave-particle interactions. Computer Physics Communications, 40(1):137–151, May 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901566.

Schatz:1987:CSD

[SAC87] George C. Schatz, B. Amaee, and J. N. L. Connor. The centrifugal sudden distorted wave method for calculating cross sections for chemical reactions: Angular distributions for Cl + HCl → ClH + Cl. Computer Physics Communications, 47(1):45–53, October 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900658.

Salvini:1982:NPC

[Sal82] Stefano A. Salvini. A new program to calculate differential and total cross sections for electron-atom or ion scattering using the momentum transfer formalism. Computer Physics Communications, 27(1):25–37, July 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900042.

Salustri:1983:PBS

[Sal83a] Carlo Salustri. Parametrization of the band structure of FCC crystals. Computer Physics Communications, 30(3):271–275, November 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900954.

Saly:1983:SBM

[Sal83b] Rudolf Saly. Soliton bag model. Computer Physics Communications, 30(4):411–419, December 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900826.

Salem:1984:CFF

[Sal84a] M. Salem. Converter of Fortran format and data statements to standard form. Computer Physics Communications, 35(1–3): C–371, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825862.

Salem:1984:IAS

[Sal84b] M. Salem. ILTHII — analysis of the spectrum of a thermal radioastronomical source. Computer Physics Communications, 35(1–3):C–305, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825333.

Salin:1984:CWF

[Sal84c] A. Salin. Calculation of wave-functions and collision matrix elements for one-electron diatomic molecules. Computer Physics Communications, 35(1–3):C–456–C–457, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826594.

Salustri:1984:PBS

[Sal84d] Carlo Salustri. Parametrization of the band structure of FCC crystals. Computer Physics Communications, 35(1–3):C–916, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584830155.

Salvini:1984:NPC

[Sal84e] Stefano A. Salvini. A new program to calculate differential and total cross sections for electron-atom or ion scattering using the momentum transfer formalism. Computer Physics Communications, 35(1–3):C–801, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829227.

Saly:1984:SBM

[Sal84f] Rudolf Saly. Soliton bag model. Computer Physics Communications, 35(1–3):C–927, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584830258.

Saly:1985:MCS

[Sal85] Rudolf Saly. Monte Carlo simulation of lattice Skyrme model. Computer Physics Communications, 36(4):417–422, June 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585900311.

Salvat:1987:ARS

[Sal87] Francesc Salvat. Algorithms for random sampling from singlevariate distributions. Computer Physics Communications, 46 (3):427–436, September 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900968.

Sameh:1985:SPA

[Sam85] A. Sameh. On some parallel algorithms on a ring of processors. Computer Physics Communications, 37(1–3):159–166, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585901481.

Sansoni:1982:HMC

[San82] A. Sansoni. Hevol2: a Monte Carlo program to calculate the evolution of structure functions with the inclusion of next to leading order effects. Computer Physics Communications, 27(4):403–414, October 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901011.

Sansoni:1984:HMC

[San84] A. Sansoni. Hevol2: a Monte Carlo program to calculate the evolution of structure functions with the inclusion of next to leading order effects. Computer Physics Communications, 35(1–3):C–826–C–827, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829410.

Saraph:1984:CSR

[Sar84a] H. E. Saraph. Collision strengths from reactance matrices. Computer Physics Communications, 35(1–3):C–29, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822973.

Saraph:1984:FSCa

[Sar84b] H. E. Saraph. Fine structure cross sections from reactance matrices. Computer Physics Communications, 35(1–3):C–139, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584823930.

Saraph:1984:FSCb

[Sar84c] H. E. Saraph. Fine structure cross sections from reactance matrices — a more versatile development of the program JA-JOM. Computer Physics Communications, 35(1–3):C–501, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826879.

Saraph:1987:POS

[Sar87] H. E. Saraph. PHOTUC: Oscillator strengths and photoionization cross sections from close coupling wave functions. Computer Physics Communications, 46(1):107–127, July 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900397.

Saulys:1985:BMS

[Sau85] Alfred C. Saulys. The BNL multiparticle spectrometer software. Computer Physics Communications, 38(2):319–322, October/ November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585900979.

Sauter:1987:NMC

[SAVV87] O. Sauter, K. Appert, L. Villard, and J. Vaclavik. Numerical modelling of the cold ion-ion hybrid resonance. Computer Physics Communications, 46(2):205–208, August 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900014.

Saxena:1984:ANVb

[Sax84a] K. M. S. Saxena. Adaptation of the new version of the reduced matrix elements (AAKP) program: Inclusion of the checking of the input data. Computer Physics Communications, 35(1–3): C–518, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827009.

Saxena:1984:ANVa

[Sax84b] K. M. S. Saxena. Adaptation of the new version of the reduced tensor matrix elements (AAKP) program: Inclusion of the evaluation of matrix elements of tensor products. Computer Physics Communications, 35(1–3):C–431, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826405.

Saxena:1984:NVA

[Sax84c] K. M. S. Saxena. A new version of AAKL (the matrix elements of spin-orbit interaction) adapted to spectroscopic notation. Computer Physics Communications, 35(1–3):C–424, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826338.

Scott:1982:CMA

[SB82] N. S. Scott and P. G. Burke. Calculation of R-matrix angular integrals on the Cray-1. Computer Physics Communications, 26(3–4):419–421, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901370.

Sjostrand:1987:LMC

[SB87] Torbjörn Sjöstrand and Mats Bengtsson. The Lund Monte Carlo for jet fragmentation and e ^{+}e^{−} physics — jetset version 6.3 — an update. Computer Physics Communications, 43 (3):367–379, February/March 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900543.

Sendall:1989:MSS

[SBB^{+}89] D. M. Sendall, C. Boissat, W. Bozzoli, P. Burkimsher, R. Jones, J-P. Matheys, G. Mornacchi, T. Nguyen, P. Vande Vyvre, A. Vascotto, and D. Weaver. Model: a software suite for data acquisition. Computer Physics Communications, 57(1–3):343–347, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902403.

Shoucri:1987:ASP

[SBFI87] M. Shoucri, P. Bertrand, M. R. Feix, and B. Izrar. The application of shape preserving splines for the modified Korteweg– de Vries equation. Computer Physics Communications, 46(1): 1–5, July 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900324.

Schnack:1986:NSR

[SBM^{+}86] D. D. Schnack, D. C. Barnes, Z. Mikic, D. S. Harned, E. J. Caramana, and R. A. Nebel. Numerical simulation of reversedfield pinch dynamics. Computer Physics Communications, 43(1):17–28, December 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900512.

Shipman:1984:HSE

[SC84] L. L. Shipman and R. E. Christoffersen. High speed evaluation of F_{0}(x). Computer Physics Communications, 35(1–3):C– 83, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584823450.

Schneider:1989:DIV

[SC89] B. I. Schneider and L. A. Collins. A direct iterative-variational method for solving large sets of linear algebraic equations. Computer Physics Communications, 53(1–3):381–392, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901744.

Sobottka:1988:FAI

[SCC^{+}88] S. E. Sobottka, R. J. Chandross, G. G. Cornick, B. D. Justice, R. S. Stewart, and J. A. Thomas. Fast algorithm for identifying clusters in digitized images. Computer Physics Communications, 51(3):295–299, November 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901427.

Seijo:1986:APA

[SCF86] Luis Seijo, Brant Coghlan, and Serafin Fraga. Association of proteins: Adaptation and coupling of two available programs. Computer Physics Communications, 41(1):169–177, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900275.

Schmalz:1981:HTE

[Sch81] R. Schmalz. High β and toroidal effects in three dimensions. Computer Physics Communications, 24(3–4):421–425, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901661.

Schrage:1982:ABA

[Sch82a] Martin H. Schrage. Architectural barriers to array processor efficiency and their analysis. Computer Physics Communications, 26(3–4):353–355, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558290128X.

Schwarz:1982:RPD

[Sch82b] Fritz Schwarz. A REDUCE package for determining Lie symmetries of ordinary and partial differential equations. Computer Physics Communications, 27(2):179–186, August 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900728.

Schmalz:1983:HTE

[Sch83] R. F. Schmalz. High-β and toroidal effects on the internal kink mode in tokamaks. Computer Physics Communications, 30(2):139–150, September/October 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900565.

Schorr:1984:PLV

[Sch84a] B. Schorr. Programs for the Landau and the Vavilov distributions and the corresponding random numbers. Computer Physics Communications, 35(1–3):C–239–C–240, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824789.

Schwarz:1984:LSK

[Sch84b] Fritz Schwarz. Lie symmetries of the von Kármán equations. Computer Physics Communications, 31(1):113–114, January 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900869.

Schwarz:1984:RPD

[Sch84c] Fritz Schwarz. A reduce package for determining lie symmetries of ordinary and partial differential equations. Computer Physics Communications, 35(1–3):C–815, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482932X.

Schwenn:1984:FVD

[Sch84d] U. Schwenn. Fourier versus difference methods in computing three-dimensional MHD equilibria. Computer Physics Communications, 31(2–3):167–199, February 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900444.

Schnupp:1985:SL

[Sch85a] Peter Schnupp. Specification languages. Computer Physics Communications, 38(2):173–179, October/November 1985.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900839.

Schrijver:1985:PDM

[Sch85b] H. Schrijver. Profiling and debugging in modern programming languages. Computer Physics Communications, 38(2): 289–293, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558590092X.

Schwandt:1985:NLI

[Sch85c] Hartmut Schwandt. Newton-like interval methods for large nonlinear systems of equations on vector computers. Computer Physics Communications, 37(1–3):223–232, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901560.

Schnack:1986:BRB

[Sch86a] Dalton D. Schnack. Book review: Numerical simulation of plasmas: Y. N. Dnestrovskii and D. P. Kostomarov, Springer Series in Computational Physics, Springer-Verlag, Berlin, 1986. 304 pages. DM 169. Computer Physics Communications, 42(3):441–442, November 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900123.

Schoitsch:1986:SEA

[Sch86b] Erwin Schoitsch. Software engineering aspects of real-time programming concepts. Computer Physics Communications, 41(2–3):327–361, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586900743.

Schwarz:1986:CPD

[Sch86c] Fritz Schwarz. Comments on the program dissys: First integrals for π-systems. Computer Physics Communications, 42(1):27, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586902262.

Schwarz:1986:RPD

[Sch86d] Fritz Schwarz. A REDUCE package for determining first integrals of autonomous systems of ordinary differential equations. Computer Physics Communications, 39(2):285–296, February/ March 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586901384.

Schilling:1987:LQC

[Sch87a] K. Schilling. Lattice QCD — a challenge in large scale computing. Computer Physics Communications, 44(3):261–269, June 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900816.

Schneider:1987:RBC

[Sch87b]  T. Schneider.  Relationship between classical diffusion

in random media and quantum localization. Computer Physics Communications, 44(3):307–308, June 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900865.

Schatz:1988:PDP

[Sch88a] George C. Schatz. A program for determining primitive semiclassical eigenvalues for vibrating/rotating nonlinear triatomic molecules. Computer Physics Communications, 51(1–2):135–147, September/October 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900677.

Schoitsch:1988:SSSa

[Sch88b] Erwin Schoitsch. Software safety and software quality assurance in real-time applications: Part 1. Software quality assurance and software safety (concepts and standardization efforts). Computer Physics Communications, 50(1–2):169–188, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901245.

Schoitsch:1988:SSSb

[Sch88c] Erwin Schoitsch. Software-safety and software quality assurance in real-time applications: Part 2: Real-time structures and languages. Computer Physics Communications, 50(1–2):189– 211, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901257.

Schaile:1989:DGT

[Sch89a] Otto Schaile. DZDISP: a graphics tool to interact with ZE-BRA data structures. Computer Physics Communications, 57 (1–3):528–531, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589902804.

Schilling:1989:GD

[Sch89b] Peter K. Schilling. Graphics at DESY. Computer Physics Communications, 57(1–3):443–446, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902592.

Schoen:1989:SSM

[Sch89c] Martin Schoen. Structure of a simple molecular dynamics FORTRAN program optimized for Cray vector processing computers. Computer Physics Communications, 52(2): 175–185, January/February 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900027.

Sciutto:1989:PPP

[Sci89] S. J. Sciutto. Polyfit — a package for polynomial fitting. Computer Physics Communications, 52(3):427–442, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901173.

Scott:1989:ILL

[Sco89] David S. Scott. Implementing Lanczos-like algorithms on hypercube architectures. Computer Physics Communications, 53(1–3):271–281, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901653.

Spadaccini:1988:MSL

[SdAD^{+}88] G. Spadaccini, R. de Asmundis, A. D’Onofrio, M. Romano, and F. Terrasi. A multiprocessor system for online data acquisition, handling and display. Computer Physics Communications, 50(1–2):263–267, July 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558890135X.

Sheldon:1984:SFI

[SDH84] E. Sheldon, D. R. Donati, and H. R. Hiddleston. ’MIA’, a Fortran-IV program for making spin and parity assignments to high-lying single and coherent twin nuclear levels from (alpha, nucleon) angular distributions in on-resonance, compound-nuclear, channel-spin-1/2 reactions. Computer Physics Communications, 35(1–3):C–274–C–275, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825060.

Seaton:1982:CFA

[Sea82] M. J. Seaton. Coulomb functions analytic in the energy. Computer Physics Communications, 25(1):87–95, January 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900479.

Sears:1984:APD

[Sea84a] Trevor J. Sears. ASYTOP — a program for detailed analysis of gas phase magnetic resonance spectra of asymmetric top molecules. Computer Physics Communications, 34(1–2):123–133, November/December 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901644.

Seaton:1984:AIJ

[Sea84b] M. J. Seaton. The accuracy of iterated JWBK approximations for Coulomb radial wave functions. Computer Physics Communications, 32(2):115–119, May 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558490064X.

Seaton:1984:CFA

[Sea84c] M. J. Seaton. Coulomb functions analytic in the energy. Computer Physics Communications, 35(1–3):C–771, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828994.

Saulys:1989:MDA

[SEF^{+}89] A. C. Saulys, A. Etkin, K. J. Foley, R. W. Hackenburg, R. S. Longrcre, W. A. Love, T. W. Morris, E. D. Platner, S. J. Lindenbaum, C. S. Chan, and M. A. Kramer. MPS data-acquisition software system. Computer Physics Communications, 57(1–3):353–357, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902427.

Seiler:1984:PCF

[Sei84] F. Seiler. A program calculating the formulae for polarization effects in nuclear reactions. Computer Physics Communications, 35(1–3):C–221, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824613.

Shayler:1984:RRF

[SF84a] P. J. Shayler and M. T. C. Fang. Radial radiative flux in cylindrically symmetric arcs. Computer Physics Communications, 35(1–3):C–527, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827083.

Sutherland:1984:CCC

[SF84b] D. C. Sutherland and W. G. Ferrier. Coulomb coefficients for complex ionic crystals. Computer Physics Communications, 35(1–3):C–396, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826077.

Shoucri:1987:AFE

[SFB87] M. Shoucri, V. Fuchs, and A. Bers. The application of the finite element library TWODEPEP for the numerical solution of the relativistic Fokker–Planck equation. Computer Physics Communications, 46(3):337–344, September 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900907.

Seijo:1986:ML

[SFP86] L. Seijo, M. Flórez, and L. Pueyo. Matrix linearization. Computer Physics Communications, 42(1):127–136, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586902377.

Scott:1981:ISM

[SG81] D. S. Scott and R. Gruber. Implementing sparse matrix techniques in the erato code. Computer Physics Communications, 23(2):115–121, July 1, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900254.

Saunders:1982:ACQ

[SG82] V. R. Saunders and M. F. Guest. Applications of the Cray-1 for quantum chemistry calculations. Computer Physics Communications, 26(3–4):389–395, June 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901333.

Sheldon:1983:PCS

[SG83] E. Sheldon and R. H. Giles. Programs ‘CELESTE’ and ‘STELLA’ for computations in special relativity: Evaluation of the celestial view from an interstellar spacecraft. Computer Physics Communications, 29(3):269–285, May 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900073.

Schulten:1984:REC

[SG84a] K. Schulten and R. G. Gordon. Recursive evaluation of 3 j and 6 j coefficients. Computer Physics Communications, 35(1–3):C–377–C–379, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825928.

Sheldon:1984:PCS

[SG84b] E. Sheldon and R. H. Giles. Programs ‘celeste’ and ‘stella’ for computations in Special Relativity: Evaluation of the celestial view from an interstellar spacecraft. Computer Physics Communications, 35(1–3):C–881, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829847.

Sotona:1984:GTB

[SG84c] M. Sotona and M. Gmitro. Generalized transformation brackets for the harmonic oscillator functions. Computer Physics Communications, 35(1–3):C–122, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823802.

Smith:1989:DBC

[SG89] Kenneth Smith and Alan H. Glasser. Data base of cross sections and reaction rates for hydrogen ion sources. Computer Physics Communications, 54(2–3):391–407, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900994.

Schastok:1989:EPG

[SGS^{+}89] J. Schastok, H. Gleixner, M. Soffel, H. Ruder, and M. Schneider. The ephemeris program GLE200. Computer Physics Communications, 54(1):167–170, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900416.

Scott:1982:MEV

[SH82] N. S. Scott and A. Hibbert. A more efficient version of the weights and NJSYM packages. Computer Physics Communications, 28(2):189–200, December 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900546.

Scott:1984:MEV

[SH84] N. S. Scott and A. Hibbert. A more efficient version of the weights and Njsym packages. Computer Physics Communications, 35(1–3):C–844–C–845, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829537.

Shapiro:1984:AS

[Sha84] Joel Shapiro. Arbitrary 3n-j symbols for SU(2). Computer Physics Communications, 35(1–3):C–25, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822924.

Shally:1987:CNL

[Sha87] Rudolph Shally. Constrained nonlinear least squares fitting. Computer Physics Communications, 46(3):437–443, September 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558790097X.

Shestakov:1984:AIF

[She84] Aleksei I. Shestakov. Application of the implicit fourierexpansion method to the calculation of three-dimensional equilibria by the iterative method. Computer Physics Communications, 31(2–3):227–233, February 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558490047X.

Shorer:1981:MDD

[Sho81] P. Shorer. R-matrix dynamic dipole polarizabilities. Computer Physics Communications, 22(4):467–472, May 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901430.

Shorer:1984:MDD

[Sho84] P. Shorer. R-matrix dynamic dipole polarizabilities. Computer Physics Communications, 35(1–3):C–698–C–699, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828477.

Shumaker:1987:EAP

[Shu87] D. E. Shumaker. EIV: Axisymmetric plasma equilibrium code. Computer Physics Communications, 44(1–2):177–196, April/ May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900269.

Silver:1984:DMBa

[Sil84a] David M. Silver. Diagrammatic many-body perturbation expansion for atoms and molecules: I. general organization. Computer Physics Communications, 35(1–3):C–448–C–449, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826545.

Silver:1984:DMBb

[Sil84b]  David M. Silver.  Diagrammatic many-body perturba-

tion expansion for atoms and molecules: II. second-order and third-order ladder energies. Computer Physics Communications, 35(1–3):C–450–C–451, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826557.

Simunek:1980:GRM

[Sim80] A. Simunek. Gilat–Raubenheimer method for k-space integration. Computer Physics Communications, 20(3):349–352, November 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900156.

Simunek:1984:GRM

[Sim84] A. Simunek. Gilat — raubenheimer method for k-space integration. Computer Physics Communications, 35(1–3):C–640, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828015.

Sinfailam:1984:ESC

[Sin84] A. L. Sinfailam. Electron scattering by closed shell diatomic molecules. Computer Physics Communications, 35(1–3):C– 52, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584823176.

Sjostrand:1982:LMC

[Sjö82] Torbjörn Sjöstrand. The lund Monte Carlo for jet fragmentation. Computer Physics Communications, 27(3):243–284, September 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901758.

Sjostrand:1983:LMC

[Sjö83] Torbjörn Sjöstrand. The lund Monte Carlo for e^{+}e^{−} jet physics. Computer Physics Communications, 28(3):229–254, January 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900413.

Sjostrand:1984:LMCb

[Sjö84a] Torbjörn Sjöstrand. The Lund Monte Carlo for e^{+}e^{−} jet physics. Computer Physics Communications, 35(1–3):C–848, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829574.

Sjostrand:1984:LMCa

[Sjö84b] Torbjörn Sjöstrand. The Lund Monte Carlo for jet fragmentation. Computer Physics Communications, 35(1–3):C–820, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829367.

Sjostrand:1986:LMC

[Sjö86] Torbjörn Sjöstrand. The Lund Monte Carlo for jet fragmentation and e^{+}e^{−} physics — jetset version 6.2. Computer Physics Communications, 39(3):347–407, April 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900962.

Spalburg:1982:PSS

[SK82] M. R. Spalburg and U. C. Klomp. A program to solve a set of linear coupled differential equations describing a collision process with several electronic and vibrational degrees of freedom. Computer Physics Communications, 28(2):207–215, December 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558290056X.

Spalburg:1984:PSS

[SK84] M. R. Spalburg and U. C. Klomp. A program to solve a set of linear coupled differential equations describing a collision process with several electronic and vibrational degrees of freedom. Computer Physics Communications, 35(1–3):C–847, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829550.

Skouras:1986:GFP

[SK86] L. D. Skouras and S. Kossionides. Generalized fractional parentage coefficients for shell-model calculations. Computer Physics Communications, 39(2):197–212, February/March 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901311.

Skala:1981:GSA

[Ská81] Lubomír Skála. Generation of symmetry-adapted functions for molecular calculations. Computer Physics Communications, 24(2):135–140, November 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900874.

Skarda:1982:MLB

[Ska82] V. Skarda. Microprocessor- and LSI-based CAMAC RAM controllers in a modular multiparameter instrumentation system. Computer Physics Communications, 26(1–2):99–111, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558290162X.

Skala:1984:GSA

[Ská84] Lubomír Skála. Generation of symmetry-adapted functions for molecular calculations. Computer Physics Communications, 35(1–3):C–752, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828830.

Skoog:1984:NEL

[Sko84] R. Skoog. Nuclear energy loss and scattering of ions penetrating thin layers of matter. Computer Physics Communications, 35(1–3):C–254, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482490X.

Skotniczny:1989:QFU

[Sko89] Zbigniew Skotniczny. Query by forms: User-oriented relational database retrieving system and its application in analysis of experiment data. Computer Physics Communications, 57(1–3):225–230, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902178.

Spagna:1983:DRT

[SL83] George F. Spagna, Jr. and Chun Ming Leung. DUSTCD: a radiative transport code for spherically symmetric dust clouds. Computer Physics Communications, 28(4):337–353, February 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900322.

Spagna:1984:DRT

[SL84] George F. Spagna, Jr. and Chun Ming Leung. Dustcd: a radiative transport code for spherically symmetric dust clouds. Computer Physics Communications, 35(1–3):C–856, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829653.

Shaub:1984:CSP

[SLH84] W. M. Shaub, S. Lemont, and A. B. Harvey. Cars spectral profiles for homonuclear diatomic molecules. Computer Physics Communications, 35(1–3):C–520–C–521, ????

1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827022.

Sliwa:1989:CEP

[Sli89] Krzysztof Sliwa. CDF’s experience with a parallel architecture multiprocessor system — ACP. Computer Physics Communications, 57(1–3):407–412, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990252X.

Soto:1981:CGT

[SM81a] M. F. Soto, Jr. and R. Mirman. Computation of group tables for the symmetric groups. Computer Physics Communications, 23(1):81–93, June 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581901314.

Soto:1981:CSG

[SM81b] M. F. Soto, Jr. and R. Mirman. Construction of symmetric group representation matrices and states. Computer Physics Communications, 23(1):95–107, June 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901326.

Soto:1982:NRM

[SM82] M. F. Soto, Jr. and R. Mirman. Number of representations and maximum dimensions for S(N). Computer Physics Communications, 27(1):57–64, July 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900078.

Savelli:1984:CCC

[SM84a] C. Savelli and M. Morando. Carla: a code to calculate the population of high spin states through compound nucleus reactions. Computer Physics Communications, 35(1–3):C–505, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826909.

Schindler:1984:FTF

[SM84b] Susan Schindler and R. Mirman. Functions on tableaux and frames of the symmetric group. Computer Physics Communications, 35(1–3):C–493, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482682X.

Schindler:1984:GCG

[SM84c] Susan Schindler and R. Mirman. Generation of the Clebsch– Gordan coefficients for S_{n}. Computer Physics Communications, 35(1–3):C–490–C–492, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826818.

Soto:1984:CGT

[SM84d] M. F. Soto, Jr. and R. Mirman. Computation of group tables for the symmetric groups. Computer Physics Communications, 35(1–3):C–705–C–706, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828520.

Soto:1984:CSG

[SM84e] M. F. Soto, Jr. and R. Mirman. Construction of symmetric group representation matrices and states. Computer Physics Communications, 35(1–3):C–707–C–710, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828532.

Soto:1984:NRM

[SM84f] M. F. Soto, Jr. and R. Mirman. Number of representations and maximum dimensions for S(N). Computer Physics Communications, 35(1–3):C–806, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829252.

Simpson:1985:RTS

[SM85a] P. Simpson and B. C. Merrifield. Real time signal processing applications of a distributed array processor. Computer Physics Communications, 37(1–3):133–140, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901456.

Soto:1985:CUG

[SM85b] M. F. Soto, Jr. and R. Mirman. Computation of unitary group representation matrices. Computer Physics Communications, 34(4):357–363, February 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900645.

Soto:1985:CCS

[SM85c] M. F. Soto, Jr. and R. Mirman. Construction of canonical states of unitary groups. Computer Physics Communications, 34(4):347–355, February 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900633.

Soto:1985:ICU

[SM85d] M. F. Soto, Jr. and R. Mirman. Invariants and commutators for unitary group representations. Computer Physics Communications, 34(4):365–370, February 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900657.

Soto:1985:NSU

[SM85e] M. F. Soto, Jr. and R. Mirman. Number of states of unitary group representations. Computer Physics Communications, 34(4):339–345, February 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900621.

Soto:1985:RAP

[SM85f] M. F. Soto, Jr. and R. Mirman. Remark on algorithm perm. Computer Physics Communications, 34(4):337–338, February 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558590061X.

Sullivan:1986:EAB

[SM86] Francis Sullivan and Raymond D. Mountain. An efficient algorithm for the Brownian dynamics simulation of aggregation. Computer Physics Communications, 42(1):43–49, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586902298.

Serzu:1989:DMF

[SM89a] Mulugeta H. Serzu and Wooil M. Moon. DIP-moveout by Fourier transform. Computer Physics Communications, 52 (3):337–344, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901094.

Serzu:1989:TDF

[SM89b] Mulugeta H. Serzu and Wooil M. Moon. Two dimensional fast Fourier transform for large data matrices. Computer Physics Communications, 52(3):333–336, March 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901082.

Silvester:1981:TFE

[SMC81] Peter P. Silvester, F. U. Minhas, and Z. J. Csendes. Tetrahedral finite element matrix primitives. Computer Physics Communications, 24(2):173–179, November 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900904.

Silvester:1984:TFE

[SMC84] Peter P. Silvester, F. U. Minhas, and Z. J. Csendes. Tetrahedral finite element matrix primitives. Computer Physics Communications, 35(1–3):C–756, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828866.

Sheldon:1984:CTDb

[SMD84] E. Sheldon, S. Mathur, and D. Donati. Computation of total, differential and double-differential cross sections for compound nuclear reactions of the type (a, b), (a, bγ) and (a, bγ − γ). (III) Fortran translations of the Algol programs ‘mandy’ and ‘barbara’. Computer Physics Communications, 35(1–3):C–91–C– 93, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584823528.

Smith:1984:STM

[Smi84a] Kenneth Smith. SUBMMW: a theoretical model to predict CW sub-millimeter wave laser performance. Computer Physics Communications, 35(1–3):C–483–C–484, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826776.

Smith:1984:HFN

[Smi84b] William R. Smith. Hauser–Feshbach nuclear scattering subroutine liana. Computer Physics Communications, 35(1–3):C–20– C–21, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584822894.

Smith:1984:NBS

[Smi84c] William R. Smith. Nuclear bound state wave functions subroutine. Computer Physics Communications, 35(1–3):C–10, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822808.

Smith:1984:NES

[Smi84d] William R. Smith. Nuclear elastic scattering program with parameter search. Computer Physics Communications, 35 (1–3):C–23–C–24, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584822912.

Smith:1984:NPP

[Smi84e] William R. Smith. Nuclear penetrability and phase shift subroutine. Computer Physics Communications, 35(1–3):C–14, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584822845.

Smith:1984:PSS

[Smi84f] William R. Smith. Parameter search subroutine. Computer Physics Communications, 35(1–3):C–17, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822869.

Scott:1989:RMP

[SMM89] T. C. Scott, R. A. Moore, and M. B. Monagan. Resolution of many particle electrodynamics by symbolic manipulation. Computer Physics Communications, 52(2):261–281, January/ February 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558990009X.

Salvat:1986:MCP

[SMMP86] F. Salvat, J. D. Martinez, R. Mayol, and J. Parellada. A Monte Carlo program to simulate the penetration and energy loss of keV electrons through matter. Computer Physics Communications, 42(1):93–104, September 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586902341.

Scott:1987:PCR

[SMR87] N. S. Scott, P. Milligan, and H. W. C. Riley. The parallel computation of Racah coefficients using transputers. Computer Physics Communications, 46(1):83–98, July 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900373.

Sutcliffe:1988:ECA

[SMT88] B. T. Sutcliffe, S. Miller, and J. Tennyson. An effective computational approach to the calculation of the vibration-rotation spectra of triatomic molecules. Computer Physics Communications, 51(1–2):73–82, September/October 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558890063X.

Savoy-Navarro:1984:UDB

[SN84] A. Savoy-Navarro. Use of a data base in the on-line environment of a large-scale high-energy physics experiment. Computer Physics Communications, 33(1–3):173–195, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901206.

Schmid:1980:VPC

[SNC80] G. Bruno Schmid, D. W. Norcross, and L. A. Collins. VLAM, a program for computing the electron-molecule static interaction potential from a Legendre expansion of the molecular charge density. Computer Physics Communications, 21(1):79–90, December 1, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580900776.

Schmid:1984:VPC

[SNC84] G. Bruno Schmid, D. W. Norcross, and L. A. Collins. VLAM, a program for computing the electron–molecule static interaction potential from a Legendre expansion of the molecular charge density. Computer Physics Communications, 35(1–3): C–651, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828118.

Snelling:1985:HAR

[Sne85] David F. Snelling. HEP applications: Real time flight simulation. Computer Physics Communications, 37(1–3):261– 271, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585901614.

Salamon:1988:SAC

[SNH^{+}88] Peter Salamon, James D. Nulton, John R. Harland, Jacob Pedersen, George Ruppeiner, and Luby Liao. Simulated annealing with constant thermodynamic speed. Computer Physics Communications, 49(3):423–428, June 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900033.

Sonnerup:1988:TSS

[Son88] B. U. Ö. Sonnerup. On the theory of steady state reconnection. Computer Physics Communications, 49(1):143–159, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902226.

Saldin:1986:CED

[SP86] D. K. Saldin and J. B. Pendry. Calculation of elastic diffuse LEED intensities from disordered adsorbates. Computer Physics Communications, 42(3):399–415, November 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900081.

Saldin:1987:CRE

[SP87] D. K. Saldin and J. B. Pendry. Calculation of the renormalised electron scattering matrix of a molecule adsorbed on a crystal surface. Computer Physics Communications, 46(1):129– 140, July 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900403.

Sphicas:1987:UES

[Sph87] P. A. Sphicas. UA1 experience with 3081/E systems. Computer Physics Communications, 45(1–3):339–343, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901718.

Singer:1988:BOD

[SPM^{+}88] C. E. Singer, D. E. Post, D. R. Mikkelsen, M. H. Redi, A. McKenney, A. Silverman, F. G. P. Seidl, P. H. Rutherford, R. J. Hawryluk, W. D. Langer, L. Foote, D. B. Heifetz, W. A. Houlberg, M. H. Hughes, R. V. Jensen, G. Lister, and J. Ogden. Baldur: a one-dimensional plasma transport code. Computer Physics Communications, 49(2):275–398, May 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900124.

Sheldon:1984:CTDc

[SR84a] E. Sheldon and V. C. Rogers. Computation of total and differential cross section for compound nuclear reactions of the type (a, a), (a, a′), (a, b), (a, γ), (a, γ − γ), (a, bγ), and (a, bγ − γ) (IV) Fortran program ‘CINDY’. Computer Physics Communications, 35(1–3):C–212–C–213, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824558.

Sokolovsk:1984:CPS

[SR84b]  T. D. Sokolovsk and L. A. Rogoschenko.  Computa-

tion of phonon spectrum from the cold neutron incoherent inelastic scattering by a polycrystal. Computer Physics Communications, 35(1–3):C–439, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826478.

Smith:1985:VHP

[SRS85] K. A. Smith, S. F. Reddaway, and D. M. Scott. Very high performance pseudo-random number generation on DAP. Computer Physics Communications, 37(1–3):239–244, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901584.

Schluter:1981:ESS

[SS81] Arnulf Schlüter and Ulrich Schwenn. Equilibrium and stability studies with the 3D MHD code tube. Computer Physics Communications, 24(3–4):263–300, December 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558190151X.

Samuel:1984:CEQ

[SS84a] M. Samuel and U. Smilansky. Calculation of electric quadrupole radial matrix elements for Coulomb excitation. Computer Physics Communications, 35(1–3):C–113, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482370X.

Sheldon:1984:CTDa

[SS84b] Eric Sheldon and Richard Michael Strang. Computation of total, differential, and double-differential cross sections for compound nuclear reactions of the type (a, b), (a, bγ) and (a, bγ −γ) (II) generalized programs “mandy” and “barbara” for arbitrary angular momenta in Hauser–Feshbach–Moldauer formalism. Computer Physics Communications, 35(1–3):C–7–C–8, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482278X.

Smend:1984:FFP

[SS84c] F. Smend and M. Schumacher. Form factor program for Rayleigh scattering of gamma rays by bound electrons. Computer Physics Communications, 35(1–3):C–253, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824893.

Smend:1984:NRF

[SS84d] F. Smend and M. Schumacher. Non-relativistic form factor program for Compton scattering of gamma rays by bound electrons. Computer Physics Communications, 35(1–3):C–386, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825990.

Sawaryn:1989:CAM

[SS89]  Andrzej Sawaryn and W. Andrzej Sokalski.  Cumula-

tive atomic multipole moments and point charge models describing molecular charge distribution. Computer Physics Communications, 52(3):397–408, March 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901148.

Shoucri:1989:QFP

[SSF^{+}89] M. Shoucri, I. Shkarofsky, V. Fuchs, K. Kupfer, A. Bers, and S. Luckhardt. A quasilinear Fokker–Planck code for the numerical solution of the lower-hybrid current-drive problem in the presence of electron cyclotron heating. Computer Physics Communications, 55(3):253–268, October 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901240.

Schonauer:1985:DPS

[SSM85] W. Schönauer, E. Schnepf, and H. Müller. Designing PDE software for vector computers as a “data flow algorithm”. Computer Physics Communications, 37(1–3):233–237, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901572.

Schaupp:1984:RSC

[SSS84] D. Schaupp, M. Schumacher, and F. Smend. Rayleigh selfconsistent relativistic form factors and modified form factors. Computer Physics Communications, 32(4):413–419, July 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900572.

Salvini:1981:VED

[ST81] S. A. Salvini and D. G. Thompson. The vibrational excitation of diatomic molecules by electron impact. Computer Physics Communications, 22(1):49–58, February/March 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900783.

Scott:1982:GPC

[ST82a] N. S. Scott and K. T. Taylor. A general program to calculate atomic continuum processes incorporating model potentials and the Breit–Pauli Hamiltonian within the R-matrix method. Computer Physics Communications, 25(4):347–387, April 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582900765.

Sullivan:1982:NIM

[ST82b] E. C. Sullivan and A. Temkin. A non-iterative method for solving PDE’s arising in electron scattering. Computer Physics Communications, 25(1):97–110, January 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900480.

Salvini:1984:VED

[ST84a] S. A. Salvini and D. G. Thompson. The vibrational excitation of diatomic molecules by electron impact. Computer Physics Communications, 35(1–3):C–684, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828350.

Schwenke:1984:OQS

[ST84b] David W. Schwenke and Donald G. Truhlar. An optimized quadrature scheme for matrix elements over the eigenfunctions of general anharmonic potentials. Computer Physics Communications, 34(1–2):57–66, November/December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901590.

Scott:1984:GPC

[ST84c] N. S. Scott and K. T. Taylor. A general program to calculate atomic continuum processes incorporating model potentials and the Breit–Pauli Hamiltonian within the R-matrix method. Computer Physics Communications, 35(1–3):C–794– C–797, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829185.

Servizi:1984:CPB

[ST84d] G. Servizi and G. Turchetti. A computer program for the Birkhoff series of area preserving maps. Computer Physics Communications, 32(2):201–207, May 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900717.

Smith:1984:PEN

[ST84e] Richard L. Smith and Donald G. Truhlar. Program for evaluation of non-exchange type integrals required in electronatom scattering theory using Slater-type orbitals as basis functions. Computer Physics Communications, 35(1–3):C–170–C– 171, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824200.

Sullivan:1984:NIM

[ST84f] E. C. Sullivan and A. Temkin. A non-iterative method for solving PDE’s arising in electron scattering. Computer Physics Communications, 35(1–3):C–772–C–773, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829008.

Slaets:1989:PCC

[ST89a] Jan F. W. Slaets and Gonzalo Travieso. Parallel computing: a case study. Computer Physics Communications, 56(1):63–67, November 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900532.

Surguladze:1989:LPM

[ST89b] L. R. Surguladze and F. V. Tkachov. LOOPS: Procedures for multiloop calculations in quantum field theory for the REDUCE system. Computer Physics Communications, 55(2):205–215, September 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900775.

Starkand:1984:SCM

[Sta84] Yair Starkand. Subroutine for calculation of matrix padé approximants. Computer Physics Communications, 35(1–3):C– 382, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825953.

Stanescu:1987:SGL

[Sta87a] Cristian Stanescu. Standard and graphics libraries in online environment. A comparative study of performance. Computer Physics Communications, 45(1–3):475–478, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901925.

Start:1987:BRB

[Sta87b] D. F. H. Start. Book review: Computational Methods for Kinetic Models of Magnetically Confined Plasmas: J. Killeen, G. f. Kerbel, M. G. McCoy and A. A. Mirin, Springer Series in Computational Physics, Springer-Verlag, Berlin, 1986. ISBN 3-540-13401-8. DM 108. Computer Physics Communications, 46(3):453–454, September 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900993.

Stanco:1989:PTR

[Sta89] Luca Stanco. Particle track reconstruction in heavy materials with the Kalman technique. Computer Physics Communications, 57(1–3):380–384, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902476.

Stedman:1984:LCF

[Ste84a] G. E. Stedman. Lanthanide crystal field fitting routine. Computer Physics Communications, 35(1–3):C–82, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823449.

Stefnescu:1984:POC

[Ste84b] A. Sabba Stefnescu. Program for optimal computation of EPR SPIN-HAMILTONIAN parameters for ions in tetragonal symmetry. Computer Physics Communications, 35(1–3): C–376, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584825916.

Stern:1984:IEP

[Ste84c] M. S. Stern. An integral equation program to calculate radial wave functions and scattering phase shifts of short-range local interactions. Computer Physics Communications, 35(1–3):C– 564, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827393.

Stevens:1984:OSD

[Ste84d] John G. Stevens. The operation of a small data centre (Mössbauer spectroscopy). Computer Physics Communications, 33(1–3):105–114, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901139.

Steuer:1987:MPP

[Ste87] M. F. Steuer. Monitoring of project-progress via distributed database. Computer Physics Communications, 45(1–3):391– 394, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587901792.

Shadwick:1989:PCV

[STN89] B. A. Shadwick, J. D. Talman, and M. R. Norman. A program to compute variationally optimized relativistic atomic potentials. Computer Physics Communications, 54(1):95–102, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900350.

Story:1989:SEI

[Sto89] C. M. Story. Software engineering in industry. Computer Physics Communications, 57(1–3):217–221, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902154.

Stringer:1981:MPR

[Str81] T. E. Stringer. MHD problems relevant to jet. Computer Physics Communications, 24(3–4):337–342, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901569.

Struven:1987:NTN

[Str87] Warren Struven. Networking through the new phone system the future of telecommunications. Computer Physics Communications, 45(1–3):443–446, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901871.

Studer:1985:KBS

[Stu85] Rudi Studer. Knowledge-based software engineering environments. Computer Physics Communications, 38(2):277– 287, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900918.

Stuart:1988:ARO

[Stu88] Robin G. Stuart. Algebraic reduction of one-loop Feynman diagrams to scalar integrals. Computer Physics Communications, 48(3):367–389, March 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588902020.

Succi:1987:CAM

[Suc87] S. Succi. Cellular automata modeling on IBM 3090/VF. Computer Physics Communications, 47(2–3):173–180, November/ December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901032.

Sufrin:1989:FMS

[Suf89] Bernard Sufrin. Formal methods in system design and implementation. Computer Physics Communications, 57(1–3): 108–117, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901987.

Sundholm:1988:BPC

[Sun88] Dage Sundholm. A block preconditioned conjugate gradient method for solving high-order finite element matrix equations. Computer Physics Communications, 49(3):409–415, June 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558890001X.

Shmakov:1989:DGI

[SUZ89] S. Yu. Shmakov, V. V. Uzhinskii, and A. M. Zadorozhny. DIA-GEN — generator of inelastic nucleus-nucleus interaction diagrams. Computer Physics Communications, 54(1):125–135, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589900386.

Stanev:1984:SSS

[SV84] T. Stanev and Ch. Vankov. A set of subroutines for simulation of electron-photon cascades. Computer Physics Communications, 35(1–3):C–547, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482723X.

Spineanu:1986:SPS

[SVP86] F. Spineanu, M. Vlad, and I. I. Popescu. Strimp — program for studying the impurity evolution in tokamak plasma. Computer Physics Communications, 41(1):155–168, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900263.

Schmidt:1980:MCC

[SW80a] H. U. Schmidt and R. Wegmann. MHD-calculations for cometary plasmas. Computer Physics Communications, 19 (3):309–326, July/August 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900867.

Stone:1980:RRF

[SW80b] A. J. Stone and C. P. Wood. Root-rational-fraction package for exact calculation of vector-coupling coefficients. Computer Physics Communications, 21(2):195–205, December 2, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900405.

Stone:1984:RRF

[SW84] A. J. Stone and C. P. Wood. Root-rational-fraction package for exact calculation of vector-coupling coefficients. Computer Physics Communications, 35(1–3):C–660, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828180.

Schmidt:1988:CGP

[SWHB88] H. U. Schmidt, R. Wegmann, W. F. Huebner, and D. C. Boice. Cometary gas and plasma flow with detailed chemistry. Computer Physics Communications, 49(1):17–59, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902147.

Swift:1988:PSP

[Swi88] Daniel W. Swift. Particle simulation of plasma processes in the Earth’s magnetotail. Computer Physics Communications, 49(1):173–183, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558890224X.

Sykes:1981:DCN

[Syk81] A. Sykes. Developments of the Culham 3D nonlinear MHD code. Computer Physics Communications, 24(3–4):437–439, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901685.

Sommer:1984:NBT

[SZ84] B. Sommer and J. G. Zabolitzky. On numerical Bessel transformation. Computer Physics Communications, 35(1–3):C–549, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584827253.

Szekely:1985:FFG

[Szé85] Géza Székely. FGM — a flexible gamma-spectrum analysis program for a small computer. Computer Physics Communications, 34(3):313–324, January 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900086.

Szymanski:1981:CCE

[Szy81] Z. Szymański. Computing collective excitations of nuclei. Computer Physics Communications, 22(2–3):177–185, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900473.

Tajuddin:1989:CAT

[TA89] Wan Ahmad Tajuddin and Wan Abdullah. Connectionist architectures for triggering and track reconstruction. Computer Physics Communications, 57(1–3):472–475, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902658.

Tajima:1986:BRB

[Taj86] T. Tajima. Book review: Plasma Physics via Computer Simulation: C. K. Birdsall and A. B. Langdon, McGraw-Hill, New York, 1985. xxiii + 479 pages. US $45. Computer Physics Communications, 42(1):151–152, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586902407.

Takemasa:1985:PCS

[Tak85a] Tadashi Takemasa. A program for calculating spectroscopic amplitudes for two-nucleon transfer reactions by projecting angular momentum. Computer Physics Communications, 36(1):79– 100, March 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585900190.

Tanaka:1985:MMR

[TAK^{+}85b] Y. Tanaka, M. Azumi, G. Kurita, T. Tsunematsu, and T. Takeda. A matrix method for resistive MHD stability analysis of axisymmetric toroidal plasma. Computer Physics Communications, 38(3):339–346, December 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901018.

Takemasa:1988:AWGa

[Tak88a] Tadashi Takemasa. Abscissae and weights for the Gauss– Hermite quadrature formula. Computer Physics Communications, 48(2):265–270, February 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558890046X.

Takemasa:1988:AWGb

[Tak88b] Tadashi Takemasa. Abscissae and weights for the Gauss– Laguerre quadrature formula. Computer Physics Communications, 52(1):133–140, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901786.

Tallents:1982:PCE

[Tal82] G. J. Tallents. Profile: a code for evaluating line profile shapes for optically thick expanding plasmas. Computer Physics Communications, 25(2):141–148, February 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900303.

Talman:1983:LSC

[Tal83] James D. Talman. LSFBTR: a subroutine for calculating spherical Bessel transforms. Computer Physics Communications, 30(1):93–99, July/August 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583901261.

Tallents:1984:PCE

[Tal84a] C. J. Tallents. Profile: a code for evaluating line profile shapes for optically thick expanding plasmas. Computer Physics Communications, 35(1–3):C–776, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829033.

Tallents:1984:CCC

[Tal84b] G. J. Tallents. Collrad: a code for calculating the quasi-steady state population densities of excited states of hydrogen-like ions. Computer Physics Communications, 35(1–3):C–398, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826090.

Talman:1984:LSC

[Tal84c] James D. Talman. LSFBTR: a subroutine for calculating spherical Bessel transforms. Computer Physics Communications, 35(1–3):C–903, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584830027.

Talman:1989:PCV

[Tal89] James D. Talman. A program to compute variationally optimized effective atomic potentials. Computer Physics Communications, 54(1):85–94, April 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900349.

Tamura:1984:AMC

[Tam84] Taro Tamura. Angular momentum coupling coefficients. Computer Physics Communications, 35(1–3):C–39, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823061.

Tarutin:1986:PCA

[Tar86] O. B. Tarutin. The programming of cross-assemblers on the ES-computer with the use of standard macroassembler. Computer Physics Communications, 41(2–3):415–417, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900822.

Tatarkiewicz:1980:IMT

[Tat80] Jakub Tatarkiewicz. On the inclusion of macroscopic theory in Monte Carlo simulation using game theory. Computer Physics Communications, 20(1):133–137, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580901198.

Taurian:1980:MPN

[Tau80] Oscar E. Taurian. A method and a program for the numerical evaluation of the Hilbert transform of a real function. Computer Physics Communications, 20(2):291–307, October 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900089.

Taurian:1981:FFP

[Tau81] Oscar E. Taurian. FSCRIPT, a full portable text formatting program. Computer Physics Communications, 22(1): 85–96, February/March 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465581900813.

Taurian:1984:EMU

[Tau84a] Oscar E. Taurian. An editor for the maintenance and use of a bank of contracted Gaussian basis set functions. Computer Physics Communications, 33(1–3):55–69, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901073.

Taurian:1984:FFP

[Tau84b] Oscar E. Taurian. Fscript, a full portable text formatting program. Computer Physics Communications, 35(1–3):C–687, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828386.

Taurian:1984:MPN

[Tau84c] Oscar E. Taurian. A method and a program for the numerical evaluation of the Hilbert transform of a real function. Computer Physics Communications, 35(1–3):C–637, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827988.

Taurian:1984:RPG

[Tau84d] Oscar E. Taurian. ROTTRA — a program for generating rotations and translations. Computer Physics Communications, 34(1–2):153–162, November/December 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558490167X.

Tausworthe:1985:CTS

[Tau85] Robert C. Tausworthe. Concepts and tools for the software life cycle. Computer Physics Communications, 38(2): 135–148, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900803.

Thompson:1985:CCF

[TB85] I. J. Thompson and A. R. Barnett. COULCC: a continuedfraction algorithm for Coulomb functions of complex order with complex arguments. Computer Physics Communications, 36(4):363–372, June 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900256.

See [Tho04b].

Thompson:1987:MBF

[TB87] I. J. Thompson and A. R. Barnett. Modified Bessel functions I_{ν}(z) and K_{ν}(z) of real order and complex argument, to selected accuracy. Computer Physics Communications, 47(2–3):245–257, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901111. See erratum [Tho04a].

Tennyson:1987:GCO

[TBB87] Jonathan Tennyson, P. G. Burke, and K. A. Berrington. The generation of continuum orbitals for molecular R-matrix calculations using Lagrange orthogonalisation. Computer Physics Communications, 47(2–3):207–212, November/December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558790107X.

Troyon:1980:ACV

[TBG80] F. Troyon, L. C. Bernard, and R. Gruber. An algorithm to compute the vacuum contribution to the ideal MHD δW in an axisymmetric configuration. Computer Physics Communications, 19(2):161–169, April/June 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900466.

Tennyson:1987:PCP

[TC87] Jonathan Tennyson and N. Chandra. PEAD — for the calculation of photoelectron angular distributions of linear molecules. Computer Physics Communications, 46(1):99–105, July 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900385.

Tamura:1984:DWB

[TCR84] T. Tamura, W. R. Coker, and F. Rybicki. Distorted wave Born approximation for nuclear reactions. Computer Physics Communications, 35(1–3):C–69, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823346.

Tanczyn:1985:APE

[TCS85] R. S. Tanczyn, G. P. Couchell, and W. A. Schier. ASYM: a program to examine fission fragment mass asymmetry in hemispherical chambers. Computer Physics Communications, 38(1):61–70, August/September 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900463.

Thiel:1981:PCA

[TDS81] J. Thiel, P. Dorio, and C. Soubry. The potential created by an alternating point charge in a Maxwellian magneto-plasma. Computer Physics Communications, 23(2):169–180, July 1, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558190031X.

Thiel:1984:PCA

[TDS84] J. Thiel, P. Dorio, and C. Soubry. The potential created by an alternating point charge in a Maxwellian magnetoplasma. Computer Physics Communications, 35(1–3):C–716, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828581.

Tellinghuisen:1984:FQM

[Tel84] Joel Tellinghuisen. A fast quadrature method for computing diatomic RKR potential curves. Computer Physics Communications, 35(1–3):C–220, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824601.

Temperton:1982:FMP

[Tem82] Clive Temperton. Fast methods on parallel and vector machines. Computer Physics Communications, 26(3–4):331–334, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901254.

Tennyson:1983:APC

[Ten83] Jonathan Tennyson. ATOMDIAT — a program for calculating variationally exact ro-vibrational levels of “floppy” triatomics. Computer Physics Communications, 29(3):307–319, May 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465583900103.

Tennyson:1984:APC

[Ten84a] Jonathan Tennyson. Atomdiat — a program for calculating variationally exact ro-vibrational levels of ‘floppy’ triatomics. Computer Physics Communications, 35(1–3):C–884, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829872.

Tennyson:1984:AGA

[Ten84b] Jonathan Tennyson. Atomdiat2 and genpot: Adaptations of atomdiat for the ro-vibrational levels of any floppy triatomic using a general potential function. Computer Physics Communications, 32(1):109–114, April 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900122.

Tennyson:1985:CME

[Ten85] Jonathan Tennyson. On the calculation of matrix elements between polynomial basis functions. Computer Physics Communications, 38(1):39–41, August/September 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900438.

Tennyson:1986:TSR

[Ten86] Jonathan Tennyson. Triatom, select and rotlev — for the calculation of the ro-vibrational levels of triatomic molecules.

Computer Physics Communications, 42(2):257–270, October 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900421.

Tepel:1984:EEN

[Tep84] J. W. Tepel. ENSDF — the evaluated nuclear structure data file. Computer Physics Communications, 33(1–3):129– 146, August/September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584901152.

Terasawa:1988:NDA

[Ter88] Toshio Terasawa. Nonlinear dynamics of Alfvén waves: Interactions between ions and shock upstream waves. Computer Physics Communications, 49(1):193–200, April 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588902263.

Thyagaraja:1989:NMF

[TF89] A. Thyagaraja and D. F. Fletcher. The nonhyperbolicity of multiphase flow equations: a nonlinear nonproblem? Computer Physics Communications, 56(2):115–127, December 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990012X.

Topacli:1988:MSP

[TGA88] C. Topaçli, Y. Gündüç, and A. Aydinli. MULTI: a simulation program for laser processing of multilayer structures. Computer Physics Communications, 52(1):65–70, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901737.

Thaler:1989:DAM

[Tha89] Jon J. Thaler. Data-acquisition modelling and simulation. Computer Physics Communications, 57(1–3):309–312, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589902336.

Taylor:1982:CCS

[THH82] D. J. Taylor, J. F. L. Hopkinson, and C. C. T. Henfrey. The Cray-1s and the Cray service provided by the SERC at the Daresbury Laboratory. Computer Physics Communications, 26(3–4):259–265, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901151.

Thomson:1984:TPD

[Tho84] R. M. Thomson. Transport properties of dilute gas mixtures. Computer Physics Communications, 35(1–3):C–582, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827551.

Thomas:1987:IMU

[Tho87] Samuel Thomas. Inner multiplicity of unitary groups — a modified version. Computer Physics Communications, 44 (1–2):221–225, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900294.

Thompson:2004:EBB

[Tho04a] I. J. Thompson. Erratum to Modified Bessel functions I nu(z) and K_{ν}(z) of real order and complex argument [Comput. Phys. Commun. 47 (1987) 245–257]. Computer Physics Communications, 159(3):243–244, June 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504001067. See [TB87].

Thompson:2004:ECC

[Tho04b] I. J. Thompson. Erratum to “COULCC: a continued-fraction algorithm for Coulomb functions of complex order with complex arguments”: [Comput. Phys. Commun. 36 (1985) 363– 372]. Computer Physics Communications, 159(3):241–242, June 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504001079. See [TB85].

Titterington:1980:CLD

[TK80] D. J. Titterington and C. G. Kinniburgh. Calculation of LEED diffracted intensities. Computer Physics Communications, 20(2):237–266, October 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900053.

Titerington:1984:CLD

[TK84] D. J. Titerington and C. G. Kinniburgh. Calculation of LEED diffracted intensities. Computer Physics Communications, 35(1–3):C–633–C–634, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827952.

Tokunaga:1985:VMC

[TKS^{+}85] Y. Tokunaga, Y. Kurita, M. Sugihara, S. Hitoki, and S. Saito. Vectorization of the Monte Carlo program dicon. Computer Physics Communications, 38(1):15–26, August/September 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900414.

Treleaven:1982:FGC

[TL82] Philip C. Treleaven and Isabel Gouveia Lima. Fifth generation computers. Computer Physics Communications, 26(3–4):277– 283, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901175.

Tamura:1984:EFRa

[TL84] T. Tamura and K. S. Low. Exact finite range DWBA calculations for heavy-ion induced nuclear reactions. Computer Physics Communications, 35(1–3):C–283–C–284, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482514X.

Tezduyar:1989:GEE

[TL89] T. E. Tezduyar and J. Liou. Grouped element-by-element iteration schemes for incompressible flow computations. Computer Physics Communications, 53(1–3):441–453, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990177X.

Tennyson:1989:PSC

[TM89] Jonathan Tennyson and Steven Miller. A program suite for the calculation of ro-vibrational spectra of triatomic molecules. Computer Physics Communications, 55(2):149–175, September 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990074X.

Tennyson:1984:RPD

[TN84] Jonathan Tennyson and Cliff J. Noble. RESON — a program for the detection and fitting of Breit–Wigner resonances. Computer Physics Communications, 33(4):421–424, October 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901474.

Tolsma:1985:SLS

[Tol85] L. D. Tolsma. Solving large sets of coupled equations iteratively by vector processing on the Cyber 205 computer. Computer Physics Communications, 37(1–3):245–249, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901596.

Tolsma:1986:RRC

[Tol86] L. D. Tolsma. Recurrence relations for Coulomb excitation electric multipole radial matrix elements. Computer Physics Communications, 41(1):41–57, July 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900202.

Tornow:1982:ESI

[Tor82] V. Tornow. An exponential spline interpolation for unequally spaced data points. Computer Physics Communications, 28(1):61–67, November 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900625.

Tornow:1984:ESI

[Tor84] V. Tornow. An exponential spline interpolation for unequally spaced data points. Computer Physics Communications, 35(1–3):C–835, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482946X.

Toussaint:1987:SQ

[Tou87] Doug Toussaint. Supercomputations in QCD. Computer Physics Communications, 45(1–3):111–120, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901457.

Toussaint:1989:IAM

[Tou89] D. Toussaint. Introduction to algorithms for Monte Carlo simulations and their application to QCD. Computer Physics Communications, 56(1):69–92, November 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900544.

Tamura:1984:CFC

[TR84] Taro Tamura and Frank Rybicki. Coulomb functions for complex energies. Computer Physics Communications, 35(1–3):C– 5, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584822766.

teRiele:1985:PSF

[tR85] H. J. J. te Riele. A program for solving first kind Fredholm integral equations by means of regularization. Computer Physics Communications, 36(4):423–432, June 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900323.

Trease:1988:TDF

[Tre88] Harold E. Trease. Three-dimensional Free-Lagrange hydrodynamics. Computer Physics Communications, 48(1):39–50, January 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588900215.

Trippe:1984:RPP

[Tri84] Thomas G. Trippe. The review-of-particle-properties system. Computer Physics Communications, 33(1–3):229–233, August/ September 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901280.

Trotman:1984:PCO

[Tro84] J. V. Trotman. A program for closed orbit minimization by analytic technique. Computer Physics Communications, 35(1–3): C–166, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824170.

Truhlar:1988:SCV

[Tru88] Donald G. Truhlar. Scientific computing on vector computers: W. Schönauer, special topics in supercomputing series (volume 2), North-Holland, Amsterdam, 1987. Dfl.180.00 (hard cover). Computer Physics Communications, 50(3):413, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901944.

Thomas:1984:IMU

[TS84a] Samuel Thomas and M. T. Sunny. Inner multiplicity of unitary groups. Computer Physics Communications, 35(1–3):C– 467, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584826673.

Turnbull:1984:CTI

[TS84b] A. D. Turnbull and L. Sparks. Compact torus-ion ring hybrid equilibria. Computer Physics Communications, 31 (2–3):235–248, February 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465584900481.

Tobimatsu:1989:RBS

[TS89] Keijiro Tobimatsu and Yoshimitsu Shimizu. Radiative Bhabha scattering in special configurations with missing final e^{+} and/ or e^{−}. Computer Physics Communications, 55(3):337–358, October 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589901318.

Thomson:1984:BCS

[TSD84] R. M. Thomson, Kenneth Smith, and A. R. Davies. Boltz: a code to solve the transport equation for electron distributions and then calculate transport coefficients and vibrational excitation rates in gases with applied fields. Computer Physics Communications, 35(1–3):C–387–C–388, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826004.

Tenner:1987:P

[TSH87] A. G. Tenner, G. J. Savonije, and J. E. Hansen. Preface. Computer Physics Communications, 44(3):v, June 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900762.

Telle:1982:FPE

[TT82] H. Telle and U. Telle. FCFRKR — a procedure to evaluate Franck–Condon type integrals for diatomic molecules. Computer Physics Communications, 28(1):1–25, November 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900595.

Telle:1984:FPE

[TT84] H. Telle and U. Telle. Fcfrkr — a procedure to evaluate Franck–Condon type integrals for diatomic molecules. Computer Physics Communications, 35(1–3):C–830–C–831, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829434.

Telle:1985:CPF

[TT85] H. H. Telle and U. Telle. Comments on the program FCFRKR. Computer Physics Communications, 36(1):109–112, March 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900219.

Turnbull:1989:TVF

[TT89] A. D. Turnbull and F. Troyon. Two variational forms of the MHD ballooning equation. Computer Physics Communications, 52(3):303–316, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901057.

Takizuka:1981:EFH

[TTAT81] T. Takizuka, S. Tokuda, M. Azumi, and T. Takeda. Effects of the finite hybrid element on MHD stability calculations in a cylindrical plasma. Computer Physics Communications, 23(1): 19–26, June 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901259.

Tucker:1988:VCI

[TTL88] Susan C. Tucker, Todd C. Thompson, and Jack G. Lauderdale. A vibrational configuration interaction program for energies and resonance widths. Computer Physics Communications, 51(1–2):233–256, September/October 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900756.

Tsunematsu:1980:CSM

[TTM^{+}80] T. Tsunematsu, T. Takeda, T. Matsuura, G. Kurita, and M. Azumi. Convergence of solutions of the MHD stability code ERATO. Computer Physics Communications, 19(2):179–183, April/June 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558090048X.

Takeda:1984:IGS

[TTT84] Tatsuoki Takeda, Toshihide Tsunematsu, and Shinji Tokuda. Integrative graphic subroutine package ARGUS-V4. Computer Physics Communications, 34(1–2):15–29, November/December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901553.

Tokuda:1987:DSG

[TTT87] Shinji Tokuda, Toshihide Tsunematsu, and Tatsuoki Takeda. Database system GAEA for a large amount of numerical data. Computer Physics Communications, 44(1–2):21–46, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587900142.

Takemasa:1984:CAM

[TTW84a] T. Takemasa, T. Tamura, and H. H. Wolter. Complex angular momentum methods for elastic scattering with an optical potential. Computer Physics Communications, 35(1–3):C–606, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827733.

Takemasa:1984:CFC

[TTW84b] T. Takemasa, T. Tamura, and H. H. Wolter. Coulomb functions with complex angular momenta. Computer Physics Communications, 35(1–3):C–562, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482737X.

Tamura:1983:DCC

[TUB83] T. Tamura, T. Udagawa, and M. Benhamou. DWBA calculations of continuum spectra of nuclear reactions. Computer Physics Communications, 29(4):391–415, June 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900176.

Tamura:1984:DCC

[TUB84] T. Tamura, T. Udagawa, and M. Benhamou. DWBA calculations of continuum spectra of nuclear reactions. Computer Physics Communications, 35(1–3):C–890, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829938.

Tamura:1984:EFRb

[TUWA84] T. Tamura, T. Udagawa, K. W. Wood, and H. Amakawa. Exact finite range dwba form factor for heavy-ion induced nuclear reactions. Computer Physics Communications, 35(1–3):C–588, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827605.

Tolsma:1986:MAS

[TV86] L. D. Tolsma and G. W. Veltkamp. Measuring the accuracy of the solution subspace obtained by numerical integration of the Schrödinger equation. Computer Physics Communications, 40(2–3):233–262, June 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586901128.

Treleaven:1989:NCO

[TV89] Philip Treleaven and Marley Vellasco. Neural computing overview. Computer Physics Communications, 57(1–3):543– 559, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589902841.

Thompson:1987:TTO

[TVD87] Murray Thompson, Ann Varda, and Gary DeClute. Two terabyte optical archival store. Computer Physics Communications, 45(1–3):403–407, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465587901810.

Turner:1981:MSJ

[TW81] M. F. Turner and J. A. Wesson. MHD simulation of jet. Computer Physics Communications, 24(3–4):343–354, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581901570.

Tiator:1983:EPV

[TW83] L. Tiator and L. E. Wright. Exact PWBA virtual photon spectrum for A(γ_{v}, D)R. Computer Physics Communications, 28(3):265–269, January 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583900437.

Tiator:1984:EPV

[TW84] L. Tiator and L. E. Wright. Exact PWBA virtual photon spectrum for a(γ_{v}, D)R. Computer Physics Communications, 35(1–3):C–850, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829598.

Talwar:1989:RTA

[TW89a] Indu Talwar and L. E. Wright. The radiation tail accompanying elastic electron scattering from the atomic nucleus. Computer Physics Communications, 55(3):367–380, October 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901331.

Tran:1989:TTD

[TW89b] T. M. Tran and J. S. Wurtele. TDA — a three-dimensional axisymmetric code for free-electron-laser (FEL) simulation. Computer Physics Communications, 54(2–3):263–272, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900908.

Ugai:1988:MSF

[Uga88] M. Ugai. MHD simulations of fast reconnection spontaneously developing in a current sheet. Computer Physics Communications, 49(1):185–192, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588902251.

Ugniewski:1980:LOT

[Ugn80] Stanislaw Ugniewski. Lateral observation of transparent non-homogeneous axisymmetrical media. Computer Physics Communications, 20(3):329–335, November 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900132.

Uchida:1985:HSV

[UI85] Keiichiro Uchida and Mikio Itoh. High speed vector processors in Japan. Computer Physics Communications, 37(1–3): 7–13, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585901316.

Vajtersic:1982:EAF

[Vaj82] Marian Vajtersic. An elimination approach for fast parallel solution of the model fourth-order elliptic problem. Computer Physics Communications, 26(3–4):357–361, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901291.

VanBinst:1981:EBL

[Van81a] Paul Van Binst. Evolution of the Brussels On-Line system since 1970. Computer Physics Communications, 22(2–3):345– 347, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581900692.

VanHove:1981:RRA

[Van81b] L. Van Hove. Review of recent advances in high energy physics. Computer Physics Communications, 22(2–3):187–197, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900485.

VanHerwijnen:1989:UTI

[Van89] Eric Van Herwijnen. The use of text interchange standards for submitting physics articles to journals. Computer Physics Communications, 57(1–3):244–250, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990221X. International Conference on

Computing in High-Energy Physics.

Varias:1989:EHC

[Var89] A. Varias. An extension of the HERA code to flux coordinate equilibria. Computer Physics Communications, 52(2): 167–174, January/February 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900015.

Vaz:1981:CFS

[Vaz81] Louis C. Vaz. The code �franpie: a semiclassical friction free model for calculating excitation functions for complete fusion of heavy ions. Computer Physics Communications, 22(4):451– 466, May 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901429.

Vaz:1984:CFS

[Vaz84] Louis C. Vaz. The code FRANPIE: a semiclassical friction free model for calculating excitation functions for complete fusion of heavy ions. Computer Physics Communications, 35(1–3):C– 697, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584828465.

Vohwinkel:1988:FCE

[VBD88] Claus Vohwinkel, Bernd A. Berg, and Alberto Devoto. A fast Cyber 205-ETA ^{10} program for SU(3) lattice gauge theory. Computer Physics Communications, 51(3):331–348, November 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901464.

Vinette:1988:PEE

[VČ88] F. Vinette and J. Čížek. Perturbation energy expansion using hypervirial theorem and symbolic computation for the Ndimensional hydrogen atom in an external spherically symmetric field. Computer Physics Communications, 52(1):35–41, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901695.

vanDam:1987:SMR

[vD87a] Andries van Dam. Solids modeling and rendering on workstations — a pictorial overview. Computer Physics Communications, 45(1–3):169–174, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901524.

vanDam:1987:CMH

[vD87b] P. H. A. van Dam. Computational methods in high energy physics. Computer Physics Communications, 44(3):255– 260, June 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900804.

vandenBos:1984:PPC

[vdB84] J. van den Bos. A programming package for the calculation of cross-sections and probabilities for charge-exchange processes. Computer Physics Communications, 35(1–3):C–232–C– 233, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584824716.

vanderMeulen:1984:TCR

[vdM84a] A. van der Meulen. Trajectory calculations for the reaction K + HBr → KBr + H in the eV-region. Computer Physics Communications, 35(1–3):C–121, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823796.

Virtue:1984:IPN

[VDM84b] C. J. Virtue, R. J. Douglas, and B. T. A. McKee. Interactive positronfit: a new version of a program for analysing positron lifetime spectra. Computer Physics Communications, 35(1–3):C–485–C–486, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826788.

vanderVorst:1989:IRM

[vdV89] Henk A. van der Vorst. ICCG and related methods for 3D problems on vector computers. Computer Physics Communications, 53(1–3):223–235, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901628.

Verkerk:1982:RCS

[Ver82] Peter Verkerk. Resolution correction: a simple and efficient algorithm with error analysis. Computer Physics Communications, 25(4):325–345, April 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900753.

Verbiest:1983:PAS

[Ver83] E. Verbiest. A program to analyze series of Mössbauer spectra. Computer Physics Communications, 29(2):131–154, April 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900693.

Verbiest:1984:PAS

[Ver84a] E. Verbiest. A program to analyze series of Mössbauer spectra. Computer Physics Communications, 35(1–3):C–869, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829744.

Vertes:1984:FPS

[Ver84b] P. Vertes. FEDGROUP-3 — a program system for processing evaluated nuclear data. Computer Physics Communications, 33(1–3):155–156, August/September 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901176.

Vertes:1989:FNT

[Ver89] P. Vertes. FEDMIX: Neutron transmission functions and lumped averaged cross-sections from standardized evaluated neutron data. Computer Physics Communications, 56(2): 199–229, December 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900209.

Vesztergombi:1989:ITA

[Ves89] G. Vesztergombi. “Iconic” tracking algorithms for high energy physics using the Trax-I massively parallel processor. Computer Physics Communications, 57(1–3):290–296, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902300.

Voevodin:1989:BSP

[VGD^{+}89] V. P. Voevodin, V. N. Govorun, A. M. Davidenko, An. V. Ekimov, N. S. Ivanova, V. I. Kovaltsov, Yu. M. Kozyaev, A. F. Lukyantsev, M. Yu. Matveev, V. A. Senko, A. N. Sytin, and G. V. Tishin. The 780/E 32-bit specialised processor-emulator. Computer Physics Communications, 57(1–3):532–535, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589902816.

Vogelsang:1987:MCF

[VH87] R. Vogelsang and C. Hoheisel. A Monte Carlo FORTRAN 200 programme for the determination of static properties of liquids vectorized to run on the CYBER 205 vector processing computer. Computer Physics Communications, 46(2):209–216, August 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900026.

Vialle:1987:GAS

[Via87] J. P. Vialle. The U_{A}1 3D graphics analysis system. Computer Physics Communications, 45(1–3):149–159, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901500.

Vielcanet:1988:IAC

[Vie88] P. Vielcanet. Influence of advanced computer-aided design techniques on space mission analysis software development. Computer Physics Communications, 49(1):235–242, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588902317.

Violino:1984:PCL

[Vio84] Paolo Violino. A program for computing level crossings and the Back–Goudsmit effect. Computer Physics Communications, 35(1–3):C–149, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824029.

Virasoro:1989:DMB

[Vir89] M. A. Virasoro. Disordered models of the brain. Computer Physics Communications, 56(1):93–100, November 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900556.

Visscher:1984:SCT

[Vis84a] P. B. Visscher. Simple compact tree storage of a multidimensional array. Computer Physics Communications, 31(4):311– 318, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046558490016X.

Visscher:1984:UFR

[Vis84b] P. B. Visscher. User-friendly routines for tree storage and retrieval of multidimensional arrays. Computer Physics Communications, 31(4):319–322, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465584900171.

Vogt:1982:MCM

[VKM^{+}82] H. Vogt, W. Kusmierz, H. Mencke, U. Schwendicke, H. Schiller, and M. Wille. A microprocessor controlled measuring stage

— part of a measuring device for track chamber pictures. Computer Physics Communications, 26(1–2):121–124, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901643.

vonMeerwall:1984:FAP

[vM84a] E. von Meerwall. A flexible, all-purpose curve-fitting program. Computer Physics Communications, 35(1–3):C–372, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825874.

vonMeerwall:1984:GPR

[vM84b] E. von Meerwall. A general-purpose routine for the analysis of spectroscopic peak shapes. Computer Physics Communications, 35(1–3):C–333, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825576.

vonMeerwall:1984:LSS

[vM84c] E. von Meerwall. A least-squares spectral curve fitting routine for strongly overlapping Lorentzians or Gaussians. Computer Physics Communications, 35(1–3):C–296, ???? 1984.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825266.

vonMeerwall:1984:FCA

[vM84d]  E. D. von Meerwall.  A Fortran code for automatic

spectrum analysis on medium-scale computers. Computer Physics Communications, 35(1–3):C–316, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825412.

vonMeerwall:1984:FPRc

[vM84e] E. D. von Meerwall. A Fortran program for routine analysis of magnetic susceptibility data. Computer Physics Communications, 35(1–3):C–500, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584826867.

vonMeerwall:1984:FPC

[vM84f] E. D. von Meerwall. A Fortran program to collect histograms of microscopic scalar interactions. Computer Physics Communications, 35(1–3):C–444, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482651X.

vonMeerwall:1984:FPI

[vM84g] E. D. von Meerwall. A Fortran program to interpret pulsed field-gradient spin-echo diffusion data. Computer Physics Communications, 35(1–3):C–557, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827320.

vonMeerwall:1984:FPS

[vM84h] E. D. von Meerwall. A Fortran program to simulate quadrupole-distorted NMR powder patterns. Computer Physics Communications, 35(1–3):C–418, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826284.

vonMeerwall:1984:SFP

[vM84i] E. D. von Meerwall. A simple Fortran program to interpret cubic X-ray powder diffraction data. Computer Physics Communications, 35(1–3):C–383, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584825965.

VanRensburg:1989:ICG

[VM89] D. J. Janse Van Rensburg and D. A. McNamara. An interactive computer-graphics-based processor for boundary element modelling of electromagnetic scattering by thin conducting wires. Computer Physics Communications, 55(3):457–461, October 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901379.

vonMeerwall:1981:FPF

[vMF81] E. D. von Meerwall and R. D. Ferguson. A Fortran program to fit diffusion models to field-gradient spin echo data. Computer Physics Communications, 21(3):421–429, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900187.

vonMeerwall:1984:FPF

[vMF84] E. D. von Meerwall and R. D. Ferguson. A Fortran program to fit diffusion models to field-gradient spin echo data. Computer Physics Communications, 35(1–3):C–677, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828301.

vonMeerwall:1984:FPRa

[vMT84] E. von Meerwall and D. Thompson. A Fortran program for reduction of NMR relaxation data. Computer Physics Communications, 31(4):385–392, March 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900213.

Vohwinkel:1988:FMG

[Voh88] Claus Vohwinkel. A fast method to gather neighbors in vectorized Monte Carlo simulations. Computer Physics Communications, 51(3):323–330, November 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901452.

Vertse:1982:GPC

[VPB82] T. Vertse, K. F. Pál, and Z. Balogh. Gamow, a program for calculating the resonant state solution of the radial Schrödinger equation in an arbitrary optical potential. Computer Physics Communications, 27(3):309–322, September 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582901783.

Vertse:1984:GPC

[VPB84] T. Vertse, K. F. Pál, and Z. Balogh. Gamow, a program for calculating the resonant state solution of the radial Schrödinger equation in an arbitrary optical potential. Computer Physics Communications, 35(1–3):C–824, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829392.

Vrba:1989:MCI

[Vrb89] Vaclav Vrba. The Monte–Carlo integration of cylindrical phase-space with leading particles. Computer Physics Communications, 56(2):165–180, December 1, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900180.

Vogelsang:1983:VMD

[VSH83] R. Vogelsang, M. Schoen, and C. Hoheisel. Vectorization of molecular dynamics Fortran programs using the Cyber 205 vector processing computer. Computer Physics Communications, 30(3):235–241, November 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900905.

Vvedensky:1986:UDC

[VSP86] D. D. Vvedensky, D. K. Saldin, and J. B. Pendry. An update of DLXANES, the calculation of X-ray absorption near-edge structure. Computer Physics Communications, 40(2–3):421– 440, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586901220.

vanTuan:1984:TRN

[vTBF84] Nguyen van Tuan, Pierre Bertrand, and Marc R. Feix. Time renormalisation numerical method for the K.d.V. equation. Computer Physics Communications, 33(4):335–342, October 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584901383.

Vermaak:1984:PSP

[VVM84] C. F. Vermaak, D. Vermaak, and H. G. Miller. A packed storage program for angular momentum coupling coefficients. Computer Physics Communications, 31(1):41–46, January 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465584900808.

Waite:1987:RNI

[Wai87] J. Waite. Routines for numerical interpolation, with 1st and 2nd order differentiation, having non-uniformly spaced points, out to 3 dimensions. Computer Physics Communications, 46(2):323–335, August 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587900087.

Walker:1989:IGM

[Wal89] Homer F. Walker. Implementations of the GMRES method. Computer Physics Communications, 53(1–3):311–320, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901689.

Wansleben:1987:UVM

[Wan87] Stephan Wansleben. Ultrafast vectorized multispin coding algorithm for the Monte Carlo simulation of the 3D Ising model.

Computer Physics Communications, 43(3):315–323, February/ March 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900488.

Warburton:1984:DVP

[War84] W. K. Warburton. DBLCON: a version of positronfit with non-Gaussian prompt for analysing positron lifetime spectra. Computer Physics Communications, 35(1–3):C–438, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826466.

Wasserman:1985:MSD

[Was85] Anthony I. Wasserman. Modern software development methodologies and their environments. Computer Physics Communications, 38(2):119–134, October/November 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900797.

Watase:1989:HEP

[Wat89] Yoshiyuki Watase. High energy physics computing in japan. Computer Physics Communications, 57(1–3):198–205, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465589902129.

Wilkinson:1982:UID

[WB82] G. G. Wilkinson and R. E. Burge. Using the ICL-DAP for satellite climatology. Computer Physics Communications, 26(3–4):469–471, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582901448.

Wood:1984:NRS

[WB84] John H. Wood and Michael Boring. A non-relativistic SCF atomic program to compute one-electron energies, total energies, and Slater integrals. Computer Physics Communications, 35(1–3):C–227, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584824662.

Wierzbicki:1988:GGC

[WB88] Andrezej Wierzbicki and Joel M. Bowman. GVSCF: a general code to perform vibrational self-consistent field calculations. Computer Physics Communications, 51(1–2):225–232, September/October 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588900744.

Weber:1984:ATD

[WBG84] Wim J. Weber, J. P. Boris, and J. H. Gardner. Alfven — a two-dimensional code based on shasta, solving the radiative, diffusive MHD equations. Computer Physics Communications, 35(1–3):C–536–C–537, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827149.

Wiegandt:1989:UOE

[WBV^{+}89] D. Wiegandt, R. Burn, W. Van Leeuwen, A. Osadzinski, R. Levine, and D. J. McKenzie. Unix in the OSF era — panel discussion. Computer Physics Communications, 57(1– 3):560–561, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589902889.

Wang:1984:BPP

[WC84] C. S. Wang and J. Callaway. BNDPKG. A package of programs for the calculation of electronic energy bands by the LCGO method. Computer Physics Communications, 35(1–3):C–473– C–474, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584826727.

Winkelmann:1985:PWS

[WC85] Klaus Winkelmann and Doris Croome. A precompiler written in spitbol applied to programs to analyze nuclear data. Computer Physics Communications, 38(2):309–313, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900955.

Wolf:1984:CSC

[WD84] D. Wolf and K. Differt. Computer simulation of correlated self-diffusion via randomly migrating vacancies in cubic crystals. Computer Physics Communications, 35(1–3):C–422, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826314.

Wolf:1984:DCF

[WDM84] D. Wolf, K. Differt, and H. Mehrer. Determination of correlation factor and NMR diffusion parameters from the computer-simulated random motion of vacancies in cubic crystals. Computer Physics Communications, 35(1–3):C–423, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826326.

Wenzel:1988:FSP

[Wen88] K-P. Wenzel. Future space plasma missions in the ESA science programme. Computer Physics Communications, 49(1):243– 252, April 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588902329.

West:1987:SE

[Wes87] D. West. Software engineering. Computer Physics Communications, 45(1–3):485, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587901986.

Weitzel:1983:EPC

[WGM83] William F. Weitzel, Thomas L. Gibson, and Michael A. Morrison. EXLAM, a program for the calculation and expansion of local model exchange potentials. Computer Physics Communications, 30(2):151–161, September/October 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465583900577.

Weitzel:1984:EPC

[WGM84] William F. Weitzel, Thomas L. Gibson, and Michael A. Morrison. Exlam, a program for the calculation and expansion of local model exchange potentials. Computer Physics Communications, 35(1–3):C–905, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584830040.

Wang:1981:ADO

[WH81] T. S. Wang and F. J. Helton. Algorithms developed to obtain particular numerical solutions to the ideal MHD equilibrium problem. Computer Physics Communications, 24(3–4):255–261, December 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901508.

Wienke:1985:PPN

[WH85] B. R. Wienke and R. E. Hiromoto. Parallel processing of numerical transport algorithms. Computer Physics Communications, 37(1–3):363–369, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585901742.

Wurtz:1989:NCS

[WH89] Diethelm Würtz and Georg Hartung. Neural computing on a system of parallel organized transputers: Software implementation and hardware configuration. Computer Physics Communications, 56(2):155–163, December 1, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900179.

Whalley:1989:DRH

[Wha89] M. R. Whalley. The Durham–RAL high energy physics database — HEPDATA. Computer Physics Communications, 57(1–3):536–537, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589902828.

Wehrum:1986:SKF

[WHD86] R. P. Wehrum, W. Hoyer, and G. Dießl. On some key features of Ada: Language and programming environment. Computer Physics Communications, 41(2–3):271–283, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900706.

White:1989:CSP

[Whi89a] Bebo White. The comparison and selection of programming languages for high energy physics applications. Computer Physics Communications, 57(1–3):538–542, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990283X.

White:1989:DDA

[Whi89b] Vicky White. Distributed data-acquisition systems (PAN-DA) for Fermilab experiments. Computer Physics Communications, 57(1–3):348–352, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902415.

Wienke:1987:ELP

[WHW87] B. R. Wienke, T. R. Hill, and P. P. Whalen. Eulerian and Lagrangian particle transport with drag. Computer Physics Communications, 43(2):171–180, January 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587902037.

Wickens:1987:CES

[Wic87] Fred Wickens. The computing environment: a summary of the Asilomar Conference on Computing in High Energy Physics. Computer Physics Communications, 45(1–3):1–8, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901354.

Wienke:1985:SSS

[Wie85] B. R. Wienke. SNEX: Semianalytic solution of the onedimensional discrete ordinates transport equation with diamond differenced angular fluxes. Computer Physics Communications, 38(3):397–402, December 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465585901079.

Wienke:1986:DCP

[Wie86a] B. R. Wienke. Decomp: Computational package for nitrogen transport modeling in tissues. Computer Physics Communications, 40(2–3):327–336, June 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465586901177.

Wienke:1986:EEM

[Wie86b] B. R. Wienke. ESECT/EMAP: Mapping algorithm for computing intersection volumes of overlaid meshes in cylindrical geometry. Computer Physics Communications, 39(2):259–266, February/March 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586901360.

Wiegandt:1989:UH

[Wie89] Dietrich Wiegandt. UNIX and HEP. Computer Physics Communications, 57(1–3):134–139, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902026.

Williams:1981:TMS

[Wil81] D. O. Williams. Trends in mass storage technology and possible impacts on high energy physics experiments. Computer Physics Communications, 22(2–3):353–363, April 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900710.

Williams:1984:PCD

[Wil84a] F. N. Williams. I. Program for calculating degenerate Raman bands of symmetric tops with an adaptation for infrared bands. Computer Physics Communications, 35(1–3):C–41–C–42, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823085.

Williams:1984:PFT

[Wil84b] I. R. Williams. Program for fitting transition energies into a level scheme according to the combination principle. Computer Physics Communications, 35(1–3):C–55, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823206.

Wilson:1984:DMB

[Wil84c] Stephen Wilson. Diagrammatic many-body perturbation expansion for atoms and molecules: III. Third-order ring energies. Computer Physics Communications, 35(1–3):C–452–C– 453, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584826569.

Williams:1986:SLM

[Wil86] David O. Williams. Software and languages for microprocessors. Computer Physics Communications, 41(2–3):295–326, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900731.

Willis:1988:GSA

[Wil88] P. J. Willis. Graphics structures and algorithms. Computer Physics Communications, 50(1–2):121–128, July 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901208.

Williams:1989:CEL

[Wil89] David O. Williams. Computing on the eve of LEP datataking: Are we ready? Computer Physics Communications, 57(1–3):8–14, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901860.

Wisinski:1987:EHE

[Wis87] Dennis E. Wisinski. An environment for high energy physics code development. Computer Physics Communications, 45 (1–3):417–421, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465587901834.

Wienke:1984:AES

[WL84] B. R. Wienke and B. L. Lathrop. Approximate and exact schemes for photon-Maxwellian electron cross sections. Computer Physics Communications, 34(1–2):77–86, November/ December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901619.

Wolfram:1984:KPN

[WMC84] Amy Wolfram, C. Fred Moore, and W. Rory Coker. Kinematical parameters of nuclear reactions. Computer Physics Communications, 35(1–3):C–110, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823681.

Wienke:1984:HMC

[WMCH84] B. R. Wienke, J. E. Morel, T. E. Cayton, and R. B. Howell. HPLAS: Multigroup cross section and reaction rate processor for coupled H, H_{2} and H_{2}^{+} transport applications in plasmas. Computer Physics Communications, 34(1–2):87–99, November/ December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901620.

Webb:1982:SLE

[WMH82] S. J. Webb, J. J. McKeown, and D. J. Hunt. The solution of linear equations on a SIMD computer using a parallel iterative algorithm. Computer Physics Communications, 26(3–4):325– 329, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901242.

Wessel:1982:AMW

[WML^{+}82] R. Wessel, E.-Ch. Müller, H. Lucius, K.-P. Pleissner, and R. Kröber. Application of a microcomputer in wide angle X-ray scattering. Computer Physics Communications, 26(1–2):209– 211, May 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465582901710.

Winkler:1985:AGM

[WMN85] Karl-Heinz A. Winkler, Dimitri Mihalas, and Michael L. Norman. Adaptive-grid methods with asymmetric timefiltering. Computer Physics Communications, 36(2):121–140, April 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585901183.

Wooff:1985:CWP

[Woo85] C. Wooff. Checking whether a point lies inside a polygon. Computer Physics Communications, 36(2):219–222, April 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585901250.

Wooff:1986:GSF

[Woo86a] C. Wooff. Generation of a smoothing function. Computer Physics Communications, 42(2):253–256, October 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558690041X.

Wooff:1986:SDL

[Woo86b] C. Wooff. Smoothing of data by least squares fitting. Computer Physics Communications, 42(2):249–251, October 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900408.

Worlton:1984:EMM

[Wor84a] T. G. Worlton. External modes of molecular crystals. Computer Physics Communications, 35(1–3):C–157, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482408X.

Worlton:1984:IMR

[Wor84b] T. G. Worlton. Irreducible multiplier representations. Computer Physics Communications, 35(1–3):C–215, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824571.

Wong:1984:CGF

[WP84] H. C. Wong and J. Paldus. Computer generation of Feynman diagrams for perturbation theory II. program description. Computer Physics Communications, 35(1–3):C–203, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824455.

Werner:1989:FHL

[WPVS89] C. M. L. Werner, M. Pimenta, J. Varela, and J. Souza. FADO 2.0: a high level tagging language. Computer Physics Communications, 57(1–3):364–369, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902440.

Wroldsen:1982:MCS

[Wro82] J. Wroldsen. MULTIQUARK calculations with SCHOONSCHIP. Computer Physics Communications, 27(1):39–48, July 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900054.

Wroldsen:1984:MCS

[Wro84] J. Wroldsen. MULTIQUARK calculations with SCHOONSCHIP. Computer Physics Communications, 35(1–3):C–802–C–803, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465584829239.

Wilson:1980:DMB

[WS80] S. Wilson and V. R. Saunders. Diagrammatic many-body perturbation expansion for atoms and molecules V. Fourthorder linked diagrams involving triply-excited states. Computer Physics Communications, 19(3):293–297, July/August 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900831.

Weniger:1982:PCS

[WS82] E. Joachim Weniger and E. Otto Steinborn. Programs for the coupling of spherical harmonics. Computer Physics Communications, 25(2):149–157, February 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0010465582900315.

Weniger:1984:PCS

[WS84a] E. Joachim Weniger and E. Otto Steinborn. Programs for the coupling of spherical harmonics. Computer Physics Communications, 35(1–3):C–777, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829045.

Wilson:1984:FLS

[WS84b] W. Wilson and L. J. Swartzendruber. A flexible least squares routine for general Mössbauer effect spectra fitting. Computer Physics Communications, 35(1–3):C–231, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824704.

Wormeester:1988:DDF

[WSV88] H. Wormeester, A. G. B. M. Sasse, and A. Van Silfhout. Deconvolution, differentiation and Fourier transformation algorithms for noise-containing data based on splines and global approximation. Computer Physics Communications, 52(1):19–27, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901671.

Wright:1980:CVP

[WV80] L. E. Wright and C. W. Soto Vargas. Calculation of the virtual photon spectrum in distorted wave analysis. Computer Physics Communications, 20(3):337–347, November 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465580900144.

Wright:1984:CVP

[WV84] L. E. Wright and C. W. Soto Vargas. Calculation of the virtual photon spectrum in distorted wave analysis. Computer Physics Communications, 35(1–3):C–639, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828003.

Wiggers:1989:UTZ

[WV89] L. W. Wiggers and J. C. Vermeulen. The use of transputers in the ZEUS online system. Computer Physics Communications, 57(1–3):316–320, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558990235X.

Winkler:1980:SNS

[WW80] R. Winkler and J. Wilhelm. Solution of the non-stationary electron Boltzmann-equation for a weakly ionized collision dominated plasma. Computer Physics Communications, 20 (1):113–118, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580901162.

Warren:1984:IVG

[WW84a] J. L. Warren and T. G. Worlton. Improved version of grouptheoretical analysis of lattice dynamics. Computer Physics Communications, 35(1–3):C–262–C–263, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824972.

Winer:1984:GWA

[WW84b] K. Winer and F. Wooten. Generalization of Weber’s adiabatic bond charge model to amorphous group IV semiconductors. Computer Physics Communications, 34(1–2):67–76, November/ December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584901607.

Worlton:1984:GTA

[WW84c]  T. G. Worlton and J. L. Warren.  Group-theoretical

analysis of lattice vibrations. Computer Physics Communications, 35(1–3):C–127–C–128, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823838.

Wei-wei:1988:CTB

[WwYpCg88] Wang Wei-wei, Gan You-ping, and Bao Cheng-guang. Calculation of transformation brackets of hyperspherical harmonics of four-body systems with arbitrary masses. Computer Physics Communications, 50(3):331–340, August 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901890.

Yates:1984:PCR

[Yat84] A. C. Yates. A program for calculating relativistic elastic electron-atom collision data. Computer Physics Communications, 35(1–3):C–79, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823425.

Yueh:1986:AEC

[YB86] F. Y. Yueh and E. J. Beiting. Analytical expressions for coherent anti-Stokes Raman spectral (CARS) profiles. Computer Physics Communications, 42(1):65–71, September 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586902328.

Youping:1987:PCR

[YFL87] Gan Youping, Liu Fuqing, and T. K. Lim. Program to calculate Raynal–Revai coefficients of a three-body system in two or three dimensions. Computer Physics Communications, 47(1):149– 157, October 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587900749.

Youssef:1989:VMC

[YMWW89] S. Youssef, W. Martin, T. C. Wan, and S. Wilderman. A vectorized Monte Carlo detector simulation program for electromagnetic interactions. Computer Physics Communications, 57(1–3):251–254, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589902221.

Youssef:1987:CLE

[You87] S. Youssef. Clustering with local equivalence relations. Computer Physics Communications, 45(1–3):423–426, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901846.

Young:1989:HOI

[You89] David M. Young. A historical overview of iterative methods. Computer Physics Communications, 53(1–3):1–17, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589901458.

Yurtsever:1986:ODV

[YP86] E. Yurtsever and M. Pehlivan. One-dimensional vibrational eigenvalue problem with numerical potentials. Computer Physics Communications, 39(3):431–437, April 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900998.

Yannouleas:1988:APS

[YP88] C. Yannouleas and J. M. Pacheco. An algebraic program for the states associated with the U(5) ⊃ O(5) ⊃ O(3) chain of groups. Computer Physics Communications, 52(1):85–92, December 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588901750.

Yannouleas:1989:AMS

[YP89] C. Yannouleas and J. M. Pacheco. Algebraic manipulation of the states associated with the U(5) ⊃ O(5) ⊃ O(3) chain of groups: orthonormalization and matrix elements. Computer Physics Communications, 54(2–3):315–328, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900945.

You-ping:1985:PCG

[YpMzCeCg85] Gan You-ping, Gong Min-zhuan, Wu Chong-en, and Bao Cheng-guang. Program to calculate generalized Talmi– Moshinsky coefficients of 3-body and 4-body systems. Computer Physics Communications, 34(4):387–393, February 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900670.

Yuba:1985:SDC

[YSHK85] Toshitsugu Yuba, Toshio Shimada, Kei Hiraki, and Hiroshi Kashiwagi. SIGMA-1: a dataflow computer for scientific computations. Computer Physics Communications, 37(1–3):141– 148, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585901468.

Yurtsever:1984:IPO

[Yur84] E. Yurtsever. An integral package for one-center integrals over slater-transform-preuss functions. Computer Physics Communications, 35(1–3):C–519, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827010.

Zacharov:1981:MCH

[Zac81] V. Zacharov. Modern computer hardware and the role of central computing facilities in particle physics. Computer Physics Communications, 22(2–3):199–207, April 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900497.

Zalewski:1985:LSM

[Zal85]  Janusz Zalewski.  Language selection for modular in-

terface systems (part 2). Computer Physics Communications, 38(2):295–300, October/November 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900931.

Zalewski:1988:STR

[Zal88] Janusz Zalewski. A step towards real time application of ada. Computer Physics Communications, 50(1–2):273–279, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465588901373.

Zehrfeld:1984:CAS

[ZC84] H. P. Zehrfeld and F. Casci. Calculation of axially symmetric resistive steady-state plasma equilibria in terms of pressure distributions. Computer Physics Communications, 31(2–3):155–165, February 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900432.

Zakrzewska:1984:NCM

[ZDN84] K. Zakrzewska, J. Dudek, and N. Nazarewicz. A numerical calculation of multidimensional integrals. Computer Physics Communications, 35(1–3):C–471, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584826703.

Zorn:1981:TMS

[ZHR81] R. Zorn, H. J. Herrmann, and C. Rebbi. Tests of the multispin-coding technique in Monte Carlo simulations of statistical systems. Computer Physics Communications, 23(4):337–342, August 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465581901740.

Zielinski:1988:SIW

[Zie88] Tomasz P. Zielinski. On a software implementation of the Wigner–Ville transform. Computer Physics Communications, 50(1–2):269–272, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465588901361.

Zangwill:1984:NPO

[ZL84] A. Zangwill and D. A. Liberman. A nonrelativistic program for optical response in atoms using a time-dependent local density approximation. Computer Physics Communications, 32(1):63– 73, April 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900080.

Zlamal:1980:FEM

[Zlá80] Milos Zlámal. Finite element method in physical and technical applications. Computer Physics Communications, 20(1):37–42, September 1980. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465580901046.

Zlokazov:1981:SPF

[Zlo81] V. B. Zlokazov. SMOOS — a program for the filtration of non-stationary statistical series. Computer Physics Communications, 21(3):373–383, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046558190014X.

Zlokazov:1982:APA

[Zlo82] V. B. Zlokazov. ACTIV — a program for automatic processing of gamma-ray spectra. Computer Physics Communications, 28(1):27–40, November 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465582900601.

Zlokazov:1984:APA

[Zlo84a] V. B. Zlokazov. ACTIV — a program for automatic processing of gamma-ray spectra. Computer Physics Communications, 35(1–3):C–832, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584829446.

Zlokazov:1984:DPA

[Zlo84b] V. B. Zlokazov. Domus — a program for the analysis of two-dimensional spectra. Computer Physics Communications, 35(1–3):C–595, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584827654.

Zlokazov:1984:SPF

[Zlo84c] V. B. Zlokazov. SMOOS — a program for the filtration of non-stationary statistical series. Computer Physics Communications, 35(1–3):C–673, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482826X.

Zlokazov:1984:USO

[Zlo84d] V. B. Zlokazov. UPEAK — spectro-oriented routine for mixture decomposition. Computer Physics Communications, 35(1–3): C–440, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046558482648X.

Zlokazov:1989:GGQ

[Zlo89] V. B. Zlokazov. GFIT — generalized quadratic approximation of functions under constraints. Computer Physics Communications, 54(2–3):371–379, June/July 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465589900970.

Zhang:1987:CSS

[ZNF87] J. F. Zhang, D. B. Newland, and M. T. C. Fang. The computation of self-similar arcs. Computer Physics Communications, 47(2–3):267–280, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901135.

Zamani-Noor:1988:CVP

[ZNO88] F. Zamani-Noor and D. S. Onley. Calculation of the virtual photon spectrum for a finite nucleus in distorted wave methods. Computer Physics Communications, 48(2):241–253, February 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465588900446.

Zohni:1984:FPC

[Zoh84] O. Zohni. A Fortran program for the computation of the generalized Talmi coefficients. Computer Physics Communications, 35(1–3):C–123–C–124, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584823814.

Zhu:1989:MDS

[ZR89] S.-B. Zhu and G. W. Robinson. Molecular dynamics study of liquid carbon monoxide. Computer Physics Communications, 52(3):317–321, March 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589901069.

Zelazny:1987:SDS

[ZS87] Roman Zelazny and Piotr Strzalkowski. A system for designing and simulating particle physics experiments. Computer Physics Communications, 43(2):231–243, January 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587902086.

Zwarts:1985:RNP

[Zwa85] D. Zwarts. RITSSCHIL, a new program for shell-model calculations. Computer Physics Communications, 38(3):365–388, December 1985. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465585901055.